A Complete Bibliography of Lecture Notes in Mathematics (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.04

Title word cross-reference

(1 ≤ p ≤ 2) [dRV00]. (2, 2) [GRKVL02d]. (2, 3) [Isa07d]. (iω_{0})^{2}iω_{1} [Lom00c]. (∂) [Li02g]. (s, μ, q) [Bil03c]. ∗ [AT09h]. 0 [Eva08i, GGM07]. 0 < p < 1 [Zva00]. 0^{2+}iω [Lom00b, Lom00a]. 1 [BSS09h, GGM07, Sim03, Wit07a, Wit07b]. 1: −2 [Efs05e]. 2 [ASWY07j, BC03, Cut02t, GM04, Jon08s, KL01i]. 3 [AP07, Bün02, Cut02u, Cut02o, DBAP08, Fla08, Jon08t, KL01d, KL01j, Riv02, Roh09g, Roh09h, Sch07i]. 3nj [Van03]. 4 [AMAF02, CT08, Hir02, Wit07f]. 5 [ABFM02, Bün02]. [M], A, P [GL05b]. [X, X]_{t} = t + ∫_{0}^{t} βX_{s^{−}}dX_{s} [CC01]. ^{2+} [San05]. A [JQQ04g]. A_{α} [HK03f]. A_{l} [Tak06k, Tak06-31, Tak06-33]. α [Ryz09]. B(ℓ_{1}) [Pis04]. B_{α} [HK03g]. B_{l} [Tak06l, Tak06-31, Tak06-33]. A_{r}(X) [Cut02b]. C [GL05f]. C∗ [HG04]. C^{∗} [Pis01i, Run02g]. C^{∞} [GdS03a]. C_{l} [Tak06m, Tak06y, Tak06-31, Tak06-33]. B_{β,ε}^{l} [Mel02d]. d(0d2) [RY09f]. d(M) = n^{2} [Isa07b]. d(M) = n^{2} − 1 [Isa07c].

D_{λ}, K [GSE07e]. Δ_{f,h}^{(0)} [HN05h]. ℓ_{p}^{N} [Zva00, dRV00, Lyu04, SZ00, STJV03]. G [Chr00d, Chr00e, Chr00f, Chr00g]. γ_{+} [Paj02g]. γd [FSU01]. GL_{n} [GSE07g]. GL_{n}(R) [JL05a]. GLn(K) [GSE07f]. GSp(4) [RS07d, Wei09f, Wei09g, Wei09k]. H^{1} [BM02, Mey02d, Yor02b]. H^{∞} [KW04]. H_{w} [Fil01e]. ∞ [Cra08]. IP_{3} [Sne05a]. K [HG04, Kie02i, Kie02l, Kie02c, Kie02q, Wan09]. K3 [JK04c, JK04m, JK04d]. K_{0} [Li02h]. K_{3} [JK04a]. L [Kac06a, Moc09h, Ste07j, Ste07p]. L^{1} [AP03, Yor02b]. L^{2} [Gil03f, Sch07d]. L^{2}(Γ \ Π) [Unt00q]. L^{[prescription−R]} [Gil03i]. L^{p} [Bob00, FSU01, HH09, Shi05, Tai09a, Tai09b]. L_{1} [AP07]. L_{2} [AP07]. L_{∞} [Mon01a]. l_{∞}^{n} [Sch07a]. L_{p} [JS04, KW04, MK07e, MK07d, Zva00]. l_{p}^{n} [Ole03]. λ [Wil08j]. ≤ 5 [Tak06x]. M [BC03, BF06d, MP04, Del06]. M_{α} [HK03e]. HP^{1} [BPF^{+}02c]. R [BM03, MR00]. R^{2} [CG03, Unt00b]. R^{d} [Cer01g]. R^{n} [Naz03a]. H_{′} [GL02c]. T [JK04i]. T (⌋, D, {D_{λ}}) [JK04g]. μ [Bil03d]. μ = S_{t}μ ∗ μ_{t} [Xu09]. N [Kho00]. n ≥ 3 [Isa07c]. ν [Cut02r, Cut02q, Cut02c, Cut02d]. O(2, l) [Bru02a, Bru02e]. p [Gil03c, GL00f, Har08i, Lyu04, Ole03, RK00]. p^{3} [Som02i]. P_{1}(R) [Unt00c]. P_{ϵ} [GL05b]. ∂ [Xia01h]. π [Gil03o]. Pos_{n} [JL05i]. Pos_{n}(R) [JL05b, JL05a, JL05f]. pq [Som02j]. Ψ_{2} [Pao03]. PSL(2, C) [ASWY07e]. PSL(2, Z) [Unt00c]. Q [Xia01d, Har08k, Har08l, Har08m, Har08n, Tar02]. R [Eva08d, Eva08j, Eva08i]. RΔf [MK07e]. Rf [MK07d]. S′ [JK04i, LH09y]. S^{3} [BPF^{+}02b]. S^{4} [BFL^{+}02]. σ [MOTV09a, MOTV09b]. SL(2, R) [Unt00l]. SU(2) [Ale04a]. SU_{q}(2) [Bia06]. Θ(τ, z) [NP01a]. u [Kar04b]. U_{q}(su(1, 1)) [Har08m]. V_{λ}, K [GSE07b]. W^{∗} [Run02g]. Z [BvdVP02m, BvdVP02n, BvdVP02p]. |L| [JK04g].

- [Run02g]. -Adic [Har08i]. -algebra [Pis01i]. -Algebras [AT09h, HG04, Run02g]. -approximation [MK07d]. -Bodies [Bou03b]. -Bradlow [Moc09h]. -Bridge [ASWY07j]. -categories [KL01i]. -Cobordisms [KL01d]. -Coefficients [Van03]. -connected [Jon08s, Jon08t]. -Continuous [MOTV09a]. -Convergence [HH09]. -convex [GL00f]. -Dependent [Sne05a]. -Differentiable [GdS03a]. -dimension [Chr00d]. -Dimensional [AP07]. -Ellipsoid [MP04]. -elliptic [Bil03d]. -equation [Xia01h]. -Estimates [Pao03]. -Estimators [BvdVP02m, BvdVP02n, BvdVP02p]. -fibrations [BF06d]. -flatness [Chr00e]. -folds [Cut02u]. -Forms [BSS09h]. -functional [KW04]. -Functions [Kac06a, Ste07j, Ste07p]. -Grassmann [Har08k]. -group [Li02h]. -growth [Bil03c]. -Hahn [Har08m]. -harmonic [Ryz09]. -Hessian [Wan09]. -Holonomic [Li02g]. -Hyperbolic [Eva08i]. -Hypergeometric [Tar02]. -injectivity [Chr00f]. -Interpolation [Har08l, Har08n]. -invariant [Kar04b, Ale04a]. -Laplace [Cra08]. -Loops [Kie02i, Kie02l, Kie02q, Kie02c]. -Manifolds [CT08, Isa07d, KL01j, Roh09h]. -norm [MK07e]. -norms [Bob00]. -orbits [Unt00c]. -projectivity [Chr00g]. -Pseudostable [Ole03]. -regularity [KW04]. -representations [ASWY07e]. -Schemes [LH09y].

-Slicely [MOTV09b]. -spaces [Sch07d]. -Stable [Del06]. -Structures [STJV03]. -Theory [HG04]. -Tree [Eva08d]. -Trees [Eva08j, Eva08i]. -Triangular [Gil03o, Gil03c]. -type [BC03]. -Variation [RK00].

/ [GGM07].

1-10 [BHK05]. 1-5 [CM06a]. 1-8 [CMI^{+}03]. 13-20 [EKMY02a]. 15-22 [AAR03]. 1997 [BGTT02]. 1998 [Ber00a, CO00, EKMY02a, Fas00a]. 1999 [BHK05, Ber02, BJS^{+}00, CMHL02, MRR02]. 1st [BJS^{+}00].

2 [Bou03b]. 2-8 [ACB^{+}03]. 2000 [ASTB03, ADD^{+}03, AZ03, CCJ^{+}03, DL04, MS00c, Tig04]. 2001 [ACB^{+}03, BKK^{+}09b, Can03, CMI^{+}03, CP04, GO08, MS03b, TZP04, VY03]. 2001-2002 [MS03b]. 2002 [BBF^{+}03, Bon08, KV03, MS03b, MS04, PV06, PP06, TW04, ZZ04]. 2002-2003 [MS04]. 2003 [BBH^{+}04, CM06a, CÇE^{+}04, CT08, Mar07, MP07, MS04, Pic05]. 2004 [BBSW07, CEKH^{+}07b, CP06, LCL07d, NS08g, SP06, TDP08]. 2005 [BM08b, DR08, Don07m, Eva08h, MS07]. 2006 [AF08a, BG08, Bra08h, Gui09w]. 2007 [GAC^{+}09, dH09q]. 21st [Yor06]. 2nd [MRR02].

3-9 [ADD^{+}03, DL04]. 30-July [BGTT02]. 30th [Yor06].

4.34 [Cao03g]. 4th [Fas00a].

6-12 [BBH^{+}04].

7-14 [MRR02].

9-15 [Can03].

A-L-G [HK03e]. A. [Bia06]. a.c.i [Van01h]. Abelian [AK01, Lee04f, Ols08c, Ols08g, Ols08i, Van01e]. Abstract [Bey07b, Doh03c, Doh03i, Füh05g, GOW04e, Jon08a, Kel00a, Kel00b, Kel00c, Kel00e, LH09a, Run02j, Doh03d]. Accretive [Chi09h, Chi09k, Chi09l, Chi09t, Chi09q]. Accretive-Type [Chi09t]. Accurate [HN05h]. Acid [Doi08a]. Action [JL05a, Sib04g, Sib04c, Sib04e, Sib04d, Sib04f]. Actions [Hsu00k, Isa07i, Mon01f]. active [HCG^{+}00]. Activity [Das09]. Adaptedness [Goh04c]. adaptive [Deu00, Nem00i]. Addition [Kie02b, Voi00a]. Additional [Phe01a]. additive [IK00, KP08, Pit06i]. Adiabatic [Hen05, Teu03g, Teu03b, Teu03c, Teu03f]. Adic [Har08i]. Adjoint [DBAP08, Dor08, Hel08, Let05g, LH09w, Gil03s]. Adjunction [Fre09a].

Admissible [Füh05e]. Admission [AGH03f]. Ado [BK07]. Adsorption [dH09i]. Advances [Cut00f]. Advantages [Sch03c]. Aerodynamic [JM00]. Affine [BPF^{+}02k, DFPS08j, Kut07a, Kut07i, Lom00h, Lus02, Van01c]. after [Kar04b]. Age [LB08, NCCFM08, Web08]. Age-Structure [NCCFM08]. Age-Structured [LB08]. Aggregation [AdlPP^{+}08, Nem00a, Nem00b]. Aging [Gui07]. Ahlfors [Paj02d]. Aims [PR07e]. aléatoire [Leu03a]. Aléatoires [Del02a, Del02b, PL03]. Algebra [Bec06, BPF^{+}02f, Chl09e, Gat00, KZ02f, Koe03, YC04g, vdH06l, CM02, GSE07f, Pis01i, SG02, Voi00e, Wen03b]. Algebraic [ASWY07a, AT09e, BM08b, BS03c, BSS09g, BF06l, Cat08, CT08, DIK00a, DFPS08a, GL02b, GL02c, GL02f, HB06, Li02c, Moc09b, Mol07e, NP01c, NP01d, NP01b, NP01i, NP01j, NP01q, Ols08b, Rob04b, Sch03h, Ter06, Tig04, Van01a, Van01d, Wal03b, dSW08f, vdH06a, NP01e, NP01l, NP01m, NP01o]. Algebras [AT09h, AMAF02, ABFM02, AM02b, Att06b, Bün02, CMHL02, Chl09b, Chl09f, Chl09g, CL02b, Fre09b, GdS03e, HG04, Let05b, Let05h, Li02b, Li02e, Li02g, Li02f, Li02h, Mol07c, Mol07f, Pfl01e, Som02n, Tia08b, Tia08c, Tia08d, Tia08e, CL02a, Fre09g, Fre09p, Fre09t, Fre09x, GR03e, Gre03e, Gre03h, KL01a, Lus02, Naz03b, Phe01h, Run02g, Run02c, Run02d, Som02g, Som02h, Som02i, Som02j, Ste03t]. Algèbre [SG02]. algèbres [Goi01]. Algebriac [Lom00i]. algébrique [Leu03b]. Algorithm [Gre03f, Sch07p]. algorithmes [MR00]. Algorithmic [Li02c, Dol03f]. Algorithms [AK04e, Koe03, Osi07f, Deu00, EK00]. allocation [AGH03q]. Allowed [dR02l]. Almost [GR03i, Jak05, Kar06, Lyu04, dS01a, GR03d, GR03e, Kun00a, GR03c]. Along [LCL07f, BCE04]. alternating [HK03h]. Alternative [Unt08a, Unt08g, Urb08d]. Amenability [KM04b, Run02o, Val08, Run02h, Run02e]. Amenability-like [Run02e]. Amenable [Pis04, Mon01f, Run02d, Run02c, Run02b]. American [BF03]. Amino [Doi08a]. Ampère [Caf03]. Analysis [AF08a, AF08b, Ber07b, Cha08a, Cow08, DH09v, DH09x, Dol03m, FJ04h, FLC08, Fri06c, Füh05g, GS09b, HN05d, HN05f, HN05g, Hel08, JP08, Jef04c, Kol00g, Kui03, Kut07a, Lin05, MS00c, MS03b, MS04, MS07, Pat05, Pri09a, Pri09b, Pri09m, Rit00a, TDP08, Uka06a, Unt08c, Unt08a, Unt08g, Unt08h, Vog08, Wen03h, WL08b, YC04l, Gui09z, GAC^{+}09, Kol00j, Moo01c, Ols03, Run02j, Wak00c, Zol00, dR02k, HN05n]. Analytic [BSS09i, GK04, INP04, JL05e, KV00b, Las04, Lom00d, Paj02h, Pfl01a, Wak00a, Yoc02, KV00a, KV00c, KV00e, MK07c, MK07e, MK07d, MK07f, MK07h, AGLT02e, GR03g, Paj02e, PV06]. Analyticity [Cer01a]. Anaplectic [Unt08c, Unt08f]. Ancestral [Tav04]. Anderson [BJ00]. Ando [Pis01b]. André [Yor06, Att06a, ÉY06]. Angiogenesis [LNH06]. Angular [Aba04]. Anharmonic [AHKM08f]. Anisotropic [Bil03j, DBAP08, Xia09b, Bil03f]. anisotropy [Bil03e]. Annihilation [Pri09c]. annulus [Pil03h]. Anomaly [Amm09]. Anosov [Cha08b, Oli02g]. Anticipations [Bau03]. Appendices [AK04n, HK03i]. Appendix [BHLV03h, Bro04h, CE05f, Efs05f, FJ04i, Goh04e, GdS03m, Grü02i, Grü02j, Grü02h, Kie02m, KZ02a, Mel02h, Mer04, NV04j, Pri09d, Rei04g, Teu03h, Wei09a, Wei09b, BPF^{+}02e, Fil01h, GR03h, WG00f]. Applicability [Sch03c]. Application [BCM09a, CM03, Cer01b, Dem08, Dor08, Fas00a, Gie07a, HN05e, Hip04, Kar04b, Lac03, LB09b, Mim03, Nua09, Phe01c, Phe01b, Tai09c, Unt08a, Wei09c, YC04l, Kun00g, Kun00h, DGP00, HN05i, Lom00b, NV04i]. Applications [All03, ACB^{+}03, AT09a, AT09i, AAFM07, AdlPP^{+}08, BBS07h, BHK05, Ben05, BG08, Bou07b, Bou08, Bra08a, Chi09a, Chu02i, Cow08, Doh03g, Doh03f, Goh04d, HH09, HN05b, Hel08, Hen05, Küm06, Las04, Lom00i, MY07a, MY07d, MV06, Mis08c, MOTV09h, Pha07d, Phe01d, Phe01e, Rös03, Sch07w, SP06, Sum09b, TF09, Tia08e, Tur00a, Utk08b, VY03, Vis04, YC04b, BS03c, BC03, Bur00, DM00, Eme00d, Fre09r, JK04e, Mik00, Mon01g, Van01h, AGH03g, CMZ02d, Dol03e, FJ04h, Pil03i, Sch03d, Som02b, Teu03d]. Applied [Coh03]. Approach [AJP06a, AJP06b, Car07, CR07, Cer01i, Éme01, GL02f, Gro07d, HH09, KS06b, MK07j, Pis07, Tai09a, Ter06, Urb08g, Hu00, Tel06l]. Approaches [Sch07b]. approche [Leu03b]. approchés [DDD02]. Approximate [Bis08a, Gro07g, Sch07e, Sch07f, Sch07d, Sch07c, Sch07w]. Approximating [Bar03]. Approximation [AZ03, BBS07f, Ber07e, BCM09r, Chi09b, Coh03, DD03a, Gau06, GL05a, HN05d, Jak05, Kaz06, Kut07d, Mel02c, Sch03h, Kol00h, MK07d, Phe01g, Rit00b, Rit00c]. Approximations [BvdVP02o, BC03, Gro07b, Yor02d, AGLT02e, DM00, FBF^{+}02c]. arbitrage [KS01b, KS05, Rás09b]. arbre [Leu03b]. arbres [Mou00]. Arising [Mim03, Mar00b]. Arithmetic [GM03, FBF^{+}02l, Unt00u]. Arithmetical [Bou07b, Har08a]. Arithmetics [FBF^{+}02m]. Arithmétique [Wit07a, Wit07b]. armed [BF03]. Arrangement [Tak06a, Tak06b]. arrival [AGH03l]. Art [Tel06a]. Article [KP01]. articles [ÉY02b]. Askey [Ter06]. Aspects [ADD^{+}03, AT09f, BF06l, BF06j, BF06o, Cha08a, Coh03, Dol03f, Fus08, MP09f, MS00c, MS03b, MS04, MS07, Mon02k, DIK00h, LTJ00]. Assessment [CLM^{+}08g]. Asset [Bac04]. Associated [BM03, Bui05, Dem09, EFL01, Gre05, HH06a, NNY05, ASWY07h, FBF^{+}02f, Fre09t, Unt00c, Unt00e]. Association [Hir02]. Associative [Li02b]. Assumptions [Wil08o]. Asymétriques [Pha01]. Asymmetric [Bac04]. Asymptotic [AK04d, AK04h, AK04l, Bis08i, Bis08j, Bor03, Cer01c, GL00a, GL00b, Grü02k, IK00, Mar00a, Mel02g, Sch04b, Uka06a, Yaf00a, MTJ00, Ver03, Yaf00p, VY03]. Asymptotically [Chi09i, Ras07d]. Asymptotics [AF08a, AP06, CM08, Daa03, Gui09a, Gui09b, Gui09w, Hab00a, HNKS06a, HN05h, Sod07, Kol00i, Pit06b]. asymptotique [Mou00, PL03]. Atom [Efs05c]. Atomique [BM02]. Atoms [Tak06-30]. Attacks [Doi08b]. Attracting [Grü02f, Grü02c, Grü02d, BvdVP02x]. Attraction [Gie07f, Grü02g]. Attractive [Mou03b]. Attractivity [Ras07a, Ras07j]. Attractors [JQQ04h, Osi07a]. Aubry [Oli02m, Sib04a]. Auditory [Bor05].

Augmentation [MP09a]. August [DR08, Fas00a, Yor06]. Auslander [Chr00a]. Automata [WG00b, WG00g, FBF^{+}02l, WG00c]. Automatic [FBF^{+}02a]. Automorphic [Lee04a, Lee04c, Lee04e, Lee04j, Unt00b, Unt00c]. Automorphism [Isa07h]. Automorphisms [Mol07c]. Autonomous [JM03, Ras07d]. Autotopism [Kie02h]. aux [Del02c, DDM02]. Average [KL00, Rit00a]. Average-Case [Rit00a]. Avoiding [Sla06d, Sla06e, Sla06i, Sla06l, Sla06q]. Axiom [Bow08a, Bow08c, Her06a, JQQ04g, QXZ09c]. Axiomatic [Kac06a]. Axiomatics [KL01e]. Azéma [MY06d].

B [Bil03h, Dol03i, Grü02i, Run02k, Oli02k]. Back [ASWY07b, AHKM08a, AT09b, BBS07a, BJS08a, BV08a, Ber07a, BCM09b, Bey07a, Bis08b, BLS07a, BvdVP02a, BvdVP02b, BvdVP02c, Bow08b, Bra08b, BSS09a, BF06a, BSPS09a, CMRT09a, CLM^{+}08a, Cer06a, Chl09a, Chu08a, Cla06a, DFT08a, DF05a, DH09r, DFPS08b, Don07a, EH08a, Eva08c, Fal08a, Fre09c, Fre09d, Fre09e, Fre09f, Fri06b, Gie07b, Gin09a, GSE07a, Gro07a, HH06b, Han07a, Har08b, Isa07a, Kac06b, Kon08a, Kut07b, LB09a, LH09b, LH09c, Lou08a, LCL07a, MY07b, Mar08a, MY02a, MMO^{+}07a, Mas08a, Mas07a, MK07a, Mie08a, MP09b, Mis08a, Moc09a, Mol07a, MOTV09c, Mon02i, Moo01a, NS08a, NS07a, Ols08a, Osi07b, Pet08a, Pha07a, PR07a, Pri09e, QXZ09d, Ras07b, Rás09a, Rez08a, Rob04a, RS07a, Roh09c, Rot09a, RY09a, SR09a, SF08a, Sch07g, ST08a, Sie09a, Sim08a, Sin08a, Sla06a, SV09a, Ste07a, Sum09a, Tai09d, Tak06c, Tel06b, Tia08a]. Back [Tof08a, Uka06b, Unt08b, Urb08a, Utk08a, Vak08a, Ves08a, Wal03a, Wal08a, Wei09d, Wil08b, Wit07c, Woj07a, WL08a, Xia09a, Yan09a, Zum07a, dH09a, dSW08a, vdH06b]. Background [Gro07e, HH06c, Hsu00a, Voi00e]. Bäcklund [BPF^{+}02m]. backprojection [Sch07p]. Backus [AGLT02c]. Backus-Gilbert [AGLT02c]. Backward [EP01, MY07d, MY07h]. Balance [Mar07]. Balanced [AGH03d]. Ball [Mal03, Ole03, SZ00, Glu04]. Balls [Woj07b, Nem00e, Nem00f, Nem00i]. Banach [Sim08f, BV08b, BV08p, BC03, Chi09f, Chi09m, Chi09v, FLC08, MTJ00, Mon01a, Pis01j, Run02c, Run02f, Run02l, Run02d, Sim08c, Sim08h, Sim08i, Sim08n]. band [BJ00]. Bandits [BF03]. Barking [Tak06g, Tak06f, Tak06w, Tak06z, Tak06e, Tak06-34]. Base [Con00]. Based [BB09, Dem08, Mos08b, EK00, JM00, vdH06d, vdH06e]. Bases [BHLV03i, Gre03c, GL05j, Li02d, Li02b, Li02c, Li02k, BHLV03f, Cao03b, Gre03h, Ste03q, Gre03e]. Basic [FBF^{+}02b, Li02c, Pet08b, San05, Son08, Tel06c, Wak00b, Yaf00b, dR02h, dH09p, vdD08a, vdDW08, BV08c, Fre09h, Li02a, Mon02b]. Basics [Gin09b, Gui09c, KZ02b]. Basin [Gie07f]. Basis [BN03, Gie07c, Gie07i, Har08j, Har08m]. Baxter [CMZ02e]. Bayes [Bis08c]. Beauty [Her06b]. Behavior [Grü02k]. Behaviors [GL02d]. Behaviour [Bou03a, Don07i, FR06a, Sch04b, Cer01c, Mar00a, YC04h]. Beliefs [SX04]. Bell [Pit06a]. Bellman [Cer01e]. Benefiting [KS06b]. Bernoulli [Bou04, Cer06b]. Bernoullian [Bob03a]. Bessel [AP07, Eis00, Pit06d, RY09f]. best [MK07d]. Beta [Har08d, Har08n]. Between [FL08, Lac08, LO00, BCM09q, Cat04b, Grü02e, KL01h, Xu08]. Beurling [Kut07j]. Beyond [Bey07b, Jon08t, Lom00i]. Bibliographic [Dol03g]. Bibliography [Doh03h, GOW04k, NV04k, Paj02i, Sch02e]. Bidimendionnelles [KP01]. bidimensionnelles [Kur01]. bifurcating [Lom00d]. Bifurcation [Ras07a, Ras07j]. Bifurcations [BHLV03i, Han07b, Han07c, Han07d, Han07g, Ras07d, Ras07c]. Big [GOW04l]. Bilinear [Gre05]. billiards [Sib04c]. Binomial [HK03g]. Biochemical [LNH06]. Biochemical/Mathematical [LNH06]. Biofilm [BJS08b]. Biological [Mim03]. Biology [MR08, Osi07e]. Biomedical [Amm09, TF09]. Biosciences [BFET05, Fri06d, Fri08, Sne05b]. Bipartite [Eva08m, Jon08b, Jon08f]. Birational [Kar04a]. Birkhoff [BHLV03d]. Birth [BBS07d, CM03]. Birth-and-growth [BBS07d]. Bishop [Phe01f]. Bivariant [Cun04]. Bivariate [AD08e]. Black [RS03]. Block [Gil03o]. Blocks [Chl09d, Chl09b, Chl09h]. BN [JK04c]. Bodies [BBS07f, BN03, BK03, Bou03b, BKM04, Cap08, LMTJ04, LMPTJ07, Piv07, GMR00, KL00, LTJ00, Pao00]. Body [Art04, MP04, SW03, GL00f]. Bohr [MK07b]. Bökstedt [LH09d]. Bol [Kie02i]. Boltzmann [GO08, Har00, SR09b, SR09e, Tos08, WG00e, WG00g]. Bond [ET07]. Bonferroni [Doh03c, Doh03i]. Boolean [BBS07g, Pri01]. Borcherds [Bru02a]. Borel [GO08, Yan06]. Boson [AHKM08d, Dha09]. bound [Cer06e]. Boundaries [dSW08f, dSW08g, dSW08e]. Boundary [AK04s, BCM09g, BB09, Kol00a, Las04, Lun04, Pis07, Ros04, SW03, Sol03, Ste03c, Tai09a, Tai09c, Tai09e, FS00a, FS00f, Gin09h, Mar00b, Phe01d, Phe01g, Phe01h, Phe01n, Pir00]. boundary-layer [Mar00b]. Bounded [CG03, GL05a, Jon08l, Kun00a, KS07, Mon01c, Pis01m, KS01a, MK07c, Mon01d, Pis01f, Pis01e, Pis01j, Pis01d, Bil03e, Gil03r]. Boundedness [DH09s]. Bounds [BvdVP02r, Led07b, AGH03n, Gin09k, NP01p, Sch02c]. bracket [Unt00f]. brackets [Unt00h, Unt00p]. Bradlow [Moc09h]. Braid [Hsu00k]. Braids [Ban09]. Brain [Amm09, Das09]. Branched [AS08]. Branches [Tak06n]. Branching [BvdVP02d, DM00, HH09, Kre09, Sla06b, Har00, Pit06d]. Branes [GH02]. Brascamp [Bar04, BCE04]. Bressanone [BBH^{+}04]. Bressanone/Brixen [BBH^{+}04]. Bridge [ASWY07j, AP06, AP07, Pit06b]. Brief [DFT08b, Ols08b, Bil03i]. Broken [Ols08h]. Brownian [AP06, BM08a, BBKM01, BP01, Bia09a, CL07, Cou07b, Dal03, ET05, Éme09, Gal03, IK00, KS00, MY06c, MY06e, MY06h, MY06i, Mei03, Mis08c, Mis08b, Mis08f, Mis08g, Mis08e, Mis08i, Mou03a, Naj07, Nou08, Pit06b, Pit06c, Pit06d, Raj06, RY09b, RY09g, Ser00, Ser07a, Ser08, Sla06j, Sla06s, SV09b]. Brownien [RVY06, Lac03]. Brownienne [Mal01, BS03b]. browniennes [BL08]. browniens [Eme00d]. Bruhat [Bro04c]. Bruhat-Tits [Bro04c]. Brunn [Bob07b]. Bryant [Ros02]. Bubbles [Mou03a]. Buchberger [Gre03f]. Buffer [GOW04g]. buffers [AGH03k]. Bunches [Tak06h]. Bundle [MMO^{+}07c, MMO^{+}07k, dS01z]. Bundles [AK01, BPF^{+}02h, Han07i, Lee04b, Lee04j, Lee04h, Kol00e]. Burkholder [FV08]. Busemann [Mil07]. BV [Bre07]. Byproducts [HN05q].

C [Dol03j, Grü02j, Run02l, Bil03i, Oli02l]. C.I.M [ADD^{+}03, CO00]. C.I.M./ C.I.M.E [CO00]. C.I.M.E [AF08a, ADD^{+}03, ACB^{+}03, AZ03, AAR03, BBH^{+}04, BBSW07, BGTT02, BM08b, BHK05, BKK^{+}09b, BG08, Bon08, BJS^{+}00, CM06a, Can03, CMI^{+}03, CT08, CMHL02, CCJ^{+}03, DR08, DL04, EKMY02a, Fas00a, GAC^{+}09, Mar07, MRR02, NS08g, PV06, TDP08, ZZ04]. C.I.M.E.-E.M.S [BBH^{+}04, BHK05]. Ca [San05]. Càdlàg [Kaj08]. Calabi [Roh09a, Roh09e]. Calcium [Sha05, Sne05a, Sne05b]. Calcul [BS03a]. Calculating [Ber09]. Calculations [Don05a, Gre03l]. Calculus [Bar06, BvdVP02k, Cou07b, Cut00c, DM08, Eis07, HN05m, Jef04d, Jef04g, Jef04f, Jef04b, Kar06, KW04, Ler08, Mas08b, Mis08f, Nua09, RV07, SF05, Gui09x, Pri01, Wak00b]. Calderón [DH09s, DH09t, Wil08c, Wil08d, Wil08e]. Cancer [Cha08a, Fri06c, Fri06d, KS06b, Lol06]. Canonical [EFL01, EE09, Hab00b, Voi00b, Pil03b]. Cantelli [Yan06]. Cantor [FBF^{+}02e]. CAP [Wei09e]. CAP-Localization [Wei09e]. Capacity [Paj02h, Paj02e, Paj02g]. Caps [Phe01m]. caractérization [Kur01]. Cardiac [Sha05, WHG05]. Carleson [Xia01f, dR02k]. Carter [GSE07b]. Cartesian [BCN03]. Cas [Mey02d]. Case [BCM09c, BSS09b, BSS09c, BHGVN06, Cut02c, Cut02d, Éme09, HH06g, Her01, Isa07d, Isa07b, Isa07c, Isa07f, Pri09g, Rit00a, Roh09g, Tak06g, Tak06f, Tak06e, Tak06-27, Urb08b, Wei09n, Wei09o, Bil03f, Bil03d, BSPS09d, Cer01a, Cer01j, Cer01k, Cer01l, Hab00a, Phe01i, Wei09i, Yaf00f, Yaf00o]. Cases [CE05e, Tak06-33, DIK00b, DIK00c]. CAT [Bou07a]. catégorie [Cha01]. Categories [CMZ02i, KL01c, KL01h, KL01k, LH09g, LH09f, NS08c, Chr00a, Fre09x, Fre09-27, Fre09-28, KL01i]. Category [Kas08, KL01d, NV04b, NV04a, Fre09u]. Cathy [How02]. Cauchy [Amb08, Paj02h, Paj02d]. Cell [Agu06, Lol06, Fri06d]. Cells [Rei05, San05, Sne05a]. Cellular [WG00g, WG00c, WG00b]. Center [BV08b, BV08n]. Centered [Cut02s]. Central [BK03, Hsu00k, MMO^{+}07s, MP09g, MP09h]. Centre [GO08]. Centro [BJS^{+}00, Fas00a, MRR02]. cercle [Leu03a]. Certain [Obł07, HK03h]. Certaines [LLP02]. certains [MR00]. Cetraro [AF08a, AZ03, AAR03, BGTT02, BM08b, BHK05, BKK^{+}09b, BG08, CT08, CCJ^{+}03, DR08, EKMY02a, Fas00a, GAC^{+}09, Mar07, NS08g, PV06]. Chacon [Cox08]. Chain [ASWY07c, Eva08m, Har08n, Osi07c]. Chains [Bey07j, DFPS08c, Eva08m, Har08h, HH01, JQQ04a, JQQ04b, Kre09, Tia08c, Cat04g, Eis03]. Challenging [Lou08b]. Chandra [Gui09a]. Change [Art04, Con00, KS07, Pes07]. Change-of-Variable [Pes07]. Changes [FG09]. channels [AGP09, Yaf00i]. Chaology [SF08b]. Chaotic [Éme01, MY06d]. Chaplygin [Mon02e]. Chapter [KZ02f, KZ02b, KZ02c, KZ02d, Li02a, Li02d, Li02b, Li02c, Li02e, Li02g, Li02f, Li02h]. Character [Cun04, Iwa06, Wei09f]. Characteristic [BSS09d, Let05g, AC02e].

Characterization [BM03, Mon01d]. Characters [Bia03, GSE07i, Sch02f]. Charged [dH09b]. Charleson [dR02n]. Chemical [Mim03]. Chemotaxis [Urb08e]. Chemotherapy [KS06b]. Chern [Bru02a, Bru02f, Cun04, Tar09]. Chernoff [AAFM07]. Chernoff-Type [AAFM07]. Childhood [Bau08]. Choice [Her06a, Her06b, Her06c, Her06f, Her06e]. Choquet [God05, Phe01d, Phe01g, Phe01f, Phe01h, Phe01i, Phe01l, Phe01n]. Choquet-Deny [God05]. Chosen [SW03]. Chow [Kar04b]. CIME [Mas08b]. Circle [BHGVN06, MK07e, MK07d, Yoc02]. Circulation [JQQ04a, JQQ04b]. CKN [Gal06]. Claims [BJR04a]. Class [BBS07d, DD03b, DFPS08f, Kac06a, PR07b, Ste07m, Ath00, Mon02f, Pil03i]. Classe [BM02]. Classes [BLS07i, BSS09d, BSS09k, BSS09l, BSS09o, Bru02a, DIK00i, GL05b, Moc09j, Xia01d, Bru02f, Pil03h, Sim08e]. Classical [BNT06, Bra08c, Hsu00b, Iof09, Kie02l, Pri08a, RB06, SF06, SF05, Sol03, Wak00c, Chr00b, Dou03, HN05d, HN05f, HN05g, Rit00e, Cao03e]. Classification [CE05b, CE05c, GL05c, Sae04c, Tak06d, Tak06-30, CE05d]. Clebsch [AC02c]. Clifford [Jef04c, JK04c, JK04g, JK04m, NV04d, Som02c]. Clifford-indices [JK04m]. Closed [BF06b, Gin09g, Gin09j, Gin09k]. Closedness [ASWY07d]. Closing [Tel06d]. Closure [Doh03d]. cluster [Cer06f, Cer06t]. Clustering [AD08a, CLM^{+}08g, Doi08a, Doi08b]. Clusters [KV00a]. CMCY [Roh09h, Roh09m]. Co [MOTV09a, Fre09g, Rab01]. Co- [MOTV09a]. Coadjoint [MMO^{+}07s]. Coagulation [Pit06e]. Coalescence [BBKM01]. coalescent [Pit06i]. Coalescing [Eva08j]. Coarse [Cer06c]. Cobordism [KL01h]. Cobordisms [KL01d, KL01k]. Cocommutative [Som02h]. Cocycles [Bha05, MMO^{+}07t]. codes [Sch02d]. codiffuseurs [Eme00f]. Codimension [Kho03]. Coefficient [GL05d]. Coefficients [Bob03a, Ste07d, Cer01g, Rab01, Unt00q, Van03]. Coends [KL01a]. coexistence [Cer06u]. Coherent [BSS09d, CLM^{+}08f, LH09v]. Cohomological [Mon01b]. Cohomologie [Kah04]. Cohomologies [NS08e, NS08f, NS08d]. Cohomology [AK01, BGTT02, Con04, Gre03c, Lee04c, Mon01c, Pfl01c, Tha08, Wei09g, Gre03l, Mon01d, Wei09b, Bro04f]. collaboration [AAR03]. Collapse [dH09j]. Collapsing [Mal03]. Collège [SG02]. colliding [O’C03]. Collisions [CL07]. colored [DH08]. Combinations [Pil03k, Pil03e, Pil03d]. Combinatorial [Doh03g, PP06]. Combinatorics [FBF^{+}02m, VY03, Hsu00c]. comme [CSM02]. Commensurate [GL02f]. Comment [Mil07, Sch00]. Comments [dR02o, Bil03i]. Common [Chi09e, Chi09c, Chi09d]. commutatifs [God05]. Commutative [Chl09e, Hir02, KZ02f, Mas03, NP01q, Gre03h, Som02g]. Commutativity [LH09e]. Commuting [Chi09d, MMO^{+}07b]. Compact [Dol03b, Füh05c, Gil03p, Gil03d, GdS03d, Mon01c, Rot09f, dS01b, CL02c, Gil03q, Gin09d, Gin09h, Kus05, Run02b, Ste03v]. Compactifying [Ols08c]. Compactness [HH01, HN05l]. Comparative [WL08b]. Comparison [Grü02i, HK03c, AGH03l]. Compartmental [Bra08i, vdD08a]. Compatibility [BG08]. Compatible [dS01c]. compléments [BL08]. Complete [BF06d, BF06c, Reb06]. Completely [Fag06, Goh04e, Pis01m, Phe01b, Van01a, Pis01f, Pis01e, Pis01j]. Completeness [BF06n, Bol08, Gre03j, Yaf00a, Yaf00p]. Completion [BF06e, GR03e]. Complex [Aba04, Fin06, For04, Hua04, Kol00b, Pil03k, Roh09a, Roh09e, Roh09i, Sae04d, Ste07o, Tak06-30, TDP08, Vog08, ZZ04, dS01a, dS01d, Bro04c, Kol00f]. Complexes [Jon08u, ASWY07h, Chr00d]. Complexity [Lio00, Ric09, Cat04e]. components [Ste03m]. Comportement [Mou00, Eme00b]. Composite [Xia01a, Pit06a]. Composition [Enr07, Unt00a, AC02d]. Compound [Tak06e]. compressible [SR09c]. compression [Cat04b, Cat04a]. Computation [Dol03m, Gre03c, Li02f, MV06, Roh09d, Sch07h, Sch07i]. Computational [BLS07b, Cha08a, Coh03, Deu00, Gau06, BJS^{+}00]. Computations [EH08b, Gin09e]. Computer [Gat00, Koe03]. computerized [Sch07i]. Computing [BHLV03i, KV00b, Wal08b, BHLV03g]. Concave [BN03, Bob07a, Fou05]. Concentrated [Sod07]. Concentration [Bob03a, Bob03b, Gui09d, Gui09e, Mal09, Mas07b, SZ00, dRV00, MP07]. Concepts [Pet08b, San05, Son08, Deu00, Yaf00b]. Concerning [Izh04a, JS04]. Concise [PR07c]. Concluding [Tai09f]. Conclusion [Cao03h, Sch07j, Sch07k, Sch07l, Sch07m]. concrete [Kel00a]. Condition [Don07k, Kot06]. conditional [Jac03]. Conditioned [Don07g]. Conditions [Bil03j, BG08, BF06n, Dol03h, Fil01h, GRKVL02f, GRKVL02a, Kut07f, Kut07g, MMO^{+}07q, Str03]. Conduction [AGLT02k, AGLT02g]. Conductivity [AK04p]. Cone [Rot09g, Unt00k]. Cone-Valued [Rot09g]. Cones [Rot09d, Ste03p, Ste03o]. Conformal [BFL^{+}02, BPF^{+}02k, Gur09, Hab00g, Vas02c, Vas02g, Hab00b, Hab00e]. Conjecture [GRKVL02b, Wei09l, Zha02e, NP01k]. conjectures [Paj02f, Ste03n]. Conley [Kun00g]. connected [BSPS09c, Eis03, Jon08s, Jon08t, Let05b]. Connection [HN05c]. Connections [Dol03h, HH06i, Lom00h]. Connes [Cun04]. Connexions [Eme00a]. Consequences [Ste07b, Sim08f]. Consistency [BvdVP02m, Fil01i]. Consistent [Fil01g, Fil01c]. Constant [Bou03b, Hab00g, Lom00e, Pao00]. Constants [Bob07a, BKM04]. Constellar [Tak06f]. Constrained [Bey07d, Rog03]. Constraints [ST03, Sus08, Oli02f]. Constructed [LH09e]. Constructing [Tia02, Gre03g]. Construction [Don05b, Gie07c, Gie07d, JQQ04c, KL01b, Roh09e, Tak06g, Tak06f, Tak06h, Tak06i, Tak06k, Tak06l, Tak06m, Tak06j, Tak06y, BvdVP02u, KL01a]. constructions [Fre09h, Fre09-31, Pit06g, Pit06k, Som02f]. Constructive [Lom00e]. Consumption [BF03, Rog03]. Contact [GRMH08, Sla06c, dS01e, dS01f]. Contemporary [Kot09]. Content [YC04a]. contents [Pis01a]. Continu [Mey02d]. Continual [Bar06]. Continuation [MY07j]. Continuity [BLM08, Urb08i, Ves08e, FSU01, Gre03a]. Continuous [Bar04, Bey07e, Bey07o, Füh05g, JQQ04b, LB08, MOTV09a, MOTV09b, Mon01c, Pri09f, Pri09l, Pri09m, Gal03, Mon01d, Unt00s, GOW04f]. Continuous-Time [LB08]. Continuum [Eva08b, Gio08, God08]. Contracting [Pat05]. Contraction [WHG05]. contractive [Chi09b, Chi09g, Chi09h, Chi09k, Chi09l]. Contrario [CLM^{+}08b, CLM^{+}08l]. contre [DDD02]. contre-exemple [DDD02]. Control [Agr08, AGH03a, Grü02k, Hip04, KS06b, Las04, MY07k, Mon02k, NS08g, Pfl01b, Utk08b, AGH03f, AGH03j, Cer01b, HCG^{+}00, JM00, Mon02h]. Controllability [Osi07d]. Controls [Las04]. Convective [HNKS06b, HNKS06e]. Convention [MY06b]. Conventions [Bey07c, Chr00c, Bil03k]. Convergence [Bec06, Bis08n, Bis08o, BvdVP02t, CLP07, EM05, HH09, Har00, Iva03, Kar06, Ryz09, Sch07n, Sla06d, Vil08, dR02p, Jac03, SST09, HK03d, Ste03x]. Conversations [Iwa06]. Convex [Art04, BBS07f, Bil03j, BN03, BK03, BKM04, Chi09o, GMR00, LMTJ04, LMPTJ07, MP04, Naz03a, Piv07, Rot09d, SW03, Bil03e, GL00f, LTJ00, Pao00, Sib04c, Wen03f]. convexes [Eme00b]. Convexity [AGH03c, Dol03h, AGH03m]. Convolution [Chu08b, Chu08f, Chu08g, Pri01]. convolutions [CC08]. Cooperative [Čhu02e, Čhu02f]. Copolymers [dH09c, dH09d]. Cores [Tak06-28]. Corners [KL01j]. Correction [KP01, Sch07b]. Correlated [Enr07]. Correlation [Woj07b]. Cortical [LB09b]. Cosets [Wei09a]. Cosiness [ÉS01]. Cosmological [AD08a]. Cost [Mec07, Ant00]. Costs [KM04c, Rás03, Rás09b]. Cotangent [MMO^{+}07c, MMO^{+}07k, dS01z, Kol00e]. Cotype [GM04]. Coulomb [Fér08]. Countable [Chi09c, Wen03c]. Countably [Mel02d, Mel02f, Mel02e]. Counting [HB06, Pil03h]. Coupled [Ste03e, Ste03h]. Coupling [WHG05]. courbes [Wit07a, Wit07b, Wit07e]. Courrège [Kol00i]. Cours [Mey02e, SG02]. Course [PR07c]. Courses [Mas08b, SG02]. covecteurs [Eme00f]. Cover [DF05b]. Coverings [BF06m, Roh09a]. Covers [AS08, Roh09f, Roh09o]. CR [ZZ04, Tum04, Yan09b]. Created [Büh05]. Creation [Pri09c, Ser08]. Credit [BJR04b, Car07]. Creeping [Don07b]. Cremona [AC02i, AC02g, AC02b]. Criteria [CS05, HN05l, NP01e, Zum07b, KS01b]. Criterion [ÉS01]. Critical [HNKS06b, HNKS06c, Mou03b, Rei04b, Rei04c]. Cross [NCCFM08]. Cross-Immunity [NCCFM08]. Crossed [Efs05c]. Crossing [Pis07]. Crustal [Tak06x]. Crusts [Tak06-27]. Cryptographic [Doi08b]. Crystal [CP06, Cer06-27]. Crystalline [Dol03m, NS08e, NS08f]. Crystallization [CM03, Cap00a]. Crystallography [Dol03i]. Cube [Glu04, Ant00, Wag00]. Cubic [CG03]. cumulants [Bia03, CC08]. cumulated [Cat04c]. Curvature [BPF^{+}02i, DD03a, Paj02h, Paj02c, PR02]. Curve [Cao03i, Fil01e, Cao03e, Cao03c, Cao03f, Cao03b, Cao03a, Cao03d]. Curves [Ade01c, Ade01d, Bro04i, CLM^{+}08d, Cut02s, ST08b, Tak06-30, Vak08b, Vas02b, Vas02g, Wer04, AC02f, JK04g, Oko03, Ste03o, Ste03m, Van01b, Wit07e]. cusp [Unt00g]. cusp-forms [Unt00g]. Cuspidal [YC04i]. Cycle [Agu06, Fri06d]. Cyclic [Con04, Cun04, GRMH08, Roh09a, Roh09f, Roh09o, Pis01i, Sch02d].

Cyclotomic [Chl09b, Sch02f]. cylinders [HCG^{+}00].

D [Bil03k, Oli02m, Run02m, DBAP08, Fla08, Sch07i]. Dalang [KS06a]. dans [CSM02, Eme00e, Leu03b, Yor02b]. d’après [KS03]. Darboux [BPF^{+}02d, dS01g]. Data [LH09u, Cat04b, Cat04a, Cer01j, Cer01k, Cer01l, Rit00f]. d’Attal [KP01]. Davis [FV08]. Dawson [BvdVP02d]. D’Azéma [Pha01, KP01, Kur01]. Decay [Bui05, HN05e]. Decaying [Bou03c]. Decision [CLM^{+}08b, Jon08c]. Decoherence [BO06]. Decomposability [App06]. Decomposition [Ade01a, Ade01b, BCN03, BM02, CJMS06, Jak05, Pil03b, Roh09j, Ste03e, Ste03h, Unt00h, Unt00r, Pil03f]. Decompositions [AP06, Ras07h, Pil03g, Run02a]. Decoupling [Cer06d, Piv07, STJ07, Teu03f]. Defaultable [BJR04a]. defect [Sch07b]. defined [CC08]. Definition [Mon01d, JK04g]. Definitions [Cut02m, Fre09h, Rit00e, Tel06c]. Deformation [Cat08, DIK00b, DIK00c, DIK00i, Man08, FS00a, FS00f, Ste03u, Ste03w, Ste03p]. Deformations [Ste03y, Tak06g, Tak06f, Tak06h, Tak06i, Tak06k, Tak06l, Tak06m, Tak06j, Tak06n, Tak06o, Tak06w, Tak06y, Tak06-29, Tak06-30, Tak06z, Ste03r, Ste03t, Ste03v, Ste03i]. Deformed [Led07b]. Degenerate [MY07h, Urb08g, Cer01a, Kol00e, Naz03b, Sie09b, Sie09g]. Degenerations [Tak06-29, Tak06-30]. degré [Wit07f]. Degree [AMAF02, ABFM02, AM02b, Bün02, Jon08l, Ols08g, Roh09g]. Degrees [NP01c, NP01d]. Dehn [Sei08]. del [Wit07f]. Delay [GL02b, GL02d, GL02f, Nus03]. Delay-Differential [GL02b, GL02d, GL02f]. Delayed [KS06a]. Delays [GL02f]. deletion [Ser08]. Deligne [Let05c, Let05f]. Denjoy [Paj02f]. Dense [Roh09i]. Denses [Yor02b]. Densities [BBS07d, KS01a, Kol00k]. Density [Doz03, Kut07a, Kut07f, Kut07g, Kut07i, Kut07j, Mas07c, Ves08c, Ves08b, Bil03h]. Denumerable [JQQ04a]. Deny [God05]. Dependence [GOW04i, KS06b, Cer01j, Cer01k, Cer01l]. Dependent [Bou03a, Sne05a, STJ07, NT00, Růž00b, Růž00c]. depinning [BvdVP02v]. Derivation [RS03, SR09i, AK04d, AK04h, AK04l]. derivations [Pis01f, Kie02d]. Derivative [HH06h]. Derivatives [Aba04, CR08, Ros05, MK07f, MK07g, MK07h]. Derived [DGMR08, Kas08, LH09g, LH09h, LH09i, LH09j, LH09f, LH09w, MP09c, NS08c, Wen03h, Wen03e]. Derrida [BK07]. Descent [Chi09n, LH09k, Sch02b]. Description [Pha01, Lom00d, Pis01a]. Descriptions [Khe09, Lac08]. Design [CO00, Sch07o, Utk08b, Pir00, Tar00]. Détaillé [Mey02a]. d’Été [ASTB03, Ber00a, Ber02, Bra08h, CP04, CP06, Don07m, Eva08h, Gui09w, LCL07d, MP07, Pic05, PP06, SP06, TZP04, TW04, dH09q]. Detection [AK04p, AK04q, AK04r, Amm09, AK04i, AK04m]. Determination [Chl09h, Gie07f]. Deterministic [AB01, vdD08a, AGH03i]. Deux [Bro03, Wit07e]. Developments [AHKM08m, AMR03, AJP06c, Don07d].

Deviation [Led07a, Cer06q, Cer06x, Gui09t, Cat04f]. Deviations [Bis08h, GOW04e, GOW04c, GOW04j, Gui09v, Pha07d, Cer06z, Fér08, Gui09u, JW00]. Diagram [LH09v, LH09y]. Diagrammatic [Sla06e]. Diagrams [LH09e, LH09j, LH09l, Ver03]. Dialgebra [Rab01]. Dialgebras [Lod01, LCFG01a]. Diameter [Art04]. dichotomies [BV08c, BV08k, BV08o]. Diffeomorphisms [Bow08a, Bow08c, Cha08b, Yoc02]. different [Phe01j]. Differentiability [FS00a]. Differentiable [Cat08, GdS03a, GdS03g, Sae04g, GdS03e, GdS03b, GdS03f, Oli02a]. Differential [AB01, BV08r, Bey07b, Bis08l, Bis08m, BE00, CE05a, CE05j, Čhu02e, Čhu02f, DM08, Deu00, EP01, Eva08a, Fag06, Gil03e, GL02b, GL02d, GL02f, JM03, JL05f, KRA09, Kho09, LCL07b, LCL07c, MY07d, MY07h, Mar00f, Mas08b, Mis08g, Mue09, Nua09, Nus03, Pfl01d, PR07c, PR07b, PR07f, RW08, Sie09f, Tai09b, Tof08b, Wu08, vdH06l, Ath00, Bas03, BC03, HN05m, Kun00j, Nem00g, NP01m, Oli02b, Sie09d, vdH06a, vdH06g, LCL07d]. Differential-Delay [Nus03]. Differentials [Sus08]. Differentiation [KMS01]. différentielles [Bro03]. diffractometry [Sch07c]. diffuseurs [Eme00f]. Diffusion [All03, Cos08, DGMR08, EFL01, JQQ04e, JQQ04c, JQQ04d, Kot06, Mel02b, Mim03, Ben03, Cap00b, Har00, Hir03, Hu00, Kol00b, Kol00e, Kol00f, Kol00h]. Diffusions [Bis08a, Bis08d, Bis08i, Bis08j, Bis08h, Bis08k, Bis08n, Eva08d, HH09, Kol00g, Kol00c, Kol00k]. Diffusive [MP03]. Digraph [Jon08d]. digression [Unt00d]. Dihedral [Dem09, Jon08e]. Dilation [Fag06]. Dilations [Bha05, Goh04a, Phe01o]. Dimension [BCM09i, Cer01i, MP09d, MP09g, Ras07c, Sod07, dH09m, dH09n, Chr00b, Chr00d, Li02e, Nou08, Riv02, RY09f, Sim03, Som02i, Som02j]. Dimensional [AF08b, AP07, BS04, Bou03c, DFPS08j, GRKVL02d, Gro07b, GL05f, Har08e, Kot06, Mee02, Mil07, Naj07, Pau06, Pha07c, Sei08, Unt08g, YYY09, BvdVP02u, BvdVP02p, BvdVP02v, BSPS09d, Gre03e, Izh04a, Lau09, LTJ00, Van01h]. Dimensionality [QXZ09f]. Dimensions [Chr00h, Cut02n, Fil01b, PR07f, Wil08k, Hab00c, Lom00b]. dimodules [CMZ02g]. d’Intégrales [Yor02d]. diophantienne [NP01h]. Diophantine [AZ03, FBF^{+}02c, NP01f, NP01h]. Dirac [BCM09c, Gin09c, Gin09d, Gin09i, HH06f, HH06j]. Direct [LH09h, LH09s, LH09w]. Directed [Jon08f]. directional [MK07g]. Directions [CLM^{+}08k, Hsu00e]. Dirichlet [Cip08, CJMS06, DGP00, Ste03c, Ste07c, Ste07d]. Disasters [Her06f, Her06e, Her06d]. Disconnected [Jon08g]. discontinuous [Han03]. Discrete [AGH03a, Bec06, BF06d, Éme01, GM07, Jon08h, MP03, Pri09g, Pri09m, Ser07b, AGH03l, Fér08, Pri08b, Sch07y, Sch07z, Unt00t, Wag00]. Discrete-Event [AGH03a]. discrete-time [AGH03l]. Discretely [Bis08d, Bis08j]. Discretization [Grü02k, Grü02a, Grü02e, Ste03d]. Discretizations [Grü02f]. Discs [Tum04]. Discussion [JQQ04d]. Disease [GLW08, Yan08]. Diseases [Bau08, JP08, MT08]. Disordered [Pas04, Ton09, DH08]. Disparition [Att06a]. Dispersal [Lou08b]. Disposition [GL05e, GL05m]. dissipation [JQQ04d]. Dissipative [HNKS06a, Tos08, Vil03]. Distance [Eva08f, Bol08, SST09]. Distances [BCM09q]. Distribution [JQQ04a, JQQ04b, Ste07o, Ste07p, Yan08, Sch07e]. Distributions [Bob03a, Bob07a, BF06k, Cox08, DH09z, GL00c, Iva03, MMO^{+}07p, Unt00e, GL00a, GL00b, GL00d, Kaj08, Unt00b, Unt00j, Unt00m, Unt00u]. d’Itô [Sim03]. Divergence [CS08, Pri09i]. Divisibility [BNT06]. Division [Agu06]. Divisors [Bru02a, EKMY02a, Bru02f, MY02b]. DNA [KL09]. Doi [CMZ02b]. Dolbeault [dS01h]. Domain [BLS07i, Ste03e, Ste03h]. Domains [BLS07c, FL08, LB09b, Grü02g]. Donaldson [Moc09b]. dont [BS03b]. Doob [Jak05]. Doppler [Sch07y]. d’ordre [Eme00e]. Double [KL01b, KL01c, KL01d, KL01h, Wei09a, CM02, HK03b]. Doubly [Urb08h]. Drag [HCG^{+}00]. drawdown [Mei03]. Drift [All03, Ser08]. Drift-Diffusion [All03]. Drifts [BvdVP02h]. Drinfel’d [Som02n, Som02g, Som02h, CMZ02e]. Driven [BJS^{+}00, LCL07b, LCL07c, LCL07d, Bas03, Nou08]. Drug [KS06b]. Dual [FJ04f, Fra06, Mey02d, Unt00n]. Duality [Chi09f, Con00, Gre03a, JL05h, LH09a, LH09l, Rog03, MS00b, Sim08b, Wen03g]. Dufresne [Bai08]. d’un [Bai08]. d’une [Leu03a, Sim03, Yor02a]. d’unicité [Bro03]. Dunkl [Dem09, MY06d, Rös03]. Duration [Deb09]. Dvoretzky [Sch07a]. Dyadic [Xia01b]. Dynamic [Car07, Rin03, Son03]. Dynamical [BO06, Cho03, Čhu02a, Čhu02b, Čhu02c, Čhu02d, Gat00, Grü02k, Han07g, HH01, JQQ04i, JQQ04k, Kun00f, MP03, Mor08, Osi07f, Osi07e, Pil03k, Pil06, Ras07a, Ric09, Ter05, FBF^{+}02f, FBF^{+}02k, Kun00g, Kun00h, Kun00d, Kun00e, Kun00i, CCJ^{+}03, EKMY02a]. Dynamics [AdlPP^{+}08, BHK05, Fla08, FBF^{+}02m, For04, Grü02a, Gui07, LB09b, LB08, Lou08b, Mie08c, Mof09, NL08, Osi07k, Sib04g, Sne05a, Sne05b, Ter05, Teu03g, dS01e, dH09o, Teu03e].

E.M.S [BBH^{+}04, BHK05]. École [ASTB03, Ber00a, Ber02, Bra08h, CP04, Don07m, Eva08h, Gui09w, LCL07d, MP07, Pic05, PP06, SP06, TZP04, dH09q, CP06, TW04]. Ecological [Cos08]. Ecology [Fri08]. editorial [Sch00]. Effective [All08a, Ste07e]. Effects [HN05g]. Efficient [BvdVP02l]. Eigenfunctions [BLS07c, HN05e, Kul09, Yaf00l]. Eigenvalue [Gui09v, Gin09g, Gin09h, Gin09k]. Eigenvalues [Gil03n, HN05h, Kul09, Led03, Led07a, Zha02b, Dou03, Gin09j, Lom00d]. Eigenvectors [BLS07h]. Einstein [Oli02l, Tel06e]. Eisenstein [JL05b, JL05c, JL05d, Unt00f, Unt00h, Unt00m, Unt00s, Unt00t]. either [Her06d]. Elastic [AK04q, dH09e, AK04k]. Elasticity [Fal08b]. Elastomers [Dol03c]. elastoplasticity [FS00f]. Elastostatics [AK04j]. Electric [Das09]. Electrical [Amm09, SW09]. Electromagnetic [AK04r, DBAP08]. Electrorheological [Růž00b, Růž00c, Růž00a, Růž00e]. Element [AMAF02, ABFM02, AM02b, Bün02, Dur08, Fal08b, Mel02i]. Élémentaire [Pha01]. Elementary [Bal04, GL05c, Her06g, Wil08f, Caf03]. Elements [Att06b, BBF08, BG08, Bra03, Dol03i, Fal08c, RV07, Kho08, Tel06p, Wei09m]. Elimination [Li02e, NP01g, NP01g]. Ellipsoid [MP04]. Ellipsoidal [Das09]. Ellipsoids [Kla04]. Elliptic [Ade01c, Ade01d, Bro04i, Lee04b, Ste03c, Tai09a, Tar09, Bil03d, Cap00b, DIK00b]. Embedded [BCM09a, PR02]. Embedding [Pis07, Xia01a, Fre09a, Zva00]. Embeddings [GL05h, GdS03h]. Emerging [JP08]. Émery [KP01]. Émile [GO08]. Empirical [BvdVP02m, BvdVP02n, Cat04e]. Encoding [CLM^{+}08i]. endofoncteur [Cha01]. Endomorphisms [QXZ09b, QXZ09c, QXZ09j, QXZ09l, QXZ09m]. Endoscopy [Wei09g, Wei09k]. energies [Bou00]. Energy [Dol03c, BHLV03c, Cer06w, Gui09y]. energy-momentum [BHLV03c]. Enhanced [Cer06e, FV08, Moc09d]. Enlargements [MY06e]. Enriques [DIK00f, DIK00g, DIK00k]. Ensemble [Del02e]. Ensembles [Del02a, Del02b, Led07b]. Entanglement [Wal08b]. Entire [Gil03k]. Entropy [DGMR08, JQQ04a, JQQ04b, JQQ04c, JQQ04d, LMPTJ07, Osi07g, QXZ09h, QXZ09j, QXZ09l, Vil08, Cha03, Gui09g, MS00b, SV00, GO08, JQQ04g, JQQ04i, JQQ04e, YC04c]. entwined [CMZ02b, CMZ02c]. Enumeration [Hir02, Sta03]. Enumerative [BM08b]. envelopes [MTJ00]. Environment [Zei04, Hir03]. Epidemic [All08b, Bra08j]. Epidemics [Bra08k, Ear08]. Epidemiology [Bra08i, BvdDW08, LB08, MR08]. Epilogue [Gre03b, Sae04e, AGLT02h, HN05j]. Equation [All08a, Amb08, BCE04, Bey07k, Bey07n, Caf03, Cra08, Dal09, GO08, HN05i, SR09e, Tos08, Wan09, CMZ02h, CMZ02f, CMZ02g, CMZ02e, CC01, Kol00b, Mar00b, Oli02l, Pri08b, SR09b, Ste03w, Xia01h, Xu09, Sim03]. Equations [AB01, All03, Arn08, BBS07d, BV08r, Ber07f, Bey07b, Bey07f, Bey07g, Bis08l, Bis08m, BE00, Bou03a, CMZ02i, CE05a, CE05j, Chi09a, Čhu02e, Čhu02f, Cla06b, Con06, Cos08, DM08, Dah03, Dem08, EP01, Eva08a, Fag06, Fil01f, Fil01b, Gal06, GL02b, HNKS06a, HNKS06b, HNKS06c, HNKS06e, HNKS06f, INP04, JM03, KRA09, Kho09, KW04, Las04, LCL07b, LCL07c, LCL07d, MY07d, MY07h, MY07i, Mar00c, Mar00f, Mas08b, Mey06, Mis08g, Mue09, Nua09, Nus03, PR07c, PR07b, PR07f, Sch04b, Sie09f, Uka06a, Wu08, Zhi01d, dSW08d, Ath00, BV08s, Bas03, BC03, Cer01e, Cer01h, Cer01g, Deu00, Kol00f, KV00c, Lom00h, Moo01c, NP01m, Yaf00k, Zhi01a, vdH06a, vdH06g, Bro03]. Equilibria [Han07b]. Equilibrium [AJPP06, Cha08b, Fin06, HN05o, Vil08]. Equity [Car07]. Equivalence [BCM09d, Hua04, Kar04a, KM04c, KM04a, Kec04, Büh05]. Equivalences [Kel00e]. equivalent [Che03, KS01a]. équivalentes [Bro03]. equivariance [BV08n]. Equivariant [Gat00, Kas08]. Ergodic [Bel06, Bow08c, Cha08b, Her01, QXZ09b, QXZ09e, Phe01c]. Ergodicity [Bou07a, Cer01d]. Errata [FJ04j, FJ04k, FJ04l, FJ04m, FJ04n]. Erratum [Mie08b, Rás09b]. Error [Ber07b, AGLT02d]. Errors [Wil08m]. Espace [DGP00]. Espaces [Yor02b]. Essential [Chl09f, MP04]. est [BS03b, Mey02d]. Estimate [Ves08g, Cer06o]. Estimates [AAFM07, Ben05, Del03, Gil03p, GL05d, HN05s, JL05e, Kho08, MK07c, MK07e, MK07d, MK07f, MK07g, MK07h, NP01q, Pao03, Sla06e, Ste03h, Tai09i, Urb08j, AGLT02d, Gin09g, Gin09h, Gui09-27, IK00, MK07b, Nem00a, Nem00b, NP01m, Tel06k, Tel06q, Tel06r]. Estimating [Bis08d, Nem00c, Nem00d, Nem00e, Nem00f, Nem00g]. Estimation [Bis08a, Bis08c, Bis08k, Bis08l, LB09b, Mas07c, Pet08f, Pet08g, Pet08h, Cat04b, Nem00i]. Estimators [Bis08h, Bis08n, Bis08o, BvdVP02m, BvdVP02n, BvdVP02l, BvdVP02p, Cat04d, Cat04e]. Estivo [BJS^{+}00, Fas00a, MRR02]. Euclidean [App05, GOW04c, GL00f, LMPTJ07, Lyu04, Rei04e, Rei04f]. Euclidian [LB09b]. Euler [VY03, BSS09j, Con06, SR09c, SR09f, Ste07c]. Euro [ADD^{+}03]. Euro-Summer [ADD^{+}03]. European [VY03]. evaluating [Lom00f]. Evaluation [Gro07g]. Evaluations [Pen04]. Even [Lyu04]. Event [AGH03a, AGH03e, AGH03f]. Events [STJ07]. everyone [Str00]. Evolution [Arn08, BBS07d, Bey07b, Bey07f, Bey07g, Bey07j, Bey07k, Bey07n, Cao03i, Fri08, INP04, Lou08b, Mey06, Sch04b, Sol03, Tia08b, Tia08c, Tia08d, Tia08e, Cao03a, Cao03d, Cao03e, Cao03f]. Evolutionary [Mie08c, Zhi01a]. Evolutions [Wer04]. Evolving [ADD^{+}03]. Exact [Dem08, QXZ09f]. Example [Chi09g, Rit00e, Ste03s]. Examples [AT09c, Bey07d, Bey07e, Bey07f, Bey07g, Gie07a, Grü02j, Hsu00d, Li02a, MMO^{+}07d, PR07e, Roh09h, Roh09i, Roh09n, Tak06v, Tak06w, Tak06z, dSW08h, ASWY07l, Fre09s, Goh04d, Pil03j, Run02d, Som02d]. Exceptional [Iwa06]. Excessive [Bui05]. Exchangeable [Pit06f]. Excitable [Sne05a]. Excitation [WHG05, Sie09d]. Excitation-Contraction [WHG05]. Exclusion [Doh03i]. Excursion [Don05a, Pis07, Sla06j]. Excursion-Theoretical [Pis07]. Excursions [Dal03, Deb09, Don07h, Hob00, Lac03, RY09f]. exemple [DDD02]. Exercises [MY06f]. Exhaust [Tak06-33]. Existence [Ale04c, AP03, Mar00d, Ser07b, Ves08c, Ves08b, BSPS09f, JK04h, Phe01f, Zhi01a]. Exit [EFL01, Pis05, Sie09b]. Expanding [QXZ09g]. Expansion [Sla06f, SP06, Sla06k, Sla06l, AGLT02b, Bru02d, Gui09f, Gui09y, Unt00f, Unt00s, Unt00t]. Expansions [AK04l, AK09, DH09y, Mel02g, Sla06g, Xia01i]. Expectations [AHKM08c, AHKM08b, Pen04]. Experimental [CLM^{+}08c, Gre03k]. Experiments [BLS07b, CLM^{+}08j, Sch07x]. Explaining [Bau08]. Explicit [Gin09e, dSW08h]. Exploration [LB09b]. Exponent [Urb08b, Sch02c]. Exponential [BV08c, Ben05, BS04, Bui05, Daa03, Eis00, Lom00f, Mas07d, Wil08g, Wil08r, BV08k, BV08o, Cat04g]. Exponential-Square [Wil08r]. Exponentially [HN05h, Mel02f]. Exponentielles [Yor02c]. Exponents [JQQ04i, JQQ04h, Osi07r, QXZ09e, Sie09f, Bou00, Kun00d, Sie09e]. Extended [Bre03, Gio08, Rug08]. Extending [Cox08, Phe01j, Phe01m]. Extension [Fre09i, Xia01g]. Extensions [DFPS08d, Füh05e, GS09b, Hsu00k, KZ02f, MMO^{+}07i, MMO^{+}07s, RVY06, vdD08a, KZ02b, Teu03d, Goh04a]. Exterior [YC04g]. External [NV04f]. Extracting [CLM^{+}08d]. Extremal [STJV03, Tum04, Vas02c, Vas02d, Vas02b, Lau09]. extrémales [Ben03]. Extreme [DFPS08f]. Eyring [FS00b].

F [Chi09t, Chi09q]. F-Accretive [Chi09q]. Faber [BLS07d, BLS07h]. Facets [GGM07]. Facilitated [CMRT09b]. factor [Wei09j]. Factorisation [Don07f]. factorization [AC02h]. Factors [DFPS08a, DFPS08g, Kel00e]. Facts [Čhu02a, Goh04e, Ler08, Fas00c]. Faithful [AMAF02, ABFM02, Bün02]. Families [AD08e, BM03, Chi09e, Chi09c, Chi09d, Chi09h, Chi09i, Chl09b, NS08f, Pet08c, Pet08d, Roh09h, Roh09i, Roh09m, Roh09o, Vas02b, Vas02g, RK00, Lee04f]. Fast [KS00]. fat [Ste03s]. Favorite [DF05b]. fBm [Mis08h]. FBSDEs [MY07a, MY07c, MY07l]. feedback [AGH03j]. Feller [Kol00i]. FEM [Mel02b]. Fenchel [Sim08b]. Fermi [DF06, Mar00c]. Few [DFPS08a, DFPS08g]. Feynman [AHKM08d, AHKM08e, AHKM08f, AHKM08l, DM00, Jef04g, RY09b]. Fibers [Roh09i, Sae04c, Sae04d, Sae04g, Tak06x, Tak06-28]. Fibrations [Kie02k, BF06d]. Fibre [Sam08]. Fibred [GdS03j]. Field [AHKM08d, KL01j, KL01e, Rob04b, Tal07, Sch02b]. Fields [Ade01b, Ade01g, Amb08, Bou07b, BSS09b, BSS09i, DBAP08, Efs05c, GH02, HH06k, HN05q, Kie02l, LB09b, Sch02f, Ste07j, Xia09b, dS01r, BSS09n, Lom00j, Oli02b]. Filtered [Li02k, NS08c, Sch07p, Li02d]. Filtered-Graded [Li02k, Li02d]. Filtering [Mis08b, DM00]. Filtration [Éme09, Fas00b, Leu03a, Mal01, Osi07h, Yor02a, BS03b, DGP00, Fas00c, Mik00, Fas00a]. Filtrations [Büh05, Éme06, Mal01, MY06e, MY06i, NS08e, NS08f, NS08d, Han03]. Finance [BBH^{+}04, BBF^{+}03, CÇE^{+}04, CEKH^{+}07b, Cut00d, Pha07d]. Financial [Bau03, Mis08c, SX04, Rás09b]. Finding [Pet08g]. fine [Naz03b, GR03d]. Finite [BBF08, BG08, Bou07b, Bra03, Cer01i, Chi09e, Chi09h, Chi09i, Dur08, Fal08b, Fal08c, FR06b, Gil03m, Gil03n, Gro07b, GL05f, GL05j, Hsu00k, JQQ04b, KH07, Let05h, PR07f, Sch02f, YC04j, Cer06v, FBF^{+}02l, Gre03e, Kuk02, PR02, GdS03l, Mel02i]. Finite-Dimensional [Gro07b, Gre03e]. finite-gap [Kuk02]. Finiteness [Bro04g]. First [Gui09f, Izh04b, JL05c, YYY09, IK00, GL02e, Teu03b]. First-order [Gui09f]. First-Out [GL02e]. Fit [Pha07c]. Fitzpatrick [Sim08j]. Fixed [Ber07b, Ber07e, Ber07i, Ber07k, Ber07l, Car07, Chi09b, Chi09e, Chi09c, Chi09d, Chi09r, Sib04d]. FK [Cer06f, Cer06r]. Flat [MRR02, Mee02, Hab00a, Hab00b]. flatness [Chr00e]. Floer [Tha08]. Floquet [Lom00e]. Flour [ASTB03, Ber00a, Ber02, Bra08h, CP04, CP06, Don07m, Eva08h, Gui09w, LCL07d, MP07, Pic05, PP06, SP06, TZP04, TW04, dH09q]. Flow [DD03a, EH08b, LvdLM00, Cao03c, EK00, HCG^{+}00, Zol00]. Flows [CM06a, GOW04h, Urb08b, Růž00b, Růž00c]. ^{≪} [Mey02d]. ^{≪}BMO^{≫} [Mey02d]. ^{≪}Générale^{≫} [Mey02a]. fluctuating [IK00]. Fluctuation [JQQ04d, KP05, Don07m]. Fluctuation-dissipation [JQQ04d]. Fluctuations [BPT^{+}08, KP08]. Fluid [Cut00b, Fla08, Mof09, Uka06a, BKK^{+}09b]. Fluids [FS00b, FS00e, Růž00a, Urb08e, Bil03i, FS00d, Růž00b, Růž00c, Růž00e]. Fock [Pau06]. Fokker [HN05i, HN05n, HN05r, HN05s, HN05o]. Folds [Cut02o, Cut02u]. Fonctions [Eme00b, Ben03, God05, Mou00]. Fondamentaux [Del02d]. Force [Pat05]. forest [Pit06c]. Forests [Jon08f, Jon08i, Pit06i, Pit06j]. Form [AHKM08h, CLP07, Han07h, dS01z, Oli02l, Voi00b]. formal [SR09b, Ste03u]. Formalism [Bow08e]. formalisms [Oli02j]. Formats [Ste03l]. Formes [DGP00]. Forms [BSS09h, Can01, Cip08, Izh04a, Lee04g, Lee04a, Lee04c, Lee04e, Lee04j, Lee04d, Nes03, Tig04, Unt00o, Unt08a, Unt08h, Vis04, Wei09l, dS01f, dS01s, Bru02b, DIK00d, Oli02b, Unt00g]. Formula [AK04h, Chy07, KS07, MY06a, Pes07, QXZ09h, QXZ09j, QXZ09l, RS03, SY07, Wil08c, Wil08d, Wil08e, Mar08d, Naz03b]. Formulae [AK04d, ETZ07, DM00]. Formulas [BSS09g]. Formules [Yor02c]. Forward [Fil01e, MY07d, Unt00k]. Forward-Backward [MY07d]. Foundation [CM06a]. Foundations [GL00c, LH09l]. Four [Dei03, Hel08, MY07e, Sei08, Moo01f, Ste03s, Efs05a, GRKVL02d]. Four-Dimensional [Sei08, GRKVL02d]. four-manifolds [Moo01f]. Fourier [Bru02d, CL02a, CL02b, HH06d, KW04, Let05f, Let05g, Let05h, Wak00b, Wil08h, Yaf00c, dR02f, dR02e, dR02d, dR02p]. Fractals [DF05b, FBF^{+}02c]. Fractional [BM08a, Bis08k, BB09, Cou07b, ET05, Mis08c, Mis08b, Mis08f, Mis08g, Mis08e, Mis08i, KS00, Nou08, Efs05e]. fragmentation [Pit06e]. fragmentations [MS03a]. Framed [KL01d]. Frames [Kut07h]. Framework [GL02b, Růž00d]. Franca [ACB^{+}03, BBSW07, Bon08, BJS^{+}00, CM06a, Can03, CMI^{+}03, CMHL02, DL04, MRR02, ZZ04]. France [SG02]. Free [BNT06, Gui09g, Gui09h, Sol03, Spe03, AHKM08k, Bia03, Gui09y]. freely [Voi00a, Voi00d]. Freeness [Gui09i]. Freidlin [Sie09g]. French [DMÉRS09, And06, Att06a, ADDD^{+}06, Bai08, BS03a, BS03b, Ben03, BM02, Bia06, Bro03, BL08, Cha01, CSM02, DDD02, Del02a, Del02b, Del02d, DGP00, DDM02, Eme00a, Eme00b, Eme00d, Eme00e, Eme00f, FG09, God05, Goi01, Kah04, KP01, Kur01, Lac03, LLP02, Leu03a, Leu03b, Mal01, Mey02e, Mey02d, Mey02a, Mey02b, Mey02c, MR00, Mic03b, Mou00, NP01g, NP01h, PL03, Pha01, RVY06, SG02, Sim03, Wit07a, Wit07b, Wit07e, Wit07f, YYY09, Yor02a, Yor02c, Yor02d, Yor02b]. Frequency [JQQ04f, SW09]. Fresnel [AHKM08g, AHKM08h]. Freter [BJS08b]. Fricke [ASWY07e]. Frobenius [BLS07h, CMZ02h, CMZ02c, CMZ02i, Kie02j]. Frobenius-Separability [CMZ02h]. Front [AMTV08, AMS^{+}08a, ASWY07f, AHKM08i, AFS08, AADH08, ACC^{+}08, AT09d, AD08b, AMS^{+}08b, BBS07b, BCM08, BV08d, BV08e, BV08f, BV08g, BV08h, BKK^{+}09a, Ber07c, BCM09e, Bey07h, BBC^{+}08a, Bis08e, Bis08f, Bis08g, BBH^{+}09, BLS07e, BBD^{+}08, BBK^{+}09, BvdVP02e, BvdVP02f, BvdVP02g, Bow08d, Bra08d, BSS09e, BSWZ07, BF06f, BF06g, BF06h, BF06i, BPSS09a, CLM^{+}08e, CEKH^{+}07a, CDM^{+}08, Cer06g, Cer06h, Cer06i, Cer06j, Cer06k, Cer06l, Cer06m, Cer06n, Chi09j, Chl09c, Chu08c, CCG^{+}04, Cou07a, CFK^{+}08, DKM^{+}09, DH09u, DFPS08e, Don07c, EKMY02b, Eva08e, FJ04o, FHL^{+}08, Fre09j, Fre09k, Fre09l, Fre09m, Fre09n, Fre09o, FHBIK06, Gie07e, Gin09f, GSE07c, Gro07c, Gui09j, Gui09k, Gui09l, Gui09m, Gui09n, Gui09o, Gui09p, Gui09q, GLM^{+}09, HH06e, Han07e, Har08c, Hol09, Isa07e, Jon08j, Kut07c, LH09m, LH09n, LH09o, LCL07e, MY07f, MMO^{+}07e, MMO^{+}07f]. Front [MMO^{+}07g, MMO^{+}07h, Mas07e, MNS^{+}03, MK07i, MP09e, Mis08d, Moc09c, Mol07b, MOTV09d, Mon02j, Moo01b, NS08b, Ols08d, Osi07i, Pet08e, Pha07b, PR07d, Pri09h, QXZ09a, Ras07e, RV08, RS07b, Roh09b, Rot09b, RY09c, SR09d, Sch07q, Sch07r, Sch07s, Sch07t, Sch07u, Sie09c, Sim08d, Sla06h, Ste07f, Tai09g, Tak06p, Tak06q, Tak06r, Tak06s, Tak06t, Tak06u, Tel06f, Tel06g, Tel06h, Tia08f, Unt08d, Urb08c, Ves08d, Wei09h, Wil08i, Wit07d, dS01i, dS01j, dS01k, dS01l, dS01m, dS01n, dS01o, dS01p, dH09f, dH09g, dSW08b, vdH06c, Wak00a]. Frontier [JQQ04k]. ^{≫} [Mey02d]. Full [AK04h, AK04l]. Funchal [ADD^{+}03]. Function [Bis08d, Deb09, Mol07c, Mol07f, Mol07e, RY09f, Unt00q, dR02k, dR02a]. Functional [INP04, MS00c, MS03b, MS04, MS07, Pfl01i, Son03, Wen03h, dSW08d, Cap00b, KW04, Vil03, Jef04d]. Functionals [GM07, Pao03, IK00, Nem00c, Nem00d]. Functions [AHKM08g, BHGVN06, CL02b, Chu08e, Cla06b, DH09z, Gie07c, Gie07d, Gie07h, Gie07i, Gil03p, Gil03l, Gil03t, Gil03k, Grü02i, JL05h, JL05i, Kac06a, KV03, Koe03, KV00b, Let05g, Let05h, Li02g, MV06, NP01b, Pfl01e, Rot09g, Ryz09, Sta03, Ste07j, Ste07p, Tak06a, Tak06b, Tar02, dS01q, dSW08j, AGH03n, AGLT02e, CL02a, CL02c, KV00a, KV00e, KV00d, Lau09, MK07c, MK07e, MK07d, MK07f, MK07g, MK07h, Nem00e, Nem00f, NP01l, NP01p, Phe01b, Rit00b, Rit00c, Rit00f, Sim08j, Unt00c, Wag00, dR02c, Gil03q]. Functor [LH09w, KL01f, Wen03g, Wen03c]. Functors [CMZ02i, Fre09w, KL01b, KL01c, LH09g, LH09i, LH09j, Wen03h, CMZ02c, Fre09i, Fre09v, Fre09y, Fre09z, Fre09-29, Wen03e]. Fundamental [EFL01, Moc09j, Wei09i, Wei09n, Wei09o, Xia01c]. Further [Bra08a, Don07d, Hsu00e, Kun00b, Sla06i, Tak06v, Tia08g, vdDW08].

G [AAR03, HK03e]. G. [Gor07]. Gabor [FLC08, Kut07j, Kut07h]. Gain [JQQ04g, JQQ04i]. Galactic [AD08a]. Galois [Roh09j, Wei09b, Wei09f]. gambling [MS00a]. Game [Mie08c]. Gamma [Har08d]. Gap [Bob03b, Fou05, DH08, Kuk02, Teu03f]. Gas [Mer06, Rin03, WG00g, DH08, Fér08, WG00c]. Gas- [Rin03]. Gases [BPT^{+}08, BCG08]. Gaussian [ABCV05, BvdVP02o, Cha03, Che03, Dou03, Eis03, EE09, FSU01, Gui09u, Gui09x, Hor03, Hu00, Kho08, Kol00c, Mas07f, Mas07g, Naz03a, PT05, Sod07, Xia09b]. Gelfand [FLC08]. Genealogical [DD03b]. General [AT09e, AT09f, Bow08e, GL00d, Gro07d, Kie02c, Mer03, Ros04, Rot09e, Bil03b, Fas00c, JK04c, Kun00j, RY09d, Sim08h, Sim08n, Čhu02a, JQQ04c, Oli02i]. Générale [Del02d]. Generalised [Cox08, EM05]. Generalities [CMZ02a, FJ04a]. Generalization [Hab00c, Her01, KL01f, Pis01b]. Generalizations [Chi09s, Gui09r, GRKVL02f]. Generalized [ABCV05, BCG08, CMZ02i, Chi09k, Chi09l, Chi09q, DM03, ETZ07, FS00e, God08, MOTV09e, QXZ09h, Sus08, Voi00b, Woj07b, Yan06, Bil03i, FS00d, Yaf00c, AK04g]. Generated [AM02b, FBF^{+}02d]. Generating [dS01q, Unt00g]. Generation [Pat05, Čhu02b]. Generators [Bey07e, Enr07, Las04, Sus08]. Generators-Applications [Las04]. Généraux [Yor02a]. Genericity [AB01]. Genetics [Tav04, Tia08d]. genre [Wit07a, Wit07b, Wit07e]. Genus [Tak06x, Wei09l, JK04d, Wit07e]. geodesics [BSPS09d]. Geometric [AAFM07, BBS07d, BBS07c, Cao03i, CC08, Chi09m, Chi09v, GM03, GM04, GK04, JL05e, Kla07, Men04, Moc09d, Mon02k, Oli02n, Pfl01a, Pfl01d, Urb08d, ASWY07a, FBF^{+}02h, MS00b, Mon02b, NP01p, Paj02a, Cao03c, Fre08, GAC^{+}09, MS00c, MS03b, MS04, MS07, Tig04]. Geometric-Arithmetic [GM03]. Geometrical [Fus08, Cao03f, Let05e]. géométrie [Eme00e, NP01h]. Geometry [Agr08, AC02i, Ale04b, AD08c, AD08e, AD08d, AD08f, BBS07c, BBSW07, BM08b, Ber00b, Bjö04, BE00, BFL^{+}02, BPF^{+}02k, CM03, Con04, Das09, DL04, GRKVL02b, GRKVL02a, GRKVL02h, GL05g, GL05i, Gur09, Har08f, Iof09, Sch02f, Sib04g, Tum04, Van01d, Yan09b, YC04l, ZZ04, dS01u, Bru02e, Cao03b, GR03g, Gin09b, Moo01f, NP01f, NP01h, Sib04e, Sib04f, Van01e, Run02i]. Gerbes [Tha08]. Gibbs [AHKM08b, Bow08f, Cat04d, Har00, Pit06a]. Gilbert [AGLT02c, BCM09f]. Ginzburg [Son03]. Given [MP04, AF08a, ADD^{+}03, ACB^{+}03, AZ03, AAR03, BBH^{+}04, BBSW07, BGTT02, BM08b, BHK05, BKK^{+}09b, BG08, Bon08, BJS^{+}00, CM06a, Can03, CMI^{+}03, CT08, CO00, CMHL02, CCJ^{+}03, DR08, DL04, EKMY02a, Fas00a, GAC^{+}09, Mar07, MRR02, Mei03, NS08g, PV06, TDP08, ZZ04]. GK [Li02e]. GK-dimension [Li02e]. Glass [Gui07, LvdLM00]. Glasses [Bol07, BB07, NS07b, NS07c, She07, Tal07]. Global [CE05d, Gie07c, Gie07f, MRR02, Moo01c, NS07b, Tak06w, Wei09k, HN05m]. Glossary [Sch03e, Sim08e, YC04k, Cut02m]. Goals [Jon08n]. going [AGP09]. Golden [DF06]. Good [Wil08j]. Good- [Wil08j]. Goodbye [Wil08j]. Gorenstein [Chr00b, Chr00h]. Gradability [NV04i]. Graded [Li02d, Li02k, NV04b, NV04a, NV04h, NV04l, NV04c, Gre03h, NV04d]. Gradient [Pri09i]. Gradients [Pri09k]. Graft [Eva08l]. Grafting [Eva08k]. graining [Cer06c]. Granular [BPT^{+}08, BCG08, CMG08, GRMH08, Gio08, God08, Mar08b, Rug08]. Graph [BLS07f, BLS07i, FJ04f, Jon08e, Jon08k, Osi07s, Bur00]. Graphs [AGH03e, BLS07h, DD03a, Jon08a, Jon08b, Jon08s, Jon08t, Jon08g, Jon08f, Jon08l, Jon08r, Jon08q, Jon08u, Osi07f, AGH03f]. Grassmann [Har08i, Har08k, Har08o]. GREM [BK07]. Grid [vdH06d, vdH06e]. Grid-based [vdH06d, vdH06e]. Grobman [BV08m]. Gröbner [BHLV03i, Gre03c, Gre03h, Li02d, Li02b, Li02c, Li02k, BHLV03f]. Gromov [Abr08, Beh02b, Eva08f, Vak08b]. Grossissement [Yor02a, DGP00]. Grothendieck [Con00, LH09a, LH09l, Pis01g]. Ground [AHKM08c]. Group [Con04, EFL01, EP01, Füh05e, Füh05f, Füh05b, Gre03c, Har08j, Har08m, HG04, Hor03, Khe09, Kie02h, MMO^{+}07i, Mas03, Roh09d, Roh09j, Bru02b, CC08, GSE07h, Li02h, Oli02k, Ols03, Pis01d, Wei09f]. Groupe [Bia06]. Grouping [CLM^{+}08f]. Groupoid [LH09p]. Groups [AK01, ASWY07j, App05, CL02b, Chu08g, Cow08, DFPS08a, DFPS08f, DFPS08g, DFPS08j, Fra06, FR06b, Füh05c, GdS03d, GK04, HN05p, Hsu00k, Isa07h, Kar04b, Kas08, KM04c, Kie02j, Kie02l, Kie02c, Lee04e, Let05b, MP09g, Mon01c, NP01c, NP01d, NP01q, Rei04i, Som02n, dSW08f, dSW08g, dSW08h, dSW08j, ASWY07g, Bia03, BS03c, Bro04g, CL02c, Hab00d, Hsu00l, Kus05, Pil03i, Run02b]. Growth [BS04, Bil03j, CM03, Cha08a, Eva08k, Sie09f, BBS07d, Bil03f, Bil03c, Bil03b, Bil03d, Bob00, Tel06n, dR02j]. GSV [BSS09f]. Guide [Mey02a]. Gundy [FV08].

H [GO08]. Hahn [Har08m, Sim08f, Sim08c]. Halmos [Pis01k]. Hamilton [Cer01e]. Hamiltonian [Agr08, AJP06a, BHK05, BHLV03i, Efs05a, Efs05h, Han07g, Jon08q, Khe09, Kol00a, Kuk05, MMO^{+}07j, Oli02j, Van01a, Van01c, Van01b, dS01r]. Hammerstein [Chi09a]. Hankel [CM02]. Hankelian [Pis01h]. Hard [Ale04b, Wei09c]. hardening [FS00b]. Hardy [dR02a]. Harish [Gui09a]. Harish-Chandra [Gui09a]. Harmonic [AHKM08c, AHKM08b, CR07, CL02a, CL02b, CL02c, Cow08, DH09v, Füh05g, HN05d, Bia09b, MK07g, Ols03, Run02j, Ryz09]. harmoniques [Eme00d, God05, Mou00]. Harnack [Tel06m]. Hartman [BV08m]. Hasse [Wit07f]. Hausdorff [AGLT02d, Eva08f]. Haydon [Phe01e]. Heat [AGLT02k, BCE04, Lan09, AGLT02g, Kol00f, LvdLM00, Pri08b]. Heath [Fil01i]. Hecke [CMHL02, CM02, Chl09b, Chl09g, Lus02, Naz03b, RS07c, Unt00h]. Hedging [BJR04a, Pri09l]. Heegner [Bro04f, Bro04e, Bro04i, Bro04d, Bru02a, Bru02f]. Heights [Mas03]. Heinz [HK03f]. Heinz-type [HK03f]. Heisenberg [Füh05f, Yaf00d]. held [AF08a, ADD^{+}03, ACB^{+}03, AZ03, AAR03, BBH^{+}04, BBSW07, BGTT02, BM08b, BHK05, BKK^{+}09b, BG08, Bon08, BJS^{+}00, CM06a, Can03, CMI^{+}03, CT08, CO00, CMHL02, CCJ^{+}03, DR08, DL04, EKMY02a, Fas00a, GAC^{+}09, Mar07, MRR02, NS08g, PV06, TDP08, VY03, ZZ04]. Henselization [GR03e]. Hermite [Fit01]. Herz [Pis01h]. Hessian [Wan09]. Heterogeneous [EH08b, SX04]. Hierarchical [CLM^{+}08g]. High [Sod07, dH09m, LTJ00]. high-dimensional [LTJ00]. Higher [Ber09, Bou03c, Har08e, Ols08g, Pet08c, Pet08d, FSU01]. Hilbert [AHKM08g, App06, BV08l, BV08s, Bor03, Cer01e, Cer01h, Gil03e, Gil03p, Gro07e, HK03j, Kui03, Lee04d, NP01i, NP01p, PR07g, Sch07f, dR02g]. Hille [Gil03h, Gil03j]. History [AHKM08d, MY06h]. Hitting [YYY09, IK00]. HJM [Car07, Fil01g, Fil01d]. Hodge [Roh09d, Roh09j, Roh09l, Shi05]. Hofer [Sib04e]. Hölder [Urb08i, Nem00e]. Holomorphic [BSS09b, ST08b, Xia01d, Unt00o, Xia01g]. holonomic [Oli02f, Li02g]. Hom [Wen03e]. homaloidal [AC02f]. Homoclinic [Lom00i, Lom00h]. Homogeneous [Bis08d, Bis08j, Bis08h, Bis08n, DH09t, DH09v, DH09x, DH09y, Isa07f, Kut07d]. Homogenization [All08a, CS08, Mik00, NV04f, NV04e, Tar00]. Homological [BSS09g, Wen03b, GR03b]. Homology [Pfl01e, Fre09g, Rab01, Run02f]. Homomorphisms [Bey07i, Pis01f, Pis01i, Run02i]. Homotopical [MP09f]. homotopy [Fre09p, Fre09v]. hook [Naz03b]. hook-length [Naz03b]. Hopf [CMZ02b, CMZ02f, Don07d, Don07f, KL01a, Som02g, Som02h, Som02i, Som02j, Som02n, Wil08a]. Hörmander [HN05k]. Host [FL08]. houses [MS00a]. hp [Mel02c, Mel02i, Mel02b]. hp-FEM [Mel02b]. hp-Finite [Mel02i]. Hulls [Dol03b]. Human [GLW08]. Hunt [dR02n]. Hybrid [Chi09n, Ste03h, Mon02f, Ste03e]. Hydrodynamic [DR08, DGMR08, Hel08, SR09e, SR09b, SR09i]. Hydrodynamical [AMR03, Khe09]. Hydrodynamics [AF08b, Tos08]. Hydrogen [Efs05c]. Hydrology [SF08c]. Hyperbolic [Bey07b, Bre07, Eva08i, Gru08, Isa07h, JQQ04i, JQQ04h, QXZ09f, DIK00c, Mar07]. Hyperbolicity [Osi07j, Oli02g]. Hypercontractive [Led07b]. Hypercontractivity [Led03]. hyperelliptic [Ste03p]. HyperExt [LH09d]. Hyperfiniteness [KM04b]. Hyperfunctions [Wak00d, Wak00c]. Hypergeometric [Van03, Tar02]. Hypergroup [BH08]. hypergroupes [God05]. Hyperstable [JQQ04f]. Hypoelliptic [HN05s]. Hypoellipticity [HN05k, HN05q, HN05p]. Hypothesis [MMO^{+}07i, Ste07l].

Idea [Ves08h]. Ideal [Mer06, KZ02c]. Ideas [Tak06w, Tak06z]. Identification [ABCV05, CLM^{+}08n]. Identities [Doh03i, Wei09f]. IDS [Ves08e]. II [Cer01l, KM04b, Tai09m, AGH03f, AGH03g, AGH03h, AGH03i, AK04q, AJP06b, BHLV03c, Del02b, Gre03j, Gre03h, Gre03i, KZ02c, Kur01, Li02b, Mey02c, Nem00b, Nem00d, NP01d, Pet08d, Rei04c, Sae04d, SF06, Sne05b, Tak06b, Wei09o, Wer04, Wil08d, Zei04, dSW08f]. iid [Kho00]. III [KM04c, AGH03j, AGH03k, AK04r, AJP06c, Fri06d, Gre03k, KZ02d, Li02c, Sae04e, Tak06-30, Wil08e, dSW08g]. Ikeda [Osi07k]. Ill [Sch07v]. Ill-posed [Sch07v]. Image [Cao03i, LH09h, LH09w, Mos08a, Osi07t, Cao03a]. Images [CLM^{+}08d, LH09s]. imaginary [Lom00d]. Imaging [Amm09, AK09, Bon08, DBAP08, Dor08, Mos08b, Amm09]. Immiscible [Urb08e]. Immunity [NCCFM08]. Immunology [CM06b]. Impacts [LL07]. Impedance [SW09]. Implementation [Ves08h]. Implies [BS04]. Improved [Doh03i]. Including [MY06d]. Inclusion [Doh03i]. Inclusions [AK04i, AK04m, AK04p, AK04q, AK04r, Kun00j]. Income [Car07]. Incomplete [Bac04, Sch04a]. incompressible [SR09f, SR09g]. Increment [SF06, SF05, MK07h]. Increments [Büh05, Jac03]. indépendants [BL08]. Independence [AD08d, AD08h, Mas03, NP01c, NP01d, NP01i, NP01j, Wal03b, NP01b, NP01e, NP01l, NP01o]. Independent [Büh05, SF06, SF05, Gui09-29, Jac03, Voi00a, Voi00d]. Index [AC02k, AK04o, AGLT02i, Bil03l, BSS09f, BSS09g, BSS09m, BSS09p, BHLV03j, Bro04j, Bru02i, BPF^{+}02n, CMZ02j, Cao03j, Cat04h, CE05g, CL02e, Čhu02g, Cut00e, Dol03k, Efs05i, FJ04p, Fil01j, Füh05h, GR03j, GRKVL02g, GL02h, Grü02l, HN05t, Izh04b, KM04d, Kie02n, KZ02h, Let05i, Li02i, Mel02j, Mon02l, Oli02o, Paj02j, Pis01n, Run02n, Sae04b, Sch03e, Sch02g, Som02m, Ste03z, Vas02f, Wen03i, WG00i, Zha02f, dR02o, JK04g, Kun00g]. Indexed [AT09c, AT09g, Iva03, Mer03, Kho00]. Indices [BSS09h, BSS09i, JK04m]. Induction [Let05c, Let05f, Let05e]. Inductions [LH09s]. Inductive [Wen03f]. Industrial [BJS^{+}00, Fas00a, Bur00, Fas00b, Lio00]. inégalité [Mic03b]. Inégalités [LLP02, Ben03]. Inequalities [Bar04, BCE04, BLS07d, Chi09n, Chi09o, Chi09p, Doh03c, Doh03i, GZ03b, HN05f, HN05c, Kla07, Led01, Led03, Led07a, Mas07b, MP07, Mas07d, Mas07i, Mec07, NP01i, Wil08q, Zha02g, Zha02d, dR02i, Cat04f, Eis00, Gui09d, Gui09e, Hu00, Lau09, Mic03a, Nem00g, Tel06j, Vil03, Zha02a]. Inequality [BS04, Bob07b, FV08, GM03, QXZ09i, Cha03, DH08, Pis01b, Pis01g, SV00, Tel06m, Wag00]. Infectious [Bau08, JP08, Yan08]. Inference [Kub08, Mis08e, Tav04]. Infinite [AF08b, BNT06, BS04, BvdVP02p, Cer01i, Fil01b, FBF^{+}02d, MT08, Pau06, Yaf00e, Lom00b, Ols03, Gil03e]. Infinite-dimensional [BvdVP02p]. Infinitesimal [Ste03r]. Infinity [CE05d, CS05]. Influence [GRMH08]. Influenza [Bra08k, NCCFM08]. Information [AD08c, AD08e, AD08d, AD08g, Bac04, Büh05, JQQ04g, JQQ04i, KS06a, Mas07d]. Information-equivalence [Büh05]. Inhomogeneities [AK04s]. Inhomogeneous [BBS07d]. Initial [Tai09c]. Initial-Boundary [Tai09c]. injectivity [Chr00f, Mon01f]. Inner [AT09i, Xia01e]. Inner-Outer [Xia01e]. Input [Son08]. Insider [KH07]. Insight [Cha08a]. Instability [Bra08e, Hel08]. Institut [GO08]. Institute [VY03]. Insurance [Hip04, Pha07d]. Integer [Lyu04]. integers [FBF^{+}02j]. Integrability [CS05, Wil08r]. Integrable [Van01c, Van01b, Van01d, Van01a]. Integral [AHKM08d, AHKM08g, AHKM08h, AHKM08e, Ale04b, AMAF02, ABFM02, AM02b, BBS07c, Bün02, Chi09a, DIK00d, Gil03h, Paj02h, Pau06, PR07g, HK03b, Paj02d, Phe01k, Str03, Wak00b, Gil03f, Gil03i]. intégrale [Lac03]. Intégrales [CSM02, DDM02, Mey02e, Mey02b, Mey02c, Yor02c]. Integrals [AHKM08f, AHKM08l, BLM08, CS05, ETZ07, EM05, HN05c, KMS01, Lom00i, PT05, RS07j, Rot09e, Rot09g, Wil08p, Del02c, AH00, Gui09a, Gui09b, JW00, Lom00f]. integrands [Bil03c, Bil03d]. Integrated [Chy07, MY06a, Sha05, Sla06j, Ves08c, Ves08b]. Integration [DP06, JL05g, LCL07f, Mis08h, Mis08i, Naj07, Rit00b, Rit00c, Kol00f]. integrators [Mon02g]. Interacting [DD03b, Rez08b, DM00]. Interaction [GLW08]. Interactions [BvdVP02y, DFT08b, Xu08, Yaf00g, Yaf00p, Yaf00q]. Interactive [BvdVP02h]. Interchange [dR02m]. Interest [Bjö04, Fil01i]. Interface [Cer06o, DF05c, dH09c, dH09d]. Interfaces [ADD^{+}03]. Interior [BCM09g]. Interlude [Ste07g]. Interludium [Van01e]. intermediate [vdH06f]. Internal [NV04e]. Internazionale [BJS^{+}00, Fas00a, MRR02]. Interpolation [Dem08, Har08l, Har08n, HK03e]. Interpolations [Har08a]. interpretation [CC08, Pri08b]. interpretations [Hsu00b, GOW04b]. Intersection [Naj07, Ros05]. Intersections [Wit07e, Wit07e]. Intertwining [Bey07i]. interval [FBF^{+}02e]. Intervals [Glu03]. Intervention [GLW08]. Intrinsic [Urb08g]. Introduction [Ade01e, AHKM08j, AHKM08l, AF08b, All08b, AGH03b, AGLT02j, AM02a, AD08f, Bal04, BV08i, Beh02a, BCM09h, Bey07l, BBC^{+}08b, Bia08, BLS07g, BvdVP02i, BvdVP02j, BvdVP02q, Bol07, Bra08f, Bra08j, Bru02g, BSPS09b, CLM^{+}08h, Cat04i, Cel00, Cer01f, CE05h, Chr00i, CL02d, Chu08d, Čhu02h, Cou07b, Da 04, DIK00e, DH09w, Don07e, Ear08, Efs05g, Eme00c, Eva08g, FJ04q, Fla08, Fre09q, Fri05, FS00c, Gie07g, Goh04f, GdS03n, GS09a, GL00e, Gre03d, GSE07d, Gui09s, HH06f, Han07f, Har08f, Hsu00f, Isa07g, JQQ04j, Jon08m, Joy06a, Joy06b, Kel00d, KL01g, Kie02o, KZ02e, Kol00d, Kun00c, Kut07e, LM08, Lee04i, Lej03, Li02j, Lin05, LH09q, LH09r, LCFG01b, Lom00g, Lun04, MY07g, Mar00e, Mas07h, Mer03, Moc09e, MOTV09f, Mon01e, Mos08a, NP01j, Oli02p, Ols08e, Osi07l, Pfl01f, Phe01k, Pri09j, Ras07f, Rit00d, Roh09k, Roh09l, Ros04, Rot09c, RY09e, Sae04a, SR09h]. Introduction [Sim08g, Som02k, Ste03f, Ste07h, Ste03-27, Tai09h, Tel06i, Ter05, Tia08h, Unt00i, Unt08e, Urb08f, Van01f, Voi00c, Wil08f, WG00g, Zhi01b, dH09h, dSW08c, FBF^{+}02l, Lej09, Ols03, Tar00, AK04a, Bil03a, BHLV03a, Bro04a, Cut02a, Doh03a, Dol03a, Fil01a, Füh05a, GR03a, GOW04e, GL02a, Grü02a, HN05a, HK03a, Jef04a, JK04b, Let05a, Mel02a, Mon02a, Pil03a, Pis01a, Rei04a, Sch03a, Sch02a, Teu03a, Vas02a, Wen03a, WG00a, YC04a]. invariance [Fre09v, VY00]. Invariant [AHKM08k, Ale04a, CLM^{+}08i, Cut02r, Cut02q, Han07c, Hen05, JQQ04c, JL05f, Kel00e, Kol00e, Kut07j, Let05h, Moc09d, Osi07m, Osi07n, Zhi01c, Kar04b, Phe01c, Sib04d, Fil01f]. Invariants [Abr08, Ade01f, AS08, Beh02b, BM08b, Ber09, Moc09b, Moc09f, Moo01d, Roh09h, Tia02]. Invasion [Lol06]. Inverse [AGLT02k, BCE04, Eng00, Kie02f, LH09h, LH09s, Sch07w, AGLT02f, AGLT02g, Sch07e, Sch07f, Sch07d, Sch07c, AC02g, Bon08]. Inversion [Füh05d, Lom00h]. Invertibility [Gil03m]. invertible [FBF^{+}02d]. Investigations [Har08a]. investment [Rog03]. Investor [LL07]. Involutions [Kie02j, DIK00j]. involved [Hsu00l]. Involving [Bou04, Mis08g]. irreducible [Sch02d]. Irregular [LCL07b]. Irreversibility [JQQ04a, JQQ04b, JQQ04e, JQQ04c]. Irreversible [BO06]. Irrigability [BCM09i]. Irrigation [BCM09a]. Ishikawa [Ber07d]. Ising [Cer06p, Cer06s, CP06, Iof09]. Isogenous [DFPS08g]. Isolated [CE05b, CE05c, Ste03x]. Isometries [GL05h, Mol07c]. Isomorphic [LMTJ04]. Isomorphism [GL05k, KL01h]. Isomorphisms [GL05h, Som02e]. Isoperimetric [BS04, Bob07a, Tel06j]. Isotopy [ST08b, Kie02g]. Isotropic [BKM04, DBAP08, Pao00]. Isotropy [Bou03b, Izh04a]. Isotropy-Constant [Bou03b]. Israel [MS00c, MS03b, MS04, MS07]. Issues [CS08, MNS09]. Itô [Chy07, ETZ07, GM07, MY06a]. Italy [AF08a, ACB^{+}03, AZ03, AAR03, BBH^{+}04, BBSW07, BGTT02, BM08b, BHK05, BKK^{+}09b, BG08, Bon08, BJS^{+}00, CM06a, Can03, CMI^{+}03, CT08, CMHL02, CCJ^{+}03, DR08, DL04, EKMY02a, Fas00a, GAC^{+}09, Mar07, MRR02, NS08g, PV06, TDP08, ZZ04]. Iterated [BM08a, AH00, Ser00]. Iteration [Ber07b, Ber07d, Ber07g, Ber07h, Ber07i, Ber07j, Ber07l, Chi09g, Chi09q]. Iterations [Chi09m]. Iterative [Ber07e, Ber07f, Chi09r, Chi09s, Chi09t, Pet08f]. Itzykson [Gui09a]. IV [Ale07, AGH03l, AGH03n, AGH03o, AGH03m, Fri08, Li02d]. Iwahori [CMHL02]. Izhboldin [Kar04a, Kar04b, Mer04].

J. [Mar08d]. Jacobi [Cer01e, Lee04g, Lee04j]. Jacobian [AC02c]. January [Yor06]. Jarrow [Fil01i]. Jessen [KMS01]. John [Kla04]. John-Type [Kla04]. Joint [AP07, Jef04e, Ste07i, ADD^{+}03, CO00]. Jones [Paj02b]. Jordan [Ber00b]. Jorgensen [ASWY07g]. Julia [Kel00b, Kel00e]. July [ADD^{+}03, AZ03, AAR03, BBH^{+}04, BGTT02, BHK05, BKK^{+}09b, BG08, CMI^{+}03, CMHL02, Mar07, MRR02, PV06, VY03, ZZ04]. jump [Kol00k]. jump-diffusions [Kol00k]. Jumps [App06, Chy07, NT00]. June [AZ03, BGTT02, BM08b, BG08, BJS^{+}00, CO00, CMHL02, CCJ^{+}03, EKMY02a, GAC^{+}09, NS08g, TDP08, ZZ04].

Kac [DM00, RY09b]. Kadison [Pis01c]. Kähler [HH06g, dS01b, dS01s]. KAM [Eva08a, Han07k, Kuk02, Kun00h]. KAM-persistence [Kuk02]. Kantorovich [Bre03]. Kernel [Doh03d, Pau06]. Kernels [BH08, Bey07j, Lan09, Sch07h, Sch07i]. Kerov [Hor03]. Kikkawa [Kie02f]. Kinetic [BCG08, Cap08, Rez08b, Rin03]. Kleinian [EFL01, Hab00d]. Kloosterman [Unt00d]. Knight [Leu03b]. Knot [ASWY07j]. Knots [Ban09, KL09]. Knotting [Sum09b]. Kobayashi [Isa07h]. Kobayashi-Hyperbolic [Isa07h]. Kodaira [Shi05]. Kohn [HN05k]. Kollár [Ste03n]. Kolmogorov [Cer01h, Cer01g]. Koppinen [CMZ02b]. Korteweg [Zhi01d]. Krahn [BLS07d, BLS07h]. Krasnoselskij [Ber07g]. Krein [Phe01b, Phe01k]. Kriloff [How02]. Krylov [Bec06]. KZ [BF06e, BF06b]. KZ-Monads [BF06e, BF06b].

L [HK03e]. Labeled [ASWY07h]. labs [Deu00]. Lace [SP06, Sla06k, Sla06l]. Lacking [Piv07]. Ladder [Don07f, Har08g]. Lagrange [Sim08f]. Lagrangian [Oli02j, dS01t]. Laguerre [Mic03b]. Landau [Son03]. Landscape [BCM09j]. Langlands [Fre08, Wei09j]. Laplace [Cra08, HN05c]. Laplacian [BLS07h, BB09]. Laplacians [BLS07f, HN05s, HN05b, HN05f, HN05g, HN05c]. laplaciens [DDD02]. Large [Bis08h, Cer06q, GOW04e, Gui09u, Gui09v, Gui09t, JP08, JW00, LL07, NP01d, Pha07d, Cer06x, Cer06z, EK00, Fér08, GOW04g, GOW04c, Gui09w]. Large-Buffer [GOW04g]. Large-Scale [JP08]. Largest [Led03, Led07a]. Lattice [Bis09, Bou03c, Bou04, BF06j, Cho03, Sla06k, Sla06m, WG00g, DH08, WG00e, WG00c]. Lattice-gas [WG00c]. Lattice-Theoretic [BF06j]. lattices [DIK00d]. Law [AP07, BM08a, BLM08, EFL01, FS00b, Gui09u, Gui09t, Jac03]. Laws [Ade01a, Mar07, Ser00, YYY09, VY00]. Lawvere [BF06k]. Lax [Unt00k]. layer [Mar00b]. LDP [Cer06t, Cer06r, Cer06s]. Leading [Gro07f]. Learning [CP04, Mas07j, Men04]. Least [Sib04g]. Leaves [Eva08d]. Lebesgue [BCM09c, Chu08e, ETZ07]. Lecture [Amb03, BvdVP02k, BvdVP02m, BvdVP02n, BvdVP02l, BvdVP02o, BvdVP02p, BvdVP02q, BvdVP02r, BvdVP02s, BvdVP02t, How02, Mor08]. Lectures [AF08a, Abr08, ADD^{+}03, ACB^{+}03, AZ03, AAR03, BBH^{+}04, BBSW07, BBF^{+}03, BGTT02, BM08b, BHK05, BG08, Bon08, BJS^{+}00, CM06a, Can03, CMI^{+}03, CÇE^{+}04, CEKH^{+}07b, CT08, CO00, CMHL02, CCJ^{+}03, DR08, Dei03, DL04, EKMY02a, Fas00a, GO08, GZ03b, Gui09w, Hel08, Isa07h, Kuk05, Mar07, Mar08c, MRR02, Moo01d, NS08g, PV06, Phe01l, Rob04b, Run02o, Sei08, ST08b, Sol03, TDP08, TW04, ZZ04, dS01u, dH09o, BKK^{+}09b, Cat04j, GAC^{+}09, Ver03, ASTB03, BKK^{+}09b, Ber00a, Ber02, Pic05, TZP04]. Leeuw [Phe01f]. Lefschetz [Ale04b, AS08, Wei09c]. Left [Kie02f, LH09s, Kie02e]. Legendre [dS01v]. Leibniz [Goi01]. Lemma [MW03, Wei09n, Wei09o, MS00b, Wei09i]. Lemmas [Yan06]. Lemmata [Han07l]. length [BS03c, Naz03b]. lengths [RY09f]. Lens [Tig04]. Leuven [KV03]. Level [CLM^{+}08i, Dal03, DD03a, MY06h, Ols08f, Cao03d]. Lévy [App05, Fra06, SF06, BNT06, BLM08, Don05a, Don05b, Don07m, Don07e, Don07g, Don07j, Éme06, EM05, FG09, Hir03, KP05, KS07, NNY05, Pis05, Pis07, SY07, YYY09]. Lie [Ber00b, Cow08, DFPS08a, EP01, GdS03d, HH06h, Kas08, Lee04e, Let05b, Let05h, dSW08i]. Lieb [Bar04, BCE04]. Life [CL08, Yor06]. Lift [EFL01, Bru02d, Bru02c]. Light [Ear08, Unt00k]. light-cone [Unt00k]. like [Mar08d, Run02e]. Likelihood [Bis08a, Bis08k, BvdVP02s]. Limit [BK03, HH01, Hor03, Mim03, Pas04, Ste07k, Tsi04, SR09c, SR09f, SR09g, Wen03c]. Limiting [GL05f, Nus03]. Limits [Kaz06, PT05, Rez08b, SR09e, SR09b, SR09i]. Line [AK01, CLM^{+}08i, Fou05, Roh09f, dR02e, Lee04b]. Linear [Bec06, Bey07b, Bey07f, Bey07g, Bey07k, Bey07n, Bil03j, Bou03a, BF06b, Fal08b, GL02f, Joy06b, Kie02c, MY07h, MY07i, Mas03, Mol07f, Mol07e, Nes03, Pao03, Ras07g, Rit00e, Rit00f, Sch03i, Sie09d, Sie09f, Wal03b, dH09c, vdH06g, Bil03f, Bil03b, Bil03d, DM00, Del03, Eli02, FBF^{+}02e, Rit00g, Rit00i, BPF^{+}02f]. linear/superlinear [Bil03f]. linearization [Yoc02]. lines [RK00]. Linking [Ber09]. Links [Cat04b]. Lins [Lac08]. Lipschitz [BV08j, Cer01j, Cer01k, Cer01l, Chi09b, Chi09h, Chi09k, Chi09l, Ves08e]. List [Cat04j, CL02e, Run02p, Sae04b, Tak06x, Teu03i, Vas02f]. Littlewood [DH09x, Wil08r, dR02a]. Living [FL08]. LLD [CLM^{+}08b, CLM^{+}08j]. Loading [GRMH08]. Local [Bis08i, Bis08j, BSS09j, Chy07, Deb09, Don07h, Eis07, ETZ07, Han07g, Har08l, Hua04, JS04, KS07, LH09t, MY06a, Mol07c, Mou03a, Naj07, NS07b, Pes07, Pri09k, RS07d, Ros05, SG02, Sie09e, Sie09f, Tak06z, Wak00e, Wei09k, dS01x, dSW08d, Kaj08, Pir00, Pit06d, RK00, Str03, Tel06l, Tel06o, Cut02l, Let05d]. locale [SG02]. Locales [DDM02]. Localic [BF06l]. localised [Kol00i]. Locality [Sie09g]. Localization [Beh02b, Bou04, Gil03t, Gil03s, GdS03k, Ton09, Wei09e, Xia01b, BJ00, Gil03n, NV04h]. Locally [Füh05c, Kus05, MOTV09e, Mon01c, Rot09d, Rot09f, CL02c, Run02b, Wen03f].

Locking [JQQ04f]. locus [JK04i]. Loeb [Cut00a, Cut00f]. Loewner [Wer04]. Log [BN03, Bob07a, NS08e, NS08f, Cha03, Fou05]. Log-Concave [BN03, Bob07a, Fou05]. log-Sobolev [Cha03]. Logarithm [BM08a, Ser00]. Logarithmic [GZ03b, Led01, FS00b, Gui09d, Lau09, Mic03a]. logarithmique [Mic03b]. Logarithms [Nes03, Wal03b]. Logically [Mor08]. loi [Bro03]. Long [Bou03a, Yaf00g, Yaf00h, Yaf00f, CMZ02g, GOW04i]. Long-range [Yaf00g, Yaf00h, Yaf00f]. Long-Time [Bou03a]. look [Unt00a, Wil08a]. loop [BC03]. Loops [Kie02i, Kie02l, Kie02f, Kie02e, Kie02q, Kie02c, Kie02k]. Lorentz [Oli02k]. Lorentzian [GRKVL02c]. lossless [Cat04a]. Lost [Wal06]. Low [Cut02n, Izh04a, Mil07, dH09n, Hab00c, JK04m, JK04d]. Low-Dimensional [Mil07, Izh04a]. Lower [BvdVP02r, Gin09g, Gin09h, MP09g, MP09h, Tel06k]. Löwner [Zha02a]. Lusztig [GSE07b, Let05c, Let05f]. Lyapounov [Bou00]. Lyapunov [BV08k, BV08l, Gie07c, Gie07d, Gie07h, JQQ04i, JQQ04h, Kun00d, Osi07r, QXZ09e, Sie09e, Sie09f].

Maass [Unt00g]. Machlup [Cap00b]. MacLane [FJ04b]. Macroparasitic [MT08]. Macroscopic [CL08, GRMH08, Gui09w, Lac08, Dol03c]. Madeira [ADD^{+}03]. Magnetic [Das09, MMO^{+}07k, SW09]. Mahler [NP01l, NP01k]. Main [ASWY07k, Cut02i, Han07k, Jon08n, RS07g, Tai09h, Sch03b]. Majorization [Zha02b]. Malliavin [Nua09]. Management [Sha05]. Mandelbrot [Kel00c, Kel00e]. Mañé [Sib04b]. Manifold [Har08i, Har08k, Har08o]. Manifolds [Ale07, BSS09c, CT08, Fil01f, GRKVL02d, GRKVL02c, GRKVL02h, GRKVL02e, GdS03b, Isa07d, Isa07h, JQQ04e, KL01j, Osi07m, QXZ09m, Roh09a, Roh09e, Tum04, dS01b, dS01d, BV08b, BV08j, BV08n, BV08q, BV08p, Gin09d, Gin09g, Gin09h, Gin09j, Gin09k, Hab00b, Moo01f, Oli02a, Oli02e, Oli02c, Rei04d, Roh09h, Ste03v]. Manis [KZ02b, KZ02f]. Mann [Ber07h, Chi09g]. Many [Kie02j, Wil08k, GOW04h]. Many-Flows [GOW04h]. Map [Ale04c, MMO^{+}07l, Ste03c, AC02g, BHLV03c, Phe01p]. Mapping [Osi07k, Vas02c, Vas02d, Pil03i]. Mappings [AP06, Chi09b, Chi09e, Chi09c, Chi09d, Chi09g, Chi09h, Chi09i, Chi09k, Chi09l, Chi09r, Chi09q, Gre05, Mal03, Vas02g, YC04h, Pit06b, YC04j]. Maps [AC02i, Bou07a, BF06c, CR07, Chi09f, Chi09s, GOW04f, Goh04e, Gui09x, MOTV09a, MOTV09b, Pis01m, QXZ09g, Sae04g, AC02b, Pil03h, Pis01e, Pis01j]. Marche [Leu03a]. Marcinkiewicz [KMS01]. Marginals [Kla07, Sod07]. Margulis [QXZ09i, Val08]. Marked [Hob00]. Markets [Bau03, Car07, Sch04a, SX04, Rás09b]. Markov [BM03, BH08, Bel06, Cat04g, Da 04, Eis03, FR06a, Goh04b, Har08h, HH01, JQQ04a, JQQ04b, Kre09, Küm06, KP08, NT00, Reb06, Tai09e, Tia08c]. Markovian [AJP06b, Dha09, Wil08a]. Martensic [Dol03e]. Martina [ACB^{+}03, BBSW07, Bon08, BJS^{+}00, CM06a, Can03, CMI^{+}03, CMHL02, DL04, MRR02, ZZ04]. Martingale [App06, KMS01, KP05, BP01, BS03b, ÉY02a, Gal03, KS01a, Kaj08, Str03].

Martingale-Valued [App06]. Martingales [BM02, CSM02, Chy07, DDM02, Fit01, FV08, HH09, Kot06, KP01, LLP02, MY06a, MY06c, MY06d, Mer03, NNY05, Pha01, Pri09f, Pri09m, Ser07a, Yor02b, CC01, Eme00a, Eme00b, Han03, Kur01, VY00, Yor02a]. Mascali [AAR03]. Mass [All08a, BPSS09b, Hab00g]. Masses [BD03]. Master [Moc09d, Moc09g]. Matches [CLM^{+}08f, CLM^{+}08j]. Matchings [Jon08o]. Matematico [BJS^{+}00, Fas00a, MRR02]. Material [Xia01c, Hsu00a]. Materials [Rug08, FS00b]. Mathematical [AHKM08l, AT09a, Arn08, AD08g, BBF^{+}03, Bau08, BCM09k, Ber05, Bey07l, Bol07, Bor05, BFET05, BvdDW08, Cap00a, CMG08, CÇE^{+}04, CEKH^{+}07b, CM06b, Cha08a, Fri06c, Fri06d, Fri08, JQQ04k, Kot09, Lol06, Lou08b, NCCFM08, Osi07e, Pat05, Růž00d, Růž00a, SR09i, Sin08b, Sne05b, Utk08b, VY03, ADD^{+}03, Amm09, AGLT02a, AAR03, CM06a, Cut00d]. Mathematicians [Wal08b]. Mathematics [BJS^{+}00]. Mather [Oli02m, Sib04a, Sib04b]. Matlab [Gau06]. Matrices [Gil03o, Gil03j, Gil03m, Gil03e, Gil03n, Gil03g, Gui09w, Jef04e, Led03, PL03, Spe03, AC02e, AGP09, CC08, Dou03, Gui09c, Gui09e, Gui09u, Gui09t, Gui09z, Gui09-27, Gui09-29, Kho08, Lom00e, O’C03, Riv02]. Matrix [Bia09a, BHGVN06, Chu08e, Chu08f, Chu08g, Dei03, Gil03l, Zha02g, Gui09b, Yaf00h, Yaf00j, Yaf00l]. Matrix-Valued [Gil03l]. Matroids [Jon08i]. Matter [AMTV08, AMS^{+}08a, ASWY07b, ASWY07f, AHKM08a, AHKM08i, AFS08, AADH08, ACC^{+}08, AT09b, AT09d, AD08b, AMS^{+}08b, BBS07a, BBS07b, BJS08a, BCM08, BV08a, BV08d, BV08e, BV08f, BV08g, BV08h, BKK^{+}09a, Ber07a, Ber07c, BCM09b, BCM09e, Bey07a, Bey07h, BBC^{+}08a, Bis08b, Bis08e, Bis08f, Bis08g, BBH^{+}09, BLS07a, BLS07e, BBD^{+}08, BBK^{+}09, BvdVP02a, BvdVP02b, BvdVP02c, BvdVP02e, BvdVP02f, BvdVP02g, Bow08b, Bow08d, Bra08b, Bra08d, BSS09a, BSS09e, BSWZ07, BF06a, BF06f, BF06g, BF06h, BF06i, BSPS09a, BPSS09a, CMRT09a, CLM^{+}08a, CLM^{+}08e, CMG08, CEKH^{+}07a, CDM^{+}08, Cer06a, Cer06g, Cer06h, Cer06i, Cer06j, Cer06k, Cer06l, Cer06m, Cer06n, Chi09j, Chl09a, Chl09c, Chu08a, Chu08c, Cla06a, CCG^{+}04, Cou07a, CFK^{+}08, DKM^{+}09, DFT08a, DF05a, DH09r, DH09u, DFPS08b, DFPS08e, Don07a, Don07c, EH08a, EKMY02b, Eva08c, Eva08e, Fal08a, FJ04o, FHL^{+}08, Fre09c]. Matter [Fre09d, Fre09e, Fre09f, Fre09j, Fre09k, Fre09l, Fre09m, Fre09n, Fre09o, FHBIK06, Fri06b, Gie07b, Gie07e, Gin09a, Gin09f, God08, GSE07a, GSE07c, Gro07a, Gro07c, Gui09j, Gui09k, Gui09l, Gui09m, Gui09n, Gui09o, Gui09p, Gui09q, GLM^{+}09, HH06b, HH06e, Han07a, Han07e, Har08b, Har08c, Hol09, Isa07a, Isa07e, Jon08j, Kac06b, Kon08a, Kut07b, Kut07c, LB09a, LH09b, LH09c, LH09m, LH09n, LH09o, Lou08a, LCL07a, LCL07e, MY07b, MY07f, Mar08a, Mar08b, MY02a, MMO^{+}07a, MMO^{+}07e, MMO^{+}07f, MMO^{+}07g, MMO^{+}07h, Mas08a, Mas07a, Mas07e, MNS^{+}03, MK07a, MK07i, Mie08a, MP09b, MP09e, Mis08a, Mis08d, Moc09a, Moc09c, Mol07a, Mol07b, MOTV09c, MOTV09d, Mon02i, Mon02j, Moo01a, Moo01b, NS08a, NS08b, NS07a, Ols08a, Ols08d, Osi07b, Osi07i, Pet08a, Pet08e, Pha07a, Pha07b, PR07a, PR07d, Pri09e, Pri09h, QXZ09a, QXZ09d, Ras07b]. Matter [Ras07e, Rás09a, Rez08a, RV08, Rob04a, RS07a, RS07b, Roh09c, Roh09b, Rot09a, Rot09b, RY09a, RY09c, SR09a, SR09d, SF08a, Sch07g, Sch07q, Sch07r, Sch07s, Sch07t, Sch07u, ST08a, Sie09a, Sie09c, Sim08a, Sim08d, Sin08a, Sla06a, Sla06h, SV09a, Ste07a, Ste07f, Sum09a, Tai09d, Tai09g, Tak06c, Tak06p, Tak06q, Tak06r, Tak06s, Tak06t, Tak06u, Tel06b, Tel06f, Tel06g, Tel06h, Tia08a, Tia08f, Tof08a, Uka06b, Unt08b, Unt08d, Urb08a, Urb08c, Utk08a, Vak08a, Ves08a, Ves08d, Wal03a, Wal08a, Wei09d, Wei09h, Wil08b, Wil08i, Wit07c, Wit07d, Woj07a, WL08a, Xia09a, Yan09a, Zum07a, dS01i, dS01j, dS01k, dS01l, dS01m, dS01n, dS01o, dS01p, dH09a, dH09f, dH09g, dSW08a, dSW08b, vdH06b, vdH06c]. Maximal [GGM07, KW04, Mas07i, Naz03a, Roh09m, STJV03, dR02d, dR02a, HN05q, HN05b, dR02i]. Maximality [Obł07]. maximally [Sim08m]. Maximisation [Sch04a]. Maximum [Bis08a, Bis08k, BvdVP02s, Deb09, Gui09v, Sus08, Cha03, Hir03]. Maxwell [BCG08, Dem08, Oli02l]. Mean [BBS07d, BPF^{+}02i, DD03a, GM03, Tal07, Cat04c, GMR00, Sch07-27]. Meaningful [CLM^{+}08d, CLM^{+}08f, CLM^{+}08j]. Means [HK03j, HK03g, HK03d, HK03e, HK03f, HK03c]. Measure [DMH03, Hor03, JQQ04c, Naz03a, dRV00, Bia09b, BvdVP02u, Che03, Fér08, FSU01, Gui09u, Gui09t, Har00, Paj02a, Pri01, RY09h, JQQ04d]. Measure-theoretic [JQQ04d]. Measurements [AK04s, Bar06]. Measures [App06, BCM09q, BN03, Bob03b, Bow08f, Bui05, Cut00a, Cut00f, Del06, FJ04g, Fou05, Har08d, JL05g, NP01l, Pen04, QXZ09f, QXZ09l, Rot09e, Rot09f, Rot09g, Sch03f, Sod07, Wal03b, Xia01f, Che03, Hu00, KS01a, Phe01c, Phe01o, Phe01q, Zhi01c]. mechanical [Mon02h, Oli02f, Oli02e]. Mechanics [AJPP06, Bar06, Cut00b, Joy06a, Oli02n, BKK^{+}09b, Oli02g, Oli02d, WG00d]. Mechanisms [WHG05]. Media [Bol07, DBAP08, EH08b, Fas00a, Gio08, SF08c, Urb08b, Fas00c, Mik00, Teu03e]. Medical [AK09]. medicine [Deu00]. Melnikov [Kun00e]. Memoriam [ÉY06]. Mendelian [Tia08d]. Menger [Paj02h, Paj02c]. meromorphic [KV00d]. Mesurable [Del02e]. mesures [BS03a]. Metamorphoses [Efs05h]. Metaplectic [Unt08f, Bru02b]. Metapopulations [MT08]. Metastability [Bov09, dH09o]. Method [AK09, CLM^{+}08b, CLM^{+}08m, Chi09n, Chi09r, Dor08, FJ04b, Gie07a, Gro07h, MY07j, MY07k, Osi07p, Rei04i, Sch07w, Urb08g, AGLT02b, Kun00e, Phe01j, Phe01m, Ste03k, Tar00, vdH06h, BHLV03c, BHLV03b]. Methodology [Fil01d]. Methods [AF08b, BBH^{+}04, Bec06, Bra03, Can03, Chi09s, Chi09t, Dah03, Dol03m, Dur08, Fal08b, For04, Fri06a, FS00e, Gat00, Gau06, GO08, INP04, Kot09, Kuk05, LM08, MY07l, Mel02i, NV04l, Pet08c, Pet08d, Pet08g, Pet08h, Pha07d, Ste03e, Ste03h, Tig04, ZZ04, Cao03e, Rás09b, Rit00g, Sch07v]. Metric [GRKVL02d, LMPTJ07, MOTV09e, Hab00b, NP01n, BPF^{+}02k]. Metrics [Hab00g]. metrizable [Phe01i]. Metropolis [DD03b, MR00]. Meyer [Yor06, Att06a, Bia06, ÉY06, Jak05]. microfunctions [Wak00a]. Microhypoellipticity [HN05b]. Microlocal [Wak00f, Wak00c].

Microparasites [FL08]. Microscopic [CL08, Lac08]. Microstructure [Dol03d, Dol03m]. Microstructure- [Dol03m]. Migration [NL08]. Milman [Phe01b, Phe01k]. Milnor [BSS09k, Goi01]. Mindlin [Fal08c]. Minicourse [KRA09]. Minimal [JQQ04c, MRR02, Mee02, Yan09b, GMR00, Gre03g, PR02, Sib04c, Sib04e, Sib04d, Sib04f]. Minimality [Cox08]. Minkowski [Bob07b, Kla00, SV00]. Minkowski-Type [Bob07b]. Miscellanea [MP09i]. Miscellaneous [Fre09r, Fre09s, Fre09t, Jon08p, Yaf00q]. Mixed [Bis08i, Dur08, Lee04j, Naz03b, BG08, Lee04a, Lee04c, Lee04e, Lee04d, Ste03d]. Möbius [Oli02k]. Mode [Utk08b]. Model [BJS08b, JL05g, MP07, Mas07c, Mas07f, BJ00, Cer06f, Cer06p, Dha09, Fre09i, Fre09u, Fre09-28, Yaf00d]. modèle [Yor06]. Modeling [Agu06, Amm09, AMR03, Bau03, BJR04b, Bor05, Bra08j, Bra08k, Cos08, MT08, Růž00e, Růž00a, Sne05a, Sne05b]. Modelling [AF08a, CM06b, Cha08a, Ear08, GRMH08, Lol06, Yan08]. Models [All08b, BBS07g, Bau08, BCM09d, BCM09h, BCM09k, BCM09n, Ber05, Bis09, Bjö04, BCG08, Bra08i, CMRT09b, CMG08, Car07, CP06, DGMR08, DD03b, DF05c, Fil01g, Fil01i, FL08, Fri06c, Gui07, Iof09, JK04a, KH07, LB08, MR08, NL08, NCCFM08, Rin03, Tal07, Ton09, Web08, WHG05, WG00e, WG00g, WL08b, Wu08, dH09p, vdD08a, vdD08b, Cap00a, EK00, JK04m, JK04k, JK04l, JK04d]. Moderate [GOW04j]. Moderately [LCL07b]. Modified [Xia01f]. Modular [Ade01d, Att06b, Lee04d, Unt00j, Unt00o, Unt08a, Unt08h, Wei09g, Bru02b, KL01f, Wei09l]. Modulation [Tof08b]. Module [CMZ02i, Bro04f, Bro04e]. Modules [Bro04i, Fre09v, Fre09w, Li02e, Li02g, Li02f, LH09j, NV04b, CMZ02b, CMZ02c, CMZ02f, CMZ02e, Fre09b, Fre09u, Fre09z, Fre09-31, GR03d, Gre03f, Gre03e, GSE07b, GSE07e, Mon01a, NV04c]. Moduli [DIK00f, Hab00g, Moc09b, Moc09g, Ols08c, Ols08g, Ols08h, Ols08i, Vak08b, Vas02g, Hab00d, Hab00e, Oko03, Vas02c, Vas02d, Vas02b, Vas02e]. Molecules [Efs05b]. mollifier [Sch07o]. Moment [AK04k, AGLT02k, AGLT02c, AGLT02d, AGLT02b]. moments [Gui09-27, Kho08]. Momentum [MMO^{+}07l, BHLV03c]. Monads [BF06e, BF06b]. Monge [Bre03, Caf03]. Monodromy [Efs05e, Hsu00g]. Monogenic [Jef04f]. Monoidal [Cut02s, KL01i, LH09e, Fre09x, Fre09-27, Fre09-28]. Monomialization [Cut02l, Cut02u, Cut02o, Cut02j, Cut02n]. Monotone [Bal04, Chu02i, Sim08h, Sim08i, SST09, Sim08m]. monotonic [Phe01b]. Monotonicity [Ath00, Lau09, Sim08c, Sim08j, AGH03j]. monster [Hsu00l]. Moore [Goi01]. Morozov [Gro07h]. Morphisms [Cut02u, Cut02n, Cut02o, GdS03l, GdS03c]. Morse [HN05f, Jon08h, Osi07o, Ras07h]. Morton [KS06a, Fil01i]. Mosaics [BBS07e]. Moser [dS01g, dS01w]. Most [Hir02, Bou00]. Motion [Bia09a, Cou07b, Mar08b, Mei03, Mis08c, Mis08f, Mis08g, Mis08e, Mis08i, Naj07, Pat05, Raj06, Sla06s, SV09b, BP01, Gal03, IK00, KS00, Nou08, RY09b]. Motions [BM08a, BBKM01, ET05]. Motivation [PR07e, Tel06l]. Motivations [Har08f, Tia08i]. Motives [Kar04b, Vis04]. Mouvement [RVY06, Lac03]. Mouvements [Eme00d]. Movements [GLW08]. Moving [Lun04, Unt00k]. MSER [CLM^{+}08j]. Much [BK07]. Multi [BF03, SW09]. Multi-Frequency [SW09]. Multidimensional [Wil07, Gal03, YC04d]. Multifractional [ABCV05]. Multifunctions [Sim08j, Sim08e, Sim08h, Sim08i, Sim08m]. multihomogène [NP01g]. Multihomogeneous [NP01g]. Multilevel [Bra03]. multimodular [AGH03n]. Multimodularity [AGH03a, AGH03c]. multiparticle [Yaf00l]. Multiple [CL07, PT05]. Multiplication [Roh09a, Roh09e, Roh09i, Voi00d, Bro04c]. Multiplicative [Ale07, Mol07f, Gil03b, KZ02c]. multiplicités [SG02]. multiplicities [SG02]. Multiplicity [NP01m, Ste03s, Li02f]. Multiplier [KW04, Shi05, Wil08h]. multipliers [Pis01h, Pis01g]. multiprojective [NP01h, NP01h]. Multiscale [CL08, Dah03, EH08b, Can03]. Multivariable [Glu04]. multivariate [Rit00b]. Mumford [Van01g]. Müntz [GL05d, GL05c, GL05i, GL05j, GL05k, GL05f, GL05g]. Muscle [Pat05]. Mutually [CL07]. My [Mer04]. Myocytes [Sha05].

Nagy [Pis01k]. Natural [Ban08, CJMS06, dSW08f, dSW08g, dSW08e]. Navier [Con06, Gal06, SR09g, Sin08b]. Near [AD08d, Lom00e, Sib04d, dH09c, dH09d]. Nearest [Mou03b]. Nearly [Bil03j]. Necessary [Han07l, Str03]. Neeman [LH09d]. Negative [Don07j, KP05, Pis05, Woj07b, KP08]. Neighborhood [dS01-28]. Neighborhoods [Tak06n, Tak06o]. Neighbors [Izh04b]. Neighbourhoods [AD08h]. Nekhoroshev [Ben05]. Nematic [Dol03c]. Networks [AGH03a, BCM09a, BCM09n, Bra08c, Bra08e, Bra08g, Bra08h, Bra08j, BPSS09b, Sam08, AGH03g, BSPS09c, AGH03k]. Neumann [Mal09, Pis01b, Ste03c]. Neural [Erm05]. Neuronal [Ter05]. Neurons [Fri05]. Neuroscience [BFET05]. Neutral [Del06]. Newforms [RS07d]. Newton [Osi07p, vdH06h]. Newtonian [Bil03i, FS00d, FS00e, Oli02d]. Nile [WL08b]. Nilpotency [DFPS08f]. Nilpotent [HN05r, HN05p, dSW08g, dSW08h]. No [CL07, KS05, AGH03k, KS01b]. No-arbitrage [KS05, KS01b]. Nodal [BLS07c, BLS07i]. Noether [AC02h]. Noetherian [LH09t]. Noise [JQQ04f, Mis08b, Tsi04]. Noises [Att06c]. noisy [AGP09, Rit00f]. Non [AHKM08h, Amb08, AM02b, AJPP06, CS08, Chi09i, DH09x, Gil03q, Hel08, Her01, Hir02, JM03, KL01j, Kun00f, LB09b, Mal03, MY06g, Pao00, RS07e, SST09, Sne05a, Tia08d, Unt00o, BS03b, Cer01k, Cer01l, DM00, Gin09d, Kah04, Kun00g, Kun00h, Kun00d, Kun00e, Kun00i, O’C03, Oli02f, Sie09g, Unt00e, Cat04c, Jon08q, Pis01d]. Non-Autonomous [JM03]. Non-Collapsing [Mal03]. non-colliding [O’C03]. Non-compact [Gil03q, Gin09d]. non-degenerate [Sie09g]. Non-Equilibrium [AJPP06]. Non-Ergodic [Her01]. Non-Euclidian [LB09b]. Non-Excitable [Sne05a]. Non-Hamiltonian [Jon08q]. Non-holomorphic [Unt00o]. non-holonomic [Oli02f]. non-linear [DM00]. non-Lipschitz [Cer01k, Cer01l]. Non-Mendelian [Tia08d].

Non-monotone [SST09]. Non-real [AM02b]. Non-self [Chi09i]. Non-Semisimple [KL01j]. Non-singular [AHKM08h]. Non-Smooth [Amb08, Kun00f, Kun00g, Kun00h, Kun00d, Kun00e, Kun00i]. Non-stopping [MY06g]. Non-supercuspidal [RS07e]. Non-symmetric [Hir02, Pao00]. Non-unitarizable [Pis01d]. non-unitary [Unt00e]. Nonautonomous [BV08r, Ras07a, Ras07h, Sch04b, BV08m, BV08s]. Noncoincident [FL08]. Noncommutative [Bia08, Cip08, Con04, Goh04g, Hor03, Li02k, Voi00e, Voi00a, Voi00d, DL04]. Noncommuting [Jef04d, Jef04i]. Nonconvective [HNKS06c, HNKS06f]. nonconvex [Kaw00]. Noncrossing [Jon08r, Voi00b]. Nonequilibrium [JQQ04k, MNS09]. Nonexpansive [Chi09e, Chi09c, Chi09d, Chi09i, Chi09r, Chi09s]. Nonholonomic [Mon02k, Mon02f, Mon02d, Mon02g, Mon02c]. Nonhomogeneous [Bis08a, Bis08i]. Nonlinear [Ber07f, CMZ02i, Chi09m, Cla06b, Coh03, DM08, HNKS06a, Kuk05, Mey06, MOTV09g, Pen04, Ras07i, Rit00g, Rit00h, Tur00b, Zhi01d, NS08g]. Nonlinearity [HNKS06g]. Nonreal [AMAF02, ABFM02]. Nonself [Gil03s]. Nonself-adjoint [Gil03s]. Nonselfadjoint [Gil03r]. nonsingular [GL00a]. Nontangential [Ryz09]. nonuniform [BV08o]. Norm [Bar03, MK07e, Gil03p, Zha02d]. Normal [Gil03d, Han07h, Pri09f, Pri09m]. Normality [Bis08i, Bis08j, BvdVP02n, DFPS08h]. normalization [BHLV03d]. normalizing [BHLV03g]. Normed [GL05l, Mel02d, Mel02f, Mel02e]. Norms [AP07, Bob00, Gil03l]. Norton [Hsu00h]. Notation [AC02k, Bil03k, Bru02i, CE05g, Dol03j, GL02h, Grü02m, MY06b, Pfl01g, Pis01n, Zhi01e]. Notations [Cut02m, Cut02p]. Note [BCN03, GM03, Gor07, KS07, Dou03, KS01b]. Notes [Amb03, FR06a, How02, Lus02, Mor08, Wag00, vdDW08, Kun00b]. Notion [ÉS01]. Notions [Ras07j, Bro03, FBF^{+}02b, Wen03b]. nous [Yor06]. Nuclei [Kie02h]. Nullstellensatz [NP01i]. Number [BSS09k, GGM07, PV06, Ste07j, vdDW08]. Numbers [Fri06a, JQQ04e, NP01n, Nes03, Sch03h, Wal03b, MS00b]. Numerical [Bec06, Can03, Cha08a, DD03a, EK00, Grü02j, LM08, MY07l, Mon02k, Rit00a, Sch07x, Cao03e].

Objects [LH09z, Pfl01d, Fre09-29]. Obnoxious [Tal07]. Observation [Gor07]. Observations [Her06g]. Observed [Bis08d, Bis08j]. obstacle [Yaf00e]. Obstruction [Ale04c, BSS09j, Moc09g]. Occupation [Doz03]. Ocone [VY00]. ODE [MT08]. Old [Yaf00m]. Oleg [Kar04b, Mer04]. Olfactory [Rei05]. One [BvdVP02l, BvdVP02v, Deb09, Don05a, Kot06, Naj07, Pha07c, Ras07c, Unt08g, YYY09, Dol03h, Lau09, Lom00e, Nou08, CE05b]. One-Dimensional [Kot06, Naj07, Pha07c, Unt08g, YYY09, BvdVP02v, Lau09]. One-Sided [Deb09, Don05a, CE05b]. One-step [BvdVP02l]. Onsager [Cap00b]. Open [AJP06a, AJP06b, AJP06c, BCM09l, BO06, DF06, MOTV09i, NS08f, Reb06, RB06, Run02q, Sim08k, FBF^{+}02g, JK04e, MY02b]. Openness [ASWY07i]. opérades [Cha01]. Operads [Fre09w, Fre09x, LCFG01a, Fre09b, Fre09g, Fre09p, Fre09t, Fre09-27, Fre09-28, Fre09z, Fre09-31]. opérateurs [Ben03, Mic03b]. Operational [Jef04g]. Operations [BCM09m, LH09u, vdH06i]. Operator [App06, Att06b, Ber07f, Bey07i, Dah03, DH09t, Gil03t, HH06a, HN05n, HN05o, HN05d, Mol07c, Mol07f, Rot09g, Run02m, Shi05, Xu08, dR02d, EK00, Paj02d, Sch07-27, Voi00e, Gil03g, Run02h]. Operator-Valued [Rot09g]. Operators [AT09g, AT09h, Bey07j, BB09, Bou03c, Bou04, Chi09t, Chu08f, Chu08g, DH09s, Doh03d, Gil03r, Gil03q, Gil03h, Gil03e, Gil03f, Gil03i, Gil03p, Gil03d, Gil03c, Gil03s, Gro07g, Gro07f, GL05j, HH06f, HH06j, HN05l, HN05r, HN05s, HN05k, HK03j, Jef04d, Jef04f, Jef04i, JL05f, Joy06b, LM08, Mol07e, Pau06, Pri09c, Pri09i, RS07c, RW08, Rös03, Tai09b, Tof08b, Ves08b, Ves08f, Wil08n, HK03h, HK03c, Wak00b, vdH06d]. Optical [Mos08b]. Optics [Bou08]. Optimal [Agr08, Amb03, BF03, BCM09j, BCM09n, BCM09o, BCM09r, BD03, BSPS09c, BSPS09d, BPSS09b, Caf03, ET07, Las04, MY07k, MMO^{+}07l, MMO^{+}07m, MMO^{+}07n, NS08g, Obł07, Pha07c, Pir00, Vil03, AP03, BSPS09g, Cer01b, Rog03, Tar00, AGH03h, AGH03i, ACB^{+}03, CO00]. optimisation [Bur00]. Optimization [AGH03c, CP04, JM00, Kaw00]. Option [Pri09l]. Optionelles [Yor02c]. Options [BF03]. Orbifolds [Abr08]. Orbit [KM04a, Kec04, MMO^{+}07m, Pfl01h]. Orbits [Han07d, Let05g, Lom00i, MMO^{+}07s, Lom00d, Unt00c, Unt00n]. Order [Ber09, Cer01i, HH06a, Pet08c, Pet08d, Som02n, Zha02a, FSU01, Gui09f, Gui09y, Rit00i, Teu03b, Čhu02c]. Order-Preserving [Čhu02c]. Ordered [Kie02l]. Ordering [AGH03o]. Orderings [Phe01o, vdH06j]. Orders [Lom00i, AGH03n]. Oriented [Sla06g, Sla06n]. Original [Ves08h]. Origins [Her06h]. Orlicz [Wil08l, Woj07b]. Ornstein [App06, Bis08o, EM05]. Orthogonal [BM03, Bec06, BHGVN06, Gau06, Har08a, Koe03, Kui03, Kuz06, Led07b, MV06, Ter06, Van03, KV03]. Oscillating [Mey06]. Oscillator [AHKM08c, AHKM08b, AHKM08f]. Oscillators [Erm05]. Oscillatory [Lom00i, Lom00f]. Osserman [GRKVL02d, GRKVL02f, GRKVL02c, GRKVL02b, GRKVL02a, GRKVL02h, GRKVL02e]. Osserman-Related [GRKVL02f]. Other [Ber07i, Gin09i, Roh09n, Eva08m, NP01k]. Outbreaks [Bau08]. Outer [Xia01e]. outline [ASWY07k]. ouverts [BL08]. Overcoming [KS06b]. Overview [Cut02o, Doh03a, Jon08m].

P. [Bia06, Paj02b]. P.-A. [Bia06]. Painlevé [BE00, Cla06b, Paj02e, Paj02g]. Pair [LH09e, Lom00d, dR02m]. Pairs [Moc09h, dR02l]. Paley [DH09x, Wil08r]. Pandemics [Bra08k]. Paper [Bia09a, Sch00]. Parabolic [KW04, Lun04, Moc09h, Mue09, Sch04b, Tel06m, DIK00b, DIK00c]. Paradoxical [Run02a]. Parameter [Bis08l, JQQ04b, ASWY07c, Sie09d]. Parameters [Bui05, Men04, STJV03, Kho00]. paramètre [Mic03b].

Parametric [Bis08m]. Paramodular [RS07f]. Paris [GO08, BBF^{+}03, CÇE^{+}04, CEKH^{+}07b]. Part [JL05c, JL05d, MK07j, Tai09l, Tai09m, MK07b, MK07c, Unt00s, Unt00t, AGH03d, AGH03f, AGH03g, AGH03l, AGH03j, AGH03k, AGH03h, AGH03i, AGH03n, AGH03o, AGH03m, AGH03e, AK04p, AK04q, AK04r, Gre03f, Gre03e, Gre03g, Gre03j, Gre03k, Gre03h, Gre03i, NP01c, NP01d, Pet08c, Pet08d, Sae04c, Sae04e, Sae04d, Tav04, Tsi04, Wer04, Zei04]. Partial [AT09h, AT09i, Bey07b, DM08, Eva08a, KRA09, Kho09, MY07h, Mas08b, Mue09, Nua09, PR07c, Wu08, Zha02a, dR02j, Ath00]. Partially [Las04]. participants [Cat04j]. Particle [BvdVP02d, Eva08j, Mou03b, Rez08b, DM00, Yaf00k]. Particles [CL07, NT01]. Particular [BBS07g, Del06]. partie [Wit07a, Wit07b]. Partitions [Vas02b, FBF^{+}02j, Pit06a, Pit06f, Pit06g, Pit06k, Voi00b]. past [HCG^{+}00]. Patch [vdD08b]. Path [AHKM08e, AHKM08f, AHKM08l, BvdVP02w, Doz03, Kol00f, LCL07h, MY06h, Raj06, Xia09b]. Paths [Lej03, LCL07c, LCL07d, LCL07f, LCL07g, RY09g, Bur00, Lej09]. Pattern [BCM09j]. Patterns [GLW08, Mey06]. Paul [Att06a, ÉY06, Yor06]. Paul-André [Att06a, ÉY06]. Pauli [AGP09]. PDE [Cer01i, Vil03]. PDEs [GAC^{+}09, Bis08c, Del03, Gur09, Kuk05, Urb08g]. Penalisation [Deb09]. Penalisations [RY09f, RVY06, RY09b, RY09d, RY09h]. Penalising [RY09g]. Penalization [DMH03]. Brixen [BBH^{+}04]. C.I.M.E [CO00]. Mathematical [LNH06]. superlinear [Bil03f]. Pencils [AS08, Wit07e]. Pendula [Efs05d]. pentagon [CMZ02f]. Percolation [CP06, Sla06g, Sla06f, Sla06n, Sla06o, Sla06p, Cer06b, Cer06r]. perfect [FS00f]. Perimeter [Naz03a, Cer06v]. Periodic [All08a, Han07d, Osi07q, Ste07d, Eli02, Kun00a, Lom00d, Lom00e, Teu03e]. Permissible [Cut02s]. Perron [BLS07h]. Persistence [JQQ04f, Kuk02]. Perspective [LNH06, She07]. perspectives [Cao03h, Sch07j, Sch07k, Sch07l, Sch07m]. Perturbation [Grü02k, Grü02a, Grü02e, Teu03g, Teu03c]. Perturbations [Eli02, Gil03r, Gil03d, Han07i, Mel02i, Yaf00g]. Perturbed [Mal09]. Perverse [Let05d]. Pesin [QXZ09j, QXZ09l]. Petersburg [VY03]. petit [Mic03b]. Petty [Mil07]. Pezzo [Wit07f]. Pfister [Izh04b]. Phase [Bis09, Cer06u, KS06b, Urb08h]. Phenomena [Bou03c, CM06b, Lom00i]. phenomenon [dRV00]. Phillips [Unt00k]. Phylogenetic [JP08, Kub08]. Physical [Ben05]. Physics [AAR03, AT09a, Kot09, Küm06, SF06, VY03]. Physiological [GLW08]. Physiology [Bor05]. Picard [Ber07j]. picture [ASWY07g]. Piecewise [FBF^{+}02e]. pinceaux [Wit07a, Wit07b, Wit07e]. Pinched [BS04]. Pinning [Ton09, dH09l, BvdVP02v]. pinning-depinning [BvdVP02v]. PIP [AT09c, AT09g, AT09h, AT09j, AT09k]. PIP-Space [AT09h]. PIP-Spaces [AT09c, AT09g, AT09j, AT09k]. Pisier [Wag00]. Pitman [Bia06, RVY06]. Planar [BHLV03b, Han07j, Kun00i, Wer04, Zum07b]. Plancherel [Füh05d, Füh05c, Hor03]. Planck [HN05i, HN05n, HN05r, HN05s, HN05o].

Plane [AC02i, Fin06, Ste07o, AC02b]. Plans [BF03, BCM09m, BCM09o, BCM09p, BCM09q, BCM09r]. Plastic [GRMH08, FS00b]. Plasticity [Ber05, FS00e, FS00a]. Plate [Fal08c]. Plücker [Ale04c]. PM [KZ02d]. PM-valuations [KZ02d]. Poincaré [BS03c, GO08, HN05c, LO00]. Point [Ade01b, Ade01g, AT09e, BBS07h, Ber07b, Ber07i, Ber07l, DFT08b, Las04, Mal03, MMO^{+}07l, Pet08f, Pet08g, Pet08h, Ste03s]. Points [Ber07e, Ber07k, CE05b, CE05c, Chi09b, Chi09e, Chi09c, Chi09d, Chi09r, DF05b, HB06, YYY09, KS00, Rei04b, Rei04c, Sib04d]. Pointwise [Wil08m, dR02p]. Poisson [AD08h, Del02c, DGP00, JL05h, JW00, Pit06g, Pri09a, Pri09k, Unt00f, Unt00h, Unt00p, Van01c]. Polar [BCN03]. Polarization [AK04g, Mos08b]. Polarization-Based [Mos08b]. Polarizations [Ols08g]. Polarized [Ols08i]. poles [KV00d]. polling [AGH03q]. Pólya [Bia09a]. polygon [vdH06h]. Polymer [dH09i, dH09j, BvdVP02u, Cap00a]. Polymers [dH09b, dH09e, dH09k, dH09q, dH09l, dH09m, dH09n]. Polynomial [Bar03, Dem08, FBF^{+}02f, GSE07g, GSE07f, HN05q, Led07b, Li02e, Li02g, Li02f, Li02h, Ste07c, Tel06n, YC04h]. Polynomials [Ade01f, BM03, Bec06, Gau06, GL05i, Har08a, KV03, Koe03, Kui03, Kuz06, MV06, NP01i, Ter06, Van03, Bia09b, Bob00, Gui09a, Pit06a]. Polytopes [BBS07f, GGM07, SW03]. Ponctuels [CSM02]. Population [AdlPP^{+}08, LB08, Lou08b, MR08, Mie08c, Tav04, Web08]. Populations [FL08]. Porous [Fas00a, SF08c, Urb08b, Fas00c, Mik00]. Portfolios [ET07]. Portugal [ADD^{+}03, CO00]. posed [Sch07v]. Posedness [AK04b]. Position [Web08]. Positive [Bou00, Don07g, Fag06, Goh04e, Gre05, FBF^{+}02j]. Positivity [Ban08, Bis09, Reb06]. Postface [And06]. Postponed [JK04j]. Potential [AHKM08e, AGLT02k, BB09, Bou03a, Bou03c, Fin06, FS08, GS09b, Pis05, Pri08a, SV09b, Tur00b, Tur00c, dH09k, AGLT02f, Tel06p, Tur00a]. Potential-theoretical [Pis05]. potentials [Yaf00h]. pour [BS03a, Bro03, Goi01, MR00, Wit07f, Yor06]. power [SV00, Cut02t]. Powers [MP09a]. Practice [Cut00f]. Prandtl [FS00b]. Pre [Ber07k]. Pre-Requisites [Ber07k]. preceding [Sch00]. Predictable [CS05, Jak05, MY06d]. Preface [Nem00h]. Preimages [Mal03]. Preliminaries [Grü02b, Hab00f, Moc09i, Moo01e, NS08c, Ols08j, Pit06h, QXZ09k, Roh09o, Tur00d, AGLT02a, Tel06c, AC02a, Bro04b, Doh03b, Gil03a, Kie02a, Pil03c, Som02a, Ste03a]. Preliminary [HNKS06d]. première [Wit07a]. Preparation [Tak06n, dS01x]. Prerequisites [Bey07m, DFPS08i, Chr00c]. Prescription [Gin09j]. Presentation [KL01k, LLP02, ÉY02b, Gre03j]. Presentations [Hsu00i]. Preserver [Mol07f, Mol07e]. Preservers [Mol07d]. Preserving [Čhu02c]. pressing [LvdLM00]. preuve [Bai08]. previously [CC08]. Pricing [Bac04]. Prime [Fri06a, Som02n]. Primer [Kho09, Ols08b]. Primes [Ade01b]. Primitive [Hir02]. Princeton [BBF^{+}03, CÇE^{+}04, CEKH^{+}07b]. Principal [DMH03, Mel02b, AC02f, Unt00e, Unt00l]. Principally [Ols08i]. Principe [Wit07f]. Principle [DGMR08, NT01, Obł07, Sib04g, Sus08, Gui09t, RY09d]. Principles [Her06c, dS01-27, Cer06x, DM03]. Prion [MT08]. Priori [Tai09i, Urb08j]. Probabilistic [Cer01i, Del03, Pri08b]. Probabilités [ASTB03, ALÉY00, AÉLY01, AÉLY03a, AÉLY03b, Ber00a, Ber02, Bra08h, CP04, CP06, DMÉRS07, DMÉRS08, DMÉRS09, Don07m, ÉLY05, ÉY06, Eva08h, Gui09w, LCL07d, MP07, Pic05, PP06, SP06, TZP04, TW04, dH09q, ÉY02a]. Probabilities [DMÉRS09, Sie09b]. Probability [ASTB03, Ber00a, Ber02, BN03, Bob03b, Bob07a, Fou05, GM04, GL00c, JQQ04k, Pic05, Pri08a, SF06, SF05, Spe03, Ste07g, TZP04, Xu08, GL00a, GL00b, GL00d, Gui09h, Kol00k, TW04, Voi00e, Eva08h]. Problem [Amb08, AK04f, AK04j, BCM09f, BBF08, Bou03b, BSPS09e, GL05k, MMO^{+}07o, Mil07, Paj02g, Pha07c, Pis01c, ST08b, ST03, Urb08h, AGLT02c, AGLT02d, Bal00, BP01, Bor03, BSPS09f, Hsu00i, Hsu00n, Kol00a, Paj02e, Pis01k, Pis01i, Sch07z, Sim08n, Unt00n, Yaf00n]. problème [DDD02]. problèmes [BL08]. Problems [Agr08, Amb03, AGLT02k, AAR03, BCM09l, Bil03j, Bol07, Bon08, BJS^{+}00, BD03, CS08, Can03, CL08, Fil01i, Fin06, FS00e, Gro07f, Hel08, Hua04, Kaz06, Las04, Lou08b, Lun04, Mal09, Mel02b, Mol07f, Mol07e, MOTV09i, Obł07, Pis01m, Pis05, Pis07, Rei04i, Rit00a, Sol03, Son03, Ste03c, Tai09a, Tai09c, Tai09e, Tal07, Tar09, Vas02c, Vas02d, Yaf00m, AGLT02b, AGLT02f, AGLT02g, Bil03e, Bil03b, Bil03d, Bur00, Cer01b, Eng00, Fas00b, FBF^{+}02g, FS00a, FS00f, Kaw00, Lio00, MY02b, Rei04e, Rei04f, Rei04d, Rit00e, Rit00f, Rit00g, Rit00h, Rit00i, Rog03, Run02q, Sch07v, Sim08k, Zhi01g]. Procedures [Ber07b, Ber07i, Ber07l, Mim03]. Process [Bis08o, BO06, GZ03a, Jak05, JQQ04c, Kho03, Sla06c, Hir03, RY09f]. Processes [App05, App06, ABCV05, BBS07d, BBS07h, BNT06, Bel06, BLM08, BvdVP02m, BvdVP02n, Bou08, CM03, CMI^{+}03, CM08, Chi09q, CJMS06, Dem09, Don05a, Don05b, Don07m, Don07e, Don07g, Don07f, Don07j, Éme06, Enr07, EM05, EE09, Fil01c, Fra06, GM07, Goh04b, Goh04g, GS09b, Gre05, Iva03, JQQ04e, JQQ04d, Kol00g, Kot06, Kul09, Küm06, KP05, KS07, Mas07g, Mis08f, NNY05, Pet08f, Pis05, Pis07, PP06, SY07, SF06, SF05, Tai09e, Yan08, YYY09, Zak05, AGH03l, Bas03, Cap00a, Cap00b, Eis00, Fas00b, Jac03, Kol00i, KP08, NT00, O’C03, Pit06d, Pit06e]. Processing [AT09j, Cao03i, Cao03a, Str00]. Processus [CSM02, Del02c, Del02d, Mey02a]. Producing [Fri06a]. Product [AT09i, Bha05, Bou07b, CR07, MMO^{+}07b, Unt00m, Unt00s, Unt00t]. Production [JQQ04a, JQQ04b, JQQ04g, JQQ04i, JQQ04e, JQQ04c, JQQ04d, Vil08]. Products [Bru02a, GdS03j, GdS03i, MMO^{+}07t, NV04g, Ste07c, Fre09y, Fre09-30, LvdLM00, Run02k, Unt00h, Unt00p, Van01b]. Profile [BvdVP02s]. Profiles [Nus03, Ser07b]. Program [Fre08]. Progress [DR08]. Progression [GLW08]. Progressivement [Del02e]. Projection [Art04, Dem08, GL05k, Ste03k]. Projection-Based [Dem08]. Projections [BPF^{+}02l, Piv07, GL00f, Mic03a]. Projective [JK04a, Osi07r, Roh09f, GR03d, JK04m, Wen03c, Wen03d, BPF^{+}02g, JK04k, JK04l, JK04d]. projectivity [Chr00g]. Proliferation [Fri06d, MT08]. Proof [Cao03g, Cut02o, Han07k, HN05k, Jon08n, KMS01, Mar08d, Pis04, Tai09l, Tai09m, Tai09j, Tai09k, Yan06, ASWY07k, FSU01, Yaf00a, Cut02i]. Proofs [Han07l, RS07g, JK04j, NP01o]. Propagation [LB09b]. Proper [Isa07i]. Properly [PR02]. Properties [Ale07, Arn08, Bel06, BLM08, BvdVP02w, Chi09m, Chi09v, Doz03, GK04, GL05a, GL05c, HN05c, HH01, Jef04i, Jon08e, Jon08d, Jon08k, MY06d, Phe01p, Sch03f, Tof08b, Ves08b, Xia09b, AGP09, BV08c, BSPS09g, Dha09, Fre09a, FS00a, GL00d, Run02e, Run02l, VY00]. Property [BH08, BK03, Éme01, KS05, Kie02f, Kut07d, LH09t, Pha07c, QXZ09e, Woj07b, Ath00, KS03, Str03]. propres [PL03]. Propriété [God05]. Prospects [DR08]. Proteolysis [Lol06]. Prüfer [KZ02b, KZ02f]. Prune [Eva08l]. Pseudo [Chi09b, Chi09g, Chi09h, Chi09k, Chi09l, HN05m, RW08, ST08b, Tai09b, Tof08b, Xia01g, Oli02c]. Pseudo-contractive [Chi09b, Chi09g, Chi09h, Chi09k, Chi09l]. Pseudo-Differential [RW08, Tai09b, Tof08b, HN05m]. Pseudo-Holomorphic [ST08b, Xia01g]. Pseudo-Riemannian [Oli02c]. Pseudodifferential [LM08, Unt08a, Unt08g, Unt08h, Wak00b]. Pseudofunctors [LH09e]. Pseudostable [Ole03]. PSI [Bou03b]. PSI-2 [Bou03b]. Puiseux [FJ04d]. Punctured [ASWY07j, ASWY07g]. Pure [Har08j, BS03b]. purely [Han03, Lom00d]. pushout [Fre09y]. pushout-products [Fre09y].

Quadratic [AHKM08h, Cut02r, JL05g, Tig04, Unt00n, Vis04, DIK00d, Kaj08, SST09, Unt00c, Efs05d]. Quadrature [Gau06]. Quadric [Li02e, Li02g, Li02f, Li02h]. Quadrics [Kar04a, Kar04b, Vis04, Wit07e]. quadriques [Kah04, Wit07e]. Qualitative [FR06a, Kut07f, Zhi01d]. Quantique [Bia06]. Quantitative [Kut07g, YC04i]. Quantities [Ber09]. Quantization [GL00c, HH06h, RW08, Unt00o, Unt00p, GL00a, GL00b, GL00d]. Quantum [AHKM08d, Arn08, AJPP06, AJP06a, AJP06b, AJP06c, Att06c, Bar06, BO06, Bou07a, Bou08, Con04, DGMR08, DF06, FR06a, Fag06, Fra06, FR06b, FS08, Har08a, Har08m, Iof09, Joy06a, KL01j, KL01e, Kon08b, Küm06, Lin05, Mer06, Mol07d, NT01, Pil06, Pri01, Reb06, Rob04b, SF08b, SF06, SF05, Teu03g, Wal08b, Xu08, AH00, CMZ02e, Kus05, AF08a, BGTT02, Teu03e]. Quarantine [NCCFM08]. Quasi [AK01, AHKM08k, Bey07b, Bey07g, Bey07n, Del03, FS00d, HH01, LH09v, LMTJ04, Eli02, Oli02l]. Quasi-Abelian [AK01]. Quasi-Coherent [LH09v]. Quasi-Compactness [HH01]. Quasi-Convex [LMTJ04]. Quasi-free [AHKM08k]. Quasi-Linear [Bey07b, Bey07g, Bey07n, Del03]. Quasi-Maxwell [Oli02l]. quasi-periodic [Eli02]. Quasi-static [FS00d]. Quasiconformal [Vas02d, Vas02g]. quasifuchsian [ASWY07g]. quasy [Gre03a]. quasy-continuity [Gre03a]. Quaternionic [Ale04c]. Quaternions [BFL^{+}02, BPF^{+}02f, BPF^{+}02a]. Quelques [RVY06, Yor02d]. Question [Bia06]. Questions [Ale04b, Don07b, GL05g, JK04e, RY09d]. Queue [GOW04d, GOW04a]. Queueing [Bra08e, Bra08g, Bra08h, GOW04f]. queues [AGH03l, AGH03i, GOW04l, O’C03]. queuing [AGH03g]. Quilts [Hsu00j, Hsu00k, Hsu00l, Hsu00b, Hsu00c, Hsu00d, Hsu00h, Hsu00m]. Quotient [Ste03j]. Quotients [LMTJ04, Mal01, GdS03d].

Radial [Dem09, Gie07c, Gie07i]. Radon [Unt00l, Unt00r]. Rainwater [Phe01e]. Ramanujan [NP01b, Wei09l]. Ramifications [Fre08]. Ramified [Han07i]. ramifiée [Kah04]. Ran [MW03]. Random [AP06, Art04, BBS07e, BBS07f, BBS07g, Bia08, Bol07, Bou03c, Čhu02e, Čhu02a, Čhu02b, Chu02i, Čhu02c, Čhu02d, Deb09, Dei03, Don05b, Eva08b, FR06b, GM04, Gui09w, JQQ04i, Kho03, Led03, LTJ00, LMTJ04, MY06e, O’C03, Oko03, Ole03, Pit06i, Pit06j, Piv07, Sch07a, Sie09f, Sla06b, Sla06r, Spe03, Sum09b, Tel06a, Ves08b, Ves08f, Wer04, Wil08m, Xia09b, Zei04, dH09d, dH09k, dH09l, BvdVP02x, CC08, Cer06f, Gru08, Gui09e, Gui09z, Har00, Hir03, Hob00, JW00, Kho08, Kho00, Pit06b, Pit06f, Pit06g, Pit06k, Voi00a, Voi00d, dH09q]. Randomized [Cat04e]. Randomly [SW03]. Randomness [AD08d, AD08h]. Range [GOW04i, NS07c, Bil03e, Yaf00f, Yaf00g, Yaf00h, Yaf00o, Yaf00p, Yaf00q]. Rank [Dol03h, GL05j]. Rank-one [Dol03h]. ranked [RY09f]. Rapport [Del02c, DDM02]. Rate [Bjö04, Fil01i]. Rates [Bis08n, Bis08o, BvdVP02t, Sie09f]. Rational [BHGVN06, HB06, KL09, Nes03]. ray [Nat08, Sch07c, Leu03b]. Raz [MW03]. Re [Eva08k, Eva08l]. Re-Graft [Eva08l]. Re-Grafting [Eva08k]. Reaction [Cos08, Mel02b, Mim03]. Reaction-Diffusion [Mim03]. Read [Pis04]. Real [AHKM08g, BSS09i, Bün02, Cat08, DIK00g, DIK00h, Eva08h, For04, Fou05, Har08n, Har08o, Hua04, Kas08, Lyu04, MK07j, Zum07b, vdH06l, AM02b, DIK00f, DIK00k, MK07b, MK07c, ZZ04]. Real-Part [MK07j]. realm [Van01d]. Receptor [Rei05]. Recognising [Éme09]. Recollections [Mer04]. Reconstruction [AK04s, CR08, Mos08a, AGLT02e, Sch07h, Sch07i, AK04e]. Recovering [Unt00q]. Rectifiability [Paj02h]. Recurrence [Sch03i, dS01y]. Recurrent [Bau08, Ear08, Osi07c]. Recursive [AP06, Tak06y, Doh03e]. Reduction [MMO^{+}07c, MMO^{+}07b, MMO^{+}07j, MMO^{+}07k, MMO^{+}07l, MMO^{+}07m, MMO^{+}07n, MMO^{+}07r, MMO^{+}07p, MMO^{+}07q, MMO^{+}07u, Wei09m, BHLV03b, Gal03, HCG^{+}00, Cut02c, Cut02d]. Reductive [Let05b, Let05h]. References [AC02j, AGH03p, AK04t, AGLT02l, Bil03m, BHLV03k, Bro04k, Bru02h, BPF^{+}02o, CMZ02k, Cao03k, Cat04k, CE05i, CL02f, Čhu02j, Cut00g, Dol03l, Efs05i, FJ04r, Fil01k, Füh05h, GR03j, GRKVL02i, GL02g, Goh04h, GdS03o, Gre03b, Grü02n, HN05t, HK03k, Jef04h, JQQ04l, JK04n, KM04e, Kie02p, KZ02g, Lee04k, Let05j, Li02l, Mel02k, Mon02m, Oli02q, Pil03l, Pis01l, Rei04h, Run02r, Sae04f, Sch03g, Sib04h, Som02l, Ste03g, Ste03-28, Teu03i, Vas02h, Wen03j, WG00h, YC04k, Zha02h, dR02q]. Refinements [AT09k]. Reflected [NNY05, Pis07]. Reflection [Bis09, BS03c]. Reflections [MY07c]. reflexive [Chr00d, Sim08i]. Reformulation [ASWY07k]. regression [Nem00e, Nem00f]. Regular [Mar00f, MMO^{+}07r, Kol00h, Mar00a, Mar00d, Paj02d, AGH03o]. Regularity [AGH03a, ABCV05, BCM09g, Urb08g, Ves08b, BV08k, BV08l, KW04, Li02h, Mel02g, Mel02e]. Regularization [RV07, AGLT02c, AGLT02d, Eng00, Sch07v, AGLT02b, AGLT02f, AGLT02g]. regularized [Bru02c]. Reissner [Fal08c]. Related [AP06, Ale04b, Doh03g, Don07b, GRKVL02f, GL05g, Kaz06, LCFG01a, Mec07, Mis08f, Mis08h, Sin08b, YC04b, Gin09i, Han03, MY02b, Wei09f]. Relation [Hir02, Tel06e, Grü02e]. Relations [Bey07i, KM04c]. Relative [AHKM08h, KL01d, Mon01f, Yaf00j]. Relatively [Gil03d, Gil03c]. Relatives [RVY06]. Relativistic [AHKM08d, NT01, Kie02b]. relativity [Oli02i, Oli02h]. relaxation [Bil03g]. Relaxed [BSPS09f]. Reliability [Doh03f]. Remark [Bob07b, Bou07a, Led03, Bal00, Rás03]. Remarkable [MY06d]. Remarks [Bob00, Dol03g, JS04, KS05, Kla00, Mec07, MW03, Tai09f, Bil03g, CC01, Lio00, NP01o, Tel06d]. Remarquable [Yor02c]. Renorming [MOTV09g]. Renormings [MOTV09e]. Repelling [CL07]. replication [ST03]. Representation [BH08, CMHL02, Cow08, Éme01, EE09, GSE07h, Har08a, Kas08, MY06d, RS07h, Ser07a, Tar02, Xia01h, CSM02, Yor02b, Bor03, FBF^{+}02h, Phe01j, Phe01k, Phe01m, Ver03, Yaf00r, AK04c, TDP08]. Representations [Che03, Füh05b, Gil03b, GL02e, GSE07g, GSE07f, Pau06, RS07e, Unt08f, Vog08, ASWY07e, ASWY07h, BS03c, Dou03, Pis01d, Wei09f]. representing [Phe01q]. Reproducing [Wil08c, Wil08d, Wil08e]. Reproduction [vdDW08]. Requisites [Ber07k]. Rescaling [CM08]. Research [Tia08g]. reservoir [EK00]. Resistance [KS06b]. Resolution [Mof09, Cut02e, Cut02f, Cut02g, Cut02h]. Resolutions [LH09d, Gre03g]. Resolvent [Ade01f, HN05l, Yaf00k]. Resolvents [Gil03b, Phe01d]. Resonance [Efs05e, SW09, Lom00b, Lom00a, Lom00c]. Resonances [Lom00j]. Respect [AHKM08c, AHKM08b, Cou07b, DP06, Mis08h, Mis08i, Naj07, Naz03a, FSU01, MK07c, MK07e, Str03]. Response [GRMH08]. Restricted [KS06a, SV00]. restriction [Fre09i]. resultant [Phe01p]. résultat [Bai08]. résultats [BL08]. Results [Bol07, CMRT09b, CLM^{+}08c, Izh04a, Jon08p, Kar06, Kar04a, KL01g, Mel02b, Mue09, NP01k, NS07c, Sin08b, Sla06i, Sla06p, Son08, Ste03b, Ste07e, Ste07g, Tai09h, Tia08g, Zhi01a, AP03, Bil03h, Gre03k, HNKS06d, Hsu00n, NP01n, Rit00i, Sim08l]. Return [HN05o]. Reversal [Dor08]. reversed [Cha03]. Reversibility [BV08n]. Reversible [Eis07, Lom00i, Mou03b, Lom00h, Lom00j]. Reversing [TF09]. Review [BJS08b, DFT08b, KP05, Pis05, Caf03]. Revisited [Fil01d, Led01, Obł07, Piv07, Unt00k]. Rham [Pfl01c]. Riccati [Las04]. Riemann [Bor03, Bru02e, Kui03, Ste07l]. Riemannian [AD08f, GRKVL02d, GRKVL02b, GRKVL02a, GRKVL02h, GRKVL02e, Hab00g, Lan09, Oli02e, Oli02c, Rei04d]. Riesz [FSU01]. Right [LH09s, LH09w, Fre09b, Fre09u, Fre09-31]. Rigid [NS08d]. Rigidity [Val08].

Rigorous [Ves08h]. Ring [GR03i, GR03c, Gre03i]. Ringed [LH09g, LH09x]. Rings [Kie02c, NV04a, NV04h, NV04l, NV04c, dSW08i, dSW08j]. Rises [MY06h]. Risk [BJR04b, Del06, Pen04, Cat04c]. Rn [BV08q]. Robust [CLM^{+}08k]. Robustness [BV08o]. Roelcke [Unt00q, Unt00r, Unt00s, Unt00t]. Role [Bau08, NCCFM08]. Romano [AAR03]. Root [Eva08k, Pet08g]. roots [ASWY07a]. Rosen [Mar08d]. Rosenthal [Ole03]. Rotation [JQQ04e]. Rotations [Zak05]. Rotund [MOTV09e]. Rough [Lej03, LCL07c, LCL07d, LCL07f, LCL07g, Lej09]. Rouquier [Chl09h]. routing [AGH03h, AGH03i]. Rudimentary [Cao03b]. Ruelle [QXZ09i]. rule [ASWY07c, DF06]. Russia [VY03].

Saint [ASTB03, Ber00a, Ber02, Bra08h, CP04, CP06, Don07m, Eva08h, Gui09w, LCL07d, MP07, Pic05, PP06, SP06, TZP04, TW04, dH09q]. Saint-Flour [ASTB03, Ber00a, Ber02, Bra08h, CP04, CP06, Don07m, Eva08h, Gui09w, LCL07d, MP07, Pic05, PP06, SP06, TZP04, TW04, dH09q]. salesman [Paj02b]. Sample [BvdVP02w, Doz03, Gre03l, Xia09b]. Sampling [GM07, Füh05f]. Sandwiched [Éme06]. satisfying [CC01, Nem00g]. Saturations [MS00a]. sauts [Sim03]. Scalar [BHGVN06, Bil03f, Rei04e]. Scale [JP08]. scales [Sie09g]. Scaling [Tsi04, Urb08g]. Scalings [GOW04g, GOW04j, GOW04h]. Scattering [AHKM08e, Yaf00m, Yaf00e, Yaf00i, Yaf00j, Yaf00l, Yaf00n]. Scenery [Kho03]. Schachermayer [KS03]. Schatten [Tof08b]. Schechtman [Gor07]. Scheme [Del03, MY07e, Ter06, Cao03f, Yaf00g]. Schemes [Doh03e, Hir02, LH09a, LH09j, LH09l, LH09p, LH09v, LH09y, Ste03d]. Scholes [RS03]. School [AF08a, ADD^{+}03, ACB^{+}03, AZ03, AAR03, BBH^{+}04, BBSW07, BGTT02, BM08b, BHK05, BKK^{+}09b, BG08, Bon08, CM06a, Can03, CMI^{+}03, CT08, CO00, CMHL02, CCJ^{+}03, DR08, DL04, EKMY02a, Mar07, NS08g, PV06, TDP08, Tig04, VY03, ZZ04, GAC^{+}09]. Schramm [Wer04]. Schrödinger [All08a, BB09, Bou03a, Bou03c, Bou04, HN05l, HN05d, Kol00b, Kol00f, Ves08b, Wil08n, Zhi01d]. Schur [GSE07f, Gui09a, Pis01h, Pis01g]. Schwartz [BSS09l, BSS09m]. Sciences [Ban08, CL08]. Scientific [Yor06]. Score [BvdVP02l]. Scores [BvdVP02k]. Scroll [Wal06, JK04i]. Scrolls [JK04a, JK04f, JK04g, JK04m, JK04k, JK04l]. SDEs [Nou08]. Seasonal [Bra08k]. Second [Cer01i, Cut02d, Gui09y, HH06a, JL05d, Rit00i]. Second-order [Gui09y, Rit00i]. seconde [Wit07b]. sections [KL00]. Sectorial [Jef04f]. Securing [CLM^{+}08l]. Segment [BCM09c]. Seiberg [Moo01c, Moo01d]. Selberg [Kac06a, Ste07m, Unt00q, Unt00r, Unt00s, Unt00t]. Selected [Mol07e, ÉY02b]. Selection [MNS09, MP07, Mas07c, Mas07f, NL08, ÉY02a, Sim08l]. Selective [dH09c, dH09d]. Self [App06, BvdVP02x, Hel08, Kre09, MS03a, Naj07, Ros05, Sla06d, Sla06e, Sla06i, Sla06l, Sla06q, Chi09i]. Self-Adjoint [Hel08]. Self-attracting [BvdVP02x]. Self-Avoiding [Sla06d, Sla06e, Sla06i, Sla06l, Sla06q]. Self-Decomposability [App06].

Self-Intersection [Naj07, Ros05]. Self-Similar [Kre09, MS03a]. Selfdual [Tar09]. Semester [GO08]. Semi [GRKVL02d, GRKVL02a, GRKVL02h, Han07g, Har08j, Tak06z, Tel06o, Yor02a, Sch07y, Sch07z, Wit07a, Wit07b, GRKVL02e, HN05d, HN05f, HN05g]. Semi-classical [HN05d, HN05f, HN05g]. semi-discrete [Sch07y, Sch07z]. Semi-Group [Har08j]. Semi-Local [Han07g, Tak06z, Tel06o]. Semi-martingales [Yor02a]. Semi-Riemannian [GRKVL02d, GRKVL02a, GRKVL02h, GRKVL02e]. semi-stables [Wit07a, Wit07b]. Semialgebraic [YC04f, YC04e]. Semiclassical [Hel08, Kol00g, Kol00i, Kol00h, Kol00j]. Semiconductor [AAR03, Rin03]. Semiconductors [AMR03]. Semiconvex [Dol03b]. Semidirect [MMO^{+}07b, MMO^{+}07t]. Semigroup [Tai09n, Cer01a]. Semigroups [Bey07e, Bey07o, Chu08b, Da 04, FR06a, Fag06, BM03]. Semilinear [Tai09c, Ath00]. Semimartingales [DP06, Eis07, Eme00e]. Séminaire [ALÉY00, AÉLY01, AÉLY03a, AÉLY03b, DMÉRS07, DMÉRS08, DMÉRS09, ÉY02a, ÉLY05, ÉY06, Mar08d, Rás09b]. Seminar [DMÉRS09, MS00c, MS03b, MS04, MS07]. Semisimple [Kas08, KL01j, Lee04e, Som02i, Som02j]. Sensitive [Sod07]. Sensitivities [DBAP08]. Separability [Bol08, CMZ02h]. Separable [AHKM08g, CMZ02i, CMZ02c]. Separation [Kuz06]. SEptember [CM06a, AF08a, ACB^{+}03, BBSW07, Bon08, Can03, CT08, DR08, DL04]. Sequence [DP06]. Sequences [AGH03d, Dem08, FBF^{+}02i, GL05c, GL05i, GL05l, Sch03i, FBF^{+}02a]. Sequential [Bis08c, Pit06k]. Series [AH00, Cut02t, FJ04d, JL05b, JL05c, JL05d, MP09c, MP09g, MP09h, Ste07c, Ste07d, Xia01i, dR02f, dR02p, BS03c, Unt00e, Unt00f, Unt00g, Unt00h, Unt00l, Unt00s, Unt00t, dR02d, vdH06e]. Server [GOW04d, GOW04a]. service [AGH03q]. Session [BJS^{+}00, Fas00a, MRR02]. Set [Del06, Iva03, Jon08a, Kel00e, MY06c, Mer03, Naz03a, Osi07c, FBF^{+}02j, Kel00c]. Set-Indexed [Iva03, Mer03]. Sets [BBS07g, BvdVP02q, Cer06v, Dal03, DD03a, Del06, Dol03b, Grü02f, Grü02c, Grü02d, Osi07n, Roh09i, Tak06x, YC04e, YC04f, BSPS09d, BSPS09g, Cao03d, FBF^{+}02e, Kel00b, KS00, Paj02d, Wak00a]. Setting [MY06e, MMO^{+}07o, Urb08d, Yaf00n, Bil03b, BSPS09e, Gui09h, Phe01n, Pis01j]. Settings [Lan09, Pri09m]. setup [Sch07y, Grü02b]. Several [Aba04, JS04, Gui09-29, Hu00, CE05e]. Shabat [Bia09b]. Shafarevich [Bro04g]. Shannon [SV00]. Shape [CLM^{+}08k, CLM^{+}08n, CR08, CO00, Zol00, JM00, Kaw00, Pir00]. Shapes [BD03, Cer06t]. Sharp [MK07j]. shear [Růž00b, Růž00c]. Sheaves [BSS09d, Let05d, LH09d, LH09g, LH09j, LH09u, LH09v, LH09y, LH09x, Bro04d]. Sheet [Dal03, Mou03a, Eis03]. Shelstad [Wei09j]. Shift [Kut07j]. Shift-Invariant [Kut07j]. Shifted [Fre09y, Fre09z, Gui09t]. Shimura [Roh09l]. Shock [Ser07b, Zum07b]. Shocks [Wil07]. Short [Ber05, NS07c, Yaf00p, Yaf00q, Cat04j, ÉY02b, Yaf00o]. Short-Range [NS07c, Yaf00p, Yaf00q, Yaf00o]. Short-Term [Ber05]. shortening [Cao03c]. side [ASWY07c]. Sided [CE05b, CE05c, Deb09, Don05a, Tel06q]. Siegel [Lee04d, Wei09g, Wei09l]. Sieve [Fri06a]. SIFT [CLM^{+}08m, CLM^{+}08l]. Sign [Art04]. Sign-Projection [Art04]. Signal [AT09j, Doi08b, Rei05, Sne05b, Str00]. Signaling [San05]. Signals [LCL07b, RW08, Nem00g]. Signature [GRKVL02d, LCL07h]. signées [BS03a]. Significance [Sch03c]. Similar [Kre09, MS03a]. Similarity [Pis01c, Pis01m, Pis01k, Pis01i]. Simons [Tar09]. Simple [FL08, Sla06r, Tak06-27, Yan06, Bai08, FSU01, Lom00d, Nou08]. Simplex [AGH03m]. Simplicial [Jon08u, Jon08v, LH09z]. Simulation [Cha08a]. Simulations [Can03, Sum09b]. Simultaneous [BCN03, Pet08c, Pet08d, Pet08h]. Sinc [AGLT02e]. Single [GOW04d, GOW04a]. Singular [BSS09d, BSS09n, BF06m, CE05b, CE05j, CE05c, Mel02i, Mim03, Pfl01d, Sae04c, Sae04d, Sae04g, Tak06x, Urb08g, Urb08h, Wil08p, AHKM08h, GL00b, JK04l, JK04i, Paj02d, Zha02c]. Singularities [BHLV03i, Gal06, Han07j, Man08, Mof09, Ste03y, Tak06a, Tak06b, Tak06-28, Ste03x, Ste03i, Ste03j]. Singularity [BHLV03e]. Singularly [Mal09]. Sites [LH09g, LH09x]. Size [MT08, Web08]. Size-Structured [MT08]. sketch [Yaf00a]. Sketches [MY06f]. Skew [BBKM01]. Skorokhod [Pis07]. Slicely [MOTV09b]. Slicing [Glu04, Ole03, Bal00]. Sliding [Utk08b]. Slope [NS08d]. Slow [Mar08b]. Small [AK04p, AK04q, AK04r, AK04s, AK09, AM02b, Don07i, EKMY02a, HN05h, Hsu00m, NP01c, MY02b, Efs05b]. Small-Time [Don07i]. Small-Volume [AK09]. Smash [NV04g]. Smooth [Amb08, BV08q, BV08p, Bou03a, Cer01j, Cer01k, Cer01l, Chi09p, Chi09u, Kun00f, NS08f, Pfl01e, Pha07c, QXZ09b, YC04l, JK04g, JK04m, JK04k, Kun00g, Kun00h, Kun00d, Kun00e, Kun00i, Yaf00g, GdS03c]. Smooth-Fit [Pha07c]. Smoothing [Wil08k, Ste03m]. Smoothings [Man08]. Smoothness [YC04j]. Snake [Ser07a, Ser00]. Snakes [Eva08d]. snug [Hsu00m]. Sobolev [Ben03, Cha03, GZ03b, Gui09d, LO00, Lau09, Led01, Mic03a, Mic03b, Nem00f, Nem00i, Tur00a, Tur00b, Tur00e]. Soft [dH09m, dH09n]. Software [Gau06]. Soils [GRMH08]. Soluble [dSW08i]. Solution [Ber07f, EK00, Zha02e]. Solutions [AK04c, Bou03a, Bre07, CE05d, Del03, Gal06, Grü02h, Mal09, MY06f, Nus03, Sol03, Urb08j, Bil03e, BSPS09f, Cer01c, FS00a, FS00f, Kuk02, Kun00a, Mar00a, Mar00d, NP01m, Oli02m, Zhi01f]. solvability [Wak00e]. Solvable [Li02e, Li02g, Li02f, Li02h]. solve [Ste03w]. Solving [Sch07z]. Some [AF08b, AHKM08m, AGLT02k, BM03, Bil03h, Bil03g, Bob03b, Bol07, Bou03c, Bou07b, Bra08a, CC01, Chi09s, Chi09v, Don05a, Fas00c, FBF^{+}02g, Goh04e, Gro07f, HN05n, Hsu00n, Izh04a, Kaw00, Kol00j, Kun00j, Ler08, Lou08b, Mal09, MY06d, MY02b, Mec07, Mey06, MW03, Mol07f, MOTV09h, MOTV09i, Mue09, NP01n, NP01o, NNY05, Pha07d, Pis07, Roh09o, RY09h, Tal07, Tar09, Tel06p, Unt00u, VY00, Wil08o, Wil08p, Yaf00m, AGLT02f, AGLT02g, Gui09b, JW00, KP05, Pis05, RY09d, Sim08f, Bru02e, JK04e, Lio00, Paj02a, WG00d, YC04b, YC04g]. Son [RVY06]. sorts [Unt00g]. Sous [Yor02b]. Sous-Espaces [Yor02b]. Space [AHKM08g, AT09h, App06, BBS07d, DH09t, Eis07, ETZ07, Fil01c, Gil03f, Gil03i, Gil03p, Gro07e, HK03j, KS07, Moc09d, Osi07r, Pri09a, Pri09b, Pri09k, Vak08b, Xu08, Zak05, Hab00e, Pis01j, Rei04e, Rei04f, Run02l, Wak00c, Teu03c]. Space-adiabatic [Teu03c]. Space-Time [ETZ07]. Spaces [AT09c, AT09g, AT09i, AT09j, AT09k, App05, Bia08, BPF^{+}02g, Chi09f, Chi09m, Chi09o, Chi09p, Chi09v, Chi09u, Chu08e, Cip08, DH09s, DH09v, DH09x, DH09y, DH09z, Eva08i, Fil01e, GOW04c, Gil03e, GdS03f, GdS03a, GdS03g, GL05f, GL05g, GL05j, GL05k, GL05l, GL05m, Hab00g, Hua04, JS04, MRR02, Mee02, Mel02d, Mel02f, Mel02e, Moc09g, Mol07e, MOTV09e, Ole03, Ols08c, Pau06, Pfl01a, Pfl01d, Pfl01h, Pfl01i, PR07g, Rot09f, STJV03, Tof08b, Tur00b, Vas02e, Wil08l, BV08b, BV08l, BV08s, BV08p, BC03, Cer01e, Cer01h, Hab00d, MTJ00, Mon01a, Run02m, Run02i, Sch07e, Sch07f, Sch07d, Sim08h, Sim08i, Sim08j, Sim08n, Tur00a, Tur00e, Wen03f, dR02c]. Spatial [BBS07h, BvdVP02y, CMI^{+}03, FL08, GLW08, Nem00i, Web08, Wu08, vdD08b]. Spatially [CLM^{+}08f, CM03, MP03]. Spatio [CM06b]. Spatio-temporal [CM06b]. Spatiotemporal [GLW08, LB09b]. SPDE [Cer01j, Cer01k, Cer01l, DR08]. Special [ASWY07l, BLS07i, CE05e, Cla06b, Cut02c, Cut02d, GO08, JL05i, KV03, Koe03, MV06, Sim08m, Sta03, AC02f, Wei09n, Wei09o, Oli02h]. Spectra [Gil03t, Gin09e, Wen03f, Wen03c, Wen03d]. Spectral [Bob03b, DH08, FBF^{+}02h, Fou05, HN05s, HN05q, Jak06, Jef04i, Fér08, Gui09u, Gui09t, Kol00j, Teu03f, HN05c]. Spectrally [Don05a, Don07j, KP05, Pis05, KP08]. spectraux [MR00]. Spectrum [Gin09c, Jef04e, Osi07o, Gin09d, Gin09i, Gil03s]. Speculation [SX04]. Speculations [NS07c]. Sphere [BPF^{+}02i]. Spherical [Das09, Efs05d, Gui07, Sch07-27, BPF^{+}02c]. Spherically [Bob03b]. Spiders [Eva08d]. Spin [Bis09, Bol07, BB07, CMRT09b, Gui07, Led01, Moo01f, NS07b, NS07c, Pas04, She07, Tal07, Dha09, Gin09b]. spin-boson [Dha09]. Spin-Glass [Gui07]. Spine [HH09]. Spinor [HH06k]. Spinors [HH06c, HH06d]. Spitzer [Don07k]. Splitting [HN05e, Tak06-29, Tak06-30, EK00]. Spread [BF06c, BF06n, GLW08]. Spreads [BF06d]. Square [Wil08g, Wil08r, Woj07b]. Squared [ET05]. SRB [QXZ09l]. SSD [Sim08j]. St [VY03]. Stability [BV08r, BV08s, Ber07l, Bra08g, Dol03d, Osi07j, Ser07b, Son08, Ste03h, Tsi04, Wil07, Zhi01f, Zum07b, AP03, Sch07n, Bra08h, Ste03b]. Stabilization [Gro07d]. Stabilized [MTJ00]. Stable [CLP07, Del06, GS09b, Gro07g, Kho03, Kul09, QXZ09m, YYY09, BV08j, BV08q, BV08p, Bas03, Jac03, Kol00k, MS03a]. stables [Wit07a, Wit07b]. Stably [Kar04a]. Stacks [Moc09b, Moc09g, Ols08b]. Stages [MMO^{+}07i, MMO^{+}07j, MMO^{+}07n, MMO^{+}07r, MMO^{+}07p, MMO^{+}07q, MMO^{+}07s, MMO^{+}07t]. Standard [ABFM02, AM02b, BHLV03f, Bün02, Deb09, Ste03q]. Standardness [ÉS01]. star [KL00]. Starting [Cox08]. State [AHKM08c, AHKM08b, Fil01c, Son08, Sus08]. stated [Mar08d]. States [AHKM08k, Cha08b, JQQ04k, Ves08c, Ves08b, Bil03i]. static [FS00d]. Stationary [Goh04g, Zhi01g, Yaf00r]. stationnaire [Leu03a]. Statistical [AJPP06, AGP09, Joy06a, Kot09, Mas07j, Mis08e, Cat04b, WG00d, CP04]. Statistics [ASTB03, AD08g, Ber00a, Ber02, Pic05, TZP04, TW04, Ver03]. Statistique [PL03]. Stay [Don07g]. Steady [JQQ04k, Růž00b, Bil03i]. Steepest [Chi09n]. Stefan [Urb08h]. Steiner [BCM09f]. Stellar [Tak06g, Tak06e, Tak06-27]. Step [MY07e, dR02h, BvdVP02l, EK00]. Stieltjes [ETZ07]. Stings [GH02]. Stochastic [AB01, All08b, AGH03a, BBS07d, BBS07c, Bar06, Bas03, Bis08c, Bis08l, Bis08m, Bou08, CM03, CMI^{+}03, CM08, CP04, CR07, CS05, CE05j, Čhu02f, Cou07b, Dal09, DP06, DF05c, EP01, Fag06, Fil01f, Fla08, Gre05, Gui09z, HH01, Hip04, Iof09, Kar06, KRA09, Kho09, Kol00g, Lin05, MY07d, MY07h, Mis08f, Mis08g, Mis08h, Mue09, Nua09, PT05, PP06, PR07c, PR07b, PR07f, PR07g, Pri09m, RV07, Sam08, SF08c, SF05, Yan08, dH09o, AGH03f, AH00, BC03, Cer01b, Kol00b, Pri01, Sie09g, Str03, AGH03e, BBH^{+}04, BBSW07, CE05a, Cut00c, Cut00b, Fil01b]. stochastique [BS03a]. Stochastiques [CSM02, Del02c, DDM02, Mey02e, Mey02b, Mey02c, Yor02c, Yor02d, Bro03]. Stokes [BBF08, Con06, Gal06, SR09g, Sin08b]. Stopping [MY06g, Obł07]. Stratifications [Han07m]. Stratified [Pfl01a, Pfl01i]. Strict [Chy07, MY06a]. Strictly [Chi09g]. String [BM08b]. Strong [CLP07, KMS01, MY06i]. Strongly [Bey07e, Bey07o, NV04c, Grü02c]. Structural [Li02a, Osi07j, Osi07s, Ric09]. Structure [Ale07, BCM09o, BCM09r, CS08, Del06, GL02c, GLW08, Har08g, LH09u, NCCFM08, Ols08f, Roh09j, SF06, Wu08, Xia01e, vdD08b, CC01, Gre03i, Hsu00i, Hsu00n, Reb06]. Structured [CM03, LB08, MR08, MT08, Web08]. Structures [Ber00b, Cat08, FJ04c, Hab00g, Li02c, Mol07d, Mol07e, Pfl01i, Roh09l, STJV03, dS01a, Fre09i, Fre09t, Hab00e, MTJ00, Pit06a]. Study [BCM09c, Éme09, Pfl01a, Tar09, GR03d, Nou08]. Sturmian [FBF^{+}02i]. Sub [Lan09]. Sub-Riemannian [Lan09]. subadditivity [Cer06u]. Subbranches [Tak06h, Tak06i, Tak06j, Tak06-31]. Subcritical [Bra08e, HNKS06e, HNKS06f]. Subgroups [DFPS08h, MP09d]. Subject [AC02k, Bru02i, CL02e, GL02h, KZ02h, Pis01n, Som02m, WG00i, dR02o]. Sublimiting [Sie09f]. Sublinear [Čhu02d]. Submanifolds [Hua04, dS01t]. submartingale [KS03]. Subordinate [SV09b]. subordination [NT00]. Subordinators [Don07l, MS03a]. Subspace [Bec06]. Subspaces [GL05e, LMTJ04, Lyu04, Zva00]. Substitutions [FBF^{+}02l, FBF^{+}02m, FBF^{+}02j, FBF^{+}02k, FBF^{+}02b, FBF^{+}02c, FBF^{+}02d, FBF^{+}02f, FBF^{+}02h]. Substructure [Cap08]. Subtree [Eva08l]. sufficient [Str03]. Suite [Yor02c]. Sum [Bou07b, Glu03, Sim08n]. Sum-Product [Bou07b]. Summary [DIK00i, Fil01h, KL01g, KZ02i, NS08d, RS07i]. Summer [AF08a, ADD^{+}03, ACB^{+}03, AZ03, AAR03, BBH^{+}04, BBSW07, BGTT02, BM08b, BHK05, BKK^{+}09b, BG08, Bon08, CM06a, Can03, CMI^{+}03, CT08, CO00, CMHL02, CCJ^{+}03, DR08, DL04, EKMY02a, Mar07, NS08g, PV06, TDP08, Tig04, VY03, ZZ04, GAC^{+}09]. Sums [Bob03a, dR02j, HK03h, Kho00, SV00, Unt00d]. Super [Sla06j, Sla06s, ST03, Val08, BP01]. Super-Brownian [Sla06j, Sla06s, BP01]. Super-replication [ST03]. Super-Rigidity [Val08]. supercuspidal [RS07e]. superhedging [Rás03]. Superlinear [Bec06]. supermartingale [Str03]. supermartingales [Gre03a]. Superprocesses [BvdVP02d, BvdVP02w]. Supplement [Tak06-32]. Support [Sim03]. Supremum [GZ03a, Kaj08]. Sure [Jak05, Kar06]. Surface [Bob07b, Cer06w, Cer06x, Cer06y, JK04i]. Surfaces [BE00, BPF^{+}02m, BFL^{+}02, BPF^{+}02c, BPF^{+}02j, BPF^{+}02b, Cat08, CT08, Cut02u, Cut02o, DIK00g, Man08, MRR02, Moc09b, Pes07, Ros02, Yan09b, ASWY07e, DIK00a, DIK00f, DIK00h, DIK00k, JK04c, JK04m, JK04d, Mee02, PR02, Wit07f, JK04f]. Switched [Mor08]. Switching [Pha07c]. Symbol [KZ02h, Som02m]. Symbolic [Osi07t, FBF^{+}02k]. Symbols [Run02p, Sae04b, CL02e, Teu03i, Unt00a, Vas02f]. Symmetric [Bob03b, Fre09-27, Fre09-28, Fre09-29, Hor03, Piv07, SY07, YYY09, Bas03, Bia03, CC08, Eis03, Fre09x, GSE07h, Hir02, Ols03, Pao00, Van01b]. Symmetries [Con04, Efs05h, Pri08b, Mon02d]. Symmetrization [BKM04, Fus08]. symmetrizations [Kla00]. Symmetry [Sch03f]. Symplectic [AS08, Can01, Cat08, HH06c, HH06d, HH06f, HH06i, HH06j, HH06k, MMO^{+}07p, MMO^{+}07u, ST08b, dS01u, dS01z, Sib04f, Tia02, CT08]. Synaptic [Ber05]. Synopsis [Chr00j]. synthesis [Rog03]. System [Bor05, Del03, Efs05e, JQQ04f, Osi07e, Sin08b, Eli02, Sie09g]. Systematic [Urb08g]. Systems [Agr08, AJP06a, AJP06b, AJP06c, Bey07d, Bha05, BO06, BvdVP02d, Bre07, BHLV03i, CCJ^{+}03, Cho03, Čhu02a, Čhu02b, Chu02i, Čhu02c, Čhu02d, CG03, Dem09, DF06, Efs05a, Efs05h, EKMY02a, Eva08j, FG09, Gat00, GL02d, GL02f, Grü02k, Han07g, HN05b, HH01, JQQ04g, JQQ04i, JQQ04k, Jon08a, Khe09, KV00c, Kun00f, Kut07j, Las04, Led01, Let05d, Lom00i, MP03, Mar07, MT08, Mim03, Mis08b, Mon02k, Mor08, Mou03b, Osi07f, Pas04, Pil03k, Pil06, Ras07a, Ras07d, Ras07g, Ras07i, Reb06, RB06, Rez08b, Ric09, Ter05, Van01d, Zum07b, DM00, FBF^{+}02f, FBF^{+}02k, Hsu00g, Hsu00h, Kol00a, Kun00g, Kun00h, Kun00d, Kun00e, Kun00i, Mon02f, Mon02h, Mon02e, Mon02c, Mon02d, Oli02g, Oli02f, Oli02e, Sie09b, Sie09d, Van01a, Van01c, Van01b, Van01g, Van01h, Yaf00k]. systems [Yaf00l]. Sz. [Pis01k]. Sz.-Nagy [Pis01k].

Table [AMAF02, ABFM02, AM02b, Bün02]. Tail [Led03, Sod07, Kaj08]. Tail-Sensitive [Sod07]. Tamarkin [Gil03h, Gil03j]. Tame [Tak06i, YC04e, YC04l, YC04f]. Tanaka [Don05b, Mar08d, SY07]. Tanaka-like [Mar08d]. Tangent [BvdVP02q, Zak05]. Tangle [KL01h, KL01k]. Tangle-Categories [KL01k]. Tangles [KL09, KL01l]. Tate [Bro04g, Har08l]. teacher [KS01b]. Technique [MOTV09g]. Techniques [CLM^{+}08l, HN05r, Jon08n, JM00, Mon01b]. technology [Deu00]. Teichmüller [Vas02e]. Telecommunication [Dor08].

Témoignages [ADDD^{+}06]. temporal [CM06b]. tension [Cer06y]. Tensor [AK04k, Fre09-30, GdS03i, Run02k, Oli02b]. Tensors [AK04g, GRKVL02d]. Term [Ber05]. Terms [CE05g, GZ03a]. Tetrahedral [Efs05b]. Their [Fri06c, Let05b, BBS07h, BCM09r, CMHL02, Eng00, HK03c, Hsu00h, MY07d, MK07c, Mof09, Roh09h, Tia08e]. Theorem [ASWY07k, Ale04b, Ben05, Cut02i, Glu04, GL05d, Han07k, Jak05, JQQ04d, KMS01, Phe01l, Pil03b, Shi05, Tai09j, Tai09k, Tak06d, Wei09c, Wil08h, dS01-28, AC02h, BV08m, Cer06-28, Gui09-28, Paj02b, Phe01b, Phe01f, Phe01i, Phe01k, Rás03, Sim08f, vdH06f, KS06a, Sch07a, Tai09l, Tai09m, dR02n]. Théorème [Bia06, Leu03b, RVY06, Goi01]. Théorémes [Del02d, Yor02a]. Theorems [All08a, BLS07h, BLS07i, Bou07b, Füh05f, HH01, KW04, MK07j, Rei04i, RS07g, Sch03b, Ste07k, Sum09b, dS01w, dR02m, AC02c, JK04h, Phe01e, Phe01j, Phe01m]. Theoretic [BF06j, JQQ04d]. Theoretical [Coh03, Pis07, Pis05]. Théorie [Del02d, LLP02, Mey02a, Lac03]. Theories [KL01j, Moc09g]. Theory [ASTB03, AHKM08l, Ale07, AGLT02k, AT09e, AT09f, AT09i, AD08g, Att06b, Bac04, BM08b, BHK05, Ber00a, Ber02, BCM09n, BB09, Bow08c, Bra08a, Bre03, CLM^{+}08n, CP04, Cha08b, CMHL02, Chu02i, Cow08, Cun04, Cut00d, Dei03, Don07m, Eva08a, Fin06, FS08, FS00e, GR03i, Gal06, Gio08, Han07h, Har08a, Har08e, HN05s, Hen05, HG04, Jak06, JQQ04k, JS04, Jon08h, Joy06b, Kac06a, KL01e, Kie02g, Kie02q, KS06b, KP05, LH09k, MMO^{+}07b, MRR02, Mel02e, Men04, Mer03, Mie08c, Moc09d, NV04d, NP01j, NS08g, PV06, Pet08f, Pfl01b, Pic05, Pri08a, QXZ09b, Rob04b, RS07h, Ros04, Rös03, Rot09e, Růž00a, Sch07w, SV09b, Spe03, Ste07g, Tai09n, Tai09b, TDP08, Tar02, TZP04, Teu03g, Tig04, Tur00b, Vak08b, Vis04, Wil08r, Xu08, Yaf00m, Yan08, Zhi01d, dS01g, dS01h, dS01x]. Theory [dSW08e, AP03, AGLT02f, Bor03, BS03c, BHLV03e, Cer06q, FBF^{+}02h, FS00a, FS00f, GR03c, GR03b, GR03f, GSE07h, JM00, Kel00a, KZ02c, Kun00g, Kun00h, Kun00j, Lom00e, Mar00b, Mik00, Nou08, Oli02m, Paj02a, Phe01g, Rás09b, Sib04a, Sib04b, Sie09g, Som02c, Ste03u, Tel06p, Tel06o, Teu03b, Teu03c, TW04, Tur00a, Tur00c, Unt00k, Ver03, Wil08a, NP01n, ÉY02a]. Thermodynamic [Bow08e, Pas04]. Thermodynamics [Rug08]. Thesis [Har08l]. theta [Bru02d, Bru02c]. Thick [KL01l]. Thm. [Cao03g]. Thomas [Mar00c]. Three [DFPS08j, Mee02, dH09o, BvdVP02u, Yaf00k]. Three-Dimensional [DFPS08j, Mee02, BvdVP02u]. three-particle [Yaf00k]. Threefolds [Wei09g]. Tikhonov [AGLT02b, Gro07h]. Time [BBS07d, Bou03a, Deb09, Don07i, Dor08, Eis07, ETZ07, FG09, KS07, LB08, Mou03a, NT00, Naj07, Pes07, Pri09f, Pri09g, Pri09l, TF09, AGH03l, EK00, IK00, Mei03, Sie09g]. Time-Dependent [Bou03a]. Time-Reversal [Dor08]. Time-Space [Eis07, KS07]. Times [DF05b, Don07h, ETZ07, MY06e, MY06g, Ros05, YYY09, Pit06d, RK00]. Tissue [Lol06]. Tissues [San05]. Titres [And06]. Tits [Bro04c]. together [dR02b]. Tomographies [Amm09]. Tomography [Nat08, SW09, Sch07i, Sch07y]. tool [Lom00f]. Tools [BBS07c, Mue09, Utk08b, Mon02b, SR09i]. Topics [AJPP06, Doh03g, Jak06, Kec04, Mis08h, Phe01a, Tia08g, YC04b, Gin09i, Kol00j, Kun00b, CMI^{+}03]. Topoi [NS08c]. Topological [Ale04c, AT09f, Ber09, BF06o, BSPS09g, CR08, KL01j, Mar08c, MMO^{+}07q, BKK^{+}09b, DIK00h, GdS03k, GdS03i]. Topologies [FJ04e]. Topology [DIK00j, DIK00k, Jon08v, Sae04g, GdS03g]. Toposes [BF06c, BF06k, BF06m]. Tori [Han07c, Sib04d]. Toric [Ols08h]. Toroidal [AK01]. Toroidalization [Cut02k]. Torsion [Ade01b, Ade01g]. Torus [ASWY07j, Han07i, Lee04j, Lee04h, ASWY07g]. total [PR02, AC02f]. Totally [CG03]. tous [Yor06]. Toy [Pau06]. TQFT [KL01b]. TQFT-Double [KL01b]. Trading [KH07, LL07, SX04]. Traffic [BCM09m, BCM09o, BCM09p, BCM09q, BCM09r]. Trajectories [Osi07q]. trajectory [Ser08]. transaction [Rás03, Rás09b]. Transcendence [Mas03, NP01c, NP01d, NP01a, FBF^{+}02a, NP01k]. Transcendental [NP01n]. Transduction [Rei05, Sne05b]. Transfer [Li02d, Li02k, MOTV09g, LvdLM00, Wei09j]. Transform [Füh05c, HH06d, dS01v, dR02e, dR02g, CM02, Unt00l, Unt00r, Yaf00c]. Transformation [Raj06, Rei04i]. Transformations [BL08, Dol03e, BHLV03g, FBF^{+}02e, HK03b, Oli02k]. Transforms [BPF^{+}02m, Cut02r, Cut02s, FSU01, Füh05g, Füh05d, Füh05b, Let05f, Let05g, Let05h, BPF^{+}02d]. Transition [Kol00k, Moc09b, Ton09]. Transitions [Bis09, Urb08h, BvdVP02v, Cat04g]. Transmission [Ber05, FL08, AK04f, AK04j]. Transport [AF08a, Amb03, Amb08, EH08b, Rin03]. Transportation [ACB^{+}03, Bal04, BCM09n, BD03, BPSS09b, Caf03, Mec07, AP03, Ant00, Vil03]. Transseries [vdH06k, vdH06l, vdH06i]. Transversality [YC04i]. Transversals [Kie02e]. Travaux [And06]. Traveling [MP03, Paj02b]. Tree [BCM09r, Eva08b, Eva08d, FJ04h, FJ04s, FJ04g, FJ04c]. Trees [Bur00, Eva08h, Eva08j, Eva08i, GZ03a, Jon08c, Kub08, Sla06k, Sla06m, Bro04c, Hob00, Oko03]. Triangular [Gil03o, Gil03c]. Triangulated [LH09f]. Tricks [Li02a]. Triples [FLC08, dS01c]. Tróia [CO00]. Trous [MR00]. true [KS03, KS05]. truncated [AGLT02b]. Trunks [Tak06o]. Tsirelson [ÉS01]. Tubes [Doh03d, Doh03c, Doh03i]. Tubular [Tak06n, Tak06o, dS01-28]. Tumour [CM06b]. Tunneling [HN05g]. Turbulence [Con06]. Turbulent [CM06a]. Tutorials [BFET05, Fri06d, Fri08, Sne05b]. twisted [Wei09i, Lee04h]. Twistor [BPF^{+}02l]. Twists [Sei08]. Two [AP06, Tel06q, Van01h, Ver03, Yan06, dH09p, AGH03i, BSPS09d, Unt00f, Unt00h, Unt00m, Unt00s, Unt00t, Wei09l, Wit07e, CE05c, JK04h, Sch02d]. Two-dimensional [Van01h, BSPS09d]. Two-sided [Tel06q, CE05c]. Two-weight [Sch02d]. Type [AAFM07, BLS07d, BLS07h, Bob07b, BCG08, Cat08, Chi09t, DH09t, DH09v, DH09y, Doh03i, Kla04, MP03, Moc09b, Roh09m, Tak06l, Tak06m, Tak06y, Van03, BC03, FS00b, FS00d, HK03f, KZ02d, Kur01, Lau09, Mar00c, MK07b, Tak06k]. typed [Har00]. Types [DFPS08g, GL05k, Man08, Tak06-31, Tak06-33, DIK00b, DIK00c].

Uhlenbeck [App06, Bis08o, EM05]. Ultrasound [Amm09]. Unbounded [Gro07g, Gro07f, Las04, Led01, Cer01g, Kun00a]. Unconditional [BN03]. Uncountable [Wen03d]. Undecomposable [Riv02]. underlying [Sie09g]. Unicité [Pha01]. Unified [Car07, MK07j, Hu00]. Unifiée [LLP02]. uniform [dRV00, Phe01h, Pri01]. Uniformity [AD08h]. Uniformly [Chi09o, Chi09p, MOTV09e, Pis01d]. Unifying [BF03]. Uniqueness [MP04, Phe01q, Rei04i, Wak00f, Dol03d, Pil03e, Pil03g, Rei04b, Rei04c]. Unit [BHGVN06, FBF^{+}02e, Wei09m]. Unital [Goh04e]. unitarizable [Pis01d]. Unitary [Vog08, Unt00e]. univariate [Rit00c, Rit00f]. Universal [FJ04f, Fre09-31, GL05m, Sae04d, Cat04a]. Universality [Ste07b, Ste07i, Ste07n]. Unstable [QXZ09m]. Unsteady [Růž00c]. Unveiling [MY06h]. Upper [Gin09k, NP01p, Tel06r, Cer06e]. Urban [BPSS09b]. Using [Gie07c]. Usual [Unt08c]. Utility [KH07, Sch04a].

V [AAR03, Li02e]. Vacancies [AGH03q]. Vakonomic [Oli02g]. Valency [Hir02]. valeurs [PL03]. Validity [CLM^{+}08g]. Valuation [BJR04b, GR03f]. Valuations [Ale04b, Ale04a, Ale07, KZ02f, KZ02b, KZ02d, FJ04d]. Valuative [FJ04s]. Value [Ste03c, Ste07o, Ste07p, Tai09a, Tai09c, Tai09e, FS00a, FS00f, Kol00a, vdH06f]. Value-Distribution [Ste07o, Ste07p]. Valued [App06, Bia09a, Gil03l, Rot09g, Wil08q, Bru02b, PT05]. Values [DMH03, Ros04, YC04i, Zha02c, NP01b]. Vanish [MY06c]. Vanishing [Bre07]. Variable [KS07, Pes07, Urb08b]. Variables [Aba04, AdlPP^{+}08, Bou04, Cut02t, Kuz06, Li02e, NS07b, Ole03, Kho00, Voi00a, Voi00d]. Variants [Jon08f]. Variation [KS07, Mar00f, NT01, Kaj08, RK00]. Variational [Bil03j, Chi09n, Dol03m, FS00e, Rei04i, Son03, Sus08, dS01-27, Bil03e, DM03, Bil03c, Bil03b, Bil03d, Rei04d]. Variations [Cut00c, DM08, Mas08b, Roh09n, YC04h, YC04d, Pir00, YC04f]. variété [Eme00e]. Variétés [Eme00f]. Varieties [AK01, BSS09d, BSS09i, BSS09n, HB06, Lee04f, Lee04b, Mas03, NS08f, Ols08c, Ols08g, Ols08h, Ols08i, Roh09l, Van01e, Van01c]. Various [BCM09d, Fas00b]. vecteurs [Eme00f]. Vector [Amb08, BSS09b, BSS09i, BSS09n, HN05q, PT05, Wil08q, dS01r, Lom00j, Bru02b, BPF^{+}02h, Oli02b, Rei04f]. Vector-Valued [Wil08q, PT05]. Vectors [Füh05e, RS07f]. Velocity [Kie02b, LB09b]. Venice [TDP08]. Versa [Cow08]. versal [Ste03p]. Vershik [ÉS01]. Version [Bar04, Glu04, Hor03, Sch07a]. Vertebrate [Rei05]. Vertex [Mar08c]. Vertices [SW03]. VI [Li02f]. Via [LH09u, BPSS09b, CM02, Doh03d, Doh03c, Doh03i, Fri06a, MMO^{+}07p, Mas07c, RV07, SV00, Xia01h]. Vibrations [Efs05b]. Vice [Cow08]. View [AT09e, BF03]. VII [Li02g]. VIII [Li02h]. Virtual [BSS09o, BSS09p, Izh04b, Moc09j, Deu00]. Virus [WL08b]. viscosities [Růž00b, Růž00c]. Viscosity [Bre07, Grü02h, Zum07b, Oli02m]. Viscous [CM06a, Wil07, Zum07b]. visible [Gre03i]. Visit [Cra08]. Vitushkin [Paj02f]. Voids [AD08a]. Volatility [LL07]. Volume [AK09, BS04, Cer06z, KL00, Tel06n]. Vortices [Tar09]. Vries [Zhi01d]. vs [NS07b].

Waerden [Zha02e]. Walk [Deb09, Sla06b, Sla06d, Sla06e, Sla06i, Sla06l, Sla06q, Sla06r]. Walks [Bia08, Don05b, FR06b, Kon08b, Tel06a, Zei04, BvdVP02x, Gru08, Pit06j]. Walsh [Cox08]. Wasserstein [Bol08, SST09]. Watanabe [BvdVP02d]. water [Lom00b]. Wave [Dal09, Wak00a]. Wavelet [Dah03, DH09s, DH09y, Füh05g, Füh05d, Kut07a, Kut07h, Füh05b]. Wavelets [DH09z]. Waves [MP03, TF09, Zum07b, Lom00b]. way [Her06d]. Weak [Bis08n, Bis08o, CJMS06, Eva08a, FS00f, Gal06, HNKS06g, Iva03, Kaz06, MY06i, Urb08j, FS00a, Zol00]. weaker [KZ02d]. Weakly [STJ07, Grü02d]. Wegner [Ves08g, Ves08h]. Weight [NS08e, NS08f, NS08d, Sch02d]. Weighted [Bob03a, Kut07i, Kut07j, Mel02f, Tak06x, Tur00b, Wil08r]. Weights [GSE07i]. Weinstein [dS01g, dS01-28]. Well [AK04b]. Well-Posedness [AK04b]. Wentzell [Sie09g]. West [WL08b]. Wetting [dH09l]. Weyl [Jef04b, Unt00a]. Whether [Éme09]. which [MY06c]. White [JQQ04f]. Wick [Ler08]. wide [Bil03e]. Width [Bün02, GMR00]. Wiener [Che03, Del02c, DGP00, DMH03, Don07d, Don07f, Mis08i, Pri09b, RY09h, Wil08a, Zak05]. Wigner [Gui09u, Gui09t, Gui09-27, Gui09-29, Gui09-28]. Wild [Eva08m, Tak06j]. Willinger [KS06a]. Willmore [BPF^{+}02m, BPF^{+}02c, BPF^{+}02j, BPF^{+}02b]. Winding [Ber09]. without [Eva08d, Her06b, Her06f, Teu03f]. Witt [Izh04b]. Witten [Abr08, Beh02b, HN05s, HN05b, HN05f, HN05g, HN05c, Moo01c, Moo01d, Vak08b]. Wong [BC03]. Words [Gui09-29, FBF^{+}02d]. Work [Yor06]. Writhing [Ber09]. Wulff [Cer06-27, CP06, Cer06-28].

X [Nat08, Sch07c]. X-ray [Nat08, Sch07c]. XL [DMÉRS07]. XLI [DMÉRS08, Rás09b]. XLII [DMÉRS09]. XXIX [Ber02]. XXVIII [Ber00a]. XXX [ASTB03]. XXXI [CP04, TZP04]. XXXII [PP06, TW04]. XXXIII [MP07, Pic05]. XXXIV [CP06, LCL07d, SP06, ALÉY00]. XXXIX [ÉY06]. XXXV [Don07m, Eva08h, AÉLY01]. XXXVI [Bra08h, Gui09w, AÉLY03a]. XXXVII [dH09q, AÉLY03b]. XXXVIII [ÉLY05, Mar08d].

Yang [CMZ02e]. Yau [Roh09a, Roh09e]. Yetter [CMZ02e, Som02g, Som02h, Som02n]. Young [Ver03].

Zakai [BC03]. Zero [NP01q]. Zeroes [MY06c]. Zeros [Chi09t, Chi09q, Kho03, KV00b, KV00e, KV00d, KV00a, Gil03k]. Zeta [JL05h, RS07j, dSW08j]. Zonoids [Pao03]. Zuber [Gui09a]. Zygmund [DH09s, DH09t, KMS01].

References

Allaire:2008:FM

[AADH08] Grégoire Allaire, Anton Arnold, Pierre Degond, and Thomas Yizhao Hou. Front matter. Lecture Notes in Mathematics, 1946:I–XIV, 2008. CODEN LNMAA2. ISBN 3-540-79573-1 (print), 3-540-79574-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-79574-2/1.pdf.

Artstein-Avidan:2007:GAC

[AAFM07] S. Artstein-Avidan, O. Friedland, and Vitali D. Milman. Geometric applications of Chernoff-type estimates. Lecture Notes in Mathematics, 1910:45–75, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72053-9_2.pdf.

Anile:2003:MPS

[AAR03] Angelo Marcello Anile, W. Allegretto, and C. A. (Christian A.) Ringhofer, editors. Mathematical Problems in Semiconductor Physics: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, July 15-22, 1998. With the collaboration of G. Mascali and V. Romano, volume 1823 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40802-9 (print), 3-540-45222-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1823. URL

http://link.springer.com/book/10.1007/b13405;   http:
//www.springerlink.com/content/978-3-540-45222-5.

Alibert:2001:GDS

[AB01] J. J. Alibert and K. Bahlali. Genericity in deterministic and stochastic differential equations. Lecture Notes in Mathematics, 1755:220–240, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44671-2_17/.

Abate:2004:ADS

[Aba04] Marco Abate. Angular derivatives in several complex variables. Lecture Notes in Mathematics, 1848:1–47, 2004. CODEN LN-MAA2. ISBN 3-540-22358-4 (print), 3-540-44487-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44487-9_
1/.

Ayache:2005:RIG

[ABCV05] Antoine Ayache, Albert Benassi, Serge Cohen, and Jacques Lévy Véhel. Regularity and identification of generalized multifractional Gaussian processes. Lecture Notes in Mathematics, 1857:290–312, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_20.pdf.

Arad:2002:SITa

[ABFM02] Z. Arad, F. Bünger, E. Fisman, and M. Muzychuk. Standard integral table algebras with a faithful nonreal element of degree 5. Lecture Notes in Mathematics, 1773:43–81, 2002. CODEN LNMAA2. ISBN 3-540-42851-8 (print), 3-540-45558-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45558-2_3/.

Abramovich:2008:LGW

[Abr08] D. Abramovich. Lectures on Gromov–Witten invariants of orbifolds. Lecture Notes in Mathematics, 1947:1–48, 2008. CO-DEN LNMAA2. ISBN 3-540-79813-7 (print), 3-540-79814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-79814-9_1.pdf.

Alberich-Carraminana:2002:P

[AC02a] Maria Alberich-Carramiñana. 1. Preliminaries. Lecture Notes in Mathematics, 1769:1–28, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45538-7_1/.

Alberich-Carraminana:2002:PCM

[AC02b] Maria Alberich-Carramiñana. 2. Plane Cremona maps. Lecture Notes in Mathematics, 1769:29–71, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45538-7_2/.

Alberich-Carraminana:2002:CTJ

[AC02c] Maria Alberich-Carramiñana. 3. Clebsch’s theorems and Jacobian. Lecture Notes in Mathematics, 1769:73–100, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45538-7_
3/.

Alberich-Carraminana:2002:C

[AC02d] Maria Alberich-Carramiñana. 4. Composition. Lecture Notes in Mathematics, 1769:101–125, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45538-7_4/.

Alberich-Carraminana:2002:CM

[AC02e] Maria Alberich-Carramiñana. 5. Characteristic matrices. Lecture Notes in Mathematics, 1769:127–176, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45538-7_5/.

Alberich-Carraminana:2002:TPS

[AC02f] Maria Alberich-Carramiñana. 6. Total principal and special homaloidal curves. Lecture Notes in Mathematics, 1769:177– 205, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45538-7_6/.

Alberich-Carraminana:2002:ICM

[AC02g] Maria Alberich-Carramiñana. 7. Inverse Cremona map. Lecture Notes in Mathematics, 1769:207–225, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45538-7_7/.

Alberich-Carraminana:2002:NFT

[AC02h] Maria Alberich-Carramiñana. 8. Noether’s factorization theorem. Lecture Notes in Mathematics, 1769:227–247, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45538-7_
8/.

Alberich-Carraminana:2002:GPC

[AC02i] Maria Alberich-Carramiñana. Geometry of the Plane Cremona Maps, volume 1769 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 257 pp. LCCN QA3 .L28 no. 1769. URL http:/

/link.springer.com/book/10.1007/b82933;   http://www.
springerlink.com/content/978-3-540-45538-7.

Alberich-Carraminana:2002:R

[AC02j] Maria Alberich-Carramiñana. References. Lecture Notes in Mathematics, 1769:249–251, 2002. CODEN LNMAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45538-7_9/.

Alberich-Carraminana:2002:SIN

[AC02k] Maria Alberich-Carramiñana. Subject and index notation. Lecture Notes in Mathematics, 1769:253–256, 2002. CODEN LN-MAA2. ISBN 3-540-42816-X (print), 3-540-45538-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45538-7_
10/.

Ambrosio:2003:OTA

[ACB^{+}03] Luigi Ambrosio, Luis A. Caffarelli, Yann Brenier, Giuseppe Buttazzo, Cedric Villani, and Sandro Salsa, editors. Optimal Transportation and Applications: Lectures given at the C.I.M.E. Summer School, held in Martina Franca, Italy, September 2-8, 2001, volume 1813 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40192-X (print), 3-540-44857-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1813. URL http:/

/link.springer.com/book/10.1007/b12016;   http://www.
springerlink.com/content/978-3-540-44857-0.

Ambrosio:2008:FM

[ACC^{+}08] Luigi Ambrosio, Luis Caffarelli, Michael G. Crandall, Lawrence C. Evans, and Nicola Fusco. Front matter. Lecture Notes in Mathematics, 1927:I–XI, 2008. CODEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-75914-0/1.pdf.

Arwini:2008:CVG

[AD08a] Khadiga A. Arwini and Christopher T. J. Dodson. Cosmological voids and galactic clustering. Lecture Notes in Mathematics, 1953:119–137, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69393-2_6.pdf.

Arwini:2008:FM

[AD08b] Khadiga A. Arwini and Christopher T. J. Dodson. Front matter. Lecture Notes in Mathematics, 1953:I–X, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-69393-2/
1.pdf.

Arwini:2008:IG

[AD08c] Khadiga A. Arwini and Christopher T. J. Dodson. Information geometry. Lecture Notes in Mathematics, 1953:31–54, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69393-2_3.pdf.

Arwini:2008:IGN

[AD08d] Khadiga A. Arwini and Christopher T. J. Dodson. Information Geometry: Near Randomness and Near Independence, volume 1953 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 253 pp. LCCN QA3 .L28 no. 1953. URL http://link.

springer.com/book/10.1007/978-3-540-69393-2;   http:/
/www.springerlink.com/content/978-3-540-69393-2.

Arwini:2008:IGB

[AD08e] Khadiga A. Arwini and Christopher T. J. Dodson. Information geometry of bivariate families. Lecture Notes in Mathematics, 1953:55–107, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69393-2_4.pdf.

Arwini:2008:IRG

[AD08f] Khadiga A. Arwini and Christopher T. J. Dodson. Introduction to Riemannian geometry. Lecture Notes in Mathematics, 1953:19– 30, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69393-2_2.pdf.

Arwini:2008:MSI

[AD08g] Khadiga A. Arwini and Christopher T. J. Dodson. Mathematical statistics and information theory. Lecture Notes in Mathematics, 1953:1–18, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69393-2_1.pdf.

Arwini:2008:NPR

[AD08h] Khadiga A. Arwini and Christopher T. J. Dodson. Neighbourhoods of Poisson randomness, independence, and uniformity. Lecture Notes in Mathematics, 1953:109–117, 2008. CO-DEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69393-2_5.pdf.

Ambrosio:2003:MAE

[ADD^{+}03] Luigi Ambrosio, Klaus Deckelnick, Gerhard Dziuk, Masayasu Mimura, Vsevolod A. Solonnikov, and Halil Mete Soner, editors. Mathematical Aspects of Evolving Interfaces: Lectures given at the C.I.M.–C.I.M.E. joint Euro-Summer School held in Madeira, Funchal, Portugal, July 3-9, 2000, volume 1812 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-14033-6 (print), 3-540-39189-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1812. URL

http://link.springer.com/book/10.1007/b11357;   http:
//www.springerlink.com/content/978-3-540-39189-0.

Azema:2006:TF

[ADDD^{+}06] Jacques Azéma, Claude Dellacherie, Catherine Doléans-Dade, Michel Émery, Yves Le Jan, Bernard Maisonneuve, Yves Meyer, Jacques Neveu, Nicolas Privault, and Daniel Revuz. Témoignages. (French) []. Lecture Notes in Mathematics, 1874:35– 46, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_4.pdf.

Adelmann:2001:DL

[Ade01a] Clemens Adelmann. Decomposition laws. Lecture Notes in Mathematics, 1761:5–24, 2001. CODEN LNMAA2. ISBN 3-540-42035-5 (print), 3-540-44949-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44949-3_2/.

Adelmann:2001:DPT

[Ade01b] Clemens Adelmann. The Decomposition of Primes in Torsion Point Fields, volume 1761 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42035-5 (paperback), 3-540-44949-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 142 pp. LCCN QA3 .L28 no. 1761; QA3 .L28 no. 1761; QA247. URL http:/

/link.springer.com/book/10.1007/b80624;   http://www.
loc.gov/catdir/enhancements/fy0816/2001031199-d.html;
 http://www.loc.gov/catdir/enhancements/fy0816/2001031199-
t.html;   http://www.springerlink.com/content/978-3-540-
44949-2.

Adelmann:2001:EC

[Ade01c] Clemens Adelmann. Elliptic curves. Lecture Notes in Mathematics, 1761:25–39, 2001. CODEN LNMAA2. ISBN 3-540-42035-5 (print), 3-540-44949-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44949-3_3/.

Adelmann:2001:EMC

[Ade01d] Clemens Adelmann. Elliptic modular curves. Lecture Notes in Mathematics, 1761:41–58, 2001. CODEN LNMAA2. ISBN 3-540-42035-5 (print), 3-540-44949-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44949-3_4/.

Adelmann:2001:I

[Ade01e] Clemens Adelmann. Introduction. Lecture Notes in Mathematics, 1761:1–4, 2001. CODEN LNMAA2. ISBN 3-540-42035-5 (print), 3-540-44949-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44949-3_1/.

Adelmann:2001:IRP

[Ade01f] Clemens Adelmann. Invariants and resolvent polynomials. Lecture Notes in Mathematics, 1761:87–106, 2001. CODEN LN-MAA2. ISBN 3-540-42035-5 (print), 3-540-44949-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44949-3_6/.

Adelmann:2001:TPF

[Ade01g] Clemens Adelmann. Torsion point fields. Lecture Notes in Mathematics, 1761:59–86, 2001. CODEN LNMAA2. ISBN 3-540-42035-5 (print), 3-540-44949-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44949-3_5/.

Auger:2008:AVA

[AdlPP^{+}08] P. Auger, R. Bravo de la Parra, J.-C. Poggiale, E. Sánchez, and T. Nguyen-Huu. Aggregation of variables and applications to population dynamics. Lecture Notes in Mathematics, 1936:209– 263, 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78273-5_5.pdf.

Azema:2001:SPX

[AÉLY01] J. Azéma, M. Émery, M. Ledoux, and M. Yor, editors. Séminaire de Probabilités XXXV, volume 1755 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1755. URL

http://link.springer.com/book/10.1007/b76885;   http:
//www.springerlink.com/content/978-3-540-44671-2.

Azema:2003:SPXa

[AÉLY03a] Jacques Azéma, Michel Émery, Michel Ledoux, and Marc Yor, editors. Séminaire de Probabilités XXXVI, volume 1801 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1801.

URL http://link.springer.com/book/10.1007/b10068;
http://www.springerlink.com/content/978-3-540-36107-7.

Azema:2003:SPXb

[AÉLY03b] Jacques Azéma, Michel Émery, Michel Ledoux, and Marc Yor, editors. Séminaire de Probabilités XXXVII, volume 1832 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1832.

URL http://link.springer.com/book/10.1007/b94376;
http://www.springerlink.com/content/978-3-540-40004-2.

Abdallah:2008:QTM

[AF08a] Naoufel Ben Abdallah and Giovanni Frosali, editors. Quantum Transport: Modelling, Analysis and Asymptotics — Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy September 11–16, 2006, volume 1946 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-79573-1 (print), 3-540-79574-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1946. URL http:

//link.springer.com/book/10.1007/978-3-540-79574-2;
http://www.springerlink.com/content/978-3-540-79574-2.

Albeverio:2008:SMI

[AF08b] Sergio Albeverio and Benedetta Ferrario. Some methods of infinite dimensional analysis in hydrodynamics: an introduction. Lecture Notes in Mathematics, 1942:1–50, 2008. CO-DEN LNMAA2. ISBN 3-540-78492-6 (print), 3-540-78493-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78493-7_1.pdf.

Albeverio:2008:FMb

[AFS08] Sergio Albeverio, Franco Flandoli, and Yakov G. Sinai. Front matter. Lecture Notes in Mathematics, 1942:i–viii, 2008. CODEN LNMAA2. ISBN 3-540-78492-6 (print), 3-540-78493-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-78493-7/
1.pdf.

Altman:2003:DEC

[AGH03a] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Discrete-Event Control of Stochastic Networks: Multimodularity and Regularity, volume 1829 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 313 pp. LCCN QA3 .L28 no. 1829. URL http:/

/link.springer.com/book/10.1007/b93837;   http://www.
springerlink.com/content/978-3-540-39705-2.

Altman:2003:I

[AGH03b] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Introduction. Lecture Notes in Mathematics, 1829:1–6, 2003. CODEN LN-MAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39705-2_
1/.

Altman:2003:MCO

[AGH03c] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Multimodularity, convexity and optimization. Lecture Notes in Mathematics, 1829: 11–38, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_2/.

Altman:2003:PBS

[AGH03d] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part I: 2. Balanced sequences. Lecture Notes in Mathematics, 1829:39– 54, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_3/.

Altman:2003:PSE

[AGH03e] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part I: 3. Stochastic event graphs. Lecture Notes in Mathematics, 1829: 55–74, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_4/.

Altman:2003:PIAa

[AGH03f] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part II: 4. Admission control in stochastic event graphs. Lecture Notes in Mathematics, 1829:79–103, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39705-2_5/.

Altman:2003:PIAb

[AGH03g] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part II: 5. Applications in queuing networks. Lecture Notes in Mathematics, 1829: 105–109, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_6/.

Altman:2003:PIOa

[AGH03h] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part II: 6. Optimal routing. Lecture Notes in Mathematics, 1829:111–118, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_7/.

Altman:2003:PIOb

[AGH03i] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part II: 7. Optimal routing in two deterministic queues. Lecture Notes in Mathematics, 1829:119–149, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39705-2_8/.

Altman:2003:PIM

[AGH03j] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part III: 10. Monotonicity of feedback control. Lecture Notes in Mathematics, 1829:205–223, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39705-2_11/.

Altman:2003:PIN

[AGH03k] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part III: 8. Networks with no buffers. Lecture Notes in Mathematics, 1829:155– 181, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_9/.

Altman:2003:PIC

[AGH03l] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part IV: 11. Comparison of queues with discrete-time arrival processes. Lecture Notes in Mathematics, 1829:229–241, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39705-2_12/.

Altman:2003:PIS

[AGH03m] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part IV: 12. Simplex convexity. Lecture Notes in Mathematics, 1829:243– 259, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_13/.

Altman:2003:PIOc

[AGH03n] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part IV: 13. Orders and bounds for multimodular functions. Lecture Notes in Mathematics, 1829:261–282, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39705-2_14/.

Altman:2003:PIR

[AGH03o] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Part IV: 14. Regular ordering. Lecture Notes in Mathematics, 1829:283– 304, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_15/.

Altman:2003:R

[AGH03p] Eitan Altman, Bruno Gaujal, and Arie Hordijk. References. Lecture Notes in Mathematics, 1829:305–310, 2003. CODEN LN-MAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39705-2_
16/.

Altman:2003:VSA

[AGH03q] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Vacancies, service allocation and polling. Lecture Notes in Mathematics, 1829: 183–204, 2003. CODEN LNMAA2. ISBN 3-540-20358-3 (print), 3-540-39705-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39705-2_10/.

Ang:2002:MP

[AGLT02a] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 1. Mathematical preliminaries. Lecture Notes in Mathematics, 1792:5–16, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45658-2_2/.

Ang:2002:RMP

[AGLT02b] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 2. Regularization of moment problems by truncated expansion and by the Tikhonov method. Lecture Notes in Mathematics, 1792:17–49, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45658-2_3/.

Ang:2002:BGR

[AGLT02c] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 3. Backus-gilbert regularization of a moment problem. Lecture Notes in Mathematics, 1792:51–81, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45658-2_4/.

Ang:2002:HMP

[AGLT02d] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 4. The Hausdorff moment problem: regularization and error estimates. Lecture Notes in Mathematics, 1792:83–97, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_5/.

Ang:2002:AFR

[AGLT02e] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 5. Analytic functions: reconstruction and sinc approximations. Lecture Notes in Mathematics, 1792:99–130, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_6/.

Ang:2002:RSIa

[AGLT02f] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 6. Regularization of some inverse problems in potential theory. Lecture Notes in Mathematics, 1792:131–146, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_7/.

Ang:2002:RSIb

[AGLT02g] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 7. Regularization of some inverse problems in heat conduction. Lecture Notes in Mathematics, 1792:147–169, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_8/.

Ang:2002:E

[AGLT02h] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. 8. Epilogue. Lecture Notes in Mathematics, 1792:171– 173, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_9/.

Ang:2002:Ib

[AGLT02i] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. Index. Lecture Notes in Mathematics, 1792:181–183, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_11/.

Ang:2002:Ia

[AGLT02j] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. Introduction. Lecture Notes in Mathematics, 1792:1– 3, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_1/.

Ang:2002:MTS

[AGLT02k] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. Moment Theory and Some Inverse Problems in Potential Theory and Heat Conduction, volume 1792 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 183 pp. LCCN QA3 .L28 no. 1792. URL http:/

/link.springer.com/book/10.1007/b84019;   http://www.
springerlink.com/content/978-3-540-45658-2.

Ang:2002:R

[AGLT02l] Dang Dinh Ang, Rudolf Gorenflo, Vy Khoi Le, and Dang Duc Trong. References. Lecture Notes in Mathematics, 1792:175– 180, 2002. CODEN LNMAA2. ISBN 3-540-44006-2 (print), 3-540-45658-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45658-2_10/.

Attal:2009:SPP

[AGP09] Stéphane Attal and Nadine Guillotin-Plantard. Statistical properties of Pauli matrices going through noisy channels. Lecture Notes in Mathematics, 1979:433–448, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01763-6_16.pdf.

Agrachev:2008:GOC

[Agr08] Andrei A. Agrachev. Geometry of optimal control problems and Hamiltonian systems. Lecture Notes in Mathematics, 1932:1–59, 2008. CODEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77653-6_1.pdf.

Aguda:2006:MCD

[Agu06] Baltazar D. Aguda. Modeling the cell division cycle. Lecture Notes in Mathematics, 1872:1–22, 2006. CODEN LN-MAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11561606_1.pdf;
 http://link.springer.com/content/pdf/bfm:978-3-540-32415-
7/1.pdf.

Attal:2000:SIQ

[AH00] Stéphane Attal and Robin L. Hudson. Series of iterated quantum stochastic integrals. Lecture Notes in Mathematics, 1729: 157–170, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103801/.

Albeverio:2008:BM

[AHKM08a] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Back matter. Lecture Notes in Mathematics, 523:141– 175, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-76956-9/1.pdf.

Albeverio:2008:ERGb

[AHKM08b] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Expectations with respect to the Gibbs state of the harmonic oscillator. Lecture Notes in Mathematics, 523:69–71, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_7.pdf.

Albeverio:2008:ERGa

[AHKM08c] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Expectations with respect to the ground state of the harmonic oscillator. Lecture Notes in Mathematics, 523:63–68, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_6.pdf.

Albeverio:2008:FHI

[AHKM08d] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. The Feynman history integral for the relativistic quantum boson field. Lecture Notes in Mathematics, 523:85–92, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_9.pdf.

Albeverio:2008:FPIa

[AHKM08e] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. The Feynman path integral in potential scattering. Lecture Notes in Mathematics, 523:19–35, 2008. CO-DEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_3.pdf.

Albeverio:2008:FPIb

[AHKM08f] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Feynman path integrals for the anharmonic oscillator. Lecture Notes in Mathematics, 523:51–62, 2008. CO-DEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_5.pdf.

Albeverio:2008:FIF

[AHKM08g] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. The Fresnel integral of functions on a separable real Hilbert space. Lecture Notes in Mathematics, 523:9–17, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_2.pdf.

Albeverio:2008:FIR

[AHKM08h] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. The Fresnel integral relative to a non-singular quadratic form. Lecture Notes in Mathematics, 523:37–50, 2008. CO-DEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_4.pdf.

Albeverio:2008:FMa

[AHKM08i] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Front matter. Lecture Notes in Mathematics, 523:i– x, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-76956-9/1.pdf.

Albeverio:2008:I

[AHKM08j] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Introduction. Lecture Notes in Mathematics, 523:1–8, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76956-9_1.pdf.

Albeverio:2008:IQF

[AHKM08k] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. The invariant quasi-free states. Lecture Notes in Mathematics, 523:73–83, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-76956-9_8.pdf.

Albeverio:2008:MTF

[AHKM08l] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Mathematical Theory of Feynman Path Integrals: An Introduction, volume 523 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 177 pp. LCCN QA3 .L28 no. 523. URL http:

//link.springer.com/book/10.1007/978-3-540-76956-9;
http://www.springerlink.com/content/978-3-540-76956-9.

Albeverio:2008:SRD

[AHKM08m] Sergio A. Albeverio, Raphael J. Høegh-Krohn, and Sonia Mazzucchi. Some recent developments. Lecture Notes in Mathematics, 523:93–140, 2008. CODEN LNMAA2. ISBN 3-540-76954-4 (print), 3-540-76956-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-76956-9_10.pdf.

Attal:2006:OQSa

[AJP06a] Stéphane Attal, Alain Joye, and Claude-Alain Pillet. Open Quantum Systems I: The Hamiltonian Approach, volume 1880 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-30991-8 (print), 3-540-33922-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 329 pp. LCCN QA3 .L28 no. 1880. URL http://

link.springer.com/book/10.1007/b128449;   http://www.
springerlink.com/content/978-3-540-33922-9.

Attal:2006:OQSb

[AJP06b] Stéphane Attal, Alain Joye, and Claude-Alain Pillet. Open Quantum Systems II: The Markovian Approach, volume 1881 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-30992-6 (print), 3-540-33966-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 239 pp. LCCN QA3 .L28 no. 1881. URL http://

link.springer.com/book/10.1007/b128451;   http://www.
springerlink.com/content/978-3-540-33966-3.

Attal:2006:OQSc

[AJP06c] Stéphane Attal, Alain Joye, and Claude-Alain Pillet. Open Quantum Systems III: Recent Developments, volume 1882 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-30993-4 (print), 3-540-33967-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 310 pp. LCCN QA3 .L28 no. 1882. URL http://

link.springer.com/book/10.1007/b128453;   http://www.
springerlink.com/content/978-3-540-33967-0.

Aschbacher:2006:TNE

[AJPP06] Walter Aschbacher, Vojkan Jakšić, Yan Pautrat, and Claude-Alain Pillet. Topics in non-equilibrium quantum statistical mechanics. Lecture Notes in Mathematics, 1882:1–66, 2006. CODEN LNMAA2. ISBN 3-540-30993-4 (print), 3-540-33967-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-33967-1_1.pdf.

Abe:2001:TGL

[AK01] Yukitaka Abe and Klaus Kopfermann. Toroidal Groups: Line Bundles, Cohomology and Quasi-Abelian Varieties, volume 1759 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-41989-6 (paperback). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 133 pp. LCCN QA3 .L28 no. 1759. URL http:/

/link.springer.com/book/10.1007/b80605;   http://www.
loc.gov/catdir/enhancements/fy0816/2001032054-d.html;
 http://www.loc.gov/catdir/enhancements/fy0816/2001032054-
t.html.

Ammari:2004:Ia

[AK04a] Habib Ammari and Hyeonbae Kang. 1. Introduction. Lecture Notes in Mathematics, 1846:1–4, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44501-2_1/.

Ammari:2004:WP

[AK04b] Habib Ammari and Hyeonbae Kang. 10. Well-posedness. Lecture Notes in Mathematics, 1846:179–183, 2004. CODEN LNMAA2.

ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44501-2_13/.

Ammari:2004:RS

[AK04c] Habib Ammari and Hyeonbae Kang. 11. Representation of solutions. Lecture Notes in Mathematics, 1846:185–195, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_14/.

Ammari:2004:DAF

[AK04d] Habib Ammari and Hyeonbae Kang. 12. Derivation of asymptotic formulae. Lecture Notes in Mathematics, 1846:197–205, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_15/.

Ammari:2004:RA

[AK04e] Habib Ammari and Hyeonbae Kang. 13. Reconstruction algorithms. Lecture Notes in Mathematics, 1846:207–214, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_16/.

Ammari:2004:TP

[AK04f] Habib Ammari and Hyeonbae Kang. 2. Transmission problem. Lecture Notes in Mathematics, 1846:11–39, 2004. CODEN LN-MAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44501-2_
3/.

Ammari:2004:GPT

[AK04g] Habib Ammari and Hyeonbae Kang. 3. Generalized polarization tensors. Lecture Notes in Mathematics, 1846:41–64, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_4/.

Ammari:2004:DFAa

[AK04h] Habib Ammari and Hyeonbae Kang. 4. Derivation of the full asymptotic formula. Lecture Notes in Mathematics, 1846:65– 78, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_5/.

Ammari:2004:DIa

[AK04i] Habib Ammari and Hyeonbae Kang. 5. Detection of inclusions. Lecture Notes in Mathematics, 1846:79–101, 2004. CODEN LN-MAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44501-2_
6/.

Ammari:2004:TPE

[AK04j] Habib Ammari and Hyeonbae Kang. 6. Transmission problem for elastostatics. Lecture Notes in Mathematics, 1846:109– 127, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_8/.

Ammari:2004:EMT

[AK04k] Habib Ammari and Hyeonbae Kang. 7. Elastic moment tensor. Lecture Notes in Mathematics, 1846:129–149, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44501-2_
9/.

Ammari:2004:DFAb

[AK04l] Habib Ammari and Hyeonbae Kang. 8. Derivation of full asymptotic expansions. Lecture Notes in Mathematics, 1846:151–157, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_10/.

Ammari:2004:DIb

[AK04m] Habib Ammari and Hyeonbae Kang. 9. Detection of inclusions. Lecture Notes in Mathematics, 1846:159–173, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44501-2_
11/.

Ammari:2004:A

[AK04n] Habib Ammari and Hyeonbae Kang. A. appendices. Lecture Notes in Mathematics, 1846:215–221, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44501-2_17/.

Ammari:2004:Ib

[AK04o] Habib Ammari and Hyeonbae Kang. Index. Lecture Notes in Mathematics, 1846:237–238, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44501-2_19.

Ammari:2004:PDS

[AK04p] Habib Ammari and Hyeonbae Kang. Part I: Detection of small conductivity inclusions. Lecture Notes in Mathematics, 1846: 5–9, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_2/.

Ammari:2004:PIDa

[AK04q] Habib Ammari and Hyeonbae Kang. Part II: Detection of small elastic inclusions. Lecture Notes in Mathematics, 1846:103– 107, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_7/.

Ammari:2004:PIDb

[AK04r] Habib Ammari and Hyeonbae Kang. Part III: Detection of small electromagnetic inclusions. Lecture Notes in Mathematics, 1846: 175–178, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44501-2_12/.

Ammari:2004:RSI

[AK04s] Habib Ammari and Hyeonbae Kang. Reconstruction of Small Inhomogeneities from Boundary Measurements, volume 1846 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 238 pp. LCCN QA3 .L28 no. 1846. URL http:/

/link.springer.com/book/10.1007/b98245;   http://www.
springerlink.com/content/978-3-540-44501-2.

Ammari:2004:R

[AK04t] Habib Ammari and Hyeonbae Kang. References. Lecture Notes in Mathematics, 1846:223–236, 2004. CODEN LNMAA2. ISBN 3-540-22483-1 (print), 3-540-44501-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44501-2_18/.

Ammari:2009:MSV

[AK09] Habib Ammari and Hyeonbae Kang. The method of small-volume expansions for medical imaging. Lecture Notes in Mathematics, 1983:99–132, 2009. CODEN LNMAA2. ISBN 3-642-03443-8 (print), 3-642-03444-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-03444-2_3.pdf.

Alesker:2004:IV

[Ale04a] S. Alesker. SU(2)-invariant valuations. Lecture Notes in Mathematics, 1850:21–29, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44489-3_3/.

Alesker:2004:HLT

[Ale04b] S. Alesker. Hard Lefschetz theorem for valuations and related questions of integral geometry. Lecture Notes in Mathematics, 1850:9–20, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_2/.

Alesker:2004:TOE

[Ale04c] S. Alesker. A topological obstruction to existence of quaternionic plücker map. Lecture Notes in Mathematics, 1850:1– 7, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_1/.

Alesker:2007:TVM

[Ale07] S. Alesker. Theory of valuations on manifolds, IV. New properties of the multiplicative structure. Lecture Notes in Mathematics, 1910:1–44, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-72053-9_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-72053-9/1.pdf.

Azema:2000:SPX

[ALÉY00] Jacques Azéma, Michel Ledoux, Michel Émery, and Marc Yor, editors. Séminaire de Probabilités XXXIV, volume 1729 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1729.

URL http://link.springer.com/book/10.1007/BFb0103797;
 http://www.springerlink.com/content/978-3-540-46413-
6.

Allegretto:2003:DDE

[All03] Walter Allegretto. Drift-diffusion equations and applications. Lecture Notes in Mathematics, 1823:57–95, 2003. CODEN LNMAA2. ISBN 3-540-40802-9 (print), 3-540-45222-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45222-5_2/.

Allaire:2008:PHE

[All08a] Grégoire Allaire. Periodic homogenization and effective mass theorems for the Schrödinger equation. Lecture Notes in Mathematics, 1946:1–44, 2008. CODEN LNMAA2. ISBN 3-540-79573-1 (print), 3-540-79574-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-79574-2_1.pdf.

Allen:2008:ISE

[All08b] Linda J. S. Allen. An introduction to stochastic epidemic models. Lecture Notes in Mathematics, 1945:81–130, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_3.pdf.

Arad:2002:I

[AM02a] Z. Arad and M. Muzychuk. Introduction. Lecture Notes in Mathematics, 1773:1–11, 2002. CODEN LNMAA2. ISBN 3-540-42851-8 (print), 3-540-45558-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45558-2_1/.

Arad:2002:SITb

[AM02b] Zvi Arad and Mikhail Muzychuk. Standard Integral Table Algebras Generated by Non-real Element of Small Degree, volume 1773 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-42851-8 (print), 3-540-45558-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 126 pp. LCCN QA3 .L28 no. 1773. URL http:/

/link.springer.com/book/10.1007/b82936;   http://www.
springerlink.com/content/978-3-540-45558-5.

Arad:2002:ITA

[AMAF02] Z. Arad, M. Muzychuk, H. Arisha, and E. Fisman. Integral table algebras with a faithful nonreal element of degree 4. Lecture Notes in Mathematics, 1773:13–41, 2002. CODEN LNMAA2. ISBN 3-540-42851-8 (print), 3-540-45558-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-45558-2_2/.

Ambrosio:2003:LNO

[Amb03] Luigi Ambrosio. Lecture notes on optimal transport problems. Lecture Notes in Mathematics, 1812:1–52, 2003. CODEN LN-MAA2. ISBN 3-540-14033-6 (print), 3-540-39189-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39189-0_

Ambrosio:2008:TEC

[Amb08] Luigi Ambrosio. Transport equation and Cauchy problem for nonsmooth vector fields. Lecture Notes in Mathematics, 1927:1–41, 2008. CODEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75914-0_1.pdf.

Ammari:2009:MMB

[Amm09] Habib Ammari, editor. Mathematical Modeling in Biomedical Imaging I: Electrical and Ultrasound Tomographies, Anomaly Detection, and Brain Imaging, volume 1983 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-03443-8 (print), 3-642-03444-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 228 pp. LCCN QA3 .L28 no. 1983; R857.M34 M38 2009. URL http://link.

springer.com/book/10.1007/978-3-642-03444-2;   http:/
/www.springerlink.com/content/978-3-642-03444-2.

Anile:2003:RDH

[AMR03] A. M. Anile, G. Mascali, and V. Romano. Recent developments in hydrodynamical modeling of semiconductors. Lecture Notes in Mathematics, 1823:1–56, 2003. CODEN LNMAA2. ISBN 3-540-40802-9 (print), 3-540-45222-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45222-5_1/.

Agrachev:2008:FM

[AMS^{+}08a] Andrei A. Agrachev, A. Stephen Morse, Eduardo D. Sontag, Héctor J. Sussmann, and Vadim I. Utkin. Front matter. Lecture Notes in Mathematics, 1932:i–xiii, 2008. CODEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-77653-6/1.pdf.

Auroux:2008:FM

[AMS^{+}08b] Denis Auroux, Marco Manetti, Paul Seidel, Bernd Siebert, and Ivan Smith. Front matter. Lecture Notes in Mathematics, 1938: i–xiii, 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-78279-7/1.pdf.

Abramovich:2008:FM

[AMTV08] Dan Abramovich, Marcos Mariño, Michael Thaddeus, and Ravi Vakil. Front matter. Lecture Notes in Mathematics, 1947:I– X, 2008. CODEN LNMAA2. ISBN 3-540-79813-7 (print), 3-540-79814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-79814-9/1.pdf.

Andre:2006:TTP

[And06] Meyer Paul André. Titres et travaux: Postface. (French) []. Lecture Notes in Mathematics, 1874:1–12, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35513-7_1.
pdf;   http://link.springer.com/content/pdf/bfm:978-3-
540-35513-7/1.pdf.

Anttila:2000:TCC

[Ant00] M. Anttila. The transportation cost for the cube. Lecture Notes in Mathematics, 1745:1–11, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0107202/.

Ambrosio:2003:ESR

[AP03] Luigi Ambrosio and Aldo Pratelli. Existence and stability results in the L^{1} theory of optimal transportation. Lecture Notes in Mathematics, 1813:123–160, 2003. CODEN LNMAA2. ISBN 3-540-40192-X (print), 3-540-44857-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44857-0_5/.

Aldous:2006:TRD

[AP06] David Aldous and Jim Pitman. Two recursive decompositions of Brownian bridge related to the asymptotics of random mappings. Lecture Notes in Mathematics, 1874:269–303, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_19.pdf.

Alili:2007:JLN

[AP07] Larbi Alili and Pierre Patie. On the joint law of the L_{1} and L_{2} norms of a 3-dimensional Bessel bridge. Lecture Notes in Mathematics, 1899:247–264, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71189-6_13.pdf.

Applebaum:2005:LPE

[App05] David Applebaum. Lévy processes in Euclidean spaces and groups. Lecture Notes in Mathematics, 1865:1–98, 2005. CODEN LNMAA2. ISBN 3-540-24406-9 (print), 3-540-31450-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11376569_1.pdf.

Applebaum:2006:MVM

[App06] David Applebaum. Martingale-valued measures, Ornstein– Uhlenbeck processes with jumps and operator self-decomposability in Hilbert space. Lecture Notes in Mathematics, 1874:171–196, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_14.pdf.

Arnold:2008:MPQ

[Arn08] Anton Arnold. Mathematical properties of quantum evolution equations. Lecture Notes in Mathematics, 1946:45–109, 2008. CODEN LNMAA2. ISBN 3-540-79573-1 (print), 3-540-79574-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-79574-2_2.pdf.

Artstein:2004:CDC

[Art04] S. Artstein. The change in the diameter of a convex body under a random sign-projection. Lecture Notes in Mathematics, 1850: 31–39, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_4/.

Auroux:2008:LPB

[AS08] Denis Auroux and Ivan Smith. Lefschetz pencils, branched covers and symplectic invariants. Lecture Notes in Mathematics, 1938:1– 53, 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78279-7_1.pdf.

Albeverio:2003:LPT

[ASTB03] Sergio Albeverio, Walter Schachermayer, Michel Talagrand, and Pierre Bernard, editors. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXX — 2000, volume 1816 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-44922-1, 3-540-40335-3. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1816. URL http://

link.springer.com/book/10.1007/3-540-44922-1;   http:
//link.springer.com/openurl?genre=book&isbn=978-3-540-
40335-7.

Akiyoshi:2007:ARG

[ASWY07a] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Algebraic roots and geometric roots. Lecture Notes in Mathematics, 1909:215–231, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71807-9_9.pdf.

Akiyoshi:2007:BM

[ASWY07b] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Back matter. Lecture Notes in Mathematics, 1909: 233–256, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-71807-9/1.pdf.

Akiyoshi:2007:CRS

[ASWY07c] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Chain rule and side parameter. Lecture Notes in Mathematics, 1909:49–99, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71807-9_4.pdf.

Akiyoshi:2007:C

[ASWY07d] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Closedness. Lecture Notes in Mathematics, 1909:171– 214, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71807-9_8.pdf.

Akiyoshi:2007:FSR

[ASWY07e] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Fricke surfaces and PSL(2, C)-representations. Lecture Notes in Mathematics, 1909:15–35, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71807-9_2.
pdf.

Akiyoshi:2007:FM

[ASWY07f] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Front matter. Lecture Notes in Mathematics, 1909: I–XLIII, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-71807-9/1.pdf.

Akiyoshi:2007:JPQ

[ASWY07g] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Jorgensen’s picture of quasifuchsian punctured torus groups. Lecture Notes in Mathematics, 1909:1–14, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71807-9_1.pdf.

Akiyoshi:2007:LRA

[ASWY07h] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Labeled representations and associated complexes. Lecture Notes in Mathematics, 1909:37–47, 2007. CO-DEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71807-9_3.pdf.

Akiyoshi:2007:O

[ASWY07i] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Openness. Lecture Notes in Mathematics, 1909:155– 169, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71807-9_7.pdf.

Akiyoshi:2007:PTG

[ASWY07j] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Punctured Torus Groups and 2-Bridge Knot Groups (I), volume 1909 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xliv + 252 pp. LCCN QA3 .L28 no. 1909. URL http:

//link.springer.com/book/10.1007/978-3-540-71807-9;
http://www.springerlink.com/content/978-3-540-71807-9.

Akiyoshi:2007:RMT

[ASWY07k] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Reformulation of main theorem 1.3.5 and outline of the proof. Lecture Notes in Mathematics, 1909:133–154, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71807-9_6.pdf.

Akiyoshi:2007:SE

[ASWY07l] Hirotaka Akiyoshi, Makoto Sakuma, Masaaki Wada, and Yasushi Yamashita. Special examples. Lecture Notes in Mathematics, 1909:101–132, 2007. CODEN LNMAA2. ISBN 3-540-71806-0 (print), 3-540-71807-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71807-9_5.pdf.

Antoine:2009:AMP

[AT09a] Jean-Pierre Antoine and Camillo Trapani. Applications in mathematical physics. Lecture Notes in Mathematics, 1986:257–292, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05136-4_7.pdf.

Antoine:2009:BM

[AT09b] Jean-Pierre Antoine and Camillo Trapani. Back matter. Lecture Notes in Mathematics, 1986:325–358, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-05136-4/1.pdf.

Antoine:2009:EIP

[AT09c] Jean-Pierre Antoine and Camillo Trapani. Examples of indexed PIP-spaces. Lecture Notes in Mathematics, 1986:103–156, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05136-4_4.pdf.

Antoine:2009:FM

[AT09d] Jean-Pierre Antoine and Camillo Trapani. Front matter. Lecture Notes in Mathematics, 1986:I–XXIX, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-05136-4/1.pdf.

Antoine:2009:GTA

[AT09e] Jean-Pierre Antoine and Camillo Trapani. General theory: Algebraic point of view. Lecture Notes in Mathematics, 1986:11–34, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05136-4_1.pdf.

Antoine:2009:GTT

[AT09f] Jean-Pierre Antoine and Camillo Trapani. General theory: Topological aspects. Lecture Notes in Mathematics, 1986:35–56, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05136-4_2.pdf.

Antoine:2009:OPS

[AT09g] Jean-Pierre Antoine and Camillo Trapani. Operators on PIPspaces and indexed PIP-spaces. Lecture Notes in Mathematics, 1986:57–101, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-05136-4_3.pdf.

Antoine:2009:PAO

[AT09h] Jean-Pierre Antoine and Camillo Trapani. Partial ∗-algebras of operators in a PIP-space. Lecture Notes in Mathematics, 1986:221–255, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-05136-4_6.pdf.

Antoine:2009:PIP

[AT09i] Jean-Pierre Antoine and Camillo Trapani. Partial Inner Product Spaces: Theory and Applications, volume 1986 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 352 pp. LCCN QA3 .L28 no.

1986. URL http://link.springer.com/book/10.1007/978-
3-642-05136-4;   http://www.springerlink.com/content/
978-3-642-05136-4.

Antoine:2009:PSS

[AT09j] Jean-Pierre Antoine and Camillo Trapani. PIP-spaces and signal processing. Lecture Notes in Mathematics, 1986:293–324, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05136-4_8.pdf.

Antoine:2009:RPS

[AT09k] Jean-Pierre Antoine and Camillo Trapani. Refinements of PIPspaces. Lecture Notes in Mathematics, 1986:157–219, 2009. CODEN LNMAA2. ISBN 3-642-05135-9 (print), 3-642-05136-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05136-4_5.pdf.

Athreya:2000:MPC

[Ath00] Siva Athreya. Monotonicity property for a class of semilinear partial differential equations. Lecture Notes in Mathematics, 1729: 388–392, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103815/.

Attal:2006:DPA

[Att06a] Stéphane Attal. Disparition de Paul-André Meyer. (French) []. Lecture Notes in Mathematics, 1874:27–34, 2006. CO-DEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_3.pdf.

Attal:2006:EOA

[Att06b] Stéphane Attal. Elements of operator algebras and modular theory. Lecture Notes in Mathematics, 1880:69–105, 2006. CODEN LNMAA2. ISBN 3-540-30991-8 (print), 3-540-33922-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-33922-1_3.pdf.

Attal:2006:QN

[Att06c] Stéphane Attal. Quantum noises. Lecture Notes in Mathematics, 1881:79–147, 2006. CODEN LNMAA2. ISBN 3-540-30992-6 (print), 3-540-33966-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-33966-3_3.pdf.

Amoroso:2003:DAL

[AZ03] Francesco Amoroso and Umberto Zannier, editors. Diophantine Approximation: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, June 28–July 6, 2000, volume 1819 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CO-DEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1819. URL http://link.springer.com/book/
10.1007/3-540-44979-5;   http://www.springerlink.com/
content/978-3-540-44979-9.

Back:2004:IAI

[Bac04] Kerry Back. Incomplete and asymmetric information in asset pricing theory. Lecture Notes in Mathematics, 1856:1–25, 2004. CODEN LNMAA2. ISBN 3-540-22953-1 (print), 3-540-44644-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44644-6_1/.

Bailleul:2008:PSR

[Bai08] Ismael Bailleul. Une preuve simple d’un résultat de Dufresne. (French) []. Lecture Notes in Mathematics, 1934:203–213, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_10.pdf.

Ball:2000:RSP

[Bal00] K. Ball. A remark on the slicing problem. Lecture Notes in Mathematics, 1745:21–26, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0107205/.

Ball:2004:EIM

[Bal04] K. Ball. An elementary introduction to monotone transportation. Lecture Notes in Mathematics, 1850:41–52, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44489-3_
5/.

Banasiak:2008:PNS

[Ban08] Jacek Banasiak. Positivity in natural sciences. Lecture Notes in Mathematics, 1940:1–89, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78362-6_1.pdf.

Bangert:2009:BK

[Ban09] Patrick D. Bangert. Braids and knots. Lecture Notes in Mathematics, 1973:1–73, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00837-5_1.pdf.

Barvinok:2003:ANP

[Bar03] Alexander Barvinok. Approximating a norm by a polynomial. Lecture Notes in Mathematics, 1807:20–26, 2003. CODEN LN-MAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36428-3_
2/.

Barthe:2004:CVB

[Bar04] F. Barthe. A continuous version of the Brascamp–Lieb inequalities. Lecture Notes in Mathematics, 1850:53–63, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44489-3_
6/.

Barchielli:2006:CMQ

[Bar06] Alberto Barchielli. Continual measurements in quantum mechanics and quantum stochastic calculus. Lecture Notes in Mathematics, 1882:207–292, 2006. CODEN LNMAA2. ISBN 3-540-30993-4 (print), 3-540-33967-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-33967-1_5.pdf.

Bass:2003:SDE

[Bas03] Richard F. Bass. Stochastic differential equations driven by symmetric stable processes. Lecture Notes in Mathematics, 1801: 302–313, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_11/.

Baudoin:2003:MAF

[Bau03] Fabrice Baudoin. Modeling anticipations on financial markets. Lecture Notes in Mathematics, 1814:43–94, 2003. CODEN LN-MAA2. ISBN 3-540-40193-8 (print), 3-540-44859-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44859-4_
2/.

Bauch:2008:RMM

[Bau08] Chris T. Bauch. The role of mathematical models in explaining recurrent outbreaks of infectious childhood diseases. Lecture Notes in Mathematics, 1945:297–319, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78911-6_11.pdf.

Bolthausen:2007:SG

[BB07] Erwin Bolthausen and Anton Bovier. Spin Glasses, volume 1900 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 179 pp. LCCN QA3 .L28 no. 1900. URL http://

link.springer.com/book/10.1007/3-540-40902-5;   http:
//www.springerlink.com/content/978-3-540-40908-3.

Bogdan:2009:BPT

[BB09] K. Bogdan and T. Byczkowski. Boundary potential theory for Schrödinger operators based on fractional Laplacian. Lecture Notes in Mathematics, 1980:25–55, 2009. CODEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-02141-1_2.
pdf.

Biane:2008:FM

[BBC^{+}08a] Philippe Biane, Luc Bouten, Fabio Cipriani, Norio Konno, Nicolas Privault, and Quanhua Xu. Front matter. Lecture Notes in Mathematics, 1954:i–xi, 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-69365-9/1.pdf.

Biane:2008:I

[BBC^{+}08b] Philippe Biane, Luc Bouten, Fabio Cipriani, Norio Konno, Nicolas Privault, and Quanhua Xu. Introduction. Lecture Notes in Mathematics, 1954:1–2, 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69365-9_1.pdf.

Boffi:2008:FM

[BBD^{+}08] Daniele Boffi, Franco Brezzi, Leszek F. Demkowicz, Ricardo G. Durán, Richard S. Falk, and Michel Fortin. Front matter. Lecture Notes in Mathematics, 1939:i–x, 2008. CODEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-78319-0/1.pdf.

Bank:2003:PPL

[BBF^{+}03] Peter Bank, Fabrice Baudoin, Hans Föllmer, L. C. G. Rogers, Mete Soner, and Nizar Touzi, editors. Paris–Princeton Lectures on Mathematical Finance 2002, volume 1814 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40193-8 (print), 3-540-44859-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1814. URL

http://link.springer.com/book/10.1007/b12041;   http:
//www.springerlink.com/content/978-3-540-44859-4.

Boffi:2008:FES

[BBF08] Daniele Boffi, Franco Brezzi, and Michel Fortin. Finite elements for the Stokes problem. Lecture Notes in Mathematics, 1939:45– 100, 2008. CODEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78319-0_2.pdf.

Back:2004:SMF

[BBH^{+}04] Kerry Back, Tomasz R. Bielecki, Christian Hipp, Shige Peng, and Walter Schachermayer, editors. Stochastic Methods in Finance: Lectures given at the C.I.M.E.-E.M.S. Summer School held in Bressanone/Brixen, Italy, July 6-12, 2003, volume 1856 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22953-1 (print), 3-540-44644-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28

no. 1856. URL http://link.springer.com/book/10.1007/
b100122;   http://www.springerlink.com/content/978-3-
540-44644-6.

Biskup:2009:FM

[BBH^{+}09] Marek Biskup, Anton Bovier, Frank Hollander, Dima Ioffe, Fabio Martinelli, Karel Netocný, and Christina Toninelli. Front matter. Lecture Notes in Mathematics, 1970:1–8, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-92796-9/
1.pdf.

Bogdan:2009:FM

[BBK^{+}09] Krzysztof Bogdan, Tomasz Byczkowski, Tadeusz Kulczycki, Michal Ryznar, Renming Song, and Zoran Vondracek. Front matter. Lecture Notes in Mathematics, 1980:1–8, 2009. CODEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-02141-1/
1.pdf.

Barlow:2001:CSB

[BBKM01] Martin Barlow, Krzysztof Burdzy, Haya Kaspi, and Avi Mandelbaum. Coalescence of skew Brownian motions. Lecture Notes in Mathematics, 1755:202–205, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44671-2_15/.

Baddeley:2007:BM

[BBS07a] Adrian Baddeley, Imre Bárány, and Rolf Schneider. Back matter. Lecture Notes in Mathematics, 1892:283–292, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-38175-4/
1.pdf.

Baddeley:2007:FM

[BBS07b] Adrian Baddeley, Imre Bárány, and Rolf Schneider. Front matter. Lecture Notes in Mathematics, 1892:I–XII, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-38175-4/
1.pdf.

Baddeley:2007:IGT

[BBS07c] Adrian Baddeley, Imre Bárány, and Rolf Schneider. Integral geometric tools for stochastic geometry. Lecture Notes in Mathematics, 1892:119–184, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-38175-4_3.pdf.

Baddeley:2007:EEM

[BBS07d] Adrian Baddeley, Imre Bárány, and Rolf Schneider. On the evolution equations of mean geometric densities for a class of space and time inhomogeneous stochastic birth-and-growth processes. Lecture Notes in Mathematics, 1892:267–281, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-38175-4_6.pdf.

Baddeley:2007:RM

[BBS07e] Adrian Baddeley, Imre Bárány, and Rolf Schneider. Random mosaics. Lecture Notes in Mathematics, 1892:247–266, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-38175-4_5.pdf.

Baddeley:2007:RPC

[BBS07f] Adrian Baddeley, Imre Bárány, and Rolf Schneider. Random polytopes, convex bodies, and approximation. Lecture Notes in Mathematics, 1892:77–118, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38175-4_2.pdf.

Baddeley:2007:RSP

[BBS07g] Adrian Baddeley, Imre Bárány, and Rolf Schneider. Random sets (in particular Boolean models). Lecture Notes in Mathematics, 1892:185–245, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-38175-4_4.pdf.

Baddeley:2007:SPP

[BBS07h] Adrian Baddeley, Imre Bárány, and Rolf Schneider. Spatial point processes and their applications. Lecture Notes in Mathematics, 1892:1–75, 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-38175-4_1.pdf.

Baddeley:2007:SGL

[BBSW07] Adrian Baddeley, Imre Bárány, Rolf Schneider, and Wolfgang Weil, editors. Stochastic Geometry: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy, September 13–18, 2004, volume 1892 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-38174-0 (print), 3-540-38175-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1892; QA273.5 .S742 2007. URL http://link.springer.

com/book/10.1007/3-540-38174-0;   http://www.loc.gov/
catdir/enhancements/fy0825/2006931679-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0825/2006931679-t.
html;   http://www.springerlink.com/content/978-3-540-
38175-4.

Brzezniak:2003:AWZ

[BC03] Zdzisław Brzeźniak and Andrew Carroll. Approximations of the Wong–Zakai type for stochastic differential equations in M-type 2 Banach spaces with applications to loop spaces. Lecture Notes in Mathematics, 1832:251–289, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40004-2_11/.

Barthe:2004:IBL

[BCE04] F. Barthe and D. Cordero-Erausquin. Inverse Brascamp–Lieb inequalities along the heat equation. Lecture Notes in Mathematics, 1850:65–71, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44489-3_7/.

Bobylev:2008:GKM

[BCG08] A. V. Bobylev, C. Cercignani, and I. M. Gamba. Generalized kinetic Maxwell type models of granular gases. Lecture Notes in Mathematics, 1937:23–57, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78277-3_2.pdf.

Banasiak:2008:FM

[BCM08] Jacek Banasiak, Mark A. J. Chaplain, and Jacek Miekisz. Front matter. Lecture Notes in Mathematics, 1940:i–xii, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-78362-6/
1.pdf.

Bernot:2009:AEI

[BCM09a] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Application: Embedded irrigation networks. Lecture Notes in Mathematics, 1955:169–177, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69315-4_14.pdf.

Bernot:2009:BM

[BCM09b] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Back matter. Lecture Notes in Mathematics, 1955:185–206, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-69315-4/
1.pdf.

Bernot:2009:DLS

[BCM09c] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Dirac to Lebesgue segment: a case study. Lecture Notes in Mathematics, 1955:151–168, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69315-4_13.pdf.

Bernot:2009:EVM

[BCM09d] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. The equivalence of various models. Lecture Notes in Mathematics, 1955:95– 104, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_9.pdf.

Bernot:2009:FM

[BCM09e] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Front matter. Lecture Notes in Mathematics, 1955:I–X, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-69315-4/
1.pdf.

Bernot:2009:GSP

[BCM09f] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. The Gilbert–Steiner problem. Lecture Notes in Mathematics, 1955: 135–149, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-69315-4_12.pdf.

Bernot:2009:IBR

[BCM09g] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Interior and boundary regularity. Lecture Notes in Mathematics, 1955:79– 93, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_8.pdf.

Bernot:2009:IM

[BCM09h] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Introduction: The models. Lecture Notes in Mathematics, 1955:1–9, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_1.pdf.

Bernot:2009:ID

[BCM09i] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Irrigability and dimension. Lecture Notes in Mathematics, 1955:105–117, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_10.pdf.

Bernot:2009:LOP

[BCM09j] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. The landscape of an optimal pattern. Lecture Notes in Mathematics, 1955:119–134, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69315-4_11.pdf.

Bernot:2009:MM

[BCM09k] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. The mathematical models. Lecture Notes in Mathematics, 1955:11–23, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_2.pdf.

Bernot:2009:OP

[BCM09l] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Open problems. Lecture Notes in Mathematics, 1955:179–183, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_15.pdf.

Bernot:2009:OTP

[BCM09m] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Operations on traffic plans. Lecture Notes in Mathematics, 1955:47–54, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_5.pdf.

Bernot:2009:OTN

[BCM09n] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Optimal Transportation Networks: Models and Theory, volume 1955 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CO-DEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 200 pp. LCCN QA3 .L28 no. 1955. URL http://link.

springer.com/book/10.1007/978-3-540-69315-4;   http:/
/www.springerlink.com/content/978-3-540-69315-4.

Bernot:2009:SOT

[BCM09o] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. The structure of optimal traffic plans. Lecture Notes in Mathematics, 1955:39–45, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69315-4_4.pdf.

Bernot:2009:TP

[BCM09p] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Traffic plans. Lecture Notes in Mathematics, 1955:25–37, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_3.pdf.

Bernot:2009:TPD

[BCM09q] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. Traffic plans and distances between measures. Lecture Notes in Mathematics, 1955:55–63, 2009. CODEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69315-4_6.pdf.

Bernot:2009:TSO

[BCM09r] Marc Bernot, Vicent Caselles, and Jean-Michel Morel. The tree structure of optimal traffic plans and their approximation. Lecture Notes in Mathematics, 1955:65–78, 2009. CO-DEN LNMAA2. ISBN 3-540-69314-9 (print), 3-540-69315-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69315-4_7.pdf.

Barthe:2003:NSP

[BCN03] F. Barthe, M. Csörnyei, and A. Naor. A note on simultaneous polar and cartesian decomposition. Lecture Notes in Mathematics, 1807:1–19, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36428-3_1/.

Buttazzo:2003:OSM

[BD03] Giuseppe Buttazzo and Luigi De Pascale. Optimal shapes and masses, and optimal transportation problems. Lecture Notes in Mathematics, 1813:11–51, 2003. CODEN LNMAA2. ISBN 3-540-40192-X (print), 3-540-44857-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44857-0_2/.

Bobenko:2000:PED

[BE00] Alexander I. Bobenko and Ulrich Eitner. Painlevé Equations in the Differential Geometry of Surfaces, volume 1753 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41414-2 (paperback), 3-540-44452-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 120 pp. LCCN QA3 .L28 no. 1753; QA3 .L28 no. 1753; QA649. URL http://

link.springer.com/book/10.1007/b76883.

Beckermann:2006:DOP

[Bec06] Bernhard Beckermann. Discrete orthogonal polynomials and superlinear convergence of Krylov subspace methods in numerical linear algebra. Lecture Notes in Mathematics, 1883:119–185, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36716-1_3.pdf.

Behrend:2002:I

[Beh02a] K. Behrend. Introduction. Lecture Notes in Mathematics, 1776: 1–2, 2002. CODEN LNMAA2. ISBN 3-540-43121-7 (print), 3-540-45617-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45617-9_1.

Behrend:2002:LGW

[Beh02b] K. Behrend. Localization and Gromov–Witten invariants. Lecture Notes in Mathematics, 1776:3–38, 2002. CODEN LNMAA2. ISBN 3-540-43121-7 (print), 3-540-45617-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45617-9_2/.

Bellet:2006:EPM

[Bel06] Luc Rey Bellet. Ergodic properties of Markov processes. Lecture Notes in Mathematics, 1881:1–39, 2006. CODEN LNMAA2. ISBN 3-540-30992-6 (print), 3-540-33966-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-33966-3_1.pdf.

Bentaleb:2003:FEI

[Ben03] Abdellatif Bentaleb. Sur les fonctions extrémales des inégalités de Sobolev des opérateurs de diffusion. (French) []. Lecture Notes in Mathematics, 1801:230–250, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36107-7_8/.

Benettin:2005:PAN

[Ben05] Giancarlo Benettin. Physical applications of Nekhoroshev theorem and exponential estimates. Lecture Notes in Mathematics, 1861:1–76, 2005. CODEN LNMAA2. ISBN 3-540-24064-0 (print), 3-540-31541-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31541-4_1.pdf.

Bernard:2000:LPT

[Ber00a] Pierre Bernard, editor. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXVIII — 1998, volume 1738 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1738. URL http:/

/link.springer.com/book/10.1007/BFb0106703;   http://
www.springerlink.com/content/978-3-540-45029-0.

Bertram:2000:GJL

[Ber00b] Wolfgang Bertram. The Geometry of Jordan and Lie Structures, volume 1754 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41426-6 (paperback), 3-540-44458-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 269 pp. LCCN QA3 .L28 no. 1754; QA3 .L28 no. 1754; QA252.5. URL http:/

/link.springer.com/book/10.1007/b76884;   http://www.
loc.gov/catdir/enhancements/fy0816/00066150-d.html;
http://www.loc.gov/catdir/enhancements/fy0816/00066150-
t.html.

Bernard:2002:LPT

[Ber02] Pierre Bernard, editor. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXIX — 1999, volume 1781 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1781. URL http:/

/link.springer.com/book/10.1007/b93152;   http://www.
springerlink.com/content/978-3-540-47944-4.

Bertram:2005:MMS

[Ber05] R. Bertram. Mathematical models of synaptic transmission and short-term plasticity. Lecture Notes in Mathematics, 1867:173– 202, 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11406501_7.pdf.

Berinde:2007:BM

[Ber07a] Vasile Berinde. Back matter. Lecture Notes in Mathematics, 1912:221–322, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-72234-2/1.pdf.

Berinde:2007:EAF

[Ber07b] Vasile Berinde. Error analysis of fixed point iteration procedures. Lecture Notes in Mathematics, 1912:199–220, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72234-2_9.pdf.

Berinde:2007:FM

[Ber07c] Vasile Berinde. Front matter. Lecture Notes in Mathematics, 1912:I–XVII, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-72234-2/1.pdf.

Berinde:2007:II

[Ber07d] Vasile Berinde. The Ishikawa iteration. Lecture Notes in Mathematics, 1912:113–134, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72234-2_5.pdf.

Berinde:2007:IAF

[Ber07e] Vasile Berinde. Iterative Approximation of Fixed Points, volume 1912 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 322 pp. LCCN QA3 .L28 no. 1912. URL http://link.

springer.com/book/10.1007/978-3-540-72234-2;   http:/
/www.springerlink.com/content/978-3-540-72234-2.

Berinde:2007:ISN

[Ber07f] Vasile Berinde. Iterative solution of nonlinear operator equations. Lecture Notes in Mathematics, 1912:179–198, 2007. CO-DEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72234-2_8.pdf.

Berinde:2007:KI

[Ber07g] Vasile Berinde. The Krasnoselskij iteration. Lecture Notes in Mathematics, 1912:63–88, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72234-2_3.pdf.

Berinde:2007:MI

[Ber07h] Vasile Berinde. The Mann iteration. Lecture Notes in Mathematics, 1912:89–112, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72234-2_4.pdf.

Berinde:2007:OFP

[Ber07i] Vasile Berinde. Other fixed point iteration procedures. Lecture Notes in Mathematics, 1912:135–156, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-72234-2_6.
pdf.

Berinde:2007:PI

[Ber07j] Vasile Berinde. The Picard iteration. Lecture Notes in Mathematics, 1912:31–62, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72234-2_2.pdf.

Berinde:2007:PRF

[Ber07k] Vasile Berinde. Pre-requisites of fixed points. Lecture Notes in Mathematics, 1912:3–30, 2007. CODEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72234-2_1.pdf.

Berinde:2007:SFP

[Ber07l] Vasile Berinde. Stability of fixed point iteration procedures. Lecture Notes in Mathematics, 1912:157–178, 2007. CO-DEN LNMAA2. ISBN 3-540-72233-5 (print), 3-540-72234-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72234-2_7.pdf.

Berger:2009:TQC

[Ber09] Mitchell A. Berger. Topological quantities: Calculating winding, writhing, linking, and higher order invariants. Lecture Notes in Mathematics, 1973:75–97, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-00837-5_2.pdf.

Beyer:2007:BM

[Bey07a] Horst Reinhard Beyer. Back matter. Lecture Notes in Mathematics, 1898:265–287, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-71129-2/1.pdf.

Beyer:2007:BPD

[Bey07b] Horst Reinhard Beyer. Beyond Partial Differential Equations: On Linear and Quasi-Linear Abstract Hyperbolic Evolution Equations, volume 1898 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 283 pp. LCCN QA3 .L28 no. 1898. URL http:

//link.springer.com/book/10.1007/978-3-540-71129-2;
http://www.springerlink.com/content/978-3-540-71129-2.

Beyer:2007:C

[Bey07c] Horst Reinhard Beyer. Conventions. Lecture Notes in Mathematics, 1898:1–3, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71129-2_1.pdf.

Beyer:2007:ECS

[Bey07d] Horst Reinhard Beyer. Examples of constrained systems. Lecture Notes in Mathematics, 1898:123–135, 2007. CO-DEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71129-2_7.pdf.

Beyer:2007:EGS

[Bey07e] Horst Reinhard Beyer. Examples of generators of strongly continuous semigroups. Lecture Notes in Mathematics, 1898:71–103, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71129-2_5.pdf.

Beyer:2007:ELE

[Bey07f] Horst Reinhard Beyer. Examples of linear evolution equations. Lecture Notes in Mathematics, 1898:177–214, 2007. CO-DEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71129-2_10.pdf.

Beyer:2007:EQL

[Bey07g] Horst Reinhard Beyer. Examples of quasi-linear evolution equations. Lecture Notes in Mathematics, 1898:235–263, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71129-2_12.pdf.

Beyer:2007:FM

[Bey07h] Horst Reinhard Beyer. Front matter. Lecture Notes in Mathematics, 1898:i–xiv, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-71129-2/1.pdf.

Beyer:2007:IRO

[Bey07i] Horst Reinhard Beyer. Intertwining relations, operator homomorphisms. Lecture Notes in Mathematics, 1898:105–121, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71129-2_6.pdf.

Beyer:2007:KCE

[Bey07j] Horst Reinhard Beyer. Kernels, chains, and evolution operators. Lecture Notes in Mathematics, 1898:137–163, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71129-2_8.pdf.

Beyer:2007:LEE

[Bey07k] Horst Reinhard Beyer. The linear evolution equation. Lecture Notes in Mathematics, 1898:165–176, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71129-2_9.
pdf.

Beyer:2007:MI

[Bey07l] Horst Reinhard Beyer. Mathematical introduction. Lecture Notes in Mathematics, 1898:5–12, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71129-2_2.pdf.

Beyer:2007:P

[Bey07m] Horst Reinhard Beyer. Prerequisites. Lecture Notes in Mathematics, 1898:13–39, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71129-2_3.pdf.

Beyer:2007:QLE

[Bey07n] Horst Reinhard Beyer. The quasi-linear evolution equation. Lecture Notes in Mathematics, 1898:215–234, 2007. CO-DEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71129-2_11.pdf.

Beyer:2007:SCS

[Bey07o] Horst Reinhard Beyer. Strongly continuous semigroups. Lecture Notes in Mathematics, 1898:41–69, 2007. CODEN LNMAA2. ISBN 3-540-71128-7 (print), 3-540-71129-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71129-2_4.
pdf.

Bank:2003:AOM

[BF03] Peter Bank and Hans Föllmer. American options, multi–armed bandits, and optimal consumption plans: a unifying view. Lecture Notes in Mathematics, 1814:1–42, 2003. CODEN LNMAA2. ISBN 3-540-40193-8 (print), 3-540-44859-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44859-4_1/.

Bunge:2006:BM

[BF06a] Marta Bunge and Jonathon Funk. Back matter. Lecture Notes in Mathematics, 1890:217–229, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-36360-6/1.pdf.

Bunge:2006:CLK

[BF06b] Marta Bunge and Jonathon Funk. Closed and linear KZmonads. Lecture Notes in Mathematics, 1890:109–127, 2006.

CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36360-6_6.pdf.

Bunge:2006:CSM

[BF06c] Marta Bunge and Jonathon Funk. Complete spread maps of toposes. Lecture Notes in Mathematics, 1890:31–54, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36360-6_2.pdf.

Bunge:2006:CSD

[BF06d] Marta Bunge and Jonathon Funk. Complete spreads as discrete M-fibrations. Lecture Notes in Mathematics, 1890:99–108, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36360-6_5.pdf.

Bunge:2006:CKM

[BF06e] Marta Bunge and Jonathon Funk. Completion KZ-monads. Lecture Notes in Mathematics, 1890:79–97, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-36360-6_4.
pdf.

Bunge:2006:FMa

[BF06f] Marta Bunge and Jonathon Funk. Front matter. Lecture Notes in Mathematics, 1890:I–XVII, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-36360-6/1.pdf.

Bunge:2006:FMb

[BF06g] Marta Bunge and Jonathon Funk. Front matter. Lecture Notes in Mathematics, 1890:8, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-36360-6/1/1.pdf.

Bunge:2006:FMc

[BF06h] Marta Bunge and Jonathon Funk. Front matter. Lecture Notes in Mathematics, 1890:78, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-36360-6/2/1.pdf.

Bunge:2006:FMd

[BF06i] Marta Bunge and Jonathon Funk. Front matter. Lecture Notes in Mathematics, 1890:130, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-36360-6/3/1.pdf.

Bunge:2006:LTA

[BF06j] Marta Bunge and Jonathon Funk. Lattice-theoretic aspects. Lecture Notes in Mathematics, 1890:131–159, 2006. CO-DEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36360-6_7.pdf.

Bunge:2006:LDT

[BF06k] Marta Bunge and Jonathon Funk. Lawvere distributions on toposes. Lecture Notes in Mathematics, 1890:9–29, 2006. CO-DEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36360-6_1.pdf.

Bunge:2006:LAA

[BF06l] Marta Bunge and Jonathon Funk. Localic and algebraic aspects. Lecture Notes in Mathematics, 1890:161–188, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36360-6_8.pdf.

Bunge:2006:SCT

[BF06m] Marta Bunge and Jonathon Funk. Singular Coverings of Toposes, volume 1890 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 225 pp. LCCN QA3 .L28 no. 1890. URL http://

link.springer.com/book/10.1007/3-540-36359-9;   http:
//www.springerlink.com/content/978-3-540-36360-6.

Bunge:2006:SCC

[BF06n] Marta Bunge and Jonathon Funk. The spread and completeness conditions. Lecture Notes in Mathematics, 1890:55–76, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36360-6_3.pdf.

Bunge:2006:TA

[BF06o] Marta Bunge and Jonathon Funk. Topological aspects. Lecture Notes in Mathematics, 1890:189–215, 2006. CODEN LNMAA2. ISBN 3-540-36359-9 (print), 3-540-36360-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-36360-6_9.
pdf.

Borisyuk:2005:TMB

[BFET05] Alla Borisyuk, Avner Friedman, Bard Ermentrout, and David Terman. Tutorials in Mathematical Biosciences I: Mathematical Neuroscience, volume 1860 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-23858-1 (print), 3-540-31544-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 170 pp. LCCN QA3 .L28 no. 1860.

URL http://link.springer.com/book/10.1007/b102786;
http://www.springerlink.com/content/978-3-540-31544-5.

Burstall:2002:CGS

[BFL^{+}02] Francis E. Burstall, Dirk Ferus, Katrin Leschke, Franz Pedit, and Ulrich Pinkall. Conformal Geometry of Surfaces in S^{4} and Quaternions, volume 1772 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 89 pp. LCCN QA3 .L28 no. 1772. URL

http://link.springer.com/book/10.1007/b82935;   http:
//www.springerlink.com/content/978-3-540-45301-7.

Boffi:2008:MFE

[BG08] Daniele Boffi and Lucia Gastaldi, editors. Mixed Finite Elements, Compatibility Conditions, and Applications: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy June 26– July 1, 2006, volume 1939 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1939. URL http://link.

springer.com/book/10.1007/978-3-540-78319-0;   http:/
/www.springerlink.com/content/978-3-540-78319-0.

Behrend:2002:QCL

[BGTT02] Kai Behrend, César Gómez, Vitaly Tarasov, and Gang Tian, editors. Quantum Cohomology: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, June 30-July 8, 1997, volume 1776 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43121-7 (print), 3-540-45617-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1776. URL http:/

/link.springer.com/book/10.1007/b82946;   http://www.
springerlink.com/content/978-3-540-45617-9.

Bakry:2008:HPR

[BH08] D. Bakry and N. Huet. The hypergroup property and representation of Markov kernels. Lecture Notes in Mathematics, 1934:295– 347, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_15.pdf.

Bhat:2005:DCP

[Bha05] B. V. Rajarama Bhat. Dilations, cocycles and product systems. Lecture Notes in Mathematics, 1865:273–291, 2005. CODEN LNMAA2. ISBN 3-540-24406-9 (print), 3-540-31450-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11376569_4.pdf.

Bultheel:2006:ORF

[BHGVN06] Adhemar Bultheel, Erik Hendriksen, Pablo González-Vera, and Olav Njåstad. Orthogonal rational functions on the unit circle: from the scalar to the matrix case. Lecture Notes in Mathematics, 1883:187–228, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-36716-1_4.pdf.

Benettin:2005:HDT

[BHK05] Giancarlo Benettin, Jacques Henrard, and Sergei Kuksin, editors. Hamiltonian Dynamics. Theory and Applications: Lectures given at the C.I.M.E.-E.M.S. Summer School held in Cetraro, Italy, July 1-10, 1999, volume 1861 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-24064-0 (print), 3-540-31541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1861. URL http://

link.springer.com/book/10.1007/b104338;   http://www.
springerlink.com/content/978-3-540-31541-4.

Broer:2003:Ia

[BHLV03a] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. 1. Introduction. Lecture Notes in Mathematics, 1806:1–18, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36398-9_1/.

Broer:2003:MPR

[BHLV03b] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. 2. Method i: Planar reduction. Lecture Notes in Mathematics, 1806: 21–44, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36398-9_2/.

Broer:2003:MIE

[BHLV03c] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. 3. Method II: The energy-momentum map. Lecture Notes in Mathematics, 1806:45–68, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36398-9_3/.

Broer:2003:BN

[BHLV03d] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. 4. Birkhoff normalization. Lecture Notes in Mathematics, 1806:71– 84, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36398-9_4/.

Broer:2003:ST

[BHLV03e] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. 5. Singularity theory. Lecture Notes in Mathematics, 1806:85– 96, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36398-9_5/.

Broer:2003:GBS

[BHLV03f] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. 6. Gröbner bases and standard bases. Lecture Notes in Mathematics, 1806:97–132, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36398-9_6/.

Broer:2003:CNT

[BHLV03g] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. 7. Computing normalizing transformations. Lecture Notes in Mathematics, 1806:133–151, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36398-9_7/.

Broer:2003:A

[BHLV03h] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. A. appendix. Lecture Notes in Mathematics, 1806:153–158, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36398-9_8/.

Broer:2003:BHS

[BHLV03i] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. Bifurcations in Hamiltonian Systems: Computing Singularities by Gröbner Bases, volume 1806 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 169 pp. LCCN QA3 .L28 no. 1806. URL

http://link.springer.com/book/10.1007/b10414;   http:
//www.springerlink.com/content/978-3-540-36398-9.

Broer:2003:Ib

[BHLV03j] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. Index. Lecture Notes in Mathematics, 1806:167–169, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36398-9_10/.

Broer:2003:R

[BHLV03k] Henk Broer, Igor Hoveijn, Gerton Lunter, and Gert Vegter. References. Lecture Notes in Mathematics, 1806:159–165, 2003. CODEN LNMAA2. ISBN 3-540-00403-3 (print), 3-540-36398-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36398-9_9/.

Biane:2003:CSG

[Bia03] Philippe Biane. Characters of symmetric groups and free cumulants. Lecture Notes in Mathematics, 1815:185–200, 2003. CO-DEN LNMAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/3-540-44890-
X_8/.

Biane:2006:TPG

[Bia06] Philippe Biane. Le théorème de Pitman, le groupe quantique SU_{q}(2), et une question de P.-A. Meyer. (French) []. Lecture Notes in Mathematics, 1874:61–75, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35513-7_6.pdf.

Biane:2008:IRW

[Bia08] Philippe Biane. Introduction to random walks on noncommutative spaces. Lecture Notes in Mathematics, 1954:61–116, 2008.

CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69365-9_3.pdf.

Biane:2009:MVB

[Bia09a] Philippe Biane. Matrix valued Brownian motion and a paper by Pólya. Lecture Notes in Mathematics, 1979:171–185, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_7.pdf.

Biane:2009:SPH

[Bia09b] Philippe Biane. Shabat polynomials and harmonic measure. Lecture Notes in Mathematics, 1979:147–151, 2009. CO-DEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_5.pdf.

Bildhauer:2003:Ia

[Bil03a] Michael Bildhauer. 1. Introduction. Lecture Notes in Mathematics, 1818:1–12, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44885-3_1/.

Bildhauer:2003:VPLa

[Bil03b] Michael Bildhauer. 2. Variational problems with linear growth: the general setting. Lecture Notes in Mathematics, 1818:13– 39, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44885-3_2/.

Bildhauer:2003:VIG

[Bil03c] Michael Bildhauer. 3. Variational integrands with (s, μ, q)-growth. Lecture Notes in Mathematics, 1818:41–96, 2003. CODEN LN-MAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44885-3_

Bildhauer:2003:VPLb

[Bil03d] Michael Bildhauer. 4. Variational problems with linear growth: the case of μ-elliptic integrands. Lecture Notes in Mathematics, 1818:97–139, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44885-3_4/.

Bildhauer:2003:BSC

[Bil03e] Michael Bildhauer. 5. Bounded solutions for convex variational problems with a wide range of anisotropy. Lecture Notes in Mathematics, 1818:141–159, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44885-3_5/.

Bildhauer:2003:ALS

[Bil03f] Michael Bildhauer. 6. Anisotropic linear/superlinear growth in the scalar case. Lecture Notes in Mathematics, 1818:161– 172, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44885-3_6/.

Bildhauer:2003:SRR

[Bil03g] Michael Bildhauer. A. some remarks on relaxation. Lecture Notes in Mathematics, 1818:173–183, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44885-3_7/.

Bildhauer:2003:BSD

[Bil03h] Michael Bildhauer. B. some density results. Lecture Notes in Mathematics, 1818:185–198, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44885-3_8/.

Bildhauer:2003:CBC

[Bil03i] Michael Bildhauer. C. Brief comments on steady states of generalized Newtonian fluids. Lecture Notes in Mathematics, 1818: 199–203, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44885-3_9/.

Bildhauer:2003:CVP

[Bil03j] Michael Bildhauer. Convex Variational Problems: Linear, Nearly Linear and Anisotropic Growth Conditions, volume 1818 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 217 pp. LCCN QA3 .L28 no. 1818. URL http:/

/link.springer.com/book/10.1007/b12308;   http://www.
springerlink.com/content/978-3-540-44885-3.

Bildhauer:2003:DNC

[Bil03k] Michael Bildhauer. D. notation and conventions. Lecture Notes in Mathematics, 1818:205–206, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44885-3_10.

Bildhauer:2003:Ib

[Bil03l] Michael Bildhauer. Index. Lecture Notes in Mathematics, 1818: 215–217, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44885-3_12/.

Bildhauer:2003:R

[Bil03m] Michael Bildhauer. References. Lecture Notes in Mathematics, 1818:207–213, 2003. CODEN LNMAA2. ISBN 3-540-40298-5 (print), 3-540-44885-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44885-3_11/.

Bishwal:2008:AML

[Bis08a] Jaya P. N. Bishwal. Approximate maximum likelihood estimation in nonhomogeneous diffusions. Lecture Notes in Mathematics, 1923:125–157, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74448-1_7.pdf.

Bishwal:2008:BM

[Bis08b] Jaya P. N. Bishwal. Back matter. Lecture Notes in Mathematics, 1923:245–264, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-74448-1/1.pdf.

Bishwal:2008:BSE

[Bis08c] Jaya P. N. Bishwal. Bayes and sequential estimation in stochastic PDEs. Lecture Notes in Mathematics, 1923:79–97, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74448-1_5.pdf.

Bishwal:2008:EFD

[Bis08d] Jaya P. N. Bishwal. Estimating function for discretely observed homogeneous diffusions. Lecture Notes in Mathematics, 1923:225– 244, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74448-1_10.pdf.

Bishwal:2008:FMa

[Bis08e] Jaya P. N. Bishwal. Front matter. Lecture Notes in Mathematics, 1923:I–XII, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-74448-1/1.pdf.

Bishwal:2008:FMb

[Bis08f] Jaya P. N. Bishwal. Front matter. Lecture Notes in Mathematics, 1923:13, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-74448-1/1/1.pdf.

Bishwal:2008:FMc

[Bis08g] Jaya P. N. Bishwal. Front matter. Lecture Notes in Mathematics, 1923:123, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-74448-1/2/1.pdf.

Bishwal:2008:LDE

[Bis08h] Jaya P. N. Bishwal. Large deviations of estimators in homogeneous diffusions. Lecture Notes in Mathematics, 1923:49–60, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74448-1_3.pdf.

Bishwal:2008:LAM

[Bis08i] Jaya P. N. Bishwal. Local asymptotic mixed normality for nonhomogeneous diffusions. Lecture Notes in Mathematics, 1923:61–78, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74448-1_4.pdf.

Bishwal:2008:LAN

[Bis08j] Jaya P. N. Bishwal. Local asymptotic normality for discretely observed homogeneous diffusions. Lecture Notes in Mathematics, 1923:201–223, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74448-1_9.pdf.

Bishwal:2008:MLE

[Bis08k] Jaya P. N. Bishwal. Maximum likelihood estimation in fractional diffusions. Lecture Notes in Mathematics, 1923:99–122, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74448-1_6.pdf.

Bishwal:2008:PES

[Bis08l] Jaya P. N. Bishwal. Parameter Estimation in Stochastic Differential Equations, volume 1923 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 264 pp. LCCN QA3 .L28 no.

1923. URL http://link.springer.com/book/10.1007/978-
3-540-74448-1;   http://www.springerlink.com/content/
978-3-540-74448-1.

Bishwal:2008:PSD

[Bis08m] Jaya P. N. Bishwal. Parametric stochastic differential equations. Lecture Notes in Mathematics, 1923:1–11, 2008. CO-DEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74448-1_1.pdf.

Bishwal:2008:RWCa

[Bis08n] Jaya P. N. Bishwal. Rates of weak convergence of estimators in homogeneous diffusions. Lecture Notes in Mathematics, 1923:15– 48, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74448-1_2.pdf.

Bishwal:2008:RWCb

[Bis08o] Jaya P. N. Bishwal. Rates of weak convergence of estimators in the Ornstein–Uhlenbeck process. Lecture Notes in Mathematics, 1923:159–200, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74448-1_8.pdf.

Biskup:2009:RPP

[Bis09] Marek Biskup. Reflection positivity and phase transitions in lattice spin models. Lecture Notes in Mathematics, 1970:1–86, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-92796-9_1.pdf.

Bourgain:2000:ALB

[BJ00] J. Bourgain and S. Jitomirskaya. Anderson localization for the band model. Lecture Notes in Mathematics, 1745:67–79, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0107208/.

Bjork:2004:GIR

[Bjö04] Tomas Björk. On the geometry of interest rate models. Lecture Notes in Mathematics, 1847:133–215, 2004. CODEN LNMAA2.

ISBN 3-540-22266-9 (print), 3-540-44468-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44468-8_2/.

Bielecki:2004:HDC

[BJR04a] Tomasz R. Bielecki, Monique Jeanblanc, and Marek Rutkowski. Hedging of defaultable claims. Lecture Notes in Mathematics, 1847:1–132, 2004. CODEN LNMAA2. ISBN 3-540-22266-9 (print), 3-540-44468-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44468-8_1/.

Bielecki:2004:MVC

[BJR04b] Tomasz R. Bielecki, Monique Jeanblanc, and Marek Rutkowski. Modeling and valuation of credit risk. Lecture Notes in Mathematics, 1856:27–126, 2004. CODEN LNMAA2. ISBN 3-540-22953-1 (print), 3-540-44644-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44644-6_2/.

Burkard:2000:CMD

[BJS^{+}00] Rainer E. Burkard, Antony Jameson, Gilbert Strang, Peter Deuflhard, Jacques-Louis Lions, Vincenzo Capasso, Jacques Periaux, and Heinz W. Engl, editors. Computational Mathematics Driven by Industrial Problems: Lectures given at the 1st Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Martina Franca, Italy, June 21–27, 1999, volume 1739 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1739.

URL http://link.springer.com/book/10.1007/BFb0103915;
 http://www.springerlink.com/content/978-3-540-44976-
8.

Ballyk:2008:BM

[BJS08a] M. Ballyk, D. Jones, and H. L. Smith. Back matter. Lecture Notes in Mathematics, 1936:303–306, 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-78273-5/1.pdf.

Ballyk:2008:BMF

[BJS08b] M. Ballyk, D. Jones, and H. L. Smith. The biofilm model of Freter: a review. Lecture Notes in Mathematics, 1936:265–302, 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78273-5_6.pdf.

Bobkov:2003:CLP

[BK03] S. G. Bobkov and A. Koldobsky. On the central limit property of convex bodies. Lecture Notes in Mathematics, 1807:44– 52, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36428-3_5/.

Bovier:2007:MAA

[BK07] Anton Bovier and Irina Kurkova. Much ado about Derrida’s GREM. Lecture Notes in Mathematics, 1900:81–115, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-40908-3_4.pdf.

Berger:2009:FM

[BKK^{+}09a] Mitchell A. Berger, Louis H. Kauffman, Boris Khesin, H. Keith Moffatt, Renzo L. Ricca, and De Witt Sumners. Front matter. Lecture Notes in Mathematics, 1973:1–12, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-00837-5/
1.pdf.

Berger:2009:LTF

[BKK^{+}09b] Mitchell A. Berger, Louis H. Kauffman, Boris Khesin, H. Keith Moffatt, Renzo L. Ricca, and De Witt Sumners, editors. Lectures on topological fluid mechanics: lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, July 2–10, 2001, volume 1973 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CO-DEN LNMAA2. ISBN 3-642-00836-4 (paperback), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1973. URL http://link.springer.com/book/
10.1007/978-3-642-00837-5;   http://www.springerlink.
com/content/978-3-642-00837-5.

Bourgain:2004:SIC

[BKM04] J. Bourgain, B. Klartag, and V. Milman. Symmetrization and isotropic constants of convex bodies. Lecture Notes in Mathematics, 1850:101–115, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44489-3_10/.

Brossard:2008:TBC

[BL08] Jean Brossard and Christophe Leuridan. Transformations browniennes et compléments indépendants: résultats et problèmes ouverts. (French) []. Lecture Notes in Mathematics, 1934:265–278, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_13.pdf.

Bertoin:2008:CPL

[BLM08] Jean Bertoin, Alexander Lindner, and Ross Maller. On continuity properties of the law of integrals of Lévy processes. Lecture Notes in Mathematics, 1934:137–159, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-77913-1_6.pdf.

Biyikogu:2007:BM

[BLS07a] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Back matter. Lecture Notes in Mathematics, 1915:93–115, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-73510-6/
1.pdf.

Biyikogu:2007:CE

[BLS07b] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Computational experiments. Lecture Notes in Mathematics, 1915:67–75, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73510-6_5.pdf.

Biyikogu:2007:END

[BLS07c] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Eigenfunctions and nodal domains. Lecture Notes in Mathematics, 1915:29– 47, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73510-6_3.pdf.

Biyikogu:2007:FKT

[BLS07d] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Faber– Krahn type inequalities. Lecture Notes in Mathematics, 1915:77– 91, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73510-6_6.pdf.

Biyikogu:2007:FM

[BLS07e] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Front matter. Lecture Notes in Mathematics, 1915:I–VIII, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-73510-6/
1.pdf.

Biyikogu:2007:GL

[BLS07f] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Graph Laplacians. Lecture Notes in Mathematics, 1915:15–27, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73510-6_2.pdf.

Biyikogu:2007:I

[BLS07g] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Introduction. Lecture Notes in Mathematics, 1915:1–14, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73510-6_1.pdf.

Biyikogu:2007:LEG

[BLS07h] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Laplacian Eigenvectors of Graphs: Perron–Frobenius and Faber–Krahn Type Theorems, volume 1915 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 115 pp. LCCN QA3 .L28 no. 1915.

URL http://link.springer.com/book/10.1007/978-3-540-
73510-6;   http://www.springerlink.com/content/978-3-
540-73510-6.

Biyikogu:2007:NDT

[BLS07i] Türker Biyikoğu, Josef Leydold, and Peter F. Stadler. Nodal domain theorems for special graph classes. Lecture Notes in Mathematics, 1915:49–65, 2007. CODEN LNMAA2. ISBN 3-540-73509-7 (print), 3-540-73510-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-73510-6_4.pdf.

Bernard:2002:DAM

[BM02] A. Bernard and B. Maisonneuve. Decomposition atomique de martingales de la classe H^{1}. (French) []. Lecture Notes in Mathematics, 1771:394–414, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_16/.

Bakry:2003:CMS

[BM03] Dominique Bakry and Olivier Mazet. Characterization of Markov semigroups on R associated to some families of orthogonal polynomials. Lecture Notes in Mathematics, 1832:60– 80, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_2/.

Baraka:2008:LIL

[BM08a] Driss Baraka and Thomas Mountford. A law of the iterated logarithm for fractional Brownian motions. Lecture Notes in Mathematics, 1934:161–179, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-77913-1_7.pdf.

Behrend:2008:EIA

[BM08b] Kai Behrend and Marco Manetti, editors. Enumerative Invariants in Algebraic Geometry and String Theory: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy June 6–11, 2005, volume 1947 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-79813-7 (print), 3-540-79814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1947. URL http:

//link.springer.com/book/10.1007/978-3-540-79814-9;
http://www.springerlink.com/content/978-3-540-79814-9.

Bobkov:2003:CBL

[BN03] S. G. Bobkov and F. L. Nazarov. On convex bodies and logconcave probability measures with unconditional basis. Lecture Notes in Mathematics, 1807:53–69, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36428-3_6/.

Barndorff-Nielsen:2006:CFI

[BNT06] Ole E. Barndorff-Nielsen and Steen Thorbjørnsen. Classical and free infinite divisibility and Lévy processes. Lecture Notes in Mathematics, 1866:33–159, 2006. CODEN LNMAA2. ISBN 3-540-24407-7 (print), 3-540-32385-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/11376637_2.pdf.

Blanchard:2006:DID

[BO06] Philippe Blanchard and Robert Olkiewicz. Decoherence as irreversible dynamical process in open quantum systems. Lecture Notes in Mathematics, 1882:117–159, 2006. CODEN LNMAA2. ISBN 3-540-30993-4 (print), 3-540-33967-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-33967-1_3.pdf.

Bobkov:2000:RGN

[Bob00] S. G. Bobkov. Remarks on the growth of L^{p}-norms of polynomials. Lecture Notes in Mathematics, 1745:27–35, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107206/.

Bobkov:2003:CDW

[Bob03a] S. G. Bobkov. Concentration of distributions of the weighted sums with Bernoullian coefficients. Lecture Notes in Mathematics, 1807:27–36, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36428-3_3/.

Bobkov:2003:SGC

[Bob03b] S. G. Bobkov. Spectral gap and concentration for some spherically symmetric probability measures. Lecture Notes in Mathematics, 1807:37–43, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36428-3_4/.

Bobkov:2007:ICL

[Bob07a] Sergey G. Bobkov. On isoperimetric constants for log-concave probability distributions. Lecture Notes in Mathematics, 1910:81– 88, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_4.pdf.

Bobkov:2007:RSB

[Bob07b] Sergey G. Bobkov. A remark on the surface Brunn–Minkowskitype inequality. Lecture Notes in Mathematics, 1910:77–79, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_3.pdf.

Bolthausen:2007:RMS

[Bol07] Erwin Bolthausen. Random media and spin glasses: an introduction into some mathematical results and problems. Lecture Notes in Mathematics, 1900:1–44, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-40908-3_1.

pdf;   http://link.springer.com/content/pdf/bfm:978-3-
540-40908-3/1.pdf.

Bolley:2008:SCW

[Bol08] F. Bolley. Separability and completeness for the Wasserstein distance. Lecture Notes in Mathematics, 1934:371–377, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_17.pdf.

Bonilla:2008:IPI

[Bon08] Luis L. Bonilla, editor. Inverse Problems and Imaging: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy September 15–21, 2002, volume 1943 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1943.

URL http://link.springer.com/book/10.1007/978-3-540-
78547-7;   http://www.springerlink.com/content/978-3-
540-78547-7.

Borodin:2003:ART

[Bor03] Alexei Borodin. Asymptotic representation theory and Riemann– Hilbert problem. Lecture Notes in Mathematics, 1815:3–19, 2003. CODEN LNMAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44890-X_1/.

Borisyuk:2005:PMM

[Bor05] Alla Borisyuk. Physiology and mathematical modeling of the auditory system. Lecture Notes in Mathematics, 1860:107–168, 2005. CODEN LNMAA2. ISBN 3-540-23858-1 (print), 3-540-31544-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31544-5_4.pdf.

Bourgain:2000:PLE

[Bou00] J. Bourgain. Positive Lyapounov exponents for most energies. Lecture Notes in Mathematics, 1745:37–66, 2000. CODEN LN-MAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107207/.

Bourgain:2003:LTB

[Bou03a] J. Bourgain. On long-time behaviour of solutions of linear Schrödinger equations with smooth time-dependent potential. Lecture Notes in Mathematics, 1807:99–113, 2003. CODEN LN-MAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36428-3_
8/.

Bourgain:2003:ICP

[Bou03b] J. Bourgain. On the isotropy-constant problem for “PSI-2”- bodies. Lecture Notes in Mathematics, 1807:114–121, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-36428-3_9/.

Bourgain:2003:RLS

[Bou03c] J. Bourgain. Random lattice Schrödinger operators with decaying potential: Some higher dimensional phenomena. Lecture Notes in Mathematics, 1807:70–98, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36428-3_7/.

Bourgain:2004:LLS

[Bou04] J. Bourgain. On localization for lattice Schrödinger operators involving Bernoulli variables. Lecture Notes in Mathematics, 1850: 77–99, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_9/.

Bourgain:2007:RQE

[Bou07a] J. Bourgain. A remark on quantum ergodicity for CAT maps. Lecture Notes in Mathematics, 1910:89–98, 2007. CO-DEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_5.pdf.

Bourgain:2007:SAA

[Bou07b] J. Bourgain. Some arithmetical applications of the sum-product theorems in finite fields. Lecture Notes in Mathematics, 1910:99– 116, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_6.pdf.

Bouten:2008:AQS

[Bou08] Lue Bouten. Applications of quantum stochastic processes in quantum optics. Lecture Notes in Mathematics, 1954:277–307, 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69365-9_6.pdf.

Bovier:2009:M

[Bov09] Anton Bovier. Metastability. Lecture Notes in Mathematics, 1970: 1–45, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-92796-9_4.pdf.

Bowen:2008:AD

[Bow08a] Rufus Bowen. Axiom a diffeomorphisms. Lecture Notes in Mathematics, 470:45–59, 2008. CODEN LNMAA2. ISBN 3-540-77605-2 (print), 3-540-77695-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77695-6_3.pdf.

Bowen:2008:BM

[Bow08b] Rufus Bowen. Back matter. Lecture Notes in Mathematics, 470: 74–78, 2008. CODEN LNMAA2. ISBN 3-540-77605-2 (print), 3-540-77695-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-77695-6/1.pdf.

Bowen:2008:ETA

[Bow08c] Rufus Bowen. Ergodic theory of axiom a diffeomorphisms. Lecture Notes in Mathematics, 470:61–73, 2008. CODEN LNMAA2. ISBN 3-540-77605-2 (print), 3-540-77695-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-77695-6_4.
pdf.

Bowen:2008:FM

[Bow08d] Rufus Bowen. Front matter. Lecture Notes in Mathematics, 470: i–x, 2008. CODEN LNMAA2. ISBN 3-540-77605-2 (print), 3-540-77695-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-77695-6/1.pdf.

Bowen:2008:GTF

[Bow08e] Rufus Bowen. General thermodynamic formalism. Lecture Notes in Mathematics, 470:29–44, 2008. CODEN LNMAA2. ISBN 3-540-77605-2 (print), 3-540-77695-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-77695-6_2.pdf.

Bowen:2008:GM

[Bow08f] Rufus Bowen. Gibbs measures. Lecture Notes in Mathematics, 470:3–27, 2008. CODEN LNMAA2. ISBN 3-540-77605-2 (print), 3-540-77695-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77695-6_1.pdf.

Bass:2001:MPS

[BP01] Richard F. Bass and Edwin A. Perkins. On the martingale problem for super-Brownian motion. Lecture Notes in Mathematics, 1755:195–201, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44671-2_14/.

Burstall:2002:Q

[BPF^{+}02a] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 1. Quaternions. Lecture Notes in Mathematics, 1772:1–4, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45301-7_1/.

Burstall:2002:WSb

[BPF^{+}02b] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 10. Willmore surfaces in S^{3}. Lecture Notes in Mathematics, 1772:61–66, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45301-7_10/.

Burstall:2002:SWS

[BPF^{+}02c] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 11. Spherical Willmore surfaces in HP^{1}. Lecture Notes in Mathematics, 1772:67–72, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45301-7_11/.

Burstall:2002:DT

[BPF^{+}02d] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 12. Darboux transforms. Lecture Notes in Mathematics, 1772:73–81, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45301-7_12/.

Burstall:2002:A

[BPF^{+}02e] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 13. Appendix. Lecture Notes in Mathematics, 1772:83–86, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45301-7_13/.

Burstall:2002:LAQ

[BPF^{+}02f] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 2. Linear algebra over the quaternions. Lecture Notes in Mathematics, 1772:5–8, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45301-7_2/.

Burstall:2002:PS

[BPF^{+}02g] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 3. Projective spaces. Lecture Notes in Mathematics, 1772:9–14, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45301-7_3/.

Burstall:2002:VB

[BPF^{+}02h] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 4. Vector bundles. Lecture Notes in Mathematics, 1772:15–22, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45301-7_4/.

Burstall:2002:MCS

[BPF^{+}02i] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 5. The mean curvature sphere. Lecture Notes in Mathematics, 1772:23–30, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45301-7_5/.

Burstall:2002:WSa

[BPF^{+}02j] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 6. Willmore surfaces. Lecture Notes in Mathematics, 1772:31–38, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45301-7_6/.

Burstall:2002:MAC

[BPF^{+}02k] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 7. Metric and affine conformal geometry. Lecture Notes in Mathematics, 1772:39–46, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45301-7_7/.

Burstall:2002:TP

[BPF^{+}02l] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 8. Twistor projections. Lecture Notes in Mathematics, 1772:47–52, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45301-7_8/.

Burstall:2002:BTW

[BPF^{+}02m] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. 9. Bäcklund transforms of Willmore surfaces. Lecture Notes in Mathematics, 1772:53–59, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45301-7_9/.

Burstall:2002:I

[BPF^{+}02n] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. Index. Lecture Notes in Mathematics, 1772: 89, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45301-7_15.

Burstall:2002:R

[BPF^{+}02o] Francis E. Burstall, Franz Pedit, Dirk Ferus, Katrin Leschke, and Ulrich Pinkall. References. Lecture Notes in Mathematics, 1772: 87, 2002. CODEN LNMAA2. ISBN 3-540-43008-3 (print), 3-540-45301-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45301-7_14.

Buttazzo:2009:FM

[BPSS09a] Giuseppe Buttazzo, Aldo Pratelli, Eugene Stepanov, and Sergio Solimini. Front matter. Lecture Notes in Mathematics, 1961: 1–8, 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-85799-0/1.pdf.

Buttazzo:2009:OUN

[BPSS09b] Giuseppe Buttazzo, Aldo Pratelli, Eugene Stepanov, and Sergio Solimini. Optimal Urban Networks via Mass Transportation, volume 1961 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 150 pp. LCCN QA3 .L28 no. 1961. URL http://link.

springer.com/book/10.1007/978-3-540-85799-0;   http:/
/www.springerlink.com/content/978-3-540-85799-0.

Barrat:2008:FGG

[BPT^{+}08] A. Barrat, A. Puglisi, E. Trizac, P. Visco, and F. van Wijland. Fluctuations in granular gases. Lecture Notes in Mathematics, 1937:125–165, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78277-3_7.pdf.

Bramble:2003:MMF

[Bra03] James H. Bramble. Multilevel methods in finite elements. Lecture Notes in Mathematics, 1825:97–151, 2003. CODEN LNMAA2. ISBN 3-540-20099-1 (print), 3-540-39810-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39810-3_3/.

Bramson:2008:ASF

[Bra08a] Maury Bramson. Applications and some further theory. Lecture Notes in Mathematics, 1950:139–173, 2008. CODEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-68896-9_5.
pdf.

Bramson:2008:BM

[Bra08b] Maury Bramson. Back matter. Lecture Notes in Mathematics, 1950:175–190, 2008. CODEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-68896-9/1.pdf.

Bramson:2008:CN

[Bra08c] Maury Bramson. The classical networks. Lecture Notes in Mathematics, 1950:17–52, 2008. CODEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-68896-9_2.pdf.

Bramson:2008:FM

[Bra08d] Maury Bramson. Front matter. Lecture Notes in Mathematics, 1950:i–viii, 2008. CODEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-68896-9/1.pdf.

Bramson:2008:ISQ

[Bra08e] Maury Bramson. Instability of subcritical queueing networks. Lecture Notes in Mathematics, 1950:53–76, 2008. CO-DEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68896-9_3.pdf.

Bramson:2008:I

[Bra08f] Maury Bramson. Introduction. Lecture Notes in Mathematics, 1950:1–16, 2008. CODEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-68896-9_1.pdf.

Bramson:2008:SQNa

[Bra08g] Maury Bramson. Stability of queueing networks. Lecture Notes in Mathematics, 1950:77–138, 2008. CODEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-68896-9_4.pdf.

Bramson:2008:SQNb

[Bra08h] Maury Bramson, editor. Stability of Queueing Networks: École d’Été de Probabilités de Saint-Flour XXXVI — 2006, volume 1950 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-68895-1 (print), 3-540-68896-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1950. URL http://link.springer.com/book/
10.1007/978-3-540-68896-9;   http://www.springerlink.
com/content/978-3-540-68896-9.

Brauer:2008:CME

[Bra08i] Fred Brauer. Compartmental models in epidemiology. Lecture Notes in Mathematics, 1945:19–79, 2008. CODEN LNMAA2.

ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78911-6_2.
pdf.

Brauer:2008:INE

[Bra08j] Fred Brauer. An introduction to networks in epidemic modeling. Lecture Notes in Mathematics, 1945:133–146, 2008. CO-DEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_4.pdf.

Brauer:2008:MIP

[Bra08k] Fred Brauer. Modeling influenza: Pandemics and seasonal epidemics. Lecture Notes in Mathematics, 1945:321–347, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_12.pdf.

Brenier:2003:EMK

[Bre03] Yann Brenier. Extended Monge–Kantorovich theory. Lecture Notes in Mathematics, 1813:91–121, 2003. CODEN LNMAA2. ISBN 3-540-40192-X (print), 3-540-44857-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44857-0_4/.

Bressan:2007:BSH

[Bre07] Alberto Bressan. BV solutions to hyperbolic systems by vanishing viscosity. Lecture Notes in Mathematics, 1911:1–77, 2007. CODEN LNMAA2. ISBN 3-540-72186-X (print), 3-540-72187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72187-1_1.pdf.

Brossard:2003:DNE

[Bro03] Jean Brossard. Deux notions équivalentes d’unicité en loi pour les équations différentielles stochastiques. (French) []. Lecture Notes in Mathematics, 1832:246–250, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40004-2_10/.

Brown:2004:Ia

[Bro04a] Martin L. Brown. 1. Introduction. Lecture Notes in Mathematics, 1849:1–11, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44475-6_1/.

Brown:2004:P

[Bro04b] Martin L. Brown. 2. Preliminaries. Lecture Notes in Mathematics, 1849:13–30, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44475-6_2/.

Brown:2004:BTT

[Bro04c] Martin L. Brown. 3. Bruhat-tits trees with complex multiplication. Lecture Notes in Mathematics, 1849:31–74, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44475-6_
3/.

Brown:2004:HS

[Bro04d] Martin L. Brown. 4. Heegner sheaves. Lecture Notes in Mathematics, 1849:75–103, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44475-6_4/.

Brown:2004:HM

[Bro04e] Martin L. Brown. 5. The Heegner module. Lecture Notes in Mathematics, 1849:105–222, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44475-6_5/.

Brown:2004:CHM

[Bro04f] Martin L. Brown. 6. Cohomology of the Heegner module. Lecture Notes in Mathematics, 1849:223–327, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44475-6_6/.

Brown:2004:FTS

[Bro04g] Martin L. Brown. 7. Finiteness of Tate–Shafarevich groups. Lecture Notes in Mathematics, 1849:329–434, 2004. CODEN LN-MAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44475-6_
7/.

Brown:2004:A

[Bro04h] Martin L. Brown. Appendix. Lecture Notes in Mathematics, 1849:435–505, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44475-6_8/.

Brown:2004:HME

[Bro04i] Martin L. Brown. Heegner Modules and Elliptic Curves, volume 1849 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 517 pp. LCCN QA3 .L28 no. 1849. URL http:/

/link.springer.com/book/10.1007/b98488;   http://www.
springerlink.com/content/978-3-540-44475-6.

Brown:2004:Ib

[Bro04j] Martin L. Brown. Index. Lecture Notes in Mathematics, 1849: 511–517, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44475-6_10/.

Brown:2004:R

[Bro04k] Martin L. Brown. References. Lecture Notes in Mathematics, 1849:507–510, 2004. CODEN LNMAA2. ISBN 3-540-22290-1 (print), 3-540-44475-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44475-6_9/.

Bruinier:2002:BPC

[Bru02a] Jan H. Bruinier. Borcherds Products on O(2, l) and Chern Classes of Heegner Divisors, volume 1780 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 152 pp. LCCN QA3 .L28 no. 1780. URL

http://link.springer.com/book/10.1007/b83278;   http:
//www.springerlink.com/content/978-3-540-45872-2.

Bruinier:2002:VVM

[Bru02b] Jan Hendrik Bruinier. 1. Vector valued modular forms for the metaplectic group. Lecture Notes in Mathematics, 1780:15– 38, 2002. CODEN LNMAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45872-2_2/.

Bruinier:2002:RTL

[Bru02c] Jan Hendrik Bruinier. 2. The regularized theta lift. Lecture Notes in Mathematics, 1780:39–61, 2002. CODEN LNMAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45872-2_3/.

Bruinier:2002:FET

[Bru02d] Jan Hendrik Bruinier. 3. The Fourier expansion of the theta lift. Lecture Notes in Mathematics, 1780:63–94, 2002. CODEN LNMAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45872-2_
4/.

Bruinier:2002:SRG

[Bru02e] Jan Hendrik Bruinier. 4. Some Riemann geometry on O(2, l). Lecture Notes in Mathematics, 1780:95–118, 2002. CODEN LN-MAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45872-2_
5/.

Bruinier:2002:CCH

[Bru02f] Jan Hendrik Bruinier. 5. Chern classes of Heegner divisors. Lecture Notes in Mathematics, 1780:119–140, 2002. CODEN LN-MAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45872-2_

Bruinier:2002:I

[Bru02g] Jan Hendrik Bruinier. Introduction. Lecture Notes in Mathematics, 1780:1–13, 2002. CODEN LNMAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45872-2_1/.

Bruinier:2002:R

[Bru02h] Jan Hendrik Bruinier. References. Lecture Notes in Mathematics, 1780:141–144, 2002. CODEN LNMAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45872-2_7/.

Bruinier:2002:SIN

[Bru02i] Jan Hendrik Bruinier. Subject index and notation index. Lecture Notes in Mathematics, 1780:145–152, 2002. CODEN LNMAA2. ISBN 3-540-43320-1 (print), 3-540-45872-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45872-2_8/.

Beghdadi-Sakrani:2003:CSP

[BS03a] Samia Beghdadi-Sakrani. Calcul stochastique pour des mesures signées. (French) []. Lecture Notes in Mathematics, 1801:366– 382, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_17/.

Beghdadi-Sakrani:2003:MNP

[BS03b] Samia Beghdadi-Sakrani. Une martingale non pure, dont la filtration est brownienne. (French) []. Lecture Notes in Mathematics, 1801:348–359, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36107-7_15/.

Bozejko:2003:ALP

[BS03c] Marek Bożejko and Ryszard Szwarc. Algebraic length and Poincaré series on reflection groups with applications to representations theory. Lecture Notes in Mathematics, 1815:201– 221, 2003. CODEN LNMAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44890-X_9/.

Benjamini:2004:PEV

[BS04] I. Benjamini and O. Schramm. Pinched exponential volume growth implies an infinite dimensional isoperimetric inequality. Lecture Notes in Mathematics, 1850:73–76, 2004. CODEN LN-MAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44489-3_
8/.

Buttazzo:2009:BM

[BSPS09a] Giuseppe Buttazzo, Sergio Solimini, Aldo Pratelli, and Eugene Stepanov. Back matter. Lecture Notes in Mathematics, 1961: 1–26, 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-85799-0/1.pdf.

Buttazzo:2009:I

[BSPS09b] Giuseppe Buttazzo, Sergio Solimini, Aldo Pratelli, and Eugene Stepanov. Introduction. Lecture Notes in Mathematics, 1961:1– 6, 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85799-0_1.pdf.

Buttazzo:2009:OCN

[BSPS09c] Giuseppe Buttazzo, Sergio Solimini, Aldo Pratelli, and Eugene Stepanov. Optimal connected networks. Lecture Notes in Mathematics, 1961:1–12, 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85799-0_3.pdf.

Buttazzo:2009:OSG

[BSPS09d] Giuseppe Buttazzo, Sergio Solimini, Aldo Pratelli, and Eugene Stepanov. Optimal sets and geodesics in the two-dimensional case. Lecture Notes in Mathematics, 1961:1–25, 2009. CO-DEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85799-0_6.pdf.

Buttazzo:2009:PS

[BSPS09e] Giuseppe Buttazzo, Sergio Solimini, Aldo Pratelli, and Eugene Stepanov. Problem setting. Lecture Notes in Mathematics, 1961: 1–17, 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-85799-0_2.pdf.

Buttazzo:2009:RPE

[BSPS09f] Giuseppe Buttazzo, Sergio Solimini, Aldo Pratelli, and Eugene Stepanov. Relaxed problem and existence of solutions. Lecture Notes in Mathematics, 1961:1–37, 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-85799-0_4.
pdf.

Buttazzo:2009:TPO

[BSPS09g] Giuseppe Buttazzo, Sergio Solimini, Aldo Pratelli, and Eugene Stepanov. Topological properties of optimal sets. Lecture Notes in Mathematics, 1961:1–29, 2009. CODEN LNMAA2. ISBN 3-540-85798-2 (print), 3-540-85799-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-85799-0_5.pdf.

Brasselet:2009:BM

[BSS09a] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. Back matter. Lecture Notes in Mathematics, 1987:215–231, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-642-05205-7/
1.pdf.

Brasselet:2009:CHV

[BSS09b] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The case of holomorphic vector fields. Lecture Notes in Mathematics, 1987:97–113, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-05205-7_6.pdf.

Brasselet:2009:CM

[BSS09c] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The case of manifolds. Lecture Notes in Mathematics, 1987:1–29, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_1.pdf.

Brasselet:2009:CCC

[BSS09d] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. Characteristic classes of coherent sheaves on singular varieties. Lecture Notes in Mathematics, 1987:201–213, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-05205-7_13.pdf.

Brasselet:2009:FM

[BSS09e] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. Front matter. Lecture Notes in Mathematics, 1987:i–xx, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-05205-7/
1.pdf.

Brasselet:2009:GI

[BSS09f] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The GSV index. Lecture Notes in Mathematics, 1987:43–69, 2009. CO-DEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_3.pdf.

Brasselet:2009:HIA

[BSS09g] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The homological index and algebraic formulas. Lecture Notes in Mathematics, 1987:115–128, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-05205-7_7.pdf.

Brasselet:2009:IF

[BSS09h] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. Indices for 1-forms. Lecture Notes in Mathematics, 1987:143–166, 2009.

CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_9.pdf.

Brasselet:2009:IVF

[BSS09i] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. Indices of vector fields on real analytic varieties. Lecture Notes in Mathematics, 1987:71–83, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-05205-7_4.pdf.

Brasselet:2009:LEO

[BSS09j] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The local Euler obstruction. Lecture Notes in Mathematics, 1987:129–141, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_8.pdf.

Brasselet:2009:MNM

[BSS09k] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. Milnor number and Milnor classes. Lecture Notes in Mathematics, 1987:193– 200, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_12.pdf.

Brasselet:2009:SC

[BSS09l] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The Schwartz classes. Lecture Notes in Mathematics, 1987:167–184, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_10.pdf.

Brasselet:2009:SI

[BSS09m] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The Schwartz index. Lecture Notes in Mathematics, 1987:31–41, 2009. CO-DEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_2.pdf.

Brasselet:2009:VFS

[BSS09n] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. Vector fields on Singular Varieties, volume 1987 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 225 pp. LCCN QA3 .L28 no.

1987. URL http://link.springer.com/book/10.1007/978-
3-642-05205-7;   http://www.springerlink.com/content/
978-3-642-05205-7.

Brasselet:2009:VC

[BSS09o] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The virtual classes. Lecture Notes in Mathematics, 1987:185–192, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_11.pdf.

Brasselet:2009:VI

[BSS09p] Jean-Paul Brasselet, José Seade, and Tatsuo Suwa. The virtual index. Lecture Notes in Mathematics, 1987:85–96, 2009. CODEN LNMAA2. ISBN 3-642-05204-5 (print), 3-642-05205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05205-7_5.pdf.

Bressan:2007:FM

[BSWZ07] Alberto Bressan, Denis Serre, Mark Williams, and Kevin Zumbrun. Front matter. Lecture Notes in Mathematics, 1911:I– XII, 2007. CODEN LNMAA2. ISBN 3-540-72186-X (print), 3-540-72187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-72187-1/1.pdf.

Buhler:2005:IEF

[Büh05] Hans Bühler. Information-equivalence: On filtrations created by independent increments. Lecture Notes in Mathematics, 1857: 195–204, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-31449-3_14.pdf.

Buiculescu:2005:EDP

[Bui05] Mioara Buiculescu. Exponential decay parameters associated with excessive measures. Lecture Notes in Mathematics, 1857: 135–144, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-31449-3_9.pdf.

Bunger:2002:SIT

[Bün02] F. Bünger. Standard integral table algebras with a faithful real element of degree 5 and width 3. Lecture Notes in Mathematics, 1773:83–103, 2002. CODEN LNMAA2. ISBN 3-540-42851-8 (print), 3-540-45558-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45558-2_4/.

Burkard:2000:TPG

[Bur00] Rainer E. Burkard. Trees and paths: graph optimisation problems with industrial applications. Lecture Notes in Mathematics, 1739:1–38, 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103916/.

Barreira:2008:BM

[BV08a] Luis Barreira and Claudia Valls. Back matter. Lecture Notes in Mathematics, 1926:277–290, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-74775-8/1.pdf.

Barreira:2008:CMB

[BV08b] Luis Barreira and Claudia Valls. Center manifolds in Banach spaces. Lecture Notes in Mathematics, 1926:171–196, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_8.pdf.

Barreira:2008:EDB

[BV08c] Luis Barreira and Claudia Valls. Exponential dichotomies and basic properties. Lecture Notes in Mathematics, 1926:19–25, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_2.pdf.

Barreira:2008:FMa

[BV08d] Luis Barreira and Claudia Valls. Front matter. Lecture Notes in Mathematics, 1926:I–XIV, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-74775-8/1.pdf.

Barreira:2008:FMb

[BV08e] Luis Barreira and Claudia Valls. Front matter. Lecture Notes in Mathematics, 1926:17, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-74775-8/1/1.pdf.

Barreira:2008:FMc

[BV08f] Luis Barreira and Claudia Valls. Front matter. Lecture Notes in Mathematics, 1926:53, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-74775-8/2/1.pdf.

Barreira:2008:FMd

[BV08g] Luis Barreira and Claudia Valls. Front matter. Lecture Notes in Mathematics, 1926:169, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-74775-8/3/1.pdf.

Barreira:2008:FMe

[BV08h] Luis Barreira and Claudia Valls. Front matter. Lecture Notes in Mathematics, 1926:217, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-74775-8/4/1.pdf.

Barreira:2008:I

[BV08i] Luis Barreira and Claudia Valls. Introduction. Lecture Notes in Mathematics, 1926:1–16, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74775-8_1.pdf.

Barreira:2008:LSM

[BV08j] Luis Barreira and Claudia Valls. Lipschitz stable manifolds. Lecture Notes in Mathematics, 1926:55–73, 2008. CO-DEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_4.pdf.

Barreira:2008:LRE

[BV08k] Luis Barreira and Claudia Valls. Lyapunov regularity and exponential dichotomies. Lecture Notes in Mathematics, 1926:219–248, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_10.pdf.

Barreira:2008:LRH

[BV08l] Luis Barreira and Claudia Valls. Lyapunov regularity in Hilbert spaces. Lecture Notes in Mathematics, 1926:249–263, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_11.pdf.

Barreira:2008:NGH

[BV08m] Luis Barreira and Claudia Valls. A nonautonomous Grobman– Hartman theorem. Lecture Notes in Mathematics, 1926:145–167, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_7.pdf.

Barreira:2008:REC

[BV08n] Luis Barreira and Claudia Valls. Reversibility and equivariance in center manifolds. Lecture Notes in Mathematics, 1926:197–215, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_9.pdf.

Barreira:2008:RNE

[BV08o] Luis Barreira and Claudia Valls. Robustness of nonuniform exponential dichotomies. Lecture Notes in Mathematics, 1926:27–51, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_3.pdf.

Barreira:2008:SSMb

[BV08p] Luis Barreira and Claudia Valls. Smooth stable manifolds in Banach spaces. Lecture Notes in Mathematics, 1926:119–143, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_6.pdf.

Barreira:2008:SSMa

[BV08q] Luis Barreira and Claudia Valls. Smooth stable manifolds in rn. Lecture Notes in Mathematics, 1926:75–117, 2008. CO-DEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74775-8_5.pdf.

Barreira:2008:SND

[BV08r] Luis Barreira and Claudia Valls. Stability of Nonautonomous Differential Equations, volume 1926 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 285 pp. LCCN QA3 .L28 no.

1926. URL http://link.springer.com/book/10.1007/978-
3-540-74775-8;   http://www.springerlink.com/content/
978-3-540-74775-8.

Barreira:2008:SNE

[BV08s] Luis Barreira and Claudia Valls. Stability of nonautonomous equations in Hilbert spaces. Lecture Notes in Mathematics, 1926:265–276, 2008. CODEN LNMAA2. ISBN 3-540-74774-5 (print), 3-540-74775-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74775-8_12.pdf.

Brauer:2008:ME

[BvdDW08] Fred Brauer, Pauline van den Driessche, and Jianhong Wu. Mathematical Epidemiology, volume 1945 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 408 pp. LCCN QA3 .L28 no.

1945. URL http://link.springer.com/book/10.1007/978-
3-540-78911-6;   http://www.springerlink.com/content/
978-3-540-78911-6.

Bolthausen:2002:BMa

[BvdVP02a] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Back matter. Lecture Notes in Mathematics, 1781:121–125, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47944-4/1/1.pdf.

Bolthausen:2002:BMb

[BvdVP02b] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Back matter. Lecture Notes in Mathematics, 1781:318–329, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47944-4/2/1.pdf.

Bolthausen:2002:BMc

[BvdVP02c] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Back matter. Lecture Notes in Mathematics, 1781:455–457, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-47944-4/3/1.pdf.

Bolthausen:2002:BPS

[BvdVP02d] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Branching particle systems and Dawson–Watanabe superprocesses. Lecture Notes in Mathematics, 1781:135–192, 2002. CO-DEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/3-540-47944-
9_6/.

Bolthausen:2002:FMa

[BvdVP02e] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Front matter. Lecture Notes in Mathematics, 1781:4, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-47944-4/
1/1.pdf.

Bolthausen:2002:FMb

[BvdVP02f] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Front matter. Lecture Notes in Mathematics, 1781:127, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-47944-4/
2/1.pdf.

Bolthausen:2002:FMc

[BvdVP02g] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Front matter. Lecture Notes in Mathematics, 1781:332, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-47944-4/
3/1.pdf.

Bolthausen:2002:ID

[BvdVP02h] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Interactive drifts. Lecture Notes in Mathematics, 1781:247–280, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-47944-9_8/.

Bolthausen:2002:Ia

[BvdVP02i] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Introduction. Lecture Notes in Mathematics, 1781:5–6, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-47944-9_1.

Bolthausen:2002:Ib

[BvdVP02j] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Introduction. Lecture Notes in Mathematics, 1781:132–134, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-47944-9_5/.

Bolthausen:2002:LCS

[BvdVP02k] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Calculus of scores. Lecture Notes in Mathematics, 1781: 357–369, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-47944-9_12/.

Bolthausen:2002:LES

[BvdVP02l] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Efficient score and one-step estimators. Lecture Notes in Mathematics, 1781:412–423, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-47944-9_16/.

Bolthausen:2002:LEPa

[BvdVP02m] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Empirical processes and consistency of Z-estimators. Lecture Notes in Mathematics, 1781:383–394, 2002. CODEN LN-MAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-47944-9_14/.

Bolthausen:2002:LEPb

[BvdVP02n] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Empirical processes and normality of Z-estimators. Lecture Notes in Mathematics, 1781:395–411, 2002. CODEN LN-MAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-47944-9_15/.

Bolthausen:2002:LGA

[BvdVP02o] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Gaussian approximations. Lecture Notes in Mathematics, 1781:370–382, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-47944-9_13/.

Bolthausen:2002:LID

[BvdVP02p] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Infinite-dimensional Z-estimators. Lecture Notes in Mathematics, 1781:446–454, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-47944-9_19/.

Bolthausen:2002:LIT

[BvdVP02q] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Introduction, tangent sets. Lecture Notes in Mathematics, 1781:336–345, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-47944-9_10/.

Bolthausen:2002:LLB

[BvdVP02r] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Lower bounds. Lecture Notes in Mathematics, 1781:346– 356, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-47944-9_11/.

Bolthausen:2002:LMP

[BvdVP02s] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Maximum and profile likelihood. Lecture Notes in Mathematics, 1781:433–445, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-47944-9_18/.

Bolthausen:2002:LRC

[BvdVP02t] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Lecture: Rates of convergence. Lecture Notes in Mathematics, 1781:

424–432, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-47944-9_17/.

Bolthausen:2002:CTD

[BvdVP02u] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. On the construction of the three-dimensional polymer measure. Lecture Notes in Mathematics, 1781:7–38, 2002. CODEN LN-MAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-47944-9_2/.

Bolthausen:2002:ODP

[BvdVP02v] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. One-dimensional pinning-depinning transitions. Lecture Notes in Mathematics, 1781:105–120, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-47944-9_4/.

Bolthausen:2002:SPP

[BvdVP02w] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Sample path properties of superprocesses. Lecture Notes in Mathematics, 1781:193–246, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-47944-9_7/.

Bolthausen:2002:SAR

[BvdVP02x] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Selfattracting random walks. Lecture Notes in Mathematics, 1781: 39–104, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-47944-9_3/.

Bolthausen:2002:SI

[BvdVP02y] Erwin Bolthausen, Aad van der Vaart, and Edwin Perkins. Spatial interactions. Lecture Notes in Mathematics, 1781:281– 317, 2002. CODEN LNMAA2. ISBN 3-540-43736-3 (print), 3-540-47944-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-47944-9_9/.

Caffarelli:2003:MAE

[Caf03] Luis A. Caffarelli. The Monge–Ampère equation and optimal transportation, an elementary review. Lecture Notes in Mathematics, 1813:1–10, 2003. CODEN LNMAA2. ISBN 3-540-40192-X (print), 3-540-44857-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44857-0_1/.

CannasdaSilva:2001:SF

[Can01] Ana Cannas da Silva. Symplectic forms. Lecture Notes in Mathematics, 1764:3–8, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_1/.

Canuto:2003:MPM

[Can03] Claudio Canuto, editor. Multiscale Problems and Methods in Numerical Simulations: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy, September 9-15, 2001, volume 1825 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20099-1 (print), 3-540-39810-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1825. URL http:/

/link.springer.com/book/10.1007/b13466;   http://www.
springerlink.com/content/978-3-540-39810-3.

Cao:2003:CEI

[Cao03a] Frédéric Cao. 1. Curve evolution and image processing. Lecture Notes in Mathematics, 1805:3–21, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36392-7_1/.

Cao:2003:RBC

[Cao03b] Frédéric Cao. 2. Rudimentary bases of curve geometry. Lecture Notes in Mathematics, 1805:23–28, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36392-7_2/.

Cao:2003:GCS

[Cao03c] Frédéric Cao. 3. Geometric curve shortening flow. Lecture Notes in Mathematics, 1805:31–53, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36392-7_3/.

Cao:2003:CEL

[Cao03d] Frédéric Cao. 4. Curve evolution and level sets. Lecture Notes in Mathematics, 1805:55–103, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36392-7_4/.

Cao:2003:CNM

[Cao03e] Frédéric Cao. 5. Classical numerical methods for curve evolution. Lecture Notes in Mathematics, 1805:107–110, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36392-7_
5/.

Cao:2003:GSC

[Cao03f] Frédéric Cao. 6. A geometrical scheme for curve evolution. Lecture Notes in Mathematics, 1805:111–166, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36392-7_6/.

Cao:2003:PT

[Cao03g] Frédéric Cao. A. proof of Thm. 4.34. Lecture Notes in Mathematics, 1805:171–176, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36392-7_8/.

Cao:2003:CP

[Cao03h] Frédéric Cao. Conclusion and perspectives. Lecture Notes in Mathematics, 1805:167–169, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36392-7_7/.

Cao:2003:GCE

[Cao03i] Frédéric Cao. Geometric Curve Evolution and Image Processing, volume 1805 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 187 pp. LCCN QA3 .L28 no. 1805. URL http:/

/link.springer.com/book/10.1007/b10404;   http://www.
springerlink.com/content/978-3-540-36392-7.

Cao:2003:I

[Cao03j] Frédéric Cao. Index. Lecture Notes in Mathematics, 1805:185– 187, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36392-7_10/.

Cao:2003:R

[Cao03k] Frédéric Cao. References. Lecture Notes in Mathematics, 1805: 177–184, 2003. CODEN LNMAA2. ISBN 3-540-00402-5 (print), 3-540-36392-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36392-7_9/.

Capasso:2000:MMP

[Cap00a] Vincenzo Capasso. Mathematical models for polymer crystallization processes. Lecture Notes in Mathematics, 1739:39–67, 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103917/.

Capitaine:2000:OMF

[Cap00b] Mireille Capitaine. On the Onsager–Machlup functional for elliptic diffusion processes. Lecture Notes in Mathematics, 1729: 313–328, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103810/.

Capriz:2008:BKS

[Cap08] Gianfranco Capriz. Bodies with kinetic substructure. Lecture Notes in Mathematics, 1937:77–90, 2008. CODEN LNMAA2.

ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78277-3_4.
pdf.

Carmona:2007:HUA

[Car07] René A. Carmona. HJM: a unified approach to dynamic models for fixed income, credit and equity markets. Lecture Notes in Mathematics, 1919:1–50, 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-73327-0_1.pdf.

Catoni:2004:ULD

[Cat04a] Olivier Catoni. 1. Universal lossless data compression. Lecture Notes in Mathematics, 1851:5–54, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44507-4_2/.

Catoni:2004:LBD

[Cat04b] Olivier Catoni. 2. Links between data compression and statistical estimation. Lecture Notes in Mathematics, 1851:55– 69, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44507-4_3/.

Catoni:2004:NCM

[Cat04c] Olivier Catoni. 3. Non cumulated mean risk. Lecture Notes in Mathematics, 1851:71–95, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44507-4_4/.

Catoni:2004:GE

[Cat04d] Olivier Catoni. 4. Gibbs estimators. Lecture Notes in Mathematics, 1851:97–154, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44507-4_5/.

Catoni:2004:REE

[Cat04e] Olivier Catoni. 5. Randomized estimators and empirical complexity. Lecture Notes in Mathematics, 1851:155–197, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44507-4_
6/.

Catoni:2004:DI

[Cat04f] Olivier Catoni. 6. Deviation inequalities. Lecture Notes in Mathematics, 1851:199–222, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44507-4_7/.

Catoni:2004:MCE

[Cat04g] Olivier Catoni. 7. Markov chains with exponential transitions. Lecture Notes in Mathematics, 1851:223–260, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44507-4_
8/.

Catoni:2004:Ib

[Cat04h] Olivier Catoni. Index. Lecture Notes in Mathematics, 1851:267– 269, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44507-4_10/.

Catoni:2004:Ia

[Cat04i] Olivier Catoni. Introduction. Lecture Notes in Mathematics, 1851: 1–4, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44507-4_1/.

Catoni:2004:LPL

[Cat04j] Olivier Catoni. List of participants and list of short lectures. Lecture Notes in Mathematics, 1851:271–273, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44507-4_
11/.

Catoni:2004:R

[Cat04k] Olivier Catoni. References. Lecture Notes in Mathematics, 1851: 261–265, 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44507-4_9/.

Catanese:2008:DDT

[Cat08] Fabrizio Catanese. Differentiable and deformation type of algebraic surfaces, real and symplectic structures. Lecture Notes in Mathematics, 1938:55–167, 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78279-7_2.pdf.

Chao:2001:SRM

[CC01] Tsung-Ming Chao and Ching-Shung Chou. Some remarks on the martingales satisfying the structure equation [X, X]_{t} = t + ∫_{0}^{t} βX_{s^{−}}dX_{s}. Lecture Notes in Mathematics, 1755:87–97, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44671-2_4/.

Capitaine:2008:GIC

[CC08] M. Capitaine and M. Casalis. Geometric interpretation of the cumulants for random matrices previously defined as convolutions on the symmetric group. Lecture Notes in Mathematics, 1934:93– 119, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_4.pdf.

Carmona:2004:PPL

[CÇE^{+}04] René A. Carmona, Erhan Çinlar, Ivar Ekeland, Elyes Jouini, José A. Scheinkman, and Nizar Touzi, editors. Paris–Princeton Lectures on Mathematical Finance 2003, volume 1847 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2.

ISBN 3-540-22266-9 (print), 3-540-44468-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1847.

URL http://link.springer.com/book/10.1007/b98353;
http://www.springerlink.com/content/978-3-540-44468-8.

Connes:2004:FM

[CCG^{+}04] Alain Connes, Joachim Cuntz, Erik Guentner, Nigel Higson, Jerome Kaminker, and John E. Roberts. Front matter. Lecture Notes in Mathematics, 1831:I–XIV, 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-39702-1/1.pdf.

Chow:2003:DSL

[CCJ^{+}03] S.-N. Chow, Roberto Conti, R. Johnson, J. Mallet-Paret, and R. Nussbaum, editors. Dynamical Systems: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, June 19–26, 2000, volume 1822 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40786-3 (print), 3-540-45204-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1822. URL http:/

/link.springer.com/book/10.1007/b13403;   http://www.
springerlink.com/content/978-3-540-45204-1.

Carpio:2008:FM

[CDM^{+}08] Ana Carpio, Oliver Dorn, Miguel Moscoso, Frank Natterer, George C. Papanicolaou, Maria Luisa Rapún, and Alessandro Teta. Front matter. Lecture Notes in Mathematics, 1943:i– xi, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-78547-7/1.pdf.

Cherny:2005:SDE

[CE05a] Alexander S. Cherny and Hans-Jürgen Engelbert. 1. Stochastic differential equations. Lecture Notes in Mathematics, 1858:5–25, 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31560-5_2.pdf.

Cherny:2005:OSC

[CE05b] Alexander S. Cherny and Hans-Jürgen Engelbert. 2. One-sided classification of isolated singular points. Lecture Notes in Mathematics, 1858:27–64, 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31560-5_3.pdf.

Cherny:2005:TSC

[CE05c] Alexander S. Cherny and Hans-Jürgen Engelbert. 3. Two-sided classification of isolated singular points. Lecture Notes in Mathematics, 1858:65–79, 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31560-5_4.pdf.

Cherny:2005:CIG

[CE05d] Alexander S. Cherny and Hans-Jürgen Engelbert. 4. Classification at infinity and global solutions. Lecture Notes in Mathematics, 1858:81–91, 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31560-5_5.pdf.

Cherny:2005:SSC

[CE05e] Alexander S. Cherny and Hans-Jürgen Engelbert. 5. Several special cases. Lecture Notes in Mathematics, 1858:93–103, 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31560-5_6.pdf.

Cherny:2005:A

[CE05f] Alexander S. Cherny and Hans-Jürgen Engelbert. Appendix. Lecture Notes in Mathematics, 1858:105–118, 2005. CO-DEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31560-5_7.pdf.

Cherny:2005:INI

[CE05g] Alexander S. Cherny and Hans-Jürgen Engelbert. Index of notation and index of terms. Lecture Notes in Mathematics, 1858:123– 128, 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31560-5_9.pdf.

Cherny:2005:I

[CE05h] Alexander S. Cherny and Hans-Jürgen Engelbert. Introduction. Lecture Notes in Mathematics, 1858:1–4, 2005. CO-DEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31560-5_1.pdf.

Cherny:2005:R

[CE05i] Alexander S. Cherny and Hans-Jürgen Engelbert. References. Lecture Notes in Mathematics, 1858:119–121, 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31560-5_8.pdf.

Cherny:2005:SSD

[CE05j] Alexander S. Cherny and Hans-Jürgen Engelbert. Singular Stochastic Differential Equations, volume 1858 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-24007-1 (print), 3-540-31560-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 128 pp. LCCN QA3 .L28

no. 1858. URL http://link.springer.com/book/10.1007/
b104187;   http://www.springerlink.com/content/978-3-
540-31560-5.

Carmona:2007:FM

[CEKH^{+}07a] René A. Carmona, Ivar Ekeland, Arturo Kohatsu-Higa, Jean-Michel Lasry, Pierre-Louis Lions, Huyên Pham, and Erik Taflin. Front matter. Lecture Notes in Mathematics, 1919:i– viii, 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-73327-0/1.pdf.

Carmona:2007:PPL

[CEKH^{+}07b] René A. Carmona, Ivar Ekeland, Arturo Kohatsu-Higa, Jean-Michel Lasry, Pierre-Louis Lions, Huyên Pham, and Erik Taflin, editors. Paris–Princeton Lectures on Mathematical Finance 2004, volume 1919 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1919. URL http://link.

springer.com/book/10.1007/978-3-540-73327-0;   http:/
/www.springerlink.com/content/978-3-540-73327-0.

Cellina:2000:I

[Cel00] Arrigo Cellina. Introduction. Lecture Notes in Mathematics, 1740: 1–5, 2000. CODEN LNMAA2. ISBN 3-540-67971-5 (print), 3-540-44486-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106740/.

Cerrai:2001:ASD

[Cer01a] Sandra Cerrai. Analyticity of the semigroup in a degenerate case. Lecture Notes in Mathematics, 1762:81–101, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45147-1_4/.

Cerrai:2001:ASO

[Cer01b] Sandra Cerrai. Application to stochastic optimal control problems. Lecture Notes in Mathematics, 1762:281–300, 2001. CO-DEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/3-540-45147-
1_11/.

Cerrai:2001:ABS

[Cer01c] Sandra Cerrai. Asymptotic behaviour of solutions. Lecture Notes in Mathematics, 1762:65–80, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-45147-1_3/.

Cerrai:2001:E

[Cer01d] Sandra Cerrai. Ergodicity. Lecture Notes in Mathematics, 1762: 221–235, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45147-1_9/.

Cerrai:2001:HJB

[Cer01e] Sandra Cerrai. Hamilton–Jacobi–Bellman equations in Hilbert spaces. Lecture Notes in Mathematics, 1762:237–279, 2001. CO-DEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/3-540-45147-
1_10/.

Cerrai:2001:I

[Cer01f] Sandra Cerrai. Introduction. Lecture Notes in Mathematics, 1762: 1–18, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45147-1_1/.

Cerrai:2001:KER

[Cer01g] Sandra Cerrai. Kolmogorov equations in R^{d} with unbounded coefficients. Lecture Notes in Mathematics, 1762:21–63, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/3-540-45147-
1_2/.

Cerrai:2001:KEH

[Cer01h] Sandra Cerrai. Kolmogorov equations in Hilbert spaces. Lecture Notes in Mathematics, 1762:143–170, 2001. CODEN LN-MAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45147-1_6/.

Cerrai:2001:SOP

[Cer01i] Sandra Cerrai. Second Order PDE’s in Finite and Infinite Dimension: A Probabilistic Approach, volume 1762 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LN-MAA2. ISBN 3-540-42136-X (paperback), 3-540-45147-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 330 pp. LCCN QA3 .L28 no. 1762; QA274.25 .C47 2001. URL http:/

/link.springer.com/book/10.1007/b80743;   http://www.
loc.gov/catdir/enhancements/fy0816/2001032282-d.html;
 http://www.loc.gov/catdir/enhancements/fy0816/2001032282-
t.html;   http://www.springerlink.com/content/978-3-540-
45147-1.

Cerrai:2001:SDDa

[Cer01j] Sandra Cerrai. Smooth dependence on data for the SPDE: the Lipschitz case. Lecture Notes in Mathematics, 1762:105– 141, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45147-1_5/.

Cerrai:2001:SDDb

[Cer01k] Sandra Cerrai. Smooth dependence on data for the SPDE: the non-Lipschitz case (I). Lecture Notes in Mathematics, 1762:171– 203, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45147-1_7/.

Cerrai:2001:SDDc

[Cer01l] Sandra Cerrai. Smooth dependence on data for the SPDE: the non-Lipschitz case (II). Lecture Notes in Mathematics, 1762:205– 220, 2001. CODEN LNMAA2. ISBN 3-540-42136-X (print), 3-540-45147-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45147-1_8/.

Cerf:2006:BM

[Cer06a] Raphaël Cerf. Back matter. Lecture Notes in Mathematics, 1878: 253–268, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-34806-1/1.pdf.

Cerf:2006:BP

[Cer06b] Raphaël Cerf. Bernoulli percolation. Lecture Notes in Mathematics, 1878:25–29, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_3.pdf.

Cerf:2006:CG

[Cer06c] Raphaël Cerf. Coarse graining. Lecture Notes in Mathematics, 1878:105–116, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_9.pdf.

Cerf:2006:D

[Cer06d] Raphaël Cerf. Decoupling. Lecture Notes in Mathematics, 1878:117–127, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_10.pdf.

Cerf:2006:EUB

[Cer06e] Raphaël Cerf. Enhanced upper bound. Lecture Notes in Mathematics, 1878:215–228, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_17.pdf.

Cerf:2006:FRC

[Cer06f] Raphaël Cerf. FK or random cluster model. Lecture Notes in Mathematics, 1878:31–42, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_4.pdf.

Cerf:2006:FMa

[Cer06g] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878: i–xiv, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/1.pdf.

Cerf:2006:FMb

[Cer06h] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878:1, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/1/1.pdf.

Cerf:2006:FMc

[Cer06i] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878: 13, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/2/1.pdf.

Cerf:2006:FMd

[Cer06j] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878: 43, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/3/1.pdf.

Cerf:2006:FMe

[Cer06k] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878: 65, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/4/1.pdf.

Cerf:2006:FMf

[Cer06l] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878: 103, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/5/1.pdf.

Cerf:2006:FMg

[Cer06m] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878: 157, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/6/1.pdf.

Cerf:2006:FMh

[Cer06n] Raphaël Cerf. Front matter. Lecture Notes in Mathematics, 1878: 201, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-34806-1/7/1.pdf.

Cerf:2006:IE

[Cer06o] Raphaël Cerf. Interface estimate. Lecture Notes in Mathematics, 1878:147–155, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_12.pdf.

Cerf:2006:IM

[Cer06p] Raphaël Cerf. Ising model. Lecture Notes in Mathematics, 1878:15–24, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_2.pdf.

Cerf:2006:LDT

[Cer06q] Raphaël Cerf. Large deviation theory. Lecture Notes in Mathematics, 1878:67–74, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_6.pdf.

Cerf:2006:LFP

[Cer06r] Raphaël Cerf. LDP for FK percolation. Lecture Notes in Mathematics, 1878:229–239, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_18.pdf.

Cerf:2006:LI

[Cer06s] Raphaël Cerf. LDP for Ising. Lecture Notes in Mathematics, 1878:241–252, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_19.pdf.

Cerf:2006:LCS

[Cer06t] Raphaël Cerf. LDP for the cluster shapes. Lecture Notes in Mathematics, 1878:203–214, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_16.pdf.

Cerf:2006:PCS

[Cer06u] Raphaël Cerf. Phase coexistence and subadditivity. Lecture Notes in Mathematics, 1878:3–12, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_1.pdf.

Cerf:2006:SFP

[Cer06v] Raphaël Cerf. Sets of finite perimeter. Lecture Notes in Mathematics, 1878:159–172, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_13.pdf.

Cerf:2006:SE

[Cer06w] Raphaël Cerf. Surface energy. Lecture Notes in Mathematics, 1878:173–188, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_14.pdf.

Cerf:2006:SLD

[Cer06x] Raphaël Cerf. Surface large deviation principles. Lecture Notes in Mathematics, 1878:75–84, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_7.pdf.

Cerf:2006:ST

[Cer06y] Raphaël Cerf. Surface tension. Lecture Notes in Mathematics, 1878:129–145, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_11.pdf.

Cerf:2006:VLD

[Cer06z] Raphaël Cerf. Volume large deviations. Lecture Notes in Mathematics, 1878:85–102, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-34806-1_8.pdf.

Cerf:2006:WC

[Cer06-27] Raphaël Cerf. The Wulff crystal. Lecture Notes in Mathematics, 1878:45–64, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_5.pdf.

Cerf:2006:WT

[Cer06-28] Raphaël Cerf. The Wulff theorem. Lecture Notes in Mathematics, 1878:189–199, 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-34806-1_15.pdf.

Cowling:2008:FM

[CFK^{+}08] Michael Cowling, Edward Frenkel, Masaki Kashiwara, Alain Valette, David A. Vogan, Jr., and Nolan R. Wallach. Front matter. Lecture Notes in Mathematics, 1931:i–xii, 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-76892-0/
1.pdf.

Conti:2003:TBC

[CG03] Roberto Conti and Marcello Galeotti. Totally bounded cubic systems in R^{2}. Lecture Notes in Mathematics, 1822:103– 171, 2003. CODEN LNMAA2. ISBN 3-540-40786-3 (print), 3-540-45204-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45204-1_2/.

Chapoton:2001:ECO

[Cha01] Frédéric Chapoton. Un endofoncteur de la catégorie des opérades. (French) []. Lecture Notes in Mathematics, 1763:105–110, 2001. CODEN LNMAA2. ISBN 3-540-42194-7 (print), 3-540-45328-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45328-8_4/.

Chafai:2003:GME

[Cha03] Djalil Chafaï. Gaussian maximum of entropy and reversed log-Sobolev inequality. Lecture Notes in Mathematics, 1801:194– 200, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_5/.

Chaplain:2008:MAC

[Cha08a] Mark A. J. Chaplain. Modelling aspects of cancer growth: Insight from mathematical and numerical analysis and computational simulation. Lecture Notes in Mathematics, 1940:147–200, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78362-6_3.pdf.

Chazottes:2008:ESE

[Cha08b] Jean-René Chazottes. Equilibrium States and the Ergodic Theory of Anosov Diffeomorphisms, volume 470 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-77605-2 (print), 3-540-77695-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 75 pp. LCCN QA3 .L28

no. 470. URL http://link.springer.com/book/10.1007/978-
3-540-77695-6;   http://www.springerlink.com/content/
978-3-540-77695-6.

Cheridito:2003:RGM

[Che03] Patrick Cheridito. Representations of Gaussian measures that are equivalent to Wiener measure. Lecture Notes in Mathematics, 1832:81–89, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40004-2_3/.

Chidume:2009:AHI

[Chi09a] Charles Chidume. Applications to Hammerstein integral equations. Lecture Notes in Mathematics, 1965:169–191, 2009. CO-DEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_13.pdf.

Chidume:2009:AFP

[Chi09b] Charles Chidume. Approximation of fixed points of Lipschitz pseudo-contractive mappings. Lecture Notes in Mathematics, 1965:151–160, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-84882-190-3_11.pdf.

Chidume:2009:CFPb

[Chi09c] Charles Chidume. Common fixed points for countable families of nonexpansive mappings. Lecture Notes in Mathematics, 1965:215–229, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-84882-190-3_16.pdf.

Chidume:2009:CFPc

[Chi09d] Charles Chidume. Common fixed points for families of commuting nonexpansive mappings. Lecture Notes in Mathematics, 1965:231–242, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-84882-190-3_17.pdf.

Chidume:2009:CFPa

[Chi09e] Charles Chidume. Common fixed points for finite families of nonexpansive mappings. Lecture Notes in Mathematics, 1965: 205–214, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-1-84882-190-3_15.pdf.

Chidume:2009:DMB

[Chi09f] Charles Chidume. Duality maps in Banach spaces. Lecture Notes in Mathematics, 1965:19–28, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-1-84882-190-3_3.pdf.

Chidume:2009:EMI

[Chi09g] Charles Chidume. An example; Mann iteration for strictly pseudo-contractive mappings. Lecture Notes in Mathematics, 1965:141–149, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-84882-190-3_10.pdf.

Chidume:2009:FFL

[Chi09h] Charles Chidume. Finite families of Lipschitz pseudo-contractive and accretive mappings. Lecture Notes in Mathematics, 1965: 243–250, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-1-84882-190-3_18.pdf.

Chidume:2009:FFN

[Chi09i] Charles Chidume. Finite families of non-self asymptotically nonexpansive mappings. Lecture Notes in Mathematics, 1965:257– 270, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-1-84882-190-3_20.pdf.

Chidume:2009:FM

[Chi09j] Charles Chidume. Front matter. Lecture Notes in Mathematics, 1965:i–xvii, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-1-84882-190-3/1.pdf.

Chidume:2009:GLA

[Chi09k] Charles Chidume. Generalized Lipschitz accretive and pseudocontractive mappings. Lecture Notes in Mathematics, 1965:161– 167, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-1-84882-190-3_12.pdf.

Chidume:2009:GLP

[Chi09l] Charles Chidume. Generalized Lipschitz pseudo-contractive and accretive mappings. Lecture Notes in Mathematics, 1965:251–256, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_19.pdf.

Chidume:2009:GPB

[Chi09m] Charles Chidume. Geometric Properties of Banach Spaces and Nonlinear Iterations, volume 1965 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 326 pp. LCCN QA3 .L28 no.

1965. URL http://link.springer.com/book/10.1007/978-
1-84882-190-3;   http://www.springerlink.com/content/
978-1-84882-190-3.

Chidume:2009:HSD

[Chi09n] Charles Chidume. Hybrid steepest descent method for variational inequalities. Lecture Notes in Mathematics, 1965:87–111, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_7.pdf.

Chidume:2009:IUC

[Chi09o] Charles Chidume. Inequalities in uniformly convex spaces. Lecture Notes in Mathematics, 1965:29–44, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-1-84882-190-3_4.
pdf.

Chidume:2009:IUS

[Chi09p] Charles Chidume. Inequalities in uniformly smooth spaces. Lecture Notes in Mathematics, 1965:45–55, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-1-84882-190-3_5.
pdf.

Chidume:2009:IPZ

[Chi09q] Charles Chidume. Iteration processes for zeros of generalized Faccretive mappings. Lecture Notes in Mathematics, 1965:129–140, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_9.pdf.

Chidume:2009:IMF

[Chi09r] Charles Chidume. Iterative method for fixed points of nonexpansive mappings. Lecture Notes in Mathematics, 1965:57–86, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_6.pdf.

Chidume:2009:IMS

[Chi09s] Charles Chidume. Iterative methods for some generalizations of nonexpansive maps. Lecture Notes in Mathematics, 1965:193–204, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_14.pdf.

Chidume:2009:IMZ

[Chi09t] Charles Chidume. Iterative methods for zeros of F – accretivetype operators. Lecture Notes in Mathematics, 1965:113–127, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_8.pdf.

Chidume:2009:SS

[Chi09u] Charles Chidume. Smooth spaces. Lecture Notes in Mathematics, 1965:11–18, 2009. CODEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-84882-190-3_2.pdf.

Chidume:2009:SGP

[Chi09v] Charles Chidume. Some geometric properties of Banach spaces. Lecture Notes in Mathematics, 1965:1–9, 2009. CO-DEN LNMAA2. ISBN 1-84882-189-1 (print), 1-84882-190-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-84882-190-3_1.pdf.

Chlouveraki:2009:BM

[Chl09a] Maria Chlouveraki. Back matter. Lecture Notes in Mathematics, 1981:1–32, 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-03064-2/1.pdf.

Chlouveraki:2009:BFC

[Chl09b] Maria Chlouveraki. Blocks and Families for Cyclotomic Hecke Algebras, volume 1981 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 160 pp. LCCN QA3 .L28 no. 1981. URL http:

//link.springer.com/book/10.1007/978-3-642-03064-2;
http://www.springerlink.com/content/978-3-642-03064-2.

Chlouveraki:2009:FM

[Chl09c] Maria Chlouveraki. Front matter. Lecture Notes in Mathematics, 1981:1–11, 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-03064-2/1.pdf.

Chlouveraki:2009:B

[Chl09d] Maria Chlouveraki. On blocks. Lecture Notes in Mathematics, 1981:21–59, 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-03064-2_2.pdf.

Chlouveraki:2009:CA

[Chl09e] Maria Chlouveraki. On commutative algebra. Lecture Notes in Mathematics, 1981:1–19, 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-03064-2_1.pdf.

Chlouveraki:2009:EA

[Chl09f] Maria Chlouveraki. On essential algebras. Lecture Notes in Mathematics, 1981:61–70, 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-03064-2_3.pdf.

Chlouveraki:2009:HA

[Chl09g] Maria Chlouveraki. On Hecke algebras. Lecture Notes in Mathematics, 1981:71–89, 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-03064-2_4.pdf.

Chlouveraki:2009:DRB

[Chl09h] Maria Chlouveraki. On the determination of the Rouquier blocks. Lecture Notes in Mathematics, 1981:91–132, 2009. CODEN LNMAA2. ISBN 3-642-03063-7 (print), 3-642-03064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-03064-2_5.pdf.

Chow:2003:LDS

[Cho03] Shui-Nee Chow. Lattice dynamical systems. Lecture Notes in Mathematics, 1822:1–102, 2003. CODEN LNMAA2. ISBN 3-540-40786-3 (print), 3-540-45204-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45204-1_1/.

Christensen:2000:AC

[Chr00a] Lars Winther Christensen. Auslander categories. Lecture Notes in Mathematics, 1747:65–90, 2000. CODEN LNMAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103986/.

Christensen:2000:CGD

[Chr00b] Lars Winther Christensen. The classical Gorenstein dimension. Lecture Notes in Mathematics, 1747:17–40, 2000. CODEN LN-MAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103984/.

Christensen:2000:CP

[Chr00c] Lars Winther Christensen. Conventions and prerequisites. Lecture Notes in Mathematics, 1747:9–15, 2000. CODEN LN-MAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103983/.

Christensen:2000:DRC

[Chr00d] Lars Winther Christensen. G-dimension and reflexive complexes. Lecture Notes in Mathematics, 1747:41–63, 2000. CODEN LN-MAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103985/.

Christensen:2000:F

[Chr00e] Lars Winther Christensen. G-flatness. Lecture Notes in Mathematics, 1747:113–134, 2000. CODEN LNMAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103988/.

Christensen:2000:Ib

[Chr00f] Lars Winther Christensen. G-injectivity. Lecture Notes in Mathematics, 1747:135–158, 2000. CODEN LNMAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103989/.

Christensen:2000:P

[Chr00g] Lars Winther Christensen. G-projectivity. Lecture Notes in Mathematics, 1747:91–112, 2000. CODEN LNMAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103987/.

Christensen:2000:GD

[Chr00h] Lars Winther Christensen. Gorenstein Dimensions, volume 1747 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 204 pp. LCCN QA3 .L28 no. 1747. URL http:/

/link.springer.com/book/10.1007/BFb0103980;   http://
www.springerlink.com/content/978-3-540-40008-0.

Christensen:2000:Ia

[Chr00i] Lars Winther Christensen. Introduction. Lecture Notes in Mathematics, 1747:1–2, 2000. CODEN LNMAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103981.

Christensen:2000:S

[Chr00j] Lars Winther Christensen. Synopsis. Lecture Notes in Mathematics, 1747:3–8, 2000. CODEN LNMAA2. ISBN 3-540-41132-1 (print), 3-540-40008-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103982/.

Chueshov:2002:GFA

[Čhu02a] Igor Čhuešhov. 1. General facts about random dynamical systems. Lecture Notes in Mathematics, 1779:9–53, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45815-9_
2/.

Chueshov:2002:GRD

[Čhu02b] Igor Čhuešhov. 2. Generation of random dynamical systems. Lecture Notes in Mathematics, 1779:55–81, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45815-9_3/.

Chueshov:2002:OPR

[Čhu02c] Igor Čhuešhov. 3. Order-preserving random dynamical systems. Lecture Notes in Mathematics, 1779:83–111, 2002. CODEN LN-MAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45815-9_
4/.

Chueshov:2002:SRD

[Čhu02d] Igor Čhuešhov. 4. Sublinear random dynamical systems. Lecture Notes in Mathematics, 1779:113–141, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45815-9_5/.

Chueshov:2002:CRD

[Čhu02e] Igor Čhuešhov. 5. Cooperative random differential equations. Lecture Notes in Mathematics, 1779:143–183, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45815-9_
6/.

Chueshov:2002:CSD

[Čhu02f] Igor Čhuešhov. 6. Cooperative stochastic differential equations. Lecture Notes in Mathematics, 1779:185–225, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45815-9_
7/.

Chueshov:2002:Ib

[Čhu02g] Igor Čhuešhov. Index. Lecture Notes in Mathematics, 1779:233– 234, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45815-9_9.

Chueshov:2002:Ia

[Čhu02h] Igor Čhuešhov. Introduction. Lecture Notes in Mathematics, 1779:1–7, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45815-9_1/.

Chueshov:2002:MRS

[Chu02i] Igor Chueshov. Monotone Random Systems Theory and Applications, volume 1779 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 234 pp. LCCN QA3 .L28 no. 1779. URL http:/

/link.springer.com/book/10.1007/b83277;   http://www.
springerlink.com/content/978-3-540-45815-9.

Chueshov:2002:R

[Čhu02j] Igor Čhuešhov. References. Lecture Notes in Mathematics, 1779: 227–231, 2002. CODEN LNMAA2. ISBN 3-540-43246-9 (print), 3-540-45815-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45815-9_8/.

Chu:2008:BM

[Chu08a] Cho-Ho Chu. Back matter. Lecture Notes in Mathematics, 1956: 101–108, 2008. CODEN LNMAA2. ISBN 3-540-69797-7 (print), 3-540-69798-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-69798-5/1.pdf.

Chu:2008:CS

[Chu08b] Cho-Ho Chu. Convolution semigroups. Lecture Notes in Mathematics, 1956:87–100, 2008. CODEN LNMAA2. ISBN 3-540-69797-7 (print), 3-540-69798-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69798-5_4.pdf.

Chu:2008:FM

[Chu08c] Cho-Ho Chu. Front matter. Lecture Notes in Mathematics, 1956: i–ix, 2008. CODEN LNMAA2. ISBN 3-540-69797-7 (print), 3-540-69798-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69798-5/1.pdf.

Chu:2008:I

[Chu08d] Cho-Ho Chu. Introduction. Lecture Notes in Mathematics, 1956: 1–4, 2008. CODEN LNMAA2. ISBN 3-540-69797-7 (print), 3-540-69798-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69798-5_1.pdf.

Chu:2008:LSM

[Chu08e] Cho-Ho Chu. Lebesgue spaces of matrix functions. Lecture Notes in Mathematics, 1956:5–19, 2008. CODEN LNMAA2. ISBN 3-540-69797-7 (print), 3-540-69798-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69798-5_2.pdf.

Chu:2008:MCOa

[Chu08f] Cho-Ho Chu. Matrix convolution operators. Lecture Notes in Mathematics, 1956:21–85, 2008. CODEN LNMAA2. ISBN 3-540-69797-7 (print), 3-540-69798-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69798-5_3.pdf.

Chu:2008:MCOb

[Chu08g] Cho-Ho Chu. Matrix Convolution Operators on Groups, volume 1956 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-69797-7 (print), 3-540-69798-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 108 pp. LCCN QA3 .L28 no. 1956. URL http://link.

springer.com/book/10.1007/978-3-540-69798-5;   http:/
/www.springerlink.com/content/978-3-540-69798-5.

Chybiryakov:2007:IIF

[Chy07] Oleksandr Chybiryakov. Itô’s integrated formula for strict local martingales with jumps. Lecture Notes in Mathematics, 1899:375– 388, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_20.pdf.

Cipriani:2008:DFN

[Cip08] Fabio Cipriani. Dirichlet forms on noncommutative spaces. Lecture Notes in Mathematics, 1954:161–276, 2008. CO-DEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69365-9_5.pdf.

Coquet:2006:NDP

[CJMS06] François Coquet, Adam Jakubowski, Jean Mémin, and Leszek Słominski. Natural decomposition of processes and weak Dirichlet processes. Lecture Notes in Mathematics, 1874:81–116, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_8.pdf.

Chu:2002:HFF

[CL02a] Cho-Ho Chu and Anthony To-Ming Lau. Harmonic functions on Fourier algebras. Lecture Notes in Mathematics, 1782:51– 89, 2002. CODEN LNMAA2. ISBN 3-540-43595-6 (print), 3-540-47793-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-47793-8_3/.

Chu:2002:HFG

[CL02b] Cho-Ho Chu and Anthony To-Ming Lau. Harmonic Functions on Groups and Fourier Algebras, volume 1782 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43595-6 (print), 3-540-47793-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 100 pp. LCCN QA3 .L28 no. 1782. URL http:/

/link.springer.com/book/10.1007/b83280;   http://www.
springerlink.com/content/978-3-540-47793-8.

Chu:2002:HFL

[CL02c] Cho-Ho Chu and Anthony To-Ming Lau. Harmonic functions on locally compact groups. Lecture Notes in Mathematics, 1782: 5–50, 2002. CODEN LNMAA2. ISBN 3-540-43595-6 (print), 3-540-47793-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-47793-8_2/.

Chu:2002:I

[CL02d] Cho-Ho Chu and Anthony To-Ming Lau. Introduction. Lecture Notes in Mathematics, 1782:1–4, 2002. CODEN LNMAA2. ISBN 3-540-43595-6 (print), 3-540-47793-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-47793-8_1/.

Chu:2002:LSS

[CL02e] Cho-Ho Chu and Anthony To-Ming Lau. List of symbols and subject index. Lecture Notes in Mathematics, 1782:98–100, 2002. CODEN LNMAA2. ISBN 3-540-43595-6 (print), 3-540-47793-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-47793-8_5/.

Chu:2002:R

[CL02f] Cho-Ho Chu and Anthony To-Ming Lau. References. Lecture Notes in Mathematics, 1782:90–97, 2002. CODEN LNMAA2. ISBN 3-540-43595-6 (print), 3-540-47793-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-47793-8_4/.

Cepa:2007:NMC

[CL07] Emmanuel Cépa and Dominique Lépingle. No multiple collisions for mutually repelling Brownian particles. Lecture Notes in Mathematics, 1899:241–246, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71189-6_12.pdf.

Capasso:2008:MPL

[CL08] Vincenzo Capasso and Mirosław Lachowicz. Multiscale Problems in the Life Sciences: From Microscopic to Macroscopic, volume 1940 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 321 pp. LCCN QA3 .L28 no. 1940. URL http://link.

springer.com/book/10.1007/978-3-540-78362-6;   http:/
/www.springerlink.com/content/978-3-540-78362-6.

Clarkson:2006:BM

[Cla06a] Peter A. Clarkson. Back matter. Lecture Notes in Mathematics, 1883:413–422, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-36716-1/1.pdf.

Clarkson:2006:PEN

[Cla06b] Peter A. Clarkson. Painlevé equations — nonlinear special functions. Lecture Notes in Mathematics, 1883:331–411, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36716-1_7.pdf.

Cao:2008:BM

[CLM^{+}08a] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Back matter. Lecture Notes in Mathematics, 1948:225–257, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-68481-7/1.pdf.

Cao:2008:CDL

[CLM^{+}08b] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. A contrario decision: the LLD method. Lecture Notes in Mathematics, 1948:81–92, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-68481-7_5.
pdf.

Cao:2008:ER

[CLM^{+}08c] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Experimental results. Lecture Notes in Mathematics, 1948:167–182, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-68481-7_9.pdf.

Cao:2008:EMC

[CLM^{+}08d] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Extracting meaningful curves from images. Lecture Notes in Mathematics, 1948:15–35, 2008. CO-DEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68481-7_2.pdf.

Cao:2008:FM

[CLM^{+}08e] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Front matter. Lecture Notes in Mathematics, 1948:i–xi, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-68481-7/1.pdf.

Cao:2008:GSC

[CLM^{+}08f] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Grouping spatially coherent meaningful matches. Lecture Notes in Mathematics, 1948:151–165, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68481-7_8.pdf.

Cao:2008:HCV

[CLM^{+}08g] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Hierarchical clustering and validity assessment. Lecture Notes in Mathematics, 1948:129–149, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68481-7_7.pdf.

Cao:2008:I

[CLM^{+}08h] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Introduction. Lecture Notes in Mathematics, 1948:1–12, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-68481-7_1.pdf.

Cao:2008:ILL

[CLM^{+}08i] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Invariant level line encoding. Lecture Notes in Mathematics, 1948:61–77, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-68481-7_4.pdf.

Cao:2008:MME

[CLM^{+}08j] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Meaningful matches: Experiments on LLD and MSER. Lecture Notes in Mathematics, 1948:93–125, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68481-7_6.pdf.

Cao:2008:RSD

[CLM^{+}08k] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Robust shape directions. Lecture Notes in Mathematics, 1948:41–59, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-68481-7_3.pdf.

Cao:2008:SSC

[CLM^{+}08l] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. Securing SIFT with a contrario techniques. Lecture Notes in Mathematics, 1948:209–224, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68481-7_11.pdf.

Cao:2008:SM

[CLM^{+}08m] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. The SIFT method. Lecture Notes in Mathematics, 1948:185–208, 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-68481-7_10.pdf.

Cao:2008:TSI

[CLM^{+}08n] Frédéric Cao, José-Luis Lisani, Jean-Michel Morel, Pablo Musé, and Frédéric Sur. A Theory of Shape Identification, volume 1948 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-68480-8 (print), 3-540-68481-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 257 pp. LCCN QA3 .L28 no. 1948. URL http://link.

springer.com/book/10.1007/978-3-540-68481-7;   http:/
/www.springerlink.com/content/978-3-540-68481-7.

Crimaldi:2007:SFS

[CLP07] Irene Crimaldi, Giorgio Letta, and Luca Pratelli. A strong form of stable convergence. Lecture Notes in Mathematics, 1899:203–225, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_9.pdf.

Cherednik:2002:HTD

[CM02] Ivan Cherednik and Yavor Markov. Hankel transform via double Hecke algebra. Lecture Notes in Mathematics, 1804:1–25, 2002. CODEN LNMAA2. ISBN 3-540-00224-3 (print), 3-540-36205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36205-0_1/.

Capasso:2003:SGS

[CM03] Vincenzo Capasso and Alessandra Micheletti. Stochastic geometry of spatially structured birth and growth processes. application to crystallization processes. Lecture Notes in Mathematics, 1802: 1–39, 2003. CODEN LNMAA2. ISBN 3-540-00295-2 (print), 3-540-36259-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36259-3_1/.

Cannone:2006:MFT

[CM06a] Marco Cannone and Tetsuro Miyakawa, editors. Mathematical Foundation of Turbulent Viscous Flows: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy, SEptember 1-5, 2003, volume 1871 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-28586-5 (print), 3-540-32454-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1871. URL http:

//link.springer.com/book/10.1007/b11545989;   http://
www.springerlink.com/content/978-3-540-32454-6.

Chaplain:2006:MMS

[CM06b] Mark Chaplain and Anastasios Matzavinos. Mathematical modelling of spatio-temporal phenomena in tumour immunology. Lecture Notes in Mathematics, 1872:131–183, 2006. CODEN LN-MAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11561606_4.pdf.

Capasso:2008:RSP

[CM08] Vincenzo Capasso and Daniela Morale. Rescaling stochastic processes: Asymptotics. Lecture Notes in Mathematics, 1940:91–146, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78362-6_2.pdf.

Capriz:2008:MMG

[CMG08] Gianfranco Capriz, Paolo Maria Mariano, and Pasquale Giovine. Mathematical Models of Granular Matter, volume 1937 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 212 pp. LCCN

QA3 .L28 no. 1937. URL http://link.springer.com/book/
10.1007/978-3-540-78277-3;   http://www.springerlink.
com/content/978-3-540-78277-3.

Cherednik:2002:IHA

[CMHL02] Ivan Cherednik, Yavor Markov, Roger Howe, and George Lusztig, editors. Iwahori–Hecke Algebras and their Representation Theory: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy, June 28–July 6, 1999, volume 1804 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-00224-3 (print), 3-540-36205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1804. URL

http://link.springer.com/book/10.1007/b10326;   http:
//www.springerlink.com/content/978-3-540-36205-0.

Capasso:2003:TSS

[CMI^{+}03] Vincenzo Capasso, Ely Merzbach, B. Gail Ivanoff, Marco Dozzi, Robert C. Dalang, and Thomas S. Mountford, editors. Topics in Spatial Stochastic Processes: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy, July 1-8, 2001, volume 1802 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00295-2 (print), 3-540-36259-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1802. URL http:/

/link.springer.com/book/10.1007/b10143;   http://www.
springerlink.com/content/978-3-540-36259-3.

Cancrini:2009:BM

[CMRT09a] N. Cancrini, F. Martinelli, C. Robert, and C. Toninelli. Back matter. Lecture Notes in Mathematics, 1970:1–9, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-92796-9/
1.pdf.

Cancrini:2009:FSM

[CMRT09b] N. Cancrini, F. Martinelli, C. Robert, and C. Toninelli. Facilitated spin models: Recent and new results. Lecture Notes in Mathematics, 1970:1–34, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-92796-9_7.pdf.

Caenepeel:2002:G

[CMZ02a] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 1. Generalities. Lecture Notes in Mathematics, 1787:3–37, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48042-6_
1/.

Caenepeel:2002:DKH

[CMZ02b] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 2. Doi– Koppinen Hopf modules and entwined modules. Lecture Notes in Mathematics, 1787:39–87, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48042-6_2/.

Caenepeel:2002:FSFa

[CMZ02c] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 3. Frobenius and separable functors for entwined modules. Lecture Notes in Mathematics, 1787:89–157, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48042-6_3/.

Caenepeel:2002:A

[CMZ02d] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 4. Applications. Lecture Notes in Mathematics, 1787:159–213, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48042-6_4/.

Caenepeel:2002:YDM

[CMZ02e] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 5. Yetter– Drinfeld modules and the quantum Yang–Baxter equation. Lecture Notes in Mathematics, 1787:217–243, 2002. CODEN LN-MAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48042-6_
5/.

Caenepeel:2002:HMP

[CMZ02f] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 6. Hopf modules and the pentagon equation. Lecture Notes in Mathematics, 1787:245–300, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-48042-6_6/.

Caenepeel:2002:LDL

[CMZ02g] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 7. Long dimodules and the Long equation. Lecture Notes in Mathematics, 1787:301–316, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-48042-6_7/.

Caenepeel:2002:FSE

[CMZ02h] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. 8. The Frobenius-separability equation. Lecture Notes in Mathematics, 1787:317–343, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-48042-6_8/.

Caenepeel:2002:FSFb

[CMZ02i] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. Frobenius and Separable Functors for Generalized Module Categories and Nonlinear Equations, volume 1787 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 354 pp. LCCN QA3 .L28 no. 1787. URL

http://link.springer.com/book/10.1007/b83849;   http:
//www.springerlink.com/content/978-3-540-48042-6.

Caenepeel:2002:I

[CMZ02j] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. Index. Lecture Notes in Mathematics, 1787:353–354, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48042-6_
10.

Caenepeel:2002:R

[CMZ02k] Stefaan Caenepeel, Gigel Militaru, and Shenglin Zhu. References. Lecture Notes in Mathematics, 1787:345–352, 2002. CODEN LNMAA2. ISBN 3-540-43782-7 (print), 3-540-48042-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48042-6_
9/.

Cellina:2000:OSD

[CO00] Arrigo Cellina and António Ornelas, editors. Optimal Shape Design: Lectures given at the joint C.I.M./C.I.M.E. Summer School held in Tróia, Portugal, June 1–6, 1998, volume 1740 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67971-5 (print), 3-540-44486-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1740.

URL http://link.springer.com/book/10.1007/BFb0106739;
 http://www.springerlink.com/content/978-3-540-44486-
2.

Cohen:2003:TAC

[Coh03] Albert Cohen. Theoretical, applied and computational aspects of nonlinear approximation. Lecture Notes in Mathematics, 1825: 1–29, 2003. CODEN LNMAA2. ISBN 3-540-20099-1 (print), 3-540-39810-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39810-3_1/.

Conrad:2000:GDB

[Con00] Brian Conrad. Grothendieck Duality and Base Change, volume 1750 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000.

CODEN LNMAA2. ISBN 3-540-41134-8 (softcover). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 296 pp. LCCN QA3 .L28 no. 1750; QA3 .L28 no. 1750; QA564. URL http://

link.springer.com/book/10.1007/b75857;   http://rave.
ohiolink.edu/ebooks/ebc/354040015X;   http://www.loc.
gov/catdir/enhancements/fy0816/00069238-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0816/00069238-t.html;
 http://www.springerlink.com/openurl.asp?genre=book&isbn=
978-3-540-41134-5;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0075-8434&volume=1750.

Connes:2004:CCN

[Con04] Alain Connes. Cyclic cohomology, noncommutative geometry and quantum group symmetries. Lecture Notes in Mathematics, 1831: 1–71, 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39702-1_1/.

Constantin:2006:EEN

[Con06] Peter Constantin. Euler equations, Navier–Stokes equations and turbulence. Lecture Notes in Mathematics, 1871:1–43, 2006. CODEN LNMAA2. ISBN 3-540-28586-5 (print), 3-540-32454-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11545989_1.pdf;   http://link.springer.com/content/
pdf/bfm:978-3-540-32454-6/1.pdf.

Cosner:2008:RDE

[Cos08] C. Cosner. Reaction–diffusion equations and ecological modeling. Lecture Notes in Mathematics, 1922:77–115, 2008. CO-DEN LNMAA2. ISBN 3-540-74328-6 (print), 3-540-74331-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74331-6_3.pdf.

Coutin:2007:FM

[Cou07a] Laure Coutin. Front matter. Lecture Notes in Mathematics, 1899: 67, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-71189-6/2/1.pdf.

Coutin:2007:ISC

[Cou07b] Laure Coutin. An introduction to (stochastic) calculus with respect to fractional Brownian motion. Lecture Notes in Mathematics, 1899:3–65, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
content/pdf/10.1007/978-3-540-71189-6_1.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-540-71189-6/
1.pdf;   http://link.springer.com/content/pdf/bfm:978-
3-540-71189-6/1/1.pdf.

Cowling:2008:ART

[Cow08] Michael Cowling. Applications of representation theory to harmonic analysis of Lie groups (and vice versa). Lecture Notes in Mathematics, 1931:1–50, 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-76892-0_1.pdf.

Cox:2008:ECW

[Cox08] A. M. G. Cox. Extending Chacon–Walsh: Minimality and generalised starting distributions. Lecture Notes in Mathematics, 1934:233–264, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77913-1_12.pdf.

Catoni:2004:SLT

[CP04] Olivier Catoni and Jean Picard, editors. Statistical Learning Theory and Stochastic Optimization: École d’Été de Probabilités de Saint-Flour XXXI — 2001, volume 1851 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22572-2 (print), 3-540-44507-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1851. URL http:/

/link.springer.com/book/10.1007/b99352;   http://www.
loc.gov/catdir/enhancements/fy0823/2004109143-d.html;
 http://www.loc.gov/catdir/enhancements/fy0823/2004109143-
t.html;   http://www.springerlink.com/content/978-3-540-
44507-4.

Cerf:2006:WCI

[CP06] Raphaël Cerf and Jean Picard, editors. The Wulff Crystal in Ising and Percolation Models: École d’Été de Probabilités de Saint-Flour XXXIV — 2004, volume 1878 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-30988-8 (print), 3-540-34806-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1878; QC175.16.P5 C44 2006. URL http://

link.springer.com/book/10.1007/b128410;   http://www.
loc.gov/catdir/enhancements/fy0663/2006921625-d.html;
 http://www.loc.gov/catdir/enhancements/fy0824/2006921625-
t.html;   http://www.springerlink.com/content/978-3-540-
34806-1.

Catuogno:2007:PHM

[CR07] Pedro J. Catuogno and Paulo R. C. Ruffino. Product of harmonic maps is harmonic: a stochastic approach. Lecture Notes in Mathematics, 1899:227–233, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71189-6_10.pdf.

Carpio:2008:TDS

[CR08] Ana Carpio and Maria Luisa Rapún. Topological derivatives for shape reconstruction. Lecture Notes in Mathematics, 1943:85– 133, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78547-7_5.pdf.

Crandall:2008:VLE

[Cra08] Michael G. Crandall. A visit with the ∞-Laplace equation. Lecture Notes in Mathematics, 1927:75–122, 2008. CO-DEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75914-0_3.pdf.

Cherny:2005:SII

[CS05] Alexander Cherny and Albert Shiryaev. On stochastic integrals up to infinity and predictable criteria for integrability. Lecture Notes in Mathematics, 1857:165–185, 2005. CO-DEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_12.pdf.

Caffarelli:2008:IHP

[CS08] Luis Caffarelli and Luis Silvestre. Issues in homogenization for problems with non divergence structure. Lecture Notes in Mathematics, 1927:43–74, 2008. CODEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75914-0_2.pdf.

Ching-Sung:2002:RMC

[CSM02] Chou Ching-Sung and P.-A. Meyer. Sur la représentation des martingales comme intégrales stochastiques dans les processus ponctuels. (French) []. Lecture Notes in Mathematics, 1771:417– 427, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_18/.

Catanese:2008:SMA

[CT08] Fabrizio Catanese and Gang Tian, editors. Symplectic 4-Manifolds and Algebraic Surfaces: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy September 2–10, 2003, volume 1938 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1938. URL http://link.

springer.com/book/10.1007/978-3-540-78279-7;   http:/
/www.springerlink.com/content/978-3-540-78279-7.

Cuntz:2004:CTB

[Cun04] Joachim Cuntz. Cyclic theory and the bivariant Chern– Connes character. Lecture Notes in Mathematics, 1831:73–135, 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39702-1_2/.

Cutland:2000:LM

[Cut00a] Nigel J. Cutland. 1. Loeb measures. Lecture Notes in Mathematics, 1751:1–28, 2000. CODEN LNMAA2. ISBN 3-540-41384-7 (print), 3-540-44531-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44531-9_1/.

Cutland:2000:SFM

[Cut00b] Nigel J. Cutland. 2. Stochastic fluid mechanics. Lecture Notes in Mathematics, 1751:29–60, 2000. CODEN LNMAA2. ISBN 3-540-41384-7 (print), 3-540-44531-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44531-9_2/.

Cutland:2000:SCV

[Cut00c] Nigel J. Cutland. 3. Stochastic calculus of variations. Lecture Notes in Mathematics, 1751:61–84, 2000. CODEN LNMAA2. ISBN 3-540-41384-7 (print), 3-540-44531-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44531-9_3/.

Cutland:2000:MFT

[Cut00d] Nigel J. Cutland. 4. Mathematical finance theory. Lecture Notes in Mathematics, 1751:85–101, 2000. CODEN LNMAA2. ISBN 3-540-41384-7 (print), 3-540-44531-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44531-9_4/.

Cutland:2000:I

[Cut00e] Nigel J. Cutland. Index. Lecture Notes in Mathematics, 1751: 109–111, 2000. CODEN LNMAA2. ISBN 3-540-41384-7 (print), 3-540-44531-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44531-9_6/.

Cutland:2000:LMP

[Cut00f] Nigel J. Cutland. Loeb Measures in Practice: Recent Advances, volume 1751 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41384-7 (print), 3-540-44531-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 111 pp. LCCN QA3 .L28 no. 1751. URL http:/

/link.springer.com/book/10.1007/b76881;   http://www.
springerlink.com/content/978-3-540-44531-9.

Cutland:2000:R

[Cut00g] Nigel J. Cutland. References. Lecture Notes in Mathematics, 1751:103–107, 2000. CODEN LNMAA2. ISBN 3-540-41384-7 (print), 3-540-44531-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44531-9_5/.

Cutkosky:2002:Ia

[Cut02a] Steven Dale Cutkosky. 1. Introduction. Lecture Notes in Mathematics, 1786:1–8, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-48030-3_1/.

Cutkosky:2002:AX

[Cut02b] Steven Dale Cutkosky. 10. A_{r}(X). Lecture Notes in Mathematics, 1786:109, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48030-3_10.

Cutkosky:2002:RSC

[Cut02c] Steven Dale Cutkosky. 11. Reduction of ν in a special case. Lecture Notes in Mathematics, 1786:110–130, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-48030-3_11/.

Cutkosky:2002:RSS

[Cut02d] Steven Dale Cutkosky. 12. Reduction of ν in a second special case. Lecture Notes in Mathematics, 1786:131–149, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48030-3_12/.

Cutkosky:2002:Ra

[Cut02e] Steven Dale Cutkosky. 13. Resolution 1. Lecture Notes in Mathematics, 1786:150–162, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_13/.

Cutkosky:2002:Rb

[Cut02f] Steven Dale Cutkosky. 14. Resolution 2. Lecture Notes in Mathematics, 1786:163–175, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_14/.

Cutkosky:2002:Rc

[Cut02g] Steven Dale Cutkosky. 15. Resolution 3. Lecture Notes in Mathematics, 1786:176–184, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_15/.

Cutkosky:2002:Rd

[Cut02h] Steven Dale Cutkosky. 16. Resolution 4. Lecture Notes in Mathematics, 1786:185–187, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_16/.

Cutkosky:2002:PMTb

[Cut02i] Steven Dale Cutkosky. 17. Proof of the main theorem. Lecture Notes in Mathematics, 1786:188, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_17.

Cutkosky:2002:M

[Cut02j] Steven Dale Cutkosky. 18. Monomialization. Lecture Notes in Mathematics, 1786:189–223, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_18/.

Cutkosky:2002:T

[Cut02k] Steven Dale Cutkosky. 19. Toroidalization. Lecture Notes in Mathematics, 1786:224–231, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_19/.

Cutkosky:2002:LM

[Cut02l] Steven Dale Cutkosky. 2. Local monomialization. Lecture Notes in Mathematics, 1786:9–10, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_2.

Cutkosky:2002:GND

[Cut02m] Steven Dale Cutkosky. 20. Glossary of notations and definitions. Lecture Notes in Mathematics, 1786:232–233, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48030-3_
20.

Cutkosky:2002:MML

[Cut02n] Steven Dale Cutkosky. 3. Monomialization of morphisms in low dimensions. Lecture Notes in Mathematics, 1786:11–13, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48030-3_3/.

Cutkosky:2002:OPM

[Cut02o] Steven Dale Cutkosky. 4. An overview of the proof of monomialization of morphisms from 3 folds to surfaces. Lecture Notes in Mathematics, 1786:14–18, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_4/.

Cutkosky:2002:N

[Cut02p] Steven Dale Cutkosky. 5. Notations. Lecture Notes in Mathematics, 1786:19, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-48030-3_5.

Cutkosky:2002:Ib

[Cut02q] Steven Dale Cutkosky. 6. The invariant ν. Lecture Notes in Mathematics, 1786:20–55, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-48030-3_6/.

Cutkosky:2002:IUQ

[Cut02r] Steven Dale Cutkosky. 7. The invariant ν under quadratic transforms. Lecture Notes in Mathematics, 1786:56–76, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-48030-3_
7/.

Cutkosky:2002:PMTa

[Cut02s] Steven Dale Cutkosky. 8. Permissible monoidal transforms centered at curves. Lecture Notes in Mathematics, 1786:77–92, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-48030-3_8/.

Cutkosky:2002:PSV

[Cut02t] Steven Dale Cutkosky. 9. Power series in 2 variables. Lecture Notes in Mathematics, 1786:93–108, 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-48030-3_9/.

Cutkosky:2002:MMF

[Cut02u] Steven Dale Cutkosky. Monomialization of Morphisms from 3folds to Surfaces, volume 1786 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43780-0 (print), 3-540-48030-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 235 pp. LCCN QA3 .L28 no. 1786. URL

http://link.springer.com/book/10.1007/b83848;   http:
//www.springerlink.com/content/978-3-540-48030-3.

DaPrato:2004:IMS

[Da 04] Giuseppe Da Prato. An introduction to Markov semigroups. Lecture Notes in Mathematics, 1855:1–63, 2004. CODEN LNMAA2. ISBN 3-540-23030-0 (print), 3-540-44653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44653-8_1/.

Daalhuis:2003:EA

[Daa03] Adri B. Olde Daalhuis. Exponential asymptotics. Lecture Notes in Mathematics, 1817:211–244, 2003. CODEN LNMAA2. ISBN 3-540-40375-2 (print), 3-540-44945-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44945-0_6/.

Dahmen:2003:MWM

[Dah03] Wolfgang Dahmen. Multiscale and wavelet methods for operator equations. Lecture Notes in Mathematics, 1825:31–96, 2003. CODEN LNMAA2. ISBN 3-540-20099-1 (print), 3-540-39810-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39810-3_2/.

Dalang:2003:LSE

[Dal03] Robert C. Dalang. Level sets and excursions of the Brownian sheet. Lecture Notes in Mathematics, 1802:167–208, 2003. CODEN LNMAA2. ISBN 3-540-00295-2 (print), 3-540-36259-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36259-3_5/.

Dalang:2009:SWE

[Dal09] Robert C. Dalang. The stochastic wave equation. Lecture Notes in Mathematics, 1962:39–71, 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-85994-9_2.pdf.

Dassios:2009:EMA

[Das09] George Dassios. Electric and magnetic activity of the brain in spherical and ellipsoidal geometry. Lecture Notes in Mathematics, 1983:133–202, 2009. CODEN LNMAA2. ISBN 3-642-03443-8 (print), 3-642-03444-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-03444-2_4.pdf.

Dorn:2008:AFS

[DBAP08] Oliver Dorn, Hugo Bertete-Aguirre, and George C. Papanicolaou. Adjoint fields and sensitivities for 3D electromagnetic imaging in isotropic and anisotropic media. Lecture Notes in Mathematics, 1943:35–65, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78547-7_3.pdf.

Deckelnick:2003:NAM

[DD03a] Klaus Deckelnick and Gerhard Dziuk. Numerical approximation of mean curvature flow of graphs and level sets. Lecture Notes in Mathematics, 1812:53–87, 2003. CODEN LNMAA2. ISBN 3-540-14033-6 (print), 3-540-39189-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39189-0_2/.

DelMoral:2003:CGI

[DD03b] Pierre Del Moral and Arnaud Doucet. On a class of genealogical and interacting Metropolis models. Lecture Notes in Mathematics, 1832:415–446, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40004-2_19/.

Delacherie:2002:CEA

[DDD02] C. Delacherie and C. Doleans-Dade. Un contre-exemple au problème des laplaciens approchés. (French) []. Lecture Notes in Mathematics, 1771:473–483, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_20/.

Doleans-Dade:2002:ISP

[DDM02] C. Doléans-Dade and P.-A. Meyer. Intégrales stochastiques par rapport aux martingales locales. (French) []. Lecture Notes in Mathematics, 1771:143–173, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_10/.

Debs:2009:PSR

[Deb09] Pierre Debs. Penalisation of the standard random walk by a function of the one-sided maximum, of the local time, or of the duration of the excursions. Lecture Notes in Mathematics, 1979:331– 363, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_12.pdf.

Deift:2003:FLR

[Dei03] Percy Deift. Four lectures on random matrix theory. Lecture Notes in Mathematics, 1815:21–52, 2003. CODEN LN-MAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44890-X_2/.

Dellacherie:2002:EAF

[Del02a] C. Dellacherie. Ensembles aléatoires I. (French) []. Lecture Notes in Mathematics, 1771:9–26, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_2/.

Dellacherie:2002:EAI

[Del02b] C. Dellacherie. Ensembles aléatoires II. (French) []. Lecture Notes in Mathematics, 1771:27–48, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_3/.

Dellacherie:2002:ISP

[Del02c] C. Dellacherie. Intégrals stochastiques par rapport aux processus de Wiener ou de Poisson. Lecture Notes in Mathematics, 1771: 415–416, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_17.

Dellacherie:2002:TFT

[Del02d] C. Dellacherie. Sur les théorémes fondamentaux de la théorie générale des processus. (French) []. Lecture Notes in Mathematics, 1771:75–84, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45530-1_5/.

Dellacherie:2002:EPM

[Del02e] C. Dellacherie. Un ensemble progressivement mesurable. Lecture Notes in Mathematics, 1771:85–87, 2002. CODEN LNMAA2.

ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45530-1_6/.

Delarue:2003:ESS

[Del03] François Delarue. Estimates of the solutions of a system of quasilinear PDEs. A probabilistic scheme. Lecture Notes in Mathematics, 1832:290–332, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40004-2_12/.

Delbaen:2006:SSS

[Del06] Freddy Delbaen. The structure of m-stable sets and in particular of the set of risk neutral measures. Lecture Notes in Mathematics, 1874:215–258, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35513-7_17.pdf.

Demkowicz:2008:PES

[Dem08] Leszek F. Demkowicz. Polynomial exact sequences and projection-based interpolation with application to Maxwell equations. Lecture Notes in Mathematics, 1939:101–158, 2008. CO-DEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78319-0_3.pdf.

Demni:2009:RDP

[Dem09] Nizar Demni. Radial Dunkl processes associated with dihedral systems. Lecture Notes in Mathematics, 1979:153–169, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_6.pdf.

Deuflhard:2000:DET

[Deu00] P. Deuflhard. Differential equations in technology and medicine: Computational concepts, adaptive algorithms, and virtual labs. Lecture Notes in Mathematics, 1739:69–125, 2000. CODEN LN-MAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103918/.

Dembo:2005:BM

[DF05a] Amir Dembo and Tadahisa Funaki. Back matter. Lecture Notes in Mathematics, 1869:??, 2005. CODEN LNMAA2. ISBN 3-540-26069-2 (print), 3-540-31537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-31537-7/1.pdf.

Dembo:2005:FPC

[DF05b] Amir Dembo and Tadahisa Funaki. Favorite points, cover times and fractals. Lecture Notes in Mathematics, 1869:1–101, 2005. CODEN LNMAA2. ISBN 3-540-26069-2 (print), 3-540-31537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11429579_1.pdf;   http://link.springer.com/content/
pdf/bfm:978-3-540-31537-7/1.pdf.

Dembo:2005:SIM

[DF05c] Amir Dembo and Tadahisa Funaki. Stochastic interface models. Lecture Notes in Mathematics, 1869:103–274, 2005. CODEN LNMAA2. ISBN 3-540-26069-2 (print), 3-540-31537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11429579_2.pdf.

Derezinski:2006:FGR

[DF06] Jan Dereziński and Rafal Früboes. Fermi Golden Rule and open quantum systems. Lecture Notes in Mathematics, 1882: 67–116, 2006. CODEN LNMAA2. ISBN 3-540-30993-4 (print), 3-540-33967-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-33967-1_2.pdf.

DiBartolo:2008:AGL

[DFPS08a] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Algebraic Groups and Lie Groups with Few Factors, volume 1944 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 207 pp. LCCN QA3 .L28 no. 1944. URL http://link.

springer.com/book/10.1007/978-3-540-78584-2;   http:/
/www.springerlink.com/content/978-3-540-78584-2.

DiBartolo:2008:BM

[DFPS08b] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Back matter. Lecture Notes in Mathematics, 1944: 199–206, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-78584-2/1.pdf.

DiBartolo:2008:C

[DFPS08c] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Chains. Lecture Notes in Mathematics, 1944:49–79, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78584-2_4.pdf.

DiBartolo:2008:E

[DFPS08d] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Extensions. Lecture Notes in Mathematics, 1944:11– 28, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78584-2_2.pdf.

DiBartolo:2008:FM

[DFPS08e] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Front matter. Lecture Notes in Mathematics, 1944: i–xvi, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-78584-2/1.pdf.

DiBartolo:2008:GEN

[DFPS08f] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Groups of extreme nilpotency class. Lecture Notes in Mathematics, 1944:29–47, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78584-2_3.pdf.

DiBartolo:2008:GFT

[DFPS08g] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Groups with few types of isogenous factors. Lecture Notes in Mathematics, 1944:81–147, 2008. CO-DEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78584-2_5.pdf.

DiBartolo:2008:NS

[DFPS08h] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Normality of subgroups. Lecture Notes in Mathematics, 1944:167–198, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78584-2_7.pdf.

DiBartolo:2008:P

[DFPS08i] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Prerequisites. Lecture Notes in Mathematics, 1944:1– 10, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78584-2_1.pdf.

DiBartolo:2008:TDA

[DFPS08j] Alfonso Di Bartolo, Giovanni Falcone, Peter Plaumann, and Karl Strambach. Three-dimensional affine groups. Lecture Notes in Mathematics, 1944:149–166, 2008. CODEN LNMAA2. ISBN 3-540-78583-3 (print), 3-540-78584-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78584-2_6.pdf.

DellAntonio:2008:BM

[DFT08a] Gianfausto Dell’Antonio, Rodolfo Figari, and Alessandro Teta. Back matter. Lecture Notes in Mathematics, 1943:191–192, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-78547-7/1.pdf.

DellAntonio:2008:BRP

[DFT08b] Gianfausto Dell’Antonio, Rodolfo Figari, and Alessandro Teta. A brief review on point interactions. Lecture Notes in Mathematics, 1943:171–189, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78547-7_7.pdf.

Degond:2008:QHD

[DGMR08] Pierre Degond, Samy Gallego, Florian Méhats, and Christian Ringhofer. Quantum hydrodynamic and diffusion models derived from the entropy principle. Lecture Notes in Mathematics, 1946:111–168, 2008. CODEN LNMAA2. ISBN 3-540-79573-1 (print), 3-540-79574-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-79574-2_3.pdf.

Denis:2000:FDE

[DGP00] Laurent Denis, Axel Grorud, and Monique Pontier. Formes de Dirichlet sur un espace de Wiener–Poisson. Application au grossissement de filtration. (French) []. Lecture Notes in Mathematics, 1729:198–217, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103804/.

Dermoune:2008:SGI

[DH08] Azzouz Dermoune and Philippe Heinrich. Spectral gap inequality for a colored disordered lattice gas. Lecture Notes in Mathematics, 1934:1–18, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-77913-1_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-77913-1/1.pdf.

denHollander:2009:BM

[dH09a] Frank den Hollander. Back matter. Lecture Notes in Mathematics, 1974:1–31, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-00333-2/1.pdf.

denHollander:2009:CP

[dH09b] Frank den Hollander. Charged polymers. Lecture Notes in Mathematics, 1974:115–127, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-00333-2_8.pdf.

denHollander:2009:CNL

[dH09c] Frank den Hollander. Copolymers near a linear selective interface. Lecture Notes in Mathematics, 1974:129–154, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00333-2_9.pdf.

denHollander:2009:CNR

[dH09d] Frank den Hollander. Copolymers near a random selective interface. Lecture Notes in Mathematics, 1974:155–179, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00333-2_10.pdf.

denHollander:2009:EP

[dH09e] Frank den Hollander. Elastic polymers. Lecture Notes in Mathematics, 1974:59–65, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00333-2_5.pdf.

denHollander:2009:FMa

[dH09f] Frank den Hollander. Front matter. Lecture Notes in Mathematics, 1974:17–18, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-00333-2/1/1.pdf.

denHollander:2009:FMb

[dH09g] Frank den Hollander. Front matter. Lecture Notes in Mathematics, 1974:113–114, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-00333-2/2/1.pdf.

denHollander:2009:I

[dH09h] Frank den Hollander. Introduction. Lecture Notes in Mathematics, 1974:1–7, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00333-2_1.pdf.

denHollander:2009:PA

[dH09i] Frank den Hollander. Polymer adsorption. Lecture Notes in Mathematics, 1974:85–112, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-00333-2_7.pdf.

denHollander:2009:PC

[dH09j] Frank den Hollander. Polymer collapse. Lecture Notes in Mathematics, 1974:67–84, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00333-2_6.pdf.

denHollander:2009:PRP

[dH09k] Frank den Hollander. Polymers in a random potential. Lecture Notes in Mathematics, 1974:205–231, 2009. CO-DEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00333-2_12.pdf.

denHollander:2009:RPW

[dH09l] Frank den Hollander. Random pinning and wetting of polymers. Lecture Notes in Mathematics, 1974:181–204, 2009. CO-DEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00333-2_11.pdf.

denHollander:2009:SPH

[dH09m] Frank den Hollander. Soft polymers in high dimension. Lecture Notes in Mathematics, 1974:41–58, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-00333-2_4.
pdf.

denHollander:2009:SPL

[dH09n] Frank den Hollander. Soft polymers in low dimension. Lecture Notes in Mathematics, 1974:19–39, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-00333-2_3.
pdf.

denHollander:2009:TLM

[dH09o] Frank den Hollander. Three lectures on metastability under stochastic dynamics. Lecture Notes in Mathematics, 1970:1–24, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-92796-9_5.pdf.

denHollander:2009:TBM

[dH09p] Frank den Hollander. Two basic models. Lecture Notes in Mathematics, 1974:9–16, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00333-2_2.pdf.

denHollander:2009:RPE

[dH09q] Frank (Wilhelmus Theodorus Franciscus) den Hollander, editor. Random Polymers: École d’Été de Probabilités de Saint-Flour XXXVII — 2007, volume 1974 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1974; QC173.4.P65 H65 2009; QC173.4.P65 R36 2009. URL http:

//link.springer.com/book/10.1007/978-3-642-00333-2;
http://www.springerlink.com/content/978-3-642-00333-2.

Deng:2009:BM

[DH09r] Donggao Deng and Yongsheng Han. Back matter. Lecture Notes in Mathematics, 1966:149–160, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-88745-4/1.pdf.

Deng:2009:BCZ

[DH09s] Donggao Deng and Yongsheng Han. The boundedness of Calderón–Zygmund operators on wavelet spaces. Lecture Notes in Mathematics, 1966:27–37, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-88745-4_3.pdf.

Deng:2009:CZO

[DH09t] Donggao Deng and Yongsheng Han. Calderón–Zygmund operator on space of homogeneous type. Lecture Notes in Mathematics, 1966:9–25, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-88745-4_2.pdf.

Deng:2009:FM

[DH09u] Donggao Deng and Yongsheng Han. Front matter. Lecture Notes in Mathematics, 1966:i–xii, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-88745-4/1.pdf.

Deng:2009:HAS

[DH09v] Donggao Deng and Yongsheng Han. Harmonic Analysis on Spaces of Homogeneous Type, volume 1966 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 154 pp. LCCN QA3 .L28 no.

1966. URL http://link.springer.com/book/10.1007/978-
3-540-88745-4;   http://www.springerlink.com/content/
978-3-540-88745-4.

Deng:2009:I

[DH09w] Donggao Deng and Yongsheng Han. Introduction. Lecture Notes in Mathematics, 1966:1–7, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-88745-4_1.pdf.

Deng:2009:LPA

[DH09x] Donggao Deng and Yongsheng Han. Littlewood–Paley analysis on non homogeneous spaces. Lecture Notes in Mathematics, 1966:137–147, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-88745-4_6.pdf.

Deng:2009:WES

[DH09y] Donggao Deng and Yongsheng Han. Wavelet expansions on spaces of homogeneous type. Lecture Notes in Mathematics, 1966:39–90, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-88745-4_4.pdf.

Deng:2009:WSF

[DH09z] Donggao Deng and Yongsheng Han. Wavelets and spaces of functions and distributions. Lecture Notes in Mathematics, 1966:91– 136, 2009. CODEN LNMAA2. ISBN 3-540-88744-X (print), 3-540-88745-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-88745-4_5.pdf.

Dhahri:2009:MPS

[Dha09] Ameur Dhahri. Markovian properties of the spin-boson model. Lecture Notes in Mathematics, 1979:397–432, 2009. CO-DEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_15.pdf.

Degtyarev:2000:AS

[DIK00a] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Algebraic surfaces. Lecture Notes in Mathematics, 1746:53– 78, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103964/.

Degtyarev:2000:DTE

[DIK00b] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Deformation types: the elliptic and parabolic cases. Lecture Notes in Mathematics, 1746:169–190, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103970/.

Degtyarev:2000:DTH

[DIK00c] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Deformation types: the hyperbolic and parabolic cases. Lecture Notes in Mathematics, 1746:145–168, 2000. CODEN LN-MAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103969/.

Degtyarev:2000:ILQ

[DIK00d] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Integral lattices and quadratic forms. Lecture Notes in Mathematics, 1746:29–52, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103963/.

Degtyarev:2000:I

[DIK00e] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Introduction. Lecture Notes in Mathematics, 1746:vii–xiii, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103961.

Degtyarev:2000:MRE

[DIK00f] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Moduli of real Enriques surfaces. Lecture Notes in Mathematics, 1746:127–144, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103968/.

Degtyarev:2000:RES

[DIK00g] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Real Enriques Surfaces, volume 1746 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 259 pp. LCCN QA3 .L28 no. 1746.

URL http://link.springer.com/book/10.1007/BFb0103960;
 http://www.springerlink.com/content/978-3-540-39948-
3.

Degtyarev:2000:RST

[DIK00h] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Real surfaces: the topological aspects. Lecture Notes in Mathematics, 1746:79–87, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103965/.

Degtyarev:2000:SDC

[DIK00i] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Summary: Deformation classes. Lecture Notes in Mathematics, 1746:88–96, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103966/.

Degtyarev:2000:TI

[DIK00j] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Topology of involutions. Lecture Notes in Mathematics, 1746: 1–28, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103962/.

Degtyarev:2000:TRE

[DIK00k] Alexander Degtyarev, Ilia Itenberg, and Viatcheslav Kharlamov. Topology of real Enriques surfaces. Lecture Notes in Mathematics, 1746:97–126, 2000. CODEN LNMAA2. ISBN 3-540-41088-0 (print), 3-540-39948-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103967/.

Dalang:2009:FM

[DKM^{+}09] Robert Dalang, Davar Khoshnevisan, Carl Mueller, David Nualart, and Yimin Xiao. Front matter. Lecture Notes in Mathematics, 1962:I–XI, 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-85994-9/1.pdf.

Doplicher:2004:NGL

[DL04] Sergio Doplicher and Roberto Longo, editors. Noncommutative Geometry: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy, September 3-9, 2000, volume 1831 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1831. URL http:/

/link.springer.com/book/10.1007/b94118;   http://www.
springerlink.com/content/978-3-540-39702-1.

DelMoral:2000:BIP

[DM00] P. Del Moral and L. Miclo. Branching and interacting particle systems approximations of Feynman–Kac formulae with applications to non-linear filtering. Lecture Notes in Mathematics, 1729: 1–145, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/chapter/10.
1007/BFb0103798/;   http://link.springer.com/content/
pdf/bfm:978-3-540-46413-6/1.pdf.

Dermoune:2003:GVP

[DM03] Azzouz Dermoune and Octave Moutsinga. Generalized variational principles. Lecture Notes in Mathematics, 1801:183– 193, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_4/.

Dacorogna:2008:CVN

[DM08] Bernard Dacorogna and Paolo Marcellini. Calculus of Variations and Nonlinear Partial Differential Equations, volume 1927 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 196 pp. LCCN QA3 .L28 no. 1927. URL http://link.

springer.com/book/10.1007/978-3-540-75914-0;   http:/
/www.springerlink.com/content/978-3-540-75914-0. With

a historical overview by Elvira Mascolo.

Donati-Martin:2007:SPX

[DMÉRS07] Catherine Donati-Martin, Michel Émery, Alain Rouault, and Christophe Stricker, editors. Séminaire de Probabilités XL, volume 1899 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1899. URL http://link.

springer.com/book/10.1007/978-3-540-71189-6;   http:/
/www.springerlink.com/content/978-3-540-71189-6.

Donati-Martin:2008:SPX

[DMÉRS08] Catherine Donati-Martin, Michel Émery, Alain Rouault, and Christophe Stricker, editors. Séminaire de Probabilités XLI, volume 1934 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1934. URL http://link.

springer.com/book/10.1007/978-3-540-77913-1;   http:/
/www.springerlink.com/content/978-3-540-77913-1.

Donati-Martin:2009:SPX

[DMÉRS09] Catherine Donati-Martin, Michel Émery, Alain Rouault, and Christophe Stricker, editors. Séminaire de Probabilités XLII. (French) [Seminar on Probabilities XLII], volume 1979 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (paperback), 3-642-01763-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1979; QA3 .L28 no.1979; QA273.A1 S46 2009.

URL http://link.springer.com/book/10.1007/978-3-642-
01763-6;   http://www.springerlink.com/content/978-3-
642-01763-6;   http://www.springerlink.com/openurl.asp?
genre=book&isbn=978-3-642-01762-9.

Donati-Martin:2003:PWM

[DMH03] Catherine Donati-Martin and Yueyun Hu. Penalization of the Wiener measure and principal values. Lecture Notes in Mathematics, 1801:251–269, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36107-7_9/.

Dohmen:2003:IO

[Doh03a] Klaus Dohmen. 1. Introduction and overview. Lecture Notes in Mathematics, 1826:1–4, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39399-3_1/.

Dohmen:2003:P

[Doh03b] Klaus Dohmen. 2. Preliminaries. Lecture Notes in Mathematics, 1826:5–8, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39399-3_2/.

Dohmen:2003:BIA

[Doh03c] Klaus Dohmen. 3. Bonferroni inequalities via abstract tubes. Lecture Notes in Mathematics, 1826:9–18, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39399-3_3/.

Dohmen:2003:ATC

[Doh03d] Klaus Dohmen. 4. Abstract tubes via closure and kernel operators. Lecture Notes in Mathematics, 1826:19–43, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39399-3_
4/.

Dohmen:2003:RS

[Doh03e] Klaus Dohmen. 5. Recursive schemes. Lecture Notes in Mathematics, 1826:44–46, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39399-3_5/.

Dohmen:2003:RA

[Doh03f] Klaus Dohmen. 6. Reliability applications. Lecture Notes in Mathematics, 1826:47–81, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39399-3_6/.

Dohmen:2003:CAR

[Doh03g] Klaus Dohmen. 7. Combinatorial applications and related topics. Lecture Notes in Mathematics, 1826:82–99, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39399-3_
7/.

Dohmen:2003:B

[Doh03h] Klaus Dohmen. Bibliography. Lecture Notes in Mathematics, 1826:100–109, 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39399-3_8/.

Dohmen:2003:IBI

[Doh03i] Klaus Dohmen. Improved Bonferroni Inequalities via Abstract Tubes: Inequalities and Identities of Inclusion–Exclusion Type, volume 1826 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20025-8 (print), 3-540-39399-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 109 pp. LCCN QA3 .L28 no. 1826; QA3 .L28 no. 1826 QA295; QA3.L471 no. 1826. URL http:/

/link.springer.com/book/10.1007/b13785;   http://www.
springerlink.com/content/978-3-540-39399-3.

Doig:2008:AAC

[Doi08a] A. J. Doig. Amino acid clustering. Lecture Notes in Mathematics, 1953:139–151, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69393-2_7.pdf.

Doig:2008:CAS

[Doi08b] A. J. Doig. Cryptographic attacks and signal clustering. Lecture Notes in Mathematics, 1953:153–159, 2008. CO-DEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69393-2_8.pdf.

Dolzmann:2003:Ia

[Dol03a] Georg Dolzmann. 1. Introduction. Lecture Notes in Mathematics, 1803:1–10, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36125-1_1/.

Dolzmann:2003:SHC

[Dol03b] Georg Dolzmann. 2. Semiconvex hulls of compact sets. Lecture Notes in Mathematics, 1803:11–68, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36125-1_2/.

Dolzmann:2003:MEN

[Dol03c] Georg Dolzmann. 3. Macroscopic energy for nematic elastomers. Lecture Notes in Mathematics, 1803:69–81, 2003. CODEN LN-MAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36125-1_
3/.

Dolzmann:2003:USM

[Dol03d] Georg Dolzmann. 4. Uniqueness and stability of microstructure. Lecture Notes in Mathematics, 1803:83–126, 2003. CODEN LN-MAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36125-1_
4/.

Dolzmann:2003:AMT

[Dol03e] Georg Dolzmann. 5. Applications to martensic transformations. Lecture Notes in Mathematics, 1803:127–152, 2003. CODEN LN-MAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36125-1_
5/.

Dolzmann:2003:AA

[Dol03f] Georg Dolzmann. 6. Algorithmic aspects. Lecture Notes in Mathematics, 1803:153–175, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36125-1_6/.

Dolzmann:2003:BR

[Dol03g] Georg Dolzmann. 7. Bibliographic remarks. Lecture Notes in Mathematics, 1803:177–182, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36125-1_7/.

Dolzmann:2003:CCR

[Dol03h] Georg Dolzmann. A. convexity conditions and rank-one connections. Lecture Notes in Mathematics, 1803:183–192, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36125-1_8/.

Dolzmann:2003:BEC

[Dol03i] Georg Dolzmann. B. elements of crystallography. Lecture Notes in Mathematics, 1803:193–196, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36125-1_9/.

Dolzmann:2003:CN

[Dol03j] Georg Dolzmann. C. notation. Lecture Notes in Mathematics, 1803:197–200, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36125-1_10/.

Dolzmann:2003:Ib

[Dol03k] Georg Dolzmann. Index. Lecture Notes in Mathematics, 1803: 211–212, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36125-1_12.

Dolzmann:2003:R

[Dol03l] Georg Dolzmann. References. Lecture Notes in Mathematics, 1803:201–209, 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36125-1_11/.

Dolzmann:2003:VMC

[Dol03m] Georg Dolzmann. Variational Methods for Crystalline Microstructure-Analysis and Computation, volume 1803 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00114-X (print), 3-540-36125-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 212 pp. LCCN QA3 .L28 no. 1803. URL http:/

/link.springer.com/book/10.1007/b10191;   http://www.
springerlink.com/content/978-3-540-36125-1.

Doney:2005:SEC

[Don05a] Ronald A. Doney. Some excursion calculations for spectrally onesided Lévy processes. Lecture Notes in Mathematics, 1857:5–15, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_2.pdf.

Doney:2005:TCR

[Don05b] Ronald A. Doney. Tanaka’s construction for random walks and Lévy processes. Lecture Notes in Mathematics, 1857:1–4, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_1.pdf.

Doney:2007:BM

[Don07a] Professor Ronald A. Doney. Back matter. Lecture Notes in Mathematics, 1897:133–150, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-48511-7/1.pdf.

Doney:2007:CRQ

[Don07b] Professor Ronald A. Doney. Creeping and related questions. Lecture Notes in Mathematics, 1897:51–64, 2007. CO-DEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48511-7_6.pdf.

Doney:2007:FM

[Don07c] Professor Ronald A. Doney. Front matter. Lecture Notes in Mathematics, 1897:I–IX, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-48511-7/1.pdf.

Doney:2007:FWH

[Don07d] Professor Ronald A. Doney. Further Wiener–Hopf developments. Lecture Notes in Mathematics, 1897:41–50, 2007. CO-DEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48511-7_5.pdf.

Doney:2007:ILP

[Don07e] Professor Ronald A. Doney. Introduction to Lévy processes. Lecture Notes in Mathematics, 1897:1–8, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-48511-7_1.
pdf.

Doney:2007:LPW

[Don07f] Professor Ronald A. Doney. Ladder processes and the Wiener– Hopf factorisation. Lecture Notes in Mathematics, 1897:25–40, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48511-7_4.pdf.

Doney:2007:LPC

[Don07g] Professor Ronald A. Doney. Lévy processes conditioned to stay positive. Lecture Notes in Mathematics, 1897:81–93, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48511-7_8.pdf.

Doney:2007:LTE

[Don07h] Professor Ronald A. Doney. Local times and excursions. Lecture Notes in Mathematics, 1897:19–24, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-48511-7_3.
pdf.

Doney:2007:STB

[Don07i] Professor Ronald A. Doney. Small-time behaviour. Lecture Notes in Mathematics, 1897:115–132, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48511-7_10.pdf.

Doney:2007:SNL

[Don07j] Professor Ronald A. Doney. Spectrally negative Lévy processes. Lecture Notes in Mathematics, 1897:95–113, 2007. CO-DEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48511-7_9.pdf.

Doney:2007:SC

[Don07k] Professor Ronald A. Doney. Spitzer’s condition. Lecture Notes in Mathematics, 1897:65–80, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48511-7_7.pdf.

Doney:2007:S

[Don07l] Professor Ronald A. Doney. Subordinators. Lecture Notes in Mathematics, 1897:9–17, 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48511-7_2.pdf.

Doney:2007:FTL

[Don07m] Ronald A. Doney, editor. Fluctuation Theory for Lévy Processes: École d’Été de Probabilités de Saint-Flour XXXV — 2005, volume 1897 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-48510-4 (print), 3-540-48511-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1897;

QA274.73 .D66 2007. URL http://link.springer.com/
book/10.1007/978-3-540-48511-7;   http://www.loc.gov/
catdir/enhancements/fy0825/2007921692-d.html;   http:
//www.loc.gov/catdir/toc/fy0716/2007921692.html;   http:
//www.springerlink.com/content/978-3-540-48511-7.

Dorn:2008:TRA

[Dor08] Oliver Dorn. Time-reversal and the adjoint imaging method with an application in telecommunication. Lecture Notes in Mathematics, 1943:135–170, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78547-7_6.pdf.

Doumerc:2003:NRE

[Dou03] Yan Doumerc. A note on representations of eigenvalues of classical Gaussian matrices. Lecture Notes in Mathematics, 1832: 370–384, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_15/.

Dozzi:2003:ODS

[Doz03] Marco Dozzi. Occupation density and sample path properties. Lecture Notes in Mathematics, 1802:127–166, 2003. CODEN LNMAA2. ISBN 3-540-00295-2 (print), 3-540-36259-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36259-3_
4/.

DeDonno:2006:SIR

[DP06] Marzia De Donno and Maurizio Pratelli. Stochastic integration with respect to a sequence of semimartingales. Lecture Notes in Mathematics, 1874:119–135, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35513-7_10.pdf.

deReyna:2002:HLM

[dR02a] Juan Arias de Reyna. 1. Hardy–Littlewood maximal function. Lecture Notes in Mathematics, 1785:3–10, 2002. CODEN LN-MAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45822-7_
1/.

deReyna:2002:AT

[dR02b] Juan Arias de Reyna. 10. All together. Lecture Notes in Mathematics, 1785:117–123, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_11/.

deReyna:2002:SF

[dR02c] Juan Arias de Reyna. 11. Spaces of functions. Lecture Notes in Mathematics, 1785:127–143, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_12/.

deReyna:2002:MOF

[dR02d] Juan Arias de Reyna. 12. The maximal operator of Fourier series. Lecture Notes in Mathematics, 1785:145–162, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45822-7_
13/.

deReyna:2002:FTL

[dR02e] Juan Arias de Reyna. 13. Fourier transform on the line. Lecture Notes in Mathematics, 1785:163–166, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45822-7_14/.

deReyna:2002:FS

[dR02f] Juan Arias de Reyna. 2. Fourier series. Lecture Notes in Mathematics, 1785:11–29, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45822-7_2/.

deReyna:2002:HT

[dR02g] Juan Arias de Reyna. 3. Hilbert transform. Lecture Notes in Mathematics, 1785:31–44, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_3/.

deReyna:2002:BS

[dR02h] Juan Arias de Reyna. 4. The basic step. Lecture Notes in Mathematics, 1785:51–72, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45822-7_5/.

deReyna:2002:MI

[dR02i] Juan Arias de Reyna. 5. Maximal inequalities. Lecture Notes in Mathematics, 1785:73–76, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_6/.

deReyna:2002:GPS

[dR02j] Juan Arias de Reyna. 6. Growth of partial sums. Lecture Notes in Mathematics, 1785:77–84, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_7/.

deReyna:2002:CAF

[dR02k] Juan Arias de Reyna. 7. Carleson analysis of the function. Lecture Notes in Mathematics, 1785:85–91, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45822-7_8/.

deReyna:2002:AP

[dR02l] Juan Arias de Reyna. 8. Allowed pairs. Lecture Notes in Mathematics, 1785:93–102, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_9/.

deReyna:2002:PIT

[dR02m] Juan Arias de Reyna. 9. Pair interchange theorems. Lecture Notes in Mathematics, 1785:103–115, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_10/.

deReyna:2002:CHT

[dR02n] Juan Arias de Reyna. The Charleson–Hunt Theorem. Lecture Notes in Mathematics, 1785:47–49, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45822-7_4/.

deReyna:2002:CSI

[dR02o] Juan Arias de Reyna. Comments and subject index. Lecture Notes in Mathematics, 1785:171–175, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45822-7_16/.

deReyna:2002:PCF

[dR02p] Juan Arias de Reyna. Pointwise Convergence of Fourier Series, volume 1785 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 175 pp. LCCN QA3 .L28 no. 1785. URL http:/

/link.springer.com/book/10.1007/b83346;   http://www.
springerlink.com/content/978-3-540-45822-7.

deReyna:2002:R

[dR02q] Juan Arias de Reyna. References. Lecture Notes in Mathematics, 1785:167–169, 2002. CODEN LNMAA2. ISBN 3-540-43270-1 (print), 3-540-45822-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45822-7_15/.

DaPrato:2008:SHR

[DR08] Giuseppe Da Prato and Michael Rückner, editors. SPDE in Hydrodynamic: Recent Progress and Prospects: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy August 29– September 3, 2005, volume 1942 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78492-6 (print), 3-540-78493-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1942. URL http://link.

springer.com/book/10.1007/978-3-540-78493-7;   http:/
/www.springerlink.com/content/978-3-540-78493-7.

Arias-de-Reyna:2000:UCM

[dRV00] J. Arias de Reyna and R. Villa. The uniform concentration of measure phenomenon in ℓ_{p}^{n} (1 ≤ p ≤ 2). Lecture Notes in Mathematics, 1745:13–18, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0107203/.

daSilva:2001:ACS

[dS01a] Ana Cannas da Silva. Almost complex structures. Lecture Notes in Mathematics, 1764:83–88, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_12/.

daSilva:2001:CKM

[dS01b] Ana Cannas da Silva. Compact Kähler manifolds. Lecture Notes in Mathematics, 1764:117–123, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_17/.

daSilva:2001:CT

[dS01c] Ana Cannas da Silva. Compatible triples. Lecture Notes in Mathematics, 1764:89–92, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_13/.

daSilva:2001:CM

[dS01d] Ana Cannas da Silva. Complex manifolds. Lecture Notes in Mathematics, 1764:101–107, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_15/.

daSilva:2001:CD

[dS01e] Ana Cannas da Silva. Contact dynamics. Lecture Notes in Mathematics, 1764:75–79, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_11/.

daSilva:2001:CF

[dS01f] Ana Cannas da Silva. Contact forms. Lecture Notes in Mathematics, 1764:69–74, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_10/.

daSilva:2001:DMW

[dS01g] Ana Cannas da Silva. Darboux–Moser–Weinstein theory. Lecture Notes in Mathematics, 1764:55–60, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45330-7_8/.

daSilva:2001:DT

[dS01h] Ana Cannas da Silva. Dolbeault theory. Lecture Notes in Mathematics, 1764:93–98, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_14/.

daSilva:2001:FMa

[dS01i] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:i–xiv, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/1.pdf.

daSilva:2001:FMb

[dS01j] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:2, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/1/1.pdf.

daSilva:2001:FMc

[dS01k] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:16, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/2/1.pdf.

daSilva:2001:FMd

[dS01l] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:40, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/3/1.pdf.

daSilva:2001:FMe

[dS01m] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:68, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/4/1.pdf.

daSilva:2001:FMf

[dS01n] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:82, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/5/1.pdf.

daSilva:2001:FMg

[dS01o] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:100, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/6/1.pdf.

daSilva:2001:FMh

[dS01p] Ana Cannas da Silva. Front matter. Lecture Notes in Mathematics, 1764:126, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45330-7/7/1.pdf.

daSilva:2001:GF

[dS01q] Ana Cannas da Silva. Generating functions. Lecture Notes in Mathematics, 1764:25–31, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_4/.

daSilva:2001:HVF

[dS01r] Ana Cannas da Silva. Hamiltonian vector fields. Lecture Notes in Mathematics, 1764:127–134, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_18/.

daSilva:2001:KF

[dS01s] Ana Cannas da Silva. Kähler forms. Lecture Notes in Mathematics, 1764:109–116, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_16/.

daSilva:2001:LS

[dS01t] Ana Cannas da Silva. Lagrangian submanifolds. Lecture Notes in Mathematics, 1764:17–23, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_3/.

daSilva:2001:LSG

[dS01u] Ana Cannas da Silva. Lectures on Symplectic Geometry, volume 1764 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 217 pp. LCCN QA3 .L28 no. 1764. URL http://link.

springer.com/book/10.1007/978-3-540-45330-7;   http:/
/www.springerlink.com/content/978-3-540-45330-7.

daSilva:2001:LT

[dS01v] Ana Cannas da Silva. Legendre transform. Lecture Notes in Mathematics, 1764:143–148, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_20/.

daSilva:2001:MT

[dS01w] Ana Cannas da Silva. Moser theorems. Lecture Notes in Mathematics, 1764:49–53, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_7/.

daSilva:2001:PLT

[dS01x] Ana Cannas da Silva. Preparation for the local theory. Lecture Notes in Mathematics, 1764:41–47, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45330-7_6/.

daSilva:2001:R

[dS01y] Ana Cannas da Silva. Recurrence. Lecture Notes in Mathematics, 1764:33–37, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45330-7_5/.

daSilva:2001:SFC

[dS01z] Ana Cannas da Silva. Symplectic form on the cotangent bundle. Lecture Notes in Mathematics, 1764:9–14, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45330-7_
2/.

daSilva:2001:VP

[dS01-27] Ana Cannas da Silva. Variational principles. Lecture Notes in Mathematics, 1764:135–142, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45330-7_19/.

daSilva:2001:WTN

[dS01-28] Ana Cannas da Silva. Weinstein tubular neighborhood theorem. Lecture Notes in Mathematics, 1764:61–66, 2001. CODEN LNMAA2. ISBN 3-540-42195-5 (print), 3-540-45330-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45330-7_
9/.

duSautoy:2008:BM

[dSW08a] Marcus du Sautoy and Luke Woodward. Back matter. Lecture Notes in Mathematics, 1925:179–212, 2008. CODEN LNMAA2.

ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-540-74776-5/1.pdf.

duSautoy:2008:FM

[dSW08b] Marcus du Sautoy and Luke Woodward. Front matter. Lecture Notes in Mathematics, 1925:I–XII, 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-74776-5/1.pdf.

duSautoy:2008:I

[dSW08c] Marcus du Sautoy and Luke Woodward. Introduction. Lecture Notes in Mathematics, 1925:1–20, 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74776-5_1.
pdf.

duSautoy:2008:LFE

[dSW08d] Marcus du Sautoy and Luke Woodward. Local functional equations. Lecture Notes in Mathematics, 1925:83–119, 2008. CO-DEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74776-5_4.pdf.

duSautoy:2008:NBT

[dSW08e] Marcus du Sautoy and Luke Woodward. Natural boundaries I: Theory. Lecture Notes in Mathematics, 1925:121–153, 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74776-5_5.pdf.

duSautoy:2008:NBIa

[dSW08f] Marcus du Sautoy and Luke Woodward. Natural boundaries II: Algebraic groups. Lecture Notes in Mathematics, 1925:155–167, 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74776-5_6.pdf.

duSautoy:2008:NBIb

[dSW08g] Marcus du Sautoy and Luke Woodward. Natural boundaries III: Nilpotent groups. Lecture Notes in Mathematics, 1925:169–177, 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74776-5_7.pdf.

duSautoy:2008:NGE

[dSW08h] Marcus du Sautoy and Luke Woodward. Nilpotent groups: Explicit examples. Lecture Notes in Mathematics, 1925:21–68, 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74776-5_2.pdf.

duSautoy:2008:SLR

[dSW08i] Marcus du Sautoy and Luke Woodward. Soluble Lie rings. Lecture Notes in Mathematics, 1925:69–82, 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74776-5_3.
pdf.

duSautoy:2008:ZFG

[dSW08j] Marcus du Sautoy and Luke Woodward. Zeta Functions of Groups and Rings, volume 1925 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-74701-X (print), 3-540-74776-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 208 pp. LCCN QA3 .L28 no.

1925. URL http://link.springer.com/book/10.1007/978-
3-540-74776-5;   http://www.springerlink.com/content/
978-3-540-74776-5.

Duran:2008:MFE

[Dur08] Ricardo G. Durán. Mixed finite element methods. Lecture Notes in Mathematics, 1939:1–44, 2008. CODEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78319-0_1.pdf.

Earn:2008:LIM

[Ear08] David J. D. Earn. A light introduction to modelling recurrent epidemics. Lecture Notes in Mathematics, 1945:3–17, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-78911-6_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-78911-6/1.pdf.

Erraoui:2009:CRG

[EE09] M. Erraoui and E. H. Essaky. Canonical representation for Gaussian processes. Lecture Notes in Mathematics, 1979:365–381, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_13.pdf.

Enriquez:2001:CLE

[EFL01] Nathanaël Enriquez, Jacques Franchi, and Yves Le Jan. Canonical lift and exit law of the fundamental diffusion associated with a Kleinian group. Lecture Notes in Mathematics, 1755:206– 219, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44671-2_16/.

Efstathiou:2005:FHS

[Efs05a] Konstantinos Efstathiou. 1. Four Hamiltonian systems. Lecture Notes in Mathematics, 1864:9–33, 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-31550-6_2.
pdf.

Efstathiou:2005:SVT

[Efs05b] Konstantinos Efstathiou. 2. Small vibrations of tetrahedral molecules. Lecture Notes in Mathematics, 1864:35–58, 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31550-6_3.pdf.

Efstathiou:2005:HAC

[Efs05c] Konstantinos Efstathiou. 3. The hydrogen atom in crossed fields. Lecture Notes in Mathematics, 1864:59–85, 2005. CO-DEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31550-6_4.pdf.

Efstathiou:2005:QSP

[Efs05d] Konstantinos Efstathiou. 4. Quadratic spherical pendula. Lecture Notes in Mathematics, 1864:87–111, 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-31550-6_5.
pdf.

Efstathiou:2005:FMR

[Efs05e] Konstantinos Efstathiou. 5. Fractional monodromy in the 1: −2 resonance system. Lecture Notes in Mathematics, 1864:113–127, 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31550-6_6.pdf.

Efstathiou:2005:A

[Efs05f] Konstantinos Efstathiou. Appendix. Lecture Notes in Mathematics, 1864:129–138, 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31550-6_7.pdf.

Efstathiou:2005:I

[Efs05g] Konstantinos Efstathiou. Introduction. Lecture Notes in Mathematics, 1864:1–8, 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31550-6_1.pdf.

Efstathiou:2005:MHS

[Efs05h] Konstantinos Efstathiou. Metamorphoses of Hamiltonian Systems with Symmetries, volume 1864 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 149 pp. LCCN QA3 .L28 no. 1864.

URL http://link.springer.com/book/10.1007/b105138;
http://www.springerlink.com/content/978-3-540-31550-6.

Efstathiou:2005:RI

[Efs05i] Konstantinos Efstathiou. References and index. Lecture Notes in Mathematics, 1864:139–149, 2005. CODEN LNMAA2. ISBN 3-540-24316-X (print), 3-540-31550-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31550-6_8.pdf.

Efendiev:2008:BM

[EH08a] Yalchin Efendiev and Thomas Yizhao Hou. Back matter. Lecture Notes in Mathematics, 1946:249–251, 2008. CODEN LNMAA2. ISBN 3-540-79573-1 (print), 3-540-79574-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-540-79574-2/1.pdf.

Efendiev:2008:MCF

[EH08b] Yalchin Efendiev and Thomas Yizhao Hou. Multiscale computations for flow and transport in heterogeneous media. Lecture Notes in Mathematics, 1946:169–248, 2008. CODEN LNMAA2. ISBN 3-540-79573-1 (print), 3-540-79574-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-79574-2_4.pdf.

Eisenbaum:2000:EIB

[Eis00] Nathalie Eisenbaum. Exponential inequalities for Bessel processes. Lecture Notes in Mathematics, 1729:146–150, 2000. CO-DEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103799/.

Eisenbaum:2003:GSC

[Eis03] Nathalie Eisenbaum. A Gaussian sheet connected to symmetric Markov chains. Lecture Notes in Mathematics, 1801:331– 334, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_13/.

Eisenbaum:2007:LTS

[Eis07] Nathalie Eisenbaum. Local time-space calculus for reversible semimartingales. Lecture Notes in Mathematics, 1899:137–146, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_6.pdf.

Espedal:2000:NSR

[EK00] Magne S. Espedal and Kenneth Hvistendahl Karlsen. Numerical solution of reservoir flow models based on large time step operator splitting algorithms. Lecture Notes in Mathematics, 1734: 9–77, 2000. CODEN LNMAA2. ISBN 3-540-67868-9 (print), 3-540-44656-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103975/.

Eliasson:2002:DSS

[EKMY02a] Hakan Eliasson, Sergei Kuksin, Stefano Marmi, and Jean-Christophe Yoccoz, editors. Dynamical Systems and Small Divisors: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, June 13-20, 1998, volume 1784 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43726-6 (print), 3-540-47928-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1784. URL

http://link.springer.com/book/10.1007/b83847;   http:
//www.springerlink.com/content/978-3-540-47928-4.

Eliasson:2002:FM

[EKMY02b] Hakan Eliasson, Sergei Kuksin, Stefano Marmi, and Jean-Christophe Yoccoz. Front matter. Lecture Notes in Mathematics, 1784:I–XIII, 2002. CODEN LNMAA2. ISBN 3-540-43726-6 (print), 3-540-47928-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-47928-4/1.pdf.

Eliasson:2002:PLQ

[Eli02] L. Hakan Eliasson. Perturbations of linear quasi-periodic system. Lecture Notes in Mathematics, 1784:1–60, 2002. CODEN LNMAA2. ISBN 3-540-43726-6 (print), 3-540-47928-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-47928-4_
1/.

Emery:2005:SPX

[ÉLY05] Michel Émery, Michel Ledoux, and Marc Yor, editors. Séminaire de Probabilités XXXVIII, volume 1857 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1857; QA273.A1 S46 2005; QA3 .L28 no.1857; QA3 .L28 2005. URL http://

link.springer.com/book/10.1007/b104072;   http://www.
springerlink.com/content/978-3-540-31449-3.

Erickson:2005:GOU

[EM05] K. Bruce Erickson and Ross A. Maller. Generalised Ornstein– Uhlenbeck processes and the convergence of Lévy integrals. Lecture Notes in Mathematics, 1857:70–94, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-31449-3_6.
pdf.

Emery:2000:CMF

[Eme00a] Michel Emery. Connexions et martingales. (French) []. Lecture Notes in Mathematics, 1738:38–51, 2000. CODEN LN-MAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0106707/.

Emery:2000:FCC

[Eme00b] Michel Emery. Fonctions convexes et comportement des martingales. (French) []. Lecture Notes in Mathematics, 1738:52– 72, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106708/.

Emery:2000:I

[Eme00c] Michel Emery. Introduction. Lecture Notes in Mathematics, 1738: 3–4, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106704.

Emery:2000:MBA

[Eme00d] Michel Emery. Mouvements browniens et applications harmoniques. (French) []. Lecture Notes in Mathematics, 1738:73– 84, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106709/.

Emery:2000:SVG

[Eme00e] Michel Emery. Semimartingales dans une variété et géométrie d’ordre 2. (French) []. Lecture Notes in Mathematics, 1738:22– 37, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106706/.

Emery:2000:VVC

[Eme00f] Michel Emery. Variétés, vecteurs, covecteurs, diffuseurs, codiffuseurs. (French) []. Lecture Notes in Mathematics, 1738:5– 21, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106705/.

Emery:2001:DAC

[Éme01] M. Émery. A discrete approach to the chaotic representation property. Lecture Notes in Mathematics, 1755:123–138, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44671-2_7/.

Emery:2006:SFL

[Éme06] Michel Émery. Sandwiched filtrations and Lévy processes. Lecture Notes in Mathematics, 1874:197–208, 2006. CO-DEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_15.pdf.

Emery:2009:RWF

[Éme09] Michel Émery. Recognising whether a filtration is Brownian: a case study. Lecture Notes in Mathematics, 1979:383–396, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_14.pdf.

Engl:2000:IPT

[Eng00] Heinz W. Engl. Inverse problems and their regularization. Lecture Notes in Mathematics, 1739:127–150, 2000. CODEN LN-MAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103919/.

Enriquez:2007:CPC

[Enr07] Nathanaül Enriquez. Correlated processes and the composition of generators. Lecture Notes in Mathematics, 1899:329–342, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_17.pdf.

Estrade:2001:BSD

[EP01] Anne Estrade and Monique Pontier. Backward stochastic differential equations in a Lie group. Lecture Notes in Mathematics, 1755:241–259, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44671-2_18/.

Ermentrout:2005:NO

[Erm05] Bard Ermentrout. Neural oscillators. Lecture Notes in Mathematics, 1860:69–106, 2005. CODEN LNMAA2. ISBN 3-540-23858-1 (print), 3-540-31544-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31544-5_3.pdf.

Emery:2001:VSC

[ÉS01] Michel Émery and Walter Schachermayer. On Vershik’s standardness criterion and Tsirelson’s notion of cosiness. Lecture Notes in Mathematics, 1755:265–305, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44671-2_20/.

Eisenbaum:2005:SFB

[ET05] Nathalie Eisenbaum and Ciprian A. Tudor. On squared fractional Brownian motions. Lecture Notes in Mathematics, 1857:282–289, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_19.pdf.

Ekeland:2007:OBP

[ET07] Ivar Ekeland and Erik Taflin. Optimal bond portfolios. Lecture Notes in Mathematics, 1919:51–102, 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-73327-0_2.
pdf.

Elworthy:2007:GIF

[ETZ07] K. David Elworthy, Aubrey Truman, and Huaizhong Zhao. Generalized Itô formulae and space-time Lebesgue–Stieltjes integrals of local times. Lecture Notes in Mathematics, 1899:117–136, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_5.pdf.

Evans:2008:WKT

[Eva08a] Lawrence C. Evans. Weak KAM theory and partial differential equations. Lecture Notes in Mathematics, 1927:123–154, 2008. CODEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75914-0_4.pdf.

Evans:2008:ACR

[Eva08b] Steven Neil Evans. Around the continuum random tree. Lecture Notes in Mathematics, 1920:9–20, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74798-7_2.
pdf.

Evans:2008:BM

[Eva08c] Steven Neil Evans. Back matter. Lecture Notes in Mathematics, 1920:163–193, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-74798-7/1.pdf.

Evans:2008:DTL

[Eva08d] Steven Neil Evans. Diffusions on a R-tree without leaves: Snakes and spiders. Lecture Notes in Mathematics, 1920:105–128, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74798-7_7.pdf.

Evans:2008:FM

[Eva08e] Steven Neil Evans. Front matter. Lecture Notes in Mathematics, 1920:I–XI, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-74798-7/1.pdf.

Evans:2008:HGH

[Eva08f] Steven Neil Evans. Hausdorff and Gromov–Hausdorff distance. Lecture Notes in Mathematics, 1920:45–68, 2008. CO-DEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74798-7_4.pdf.

Evans:2008:I

[Eva08g] Steven Neil Evans. Introduction. Lecture Notes in Mathematics, 1920:1–8, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74798-7_1.pdf.

Evans:2008:PRT

[Eva08h]  Steven Neil Evans, editor.  Probability and Real Trees:

École d’Été de Probabilités de Saint-Flour XXXV — 2005, volume 1920 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no.

1920; QA166.2 .E94 2008. URL http://link.springer.com/
book/10.1007/978-3-540-74798-7;   http://www.loc.gov/
catdir/enhancements/fy0905/2007934014-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0905/2007934014-d.
html;   http://www.loc.gov/catdir/enhancements/fy0905/
2007934014-t.html;   http://www.springerlink.com/content/
978-3-540-74798-7.

Evans:2008:THS

[Eva08i] Steven Neil Evans. R-trees and 0-hyperbolic spaces. Lecture Notes in Mathematics, 1920:21–44, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74798-7_3.pdf.

Evans:2008:TCP

[Eva08j] Steven Neil Evans. R-trees from coalescing particle systems. Lecture Notes in Mathematics, 1920:129–141, 2008. CO-DEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74798-7_8.pdf.

Evans:2008:RGR

[Eva08k] Steven Neil Evans. Root growth with re-grafting. Lecture Notes in Mathematics, 1920:69–86, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74798-7_5.pdf.

Evans:2008:SPR

[Eva08l] Steven Neil Evans. Subtree prune and re-graft. Lecture Notes in Mathematics, 1920:143–162, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74798-7_9.pdf.

Evans:2008:WCO

[Eva08m] Steven Neil Evans. The wild chain and other bipartite chains. Lecture Notes in Mathematics, 1920:87–103, 2008. CODEN LNMAA2. ISBN 3-540-74797-4 (print), 3-540-74798-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74798-7_6.pdf.

Emery:2002:SPSb

[ÉY02a] Michel Émery and Marc Yor, editors. Séminaire de probabilités 1967–1980: A selection in martingale theory, volume 1771 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CO-DEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1771; QA3 .L28 no. 1771 QA274.5. URL http:/

/link.springer.com/book/10.1007/b82894;   http://www.
springerlink.com/content/978-3-540-45530-1.

Emery:2002:SPSa

[ÉY02b] Michel Émery and Marc Yor. A short presentation of the selected articles. Lecture Notes in Mathematics, 1771:1–8, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_1/.

Emery:2006:MPA

[ÉY06] Michel Émery and Marc Yor, editors. In Memoriam Paul-André Meyer: Séminaire de Probabilités XXXIX, volume 1874 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LN-MAA2. ISBN 3-540-30994-2 (paperback), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28

no. 1874. URL http://www.springerlink.com/content/978-
3-540-35513-7.

Fagnola:2006:QSD

[Fag06] Franco Fagnola. Quantum stochastic differential equations and dilation of completely positive semigroups. Lecture Notes in Mathematics, 1881:183–220, 2006. CODEN LNMAA2. ISBN 3-540-30992-6 (print), 3-540-33966-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-33966-3_5.pdf.

Falk:2008:BM

[Fal08a] Richard S. Falk. Back matter. Lecture Notes in Mathematics, 1939:233–235, 2008. CODEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-78319-0/1.pdf.

Falk:2008:FEM

[Fal08b] Richard S. Falk. Finite element methods for linear elasticity. Lecture Notes in Mathematics, 1939:159–194, 2008. CO-DEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78319-0_4.pdf.

Falk:2008:FER

[Fal08c] Richard S. Falk. Finite elements for the Reissner–Mindlin plate. Lecture Notes in Mathematics, 1939:195–232, 2008. CODEN LNMAA2. ISBN 3-540-78314-8 (print), 3-540-78319-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78319-0_5.pdf.

Fasano:2000:FPM

[Fas00a] A. Fasano, editor. Filtration in Porous Media and Industrial Application: Lectures given at the 4th Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Cetraro, Italy August 24–29, 1998, volume 1734 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67868-9 (print), 3-540-44656-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1734. URL http:/

/link.springer.com/book/10.1007/BFb0103973;   http://
www.springerlink.com/content/978-3-540-44656-9.

Fasano:2000:FPV

[Fas00b] Antonio Fasano. Filtration problems in various industrial processes. Lecture Notes in Mathematics, 1734:79–126, 2000. CO-DEN LNMAA2. ISBN 3-540-67868-9 (print), 3-540-44656-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103976/.

Fasano:2000:SGF

[Fas00c] Antonio Fasano. Some general facts about filtration through porous media. Lecture Notes in Mathematics, 1734:1–8, 2000. CODEN LNMAA2. ISBN 3-540-67868-9 (print), 3-540-44656-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103974/.

Fogg:2002:AST

[FBF^{+}02a] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Automatic sequences and transcendence. Lecture Notes in Mathematics, 1794:53–80, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45714-3_3/.

Fogg:2002:BNS

[FBF^{+}02b] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Basic notions on substitutions. Lecture Notes in Mathematics, 1794:1–32, 2002. CODEN LN-MAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45714-3_1/.

Fogg:2002:DAS

[FBF^{+}02c] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Diophantine approximations, substitutions, and fractals. Lecture Notes in Mathematics, 1794:253– 292, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45714-3_8/.

Fogg:2002:IWG

[FBF^{+}02d] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Infinite words generated by invertible substitutions. Lecture Notes in Mathematics, 1794:295– 320, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45714-3_9/.

Fogg:2002:PLT

[FBF^{+}02e] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Piecewise linear transformations of the unit interval and Cantor sets. Lecture Notes in Mathematics, 1794:343–361, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45714-3_11/.

Fogg:2002:PDS

[FBF^{+}02f] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Polynomial dynamical systems associated with substitutions. Lecture Notes in Mathematics, 1794: 321–342, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45714-3_10/.

Fogg:2002:SOP

[FBF^{+}02g] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Some open problems. Lecture Notes in Mathematics, 1794:363–374, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-45714-3_12/.

Fogg:2002:STG

[FBF^{+}02h] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Spectral theory and geometric representation of substitutions. Lecture Notes in Mathematics, 1794: 199–252, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45714-3_7/.

Fogg:2002:SS

[FBF^{+}02i] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Sturmian sequences. Lecture Notes in Mathematics, 1794:143–198, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-45714-3_6/.

Fogg:2002:SPS

[FBF^{+}02j] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Substitutions and partitions of the set of positive integers. Lecture Notes in Mathematics, 1794: 81–98, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45714-3_4/.

Fogg:2002:SSD

[FBF^{+}02k] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Substitutions and symbolic dynamical systems. Lecture Notes in Mathematics, 1794:101–142, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45714-3_5/.

Fogg:2002:SAF

[FBF^{+}02l] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Substitutions, arithmetic and finite automata: an introduction. Lecture Notes in Mathematics, 1794: 35–52, 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45714-3_2/.

Fogg:2002:SDA

[FBF^{+}02m] N. Pytheas Fogg, Valéré Berthé, Sébastien Ferenczi, Christian Mauduit, and Anne Siegel. Substitutions in Dynamics, Arithmetics and Combinatorics, volume 1794 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-44141-7 (paperback), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvii + 402 pp. LCCN QA3 .L28 no. 1794. URL http:/

/link.springer.com/book/10.1007/b13861;   http://www.
loc.gov/catdir/enhancements/fy0817/2002030325-d.html;
 http://www.loc.gov/catdir/enhancements/fy0817/2002030325-
t.html.

Feral:2008:LDS

[Fér08] D. Féral. On large deviations for the spectral measure of discrete Coulomb gas. Lecture Notes in Mathematics, 1934:19–49, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_2.pdf.

Fitzsimmons:2009:LST

[FG09] P. J. Fitzsimmons and R. K. Getoor. Lévy systems and time changes. (French) []. Lecture Notes in Mathematics, 1979:229– 259, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_9.pdf.

Friedlander:2006:FM

[FHBIK06] J. B. Friedlander, D. R. Heath-Brown, H. Iwaniec, and J. Kaczorowski. Front matter. Lecture Notes in Mathematics, 1891: I–XI, 2006. CODEN LNMAA2. ISBN 3-540-36363-7 (print), 3-540-36364-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-36364-4/1.pdf.

Feichtinger:2008:FM

[FHL^{+}08] Hans G. Feichtinger, Bernard Helffer, Michael P. Lamoureux, Nicolas Lerner, and Joachim Toft. Front matter. Lecture Notes in Mathematics, 1949:i–xxiv, 2008. CODEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-68268-4/1.pdf.

Filipovic:2001:Ia

[Fil01a] Damir Filipović. 1. Introduction. Lecture Notes in Mathematics, 1760:1–11, 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44548-7_1/.

Filipovic:2001:SEI

[Fil01b] Damir Filipović. 2. Stochastic equations in infinite dimensions. Lecture Notes in Mathematics, 1760:13–27, 2001. CODEN LN-MAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44548-7_
2/.

Filipovic:2001:CSS

[Fil01c] Damir Filipović. 3. Consistent state space processes. Lecture Notes in Mathematics, 1760:29–56, 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44548-7_3/.

Filipovic:2001:HMR

[Fil01d] Damir Filipović. 4. The HJM methodology revisited. Lecture Notes in Mathematics, 1760:57–73, 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44548-7_4/.

Filipovic:2001:FCS

[Fil01e] Damir Filipović. 5. The forward curve spaces H_{w}. Lecture Notes in Mathematics, 1760:75–94, 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44548-7_5/.

Filipovic:2001:IMS

[Fil01f] Damir Filipović. 6. Invariant manifolds for stochastic equations. Lecture Notes in Mathematics, 1760:95–111, 2001. CODEN LN-MAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44548-7_
6/.

Filipovic:2001:CHM

[Fil01g] Damir Filipović. 7. Consistent HJM models. Lecture Notes in Mathematics, 1760:113–125, 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44548-7_7/.

Filipovic:2001:ASC

[Fil01h] Damir Filipović. 8. Appendix: a summary of conditions. Lecture Notes in Mathematics, 1760:127–128, 2001. CODEN LN-MAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-44548-7_8.

Filipovic:2001:CPH

[Fil01i] Damir Filipović. Consistency Problems for Heath–Jarrow–Morton Interest Rate Models, volume 1760 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 134 pp. LCCN QA3 .L28 no. 1760. URL

http://link.springer.com/book/10.1007/b76888;   http:
//www.springerlink.com/content/978-3-540-44548-7.

Filipovic:2001:Ib

[Fil01j] Damir Filipović. Index. Lecture Notes in Mathematics, 1760: 133–134, 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44548-7_10.

Filipovic:2001:R

[Fil01k] Damir Filipović. References. Lecture Notes in Mathematics, 1760: 129–131, 2001. CODEN LNMAA2. ISBN 3-540-41493-2 (print), 3-540-44548-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44548-7_9/.

Finkelshtein:2006:EPP

[Fin06] Andrei Martínez Finkelshtein. Equilibrium problems of potential theory in the complex plane. Lecture Notes in Mathematics, 1883:79–117, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-36716-1_2.pdf.

Fitzsimmons:2001:HM

[Fit01] P. J. Fitzsimmons. Hermite martingales. Lecture Notes in Mathematics, 1755:153–157, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44671-2_10/.

Favre:2004:G

[FJ04a] Charles Favre and Mattias Jonsson. 1. Generalities. Lecture Notes in Mathematics, 1853:9–24, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_2/.

Favre:2004:MM

[FJ04b] Charles Favre and Mattias Jonsson. 2. MacLane’s method. Lecture Notes in Mathematics, 1853:25–42, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44646-0_3/.

Favre:2004:TS

[FJ04c] Charles Favre and Mattias Jonsson. 3. Tree structures. Lecture Notes in Mathematics, 1853:43–80, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44646-0_4/.

Favre:2004:VTP

[FJ04d] Charles Favre and Mattias Jonsson. 4. Valuations through Puiseux series. Lecture Notes in Mathematics, 1853:81–96, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44646-0_5/.

Favre:2004:T

[FJ04e] Charles Favre and Mattias Jonsson. 5. Topologies. Lecture Notes in Mathematics, 1853:97–110, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_6/.

Favre:2004:UDG

[FJ04f] Charles Favre and Mattias Jonsson. 6. The universal dual graph. Lecture Notes in Mathematics, 1853:111–150, 2004. CODEN LN-MAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44646-0_

Favre:2004:TM

[FJ04g] Charles Favre and Mattias Jonsson. 7. Tree measures. Lecture Notes in Mathematics, 1853:151–192, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44646-0_8/.

Favre:2004:ATA

[FJ04h] Charles Favre and Mattias Jonsson. 8. Applications of the tree analysis. Lecture Notes in Mathematics, 1853:193–209, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44646-0_9/.

Favre:2004:A

[FJ04i] Charles Favre and Mattias Jonsson. Appendix. Lecture Notes in Mathematics, 1853:211–225, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_10/.

Favre:2004:Ea

[FJ04j] Charles Favre and Mattias Jonsson. Errata. Lecture Notes in Mathematics, 1853:241, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_13.

Favre:2004:Eb

[FJ04k] Charles Favre and Mattias Jonsson. Errata. Lecture Notes in Mathematics, 1853:241–243, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_14/.

Favre:2004:Ec

[FJ04l] Charles Favre and Mattias Jonsson. Errata. Lecture Notes in Mathematics, 1853:243, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_15.

Favre:2004:Ed

[FJ04m] Charles Favre and Mattias Jonsson. Errata. Lecture Notes in Mathematics, 1853:243, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_16.

Favre:2004:Ee

[FJ04n] Charles Favre and Mattias Jonsson. Errata. Lecture Notes in Mathematics, 1853:243–244, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_17.

Favre:2004:FM

[FJ04o] Charles Favre and Mattias Jonsson. Front matter. Lecture Notes in Mathematics, 1853:I–XV, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-44646-0/1.pdf.

Favre:2004:Ib

[FJ04p] Charles Favre and Mattias Jonsson. Index. Lecture Notes in Mathematics, 1853:231–234, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_12/.

Favre:2004:Ia

[FJ04q] Charles Favre and Mattias Jonsson. Introduction. Lecture Notes in Mathematics, 1853:1–7, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_1/.

Favre:2004:R

[FJ04r] Charles Favre and Mattias Jonsson. References. Lecture Notes in Mathematics, 1853:227–229, 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44646-0_11/.

Favre:2004:VT

[FJ04s] Charles Favre and Mattias Jonsson. The Valuative Tree, volume 1853 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22984-1 (print), 3-540-44646-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 234 pp. LCCN QA3 .L28 no. 1853. URL http://

link.springer.com/book/10.1007/b100262;   http://www.
springerlink.com/content/978-3-540-44646-0.

Fitzgibbon:2008:SMT

[FL08] W.-E. Fitzgibbon and M. Langlais. Simple models for the transmission of microparasites between host populations living on noncoincident spatial domains. Lecture Notes in Mathematics, 1936:115–164, 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78273-5_3.pdf.

Flandoli:2008:ISF

[Fla08] Franco Flandoli. An introduction to 3D stochastic fluid dynamics. Lecture Notes in Mathematics, 1942:51–150, 2008. CODEN LNMAA2. ISBN 3-540-78492-6 (print), 3-540-78493-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78493-7_2.pdf.

Feichtinger:2008:BGT

[FLC08] H. Feichtinger, F. Luef, and E. Cordero. Banach Gelfand triples for Gabor analysis. Lecture Notes in Mathematics, 1949:1–33, 2008. CODEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68268-4_1.pdf.

Fornaess:2004:RMC

[For04] John Erik Fornæss. Real methods in complex dynamics. Lecture Notes in Mathematics, 1848:49–107, 2004. CODEN LNMAA2. ISBN 3-540-22358-4 (print), 3-540-44487-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44487-9_2/.

Fougeres:2005:SGL

[Fou05] Pierre Fougères. Spectral gap for log-concave probability measures on the real line. Lecture Notes in Mathematics, 1857:95–123, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_7.pdf.

Fagnola:2006:NQB

[FR06a] Franco Fagnola and Rolando Rebolledo. Notes on the qualitative behaviour of quantum Markov semigroups. Lecture Notes in Mathematics, 1882:161–205, 2006. CODEN LNMAA2. ISBN 3-540-30993-4 (print), 3-540-33967-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-33967-1_4.pdf.

Franz:2006:RWF

[FR06b] Uwe Franz and Rolf Rolf. Random walks on finite quantum groups. Lecture Notes in Mathematics, 1866:1–32, 2006. CODEN LNMAA2. ISBN 3-540-24407-7 (print), 3-540-32385-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11376637_1.pdf.

Franz:2006:LPQ

[Fra06] Uwe Franz. Lévy processes on quantum groups and dual groups. Lecture Notes in Mathematics, 1866:161–257, 2006. CODEN LNMAA2. ISBN 3-540-24407-7 (print), 3-540-32385-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11376637_3.pdf.

Frenkel:2008:RGL

[Fre08] Edward Frenkel. Ramifications of the Geometric Langlands Program. Lecture Notes in Mathematics, 1931:51–135, 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76892-0_2.pdf.

Fresse:2009:AEP

[Fre09a] Benoit Fresse. Adjunction and embedding properties. Lecture Notes in Mathematics, 1967:121–128, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89056-0_8.
pdf.

Fresse:2009:ARM

[Fre09b] Benoit Fresse. Algebras in right modules over operads. Lecture Notes in Mathematics, 1967:129–138, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89056-0_9.
pdf.

Fresse:2009:BMa

[Fre09c] Benoit Fresse. Back matter. Lecture Notes in Mathematics, 1967: 95–96, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-89056-0/1/1.pdf.

Fresse:2009:BMb

[Fre09d] Benoit Fresse. Back matter. Lecture Notes in Mathematics, 1967: 149, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-89056-0/2/1.pdf.

Fresse:2009:BMc

[Fre09e] Benoit Fresse. Back matter. Lecture Notes in Mathematics, 1967: 215–216, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-89056-0/3/1.pdf.

Fresse:2009:BMd

[Fre09f] Benoit Fresse. Back matter. Lecture Notes in Mathematics, 1967: 261, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-89056-0/4/1.pdf.

Fresse:2009:CHA

[Fre09g] Benoit Fresse. The (co)homology of algebras over operads. Lecture Notes in Mathematics, 1967:203–214, 2009. CO-DEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_13.pdf.

Fresse:2009:DBC

[Fre09h] Benoit Fresse. Definitions and basic constructions. Lecture Notes in Mathematics, 1967:99–106, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89056-0_5.pdf.

Fresse:2009:ERF

[Fre09i] Benoit Fresse. Extension and restriction functors and model structures. Lecture Notes in Mathematics, 1967:235–239, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_16.pdf.

Fresse:2009:FMa

[Fre09j] Benoit Fresse. Front matter. Lecture Notes in Mathematics, 1967: 1–2, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-89056-0/5/1.pdf.

Fresse:2009:FMb

[Fre09k] Benoit Fresse. Front matter. Lecture Notes in Mathematics, 1967: 1–8, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-89056-0/1.pdf.

Fresse:2009:FMc

[Fre09l] Benoit Fresse. Front matter. Lecture Notes in Mathematics, 1967: 17–20, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-89056-0/1/1.pdf.

Fresse:2009:FMd

[Fre09m] Benoit Fresse. Front matter. Lecture Notes in Mathematics, 1967: 98, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-89056-0/2/1.pdf.

Fresse:2009:FMe

[Fre09n] Benoit Fresse. Front matter. Lecture Notes in Mathematics, 1967: 152, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-89056-0/3/1.pdf.

Fresse:2009:FMf

[Fre09o] Benoit Fresse. Front matter. Lecture Notes in Mathematics, 1967: 218, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-89056-0/4/1.pdf.

Fresse:2009:HAO

[Fre09p] Benoit Fresse. The homotopy of algebras over operads. Lecture Notes in Mathematics, 1967:185–202, 2009. CO-DEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_12.pdf.

Fresse:2009:I

[Fre09q] Benoit Fresse. Introduction. Lecture Notes in Mathematics, 1967: 1–13, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-89056-0_0.pdf.

Fresse:2009:MA

[Fre09r] Benoit Fresse. Miscellaneous applications. Lecture Notes in Mathematics, 1967:241–259, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89056-0_17.pdf.

Fresse:2009:ME

[Fre09s] Benoit Fresse. Miscellaneous examples. Lecture Notes in Mathematics, 1967:139–147, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89056-0_10.pdf.

Fresse:2009:MSA

[Fre09t] Benoit Fresse. Miscellaneous structures associated to algebras over operads. Lecture Notes in Mathematics, 1967:77–93, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_4.pdf.

Fresse:2009:MCR

[Fre09u] Benoit Fresse. The model category of right modules. Lecture Notes in Mathematics, 1967:219–223, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89056-0_14.
pdf.

Fresse:2009:MHI

[Fre09v] Benoit Fresse. Modules and homotopy invariance of functors. Lecture Notes in Mathematics, 1967:225–233, 2009. CO-DEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_15.pdf.

Fresse:2009:MOF

[Fre09w] Benoit Fresse. Modules over Operads and Functors, volume 1967 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 308 pp. LCCN QA3 .L28 no. 1967. URL http://link.

springer.com/book/10.1007/978-3-540-89056-0;   http:/
/www.springerlink.com/content/978-3-540-89056-0.

Fresse:2009:OAS

[Fre09x] Benoit Fresse. Operads and algebras in symmetric monoidal categories. Lecture Notes in Mathematics, 1967:53–76, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_3.pdf.

Fresse:2009:SFP

[Fre09y] Benoit Fresse. Shifted functors and pushout-products. Lecture Notes in Mathematics, 1967:277–286, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89056-0_19.
pdf.

Fresse:2009:SMO

[Fre09z] Benoit Fresse. Shifted modules over operads and functors. Lecture Notes in Mathematics, 1967:267–276, 2009. CO-DEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_18.pdf.

Fresse:2009:SMC

[Fre09-27] Benoit Fresse. Symmetric monoidal categories for operads. Lecture Notes in Mathematics, 1967:21–34, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89056-0_1.
pdf.

Fresse:2009:SMM

[Fre09-28] Benoit Fresse. Symmetric monoidal model categories for operads. Lecture Notes in Mathematics, 1967:153–184, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_11.pdf.

Fresse:2009:SOF

[Fre09-29] Benoit Fresse. Symmetric objects and functors. Lecture Notes in Mathematics, 1967:1–18, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89056-0_2.pdf.

Fresse:2009:TP

[Fre09-30] Benoit Fresse. Tensor products. Lecture Notes in Mathematics, 1967:107–112, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-89056-0_6.pdf.

Fresse:2009:UCR

[Fre09-31] Benoit Fresse. Universal constructions on right modules over operads. Lecture Notes in Mathematics, 1967:113–119, 2009. CODEN LNMAA2. ISBN 3-540-89055-6 (print), 3-540-89056-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89056-0_7.pdf.

Friedman:2005:IN

[Fri05] Avner Friedman. Introduction to neurons. Lecture Notes in Mathematics, 1860:1–20, 2005. CODEN LNMAA2. ISBN 3-540-23858-1 (print), 3-540-31544-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31544-5_1.pdf.

Friedlander:2006:PPN

[Fri06a] John B. Friedlander. Producing prime numbers via sieve methods. Lecture Notes in Mathematics, 1891:1–49, 2006. CO-DEN LNMAA2. ISBN 3-540-36363-7 (print), 3-540-36364-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36364-4_1.pdf.

Friedman:2006:BM

[Fri06b] Avner Friedman. Back matter. Lecture Notes in Mathematics, 1872:??, 2006. CODEN LNMAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-32415-7/1.pdf.

Friedman:2006:CMT

[Fri06c] Avner Friedman. Cancer models and their mathematical analysis. Lecture Notes in Mathematics, 1872:223–246, 2006. CODEN LNMAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11561606_6.pdf.

Friedman:2006:TMB

[Fri06d] Avner Friedman. Tutorials in Mathematical Biosciences III: Cell Cycle, Proliferation, and Cancer, volume 1872 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LN-MAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 246 pp.

LCCN QA3 .L28 no. 1872. URL http://link.springer.com/
book/10.1007/11561606;   http://www.springerlink.com/
content/978-3-540-32415-7.

Friedman:2008:TMB

[Fri08] Avner Friedman. Tutorials in Mathematical Biosciences IV: Evolution and Ecology, volume 1922 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-74328-6 (print), 3-540-74331-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 205 pp. LCCN QA3 .L28 no.

1922. URL http://link.springer.com/book/10.1007/978-
3-540-74331-6;   http://www.springerlink.com/content/
978-3-540-74331-6.

Fuchs:2000:DPW

[FS00a] Martin Fuchs and Gregory Seregin. Differentiability properties of weak solutions to boundary value problems in the deformation theory of plasticity. Lecture Notes in Mathematics, 1749: 40–130, 2000. CODEN LNMAA2. ISBN 3-540-41397-9 (print), 3-540-44442-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103754/.

Fuchs:2000:FPE

[FS00b] Martin Fuchs and Gregory Seregin. Fluids of Prandtl–Eyring type and plastic materials with logarithmic hardening law. Lecture Notes in Mathematics, 1749:207–259, 2000. CODEN LN-MAA2. ISBN 3-540-41397-9 (print), 3-540-44442-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103756/.

Fuchs:2000:I

[FS00c] Martin Fuchs and Gregory Seregin. Introduction. Lecture Notes in Mathematics, 1749:1–4, 2000. CODEN LNMAA2. ISBN 3-540-41397-9 (print), 3-540-44442-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103752/.

Fuchs:2000:QSF

[FS00d] Martin Fuchs and Gregory Seregin. Quasi-static fluids of generalized Newtonian type. Lecture Notes in Mathematics, 1749: 131–206, 2000. CODEN LNMAA2. ISBN 3-540-41397-9 (print), 3-540-44442-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103755/.

Fuchs:2000:VMP

[FS00e] Martin Fuchs and Gregory Seregin. Variational Methods for Problems from Plasticity Theory and for Generalized Newtonian Fluids, volume 1749 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41397-9 (print), 3-540-44442-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 269 pp. LCCN QA3 .L28 no. 1749. URL http:/

/link.springer.com/book/10.1007/BFb0103751;   http://
www.springerlink.com/content/978-3-540-44442-8.

Fuchs:2000:WSB

[FS00f] Martin Fuchs and Gregory Seregin. Weak solutions to boundary value problems in the deformation theory of perfect elastoplasticity. Lecture Notes in Mathematics, 1749:5–39, 2000. CODEN LNMAA2. ISBN 3-540-41397-9 (print), 3-540-44442-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103753/.

Franz:2008:QPT

[FS08] Uwe Franz and Michael Schürmann. Quantum Potential Theory, volume 1954 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 457 pp. LCCN QA3 .L28 no. 1954. URL http://link.

springer.com/book/10.1007/978-3-540-69365-9;   http:/
/www.springerlink.com/content/978-3-540-69365-9.

Forzani:2001:SPC

[FSU01] Liliana Forzani, Roberto Scotto, and Wilfredo Urbina. A simple proof of the L^{p} continuity of the higher order Riesz transforms with respect to the Gaussian measure γd. Lecture Notes in Mathematics, 1755:162–166, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44671-2_12/.

Fuhr:2005:I

[Füh05a] Hartmut Führ. 1. Introduction. Lecture Notes in Mathematics, 1863:1–13, 2005. CODEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31552-0_1.pdf.

Fuhr:2005:WTG

[Füh05b] Hartmut Führ. 2. Wavelet transforms and group representations. Lecture Notes in Mathematics, 1863:15–58, 2005. CO-DEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31552-0_2.pdf.

Fuhr:2005:PTL

[Füh05c] Hartmut Führ. 3. The Plancherel transform for locally compact groups. Lecture Notes in Mathematics, 1863:59–103, 2005. CODEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31552-0_3.pdf.

Fuhr:2005:PIW

[Füh05d] Hartmut Führ. 4. Plancherel inversion and wavelet transforms. Lecture Notes in Mathematics, 1863:105–138, 2005. CODEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31552-0_4.pdf.

Fuhr:2005:AVG

[Füh05e] Hartmut Führ. 5. Admissible vectors for group extensions. Lecture Notes in Mathematics, 1863:139–168, 2005. CO-DEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31552-0_5.pdf.

Fuhr:2005:STH

[Füh05f] Hartmut Führ. 6. Sampling theorems for the Heisenberg group. Lecture Notes in Mathematics, 1863:169–184, 2005. CODEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31552-0_6.pdf.

Fuhr:2005:AHA

[Füh05g] Hartmut Führ. Abstract Harmonic Analysis of Continuous Wavelet Transforms, volume 1863 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 193 pp. LCCN QA3 .L28 no. 1863.

URL http://link.springer.com/book/10.1007/b104912;
http://www.springerlink.com/content/978-3-540-31552-0.

Fuhr:2005:RI

[Füh05h] Hartmut Führ. References and index. Lecture Notes in Mathematics, 1863:185–193, 2005. CODEN LNMAA2. ISBN 3-540-24259-7 (print), 3-540-31552-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31552-0_7.pdf.

Fusco:2008:GAS

[Fus08] Nicola Fusco. Geometrical aspects of symmetrization. Lecture Notes in Mathematics, 1927:155–181, 2008. CODEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-75914-0_5.
pdf.

Friz:2008:BDG

[FV08] Peter Friz and Nicolas Victoir. The Burkholder–Davis–Gundy inequality for enhanced martingales. Lecture Notes in Mathematics, 1934:421–438, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77913-1_20.pdf.

Gursky:2009:GAP

[GAC^{+}09] Matthew J. Gursky, A. (Antonio) Ambrosetti, Sun-Yung A. Chang, Andrea Malchiodi, et al., editors. Geometric analysis and PDEs: lectures given at the C.I.M.E. summer school held in Cetraro, Italy June 11–16, 2007, volume 1977 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no.

1977; QA377 .G4573 2009. URL http://link.springer.com/
book/10.1007/978-3-642-01674-5;   http://www.loc.gov/
catdir/enhancements/fy1108/2009926187-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1108/2009926187-t.
html;   http://www.springerlink.com/content/978-3-642-
01674-5.

Galtchouk:2003:RMC

[Gal03] Leonid Galtchouk. On the reduction of a multidimensional continuous martingale to a Brownian motion. Lecture Notes in Mathematics, 1832:90–93, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40004-2_4/.

Gallavotti:2006:CTS

[Gal06] Giovanni Gallavotti. Ckn theory of singularities of weak solutions of the Navier–Stokes equations. Lecture Notes in Mathematics, 1871:45–74, 2006. CODEN LNMAA2. ISBN 3-540-28586-5 (print), 3-540-32454-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/11545989_2.pdf.

Gatermann:2000:CAM

[Gat00] Karin Gatermann. Computer Algebra Methods for Equivariant Dynamical Systems, volume 1728 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67161-7. ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 153 pp. LCCN QA3 .L28 no. 1728; QA3 .L28 no. 1728

QA155.7.E4. URL http://link.springer.com/book/10.1007/
BFb0104059;   http://www.loc.gov/catdir/enhancements/
fy0816/00028453-d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00028453-t.html.

Gautschi:2006:OPQ

[Gau06] Walter Gautschi. Orthogonal polynomials, quadrature, and approximation: Computational methods and software (in matlab). Lecture Notes in Mathematics, 1883:1–77, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-36716-1_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-36716-1/1.pdf.

Gonzalez:2003:DSb

[GdS03a] Juan A. Navarro González and Juan B. Sancho de Salas. C^{∞}- Differentiable Spaces, volume 1824 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 183 (est.) pp. LCCN QA3 .L28 no. 1824. URL

http://link.springer.com/book/10.1007/b13465;   http:
//www.springerlink.com/content/978-3-540-39665-9.

Gonzalez:2003:DM

[GdS03b] Juan A. Navarro González and Juan B. Sancho de Salas. 1. Differentiable manifolds. Lecture Notes in Mathematics, 1824: 7–20, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_2/.

Gonzalez:2003:SM

[GdS03c] Juan A. Navarro González and Juan B. Sancho de Salas. 10. Smooth morphisms. Lecture Notes in Mathematics, 1824:113– 125, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_11/.

Gonzalez:2003:QCL

[GdS03d] Juan A. Navarro González and Juan B. Sancho de Salas. 11. Quotients by compact Lie groups. Lecture Notes in Mathematics, 1824:127–150, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39665-9_12/.

Gonzalez:2003:DA

[GdS03e] Juan A. Navarro González and Juan B. Sancho de Salas. 2. Differentiable algebras. Lecture Notes in Mathematics, 1824:21– 38, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_3/.

Gonzalez:2003:DSa

[GdS03f] Juan A. Navarro González and Juan B. Sancho de Salas. 3. Differentiable spaces. Lecture Notes in Mathematics, 1824:39– 49, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_4/.

Gonzalez:2003:TDS

[GdS03g] Juan A. Navarro González and Juan B. Sancho de Salas. 4. Topology of differentiable spaces. Lecture Notes in Mathematics, 1824: 51–56, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_5/.

Gonzalez:2003:E

[GdS03h] Juan A. Navarro González and Juan B. Sancho de Salas. 5. Embeddings. Lecture Notes in Mathematics, 1824:57–68, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_6/.

Gonzalez:2003:TTP

[GdS03i] Juan A. Navarro González and Juan B. Sancho de Salas. 6. Topological tensor products. Lecture Notes in Mathematics, 1824: 69–77, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_7/.

Gonzalez:2003:FP

[GdS03j] Juan A. Navarro González and Juan B. Sancho de Salas. 7. Fibred products. Lecture Notes in Mathematics, 1824:79–87, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_8/.

Gonzalez:2003:TL

[GdS03k] Juan A. Navarro González and Juan B. Sancho de Salas. 8. Topological localization. Lecture Notes in Mathematics, 1824: 89–97, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_9/.

Gonzalez:2003:FM

[GdS03l] Juan A. Navarro González and Juan B. Sancho de Salas. 9. Finite morphisms. Lecture Notes in Mathematics, 1824:99–111, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_10/.

Gonzalez:2003:A

[GdS03m] Juan A. Navarro González and Juan B. Sancho de Salas. Appendix. Lecture Notes in Mathematics, 1824:151–179, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_13/.

Gonzalez:2003:I

[GdS03n] Juan A. Navarro González and Juan B. Sancho de Salas. Introduction. Lecture Notes in Mathematics, 1824:1–5, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39665-9_
1/.

Gonzalez:2003:R

[GdS03o] Juan A. Navarro González and Juan B. Sancho de Salas. References. Lecture Notes in Mathematics, 1824:181–183, 2003. CODEN LNMAA2. ISBN 3-540-20072-X (print), 3-540-39665-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39665-9_14/.

Gatzouras:2007:MNF

[GGM07] D. Gatzouras, A. Giannopoulos, and N. Markoulakis. On the maximal number of facets of 0 / 1 polytopes. Lecture Notes in Mathematics, 1910:117–125, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72053-9_7.pdf.

Gomez:2002:FSB

[GH02] César Gómez and Rafael Hernández. Fields, stings and branes. Lecture Notes in Mathematics, 1776:39–191, 2002. CODEN LN-MAA2. ISBN 3-540-43121-7 (print), 3-540-45617-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45617-9_
3/.

Giesl:2007:AME

[Gie07a] Peter Giesl. Application of the method: Examples. Lecture Notes in Mathematics, 1904:133–147, 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69909-5_6.pdf.

Giesl:2007:BM

[Gie07b] Peter Giesl. Back matter. Lecture Notes in Mathematics, 1904: 149–170, 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-69909-5/1.pdf.

Giesl:2007:CGL

[Gie07c] Peter Giesl. Construction of Global Lyapunov Functions Using Radial Basis Functions, volume 1904 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 166 pp. LCCN QA3 .L28 no.

1904. URL http://link.springer.com/book/10.1007/978-
3-540-69909-5;   http://www.springerlink.com/content/
978-3-540-69909-5.

Giesl:2007:CLF

[Gie07d] Peter Giesl. Construction of Lyapunov functions. Lecture Notes in Mathematics, 1904:99–114, 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69909-5_4.pdf.

Giesl:2007:FM

[Gie07e] Peter Giesl. Front matter. Lecture Notes in Mathematics, 1904: I–VIII, 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69909-5/1.pdf.

Giesl:2007:GDB

[Gie07f] Peter Giesl. Global determination of the basin of attraction. Lecture Notes in Mathematics, 1904:115–132, 2007. CO-DEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69909-5_5.pdf.

Giesl:2007:I

[Gie07g] Peter Giesl. Introduction. Lecture Notes in Mathematics, 1904:1– 10, 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69909-5_1.pdf.

Giesl:2007:LF

[Gie07h] Peter Giesl. Lyapunov functions. Lecture Notes in Mathematics, 1904:11–59, 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69909-5_2.pdf.

Giesl:2007:RBF

[Gie07i] Peter Giesl. Radial basis functions. Lecture Notes in Mathematics, 1904:61–98, 2007. CODEN LNMAA2. ISBN 3-540-69907-4 (print), 3-540-69909-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69909-5_3.pdf.

Gil:2003:P

[Gil03a] Michael I. Gil’. 1. Preliminaries. Lecture Notes in Mathematics, 1830:1–9, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45225-6_1/.

Gil:2003:MRR

[Gil03b] Michael I. Gil’. 10. Multiplicative representations of resolvents. Lecture Notes in Mathematics, 1830:151–161, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45225-6_
10/.

Gil:2003:RTO

[Gil03c] Michael I. Gil’. 11. Relatively p-triangular operators. Lecture Notes in Mathematics, 1830:163–172, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45225-6_11/.

Gil:2003:RCP

[Gil03d] Michael I. Gil’. 12. Relatively compact perturbations of normal operators. Lecture Notes in Mathematics, 1830:173–180, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45225-6_12/.

Gil:2003:IMH

[Gil03e] Michael I. Gil’. 13. Infinite matrices in Hilbert spaces and differential operators. Lecture Notes in Mathematics, 1830:181– 188, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45225-6_13/.

Gil:2003:IOSa

[Gil03f] Michael I. Gil’. 14. Integral operators in space L^{2}. Lecture Notes in Mathematics, 1830:189–197, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45225-6_14/.

Gil:2003:OM

[Gil03g] Michael I. Gil’. 15. Operator matrices. Lecture Notes in Mathematics, 1830:199–213, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45225-6_15/.

Gil:2003:HTI

[Gil03h] Michael I. Gil’. 16. Hille–Tamarkin integral operators. Lecture Notes in Mathematics, 1830:215–226, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45225-6_16/.

Gil:2003:IOSb

[Gil03i] Michael I. Gil’. 17. Integral operators in space L^{[prescription−R]}. Lecture Notes in Mathematics, 1830:227–234, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45225-6_
17/.

Gil:2003:HTM

[Gil03j] Michael I. Gil’. 18. Hille–Tamarkin matrices. Lecture Notes in Mathematics, 1830:235–241, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45225-6_18/.

Gil:2003:ZEF

[Gil03k] Michael I. Gil’. 19. Zeros of entire functions. Lecture Notes in Mathematics, 1830:243–252, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45225-6_19/.

Gil:2003:NMV

[Gil03l] Michael I. Gil’. 2. Norms of matrix-valued functions. Lecture Notes in Mathematics, 1830:11–34, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45225-6_2/.

Gil:2003:IFM

[Gil03m] Michael I. Gil’. 3. Invertibility of finite matrices. Lecture Notes in Mathematics, 1830:35–48, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45225-6_3/.

Gil:2003:LEF

[Gil03n] Michael I. Gil’. 4. Localization of eigenvalues of finite matrices. Lecture Notes in Mathematics, 1830:49–63, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45225-6_
4/.

Gil:2003:BMT

[Gil03o] Michael I. Gil’. 5. Block matrices and π-triangular matrices. Lecture Notes in Mathematics, 1830:65–74, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45225-6_5/.

Gil:2003:NEF

[Gil03p] Michael I. Gil’. 6. Norm estimates for functions of compact operators in a Hilbert space. Lecture Notes in Mathematics, 1830: 75–96, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45225-6_6/.

Gil:2003:FNC

[Gil03q] Michael I. Gil’. 7. Functions of non-compact operators. Lecture Notes in Mathematics, 1830:97–121, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45225-6_7/.

Gil:2003:BPN

[Gil03r] Michael I. Gil’. 8. Bounded perturbations of nonselfadjoint operators. Lecture Notes in Mathematics, 1830:123–134, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45225-6_
8/.

Gil:2003:SLN

[Gil03s] Michael I. Gil’. 9. Spectrum localization of nonself-adjoint operators. Lecture Notes in Mathematics, 1830:135–149, 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45225-6_
9/.

Gil:2003:OFL

[Gil03t] Michael I. Gil’. Operator Functions and Localization of Spectra, volume 1830 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20246-3 (print), 3-540-45225-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 256 pp. LCCN QA3 .L28 no. 1830. URL http:/

/link.springer.com/book/10.1007/b93845;   http://www.
springerlink.com/content/978-3-540-45225-6.

Ginoux:2009:BM

[Gin09a] Nicolas Ginoux. Back matter. Lecture Notes in Mathematics, 1976:1–32, 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-01570-0/1.pdf.

Ginoux:2009:BSG

[Gin09b] Nicolas Ginoux. Basics of spin geometry. Lecture Notes in Mathematics, 1976:1–27, 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01570-0_1.pdf.

Ginoux:2009:DS

[Gin09c] Nicolas Ginoux. The Dirac Spectrum, volume 1976 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 156 pp. LCCN

QA3 .L28 no. 1976. URL http://link.springer.com/book/
10.1007/978-3-642-01570-0;   http://www.springerlink.
com/content/978-3-642-01570-0.

Ginoux:2009:DSN

[Gin09d] Nicolas Ginoux. The Dirac spectrum on non-compact manifolds. Lecture Notes in Mathematics, 1976:103–111, 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01570-0_7.pdf.

Ginoux:2009:ECS

[Gin09e] Nicolas Ginoux. Explicit computations of spectra. Lecture Notes in Mathematics, 1976:29–39, 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01570-0_2.pdf.

Ginoux:2009:FM

[Gin09f] Nicolas Ginoux. Front matter. Lecture Notes in Mathematics, 1976:1–11, 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-01570-0/1.pdf.

Ginoux:2009:LEEa

[Gin09g] Nicolas Ginoux. Lower eigenvalue estimates on closed manifolds. Lecture Notes in Mathematics, 1976:41–68, 2009. CO-DEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01570-0_3.pdf.

Ginoux:2009:LEEb

[Gin09h] Nicolas Ginoux. Lower eigenvalue estimates on compact manifolds with boundary. Lecture Notes in Mathematics, 1976:69–75, 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01570-0_4.pdf.

Ginoux:2009:OTR

[Gin09i] Nicolas Ginoux. Other topics related with the Dirac spectrum. Lecture Notes in Mathematics, 1976:113–129, 2009. CODEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01570-0_8.pdf.

Ginoux:2009:PEC

[Gin09j] Nicolas Ginoux. Prescription of eigenvalues on closed manifolds. Lecture Notes in Mathematics, 1976:93–101, 2009. CO-DEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01570-0_6.pdf.

Ginoux:2009:UEB

[Gin09k] Nicolas Ginoux. Upper eigenvalue bounds on closed manifolds. Lecture Notes in Mathematics, 1976:77–92, 2009. CO-DEN LNMAA2. ISBN 3-642-01569-7 (print), 3-642-01570-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01570-0_5.pdf.

Giovine:2008:ECT

[Gio08] Pasquale Giovine. An extended continuum theory for granular media. Lecture Notes in Mathematics, 1937:167–192, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78277-3_8.pdf.

Guentner:2004:GAP

[GK04] Erik Guentner and Jerome Kaminker. Geometric and analytic properties of groups. Lecture Notes in Mathematics, 1831:253– 262, 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39702-1_4/.

Graf:2000:AQN

[GL00a] Siegfried Graf and Harald Luschgy. Asymptotic quantization for nonsingular probability distributions. Lecture Notes in Mathematics, 1730:77–154, 2000. CODEN LNMAA2. ISBN 3-540-67394-6 (print), 3-540-45577-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103948/.

Graf:2000:AQS

[GL00b] Siegfried Graf and Harald Luschgy. Asymptotic quantization for singular probability distributions. Lecture Notes in Mathematics, 1730:155–207, 2000. CODEN LNMAA2. ISBN 3-540-67394-6 (print), 3-540-45577-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103949/.

Graf:2000:FQP

[GL00c] Siegfried Graf and Harald Luschgy. Foundations of Quantization for Probability Distributions, volume 1730 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67394-6 (print), 3-540-45577-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 230 pp. LCCN QA3 .L28 no. 1730; QA3.L28 no. 1730 QA273.6. URL http://lcweb.loc.gov/

catdir/toc/00038817.html;   http://link.springer.com/
book/10.1007/BFb0103945;   http://www.loc.gov/catdir/
enhancements/fy0816/00038817-d.html;   http://www.springerlink.
com/content/978-3-540-45577-6.

Graf:2000:GPQ

[GL00d] Siegfried Graf and Harald Luschgy. General properties of the quantization for probability distributions. Lecture Notes in Mathematics, 1730:7–75, 2000. CODEN LNMAA2. ISBN 3-540-67394-6 (print), 3-540-45577-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103947/.

Graf:2000:I

[GL00e] Siegfried Graf and Harald Luschgy. Introduction. Lecture Notes in Mathematics, 1730:1–5, 2000. CODEN LNMAA2. ISBN 3-540-67394-6 (print), 3-540-45577-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103946/.

Guedon:2000:EPC

[GL00f] O. Guédon and A. E. Litvak. Euclidean projections of a p-convex body. Lecture Notes in Mathematics, 1745:95–108, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107210/.

Gluesing-Luerssen:2002:I

[GL02a] Heide Gluesing-Luerssen. 1. Introduction. Lecture Notes in Mathematics, 1770:1–5, 2002. CODEN LNMAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45543-1_1/.

Gluesing-Luerssen:2002:AFD

[GL02b] Heide Gluesing-Luerssen. 2. The algebraic framework for delaydifferential equations. Lecture Notes in Mathematics, 1770:7– 21, 2002. CODEN LNMAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45543-1_2/.

Gluesing-Luerssen:2002:AS

[GL02c] Heide Gluesing-Luerssen. 3. The algebraic structure of H_{′}. Lecture Notes in Mathematics, 1770:23–72, 2002. CODEN LNMAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45543-1_3/.

Gluesing-Luerssen:2002:BDD

[GL02d] Heide Gluesing-Luerssen. 4. Behaviors of delay-differential systems. Lecture Notes in Mathematics, 1770:73–134, 2002. CODEN LNMAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45543-1_

Gluesing-Luerssen:2002:FR

[GL02e] Heide Gluesing-Luerssen. 5. First-out representations. Lecture Notes in Mathematics, 1770:135–167, 2002. CODEN LNMAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45543-1_5/.

Gluesing-Luerssen:2002:LDD

[GL02f] Heide Gluesing-Luerssen. Linear Delay-Differential Systems with Commensurate Delays: an Algebraic Approach, volume 1770 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 176 pp. LCCN QA3 .L28 no. 1770. URL http:

//link.springer.com/book/10.1007/82934;   http://www.
springerlink.com/content/978-3-540-45543-1.

Gluesing-Luerssen:2002:R

[GL02g] Heide Gluesing-Luerssen. References. Lecture Notes in Mathematics, 1770:169–174, 2002. CODEN LNMAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45543-1_6/.

Gluesing-Luerssen:2002:SIN

[GL02h] Heide Gluesing-Luerssen. Subject index and notation index. Lecture Notes in Mathematics, 1770:175–176, 2002. CODEN LN-MAA2. ISBN 3-540-42821-6 (print), 3-540-45543-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-45543-1_7.

Gurariy:2005:BAP

[GL05a] Vladimir I. Gurariy and Wolfgang Lusky. Bounded approximation properties. Lecture Notes in Mathematics, 1870:61– 69, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11551621_5.pdf.

Gurariy:2005:C

[GL05b] Vladimir I. Gurariy and Wolfgang Lusky. The classes [M], A, P and P_{ϵ}. Lecture Notes in Mathematics, 1870:147–154, 2005.

CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11551621_11.pdf.

Gurariy:2005:CEP

[GL05c] Vladimir I. Gurariy and Wolfgang Lusky. Classification and elementary properties of müntz sequences. Lecture Notes in Mathematics, 1870:93–103, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/11551621_7.pdf.

Gurariy:2005:CEM

[GL05d] Vladimir I. Gurariy and Wolfgang Lusky. Coefficient estimates and the müntz theorem. Lecture Notes in Mathematics, 1870: 71–92, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11551621_6.pdf.

Gurariy:2005:DS

[GL05e] Vladimir I. Gurariy and Wolfgang Lusky. Disposition of subspaces. Lecture Notes in Mathematics, 1870:1–21, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11551621_1.pdf.

Gurariy:2005:FDM

[GL05f] Vladimir I. Gurariy and Wolfgang Lusky. Finite dimensional Müntz limiting spaces in C. Lecture Notes in Mathematics, 1870:155–161, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/11551621_12.pdf.

Gurariy:2005:GMS

[GL05g] Vladimir I. Gurariy and Wolfgang Lusky. Geometry of Müntz Spaces and Related Questions, volume 1870 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 172 pp. LCCN

QA3 .L28 no. 1870. URL http://link.springer.com/book/10.
1007/11551621;   http://www.springerlink.com/content/
978-3-540-31546-9.

Gurariy:2005:IIE

[GL05h] Vladimir I. Gurariy and Wolfgang Lusky. Isomorphisms, isometries and embeddings. Lecture Notes in Mathematics, 1870:45– 51, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11551621_3.pdf.

Gurariy:2005:MGM

[GL05i] Vladimir I. Gurariy and Wolfgang Lusky. More on the geometry of müntz sequences and müntz polynomials. Lecture Notes in Mathematics, 1870:105–116, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/11551621_8.pdf.

Gurariy:2005:OFR

[GL05j] Vladimir I. Gurariy and Wolfgang Lusky. Operators of finite rank and bases in müntz spaces. Lecture Notes in Mathematics, 1870:117–136, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/11551621_9.pdf.

Gurariy:2005:PTI

[GL05k] Vladimir I. Gurariy and Wolfgang Lusky. Projection types and the isomorphism problem for müntz spaces. Lecture Notes in Mathematics, 1870:137–145, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/11551621_10.pdf.

Gurariy:2005:SNS

[GL05l] Vladimir I. Gurariy and Wolfgang Lusky. Sequences in normed spaces. Lecture Notes in Mathematics, 1870:23–43, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11551621_2.pdf.

Gurariy:2005:SUD

[GL05m] Vladimir I. Gurariy and Wolfgang Lusky. Spaces of universal disposition. Lecture Notes in Mathematics, 1870:53–60, 2005. CODEN LNMAA2. ISBN 3-540-28800-7 (print), 3-540-31546-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11551621_4.pdf.

Gursky:2009:FM

[GLM^{+}09] Matthew J. Gursky, Ermanno Lanconelli, Andrea Malchiodi, Gabriella Tarantello, Xu-Jia Wang, and Paul C. Yang. Front matter. Lecture Notes in Mathematics, 1977:1–12, 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-01674-5/
1.pdf.

Gluskin:2003:SI

[Glu03] E. D. Gluskin. On the sum of intervals. Lecture Notes in Mathematics, 1807:122–130, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36428-3_10/.

Gluskin:2004:MVB

[Glu04] E. Gluskin. On the multivariable version of Ball’s slicing cube theorem. Lecture Notes in Mathematics, 1850:117–121, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_11/.

Gourley:2008:SPD

[GLW08] S. A. Gourley, R. Liu, and J. Wu. Spatiotemporal patterns of disease spread: Interaction of physiological structure, spatial movements, disease progression and human intervention. Lecture Notes in Mathematics, 1936:165–208, 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78273-5_4.pdf.

Gluskin:2003:NGA

[GM03] E. Gluskin and V. Milman. Note on the geometric-arithmetic mean inequality. Lecture Notes in Mathematics, 1807:131–135, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36428-3_11/.

Gluskin:2004:GPR

[GM04] E. Gluskin and V. Milman. Geometric probability and random cotype 2. Lecture Notes in Mathematics, 1850:123–138, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_12/.

Gobet:2007:DSF

[GM07] Emmanuel Gobet and Stéphane Menozzi. Discrete sampling of functionals of Itô processes. Lecture Notes in Mathematics, 1899:355–374, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71189-6_19.pdf.

Giannopoulos:2000:CBM

[GMR00] A. A. Giannopoulos, V. D. Milman, and M. Rudelson. Convex bodies with minimal mean width. Lecture Notes in Mathematics, 1745:81–93, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0107209/.

Golse:2008:EMB

[GO08] François Golse and Stefano Olla, editors. Entropy Methods for the Boltzmann Equation: Lectures from a Special Semester at the Centre Émile Borel, Institut H. Poincaré, Paris, 2001, volume 1916 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-73704-9 (print), 3-540-73705-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1916. URL http://link.springer.com/book/
10.1007/978-3-540-73705-6;   http://www.springerlink.
com/content/978-3-540-73705-6.

Godefroy:2005:PCD

[God05] Laurent Godefroy. Propriété de Choquet-Deny et fonctions harmoniques sur les hypergroupes commutatifs. (French) []. Lecture Notes in Mathematics, 1857:124–134, 2005. CO-DEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_8.pdf.

Goddard:2008:GMG

[God08] J. D. Goddard. From granular matter to generalized continuum. Lecture Notes in Mathematics, 1937:1–22, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-78277-3_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-78277-3/1.pdf.

Gohm:2004:ED

[Goh04a] Rolf Gohm. 1. Extensions and dilations. Lecture Notes in Mathematics, 1839:9–36, 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39902-5_2/.

Gohm:2004:MP

[Goh04b] Rolf Gohm. 2. Markov processes. Lecture Notes in Mathematics, 1839:37–71, 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39902-5_3/.

Gohm:2004:A

[Goh04c] Rolf Gohm. 3. Adaptedness. Lecture Notes in Mathematics, 1839: 73–111, 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39902-5_4/.

Gohm:2004:EA

[Goh04d] Rolf Gohm. 4. Examples and applications. Lecture Notes in Mathematics, 1839:113–147, 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39902-5_5/.

Gohm:2004:ASF

[Goh04e] Rolf Gohm. Appendix a: Some facts about unital completely positive maps. Lecture Notes in Mathematics, 1839:149–163, 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39902-5_6/.

Gohm:2004:I

[Goh04f] Rolf Gohm. Introduction. Lecture Notes in Mathematics, 1839: 1–7, 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39902-5_1/.

Gohm:2004:NSP

[Goh04g] Rolf Gohm. Noncommutative Stationary Processes, volume 1839 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 170 pp. LCCN QA3 .L28 no. 1839. URL http:/

/link.springer.com/book/10.1007/b95349;   http://www.
springerlink.com/content/978-3-540-39902-5.

Gohm:2004:R

[Goh04h] Rolf Gohm. References. Lecture Notes in Mathematics, 1839: 165–168, 2004. CODEN LNMAA2. ISBN 3-540-20926-3 (print), 3-540-39902-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39902-5_7/.

Goichot:2001:TMM

[Goi01] François Goichot. Un théoréme de Milnor–Moore pour les algèbres de Leibniz. (French) []. Lecture Notes in Mathematics, 1763:111–133, 2001. CODEN LNMAA2. ISBN 3-540-42194-7 (print), 3-540-45328-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45328-8_5/.

Gordon:2007:NOG

[Gor07] Y. Gordon. A note on an observation of G. Schechtman. Lecture Notes in Mathematics, 1910:127–132, 2007. CO-DEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_8.pdf.

Ganesh:2004:SSQ

[GOW04a] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 1. The single server queue. Lecture Notes in Mathematics, 1838:1– 21, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_1/.

Ganesh:2004:I

[GOW04b] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 10. Interpretations. Lecture Notes in Mathematics, 1838:211–238, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_10/.

Ganesh:2004:LDE

[GOW04c] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 2. Large deviations in Euclidean spaces. Lecture Notes in Mathematics, 1838:23–45, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39889-9_2/.

Ganesh:2004:MSS

[GOW04d] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 3. More on the single server queue. Lecture Notes in Mathematics, 1838: 47–55, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_3/.

Ganesh:2004:IAL

[GOW04e] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 4. Introduction to abstract large deviations. Lecture Notes in Mathematics, 1838:57–76, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39889-9_4/.

Ganesh:2004:CQM

[GOW04f] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 5. Continuous queueing maps. Lecture Notes in Mathematics, 1838:77– 104, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_5/.

Ganesh:2004:LBS

[GOW04g] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 6. Large-buffer scalings. Lecture Notes in Mathematics, 1838:105– 150, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_6/.

Ganesh:2004:MFS

[GOW04h] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 7. Many-flows scalings. Lecture Notes in Mathematics, 1838:151– 181, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_7/.

Ganesh:2004:LRD

[GOW04i] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 8. Long range dependence. Lecture Notes in Mathematics, 1838:183– 198, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_8/.

Ganesh:2004:MDS

[GOW04j] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. 9. Moderate deviations scalings. Lecture Notes in Mathematics, 1838: 199–209, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39889-9_9/.

Ganesh:2004:B

[GOW04k] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. Bibliography. Lecture Notes in Mathematics, 1838:239–248, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39889-9_
11/.

Ganesh:2004:BQ

[GOW04l] Ayalvadi Ganesh, Neil O’Connell, and Damon Wischik. Big queues, volume 1838 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20912-3. ISSN 0075-8434. xi + 254 pp. LCCN QA3 .L28 no. 1838; QA3 .L28 no. 1838; QA274.8. URL

http://rave.ohiolink.edu/ebooks/ebc/10983423;   http:
//www.springerlink.com/openurl.asp?genre=book&isbn=978-
3-540-20912-6;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0075-8434&volume=1838.

Gabber:2003:I

[GR03a] Ofer Gabber and Lorenzo Ramero. 1. Introduction. Lecture Notes in Mathematics, 1800:1–10, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45096-2_1/.

Gabber:2003:HT

[GR03b] Ofer Gabber and Lorenzo Ramero. 2. Homological theory. Lecture Notes in Mathematics, 1800:11–49, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45096-2_2/.

Gabber:2003:ARTa

[GR03c] Ofer Gabber and Lorenzo Ramero. 3. Almost ring theory. Lecture Notes in Mathematics, 1800:50–91, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45096-2_3/.

Gabber:2003:FSA

[GR03d] Ofer Gabber and Lorenzo Ramero. 4. Fine study of almost projective modules. Lecture Notes in Mathematics, 1800:92– 129, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45096-2_4/.

Gabber:2003:HCA

[GR03e] Ofer Gabber and Lorenzo Ramero. 5. Henselization and completion of almost algebras. Lecture Notes in Mathematics, 1800: 130–194, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45096-2_5/.

Gabber:2003:VT

[GR03f] Ofer Gabber and Lorenzo Ramero. 6. Valuation theory. Lecture Notes in Mathematics, 1800:195–241, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45096-2_6/.

Gabber:2003:AG

[GR03g] Ofer Gabber and Lorenzo Ramero. 7. Analytic geometry. Lecture Notes in Mathematics, 1800:242–286, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45096-2_7/.

Gabber:2003:A

[GR03h] Ofer Gabber and Lorenzo Ramero. 8. Appendix. Lecture Notes in Mathematics, 1800:287–300, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45096-2_8/.

Gabber:2003:ARTb

[GR03i] Ofer Gabber and Lorenzo Ramero. Almost Ring Theory, volume 1800 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 307 pp. LCCN QA3 .L28 no. 1800. URL http:/

/link.springer.com/book/10.1007/b10047;   http://www.
springerlink.com/content/978-3-540-45096-2.

Gabber:2003:RI

[GR03j] Ofer Gabber and Lorenzo Ramero. References and index. Lecture Notes in Mathematics, 1800:301–303, 2003. CODEN LNMAA2. ISBN 3-540-40594-1 (print), 3-540-45096-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45096-2_9/.

Grecea:2003:DQC

[Gre03a] Valentin Grecea. Duality and quasy-continuity for supermartingales. Lecture Notes in Mathematics, 1801:402–412, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_19/.

Green:2003:ER

[Gre03b] David J. Green. Epilogue and references. Lecture Notes in Mathematics, 1828:131–135, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39680-2_10/.

Green:2003:GBC

[Gre03c] David J. Green. Gröbner Bases and the Computation of Group Cohomology, volume 1828 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 138 pp. LCCN QA3 .L28 no. 1828; QA251.5 .G733 2003. URL http:/

/link.springer.com/book/10.1007/b93836;   http://www.
springerlink.com/content/978-3-540-39680-2.

Green:2003:I

[Gre03d] David J. Green. Introduction. Lecture Notes in Mathematics, 1828:1–9, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39680-2_1/.

Green:2003:PBF

[Gre03e] David J. Green. Part I: 1. Bases for finite-dimensional algebras and modules. Lecture Notes in Mathematics, 1828:13– 20, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39680-2_2/.

Green:2003:PBA

[Gre03f] David J. Green. Part I: 2. The Buchberger algorithm for modules. Lecture Notes in Mathematics, 1828:21–32, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39680-2_
3/.

Green:2003:PCM

[Gre03g] David J. Green. Part I: 3. Constructing minimal resolutions. Lecture Notes in Mathematics, 1828:33–46, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39680-2_4/.

Green:2003:PIG

[Gre03h] David J. Green. Part II: 4. Gröbner bases for graded commutative algebras. Lecture Notes in Mathematics, 1828:49– 65, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39680-2_5/.

Green:2003:PIV

[Gre03i] David J. Green. Part II: 5. The visible ring structure. Lecture Notes in Mathematics, 1828:67–80, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39680-2_6/.

Green:2003:PIC

[Gre03j] David J. Green. Part II: 6. The completeness of the presentation. Lecture Notes in Mathematics, 1828:81–90, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39680-2_

Green:2003:PIE

[Gre03k] David J. Green. Part III: 7. Experimental results. Lecture Notes in Mathematics, 1828:93–100, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39680-2_8/.

Green:2003:SCC

[Gre03l] David J. Green. A sample cohomology calculations. Lecture Notes in Mathematics, 1828:101–130, 2003. CODEN LNMAA2. ISBN 3-540-20339-7 (print), 3-540-39680-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39680-2_9/.

Grecea:2005:PBM

[Gre05] Valentin Grecea. Positive bilinear mappings associated with stochastic processes. Lecture Notes in Mathematics, 1857:145– 157, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_10.pdf.

Garcia-Rio:2002:OCS

[GRKVL02a] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. 1. The Osserman conditions in semi-Riemannian geometry. Lecture Notes in Mathematics, 1777:1–20, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45629-2_
1/.

Garcia-Rio:2002:OCR

[GRKVL02b] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. 2. The Osserman conjecture in Riemannian geometry. Lecture Notes in Mathematics, 1777:21–37, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45629-2_
2/.

Garcia-Rio:2002:LOM

[GRKVL02c] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. 3. Lorentzian–Osserman manifolds. Lecture Notes in Mathematics, 1777:39–61, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45629-2_3/.

Garcia-Rio:2002:FDS

[GRKVL02d] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. 4. Four-dimensional semi-Riemannian Osserman manifolds with metric tensors of signature (2, 2). Lecture Notes in Mathematics, 1777:63–94, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45629-2_4/.

Garcia-Rio:2002:SRO

[GRKVL02e] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. 5. Semi-Riemannian Osserman manifolds. Lecture Notes in Mathematics, 1777:95–136, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45629-2_5/.

Garcia-Rio:2002:GOR

[GRKVL02f] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. 6. Generalizations and Osserman-related conditions. Lecture Notes in Mathematics, 1777:137–156, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45629-2_
6/.

Garcia-Rio:2002:I

[GRKVL02g] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. Index. Lecture Notes in Mathematics, 1777:165–166, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45629-2_8.

Garcia-Rio:2002:OMS

[GRKVL02h] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. Osserman Manifolds in Semi-Riemannian Geometry, volume 1777 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 166 pp. LCCN QA3 .L28 no. 1777. URL http:/

/link.springer.com/book/10.1007/b83213;   http://www.
springerlink.com/content/978-3-540-45629-2.

Garcia-Rio:2002:R

[GRKVL02i] Eduardo García-Río, Demir N. Kupeli, and Ramón Vázquez-Lorenzo. References. Lecture Notes in Mathematics, 1777:157– 163, 2002. CODEN LNMAA2. ISBN 3-540-43144-6 (print), 3-540-45629-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45629-2_7/.

Garcia-Rojo:2008:ICM

[GRMH08] R. García-Rojo, S. McNamara, and H. J. Herrmann. Influence of contact modelling on the macroscopic plastic response of granular soils under cyclic loading. Lecture Notes in Mathematics, 1937:109–124, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78277-3_6.pdf.

Groetsch:2007:BM

[Gro07a] Charles W. Groetsch. Back matter. Lecture Notes in Mathematics, 1894:121–131, 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-39943-8/1.pdf.

Groetsch:2007:FDA

[Gro07b] Charles W. Groetsch. Finite-dimensional approximations. Lecture Notes in Mathematics, 1894:101–119, 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-39943-8_5.
pdf.

Groetsch:2007:FM

[Gro07c] Charles W. Groetsch. Front matter. Lecture Notes in Mathematics, 1894:I–X, 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-39943-8/1.pdf.

Groetsch:2007:GAS

[Gro07d] Charles W. Groetsch. A general approach to stabilization. Lecture Notes in Mathematics, 1894:53–75, 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-39943-8_3.
pdf.

Groetsch:2007:HSB

[Gro07e] Charles W. Groetsch. Hilbert space background. Lecture Notes in Mathematics, 1894:19–51, 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-39943-8_2.pdf.

Groetsch:2007:SPL

[Gro07f] Charles W. Groetsch. Some problems leading to unbounded operators. Lecture Notes in Mathematics, 1894:1–17, 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-39943-8_1.pdf.

Groetsch:2007:SAE

[Gro07g] Charles W. Groetsch. Stable Approximate Evaluation of Unbounded Operators, volume 1894 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 127 pp. LCCN QA3 .L28 no.

1894. URL http://link.springer.com/book/10.1007/3-540-
39942-9;   http://www.springerlink.com/content/978-3-
540-39943-8.

Groetsch:2007:TMM

[Gro07h] Charles W. Groetsch. The Tikhonov–Morozov method. Lecture Notes in Mathematics, 1894:77–99, 2007. CODEN LNMAA2. ISBN 3-540-39942-9 (print), 3-540-39943-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-39943-8_4.
pdf.

Grune:2002:IDP

[Grü02a] Lars Grüne. 1. Introduction: Dynamics, perturbation and discretization. Lecture Notes in Mathematics, 1783:1–12, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36784-0_1/.

Grune:2002:SP

[Grü02b] Lars Grüne. 2. Setup and preliminaries. Lecture Notes in Mathematics, 1783:13–25, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36784-0_2/.

Grune:2002:SAS

[Grü02c] Lars Grüne. 3. Strongly attracting sets. Lecture Notes in Mathematics, 1783:27–68, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36784-0_3/.

Grune:2002:WAS

[Grü02d] Lars Grüne. 4. Weakly attracting sets. Lecture Notes in Mathematics, 1783:69–112, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36784-0_4/.

Grune:2002:RBD

[Grü02e] Lars Grüne. 5. Relation between discretization and perturbation. Lecture Notes in Mathematics, 1783:113–136, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36784-0_
5/.

Grune:2002:DAS

[Grü02f] Lars Grüne. 6. Discretizations of attracting sets. Lecture Notes in Mathematics, 1783:137–156, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36784-0_6/.

Grune:2002:DA

[Grü02g] Lars Grüne. 7. Domains of attraction. Lecture Notes in Mathematics, 1783:157–194, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36784-0_7/.

Grune:2002:AVS

[Grü02h] Lars Grüne. Appendix a: Viscosity solutions. Lecture Notes in Mathematics, 1783:195–200, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36784-0_8/.

Grune:2002:ABC

[Grü02i] Lars Grüne. Appendix b: Comparison functions. Lecture Notes in Mathematics, 1783:201–205, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36784-0_9/.

Grune:2002:ACN

[Grü02j] Lars Grüne. Appendix c: Numerical examples. Lecture Notes in Mathematics, 1783:207–217, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36784-0_10/.

Grune:2002:ABD

[Grü02k] Lars Grüne. Asymptotic Behavior of Dynamical and Control Systems under Perturbation and Discretization, volume 1783 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 231 pp. LCCN QA3 .L28 no. 1783. URL http:/

/link.springer.com/book/10.1007/b83677;   http://www.
springerlink.com/content/978-3-540-36784-0.

Grune:2002:I

[Grü02l] Lars Grüne. Index. Lecture Notes in Mathematics, 1783:229– 231, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36784-0_13/.

Grune:2002:N

[Grü02m] Lars Grüne. Notation. Lecture Notes in Mathematics, 1783:219– 220, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36784-0_11.

Grune:2002:R

[Grü02n] Lars Grüne. References. Lecture Notes in Mathematics, 1783: 221–227, 2002. CODEN LNMAA2. ISBN 3-540-43391-0 (print), 3-540-36784-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36784-0_12/.

Gruet:2008:HRW

[Gru08] Jean-Claude Gruet. Hyperbolic random walks. Lecture Notes in Mathematics, 1934:279–294, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-77913-1_14.pdf.

Graczyk:2009:I

[GS09a] Piotr Graczyk and Andrzej Stos. Introduction. Lecture Notes in Mathematics, 1980:1–24, 2009. CODEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-02141-1_1.pdf.

Graczyk:2009:PAS

[GS09b] Piotr Graczyk and Andrzej Stos. Potential Analysis of Stable Processes and its Extensions, volume 1980 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 187 pp. LCCN QA3 .L28 no.

1980. URL http://link.springer.com/book/10.1007/978-
3-642-02141-1;   http://www.springerlink.com/content/
978-3-642-02141-1.

Green:2007:BM

[GSE07a] James A. Green, Manfred Schocker, and Karin Erdmann. Back matter. Lecture Notes in Mathematics, 830:72–163, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-46959-9/1.pdf.

Green:2007:CLM

[GSE07b] James A. Green, Manfred Schocker, and Karin Erdmann. The Carter–Lusztig modules V_{λ}, K. Lecture Notes in Mathematics, 830:43–52, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-46959-9_5.pdf.

Green:2007:FM

[GSE07c] James A. Green, Manfred Schocker, and Karin Erdmann. Front matter. Lecture Notes in Mathematics, 830:I–IX, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-46959-9/
1.pdf.

Green:2007:I

[GSE07d] James A. Green, Manfred Schocker, and Karin Erdmann. Introduction. Lecture Notes in Mathematics, 830:1–10, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-46959-9_1.pdf.

Green:2007:M

[GSE07e] James A. Green, Manfred Schocker, and Karin Erdmann. The modules D_{λ}, K. Lecture Notes in Mathematics, 830:33–42, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-46959-9_4.pdf.

Green:2007:PRS

[GSE07f] James A. Green, Manfred Schocker, and Karin Erdmann. Polynomial representations of GLn(K): The Schur algebra. Lecture Notes in Mathematics, 830:11–22, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-46959-9_2.
pdf.

Green:2007:PR

[GSE07g] James A. Green, Manfred Schocker, and Karin Erdmann. Polynomial Representations of GL_{n}, volume 830 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 161 pp. LCCN QA3 .L28 no.

830. URL http://link.springer.com/book/10.1007/3-540-
46944-3;   http://www.springerlink.com/content/978-3-
540-46959-9.

Green:2007:RTS

[GSE07h] James A. Green, Manfred Schocker, and Karin Erdmann. Representation theory of the symmetric group. Lecture Notes in Mathematics, 830:53–70, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-46959-9_6.pdf.

Green:2007:WC

[GSE07i] James A. Green, Manfred Schocker, and Karin Erdmann. Weights and characters. Lecture Notes in Mathematics, 830:23–31, 2007. CODEN LNMAA2. ISBN 3-540-46944-3 (print), 3-540-46959-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-46959-9_3.pdf.

Guionnet:2007:DSM

[Gui07] Alice Guionnet. Dynamics for spherical models of spin-glass and aging. Lecture Notes in Mathematics, 1900:117–144, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-40908-3_5.pdf.

Guionnet:2009:AHC

[Gui09a] Alice Guionnet. Asymptotics of Harish-Chandra–Itzykson–Zuber integrals and of Schur polynomials. Lecture Notes in Mathematics, 1957:1–6, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_15.pdf.

Guionnet:2009:ASM

[Gui09b] Alice Guionnet. Asymptotics of some matrix integrals. Lecture Notes in Mathematics, 1957:1–8, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69897-5_16.pdf.

Guionnet:2009:BM

[Gui09c] Alice Guionnet. Basics of matrices. Lecture Notes in Mathematics, 1957:1–3, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_20.pdf.

Guionnet:2009:CIL

[Gui09d] Alice Guionnet. Concentration inequalities and logarithmic Sobolev inequalities. Lecture Notes in Mathematics, 1957:1–9, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69897-5_5.pdf.

Guionnet:2009:CIR

[Gui09e] Alice Guionnet. Concentration inequalities for random matrices. Lecture Notes in Mathematics, 1957:1–23, 2009. CO-DEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69897-5_7.pdf.

Guionnet:2009:FOE

[Gui09f] Alice Guionnet. First-order expansion. Lecture Notes in Mathematics, 1957:1–11, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_9.pdf.

Guionnet:2009:FE

[Gui09g] Alice Guionnet. Free entropy. Lecture Notes in Mathematics, 1957:1–15, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_19.pdf.

Guionnet:2009:FPS

[Gui09h] Alice Guionnet. Free probability setting. Lecture Notes in Mathematics, 1957:1–4, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_17.pdf.

Guionnet:2009:F

[Gui09i] Alice Guionnet. Freeness. Lecture Notes in Mathematics, 1957:1– 14, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69897-5_18.pdf.

Guionnet:2009:FMa

[Gui09j] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/1/1.pdf.

Guionnet:2009:FMb

[Gui09k] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/5/1.pdf.

Guionnet:2009:FMc

[Gui09l] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/6/1.pdf.

Guionnet:2009:FMd

[Gui09m] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/7/1.pdf.

Guionnet:2009:FMe

[Gui09n] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1–2, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/2/1.pdf.

Guionnet:2009:FMf

[Gui09o] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1–2, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/4/1.pdf.

Guionnet:2009:FMg

[Gui09p] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1–3, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/3/1.pdf.

Guionnet:2009:FMh

[Gui09q] Alice Guionnet. Front matter. Lecture Notes in Mathematics, 1957:1–11, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-69897-5/1.pdf.

Guionnet:2009:G

[Gui09r] Alice Guionnet. Generalizations. Lecture Notes in Mathematics, 1957:1–6, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_6.pdf.

Guionnet:2009:I

[Gui09s] Alice Guionnet. Introduction. Lecture Notes in Mathematics, 1957:1–4, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_1.pdf.

Guionnet:2009:LDP

[Gui09t] Alice Guionnet. Large deviation principle for the law of the spectral measure of shifted Wigner matrices. Lecture Notes in Mathematics, 1957:1–27, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_14.pdf.

Guionnet:2009:LDL

[Gui09u] Alice Guionnet. Large deviations for the law of the spectral measure of Gaussian Wigner’s matrices. Lecture Notes in Mathematics, 1957:1–10, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_11.pdf.

Guionnet:2009:LDM

[Gui09v] Alice Guionnet. Large deviations of the maximum eigenvalue. Lecture Notes in Mathematics, 1957:1–5, 2009. CO-DEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69897-5_12.pdf.

Guionnet:2009:LRM

[Gui09w]  Alice Guionnet, editor.  Large Random Matrices: Lec-

tures on Macroscopic Asymptotics: École d’Été de Probabilités de Saint-Flour XXXVI — 2006, volume 1957 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1957; QA3 .L28 no.1957.

URL http://link.springer.com/book/10.1007/978-3-540-
69897-5;   http://www.springerlink.com/content/978-3-
540-69897-5;   http://www.springerlink.com/openurl.asp?
genre=issue&issn=0075-8434&volume=1957.

Guionnet:2009:MGC

[Gui09x] Alice Guionnet. Maps and Gaussian calculus. Lecture Notes in Mathematics, 1957:1–15, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69897-5_8.pdf.

Guionnet:2009:SOE

[Gui09y] Alice Guionnet. Second-order expansion for the free energy. Lecture Notes in Mathematics, 1957:1–25, 2009. CO-DEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69897-5_10.pdf.

Guionnet:2009:SAR

[Gui09z] Alice Guionnet. Stochastic analysis for random matrices. Lecture Notes in Mathematics, 1957:1–15, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-69897-5_13.
pdf.

Guionnet:2009:WMM

[Gui09-27] Alice Guionnet. Wigner’s matrices; more moments estimates. Lecture Notes in Mathematics, 1957:1–12, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-69897-5_3.
pdf.

Guionnet:2009:WT

[Gui09-28] Alice Guionnet. Wigner’s theorem. Lecture Notes in Mathematics, 1957:1–22, 2009. CODEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69897-5_2.pdf.

Guionnet:2009:WSI

[Gui09-29] Alice Guionnet. Words in several independent Wigner matrices. Lecture Notes in Mathematics, 1957:1–6, 2009. CO-DEN LNMAA2. ISBN 3-540-69896-5 (print), 3-540-69897-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69897-5_4.pdf.

Gursky:2009:PCG

[Gur09] Matthew J. Gursky. PDEs in conformal geometry. Lecture Notes in Mathematics, 1977:1–33, 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01674-5_1.pdf.

Guedon:2003:SPT

[GZ03a] Olivier Guédon and Artem Zvavitch. Supremum of a process in terms of trees. Lecture Notes in Mathematics, 1807:136– 147, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36428-3_12/.

Guionnet:2003:LLS

[GZ03b] A. Guionnet and B. Zegarlinksi. Lectures on logarithmic Sobolev inequalities. Lecture Notes in Mathematics, 1801:1– 134, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_1/.

Habermann:2000:AFC

[Hab00a] Lutz Habermann. Asymptotics: The flat case. Lecture Notes in Mathematics, 1743:55–82, 2000. CODEN LNMAA2. ISBN 3-540-67987-1 (print), 3-540-44443-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103868/.

Habermann:2000:CMF

[Hab00b] Lutz Habermann. A canonical metric for flat conformal manifolds. Lecture Notes in Mathematics, 1743:11–31, 2000. CODEN LNMAA2. ISBN 3-540-67987-1 (print), 3-540-44443-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103866/.

Habermann:2000:GLD

[Hab00c] Lutz Habermann. Generalization in low dimensions. Lecture Notes in Mathematics, 1743:83–100, 2000. CODEN LN-MAA2. ISBN 3-540-67987-1 (print), 3-540-44443-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103869/.

Habermann:2000:KGM

[Hab00d] Lutz Habermann. Kleinian groups and moduli spaces. Lecture Notes in Mathematics, 1743:33–53, 2000. CODEN LN-MAA2. ISBN 3-540-67987-1 (print), 3-540-44443-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103867/.

Habermann:2000:MSA

[Hab00e] Lutz Habermann. The moduli space of all conformal structures. Lecture Notes in Mathematics, 1743:101–107, 2000. CODEN LNMAA2. ISBN 3-540-67987-1 (print), 3-540-44443-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103870/.

Habermann:2000:P

[Hab00f] Lutz Habermann. Preliminaries. Lecture Notes in Mathematics, 1743:1–9, 2000. CODEN LNMAA2. ISBN 3-540-67987-1 (print), 3-540-44443-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103865/.

Habermann:2000:RMC

[Hab00g] Lutz Habermann. Riemannian Metrics of Constant Mass and Moduli Spaces of Conformal Structures, volume 1743 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67987-1. ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 116 pp. LCCN QA3 .L28 no. 1743; QA3 .L28 no. 1743; QA609. URL http:/

/link.springer.com/book/10.1007/BFb0103864;   http://
rave.ohiolink.edu/ebooks/ebc/3540679871;   http://www.
loc.gov/catdir/enhancements/fy0816/00064118-d.html;

http://www.loc.gov/catdir/enhancements/fy0816/00064118-
t.html;   http://www.springerlink.com/openurl.asp?genre=
book&isbn=978-3-540-67987-5;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0075-8434&volume=1743.

Hannig:2003:FRP

[Han03] Jan Hannig. On filtrations related to purely discontinuous martingales. Lecture Notes in Mathematics, 1801:360–365, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_16/.

Hanssmann:2007:BM

[Han07a] Heinz Hanßmann. Back matter. Lecture Notes in Mathematics, 1893:207–241, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-38896-8/1.pdf.

Hanssmann:2007:BE

[Han07b] Heinz Hanßmann. Bifurcations of equilibria. Lecture Notes in Mathematics, 1893:17–89, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38896-8_2.pdf.

Hanssmann:2007:BIT

[Han07c] Heinz Hanßmann. Bifurcations of invariant tori. Lecture Notes in Mathematics, 1893:109–142, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38896-8_4.pdf.

Hanssmann:2007:BPO

[Han07d] Heinz Hanßmann. Bifurcations of periodic orbits. Lecture Notes in Mathematics, 1893:91–107, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38896-8_3.pdf.

Hanssmann:2007:FM

[Han07e] Heinz Hanßmann. Front matter. Lecture Notes in Mathematics, 1893:I–XV, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-38896-8/1.pdf.

Hanssmann:2007:I

[Han07f] Heinz Hanßmann. Introduction. Lecture Notes in Mathematics, 1893:1–15, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-38896-8_1.pdf.

Hanssmann:2007:LSL

[Han07g] Heinz Hanßmann. Local and Semi-Local Bifurcations in Hamiltonian Dynamical Systems, volume 1893 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 237 pp. LCCN QA3 .L28 no.

1893. URL http://link.springer.com/book/10.1007/3-540-
38894-X;   http://www.springerlink.com/content/978-3-
540-38896-8.

Hanssmann:2007:NFT

[Han07h] Heinz Hanßmann. Normal form theory. Lecture Notes in Mathematics, 1893:173–184, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38896-8_8.pdf.

Hanssmann:2007:PRT

[Han07i] Heinz Hanßmann. Perturbations of ramified torus bundles. Lecture Notes in Mathematics, 1893:143–159, 2007. CO-DEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-38896-8_5.pdf.

Hanssmann:2007:PS

[Han07j] Heinz Hanßmann. Planar singularities. Lecture Notes in Mathematics, 1893:161–165, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38896-8_6.pdf.

Hanssmann:2007:PMK

[Han07k] Heinz Hanßmann. Proof of the main KAM theorem. Lecture Notes in Mathematics, 1893:185–200, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38896-8_9.pdf.

Hanssmann:2007:PNL

[Han07l] Heinz Hanßmann. Proofs of the necessary lemmata. Lecture Notes in Mathematics, 1893:201–206, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-38896-8_10.pdf.

Hanssmann:2007:S

[Han07m] Heinz Hanßmann. Stratifications. Lecture Notes in Mathematics, 1893:167–171, 2007. CODEN LNMAA2. ISBN 3-540-38894-X (print), 3-540-38896-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-38896-8_7.pdf.

Harris:2000:CGB

[Har00] Simon C. Harris. Convergence of a ‘Gibbs–Boltzmann’ random measure for a typed branching diffusion. Lecture Notes in Mathematics, 1729:239–256, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103806/.

Haran:2008:AIR

[Har08a] Shai M. J. Haran. Arithmetical Investigations: Representation Theory, Orthogonal Polynomials, and Quantum Interpolations, volume 1941 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (paperback), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 217 pp. LCCN QA3 .L28 no. 1941; QA3 .L28

1941; QA241 .H273 2008. URL http://link.springer.com/

book/10.1007/978-3-540-78379-4;   http://www.loc.gov/
catdir/toc/fy0805/2008921367.html.

Haran:2008:BM

[Har08b] Shai M. J. Haran. Back matter. Lecture Notes in Mathematics, 1941:199–222, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-78379-4/1.pdf.

Haran:2008:FM

[Har08c] Shai M. J. Haran. Front matter. Lecture Notes in Mathematics, 1941:I–XII, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-78379-4/1.pdf.

Haran:2008:GBM

[Har08d] Shai M. J. Haran. Gamma and beta measures. Lecture Notes in Mathematics, 1941:19–31, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78379-4_2.pdf.

Haran:2008:HDT

[Har08e] Shai M. J. Haran. Higher dimensional theory. Lecture Notes in Mathematics, 1941:131–142, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78379-4_8.pdf.

Haran:2008:IMG

[Har08f] Shai M. J. Haran. Introduction: Motivations from geometry. Lecture Notes in Mathematics, 1941:1–18, 2008. CO-DEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78379-4_1.pdf.

Haran:2008:LS

[Har08g] Shai M. J. Haran. Ladder structure. Lecture Notes in Mathematics, 1941:63–93, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78379-4_5.pdf.

Haran:2008:MC

[Har08h] Shai M. J. Haran. Markov chains. Lecture Notes in Mathematics, 1941:33–46, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78379-4_3.pdf.

Haran:2008:AGM

[Har08i] Shai M. J. Haran. p-Adic Grassmann manifold. Lecture Notes in Mathematics, 1941:157–171, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78379-4_10.pdf.

Haran:2008:PBS

[Har08j] Shai M. J. Haran. Pure basis and semi-group. Lecture Notes in Mathematics, 1941:117–130, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78379-4_7.pdf.

Haran:2008:GM

[Har08k] Shai M. J. Haran. q-Grassmann manifold. Lecture Notes in Mathematics, 1941:173–184, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78379-4_11.pdf.

Haran:2008:ILT

[Har08l] Shai M. J. Haran. q-interpolation of local Tate thesis. Lecture Notes in Mathematics, 1941:95–115, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78379-4_6.
pdf.

Haran:2008:QGU

[Har08m] Shai M. J. Haran. Quantum group U_{q}(su(1, 1)) and the q-Hahn basis. Lecture Notes in Mathematics, 1941:185–197, 2008.

CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78379-4_12.pdf.

Haran:2008:RBC

[Har08n] Shai M. J. Haran. Real beta chain and q-interpolation. Lecture Notes in Mathematics, 1941:47–62, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78379-4_4.
pdf.

Haran:2008:RGM

[Har08o] Shai M. J. Haran. Real Grassmann manifold. Lecture Notes in Mathematics, 1941:143–156, 2008. CODEN LNMAA2. ISBN 3-540-78378-4 (print), 3-540-78379-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78379-4_9.pdf.

Heath-Brown:2006:CRP

[HB06] D. R. Heath-Brown. Counting rational points on algebraic varieties. Lecture Notes in Mathematics, 1891:51–95, 2006. CODEN LNMAA2. ISBN 3-540-36363-7 (print), 3-540-36364-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36364-4_2.pdf.

He:2000:DRA

[HCG^{+}00] J.-W. He, M. Chevalier, R. Glowinski, R. Metcalfe, A. Nordlander, and J. Periaux. Drag reduction by active control for flow past cylinders. Lecture Notes in Mathematics, 1739:287–363, 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103923/.

Helffer:2008:FLS

[Hel08] B. Helffer. Four lectures in semiclassical analysis for non selfadjoint problems with applications to hydrodynamic instability. Lecture Notes in Mathematics, 1949:35–77, 2008. CO-DEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68268-4_2.pdf.

Henrard:2005:AIT

[Hen05] Jacques Henrard. The adiabatic invariant theory and applications. Lecture Notes in Mathematics, 1861:77–141, 2005. CODEN LNMAA2. ISBN 3-540-24064-0 (print), 3-540-31541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31541-4_2.pdf.

Herve:2001:GNE

[Her01] Loïc Herve. Generalization to the non-ergodic case. Lecture Notes in Mathematics, 1766:115–140, 2001. CODEN LNMAA2. ISBN 3-540-42415-6 (print), 3-540-44623-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44623-0_15/.

Herrlich:2006:AC

[Her06a] Horst Herrlich. Axiom of Choice, volume 1876 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 194 pp. LCCN

QA3 .L28 no. 1876. URL http://link.springer.com/book/10.
1007/11601562;   http://www.springerlink.com/content/
978-3-540-34268-7.

Herrlich:2006:BC

[Her06b] Horst Herrlich. Beauty without choice. Lecture Notes in Mathematics, 1876:143–157, 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-34268-0_7.pdf.

Herrlich:2006:CP

[Her06c] Horst Herrlich. Choice principles. Lecture Notes in Mathematics, 1876:9–20, 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-34268-0_2.pdf.

Herrlich:2006:DEW

[Her06d] Horst Herrlich. Disasters either way. Lecture Notes in Mathematics, 1876:137–141, 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-34268-0_6.pdf.

Herrlich:2006:DCb

[Her06e] Horst Herrlich. Disasters with choice. Lecture Notes in Mathematics, 1876:117–136, 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-34268-0_5.pdf.

Herrlich:2006:DCa

[Her06f] Horst Herrlich. Disasters without choice. Lecture Notes in Mathematics, 1876:43–116, 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-34268-0_4.pdf.

Herrlich:2006:EO

[Her06g] Horst Herrlich. Elementary observations. Lecture Notes in Mathematics, 1876:21–42, 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-34268-0_3.pdf.

Herrlich:2006:O

[Her06h] Horst Herrlich. Origins. Lecture Notes in Mathematics, 1876: 1–8, 2006. CODEN LNMAA2. ISBN 3-540-30989-6 (print), 3-540-34268-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-34268-0_1.pdf.

Higson:2004:GAT

[HG04] Nigel Higson and Erik Guentner. Group C∗-algebras and Ktheory. Lecture Notes in Mathematics, 1831:137–251, 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-39702-1_3/.

Hennion:2001:LTM

[HH01] Hubert Hennion and Loïc Hervé. Limit Theorems for Markov Chains and Stochastic Properties of Dynamical Systems by Quasi-Compactness, volume 1766 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42415-6 (print), 3-540-44623-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 145 pp. LCCN QA3 .L28 no. 1766. URL

http://link.springer.com/book/10.1007/b87874;   http:
//www.springerlink.com/content/978-3-540-44623-1.

Habermann:2006:ASO

[HH06a] Katharina Habermann and Lutz Habermann. An associated second order operator. Lecture Notes in Mathematics, 1887:67–79, 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_5.pdf.

Habermann:2006:BM

[HH06b] Katharina Habermann and Lutz Habermann. Back matter. Lecture Notes in Mathematics, 1887:115–124, 2006. CODEN LN-MAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-33421-7/
1.pdf.

Habermann:2006:BSS

[HH06c] Katharina Habermann and Lutz Habermann. Background on symplectic spinors. Lecture Notes in Mathematics, 1887:1–19, 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_1.pdf.

Habermann:2006:FTS

[HH06d] Katharina Habermann and Lutz Habermann. Fourier transform for symplectic spinors. Lecture Notes in Mathematics, 1887:97– 100, 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_7.pdf.

Habermann:2006:FM

[HH06e] Katharina Habermann and Lutz Habermann. Front matter. Lecture Notes in Mathematics, 1887:i–xii, 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-33421-7/1.pdf.

Habermann:2006:ISD

[HH06f] Katharina Habermann and Lutz Habermann. Introduction to Symplectic Dirac Operators, volume 1887 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 124 (est.) pp. LCCN QA3 .L28

no. 1887. URL http://link.springer.com/book/10.1007/
b138212;   http://www.springerlink.com/content/978-3-
540-33421-7.

Habermann:2006:KC

[HH06g] Katharina Habermann and Lutz Habermann. The Kähler case. Lecture Notes in Mathematics, 1887:81–96, 2006. CO-DEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_6.pdf.

Habermann:2006:LDQ

[HH06h] Katharina Habermann and Lutz Habermann. Lie derivative and quantization. Lecture Notes in Mathematics, 1887:101–113, 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_8.pdf.

Habermann:2006:SC

[HH06i] Katharina Habermann and Lutz Habermann. Symplectic connections. Lecture Notes in Mathematics, 1887:21–34, 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_2.pdf.

Habermann:2006:SDO

[HH06j] Katharina Habermann and Lutz Habermann. Symplectic Dirac operators. Lecture Notes in Mathematics, 1887:49–66, 2006. CODEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_4.pdf.

Habermann:2006:SSF

[HH06k] Katharina Habermann and Lutz Habermann. Symplectic spinor fields. Lecture Notes in Mathematics, 1887:35–48, 2006. CO-DEN LNMAA2. ISBN 3-540-33420-3 (print), 3-540-33421-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33421-7_3.pdf.

Hardy:2009:SAB

[HH09] Robert Hardy and Simon C. Harris. A spine approach to branching diffusions with applications to L^{p}-convergence of martingales. Lecture Notes in Mathematics, 1979:281–330, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_11.pdf.

Hipp:2004:SCA

[Hip04] Christian Hipp. Stochastic control with application in insurance. Lecture Notes in Mathematics, 1856:127–164, 2004. CODEN LNMAA2. ISBN 3-540-22953-1 (print), 3-540-44644-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44644-6_
3/.

Hirasaka:2002:EPC

[Hir02] Mitsugu Hirasaka. The enumeration of primitive commutative association schemes with a non-symmetric relation of valency, at most 4. Lecture Notes in Mathematics, 1773:105– 119, 2002. CODEN LNMAA2. ISBN 3-540-42851-8 (print), 3-540-45558-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45558-2_5/.

Hirano:2003:MDP

[Hir03] Katsuhiro Hirano. On the maximum of a diffusion process in a random Lévy environment. Lecture Notes in Mathematics, 1832: 216–235, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_8/.

Hiai:2003:I

[HK03a] Fumio Hiai and Hideki Kosaki. 1. Introduction. Lecture Notes in Mathematics, 1820:1–6, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45152-5_1/.

Hiai:2003:DIT

[HK03b] Fumio Hiai and Hideki Kosaki. 2. Double integral transformations. Lecture Notes in Mathematics, 1820:7–32, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45152-5_
2/.

Hiai:2003:MOT

[HK03c] Fumio Hiai and Hideki Kosaki. 3. Means of operators and their comparison. Lecture Notes in Mathematics, 1820:33–55, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45152-5_3/.

Hiai:2003:CM

[HK03d] Fumio Hiai and Hideki Kosaki. 4. Convergence of means. Lecture Notes in Mathematics, 1820:57–63, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45152-5_4/.

Hiai:2003:GIM

[HK03e] Fumio Hiai and Hideki Kosaki. 5. A-L-G interpolation means M_{α}. Lecture Notes in Mathematics, 1820:65–78, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45152-5_
5/.

Hiai:2003:HTM

[HK03f] Fumio Hiai and Hideki Kosaki. 6. Heinz-type means A_{α}. Lecture Notes in Mathematics, 1820:79–87, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45152-5_6/.

Hiai:2003:BM

[HK03g] Fumio Hiai and Hideki Kosaki. 7. Binomial means B_{α}. Lecture Notes in Mathematics, 1820:89–104, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45152-5_7/.

Hiai:2003:CAS

[HK03h] Fumio Hiai and Hideki Kosaki. 8. Certain alternating sums of operators. Lecture Notes in Mathematics, 1820:105–121, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45152-5_8/.

Hiai:2003:A

[HK03i] Fumio Hiai and Hideki Kosaki. A appendices. Lecture Notes in Mathematics, 1820:123–139, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45152-5_9/.

Hiai:2003:MHS

[HK03j] Fumio Hiai and Hideki Kosaki. Means of Hilbert Space Operators, volume 1820 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 148 pp. LCCN QA3 .L28 no. 1820. URL http:/

/link.springer.com/book/10.1007/b13213;   http://www.
springerlink.com/content/978-3-540-45152-5.

Hiai:2003:R

[HK03k] Fumio Hiai and Hideki Kosaki. References. Lecture Notes in Mathematics, 1820:141–144, 2003. CODEN LNMAA2. ISBN 3-540-40680-8 (print), 3-540-45152-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45152-5_10/.

Helffer:2005:I

[HN05a] Bernard Helffer and Francis Nier. 1. Introduction. Lecture Notes in Mathematics, 1862:1–9, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31553-7_1.pdf.

Helffer:2005:MMS

[HN05b] Bernard Helffer and Francis Nier. 10. Maximal microhypoellipticity for systems and applications to Witten Laplacians. Lecture Notes in Mathematics, 1862:97–112, 2005. CO-DEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_10.pdf.

Helffer:2005:SPW

[HN05c] Bernard Helffer and Francis Nier. 11. Spectral properties of the Witten–Laplacians in connection with Poincaré inequalities for Laplace integrals. Lecture Notes in Mathematics, 1862:113–131, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_11.pdf.

Helffer:2005:SCAa

[HN05d] Bernard Helffer and Francis Nier. 12. Semi-classical analysis for the Schrödinger operator: Harmonic approximation. Lecture Notes in Mathematics, 1862:133–145, 2005. CO-DEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_12.pdf.

Helffer:2005:DEA

[HN05e] Bernard Helffer and Francis Nier. 13. Decay of eigenfunctions and application to the splitting. Lecture Notes in Mathematics, 1862:147–161, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31553-7_13.pdf.

Helffer:2005:SCAb

[HN05f] Bernard Helffer and Francis Nier. 14. Semi-classical analysis and Witten Laplacians: Morse inequalities. Lecture Notes in Mathematics, 1862:163–172, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31553-7_14.pdf.

Helffer:2005:SCAc

[HN05g] Bernard Helffer and Francis Nier. 15. Semi-classical analysis and Witten Laplacians: Tunneling effects. Lecture Notes in Mathematics, 1862:173–180, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31553-7_15.pdf.

Helffer:2005:AAE

[HN05h] Bernard Helffer and Francis Nier. 16. Accurate asymptotics for the exponentially small eigenvalues of Δ_{f,h}^{(0)}. Lecture Notes in Mathematics, 1862:181–188, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31553-7_16.pdf.

Helffer:2005:AFP

[HN05i] Bernard Helffer and Francis Nier. 17. Application to the Fokker– Planck equation. Lecture Notes in Mathematics, 1862:189–191, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_17.pdf.

Helffer:2005:E

[HN05j] Bernard Helffer and Francis Nier. 18. Epilogue. Lecture Notes in Mathematics, 1862:193, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31553-7_18.pdf.

Helffer:2005:KPH

[HN05k] Bernard Helffer and Francis Nier. 2. Kohn’s proof of the hypoellipticity of the Hörmander operators. Lecture Notes in Mathematics, 1862:11–18, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31553-7_2.pdf.

Helffer:2005:CCR

[HN05l] Bernard Helffer and Francis Nier. 3. Compactness criteria for the resolvent of Schrödinger operators. Lecture Notes in Mathematics, 1862:19–26, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31553-7_3.pdf.

Helffer:2005:GPD

[HN05m] Bernard Helffer and Francis Nier. 4. Global pseudo-differential calculus. Lecture Notes in Mathematics, 1862:27–42, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_4.pdf.

Helffer:2005:ASF

[HN05n] Bernard Helffer and Francis Nier. 5. Analysis of some Fokker– Planck operator. Lecture Notes in Mathematics, 1862:43–64, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_5.pdf.

Helffer:2005:REF

[HN05o] Bernard Helffer and Francis Nier. 6. Return to equilibrium for the Fokker–Planck operator. Lecture Notes in Mathematics, 1862:65– 72, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_6.pdf.

Helffer:2005:HNG

[HN05p] Bernard Helffer and Francis Nier. 7. Hypoellipticity and nilpotent groups. Lecture Notes in Mathematics, 1862:73–78, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_7.pdf.

Helffer:2005:MHP

[HN05q] Bernard Helffer and Francis Nier. 8. Maximal hypoellipticity for polynomial of vector fields and spectral byproducts. Lecture Notes in Mathematics, 1862:79–87, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31553-7_8.pdf.

Helffer:2005:FPO

[HN05r] Bernard Helffer and Francis Nier. 9. On Fokker–Planck operators and nilpotent techniques. Lecture Notes in Mathematics, 1862:89– 95, 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_9.pdf.

Helffer:2005:HES

[HN05s] Bernard Helffer and Francis Nier. Hypoelliptic Estimates and Spectral Theory for Fokker–Planck Operators and Witten Laplacians, volume 1862 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 209 pp. LCCN QA3 .L28 no. 1862. URL http://

link.springer.com/book/10.1007/b104762;   http://www.
springerlink.com/content/978-3-540-31553-7.

Helffer:2005:RI

[HN05t] Bernard Helffer and Francis Nier. References and index. Lecture Notes in Mathematics, 1862:195–209, 2005. CO-DEN LNMAA2. ISBN 3-540-24200-7 (print), 3-540-31553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31553-7_19.pdf.

Hayashi:2006:ADN

[HNKS06a] Nakao Hayashi, Pavel I. Naumkin, Elena I. Kaikina, and Ilya A. Shishmarev. Asymptotics for Dissipative Nonlinear Equations, volume 1884 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-32059-8 (print), 3-540-32060-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 540 (est.) pp. LCCN QA3 .L28 no. 1884. URL http://

link.springer.com/book/10.1007/b133345;   http://www.
springerlink.com/content/978-3-540-32060-9.

Hayashi:2006:CCE

[HNKS06b] Nakao Hayashi, Pavel I. Naumkin, Elena I. Kaikina, and Ilya A. Shishmarev. Critical convective equations. Lecture Notes in Mathematics, 1884:323–429, 2006. CODEN LNMAA2. ISBN 3-540-32059-8 (print), 3-540-32060-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-32060-1_4.pdf.

Hayashi:2006:CNE

[HNKS06c] Nakao Hayashi, Pavel I. Naumkin, Elena I. Kaikina, and Ilya A. Shishmarev. Critical nonconvective equations. Lecture Notes in Mathematics, 1884:179–322, 2006. CODEN LNMAA2. ISBN 3-540-32059-8 (print), 3-540-32060-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-32060-1_3.pdf.

Hayashi:2006:PR

[HNKS06d] Nakao Hayashi, Pavel I. Naumkin, Elena I. Kaikina, and Ilya A. Shishmarev. Preliminary results. Lecture Notes in Mathematics, 1884:1–50, 2006. CODEN LNMAA2. ISBN 3-540-32059-8 (print), 3-540-32060-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-32060-1_1.pdf.

Hayashi:2006:SCE

[HNKS06e] Nakao Hayashi, Pavel I. Naumkin, Elena I. Kaikina, and Ilya A. Shishmarev. Subcritical convective equations. Lecture Notes in Mathematics, 1884:513–540, 2006. CODEN LNMAA2. ISBN 3-540-32059-8 (print), 3-540-32060-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-32060-1_6.pdf.

Hayashi:2006:SNE

[HNKS06f] Nakao Hayashi, Pavel I. Naumkin, Elena I. Kaikina, and Ilya A. Shishmarev. Subcritical nonconvective equations. Lecture Notes in Mathematics, 1884:431–512, 2006. CODEN LNMAA2. ISBN 3-540-32059-8 (print), 3-540-32060-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-32060-1_5.pdf.

Hayashi:2006:WN

[HNKS06g] Nakao Hayashi, Pavel I. Naumkin, Elena I. Kaikina, and Ilya A. Shishmarev. Weak nonlinearity. Lecture Notes in Mathematics, 1884:51–178, 2006. CODEN LNMAA2. ISBN 3-540-32059-8 (print), 3-540-32060-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-32060-1_2.pdf.

Hobson:2000:MER

[Hob00] David G. Hobson. Marked excursions and random trees. Lecture Notes in Mathematics, 1729:289–301, 2000. CODEN LN-MAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103808/.

Hollander:2009:FM

[Hol09] Frank Hollander. Front matter. Lecture Notes in Mathematics, 1974:1–13, 2009. CODEN LNMAA2. ISBN 3-642-00332-X (print), 3-642-00333-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-00333-2/1.pdf.

Hora:2003:NVK

[Hor03] Akihito Hora. A noncommutative version of Kerov’s Gaussian limit for the Plancherel measure of the symmetric group. Lecture Notes in Mathematics, 1815:77–88, 2003. CODEN LN-MAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44890-X_4/.

Howe:2002:LNC

[How02] Roger Howe. Lecture notes by Cathy Kriloff. Lecture Notes in Mathematics, 1804:27–69, 2002. CODEN LNMAA2. ISBN 3-540-00224-3 (print), 3-540-36205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36205-0_2/.

Hsu:2000:BM

[Hsu00a] Tim Hsu. Background material. Lecture Notes in Mathematics, 1731:9–26, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103894/.

Hsu:2000:CIQ

[Hsu00b] Tim Hsu. Classical interpretations of quilts. Lecture Notes in Mathematics, 1731:93–98, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103899/.

Hsu:2000:CQ

[Hsu00c] Tim Hsu. The combinatorics of quilts. Lecture Notes in Mathematics, 1731:81–91, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103898/.

Hsu:2000:EQ

[Hsu00d] Tim Hsu. Examples of quilts. Lecture Notes in Mathematics, 1731:71–79, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103897/.

Hsu:2000:FD

[Hsu00e] Tim Hsu. Further directions. Lecture Notes in Mathematics, 1731:157–178, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103905/.

Hsu:2000:I

[Hsu00f] Tim Hsu. Introduction. Lecture Notes in Mathematics, 1731: 1–8, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103893/.

Hsu:2000:MS

[Hsu00g] Tim Hsu. Monodromy systems. Lecture Notes in Mathematics, 1731:127–131, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103902/.

Hsu:2000:NST

[Hsu00h] Tim Hsu. Norton systems and their quilts. Lecture Notes in Mathematics, 1731:55–70, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103896/.

Hsu:2000:PSP

[Hsu00i] Tim Hsu. Presentations and the structure problem. Lecture Notes in Mathematics, 1731:100–117, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103900/.

Hsu:2000:Q

[Hsu00j] Tim Hsu. Quilts. Lecture Notes in Mathematics, 1731:28– 53, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103895/.

Hsu:2000:QCE

[Hsu00k] Tim Hsu. Quilts: Central Extensions, Braid Actions, and Finite Groups, volume 1731 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 185 pp. LCCN QA3 .L28 no. 1731; QA3 .L28 v. 1731; QA177. URL http:/

/link.springer.com/book/10.1007/BFb0103892;   http://
www.springerlink.com/content/978-3-540-45580-6.

Hsu:2000:QGI

[Hsu00l] Tim Hsu. Quilts for groups involved in the monster. Lecture Notes in Mathematics, 1731:133–149, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103903/.

Hsu:2000:SSQ

[Hsu00m] Tim Hsu. Small snug quilts. Lecture Notes in Mathematics, 1731: 119–125, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103901/.

Hsu:2000:SRS

[Hsu00n] Tim Hsu. Some results on the structure problem. Lecture Notes in Mathematics, 1731:151–156, 2000. CODEN LNMAA2. ISBN 3-540-67397-0 (print), 3-540-45580-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103904/.

Hu:2000:UAS

[Hu00] Yaozhong Hu. A unified approach to several inequalities for Gaussian and diffusion measures. Lecture Notes in Mathematics, 1729: 329–335, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103811/.

Huang:2004:LEP

[Hua04] Xiaojun Huang. Local equivalence problems for real submanifolds in complex spaces. Lecture Notes in Mathematics, 1848: 109–163, 2004. CODEN LNMAA2. ISBN 3-540-22358-4 (print), 3-540-44487-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44487-9_3/.

Isozaki:2000:AEF

[IK00] Y. Isozaki and S. Kotani. Asymptotic estimates for the first hitting time of fluctuating additive functionals of Brownian motion. Lecture Notes in Mathematics, 1729:374–387, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103814/.

Iannelli:2004:FAM

[INP04] Mimmo Iannelli, Rainer Nagel, and Susanna Piazzera. Functional Analytic Methods for Evolution Equations, volume 1855 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-23030-0 (print), 3-540-44653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 472 pp. LCCN QA3 .L28 no. 1855. URL http://

link.springer.com/book/10.1007/b100449;   http://www.
springerlink.com/content/978-3-540-44653-8.

Ioffe:2009:SGC

[Iof09] Dmitry Ioffe. Stochastic geometry of classical and quantum Ising models. Lecture Notes in Mathematics, 1970:1–41, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-92796-9_2.pdf.

Isaev:2007:BM

[Isa07a] Alexander Isaev. Back matter. Lecture Notes in Mathematics, 1902:131–143, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-69153-2/1.pdf.

Isaev:2007:Ca

[Isa07b] Alexander Isaev. The case d(M) = n^{2}. Lecture Notes in Mathematics, 1902:29–50, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69153-2_3.pdf.

Isaev:2007:Cb

[Isa07c] Alexander Isaev. The case d(M) = n^{2} − 1, n ≥ 3. Lecture Notes in Mathematics, 1902:51–60, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69153-2_4.pdf.

Isaev:2007:CM

[Isa07d] Alexander Isaev. The case of (2, 3)-manifolds. Lecture Notes in Mathematics, 1902:61–119, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69153-2_5.pdf.

Isaev:2007:FM

[Isa07e] Alexander Isaev. Front matter. Lecture Notes in Mathematics, 1902:I–VIII, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-69153-2/1.pdf.

Isaev:2007:HC

[Isa07f] Alexander Isaev. The homogeneous case. Lecture Notes in Mathematics, 1902:23–28, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69153-2_2.pdf.

Isaev:2007:I

[Isa07g] Alexander Isaev. Introduction. Lecture Notes in Mathematics, 1902:1–22, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69153-2_1.pdf.

Isaev:2007:LAG

[Isa07h] Alexander Isaev. Lectures on the Automorphism Groups of Kobayashi-Hyperbolic Manifolds, volume 1902 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 139 pp. LCCN QA3 .L28 no.

1902. URL http://link.springer.com/book/10.1007/3-540-
69151-0;   http://www.springerlink.com/content/978-3-
540-69153-2.

Isaev:2007:PA

[Isa07i] Alexander Isaev. Proper actions. Lecture Notes in Mathematics, 1902:121–130, 2007. CODEN LNMAA2. ISBN 3-540-69151-0 (print), 3-540-69153-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69153-2_6.pdf.

Ivanoff:2003:SIP

[Iva03] B. Gail Ivanoff. Set-indexed processes: Distributions and weak convergence. Lecture Notes in Mathematics, 1802:85–125, 2003. CODEN LNMAA2. ISBN 3-540-00295-2 (print), 3-540-36259-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36259-3_3/.

Iwaniec:2006:CEC

[Iwa06] Henryk Iwaniec. Conversations on the exceptional character. Lecture Notes in Mathematics, 1891:97–132, 2006. CO-DEN LNMAA2. ISBN 3-540-36363-7 (print), 3-540-36364-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36364-4_3.pdf.

Izhboldin:2004:SNR

[Izh04a] Oleg T. Izhboldin. Some new results concerning isotropy of lowdimensional forms. Lecture Notes in Mathematics, 1835:143– 150, 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40990-8_5/.

Izhboldin:2004:VPN

[Izh04b] Oleg T. Izhboldin. Virtual Pfister neighbors and first Witt index. Lecture Notes in Mathematics, 1835:131–142, 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40990-8_
4/.

Jacod:2003:PCI

[Jac03] Jean Jacod. On processes with conditional independent increments and stable convergence in law. Lecture Notes in Mathematics, 1801:383–401, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36107-7_18/.

Jakubowski:2005:ASA

[Jak05] Adam Jakubowski. An almost sure approximation for the predictable process in the Doob–Meyer decomposition theorem. Lecture Notes in Mathematics, 1857:158–164, 2005. CO-DEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_11.pdf.

Jaksic:2006:TST

[Jak06] Vojkan Jakšić. Topics in spectral theory. Lecture Notes in Mathematics, 1880:235–312, 2006. CODEN LNMAA2. ISBN 3-540-30991-8 (print), 3-540-33922-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-33922-1_6.pdf.

Jefferies:2004:I

[Jef04a] Brian Jefferies. 1. Introduction. Lecture Notes in Mathematics, 1843:1–11, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-70746-2_1/.

Jefferies:2004:WC

[Jef04b] Brian Jefferies. 2. Weyl calculus. Lecture Notes in Mathematics, 1843:13–25, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-70746-2_2/.

Jefferies:2004:CA

[Jef04c] Brian Jefferies. 3. Clifford analysis. Lecture Notes in Mathematics, 1843:27–38, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-70746-2_3/.

Jefferies:2004:FCN

[Jef04d] Brian Jefferies. 4. Functional calculus for noncommuting operators. Lecture Notes in Mathematics, 1843:39–66, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-70746-2_
4/.

Jefferies:2004:JSM

[Jef04e] Brian Jefferies. 5. The joint spectrum of matrices. Lecture Notes in Mathematics, 1843:67–121, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70746-2_5/.

Jefferies:2004:MCS

[Jef04f] Brian Jefferies. 6. The monogenic calculus for sectorial operators. Lecture Notes in Mathematics, 1843:123–155, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-70746-2_
6/.

Jefferies:2004:FOC

[Jef04g] Brian Jefferies. 7. Feynman’s operational calculus. Lecture Notes in Mathematics, 1843:157–171, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-70746-2_7/.

Jefferies:2004:R

[Jef04h] Brian Jefferies. References. Lecture Notes in Mathematics, 1843: 173–179, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-70746-2_8/.

Jefferies:2004:SPN

[Jef04i] Brian Jefferies. Spectral Properties of Noncommuting Operators, volume 1843 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21923-4 (paperback). ISSN 0075-8434. viii + 184 pp. LCCN QA3 .L28 no. 1843; analytic 051102 rcv; QA3 .L28 no.1843.

Johnsen:2004:PMSd

[JK04a] Trygve Johnsen and Andreas Leopold Knutsen. K_{3} Projective Models in Scrolls, volume 1842 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 164 pp. LCCN QA3 .L28 no. 1842. URL

http://link.springer.com/book/10.1007/b97183;   http:
//www.springerlink.com/content/978-3-540-40898-7.

Johnsen:2004:I

[JK04b] Trygve Johnsen and Andreas Leopold Knutsen. 1. Introduction. Lecture Notes in Mathematics, 1842:1–14, 2004. CODEN LN-MAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40898-7_
1/.

Johnsen:2004:BGC

[JK04c] Trygve Johnsen and Andreas Leopold Knutsen. 10. BN general and Clifford general K3 surfaces. Lecture Notes in Mathematics, 1842:121–128, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40898-7_10/. Check math in title??

Johnsen:2004:PMSc

[JK04d] Trygve Johnsen and Andreas Leopold Knutsen. 11. Projective models of K3 surfaces of low genus. Lecture Notes in Mathematics, 1842:129–154, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40898-7_11/.

Johnsen:2004:SAO

[JK04e] Trygve Johnsen and Andreas Leopold Knutsen. 12. Some applications and open questions. Lecture Notes in Mathematics, 1842: 155–158, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40898-7_12/.

Johnsen:2004:SS

[JK04f] Trygve Johnsen and Andreas Leopold Knutsen. 2. Surfaces in scrolls. Lecture Notes in Mathematics, 1842:15–18, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40898-7_
2/.

Johnsen:2004:CIS

[JK04g] Trygve Johnsen and Andreas Leopold Knutsen. 3. The Clifford index of smooth curves in |L| and the definition of the scrolls T (⌋, D, {D_{λ}}). Lecture Notes in Mathematics, 1842:19– 29, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40898-7_3/.

Johnsen:2004:TET

[JK04h] Trygve Johnsen and Andreas Leopold Knutsen. 4. Two existence theorems. Lecture Notes in Mathematics, 1842:31–33, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40898-7_4/.

Johnsen:2004:SLS

[JK04i] Trygve Johnsen and Andreas Leopold Knutsen. 5. The singular locus of the surface S′ and the scroll T . Lecture Notes in Mathematics, 1842:35–45, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40898-7_5/.

Johnsen:2004:PP

[JK04j] Trygve Johnsen and Andreas Leopold Knutsen. 6. Postponed proofs. Lecture Notes in Mathematics, 1842:47–57, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40898-7_6/.

Johnsen:2004:PMSa

[JK04k] Trygve Johnsen and Andreas Leopold Knutsen. 7. Projective models in smooth scrolls. Lecture Notes in Mathematics, 1842: 59–61, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40898-7_7/.

Johnsen:2004:PMSb

[JK04l] Trygve Johnsen and Andreas Leopold Knutsen. 8. Projective models in singular scrolls. Lecture Notes in Mathematics, 1842: 63–98, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40898-7_8/.

Johnsen:2004:MPM

[JK04m] Trygve Johnsen and Andreas Leopold Knutsen. 9. More on projective models in smooth scrolls of K3 surfaces of low Clifford-indices. Lecture Notes in Mathematics, 1842:99–120, 2004. CODEN LNMAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40898-7_9/.

Johnsen:2004:R

[JK04n] Trygve Johnsen and Andreas Leopold Knutsen. References. Lecture Notes in Mathematics, 1842:159–162, 2004. CODEN LN-MAA2. ISBN 3-540-21505-0 (print), 3-540-40898-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40898-7_
13/.

Jorgenson:2005:A

[JL05a] Jay Jorgenson and Serge Lang. GL_{n}(R) action on Pos_{n}(R). Lecture Notes in Mathematics, 1868:1–22, 2005. CODEN LN-MAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11422372_1.pdf.

Jorgenson:2005:ES

[JL05b] Jay Jorgenson and Serge Lang. Pos_{n}(R) and Eisenstein Series, volume 1868 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 168 pp. LCCN QA3 .L28 no. 1868. URL http://

link.springer.com/book/10.1007/b136063;   http://www.
springerlink.com/content/978-3-540-31548-3.

Jorgenson:2005:ESF

[JL05c] Jay Jorgenson and Serge Lang. Eisenstein series first part. Lecture Notes in Mathematics, 1868:107–120, 2005. CODEN LN-MAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11422372_6.pdf.

Jorgenson:2005:ESS

[JL05d] Jay Jorgenson and Serge Lang. Eisenstein series second part. Lecture Notes in Mathematics, 1868:133–162, 2005. CODEN LN-MAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11422372_8.pdf.

Jorgenson:2005:GAE

[JL05e] Jay Jorgenson and Serge Lang. Geometric and analytic estimates. Lecture Notes in Mathematics, 1868:121–132, 2005. CODEN LN-MAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11422372_7.pdf.

Jorgenson:2005:IDO

[JL05f] Jay Jorgenson and Serge Lang. Invariant differential operators on Pos_{n}(R). Lecture Notes in Mathematics, 1868:75–94, 2005. CODEN LNMAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11422372_4.pdf.

Jorgenson:2005:MIQ

[JL05g] Jay Jorgenson and Serge Lang. Measures, integration and quadratic model. Lecture Notes in Mathematics, 1868:23–47, 2005. CODEN LNMAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11422372_2.pdf.

Jorgenson:2005:PDZ

[JL05h] Jay Jorgenson and Serge Lang. Poisson duality and zeta functions. Lecture Notes in Mathematics, 1868:95–106, 2005. CODEN LNMAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11422372_5.pdf.

Jorgenson:2005:SF

[JL05i] Jay Jorgenson and Serge Lang. Special functions on Pos_{n}. Lecture Notes in Mathematics, 1868:49–74, 2005. CODEN LN-MAA2. ISBN 3-540-25787-X (print), 3-540-31548-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11422372_3.pdf.

Jameson:2000:ASO

[JM00] Antony Jameson and Luigi Martinelli. Aerodynamic shape optimization techniques based on control theory. Lecture Notes in Mathematics, 1739:151–221, 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103920/.

Johnson:2003:NAD

[JM03] Russell Johnson and Francesca Mantellini. Non-autonomous differential equations. Lecture Notes in Mathematics, 1822:173– 229, 2003. CODEN LNMAA2. ISBN 3-540-40786-3 (print), 3-540-45204-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45204-1_3/.

Jonsson:2008:AGS

[Jon08a] Jakob Jonsson. Abstract graphs and set systems. Lecture Notes in Mathematics, 1928:19–28, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75859-4_2.pdf.

Jonsson:2008:BG

[Jon08b] Jakob Jonsson. Bipartite graphs. Lecture Notes in Mathematics, 1928:189–204, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_14.pdf.

Jonsson:2008:DT

[Jon08c] Jakob Jonsson. Decision trees. Lecture Notes in Mathematics, 1928:67–86, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_5.pdf.

Jonsson:2008:DP

[Jon08d] Jakob Jonsson. Digraph properties. Lecture Notes in Mathematics, 1928:113–118, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_9.pdf.

Jonsson:2008:DGP

[Jon08e] Jakob Jonsson. Dihedral graph properties. Lecture Notes in Mathematics, 1928:107–112, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75859-4_8.pdf.

Jonsson:2008:DVF

[Jon08f] Jakob Jonsson. Directed variants of forests and bipartite graphs. Lecture Notes in Mathematics, 1928:205–215, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75859-4_15.pdf.

Jonsson:2008:DG

[Jon08g] Jakob Jonsson. Disconnected graphs. Lecture Notes in Mathematics, 1928:245–262, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_18.pdf.

Jonsson:2008:DMT

[Jon08h] Jakob Jonsson. Discrete Morse theory. Lecture Notes in Mathematics, 1928:51–66, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_4.pdf.

Jonsson:2008:FMb

[Jon08i] Jakob Jonsson. Forests and matroids. Lecture Notes in Mathematics, 1928:171–188, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_13.pdf.

Jonsson:2008:FMa

[Jon08j] Jakob Jonsson. Front matter. Lecture Notes in Mathematics, 1928:i–xiv, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-75859-4/1.pdf.

Jonsson:2008:GP

[Jon08k] Jakob Jonsson. Graph properties. Lecture Notes in Mathematics, 1928:99–106, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_7.pdf.

Jonsson:2008:GBD

[Jon08l] Jakob Jonsson. Graphs of bounded degree. Lecture Notes in Mathematics, 1928:151–168, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75859-4_12.pdf.

Jonsson:2008:IO

[Jon08m] Jakob Jonsson. Introduction and overview. Lecture Notes in Mathematics, 1928:3–17, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75859-4_1.pdf.

Jonsson:2008:MGP

[Jon08n] Jakob Jonsson. Main goals and proof techniques. Lecture Notes in Mathematics, 1928:119–124, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75859-4_10.pdf.

Jonsson:2008:M

[Jon08o] Jakob Jonsson. Matchings. Lecture Notes in Mathematics, 1928:127–149, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_11.pdf.

Jonsson:2008:MR

[Jon08p] Jakob Jonsson. Miscellaneous results. Lecture Notes in Mathematics, 1928:87–95, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_6.pdf.

Jonsson:2008:NHG

[Jon08q] Jakob Jonsson. Non-Hamiltonian graphs. Lecture Notes in Mathematics, 1928:233–242, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75859-4_17.pdf.

Jonsson:2008:NG

[Jon08r] Jakob Jonsson. Noncrossing graphs. Lecture Notes in Mathematics, 1928:217–231, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_16.pdf.

Jonsson:2008:CG

[Jon08s] Jakob Jonsson. Not 2-connected graphs. Lecture Notes in Mathematics, 1928:263–273, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75859-4_19.pdf.

Jonsson:2008:CGB

[Jon08t] Jakob Jonsson. Not 3-connected graphs and beyond. Lecture Notes in Mathematics, 1928:275–290, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-75859-4_20.
pdf.

Jonsson:2008:SCG

[Jon08u] Jakob Jonsson. Simplicial Complexes of Graphs, volume 1928 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 378 pp. LCCN QA3 .L28 no. 1928. URL http://link.

springer.com/book/10.1007/978-3-540-75859-4;   http:/
/www.springerlink.com/content/978-3-540-75859-4.

Jonsson:2008:ST

[Jon08v] Jakob Jonsson. Simplicial topology. Lecture Notes in Mathematics, 1928:29–47, 2008. CODEN LNMAA2. ISBN 3-540-75858-5 (print), 3-540-75859-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75859-4_3.pdf.

Joye:2006:IQS

[Joy06a] Alain Joye. Introduction to quantum statistical mechanics. Lecture Notes in Mathematics, 1880:41–67, 2006. CODEN LNMAA2. ISBN 3-540-30991-8 (print), 3-540-33922-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-33922-1_2.pdf.

Joye:2006:ITL

[Joy06b] Alain Joye. Introduction to the theory of linear operators. Lecture Notes in Mathematics, 1880:1–40, 2006. CODEN LNMAA2. ISBN 3-540-30991-8 (print), 3-540-33922-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-33922-1_1.pdf.

Janies:2008:LSP

[JP08] D. Janies and D. Pol. Large-scale phylogenetic analysis of emerging infectious diseases. Lecture Notes in Mathematics, 1922:39– 76, 2008. CODEN LNMAA2. ISBN 3-540-74328-6 (print), 3-540-74331-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74331-6_2.pdf.

Jiang:2004:CDEa

[JQQ04a] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 1. Circulation distribution, entropy production and irreversibility of denumerable Markov chains. Lecture Notes in Mathematics, 1833:11– 44, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40957-1_2/.

Jiang:2004:CDEb

[JQQ04b] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 2. Circulation distribution, entropy production and irreversibility of finite Markov chains with continuous parameter. Lecture Notes in Mathematics, 1833:45–66, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40957-1_3/.

Jiang:2004:GMD

[JQQ04c] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 3. General minimal diffusion process: its construction, invariant measure, entropy production and irreversibility. Lecture Notes in Mathematics, 1833:67–92, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40957-1_4/.

Jiang:2004:MTD

[JQQ04d] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 4. Measuretheoretic discussion on entropy production of diffusion processes and fluctuation-dissipation theorem. Lecture Notes in Mathematics, 1833:93–120, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40957-1_5/.

Jiang:2004:EPR

[JQQ04e] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 5. Entropy production, rotation numbers and irreversibility of diffusion processes on manifolds. Lecture Notes in Mathematics, 1833:121– 148, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40957-1_6/.

Jiang:2004:SHF

[JQQ04f] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 6. On a system of hyperstable frequency locking persistence under white noise. Lecture Notes in Mathematics, 1833:149–158, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40957-1_7/.

Jiang:2004:EPIa

[JQQ04g] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 7. Entropy production and information gain in axiom A systems. Lecture Notes in Mathematics, 1833:159–188, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40957-1_8/.

Jiang:2004:LEH

[JQQ04h] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 8. Lyapunov exponents of hyperbolic attractors. Lecture Notes in Mathematics, 1833:189–214, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40957-1_9/.

Jiang:2004:EPIb

[JQQ04i] Da-Quan Jiang, Min Qian, and Min-Ping Qian. 9. Entropy production, information gain and Lyapunov exponents of random hyperbolic dynamical systems. Lecture Notes in Mathematics, 1833: 215–252, 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40957-1_10/.

Jiang:2004:I

[JQQ04j] Da-Quan Jiang, Min Qian, and Min-Ping Qian. Introduction. Lecture Notes in Mathematics, 1833:1–10, 2004. CODEN LN-MAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40957-1_

Jiang:2004:MTN

[JQQ04k] Da-Quan Jiang, Min Qian, and Min-Ping Qian. Mathematical Theory of Nonequilibrium Steady States: On the Frontier of Probability and Dynamical Systems, volume 1833 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 280 pp. LCCN QA3 .L28 no. 1833. URL http:/

/link.springer.com/book/10.1007/b94615;   http://www.
springerlink.com/content/978-3-540-40957-1.

Jiang:2004:R

[JQQ04l] Da-Quan Jiang, Min Qian, and Min-Ping Qian. References. Lecture Notes in Mathematics, 1833:253–276, 2004. CODEN LN-MAA2. ISBN 3-540-20611-6 (print), 3-540-40957-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40957-1_
11/.

Johnson:2004:SRC

[JS04] W. B. Johnson and G. Schechtman. Several remarks concerning the local theory of L_{p} spaces. Lecture Notes in Mathematics, 1850:139–148, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44489-3_13/.

Jurek:2000:LDS

[JW00] Zbigniew J. Jurek and Liming Wu. Large deviations for some Poisson random integrals. Lecture Notes in Mathematics, 1729: 185–197, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103803/.

Kaczorowski:2006:ATF

[Kac06a] Jerzy Kaczorowski. Axiomatic theory of L-functions: the Selberg class. Lecture Notes in Mathematics, 1891:133–209, 2006. CODEN LNMAA2. ISBN 3-540-36363-7 (print), 3-540-36364-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36364-4_4.pdf.

Kaczorowski:2006:BM

[Kac06b] Jerzy Kaczorowski. Back matter. Lecture Notes in Mathematics, 1891:211–216, 2006. CODEN LNMAA2. ISBN 3-540-36363-7 (print), 3-540-36364-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-36364-4/1.pdf.

Kahn:2004:CNR

[Kah04] Bruno Kahn. Cohomologie non ramifiée des quadriques. (French) []. Lecture Notes in Mathematics, 1835:1–23, 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40990-8_
1/.

Kaji:2008:TDS

[Kaj08] Shunsuke Kaji. On the tail distributions of the supremum and the quadratic variation of a Càdlàg local martingale. Lecture Notes in Mathematics, 1934:401–420, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-77913-1_19.pdf.

Karpenko:2004:IRS

[Kar04a] Nikita A. Karpenko. Izhboldin’s results on stably birational equivalence of quadrics. Lecture Notes in Mathematics, 1835:151– 183, 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40990-8_6/.

Karpenko:2004:MCG

[Kar04b] Nikita A. Karpenko. Motives and Chow groups of quadrics with application to the u-invariant (after Oleg Izhboldin). Lecture Notes in Mathematics, 1835:103–129, 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40990-8_3/.

Karandikar:2006:ASC

[Kar06] Rajeeva L. Karandikar. On almost sure convergence results in stochastic calculus. Lecture Notes in Mathematics, 1874:137–147, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_11.pdf.

Kashiwara:2008:EDC

[Kas08] Masaki Kashiwara. Equivariant derived category and representation of real semisimple Lie groups. Lecture Notes in Mathematics, 1931:137–234, 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-76892-0_3.pdf.

Kawohl:2000:SNS

[Kaw00] Bernd Kawohl. Some nonconvex shape optimization problems. Lecture Notes in Mathematics, 1740:7–46, 2000. CODEN LN-MAA2. ISBN 3-540-67971-5 (print), 3-540-44486-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0106741/.

Kazhikhov:2006:AWL

[Kaz06] Alexandre V. Kazhikhov. Approximation of weak limits and related problems. Lecture Notes in Mathematics, 1871:75–100, 2006. CODEN LNMAA2. ISBN 3-540-28586-5 (print), 3-540-32454-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11545989_3.pdf.

Kechris:2004:TOE

[Kec04] Alexander S. Kechris. Topics in Orbit Equivalence, volume 1852 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22603-6 (print), 3-540-44508-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 134 pp. LCCN QA3 .L28 no. 1852. URL http:/

/link.springer.com/book/10.1007/b99421;   http://www.
springerlink.com/content/978-3-540-44508-1.

Keller:2000:ACT

[Kel00a] Karsten Keller. Abstract and concrete theory. Lecture Notes in Mathematics, 1732:141–180, 2000. CODEN LNMAA2. ISBN 3-540-67434-9 (print), 3-540-45589-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104003/.

Keller:2000:AJS

[Kel00b] Karsten Keller. Abstract Julia sets. Lecture Notes in Mathematics, 1732:25–71, 2000. CODEN LNMAA2. ISBN 3-540-67434-9 (print), 3-540-45589-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0104001/.

Keller:2000:AMS

[Kel00c] Karsten Keller. The abstract Mandelbrot set. Lecture Notes in Mathematics, 1732:73–139, 2000. CODEN LNMAA2. ISBN 3-540-67434-9 (print), 3-540-45589-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104002/.

Keller:2000:I

[Kel00d] Karsten Keller. Introduction. Lecture Notes in Mathematics, 1732:1–23, 2000. CODEN LNMAA2. ISBN 3-540-67434-9 (print), 3-540-45589-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104000/.

Keller:2000:IFJ

[Kel00e] Karsten Keller. Invariant Factors, Julia Equivalences and the (Abstract) Mandelbrot Set, volume 1732 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67434-9 (print), 3-540-45589-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 206 pp. LCCN QA3 .L28 no. 1732.

URL http://link.springer.com/book/10.1007/BFb0103999;
 http://www.springerlink.com/content/978-3-540-45589-
9.

Kohatsu-Higa:2007:MIT

[KH07] Arturo Kohatsu-Higa. Models for insider trading with finite utility. Lecture Notes in Mathematics, 1919:103–171, 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73327-0_3.pdf.

Khesin:2009:GHD

[Khe09] Boris Khesin. The group and Hamiltonian descriptions of hydrodynamical systems. Lecture Notes in Mathematics, 1973:139–155, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00837-5_4.pdf.

Khoshnevisan:2000:SIR

[Kho00] D. Khoshnevisan. On sums of iid random variables indexed by N parameters. Lecture Notes in Mathematics, 1729:151–156, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103800/.

Khoshnevisan:2003:CZS

[Kho03] Davar Khoshnevisan. The codimension of the zeros of a stable process in random scenery. Lecture Notes in Mathematics, 1832: 236–245, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_9/.

Khorunzhiy:2008:EMR

[Kho08] Oleksiy Khorunzhiy. Estimates for moments of random matrices with Gaussian elements. Lecture Notes in Mathematics, 1934:51– 92, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_3.pdf.

Khoshnevisan:2009:PSP

[Kho09] Davar Khoshnevisan. A primer on stochastic partial differential equations. Lecture Notes in Mathematics, 1962:1–38, 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85994-9_1.pdf.

Kiechle:2002:P

[Kie02a] Hubert Kiechle. 1. Preliminaries. Lecture Notes in Mathematics, 1778:7–22, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45817-3_2/.

Kiechle:2002:RVA

[Kie02b] Hubert Kiechle. 10. Relativistic velocity addition. Lecture Notes in Mathematics, 1778:137–142, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45817-3_11/.

Kiechle:2002:LGL

[Kie02c] Hubert Kiechle. 11. K-loops from the general linear groups over rings. Lecture Notes in Mathematics, 1778:143–150, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45817-3_12/.

Kiechle:2002:D

[Kie02d] Hubert Kiechle. 12. Derivations. Lecture Notes in Mathematics, 1778:151–164, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45817-3_13/.

Kiechle:2002:LLT

[Kie02e] Hubert Kiechle. 2. Left loops and transversals. Lecture Notes in Mathematics, 1778:23–42, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45817-3_3/.

Kiechle:2002:LIP

[Kie02f] Hubert Kiechle. 3. The left inverse property and Kikkawa loops. Lecture Notes in Mathematics, 1778:43–52, 2002. CODEN LN-MAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45817-3_
4/.

Kiechle:2002:IT

[Kie02g] Hubert Kiechle. 4. Isotopy theory. Lecture Notes in Mathematics, 1778:53–58, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45817-3_5/.

Kiechle:2002:NAG

[Kie02h] Hubert Kiechle. 5. Nuclei and the autotopism group. Lecture Notes in Mathematics, 1778:59–64, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45817-3_6/.

Kiechle:2002:BLL

[Kie02i] Hubert Kiechle. 6. Bol loops and K-loops. Lecture Notes in Mathematics, 1778:65–81, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45817-3_7/.

Kiechle:2002:FGM

[Kie02j] Hubert Kiechle. 7. Frobenius groups with many involutions. Lecture Notes in Mathematics, 1778:83–102, 2002. CODEN LN-MAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45817-3_
8/.

Kiechle:2002:LF

[Kie02k] Hubert Kiechle. 8. Loops with fibrations. Lecture Notes in Mathematics, 1778:103–106, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45817-3_9/.

Kiechle:2002:LCG

[Kie02l] Hubert Kiechle. 9. K-loops from classical groups over ordered fields. Lecture Notes in Mathematics, 1778:107–136, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45817-3_10/.

Kiechle:2002:A

[Kie02m] Hubert Kiechle. Appendix. Lecture Notes in Mathematics, 1778: 165–170, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45817-3_14/.

Kiechle:2002:Ib

[Kie02n] Hubert Kiechle. Index. Lecture Notes in Mathematics, 1778: 181–186, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45817-3_16/.

Kiechle:2002:Ia

[Kie02o] Hubert Kiechle. Introduction. Lecture Notes in Mathematics, 1778:1–5, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45817-3_1/.

Kiechle:2002:R

[Kie02p] Hubert Kiechle. References. Lecture Notes in Mathematics, 1778: 171–180, 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45817-3_15/.

Kiechle:2002:TL

[Kie02q] Hubert Kiechle. Theory of K-Loops, volume 1778 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43262-0 (print), 3-540-45817-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 186 pp. LCCN QA3 .L28 no. 1778; QA3 L4 V.1778. URL http://catdir.loc.gov/

catdir/enhancements/fy0817/2002022101-d.html;   http:
//catdir.loc.gov/catdir/enhancements/fy0817/2002022101-
t.html;   http://link.springer.com/book/10.1007/b83276;
 http://www.springerlink.com/content/978-3-540-45817-
3;   http://www.springerlink.com/openurl.asp?genre=book\
&isbn=978-3-540-43262-3;   http://www.springerlink.com/
openurl.asp?genre=issue&issn=0075-8434&volume=1778.

Koldobsky:2000:AVS

[KL00] A. Koldobsky and M. Lifshits. Average volume of sections of star bodies. Lecture Notes in Mathematics, 1745:119–146, 2000.

CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0107212/.

Kerler:2001:CCH

[KL01a] Thomas Kerler and Volodymyr V. Lyubashenko. Coends and construction of Hopf algebras. Lecture Notes in Mathematics, 1765:261–282, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44625-7_6/.

Kerler:2001:CTD

[KL01b] Thomas Kerler and Volodymyr V. Lyubashenko. Construction of TQFT-double functors. Lecture Notes in Mathematics, 1765: 283–311, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44625-7_7/.

Kerler:2001:DCD

[KL01c] Thomas Kerler and Volodymyr V. Lyubashenko. Double categories and double functors. Lecture Notes in Mathematics, 1765: 343–352, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44625-7_10/.

Kerler:2001:DCF

[KL01d] Thomas Kerler and Volodymyr V. Lyubashenko. The double category of framed, relative 3-cobordisms. Lecture Notes in Mathematics, 1765:15–95, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44625-7_2/.

Kerler:2001:QFT

[KL01e] Thomas Kerler and Volodymyr V. Lyubashenko. From quantum field theory to axiomatics. Lecture Notes in Mathematics, 1765: 335–342, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44625-7_9/.

Kerler:2001:GMF

[KL01f] Thomas Kerler and Volodymyr V. Lyubashenko. Generalization of a modular functor. Lecture Notes in Mathematics, 1765:313– 334, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44625-7_8/.

Kerler:2001:ISR

[KL01g] Thomas Kerler and Volodymyr V. Lyubashenko. Introduction and summary of results. Lecture Notes in Mathematics, 1765: 1–14, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44625-7_1/.

Kerler:2001:IBT

[KL01h] Thomas Kerler and Volodymyr V. Lyubashenko. Isomorphism between tangle and Cobordism double categories. Lecture Notes in Mathematics, 1765:173–215, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44625-7_4/.

Kerler:2001:MCM

[KL01i] Thomas Kerler and Volodymyr V. Lyubashenko. Monoidal categories and monoidal 2-categories. Lecture Notes in Mathematics, 1765:217–259, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44625-7_5/.

Kerler:2001:NST

[KL01j] Thomas Kerler and Volodymyr V. Lyubashenko. Non-Semisimple Topological Quantum Field Theories for 3-Manifolds with Corners, volume 1765 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 379 pp. LCCN QA3 .L28 no. 1765. URL http:/

/link.springer.com/book/10.1007/b82618;   http://www.
springerlink.com/content/978-3-540-44625-5.

Kerler:2001:TCP

[KL01k] Thomas Kerler and Volodymyr V. Lyubashenko. Tanglecategories and presentation of cobordisms. Lecture Notes in Mathematics, 1765:97–172, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44625-7_3/.

Kerler:2001:TT

[KL01l] Thomas Kerler and Volodymyr V. Lyubashenko. Thick tangles. Lecture Notes in Mathematics, 1765:353–368, 2001. CODEN LNMAA2. ISBN 3-540-42416-4 (print), 3-540-44625-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44625-7_11/.

Kauffman:2009:TRK

[KL09] Louis H. Kauffman and Sofia Lambropoulou. Tangles, rational knots and DNA. Lecture Notes in Mathematics, 1973:99–138, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00837-5_3.pdf.

Klartag:2000:RMS

[Kla00] B. Klartag. Remarks on Minkowski symmetrizations. Lecture Notes in Mathematics, 1745:109–117, 2000. CODEN LN-MAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107211/.

Klartag:2004:JTE

[Kla04] B. Klartag. On John-type ellipsoids. Lecture Notes in Mathematics, 1850:149–158, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44489-3_14/.

Klartag:2007:MGI

[Kla07] Boaz Klartag. Marginals of geometric inequalities. Lecture Notes in Mathematics, 1910:133–166, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72053-9_9.pdf.

Kechris:2004:OE

[KM04a] Alexander S. Kechris and Benjamin D. Miller. I. orbit equivalence. Lecture Notes in Mathematics, 1852:1–6, 2004. CODEN LNMAA2. ISBN 3-540-22603-6 (print), 3-540-44508-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44508-1_
1/.

Kechris:2004:IAH

[KM04b] Alexander S. Kechris and Benjamin D. Miller. Ii. amenability and hyperfiniteness. Lecture Notes in Mathematics, 1852: 7–53, 2004. CODEN LNMAA2. ISBN 3-540-22603-6 (print), 3-540-44508-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44508-1_2/.

Kechris:2004:ICE

[KM04c] Alexander S. Kechris and Benjamin D. Miller. Iii. costs of equivalence relations and groups. Lecture Notes in Mathematics, 1852: 55–128, 2004. CODEN LNMAA2. ISBN 3-540-22603-6 (print), 3-540-44508-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44508-1_3/.

Kechris:2004:I

[KM04d] Alexander S. Kechris and Benjamin D. Miller. Index. Lecture Notes in Mathematics, 1852:131–134, 2004. CODEN LNMAA2. ISBN 3-540-22603-6 (print), 3-540-44508-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44508-1_5/.

Kechris:2004:R

[KM04e] Alexander S. Kechris and Benjamin D. Miller. References. Lecture Notes in Mathematics, 1852:129–130, 2004. CODEN LN-MAA2. ISBN 3-540-22603-6 (print), 3-540-44508-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-540-44508-1_4.

Kuchta:2001:MPT

[KMS01] Małgorzata Kuchta, Michał Morayne, and Sławomir Solecki. A martingale proof of the theorem by Jessen, Marcinkiewicz and Zygmund on strong differentiation of integrals. Lecture Notes in Mathematics, 1755:158–161, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44671-2_11/.

Koepf:2003:CAA

[Koe03] Wolfram Koepf. Computer algebra algorithms for orthogonal polynomials and special functions. Lecture Notes in Mathematics, 1817:1–24, 2003. CODEN LNMAA2. ISBN 3-540-40375-2 (print), 3-540-44945-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44945-0_1/.

Kolokoltsov:2000:BVP

[Kol00a] Vassili N. Kolokoltsov. Boundary value problem for Hamiltonian systems. Lecture Notes in Mathematics, 1724:40–96, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112491/.

Kolokoltsov:2000:CSD

[Kol00b] Vassili N. Kolokoltsov. Complex stochastic diffusion or stochastic Schrödinger equation. Lecture Notes in Mathematics, 1724: 223–238, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112496/.

Kolokoltsov:2000:GD

[Kol00c] Vassili N. Kolokoltsov. Gaussian diffusions. Lecture Notes in Mathematics, 1724:17–39, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112490/.

Kolokoltsov:2000:I

[Kol00d] Vassili N. Kolokoltsov. Introduction. Lecture Notes in Mathematics, 1724:1–16, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112489/.

Kolokoltsov:2000:IDD

[Kol00e] Vassili N. Kolokoltsov. Invariant degenerate diffusion on cotangent bundles. Lecture Notes in Mathematics, 1724:136–145, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112493/.

Kolokoltsov:2000:PIS

[Kol00f] Vassili N. Kolokoltsov. Path integration for the Schrödinger, heat and complex diffusion equations. Lecture Notes in Mathematics, 1724:255–279, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112498/.

Kolokoltsov:2000:SAD

[Kol00g] Vassili N. Kolokoltsov. Semiclassical Analysis for Diffusions and Stochastic Processes, volume 1724 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 347 pp. LCCN QA3 .L28 no. 1724.

URL http://link.springer.com/book/10.1007/BFb0112488;
 http://www.springerlink.com/content/978-3-540-46587-
4.

Kolokoltsov:2000:SAR

[Kol00h] Vassili N. Kolokoltsov. Semiclassical approximation for regular diffusion. Lecture Notes in Mathematics, 1724:97–135, 2000. CO-DEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112492/.

Kolokoltsov:2000:SAL

[Kol00i] Vassili N. Kolokoltsov. Semiclassical asymptotics for the localised Feller–Courrège processes. Lecture Notes in Mathematics, 1724: 191–222, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112495/.

Kolokoltsov:2000:STS

[Kol00j] Vassili N. Kolokoltsov. Some topics in semiclassical spectral analysis. Lecture Notes in Mathematics, 1724:239–254, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112497/.

Kolokoltsov:2000:TPD

[Kol00k] Vassili N. Kolokoltsov. Transition probability densities for stable jump-diffusions. Lecture Notes in Mathematics, 1724:146– 190, 2000. CODEN LNMAA2. ISBN 3-540-66972-8 (print), 3-540-46587-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0112494/.

Konno:2008:BM

[Kon08a] Norie Konno. Back matter. Lecture Notes in Mathematics, 1954: 453–463, 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-69365-9/1.pdf.

Konno:2008:QW

[Kon08b] Norie Konno. Quantum walks. Lecture Notes in Mathematics, 1954:309–452, 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69365-9_7.pdf.

Kotani:2006:COD

[Kot06] Shinichi Kotani. On a condition that one-dimensional diffusion processes are martingales. Lecture Notes in Mathematics, 1874:149–156, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35513-7_12.pdf.

Kotecky:2009:MCM

[Kot09] Roman Kotecký. Methods of Contemporary Mathematical Statistical Physics, volume 1970 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 343 pp. LCCN QA3 .L28 no. 1970.

URL http://link.springer.com/book/10.1007/978-3-540-
92796-9;   http://www.springerlink.com/content/978-3-
540-92796-9.

Kurtz:2001:CAA

[KP01] David Kurtz and Anthony Phan. Correction à un article d’Attal et Émery sur les martingales d’Azéma bidimendionnelles. (French) []. Lecture Notes in Mathematics, 1755:120– 122, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44671-2_6/.

Kyprianou:2005:MRS

[KP05] Andreas E. Kyprianou and Zbigniew Palmowski. A martingale review of some fluctuation theory for spectrally negative Lévy processes. Lecture Notes in Mathematics, 1857:16–29, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_3.pdf.

Kyprianou:2008:FSN

[KP08] Andreas E. Kyprianou and Zbigniew Palmowski. Fluctuations of spectrally negative Markov additive processes. Lecture Notes in Mathematics, 1934:121–135, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-77913-1_5.pdf.

Khoshnevisan:2009:MSP

[KRA09] Davar Khoshnevisan and Firas Rassoul-Agha. A Minicourse on Stochastic Partial Differential Equations, volume 1962 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 216 pp. LCCN

QA3 .L28 no. 1962. URL http://link.springer.com/book/
10.1007/978-3-540-85994-9;   http://www.springerlink.
com/content/978-3-540-85994-9.

Krell:2009:SSB

[Kre09] Nathalie Krell. Self-similar branching Markov chains. Lecture Notes in Mathematics, 1979:261–280, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-01763-6_10.
pdf.

Khoshnevisan:2000:FSP

[KS00] Davar Khoshnevisan and Zhan Shi. Fast sets and points for fractional Brownian motion. Lecture Notes in Mathematics, 1729: 393–416, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103816/.

Kabanov:2001:EMM

[KS01a] Yuri Kabanov and Christophe Sticker. On equivalent martingale measures with bounded densities. Lecture Notes in Mathematics, 1755:139–148, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44671-2_8/.

Kabanov:2001:TNN

[KS01b] Yuri Kabanov and Christophe Sticker. A teacher’s note on no-arbitrage criteria. Lecture Notes in Mathematics, 1755:149– 152, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44671-2_9/.

Kabanov:2003:TSP

[KS03] Yuri Kabanov and Christophe Stricker. On the true submartingale property, d’après Schachermayer. Lecture Notes in Mathematics, 1801:413–414, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36107-7_20.

Kabanov:2005:RTN

[KS05] Yuri Kabanov and Christophe Stricker. Remarks on the true noarbitrage property. Lecture Notes in Mathematics, 1857:186–194, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_13.pdf.

Kabanov:2006:DMW

[KS06a] Yuri Kabanov and Christophe Stricker. The Dalang–Morton– Willinger Theorem under delayed and restricted information. Lecture Notes in Mathematics, 1874:209–213, 2006. CO-DEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_16.pdf.

Kimmel:2006:CTA

[KS06b] Marek Kimmel and Andrzej Swierniak. Control theory approach to cancer chemotherapy: Benefiting from phase dependence and overcoming drug resistance. Lecture Notes in Mathematics, 1872: 185–221, 2006. CODEN LNMAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11561606_5.pdf.

Kyprianou:2007:NCV

[KS07] Andreas E. Kyprianou and Budhi A. Surya. A note on a change of variable formula with local time-space for Lévy processes of bounded variation. Lecture Notes in Mathematics, 1899:97–104, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_3.pdf.

Kubatko:2008:IPT

[Kub08] L. S. Kubatko. Inference of phylogenetic trees. Lecture Notes in Mathematics, 1922:1–38, 2008. CODEN LNMAA2. ISBN 3-540-74328-6 (print), 3-540-74331-6 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
content/pdf/10.1007/978-3-540-74331-6_1.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-540-74331-6/
1.pdf.

Kuijlaars:2003:RHA

[Kui03] Arno B. J. Kuijlaars. Riemann–Hilbert analysis for orthogonal polynomials. Lecture Notes in Mathematics, 1817:167– 210, 2003. CODEN LNMAA2. ISBN 3-540-40375-2 (print), 3-540-44945-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44945-0_5/.

Kuksin:2002:KPF

[Kuk02] Sergei B. Kuksin. KAM-persistence of finite-gap solutions. Lecture Notes in Mathematics, 1784:61–123, 2002. CODEN LN-MAA2. ISBN 3-540-43726-6 (print), 3-540-47928-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-47928-4_
2/.

Kuksin:2005:LHM

[Kuk05] Sergei Kuksin. Lectures on Hamiltonian methods in nonlinear PDEs. Lecture Notes in Mathematics, 1861:143–164, 2005. CODEN LNMAA2. ISBN 3-540-24064-0 (print), 3-540-31541-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31541-4_3.pdf.

Kulczycki:2009:EES

[Kul09] T. Kulczycki. Eigenvalues and eigenfunctions for stable processes. Lecture Notes in Mathematics, 1980:73–86, 2009. CO-DEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-02141-1_4.pdf.

Kummerer:2006:QMP

[Küm06] Burkhard Kümmerer. Quantum Markov processes and applications in physics. Lecture Notes in Mathematics, 1866:259– 330, 2006. CODEN LNMAA2. ISBN 3-540-24407-7 (print), 3-540-32385-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11376637_4.pdf.

Kunze:2000:BUP

[Kun00a] Markus Kunze. Bounded, unbounded, periodic, and almost periodic solutions. Lecture Notes in Mathematics, 1744:19–61, 2000.

CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103846/.

Kunze:2000:FTN

[Kun00b] Markus Kunze. Further topics and notes. Lecture Notes in Mathematics, 1744:203–209, 2000. CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103852/.

Kunze:2000:I

[Kun00c] Markus Kunze. Introduction. Lecture Notes in Mathematics, 1744:1–6, 2000. CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103844/.

Kunze:2000:LEN

[Kun00d] Markus Kunze. Lyapunov exponents for non-smooth dynamical systems. Lecture Notes in Mathematics, 1744:63–140, 2000. CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103847/.

Kunze:2000:MMN

[Kun00e] Markus Kunze. Melnikov’s method for non-smooth dynamical systems. Lecture Notes in Mathematics, 1744:197–201, 2000. CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103851/.

Kunze:2000:NSD

[Kun00f] Markus Kunze. Non-Smooth Dynamical Systems, volume 1744 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 228 pp. LCCN QA3 .L28 no. 1744; QA3 .L28 vol.

1744 QA845. URL http://link.springer.com/book/10.1007/
BFb0103843;   http://www.loc.gov/catdir/enhancements/
fy0816/00064115-d.html;   http://www.loc.gov/catdir/enhancements/

fy0816/00064115-t.html;   http://www.springerlink.com/
content/978-3-540-44441-1.

Kunze:2000:ACI

[Kun00g] Markus Kunze. On the application of Conley index theory to non-smooth dynamical systems. Lecture Notes in Mathematics, 1744:141–162, 2000. CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103848/.

Kunze:2000:AKT

[Kun00h] Markus Kunze. On the application of KAM theory to non-smooth dynamical systems. Lecture Notes in Mathematics, 1744:163– 184, 2000. CODEN LNMAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103849/.

Kunze:2000:PNS

[Kun00i] Markus Kunze. Planar non-smooth dynamical systems. Lecture Notes in Mathematics, 1744:185–196, 2000. CODEN LN-MAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103850/.

Kunze:2000:SGT

[Kun00j] Markus Kunze. Some general theory of differential inclusions. Lecture Notes in Mathematics, 1744:7–18, 2000. CODEN LN-MAA2. ISBN 3-540-67993-6 (print), 3-540-44441-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103845/.

Kurtz:2001:CMA

[Kur01] David Kurtz. Une caractérization des martingales d’Azéma bidimensionnelles de type (II). (French) []. Lecture Notes in Mathematics, 1755:98–119, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44671-2_5/.

Kustermans:2005:LCQ

[Kus05] Johan Kustermans. Locally compact quantum groups. Lecture Notes in Mathematics, 1865:99–180, 2005. CODEN LN-MAA2. ISBN 3-540-24406-9 (print), 3-540-31450-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11376569_2.pdf.

Kutyniok:2007:ADW

[Kut07a] Gitta Kutyniok. Affine Density in Wavelet Analysis, volume 1914 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 138 pp. LCCN QA3 .L28 no. 1914. URL http://link.

springer.com/book/10.1007/978-3-540-72949-5;   http:/
/www.springerlink.com/content/978-3-540-72949-5.

Kutyniok:2007:BM

[Kut07b] Gitta Kutyniok. Back matter. Lecture Notes in Mathematics, 1914:127–142, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-72949-5/1.pdf.

Kutyniok:2007:FM

[Kut07c] Gitta Kutyniok. Front matter. Lecture Notes in Mathematics, 1914:I–XII, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-72949-5/1.pdf.

Kutyniok:2007:HAP

[Kut07d] Gitta Kutyniok. Homogeneous approximation property. Lecture Notes in Mathematics, 1914:87–104, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-72949-5_6.
pdf.

Kutyniok:2007:I

[Kut07e] Gitta Kutyniok. Introduction. Lecture Notes in Mathematics, 1914:1–10, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72949-5_1.pdf.

Kutyniok:2007:QDCa

[Kut07f] Gitta Kutyniok. Qualitative density conditions. Lecture Notes in Mathematics, 1914:35–57, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72949-5_4.pdf.

Kutyniok:2007:QDCb

[Kut07g] Gitta Kutyniok. Quantitative density conditions. Lecture Notes in Mathematics, 1914:59–86, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72949-5_5.pdf.

Kutyniok:2007:WGF

[Kut07h] Gitta Kutyniok. Wavelet and Gabor frames. Lecture Notes in Mathematics, 1914:11–20, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72949-5_2.pdf.

Kutyniok:2007:WAD

[Kut07i] Gitta Kutyniok. Weighted affine density. Lecture Notes in Mathematics, 1914:21–33, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72949-5_3.pdf.

Kutyniok:2007:WBD

[Kut07j] Gitta Kutyniok. Weighted Beurling density and shift-invariant Gabor systems. Lecture Notes in Mathematics, 1914:105–125, 2007. CODEN LNMAA2. ISBN 3-540-72916-X (print), 3-540-72949-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72949-5_7.pdf.

Kuznetsov:2006:OPS

[Kuz06] Vadim B. Kuznetsov. Orthogonal polynomials and separation of variables. Lecture Notes in Mathematics, 1883:229–254, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36716-1_5.pdf.

Kravanja:2000:CZAa

[KV00a] Peter Kravanja and Marc Van Barel. Clusters of zeros of analytic functions. Lecture Notes in Mathematics, 1727:61–81, 2000. CODEN LNMAA2. ISBN 3-540-67162-5 (print), 3-540-46518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103929/.

Kravanja:2000:CZAb

[KV00b] Peter Kravanja and Marc Van Barel. Computing the Zeros of Analytic Functions, volume 1727 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67162-5 (print), 3-540-46518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 111 pp. LCCN QA3 .L28 no. 1727.

URL http://link.springer.com/book/10.1007/BFb0103927;
 http://www.springerlink.com/content/978-3-540-46518-
8.

Kravanja:2000:SAE

[KV00c] Peter Kravanja and Marc Van Barel. Systems of analytic equations. Lecture Notes in Mathematics, 1727:91–103, 2000. CODEN LNMAA2. ISBN 3-540-67162-5 (print), 3-540-46518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103931/.

Kravanja:2000:ZPM

[KV00d] Peter Kravanja and Marc Van Barel. Zeros and poles of meromorphic functions. Lecture Notes in Mathematics, 1727:83–89, 2000. CODEN LNMAA2. ISBN 3-540-67162-5 (print), 3-540-46518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103930/.

Kravanja:2000:ZAF

[KV00e] Peter Kravanja and Marc Van Barel. Zeros of analytic functions. Lecture Notes in Mathematics, 1727:1–59, 2000. CODEN LNMAA2. ISBN 3-540-67162-5 (print), 3-540-46518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103928/.

Koelink:2003:OPS

[KV03] Erik Koelink and Walter Van Assche, editors. Orthogonal Polynomials and Special Functions: Leuven 2002, volume 1817 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40375-2 (print), 3-540-44945-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1817.

URL http://link.springer.com/book/10.1007/b12166;
http://www.springerlink.com/content/978-3-540-44945-4.

Kunstmann:2004:MRP

[KW04] Peer C. Kunstmann and Lutz Weis. Maximal L_{p}-regularity for parabolic equations, Fourier multiplier theorems and H^{∞}- functional calculus. Lecture Notes in Mathematics, 1855:65– 311, 2004. CODEN LNMAA2. ISBN 3-540-23030-0 (print), 3-540-44653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44653-8_2/.

Knebusch:2002:A

[KZ02a] Manfred Knebusch and Digen Zhang. Appendix. Lecture Notes in Mathematics, 1791:251–256, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45625-4_6/.

Knebusch:2002:CBM

[KZ02b] Manfred Knebusch and Digen Zhang. Chapter I: Basics on Manis valuations and Prüfer extensions. Lecture Notes in Mathematics, 1791:9–81, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45625-4_3/.

Knebusch:2002:CIM

[KZ02c] Manfred Knebusch and Digen Zhang. Chapter II: Multiplicative ideal theory. Lecture Notes in Mathematics, 1791:83–176, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45625-4_4/.

Knebusch:2002:CIP

[KZ02d] Manfred Knebusch and Digen Zhang. Chapter III: PM-valuations and valuations of weaker type. Lecture Notes in Mathematics, 1791:177–250, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45625-4_5/.

Knebusch:2002:I

[KZ02e] Manfred Knebusch and Digen Zhang. Introduction. Lecture Notes in Mathematics, 1791:1–6, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45625-4_1/.

Knebusch:2002:MVP

[KZ02f] Manfred Knebusch and Digen Zhang. Manis Valuations and Prüfer Extensions I: A New Chapter in Commutative Algebra, volume 1791 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 267 pp. LCCN QA3 .L28 no. 1791. URL http:/

/link.springer.com/book/10.1007/b84018;   http://www.
springerlink.com/content/978-3-540-45625-4.

Knebusch:2002:R

[KZ02g] Manfred Knebusch and Digen Zhang. References. Lecture Notes in Mathematics, 1791:257–262, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45625-4_7/.

Knebusch:2002:SSI

[KZ02h] Manfred Knebusch and Digen Zhang. Subject and symbol index. Lecture Notes in Mathematics, 1791:263–267, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45625-4_
8/.

Knebusch:2002:S

[KZ02i] Manfred Knebusch and Digen Zhang. Summary. Lecture Notes in Mathematics, 1791:7, 2002. CODEN LNMAA2. ISBN 3-540-43951-X (print), 3-540-45625-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45625-4_2.

Lachal:2003:ATE

[Lac03] Aimé Lachal. Application de la théorie des excursions à l’intégrale du mouvement brownien. (French) []. Lecture Notes in Mathematics, 1832:109–195, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40004-2_6/.

Lachowicz:2008:LBM

[Lac08] Mirosław Lachowicz. Lins between microscopic and macroscopic descriptions. Lecture Notes in Mathematics, 1940:201–267, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78362-6_4.pdf.

Lanconelli:2009:HKS

[Lan09] Ermanno Lanconelli. Heat kernels in sub-Riemannian settings. Lecture Notes in Mathematics, 1977:35–61, 2009. CO-DEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01674-5_2.pdf.

Lasiecka:2004:OCP

[Las04] Irena Lasiecka. Optimal control problems and Riccati equations for systems with unbounded controls and partially analytic generators-applications to boundary and point control problems. Lecture Notes in Mathematics, 1855:313–369, 2004. CODEN LNMAA2. ISBN 3-540-23030-0 (print), 3-540-44653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44653-8_
3/.

Laurent:2009:MEF

[Lau09] Miclo Laurent. Monotonicity of the extremal functions for onedimensional inequalities of logarithmic Sobolev type. Lecture Notes in Mathematics, 1979:103–130, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-01763-6_2.
pdf.

Li:2008:CTA

[LB08] Jia Li and Fred Brauer. Continuous-time age-structured models in population dynamics and epidemiology. Lecture Notes in Mathematics, 1945:205–227, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78911-6_9.pdf.

Lefevre:2009:BM

[LB09a] Julien Lefèvre and Sylvain Baillet. Back matter. Lecture Notes in Mathematics, 1983:1–8, 2009. CODEN LNMAA2. ISBN 3-642-03443-8 (print), 3-642-03444-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-03444-2/1.pdf.

Lefevre:2009:EVF

[LB09b] Julien Lefèvre and Sylvain Baillet. Estimation of velocity fields and propagation on non-Euclidian domains: Application to the exploration of cortical spatiotemporal dynamics. Lecture Notes in Mathematics, 1983:203–226, 2009. CODEN LNMAA2. ISBN 3-642-03443-8 (print), 3-642-03444-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-03444-2_5.pdf.

Loday:2001:DRO

[LCFG01a] Jean-Louis Loday, Frédéric Chapoton, Alessandra Frabetti, and François Goichot. Dialgebras and Related Operads, volume 1763 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42194-7 (print), 3-540-45328-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 133 pp. LCCN QA3 .L28 no. 1763. URL http:/

/link.springer.com/book/10.1007/b80864;   http://www.
springerlink.com/content/978-3-540-45328-4.

Loday:2001:I

[LCFG01b] Jean-Louis Loday, Frédéric Chapoton, Alessandra Frabetti, and François Goichot. Introduction. Lecture Notes in Mathematics, 1763:1–6, 2001. CODEN LNMAA2. ISBN 3-540-42194-7 (print), 3-540-45328-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45328-8_1/.

Lyons:2007:BM

[LCL07a] Terry J. Lyons, Michael Caruana, and Thierry Lévy. Back matter. Lecture Notes in Mathematics, 1908:95–115, 2007. CODEN LNMAA2. ISBN 3-540-71284-4 (print), 3-540-71285-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-71285-5/
1.pdf.

Lyons:2007:DEDa

[LCL07b] Terry J. Lyons, Michael Caruana, and Thierry Lévy. Differential equations driven by moderately irregular signals. Lecture Notes in Mathematics, 1908:1–24, 2007. CODEN LNMAA2. ISBN 3-540-71284-4 (print), 3-540-71285-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71285-5_1.pdf.

Lyons:2007:DEDb

[LCL07c] Terry J. Lyons, Michael Caruana, and Thierry Lévy. Differential equations driven by rough paths. Lecture Notes in Mathematics, 1908:81–93, 2007. CODEN LNMAA2. ISBN 3-540-71284-4 (print), 3-540-71285-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71285-5_5.pdf.

Lyons:2007:DEDc

[LCL07d] Terry J. Lyons, Michael Caruana, and Thierry Lévy, editors. Differential Equations Driven by Rough Paths: École d’Été de Probabilités de Saint-Flour XXXIV — 2004, volume 1908 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-71284-4 (paperback), 3-540-71285-2. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1908. URL http:

//link.springer.com/book/10.1007/978-3-540-71285-5;
http://swbplus.bsz-bw.de/bsz264672844cov.htm;   http:
//swbplus.bsz-bw.de/bsz264672844inh.htm;   http://www.
loc.gov/catdir/enhancements/fy0826/2007922938-d.html;
 http://www.loc.gov/catdir/enhancements/fy0826/2007922938-
t.html;   http://www.springerlink.com/content/978-3-540-
71285-5;   http://www.springerlink.com/content/v07h607836t8/
?p=b004e9536b4343da8bc1af8e33aaf2f7&pi=0;   http://www.
springerlink.com/openurl.asp?genre=issue&issn=0075-8434&
volume=1908.

Lyons:2007:FM

[LCL07e] Terry J. Lyons, Michael Caruana, and Thierry Lévy. Front matter. Lecture Notes in Mathematics, 1908:I–XVIII, 2007. CODEN LNMAA2. ISBN 3-540-71284-4 (print), 3-540-71285-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-71285-5/
1.pdf.

Lyons:2007:IAR

[LCL07f] Terry J. Lyons, Michael Caruana, and Thierry Lévy. Integration along rough paths. Lecture Notes in Mathematics, 1908:63–79, 2007. CODEN LNMAA2. ISBN 3-540-71284-4 (print), 3-540-71285-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71285-5_4.pdf.

Lyons:2007:RP

[LCL07g] Terry J. Lyons, Michael Caruana, and Thierry Lévy. Rough paths. Lecture Notes in Mathematics, 1908:41–61, 2007. CO-DEN LNMAA2. ISBN 3-540-71284-4 (print), 3-540-71285-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71285-5_3.pdf.

Lyons:2007:SP

[LCL07h] Terry J. Lyons, Michael Caruana, and Thierry Lévy. The signature of a path. Lecture Notes in Mathematics, 1908:25–40, 2007. CODEN LNMAA2. ISBN 3-540-71284-4 (print), 3-540-71285-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71285-5_2.pdf.

Ledoux:2001:LSI

[Led01] M. Ledoux. Logarithmic Sobolev inequalities for unbounded spin systems revisited. Lecture Notes in Mathematics, 1755:167– 194, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44671-2_13/.

Ledoux:2003:RHT

[Led03] Michel Ledoux. A remark on hypercontractivity and tail inequalities for the largest eigenvalues of random matrices. Lecture Notes in Mathematics, 1832:360–369, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40004-2_14/.

Ledoux:2007:DIL

[Led07a] M. Ledoux. Deviation inequalities on largest eigenvalues. Lecture Notes in Mathematics, 1910:167–219, 2007. CO-DEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_10.pdf.

Ledoux:2007:MHB

[Led07b] Michel Ledoux. More hypercontractive bounds for deformed orthogonal polynomial ensembles. Lecture Notes in Mathematics, 1899:235–240, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71189-6_11.pdf.

Lee:2004:MAFa

[Lee04a] Min Ho Lee. 1. Mixed automorphic forms. Lecture Notes in Mathematics, 1845:11–34, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40978-6_2/.

Lee:2004:LBE

[Lee04b] Min Ho Lee. 2. Line bundles and elliptic varieties. Lecture Notes in Mathematics, 1845:35–58, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40978-6_3/.

Lee:2004:MAFb

[Lee04c] Min Ho Lee. 3. Mixed automorphic forms and cohomology. Lecture Notes in Mathematics, 1845:59–82, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40978-6_4/.

Lee:2004:MHS

[Lee04d] Min Ho Lee. 4. Mixed Hilbert and Siegel modular forms. Lecture Notes in Mathematics, 1845:83–107, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40978-6_5/.

Lee:2004:MAFc

[Lee04e] Min Ho Lee. 5. Mixed automorphic forms on semisimple Lie groups. Lecture Notes in Mathematics, 1845:109–139, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40978-6_6/.

Lee:2004:FAV

[Lee04f] Min Ho Lee. 6. Families of Abelian varieties. Lecture Notes in Mathematics, 1845:141–175, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40978-6_7/.

Lee:2004:JF

[Lee04g] Min Ho Lee. 7. Jacobi forms. Lecture Notes in Mathematics, 1845:177–207, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40978-6_8/.

Lee:2004:TTB

[Lee04h] Min Ho Lee. 8. Twisted torus bundles. Lecture Notes in Mathematics, 1845:209–230, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40978-6_9/.

Lee:2004:I

[Lee04i] Min Ho Lee. Introduction. Lecture Notes in Mathematics, 1845: 1–9, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40978-6_1/.

Lee:2004:MAFd

[Lee04j] Min Ho Lee. Mixed Automorphic Forms, Torus Bundles, and Jacobi Forms, volume 1845 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 239 pp. LCCN QA3 .L28 no. 1845. URL

http://link.springer.com/book/10.1007/b97773;   http:
//www.springerlink.com/content/978-3-540-40978-6.

Lee:2004:R

[Lee04k] Min Ho Lee. References. Lecture Notes in Mathematics, 1845: 231–236, 2004. CODEN LNMAA2. ISBN 3-540-21922-6 (print), 3-540-40978-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40978-6_10/.

Lejay:2003:IRP

[Lej03] Antoine Lejay. An introduction to rough paths. Lecture Notes in Mathematics, 1832:1–59, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40004-2_1/.

Lejay:2009:YAI

[Lej09] Antoine Lejay. Yet another introduction to rough paths. Lecture Notes in Mathematics, 1979:1–101, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-01763-6_1.
pdf;   http://link.springer.com/content/pdf/bfm:978-3-
642-01763-6/1.pdf.

Lerner:2008:SFA

[Ler08] N. Lerner. Some facts about the wick calculus. Lecture Notes in Mathematics, 1949:135–174, 2008. CODEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-68268-4_4.pdf.

Letellier:2005:Ia

[Let05a] Emmanuel Letellier. 1. Introduction. Lecture Notes in Mathematics, 1859:1–4, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31561-2_1.pdf.

Letellier:2005:CRG

[Let05b] Emmanuel Letellier. 2. Connected reductive groups and their Lie algebras. Lecture Notes in Mathematics, 1859:5–31, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31561-2_2.pdf.

Letellier:2005:DLIa

[Let05c] Emmanuel Letellier. 3. Deligne–Lusztig induction. Lecture Notes in Mathematics, 1859:33–43, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31561-2_3.pdf.

Letellier:2005:LSP

[Let05d] Emmanuel Letellier. 4. Local systems and perverse sheaves. Lecture Notes in Mathematics, 1859:45–60, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-31561-2_4.
pdf.

Letellier:2005:GI

[Let05e] Emmanuel Letellier. 5. Geometrical induction. Lecture Notes in Mathematics, 1859:61–113, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31561-2_5.pdf.

Letellier:2005:DLIb

[Let05f] Emmanuel Letellier. 6. Deligne–Lusztig induction and Fourier transforms. Lecture Notes in Mathematics, 1859:115–149, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31561-2_6.pdf.

Letellier:2005:FTC

[Let05g] Emmanuel Letellier. 7. Fourier transforms of the characteristic functions of the adjoint orbits. Lecture Notes in Mathematics, 1859:151–158, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31561-2_7.pdf.

Letellier:2005:FTI

[Let05h] Emmanuel Letellier. Fourier Transforms of Invariant Functions on Finite Reductive Lie Algebras, volume 1859 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 165 pp. LCCN QA3 .L28

no. 1859. URL http://link.springer.com/book/10.1007/
b104209;   http://www.springerlink.com/content/978-3-
540-31561-2.

Letellier:2005:Ib

[Let05i] Emmanuel Letellier. Index. Lecture Notes in Mathematics, 1859:163–165, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31561-2_9.pdf.

Letellier:2005:R

[Let05j] Emmanuel Letellier. References. Lecture Notes in Mathematics, 1859:159–162, 2005. CODEN LNMAA2. ISBN 3-540-24020-9 (print), 3-540-31561-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31561-2_8.pdf.

Leuridan:2003:FMA

[Leu03a] Christophe Leuridan. Filtration d’une Marche aléatoire stationnaire sur le cercle. (French) []. Lecture Notes in Mathematics, 1801:335–347, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36107-7_14/.

Leuridan:2003:TRK

[Leu03b] Christophe Leuridan. Théorème de Ray–Knight dans un arbre: Une approche algébrique. (French) []. Lecture Notes in Mathematics, 1801:270–301, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36107-7_10/.

Lipman:2009:AGD

[LH09a] Joseph Lipman and Mitsuyasu Hashimoto. Abstract Grothendieck duality for schemes. Lecture Notes in Mathematics, 1960:159–252, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_5.pdf.

Lipman:2009:BMa

[LH09b] Joseph Lipman and Mitsuyasu Hashimoto. Back matter. Lecture Notes in Mathematics, 1960:253–259, 2009. CODEN LN-MAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-85420-3/
1/1.pdf.

Lipman:2009:BMb

[LH09c] Joseph Lipman and Mitsuyasu Hashimoto. Back matter. Lecture Notes in Mathematics, 1960:467–478, 2009. CODEN LN-MAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-85420-3/
2/1.pdf.

Lipman:2009:BNR

[LH09d] Joseph Lipman and Mitsuyasu Hashimoto. Bökstedt–Neeman resolutions and HyperExt sheaves. Lecture Notes in Mathematics, 1960:373–376, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85420-3_19.pdf.

Lipman:2009:CDC

[LH09e] Joseph Lipman and Mitsuyasu Hashimoto. Commutativity of diagrams constructed from a monoidal pair of pseudofunctors. Lecture Notes in Mathematics, 1960:263–278, 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_7.pdf.

Lipman:2009:DTC

[LH09f] Joseph Lipman and Mitsuyasu Hashimoto. Derived and triangulated categories. Lecture Notes in Mathematics, 1960:11–42, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_2.pdf.

Lipman:2009:DCD

[LH09g] Joseph Lipman and Mitsuyasu Hashimoto. Derived categories and derived functors of sheaves on ringed sites. Lecture Notes in Mathematics, 1960:303–312, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-85420-3_9.pdf.

Lipman:2009:DDI

[LH09h] Joseph Lipman and Mitsuyasu Hashimoto. Derived direct and inverse image. Lecture Notes in Mathematics, 1960:83–158, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_4.pdf.

Lipman:2009:DF

[LH09i] Joseph Lipman and Mitsuyasu Hashimoto. Derived functors. Lecture Notes in Mathematics, 1960:43–81, 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_3.pdf.

Lipman:2009:DFF

[LH09j] Joseph Lipman and Mitsuyasu Hashimoto. Derived functors of functors on sheaves of modules over diagrams of schemes. Lecture Notes in Mathematics, 1960:343–349, 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_14.pdf.

Lipman:2009:DT

[LH09k] Joseph Lipman and Mitsuyasu Hashimoto. Descent theory. Lecture Notes in Mathematics, 1960:355–361, 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_16.pdf.

Lipman:2009:FGD

[LH09l] Joseph Lipman and Mitsuyasu Hashimoto. Foundations of Grothendieck Duality for Diagrams of Schemes, volume 1960 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 478 pp. LCCN QA3 .L28 no. 1960. URL http://link.

springer.com/book/10.1007/978-3-540-85420-3;   http:/
/www.springerlink.com/content/978-3-540-85420-3.

Lipman:2009:FMa

[LH09m] Joseph Lipman and Mitsuyasu Hashimoto. Front matter. Lecture Notes in Mathematics, 1960:i–x, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-85420-3/1.pdf.

Lipman:2009:FMb

[LH09n] Joseph Lipman and Mitsuyasu Hashimoto. Front matter. Lecture Notes in Mathematics, 1960:1–3, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-85420-3/1/1.pdf.

Lipman:2009:FMc

[LH09o] Joseph Lipman and Mitsuyasu Hashimoto. Front matter. Lecture Notes in Mathematics, 1960:253–257, 2009. CODEN LN-MAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-85420-3/
2/1.pdf.

Lipman:2009:GS

[LH09p] Joseph Lipman and Mitsuyasu Hashimoto. Groupoid of schemes. Lecture Notes in Mathematics, 1960:367–371, 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_18.pdf.

Lipman:2009:Ia

[LH09q] Joseph Lipman and Mitsuyasu Hashimoto. Introduction. Lecture Notes in Mathematics, 1960:5–10, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-85420-3_1.
pdf.

Lipman:2009:Ib

[LH09r] Joseph Lipman and Mitsuyasu Hashimoto. Introduction. Lecture Notes in Mathematics, 1960:259–262, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-85420-3_6.
pdf.

Lipman:2009:LRI

[LH09s] Joseph Lipman and Mitsuyasu Hashimoto. The left and right inductions and the direct and inverse images. Lecture Notes in Mathematics, 1960:319–321, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-85420-3_11.pdf.

Lipman:2009:LNP

[LH09t] Joseph Lipman and Mitsuyasu Hashimoto. Local Noetherian property. Lecture Notes in Mathematics, 1960:363–365, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_17.pdf.

Lipman:2009:OSS

[LH09u] Joseph Lipman and Mitsuyasu Hashimoto. Operations on sheaves via the structure data. Lecture Notes in Mathematics, 1960:323– 336, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_12.pdf.

Lipman:2009:QCS

[LH09v] Joseph Lipman and Mitsuyasu Hashimoto. Quasi-coherent sheaves over a diagram of schemes. Lecture Notes in Mathematics, 1960:337–342, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85420-3_13.pdf.

Lipman:2009:RAD

[LH09w] Joseph Lipman and Mitsuyasu Hashimoto. The right adjoint of the derived direct image functor. Lecture Notes in Mathematics, 1960:377–383, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85420-3_20.pdf.

Lipman:2009:SRS

[LH09x] Joseph Lipman and Mitsuyasu Hashimoto. Sheaves on ringed sites. Lecture Notes in Mathematics, 1960:279–302, 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_8.pdf.

Lipman:2009:SDS

[LH09y] Joseph Lipman and Mitsuyasu Hashimoto. Sheaves over a diagram of S-schemes. Lecture Notes in Mathematics, 1960:313–317, 2009. CODEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_10.pdf.

Lipman:2009:SO

[LH09z] Joseph Lipman and Mitsuyasu Hashimoto. Simplicial objects. Lecture Notes in Mathematics, 1960:351–353, 2009. CO-DEN LNMAA2. ISBN 3-540-85419-3 (print), 3-540-85420-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85420-3_15.pdf.

Li:2002:CBS

[Li02a] Huishi Li. Chapter I: Basic structural tricks and examples. Lecture Notes in Mathematics, 1795:5–32, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45765-7_2/.

Li:2002:CIGa

[Li02b] Huishi Li. Chapter II: Gröbner bases in associative algebras. Lecture Notes in Mathematics, 1795:33–65, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45765-7_3/.

Li:2002:CIGb

[Li02c] Huishi Li. Chapter III: Gröbner bases and basic algebraic– algorithmic structures. Lecture Notes in Mathematics, 1795:67– 90, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45765-7_4/.

Li:2002:CIF

[Li02d] Huishi Li. Chapter IV: Filtered-graded transfer of Gröbner bases. Lecture Notes in Mathematics, 1795:91–105, 2002. CODEN LN-MAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45765-7_
5/.

Li:2002:CVG

[Li02e] Huishi Li. Chapter V: GK-dimension of modules over quadric solvable polynomial algebras and elimination of variables. Lecture Notes in Mathematics, 1795:107–132, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45765-7_6/.

Li:2002:CVM

[Li02f] Huishi Li. Chapter VI: Multiplicity computation of modules over quadric solvable polynomial algebras. Lecture Notes in Mathematics, 1795:133–151, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45765-7_7/.

Li:2002:CVH

[Li02g] Huishi Li. Chapter VII: (∂)-Holonomic modules and functions over quadric solvable polynomial algebras. Lecture Notes in Mathematics, 1795:153–173, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45765-7_8/.

Li:2002:CVR

[Li02h] Huishi Li. Chapter VIII: Regularity and K_{0}-group of quadric solvable polynomial algebras. Lecture Notes in Mathematics, 1795: 175–186, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45765-7_9/.

Li:2002:Ib

[Li02i] Huishi Li. Index. Lecture Notes in Mathematics, 1795:195– 197, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45765-7_11/.

Li:2002:Ia

[Li02j] Huishi Li. Introduction. Lecture Notes in Mathematics, 1795: 1–4, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45765-7_1/.

Li:2002:NGB

[Li02k] Huishi Li. Noncommutative Gröbner Bases and Filtered-Graded Transfer, volume 1795 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 197 pp. LCCN QA3 .L28 no. 1795. URL http:/

/link.springer.com/book/10.1007/b84211;   http://www.
springerlink.com/content/978-3-540-45765-7.

Li:2002:R

[Li02l] Huishi Li. References. Lecture Notes in Mathematics, 1795:187– 193, 2002. CODEN LNMAA2. ISBN 3-540-44196-4 (print), 3-540-45765-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45765-7_10/.

Lindsay:2005:QSA

[Lin05] J. Martin Lindsay. Quantum stochastic analysis – an introduction. Lecture Notes in Mathematics, 1865:181–271, 2005. CODEN LNMAA2. ISBN 3-540-24406-9 (print), 3-540-31450-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11376569_3.pdf.

Lions:2000:CIP

[Lio00] J.-L. Lions. Complexity in industrial problems. Some remarks. Lecture Notes in Mathematics, 1739:223–266, 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103921/.

Lions:2007:LIT

[LL07] Pierre-Louis Lions and Jean-Michel Lasry. Large investor trading impacts on volatility. Lecture Notes in Mathematics, 1919:173– 190, 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73327-0_4.pdf.

Lenglart:2002:PUC

[LLP02] E. Lenglart, D. Lepingle, and M. Pratelli. Présentation unifiée de certaines inégalités de la théorie des martingales. (French) []. Lecture Notes in Mathematics, 1771:361–383, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_14/.

Lamoureux:2008:INM

[LM08] M. P. Lamoureux and G. F. Margrave. An introduction to numerical methods of pseudodifferential operators. Lecture Notes in Mathematics, 1949:79–133, 2008. CODEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-68268-4_3.pdf.

Litvak:2007:EME

[LMPTJ07] A. E. Litvak, Vitali D. Milman, A. Pajor, and N. Tomczak-Jaegermann. On the Euclidean metric entropy of convex bodies. Lecture Notes in Mathematics, 1910:221–235, 2007. CO-DEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_11.pdf.

Litvak:2004:IRS

[LMTJ04] A. E. Litvak, V. D. Milman, and N. Tomczak-Jaegermann. Isomorphic random subspaces and quotients of convex and quasiconvex bodies. Lecture Notes in Mathematics, 1850:159–178, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_15/.

Levine:2006:ABM

[LNH06] Howard A. Levine and Marit Nilsen-Hamilton. Angiogenesis — a biochemical/mathematical perspective. Lecture Notes in Mathematics, 1872:23–76, 2006. CODEN LNMAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/11561606_2.pdf.

Latala:2000:BSP

[LO00] R. Latała and K. Oleszkiewicz. Between Sobolev and Poincaré. Lecture Notes in Mathematics, 1745:147–168, 2000. CODEN LN-MAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107213/.

Loday:2001:D

[Lod01] Jean-Louis Loday. Dialgebras. Lecture Notes in Mathematics, 1763:7–66, 2001. CODEN LNMAA2. ISBN 3-540-42194-7 (print), 3-540-45328-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45328-8_2/.

Lolas:2006:MMP

[Lol06] Georgios Lolas. Mathematical modelling of proteolysis and cancer cell invasion of tissue. Lecture Notes in Mathematics, 1872: 77–129, 2006. CODEN LNMAA2. ISBN 3-540-29162-8 (print), 3-540-32415-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11561606_3.pdf.

Lombardi:2000:Ra

[Lom00a] Eric Lombardi. The 0^{2+}iω resonance. Lecture Notes in Mathematics, 1741:186–325, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0104109/.

Lombardi:2000:RID

[Lom00b] Eric Lombardi. The 0^{2+}iω resonance in infinite dimensions. Application to water waves. Lecture Notes in Mathematics, 1741: 327–357, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104110/.

Lombardi:2000:Rb

[Lom00c] Eric Lombardi. The (iω_{0})^{2}iω_{1} resonance. Lecture Notes in Mathematics, 1741:359–403, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104111/.

Lombardi:2000:ADP

[Lom00d] Eric Lombardi. Analytic description of periodic orbits bifurcating from a pair of simple purely imaginary eigenvalues. Lecture Notes in Mathematics, 1741:101–122, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104106/.

Lombardi:2000:CFT

[Lom00e] Eric Lombardi. Constructive Floquet theory for periodic matrices near a constant one. Lecture Notes in Mathematics, 1741: 123–134, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104107/.

Lombardi:2000:ETE

[Lom00f] Eric Lombardi. ‘exponential tools’ for evaluating oscillatory integrals. Lecture Notes in Mathematics, 1741:22–76, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104104/.

Lombardi:2000:I

[Lom00g] Eric Lombardi. Introduction. Lecture Notes in Mathematics, 1741:1–19, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104103/.

Lombardi:2000:IAE

[Lom00h] Eric Lombardi. Inversion of affine equations around reversible homoclinic connections. Lecture Notes in Mathematics, 1741:135– 184, 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104108/.

Lombardi:2000:OIP

[Lom00i] Eric Lombardi. Oscillatory Integrals and Phenomena Beyond all Algebriac Orders: With Applications to Homoclinic Orbits in Reversible Systems, volume 1741 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 413 pp. LCCN QA3 .L28 no. 1741.

URL http://link.springer.com/book/10.1007/BFb0104102;
 http://www.springerlink.com/content/978-3-540-44971-

Lombardi:2000:RRV

[Lom00j] Eric Lombardi. Resonances of reversible vector fields. Lecture Notes in Mathematics, 1741:78–100, 2000. CODEN LN-MAA2. ISBN 3-540-67785-2 (print), 3-540-44971-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104105/.

Lou:2008:BM

[Lou08a] Y. Lou. Back matter. Lecture Notes in Mathematics, 1922:206– 209, 2008. CODEN LNMAA2. ISBN 3-540-74328-6 (print), 3-540-74331-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-74331-6/1.pdf.

Lou:2008:SCM

[Lou08b] Y. Lou. Some challenging mathematical problems in evolution of dispersal and population dynamics. Lecture Notes in Mathematics, 1922:171–205, 2008. CODEN LNMAA2. ISBN 3-540-74328-6 (print), 3-540-74331-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74331-6_5.pdf.

Litvak:2000:RAH

[LTJ00] A. E. Litvak and N. Tomczak-Jaegermann. Random aspects of high-dimensional convex bodies. Lecture Notes in Mathematics, 1745:169–190, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0107214/.

Lunardi:2004:IPM

[Lun04] Alessandra Lunardi. An introduction to parabolic moving boundary problems. Lecture Notes in Mathematics, 1855:371–399, 2004. CODEN LNMAA2. ISBN 3-540-23030-0 (print), 3-540-44653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44653-8_4/.

Lusztig:2002:NAH

[Lus02] George Lusztig. Notes on affine Hecke algebras. Lecture Notes in Mathematics, 1804:71–103, 2002. CODEN LNMAA2. ISBN 3-540-00224-3 (print), 3-540-36205-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36205-0_3/.

Laevsky:2000:FHT

[LvdLM00] K. Laevsky, B. J. van der Linden, and R. M. M. Mattheij. Flow and heat transfer in pressing of glass products. Lecture Notes in Mathematics, 1739:267–285, 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103922/.

Lyubich:2004:AES

[Lyu04] Yu I. Lyubich. Almost Euclidean subspaces of real ℓ_{p}^{n} with p an even integer. Lecture Notes in Mathematics, 1850:179– 192, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_16/.

Malric:2001:FQF

[Mal01] M. Malric. Filtrations quotients de la filtration brownienne. (French) []. Lecture Notes in Mathematics, 1755:260–264, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44671-2_19/.

Maleva:2003:PPU

[Mal03] Olga Maleva. Point preimages under ball non-collapsing mappings. Lecture Notes in Mathematics, 1807:148–157, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-36428-3_13/.

Malchiodi:2009:CSS

[Mal09] Andrea Malchiodi. Concentration of solutions for some singularly perturbed Neumann problems. Lecture Notes in Mathematics, 1977:63–115, 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01674-5_3.pdf.

Manetti:2008:SSD

[Man08] Marco Manetti. Smoothings of singularities and deformation types of surfaces. Lecture Notes in Mathematics, 1938:169–230, 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78279-7_3.pdf.

Maric:2000:ABR

[Mar00a] Vojislav Marić. Asymptotic behaviour of regular solutions. Lecture Notes in Mathematics, 1726:49–70, 2000. CODEN LN-MAA2. ISBN 3-540-67160-9 (print), 3-540-46520-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103955/.

Maric:2000:EAB

[Mar00b] Vojislav Marić. An equation arising in boundary-layer theory. Lecture Notes in Mathematics, 1726:105–114, 2000. CODEN LNMAA2. ISBN 3-540-67160-9 (print), 3-540-46520-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103957/.

Maric:2000:ETF

[Mar00c] Vojislav Marić. Equations of Thomas–Fermi type. Lecture Notes in Mathematics, 1726:71–104, 2000. CODEN LNMAA2. ISBN 3-540-67160-9 (print), 3-540-46520-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103956/.

Maric:2000:ERS

[Mar00d] Vojislav Marić. Existence of regular solutions. Lecture Notes in Mathematics, 1726:9–47, 2000. CODEN LNMAA2. ISBN 3-540-67160-9 (print), 3-540-46520-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103954/.

Maric:2000:I

[Mar00e] Vojislav Marić. Introduction. Lecture Notes in Mathematics, 1726:1–8, 2000. CODEN LNMAA2. ISBN 3-540-67160-9 (print), 3-540-46520-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103953/.

Maric:2000:RVD

[Mar00f] Vojislav Marić. Regular Variation and Differential Equations, volume 1726 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67160-9 (print), 3-540-46520-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 127 pp. LCCN QA3 .L28 no. 1726. URL http:

//link.springer.com/book/10.1007/BFb0103952;   http:/
/www.springerlink.com/content/978-3-540-46520-1.

Marcati:2007:HSB

[Mar07] Pierangelo Marcati, editor. Hyperbolic Systems of Balance Laws: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, July 14–21, 2003, volume 1911 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-72186-X (print), 3-540-72187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1911. URL http:

//link.springer.com/book/10.1007/978-3-540-72187-1;
http://www.springerlink.com/content/978-3-540-72187-1.

Mariano:2008:BM

[Mar08a] Paolo Maria Mariano. Back matter. Lecture Notes in Mathematics, 1937:211–212, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-78277-3/1.pdf.

Mariano:2008:SMG

[Mar08b] Paolo Maria Mariano. Slow motion in granular matter. Lecture Notes in Mathematics, 1937:193–210, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78277-3_9.
pdf.

Marino:2008:LTV

[Mar08c] M. Mariño. Lectures on the topological vertex. Lecture Notes in Mathematics, 1947:49–104, 2008. CODEN LNMAA2. ISBN 3-540-79813-7 (print), 3-540-79814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-79814-9_2.pdf.

Markowsky:2008:PTL

[Mar08d] Greg Markowsky. Proof of a Tanaka-like formula stated by J. Rosen in Séminaire XXXVIII. Lecture Notes in Mathematics, 1934:199–202, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77913-1_9.pdf.

Masser:2003:HTL

[Mas03] David Masser. Heights, transcendence, and linear independence on commutative group varieties. Lecture Notes in Mathematics, 1819:1–51, 2003. CODEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44979-5_1/.

Massart:2007:BM

[Mas07a] Professor Pascal Massart. Back matter. Lecture Notes in Mathematics, 1896:319–341, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-48503-2/1.pdf.

Massart:2007:CI

[Mas07b] Professor Pascal Massart. Concentration inequalities. Lecture Notes in Mathematics, 1896:147–181, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-48503-2_5.
pdf.

Massart:2007:DEM

[Mas07c] Professor Pascal Massart. Density estimation via model selection. Lecture Notes in Mathematics, 1896:201–277, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48503-2_7.pdf.

Massart:2007:EII

[Mas07d] Professor Pascal Massart. Exponential and information inequalities. Lecture Notes in Mathematics, 1896:15–51, 2007.

CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48503-2_2.pdf.

Massart:2007:FM

[Mas07e] Professor Pascal Massart. Front matter. Lecture Notes in Mathematics, 1896:I–XIV, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-48503-2/1.pdf.

Massart:2007:GMS

[Mas07f] Professor Pascal Massart. Gaussian model selection. Lecture Notes in Mathematics, 1896:83–146, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-48503-2_4.
pdf.

Massart:2007:GP

[Mas07g] Professor Pascal Massart. Gaussian processes. Lecture Notes in Mathematics, 1896:53–82, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48503-2_3.pdf.

Massart:2007:I

[Mas07h] Professor Pascal Massart. Introduction. Lecture Notes in Mathematics, 1896:1–13, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-48503-2_1.pdf.

Massart:2007:MI

[Mas07i] Professor Pascal Massart. Maximal inequalities. Lecture Notes in Mathematics, 1896:183–199, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48503-2_6.pdf.

Massart:2007:SL

[Mas07j] Professor Pascal Massart. Statistical learning. Lecture Notes in Mathematics, 1896:279–318, 2007. CODEN LNMAA2. ISBN 3-540-48497-3 (print), 3-540-48503-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48503-2_8.pdf.

Mascolo:2008:BM

[Mas08a] Elvira Mascolo. Back matter. Lecture Notes in Mathematics, 1927:193–204, 2008. CODEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-75914-0/1.pdf.

Mascolo:2008:CCP

[Mas08b] Elvira Mascolo. CIME courses on partial differential equations and calculus of variations. Lecture Notes in Mathematics, 1927:183–189, 2008. CODEN LNMAA2. ISBN 3-540-75913-1 (print), 3-540-75914-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75914-0_6.pdf.

Meckes:2007:SRT

[Mec07] M. Meckes. Some remarks on transportation cost and related inequalities. Lecture Notes in Mathematics, 1910:237–244, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_12.pdf.

Meeks:2002:MSF

[Mee02] William H. Meeks III. Minimal surfaces in flat three-dimensional spaces. Lecture Notes in Mathematics, 1775:1–14, 2002. CODEN LNMAA2. ISBN 3-540-43120-9 (print), 3-540-45609-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45609-4_
1/.

Meilijson:2003:TGD

[Mei03] Isaac Meilijson. The time to a given drawdown in Brownian motion. Lecture Notes in Mathematics, 1832:94–108, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_5/.

Melenk:2002:Ia

[Mel02a] Jens M. Melenk. 1. Introduction. Lecture Notes in Mathematics, 1796:1–20, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45781-7_1/.

Melenk:2002:HFR

[Mel02b] Jens M. Melenk. 2. hp-FEM for reaction diffusion problems: Principal results. Lecture Notes in Mathematics, 1796:23–72, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45781-7_2/.

Melenk:2002:HA

[Mel02c] Jens M. Melenk. 3. hp approximation. Lecture Notes in Mathematics, 1796:73–138, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45781-7_3/.

Melenk:2002:CNS

[Mel02d] Jens M. Melenk. 4. The countably normed spaces B_{β,ε}^{l}. Lecture Notes in Mathematics, 1796:141–168, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45781-7_4/.

Melenk:2002:RTC

[Mel02e] Jens M. Melenk. 5. Regularity theory in countably normed spaces. Lecture Notes in Mathematics, 1796:169–224, 2002. CODEN LN-MAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45781-7_

Melenk:2002:EWC

[Mel02f] Jens M. Melenk. 6. Exponentially weighted countably normed spaces. Lecture Notes in Mathematics, 1796:227–254, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-45781-7_6/.

Melenk:2002:RTA

[Mel02g] Jens M. Melenk. 7. Regularity through asymptotic expansions. Lecture Notes in Mathematics, 1796:255–295, 2002. CODEN LN-MAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45781-7_
7/.

Melenk:2002:A

[Mel02h] Jens M. Melenk. Appendix. Lecture Notes in Mathematics, 1796: 297–310, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45781-7_8/.

Melenk:2002:HFE

[Mel02i] Jens M. Melenk. hp-Finite Element Methods for Singular Perturbations, volume 1796 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 318 pp. LCCN QA3 .L28 no. 1796. URL http:/

/link.springer.com/book/10.1007/b84212;   http://www.
springerlink.com/content/978-3-540-45781-7.

Melenk:2002:Ib

[Mel02j] Jens M. Melenk. Index. Lecture Notes in Mathematics, 1796: 317–318, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45781-7_10.

Melenk:2002:R

[Mel02k] Jens M. Melenk. References. Lecture Notes in Mathematics, 1796: 311–316, 2002. CODEN LNMAA2. ISBN 3-540-44201-4 (print), 3-540-45781-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45781-7_9/.

Mendelson:2004:GPL

[Men04] S. Mendelson. Geometric parameters in learning theory. Lecture Notes in Mathematics, 1850:193–235, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44489-3_17/.

Merzbach:2003:IGT

[Mer03] Ely Merzbach. An introduction to the general theory of setindexed martingales. Lecture Notes in Mathematics, 1802:41– 84, 2003. CODEN LNMAA2. ISBN 3-540-00295-2 (print), 3-540-36259-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36259-3_2/.

Merkurjev:2004:AMR

[Mer04] Alexander S. Merkurjev. Appendix: My recollections about Oleg Izhboldin. Lecture Notes in Mathematics, 1835:185–187, 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40990-8_7/.

Merkli:2006:IQG

[Mer06] Marco Merkli. The ideal quantum gas. Lecture Notes in Mathematics, 1880:183–233, 2006. CODEN LNMAA2. ISBN 3-540-30991-8 (print), 3-540-33922-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-33922-1_5.pdf.

Meyer:2002:GDT

[Mey02a] P.-A. Meyer. Guide détaillé de la théorie ^{≪}générale^{≫} des processus. (French) []. Lecture Notes in Mathematics, 1771:49– 74, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_4/.

Meyer:2002:ISF

[Mey02b] P.-A. Meyer. Intégrales stochastiques I. (French) []. Lecture Notes in Mathematics, 1771:97–119, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_8/.

Meyer:2002:ISI

[Mey02c] P.-A. Meyer. Intégrales stochastiques II. (French) []. Lecture Notes in Mathematics, 1771:120–142, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_9/.

Meyer:2002:DFF

[Mey02d] P.-A. Meyer. Le dual de ^{≪}H^{1}^{≫} est ^{≪}BMO^{≫} (cas continu). (French) []. Lecture Notes in Mathematics, 1771:384–393, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_15/.

Meyer:2002:CIS

[Mey02e] P.-A. Meyer. Un cours sur les intégrales stochastiques. (French) []. Lecture Notes in Mathematics, 1771:174–329, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45530-1_
11/.

Meyer:2006:OPS

[Mey06] Yves Meyer. Oscillating patterns in some nonlinear evolution equations. Lecture Notes in Mathematics, 1871:101–187, 2006. CODEN LNMAA2. ISBN 3-540-28586-5 (print), 3-540-32454-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11545989_4.pdf.

Miclo:2003:APL

[Mic03a] Laurent Miclo. About projections of logarithmic Sobolev inequalities. Lecture Notes in Mathematics, 1801:201–221, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_6/.

Miclo:2003:ISL

[Mic03b] Laurent Miclo. Sur l’inégalité de Sobolev logarithmique des opérateurs de Laguerre à petit paramètre. (French) []. Lecture Notes in Mathematics, 1801:222–229, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36107-7_7/.

Miekisz:2008:BM

[Mie08a] Jacek Miekisz. Back matter. Lecture Notes in Mathematics, 1940: 317–321, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-78362-6/1.pdf.

Miekisz:2008:E

[Mie08b] Jacek Miekisz. Erratum. Lecture Notes in Mathematics, 1940: 323, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78362-6_6.pdf.

Miekisz:2008:EGT

[Mie08c] Jacek Miekisz. Evolutionary game theory and population dynamics. Lecture Notes in Mathematics, 1940:269–316, 2008. CODEN LNMAA2. ISBN 3-540-78360-1 (print), 3-540-78362-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78362-6_5.pdf.

Mikelic:2000:HTA

[Mik00] Andro Mikelić. Homogenization theory and applications to filtration through porous media. Lecture Notes in Mathematics, 1734: 127–214, 2000. CODEN LNMAA2. ISBN 3-540-67868-9 (print), 3-540-44656-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103977/.

Milman:2007:CLD

[Mil07] E. Milman. A comment on the low-dimensional Busemann– Petty problem. Lecture Notes in Mathematics, 1910:245–253, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_13.pdf.

Mimura:2003:RDS

[Mim03] Masayasu Mimura. Reaction-diffusion systems arising in biological and chemical systems: Application of singular limit procedures. Lecture Notes in Mathematics, 1812:89–121, 2003. CODEN LNMAA2. ISBN 3-540-14033-6 (print), 3-540-39189-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39189-0_
3/.

Mishura:2008:BM

[Mis08a] Yuliya S. Mishura. Back matter. Lecture Notes in Mathematics, 1929:363–393, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-75873-0/1.pdf.

Mishura:2008:FSF

[Mis08b] Yuliya S. Mishura. Filtering in systems with fractional Brownian noise. Lecture Notes in Mathematics, 1929:291–299, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75873-0_4.pdf.

Mishura:2008:FAF

[Mis08c] Yuliya S. Mishura. Financial applications of fractional Brownian motion. Lecture Notes in Mathematics, 1929:301–326, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75873-0_5.pdf.

Mishura:2008:FM

[Mis08d] Yuliya S. Mishura. Front matter. Lecture Notes in Mathematics, 1929:I–XVII, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-75873-0/1.pdf.

Mishura:2008:SIF

[Mis08e] Yuliya S. Mishura. Statistical inference with fractional Brownian motion. Lecture Notes in Mathematics, 1929:327–362, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75873-0_6.pdf.

Mishura:2008:SCF

[Mis08f] Yuliya S. Mishura. Stochastic Calculus for Fractional Brownian Motion and Related Processes, volume 1929 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 393 pp. LCCN

QA3 .L28 no. 1929. URL http://link.springer.com/book/
10.1007/978-3-540-75873-0;   http://www.springerlink.
com/content/978-3-540-75873-0.

Mishura:2008:SDE

[Mis08g] Yuliya S. Mishura. Stochastic differential equations involving fractional Brownian motion. Lecture Notes in Mathematics, 1929:197–290, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75873-0_3.pdf.

Mishura:2008:SIR

[Mis08h] Yuliya S. Mishura. Stochastic integration with respect to fBm and related topics. Lecture Notes in Mathematics, 1929:123–196, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75873-0_2.pdf.

Mishura:2008:WIR

[Mis08i] Yuliya S. Mishura. Wiener integration with respect to fractional Brownian motion. Lecture Notes in Mathematics, 1929:1–121, 2008. CODEN LNMAA2. ISBN 3-540-75872-0 (print), 3-540-75873-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75873-0_1.pdf.

Mazya:2007:BM

[MK07a] Vladimir Maz’ya and Gershon Kresin. Back matter. Lecture Notes in Mathematics, 1903:129–144, 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-69574-5/1.pdf.

Mazya:2007:BTR

[MK07b] Vladimir Maz’ya and Gershon Kresin. Bohr’s type real part estimates. Lecture Notes in Mathematics, 1903:99–114, 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69574-5_6.pdf.

Mazya:2007:EAFa

[MK07c] Vladimir Maz’ya and Gershon Kresin. Estimates for analytic functions bounded with respect to their real part. Lecture Notes in Mathematics, 1903:1–16, 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69574-5_1.pdf.

Mazya:2007:EAFc

[MK07d] Vladimir Maz’ya and Gershon Kresin. Estimates for analytic functions by the best L_{p}-approximation of Rf on the circle. Lecture Notes in Mathematics, 1903:37–55, 2007. CO-DEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69574-5_3.pdf.

Mazya:2007:EAFb

[MK07e] Vladimir Maz’ya and Gershon Kresin. Estimates for analytic functions with respect to the L_{p}-norm of RΔf on the circle. Lecture Notes in Mathematics, 1903:17–35, 2007. CO-DEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69574-5_2.pdf.

Mazya:2007:EDA

[MK07f] Vladimir Maz’ya and Gershon Kresin. Estimates for derivatives of analytic functions. Lecture Notes in Mathematics, 1903:69–98, 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69574-5_5.pdf.

Mazya:2007:EDD

[MK07g] Vladimir Maz’ya and Gershon Kresin. Estimates for directional derivatives of harmonic functions. Lecture Notes in Mathematics, 1903:57–67, 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69574-5_4.pdf.

Mazya:2007:EID

[MK07h] Vladimir Maz’ya and Gershon Kresin. Estimates for the increment of derivatives of analytic functions. Lecture Notes in Mathematics, 1903:115–128, 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69574-5_7.pdf.

Mazya:2007:FM

[MK07i] Vladimir Maz’ya and Gershon Kresin. Front matter. Lecture Notes in Mathematics, 1903:I–XV, 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-69574-5/1.pdf.

Mazya:2007:SRP

[MK07j] Vladimir Maz’ya and Gershon Kresin. Sharp Real-Part Theorems: A Unified Approach, volume 1903 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-69573-7 (print), 3-540-69574-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 140 pp. LCCN QA3 .L28 no.

1903. URL http://link.springer.com/book/10.1007/3-540-

69573-7;   http://www.springerlink.com/content/978-3-
540-69574-5.

Marsden:2007:BM

[MMO^{+}07a] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Back matter. Lecture Notes in Mathematics, 1913:483–523, 2007. CODEN LN-MAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-72470-4/
1.pdf.

Marsden:2007:CRS

[MMO^{+}07b] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Commuting reduction and semidirect product theory. Lecture Notes in Mathematics, 1913:113–142, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72470-4_4.pdf.

Marsden:2007:CBR

[MMO^{+}07c] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Cotangent bundle reduction. Lecture Notes in Mathematics, 1913:43–99, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_2.pdf.

Marsden:2007:E

[MMO^{+}07d] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Examples. Lecture Notes in Mathematics, 1913:251–283, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-72470-4_9.
pdf.

Marsden:2007:FMa

[MMO^{+}07e] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Front matter. Lecture Notes in Mathematics, 1913:I–XV, 2007. CODEN LN-MAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-72470-4/
1.pdf.

Marsden:2007:FMb

[MMO^{+}07f] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Front matter. Lecture Notes in Mathematics, 1913:1, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-72470-4/1/1.pdf.

Marsden:2007:FMc

[MMO^{+}07g] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Front matter. Lecture Notes in Mathematics, 1913:111, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-72470-4/2/1.pdf.

Marsden:2007:FMd

[MMO^{+}07h] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Front matter. Lecture Notes in Mathematics, 1913:421–422, 2007. CODEN LN-MAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-72470-4/
3/1.pdf.

Marsden:2007:GES

[MMO^{+}07i] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Group extensions and the stages hypothesis. Lecture Notes in Mathematics, 1913:177– 210, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_6.pdf.

Marsden:2007:HRS

[MMO^{+}07j] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Hamiltonian Reduction by Stages, volume 1913 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 519 pp. LCCN QA3 .L28 no. 1913.

URL http://link.springer.com/book/10.1007/978-3-540-
72470-4;   http://www.springerlink.com/content/978-3-
540-72470-4.

Marsden:2007:MCB

[MMO^{+}07k] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Magnetic cotangent bundle reduction. Lecture Notes in Mathematics, 1913:211–237, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_7.pdf.

Marsden:2007:OMM

[MMO^{+}07l] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. The optimal momentum map and point reduction. Lecture Notes in Mathematics, 1913:423–436, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72470-4_13.pdf.

Marsden:2007:OOR

[MMO^{+}07m] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Optimal orbit reduction. Lecture Notes in Mathematics, 1913:437–459, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_14.pdf.

Marsden:2007:ORS

[MMO^{+}07n] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Optimal reduction by stages. Lecture Notes in Mathematics, 1913:461–481, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_15.pdf.

Marsden:2007:PS

[MMO^{+}07o] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. The problem setting. Lecture Notes in Mathematics, 1913:101–109, 2007. CO-DEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_3.pdf.

Marsden:2007:RSS

[MMO^{+}07p] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Reduction by stages via symplectic distributions. Lecture Notes in Mathematics, 1913: 397–407, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-72470-4_11.pdf.

Marsden:2007:RST

[MMO^{+}07q] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Reduction by stages with topological conditions. Lecture Notes in Mathematics, 1913: 409–420, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-72470-4_12.pdf.

Marsden:2007:RRS

[MMO^{+}07r] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Regular reduction by stages. Lecture Notes in Mathematics, 1913:143–175, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_5.pdf.

Marsden:2007:SCO

[MMO^{+}07s] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Stages and coadjoint orbits of central extensions. Lecture Notes in Mathematics, 1913:239–250, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72470-4_8.pdf.

Marsden:2007:SSP

[MMO^{+}07t] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Stages and semidirect products with cocycles. Lecture Notes in Mathematics, 1913:285– 396, 2007. CODEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_10.pdf.

Marsden:2007:SR

[MMO^{+}07u] Jerrold E. Marsden, Gerard Misiolek, Juan-Pablo Ortega, Matthew Perlmutter, and Tudor S. Raţiu. Symplectic reduction. Lecture Notes in Mathematics, 1913:3–42, 2007. CO-DEN LNMAA2. ISBN 3-540-72469-9 (print), 3-540-72470-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72470-4_1.pdf.

Masser:2003:FM

[MNS^{+}03] David Masser, Yuri V. Nesterenko, Hans Peter Schlickewei, Wolfgang Schmidt, and Michel Waldschmidt. Front matter. Lecture Notes in Mathematics, 1819:I–XI, 2003. CODEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-44979-9/1.pdf.

Maes:2009:SNI

[MNS09] Christian Maes, Karel Netočný, and Bidzina Shergelashvili. A selection of nonequilibrium issues. Lecture Notes in Mathematics, 1970:1–60, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-92796-9_6.pdf.

Mochizuki:2009:BM

[Moc09a] Takuro Mochizuki. Back matter. Lecture Notes in Mathematics, 1972:1–48, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-93913-9/1.pdf.

Mochizuki:2009:DTI

[Moc09b] Takuro Mochizuki. Donaldson Type Invariants for Algebraic Surfaces: Transition of Moduli Stacks, volume 1972 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxiv + 383 pp. LCCN

QA3 .L28 no. 1972. URL http://link.springer.com/book/
10.1007/978-3-540-93913-9;   http://www.springerlink.
com/content/978-3-540-93913-9.

Mochizuki:2009:FM

[Moc09c] Takuro Mochizuki. Front matter. Lecture Notes in Mathematics, 1972:1–20, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-93913-9/1.pdf.

Mochizuki:2009:GIT

[Moc09d] Takuro Mochizuki. Geometric invariant theory and enhanced master space. Lecture Notes in Mathematics, 1972:1–47, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-93913-9_4.pdf.

Mochizuki:2009:Ia

[Moc09e] Takuro Mochizuki. Introduction. Lecture Notes in Mathematics, 1972:1–23, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-93913-9_1.pdf.

Mochizuki:2009:Ib

[Moc09f] Takuro Mochizuki. Invariants. Lecture Notes in Mathematics, 1972:1–77, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-93913-9_7.pdf.

Mochizuki:2009:OTM

[Moc09g] Takuro Mochizuki. Obstruction theories of moduli stacks and master spaces. Lecture Notes in Mathematics, 1972:1–67, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-93913-9_5.pdf.

Mochizuki:2009:PBP

[Moc09h] Takuro Mochizuki. Parabolic L-Bradlow pairs. Lecture Notes in Mathematics, 1972:1–33, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-93913-9_3.pdf.

Mochizuki:2009:P

[Moc09i] Takuro Mochizuki. Preliminaries. Lecture Notes in Mathematics, 1972:1–38, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-93913-9_2.pdf.

Mochizuki:2009:VFC

[Moc09j] Takuro Mochizuki. Virtual fundamental classes. Lecture Notes in Mathematics, 1972:1–50, 2009. CODEN LNMAA2. ISBN 3-540-93912-1 (print), 3-540-93913-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-93913-9_6.pdf.

Moffatt:2009:SFD

[Mof09] H. K. Moffatt. Singularities in fluid dynamics and their resolution. Lecture Notes in Mathematics, 1973:157–166, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00837-5_5.pdf.

Molnar:2007:BM

[Mol07a] Lajos Molnár. Back matter. Lecture Notes in Mathematics, 1895: 205–236, 2007. CODEN LNMAA2. ISBN 3-540-39944-5 (print), 3-540-39946-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-39946-9/1.pdf.

Molnar:2007:FM

[Mol07b] Lajos Molnár. Front matter. Lecture Notes in Mathematics, 1895: I–IXL, 2007. CODEN LNMAA2. ISBN 3-540-39944-5 (print), 3-540-39946-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-39946-9/1.pdf.

Molnar:2007:LAL

[Mol07c] Lajos Molnár. Local automorphisms and local isometries of operator algebras and function algebras. Lecture Notes in Mathematics, 1895:159–204, 2007. CODEN LNMAA2. ISBN 3-540-39944-5 (print), 3-540-39946-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-39946-9_3.pdf.

Molnar:2007:PQS

[Mol07d] Lajos Molnár. Preservers on quantum structures. Lecture Notes in Mathematics, 1895:65–157, 2007. CODEN LNMAA2. ISBN 3-540-39944-5 (print), 3-540-39946-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-39946-9_2.pdf.

Molnar:2007:SPP

[Mol07e] Lajos Molnár. Selected Preserver Problems on Algebraic Structures of Linear Operators and on Function Spaces, volume 1895 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-39944-5 (print), 3-540-39946-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 232 pp. LCCN QA3 .L28 no. 1895. URL http://

link.springer.com/book/10.1007/3-540-39944-5;   http:
//www.springerlink.com/content/978-3-540-39946-9.

Molnar:2007:SLM

[Mol07f] Lajos Molnár. Some linear and multiplicative preserver problems on operator algebras and function algebras. Lecture Notes in Mathematics, 1895:29–64, 2007. CODEN LNMAA2. ISBN 3-540-39944-5 (print), 3-540-39946-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-39946-9_1.pdf.

Monod:2001:BMT

[Mon01a] Nicolas Monod. Banach modules, L_{∞} spaces. Lecture Notes in Mathematics, 1758:9–30, 2001. CODEN LNMAA2. ISBN 3-540-42054-1 (print), 3-540-44962-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44962-0_2/.

Monod:2001:CT

[Mon01b] Nicolas Monod. Cohomological techniques. Lecture Notes in Mathematics, 1758:129–168, 2001. CODEN LNMAA2. ISBN 3-540-42054-1 (print), 3-540-44962-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44962-0_5/.

Monod:2001:CBC

[Mon01c] Nicolas Monod. Continuous Bounded Cohomology of Locally Compact Groups, volume 1758 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42054-1 (paperback), 3-540-44962-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 214 pp. LCCN QA3 .L28 no. 1758; QA3 .L28 no. 1758; QA387. URL http:/

/link.springer.com/book/10.1007/b80626;   http://www.
loc.gov/catdir/enhancements/fy0816/2001032279-d.html;
 http://www.loc.gov/catdir/enhancements/fy0816/2001032279-
t.html.

Monod:2001:DCC

[Mon01d] Nicolas Monod. Definition and characterization of continuous bounded cohomology. Lecture Notes in Mathematics, 1758:61– 127, 2001. CODEN LNMAA2. ISBN 3-540-42054-1 (print), 3-540-44962-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44962-0_4/.

Monod:2001:I

[Mon01e] Nicolas Monod. Introduction. Lecture Notes in Mathematics, 1758:1–7, 2001. CODEN LNMAA2. ISBN 3-540-42054-1 (print), 3-540-44962-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44962-0_1/.

Monod:2001:RIA

[Mon01f] Nicolas Monod. Relative injectivity and amenable actions. Lecture Notes in Mathematics, 1758:31–60, 2001. CODEN LN-MAA2. ISBN 3-540-42054-1 (print), 3-540-44962-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44962-0_3/.

Monod:2001:TA

[Mon01g] Nicolas Monod. Towards applications. Lecture Notes in Mathematics, 1758:169–201, 2001. CODEN LNMAA2. ISBN 3-540-42054-1 (print), 3-540-44962-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44962-0_6/.

Monforte:2002:Ia

[Mon02a] Jorge Cortés Monforte. 1. Introduction. Lecture Notes in Mathematics, 1793:1–12, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45730-5_1/.

Monforte:2002:BGT

[Mon02b] Jorge Cortés Monforte. 2. Basic geometric tools. Lecture Notes in Mathematics, 1793:13–37, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45730-5_2/.

Monforte:2002:NS

[Mon02c] Jorge Cortés Monforte. 3. Nonholonomic systems. Lecture Notes in Mathematics, 1793:39–61, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45730-5_3/.

Monforte:2002:SNS

[Mon02d] Jorge Cortés Monforte. 4. Symmetries of nonholonomic systems. Lecture Notes in Mathematics, 1793:63–102, 2002. CODEN LN-MAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45730-5_

Monforte:2002:CS

[Mon02e] Jorge Cortés Monforte. 5. Chaplygin systems. Lecture Notes in Mathematics, 1793:103–120, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45730-5_5/.

Monforte:2002:CHN

[Mon02f] Jorge Cortés Monforte. 6. A class of hybrid nonholonomic systems. Lecture Notes in Mathematics, 1793:121–140, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-45730-5_6/.

Monforte:2002:NI

[Mon02g] Jorge Cortés Monforte. 7. Nonholonomic integrators. Lecture Notes in Mathematics, 1793:141–170, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45730-5_7/.

Monforte:2002:CMS

[Mon02h] Jorge Cortés Monforte. 8. Control of mechanical systems. Lecture Notes in Mathematics, 1793:171–202, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45730-5_8/.

Monforte:2002:BM

[Mon02i] Jorge Cortés Monforte. Back matter. Lecture Notes in Mathematics, 1793:203–224, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-45730-5/1.pdf.

Monforte:2002:FM

[Mon02j] Jorge Cortés Monforte. Front matter. Lecture Notes in Mathematics, 1793:I–XV, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-45730-5/1.pdf.

Monforte:2002:GCN

[Mon02k] Jorge Cortés Monforte. Geometric, Control and Numerical Aspects of Nonholonomic Systems, volume 1793 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 219 pp. LCCN QA3 .L28 no. 1793. URL http:/

/link.springer.com/book/10.1007/b84020;   http://www.
springerlink.com/content/978-3-540-45730-5.

Monforte:2002:Ib

[Mon02l] Jorge Cortés Monforte. Index. Lecture Notes in Mathematics, 1793:217–219, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45730-5_10/.

Monforte:2002:R

[Mon02m] Jorge Cortés Monforte. References. Lecture Notes in Mathematics, 1793:203–216, 2002. CODEN LNMAA2. ISBN 3-540-44154-9 (print), 3-540-45730-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45730-5_9/.

Moore:2001:BM

[Moo01a] John Douglas Moore. Back matter. Lecture Notes in Mathematics, 1629:117–121, 2001. CODEN LNMAA2. ISBN 3-540-41221-2 (print), 3-540-40952-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-40952-6/1.pdf.

Moore:2001:FM

[Moo01b] John Douglas Moore. Front matter. Lecture Notes in Mathematics, 1629:i–viii, 2001. CODEN LNMAA2. ISBN 3-540-41221-2 (print), 3-540-40952-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-40952-6/1.pdf.

Moore:2001:GAS

[Moo01c] John Douglas Moore. Global analysis of the Seiberg–Witten equations. Lecture Notes in Mathematics, 1629:73–116, 2001. CODEN LNMAA2. ISBN 3-540-41221-2 (print), 3-540-40952-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40952-6_
3/.

Moore:2001:LSW

[Moo01d] John Douglas Moore. Lectures on Seiberg–Witten Invariants, volume 1629 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-41221-2 (print), 3-540-40952-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 121 pp. LCCN QA3 .L28 no. 1629. URL http://link.

springer.com/book/10.1007/978-3-540-40952-6;   http:/
/www.springerlink.com/content/978-3-540-40952-6.

Moore:2001:P

[Moo01e] John Douglas Moore. Preliminaries. Lecture Notes in Mathematics, 1629:1–44, 2001. CODEN LNMAA2. ISBN 3-540-41221-2 (print), 3-540-40952-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40952-6_1/.

Moore:2001:SGF

[Moo01f] John Douglas Moore. Spin geometry on four-manifolds. Lecture Notes in Mathematics, 1629:45–72, 2001. CODEN LNMAA2. ISBN 3-540-41221-2 (print), 3-540-40952-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40952-6_2/.

Morse:2008:LNL

[Mor08] A. Stephen Morse. Lecture notes on logically switched dynamical systems. Lecture Notes in Mathematics, 1932:61–161, 2008. CODEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77653-6_2.pdf.

Moscoso:2008:IIR

[Mos08a] Miguel Moscoso. Introduction to image reconstruction. Lecture Notes in Mathematics, 1943:1–16, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78547-7_1.
pdf.

Moscoso:2008:PBO

[Mos08b] Miguel Moscoso. Polarization-based optical imaging. Lecture Notes in Mathematics, 1943:67–83, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78547-7_4.
pdf.

Molto:2009:CCC

[MOTV09a] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. σ-continuous and co- σ-continuous maps. Lecture Notes in Mathematics, 1951:13–47, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-85031-1_2.pdf.

Molto:2009:SCM

[MOTV09b] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. σ-slicely continuous maps. Lecture Notes in Mathematics, 1951:73–99, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85031-1_4.pdf.

Molto:2009:BM

[MOTV09c] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. Back matter. Lecture Notes in Mathematics, 1951: 131–148, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-85031-1/1.pdf.

Molto:2009:FM

[MOTV09d] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. Front matter. Lecture Notes in Mathematics, 1951: I–XI, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-85031-1/1.pdf.

Molto:2009:GMS

[MOTV09e] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. Generalized metric spaces and locally uniformly rotund renormings. Lecture Notes in Mathematics, 1951:49–72, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85031-1_3.pdf.

Molto:2009:I

[MOTV09f] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. Introduction. Lecture Notes in Mathematics, 1951:1– 11, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85031-1_1.pdf.

Molto:2009:NTT

[MOTV09g] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. A Nonlinear Transfer Technique for Renorming, volume 1951 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 142 pp. LCCN QA3 .L28 no. 1951. URL http://link.

springer.com/book/10.1007/978-3-540-85031-1;   http:/
/www.springerlink.com/content/978-3-540-85031-1.

Molto:2009:SA

[MOTV09h] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. Some applications. Lecture Notes in Mathematics, 1951:101–116, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85031-1_5.pdf.

Molto:2009:SOP

[MOTV09i] Aníbal Moltó, José Orihuela, Stanimir Troyanski, and Manuel Valdivia. Some open problems. Lecture Notes in Mathematics, 1951:117–129, 2009. CODEN LNMAA2. ISBN 3-540-85030-9 (print), 3-540-85031-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85031-1_6.pdf.

Mouton:2000:CAF

[Mou00] Frédéric Mouton. Comportement asymptotique des fonctions harmoniques sur les arbres. (French) []. Lecture Notes in Mathematics, 1729:353–373, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103813/.

Mountford:2003:BSL

[Mou03a] Thomas S. Mountford. Brownian sheet local time and bubbles. Lecture Notes in Mathematics, 1832:196–215, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40004-2_
7/.

Mountford:2003:CRA

[Mou03b] Thomas S. Mountford. Critical reversible attractive nearest particle systems. Lecture Notes in Mathematics, 1802:209–241, 2003. CODEN LNMAA2. ISBN 3-540-00295-2 (print), 3-540-36259-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36259-3_6/.

Mallet-Paret:2003:TWS

[MP03] John Mallet-Paret. Traveling waves in spatially discrete dynamical systems of diffusive type. Lecture Notes in Mathematics, 1822: 231–298, 2003. CODEN LNMAA2. ISBN 3-540-40786-3 (print), 3-540-45204-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45204-1_4/.

Milman:2004:EUE

[MP04] V. D. Milman and A. Pajor. Essential uniqueness of an Mellipsoid of a given convex body. Lecture Notes in Mathematics, 1850:237–241, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44489-3_18/.

Massart:2007:CIM

[MP07] Pascal Massart and Jean Picard, editors. Concentration Inequalities and Model Selection: École d’Été de Probabilités de Saint-Flour XXXIII — 2003, volume 1896 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LN-MAA2. ISBN 3-540-48497-3 (paperback), 3-540-48503-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1896.  URL http://link.springer.com/
book/10.1007/978-3-540-48503-2;   http://www.loc.gov/
catdir/enhancements/fy0825/2007921691-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0825/2007921691-d.
html;   http://www.loc.gov/catdir/enhancements/fy0825/
2007921691-t.html;   http://www.springerlink.com/content/
978-3-540-48503-2.

Mikhailov:2009:AP

[MP09a] Roman Mikhailov and Inder Bir Singh Passi. Augmentation powers. Lecture Notes in Mathematics, 1952:187–227, 2009. CODEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85818-8_4.pdf.

Mikhailov:2009:BM

[MP09b] Roman Mikhailov and Inder Bir Singh Passi. Back matter. Lecture Notes in Mathematics, 1952:299–352, 2009. CODEN LN-MAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-85818-8/
1.pdf.

Mikhailov:2009:DSb

[MP09c] Roman Mikhailov and Inder Bir Singh Passi. Derived series. Lecture Notes in Mathematics, 1952:165–185, 2009. CO-DEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85818-8_3.pdf.

Mikhailov:2009:DSa

[MP09d] Roman Mikhailov and Inder Bir Singh Passi. Dimension subgroups. Lecture Notes in Mathematics, 1952:101–164, 2009. CODEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85818-8_2.pdf.

Mikhailov:2009:FM

[MP09e] Roman Mikhailov and Inder Bir Singh Passi. Front matter. Lecture Notes in Mathematics, 1952:i–xxi, 2009. CODEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-85818-8/1.pdf.

Mikhailov:2009:HA

[MP09f] Roman Mikhailov and Inder Bir Singh Passi. Homotopical aspects. Lecture Notes in Mathematics, 1952:229–290, 2009. CODEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85818-8_5.pdf.

Mikhailov:2009:LCD

[MP09g] Roman Mikhailov and Inder Bir Singh Passi. Lower Central and Dimension Series of Groups, volume 1952 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 346 pp. LCCN

QA3 .L28 no. 1952. URL http://link.springer.com/book/
10.1007/978-3-540-85818-8;   http://www.springerlink.
com/content/978-3-540-85818-8.

Mikhailov:2009:LCS

[MP09h] Roman Mikhailov and Inder Bir Singh Passi. Lower central series. Lecture Notes in Mathematics, 1952:1–100, 2009. CODEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85818-8_1.pdf.

Mikhailov:2009:M

[MP09i] Roman Mikhailov and Inder Bir Singh Passi. Miscellanea. Lecture Notes in Mathematics, 1952:291–297, 2009. CO-DEN LNMAA2. ISBN 3-540-85817-2 (print), 3-540-85818-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85818-8_6.pdf.

Miclo:2000:TSP

[MR00] Laurent Miclo and Cyril Roberto. Trous spectraux pour certains algorithmes de Metropolis sur R. (French) []. Lecture Notes in Mathematics, 1729:336–352, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103812/.

Magal:2008:SPM

[MR08] Pierre Magal and Shigui Ruan. Structured Population Models in Biology and Epidemiology, volume 1936 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 302 pp. LCCN QA3 .L28 no.

1936. URL http://link.springer.com/book/10.1007/978-
3-540-78273-5;   http://www.springerlink.com/content/
978-3-540-78273-5.

Meeks:2002:GTM

[MRR02] William H. Meeks III, Antonio Ros, and Harold Rosenberg, editors. The Global Theory of Minimal Surfaces in Flat Spaces: Lectures given at the 2nd Session of the Centro Internazionale Matematico Estivo (C.I.M.E.) held in Martina Franca, Italy July 7-14, 1999, volume 1775 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43120-9 (print), 3-540-45609-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1775. URL http:/

/link.springer.com/book/10.1007/b83168;   http://www.
springerlink.com/content/978-3-540-45609-4.

Maitra:2000:SGH

[MS00a] A. Maitra and W. Sudderth. Saturations of gambling houses. Lecture Notes in Mathematics, 1729:218–238, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103805/.

Milman:2000:GLD

[MS00b] V. D. Milman and S. J. Szarek. A geometric lemma and duality of entropy numbers. Lecture Notes in Mathematics, 1745:

191–222, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0107215/.

Milman:2000:GAF

[MS00c] Vitali D. Milman and Gideon Schechtman, editors. Geometric Aspects of Functional Analysis: Israel Seminar 1996– 2000, volume 1745 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1745. URL http:/

/link.springer.com/book/10.1007/BFb0107201;   http://
www.springerlink.com/content/978-3-540-45392-5.

Miermont:2003:SSF

[MS03a] Grégory Miermont and Jason Schweinsberg. Self-similar fragmentations and stable subordinators. Lecture Notes in Mathematics, 1832:333–359, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40004-2_13/.

Milman:2003:GAF

[MS03b] Vitali D. Milman and Gideon Schechtman, editors. Geometric Aspects of Functional Analysis: Israel Seminar 2001-2002, volume 1807 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1807. URL http:/

/link.springer.com/book/10.1007/b10415;   http://www.
springerlink.com/content/978-3-540-36428-3.

Milman:2004:GAF

[MS04] Vitali D. Milman and Gideon Schechtman, editors. Geometric Aspects of Functional Analysis: Israel Seminar 2002-2003, volume 1850 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1850. URL http:/

/link.springer.com/book/10.1007/b98686;   http://www.
springerlink.com/content/978-3-540-44489-3.

Milman:2007:GAF

[MS07] Vitali D. Milman and Gideon Schechtman, editors. Geometric Aspects of Functional Analysis: Israel Seminar 2004–2005, volume 1910 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1910. URL http://link.springer.com/book/
10.1007/978-3-540-72053-9;   http://www.springerlink.
com/content/978-3-540-72053-9.

Martcheva:2008:IOS

[MT08] M. Martcheva and H. R. Thieme. Infinite ODE systems modeling size-structured metapopulations, macroparasitic diseases, and prion proliferation. Lecture Notes in Mathematics, 1936:51–113, 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78273-5_2.pdf.

Milman:2000:SAS

[MTJ00] V. D. Milman and N. Tomczak-Jaegermann. Stabilized asymptotic structures and envelopes in Banach spaces. Lecture Notes in Mathematics, 1745:223–237, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0107216/.

Mueller:2009:STR

[Mue09] Carl Mueller. Some tools and results for parabolic stochastic partial differential equations. Lecture Notes in Mathematics, 1962:111–144, 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85994-9_4.pdf.

Marcellan:2006:OPS

[MV06] Francisco Marcellán and Walter Van Assche. Orthogonal Polynomials and Special Functions: Computation and Applications, volume 1883 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 418 pp. LCCN QA3 .L28 no. 1883; QA3 .L28 no. 1883; QA404.5 .O735 2006. URL http://

link.springer.com/book/10.1007/b128597;   http://www.
springerlink.com/content/978-3-540-36716-1;   http://
www.springerlink.com/openurl.asp?genre=issue&issn=0075-
8434&volume=1883.

Milman:2003:SRL

[MW03] Vitali Milman and Roy Wagner. Some remarks on a lemma of Ran Raz. Lecture Notes in Mathematics, 1807:158–168, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36428-3_14/.

Marmi:2002:BM

[MY02a] Stefano Marmi and Jean-Christophe Yoccoz. Back matter. Lecture Notes in Mathematics, 1784:193–199, 2002. CODEN LN-MAA2. ISBN 3-540-43726-6 (print), 3-540-47928-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-47928-4/
1.pdf.

Marmi:2002:SOP

[MY02b] Stefano Marmi and Jean-Christophe Yoccoz. Some open problems related to small divisors. Lecture Notes in Mathematics, 1784:175–191, 2002. CODEN LNMAA2. ISBN 3-540-43726-6 (print), 3-540-47928-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-47928-4_4/.

Madan:2006:IIF

[MY06a] Dilip B. Madan and Marc Yor. Itô’s integrated formula for strict local martingales. Lecture Notes in Mathematics, 1874:157–170, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_13.pdf.

Mansuy:2006:NC

[MY06b] Roger Mansuy and Marc Yor. Notation and convention. Lecture Notes in Mathematics, 1873:1–39, 2006. CODEN LNMAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-32416-X_1.pdf.

Mansuy:2006:MWV

[MY06c] Roger Mansuy and Marc Yor. On the martingales which vanish on the set of Brownian zeroes. Lecture Notes in Mathematics, 1873:53–69, 2006. CODEN LNMAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-32416-X_3.pdf.

Mansuy:2006:PCR

[MY06d] Roger Mansuy and Marc Yor. Predictable and chaotic representation properties for some remarkable martingales including the Azéma and the Dunkl martingales. Lecture Notes in Mathematics, 1873:71–86, 2006. CODEN LNMAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-32416-X_4.pdf.

Mansuy:2006:RTE

[MY06e] Roger Mansuy and Marc Yor. Random Times and Enlargements of Filtrations in a Brownian Setting, volume 1873 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LN-MAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 158 pp.

LCCN QA3 .L28 no. 1873. URL http://link.springer.com/
book/10.1007/11415558;   http://www.springerlink.com/
content/978-3-540-32416-4.

Mansuy:2006:SSE

[MY06f] Roger Mansuy and Marc Yor. Sketches of solutions for the exercises. Lecture Notes in Mathematics, 1873:117–139, 2006. CODEN LNMAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-32416-X_7.pdf.

Mansuy:2006:SNS

[MY06g] Roger Mansuy and Marc Yor. Stopping and non-stopping times. Lecture Notes in Mathematics, 1873:41–51, 2006. CODEN LN-MAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/3-540-32416-X_
2.pdf.

Mansuy:2006:UBP

[MY06h] Roger Mansuy and Marc Yor. Unveiling the Brownian path (or history) as the level rises. Lecture Notes in Mathematics, 1873:87–102, 2006. CODEN LNMAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-32416-X_5.pdf.

Mansuy:2006:WSB

[MY06i] Roger Mansuy and Marc Yor. Weak and strong Brownian filtrations. Lecture Notes in Mathematics, 1873:103–116, 2006. CODEN LNMAA2. ISBN 3-540-29407-4 (print), 3-540-32416-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-32416-X_6.pdf.

Ma:2007:AF

[MY07a] Jin Ma and Jiongmin Yong. Applications of FBSDEs. Lecture Notes in Mathematics, 1702:193–234, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-48831-6_8.
pdf.

Ma:2007:BM

[MY07b] Jin Ma and Jiongmin Yong. Back matter. Lecture Notes in Mathematics, 1702:257–274, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-48831-6/1.pdf.

Ma:2007:FR

[MY07c] Jin Ma and Jiongmin Yong. FBSDEs with reflections. Lecture Notes in Mathematics, 1702:169–192, 2007. CODEN LNMAA2.

ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-48831-6_7.
pdf.

Ma:2007:FBS

[MY07d] Jin Ma and Jiongmin Yong. Forward-Backward Stochastic Differential Equations and their Applications, volume 1702 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 274 (est.) pp. LCCN

QA3 .L28 no. 1702. URL http://link.springer.com/book/
10.1007/978-3-540-48831-6;   http://www.springerlink.
com/content/978-3-540-48831-6.

Ma:2007:FSS

[MY07e] Jin Ma and Jiongmin Yong. Four step scheme. Lecture Notes in Mathematics, 1702:80–102, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48831-6_4.pdf.

Ma:2007:FM

[MY07f] Jin Ma and Jiongmin Yong. Front matter. Lecture Notes in Mathematics, 1702:i–xiii, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-48831-6/1.pdf.

Ma:2007:I

[MY07g] Jin Ma and Jiongmin Yong. Introduction. Lecture Notes in Mathematics, 1702:1–24, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-48831-6_1.pdf.

Ma:2007:LDB

[MY07h] Jin Ma and Jiongmin Yong. Linear, degenerate backward stochastic partial differential equations. Lecture Notes in Mathematics, 1702:103–136, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-48831-6_5.pdf.

Ma:2007:LE

[MY07i] Jin Ma and Jiongmin Yong. Linear equations. Lecture Notes in Mathematics, 1702:25–50, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-48831-6_2.pdf.

Ma:2007:MC

[MY07j] Jin Ma and Jiongmin Yong. The method of continuation. Lecture Notes in Mathematics, 1702:137–168, 2007. CO-DEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48831-6_6.pdf.

Ma:2007:MOC

[MY07k] Jin Ma and Jiongmin Yong. Method of optimal control. Lecture Notes in Mathematics, 1702:51–79, 2007. CODEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-48831-6_3.
pdf.

Ma:2007:NMF

[MY07l] Jin Ma and Jiongmin Yong. Numerical methods for FBS-DEs. Lecture Notes in Mathematics, 1702:235–256, 2007. CO-DEN LNMAA2. ISBN 3-540-65960-9 (print), 3-540-48831-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-48831-6_9.pdf.

Najnudel:2007:IRS

[Naj07] Joseph Najnudel. Integration with respect to self-intersection local time of a one-dimensional Brownian motion. Lecture Notes in Mathematics, 1899:105–116, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71189-6_4.pdf.

Natterer:2008:XRT

[Nat08] Frank Natterer. X-ray tomography. Lecture Notes in Mathematics, 1943:17–34, 2008. CODEN LNMAA2. ISBN 3-540-78545-0 (print), 3-540-78547-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78547-7_2.pdf.

Nazarov:2003:MPC

[Naz03a] Fedor Nazarov. On the maximal perimeter of a convex set in R^{n} with respect to a Gaussian measure. Lecture Notes in Mathematics, 1807:169–187, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36428-3_15/.

Nazarov:2003:MHL

[Naz03b] Maxim Nazarov. Mixed hook-length formula for degenerate a fine Hecke algebras. Lecture Notes in Mathematics, 1815:223– 236, 2003. CODEN LNMAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44890-X_10/.

Nuno:2008:MMI

[NCCFM08] M. Nuño, C. Castillo-Chavez, Z. Feng, and M. Martcheva. Mathematical models of influenza: The role of cross-immunity, quarantine and age-structure. Lecture Notes in Mathematics, 1945:349– 364, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_13.pdf.

Nemirovski:2000:AE

[Nem00a] Arkadi Nemirovski. Aggregation of estimates, I. Lecture Notes in Mathematics, 1738:183–206, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0106715/.

Nemirovski:2000:AEI

[Nem00b] Arkadi Nemirovski. Aggregation of estimates, II. Lecture Notes in Mathematics, 1738:207–227, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0106716/.

Nemirovski:2000:EF

[Nem00c] Arkadi Nemirovski. Estimating functionals, I. Lecture Notes in Mathematics, 1738:228–257, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0106717/.

Nemirovski:2000:EFI

[Nem00d] Arkadi Nemirovski. Estimating functionals, II. Lecture Notes in Mathematics, 1738:258–277, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0106718/.

Nemirovski:2000:ERFa

[Nem00e] Arkadi Nemirovski. Estimating regression functions from Hölder balls. Lecture Notes in Mathematics, 1738:89–112, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0106711/.

Nemirovski:2000:ERFb

[Nem00f] Arkadi Nemirovski. Estimating regression functions from Sobolev balls. Lecture Notes in Mathematics, 1738:113–131, 2000. CO-DEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0106712/.

Nemirovski:2000:ESS

[Nem00g] Arkadi Nemirovski. Estimating signals satisfying differential inequalities. Lecture Notes in Mathematics, 1738:155–182, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0106714/.

Nemirovski:2000:P

[Nem00h] Arkadi Nemirovski. Preface. Lecture Notes in Mathematics, 1738: 88, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106710.

Nemirovski:2000:SAE

[Nem00i] Arkadi Nemirovski. Spatial adaptive estimation on Sobolev balls. Lecture Notes in Mathematics, 1738:132–154, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0106713/.

Nesterenko:2003:LFL

[Nes03] Yuri Nesterenko. Linear forms in logarithms of rational numbers. Lecture Notes in Mathematics, 1819:53–106, 2003. CODEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44979-5_2/.

Nagylaki:2008:DMS

[NL08] T. Nagylaki and Y. Lou. The dynamics of migration–selection models. Lecture Notes in Mathematics, 1922:117–170, 2008. CODEN LNMAA2. ISBN 3-540-74328-6 (print), 3-540-74331-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74331-6_4.pdf.

Nguyen-Ngoc:2005:SMA

[NNY05] Laurent Nguyen-Ngoc and Marc Yor. Some martingales associated to reflected Lévy processes. Lecture Notes in Mathematics, 1857:42–69, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-31449-3_5.pdf.

Nourdin:2008:STS

[Nou08] Ivan Nourdin. A simple theory for the study of SDEs driven by a fractional Brownian motion, in dimension one. Lecture Notes in Mathematics, 1934:181–197, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-77913-1_8.pdf.

Nesterenko:2001:TTZ

[NP01a] Yuri V. Nesterenko and Patrice Philippon. Θ(τ, z) and transcendence. Lecture Notes in Mathematics, 1752:1–11, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44550-1_1/.

Nesterenko:2001:AIV

[NP01b] Yuri V. Nesterenko and Patrice Philippon. Algebraic independence for values of Ramanujan functions. Lecture Notes in Mathematics, 1752:27–46, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44550-1_3/.

Nesterenko:2001:AIAa

[NP01c] Yuri V. Nesterenko and Patrice Philippon. Algebraic independence in algebraic groups. Part I: Small transcendence degrees. Lecture Notes in Mathematics, 1752:199–211, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44550-1_13/.

Nesterenko:2001:AIAb

[NP01d] Yuri V. Nesterenko and Patrice Philippon. Algebraic independence in algebraic groups. Part II: Large transcendence degrees. Lecture Notes in Mathematics, 1752:213–225, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44550-1_14/.

Nesterenko:2001:CAI

[NP01e] Yuri V. Nesterenko and Patrice Philippon. Criteria for algebraic independence. Lecture Notes in Mathematics, 1752:133– 141, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44550-1_8/.

Nesterenko:2001:DG

[NP01f] Yuri V. Nesterenko and Patrice Philippon. Diophantine geometry. Lecture Notes in Mathematics, 1752:83–94, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44550-1_6/.

Nesterenko:2001:EM

[NP01g] Yuri V. Nesterenko and Patrice Philippon. Élimination multihomogène. (French) [multihomogeneous elimination]. Lecture Notes in Mathematics, 1752:53–81, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44550-1_5/.

Nesterenko:2001:GDM

[NP01h] Yuri V. Nesterenko and Patrice Philippon. Géométrie diophantienne multiprojective. (French) [Multiprojective Diophantine geometry]. Lecture Notes in Mathematics, 1752:95–131, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44550-1_7/.

Nesterenko:2001:HNI

[NP01i] Yuri V. Nesterenko and Patrice Philippon. The Hilbert Nullstellensatz, inequalities for polynomials, and algebraic independence. Lecture Notes in Mathematics, 1752:239–248, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44550-1_16/.

Nesterenko:2001:IAI

[NP01j] Yuri V. Nesterenko and Patrice Philippon. Introduction to Algebraic Independence Theory, volume 1752 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 256 pp. LCCN QA3 .L28 no. 1752. URL http:/

/link.springer.com/book/10.1007/b76882;   http://www.
springerlink.com/content/978-3-540-44550-0.

Nesterenko:2001:MCO

[NP01k] Yuri V. Nesterenko and Patrice Philippon. Mahler’s conjecture and other transcendence results. Lecture Notes in Mathematics, 1752:13–26, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44550-1_2/.

Nesterenko:2001:MAI

[NP01l] Yuri V. Nesterenko and Patrice Philippon. Measures of algebraic independence for Mahler functions. Lecture Notes in Mathematics, 1752:187–197, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44550-1_12/.

Nesterenko:2001:MES

[NP01m] Yuri V. Nesterenko and Patrice Philippon. Multiplicity estimates for solutions of algebraic differential equations. Lecture Notes in Mathematics, 1752:149–165, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44550-1_10/.

Nesterenko:2001:SMR

[NP01n] Yuri V. Nesterenko and Patrice Philippon. Some metric results in Transcendental Numbers Theory. Lecture Notes in Mathematics, 1752:227–237, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44550-1_15/.

Nesterenko:2001:SRP

[NP01o] Yuri V. Nesterenko and Patrice Philippon. Some remarks on proofs of algebraic independence. Lecture Notes in Mathematics, 1752:47–51, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44550-1_4/.

Nesterenko:2001:UBG

[NP01p] Yuri V. Nesterenko and Patrice Philippon. Upper bounds for (geometric) Hilbert functions. Lecture Notes in Mathematics, 1752: 143–148, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44550-1_9/.

Nesterenko:2001:ZEC

[NP01q] Yuri V. Nesterenko and Patrice Philippon. Zero estimates on commutative algebraic groups. Lecture Notes in Mathematics, 1752:167–185, 2001. CODEN LNMAA2. ISBN 3-540-41496-7 (print), 3-540-44550-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44550-1_11/.

Newman:2007:BM

[NS07a] Charles M. Newman and Daniel L. Stein. Back matter. Lecture Notes in Mathematics, 1900:177–182, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-540-40908-3/1.pdf.

Newman:2007:LVG

[NS07b] Charles M. Newman and Daniel L. Stein. Local vs. global variables for spin glasses. Lecture Notes in Mathematics, 1900:145– 158, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-40908-3_6.pdf.

Newman:2007:SRS

[NS07c] Charles M. Newman and Daniel L. Stein. Short-range spin glasses: Results and speculations. Lecture Notes in Mathematics, 1900:159–175, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-40908-3_7.pdf.

Nakkajima:2008:BM

[NS08a] Yukiyoshi Nakkajima and Atsushi Shiho. Back matter. Lecture Notes in Mathematics, 1959:249–266, 2008. CODEN LNMAA2. ISBN 3-540-70564-3 (print), 3-540-70565-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-540-70565-9/1.pdf.

Nakkajima:2008:FM

[NS08b] Yukiyoshi Nakkajima and Atsushi Shiho. Front matter. Lecture Notes in Mathematics, 1959:i–xxiii, 2008. CODEN LNMAA2. ISBN 3-540-70564-3 (print), 3-540-70565-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-70565-9/1.pdf.

Nakkajima:2008:PFD

[NS08c] Yukiyoshi Nakkajima and Atsushi Shiho. Preliminaries on filtered derived categories and topoi. Lecture Notes in Mathematics, 1959:15–53, 2008. CODEN LNMAA2. ISBN 3-540-70564-3 (print), 3-540-70565-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-70565-9_1.pdf.

Nakkajima:2008:WFS

[NS08d] Yukiyoshi Nakkajima and Atsushi Shiho. Weight filtrations and slope filtrations on rigid cohomologies (summary). Lecture Notes in Mathematics, 1959:219–248, 2008. CODEN LNMAA2. ISBN 3-540-70564-3 (print), 3-540-70565-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-70565-9_3.pdf.

Nakkajima:2008:WFLa

[NS08e] Yukiyoshi Nakkajima and Atsushi Shiho. Weight filtrations on log crystalline cohomologies. Lecture Notes in Mathematics, 1959:55– 217, 2008. CODEN LNMAA2. ISBN 3-540-70564-3 (print), 3-540-70565-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-70565-9_2.pdf.

Nakkajima:2008:WFLb

[NS08f] Yukiyoshi Nakkajima and Atsushi Shiho. Weight Filtrations on Log Crystalline Cohomologies of Families of Open Smooth Varieties, volume 1959 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-70564-3 (print), 3-540-70565-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 266 pp. LCCN QA3 .L28 no. 1959. URL http:

//link.springer.com/book/10.1007/978-3-540-70565-9;
http://www.springerlink.com/content/978-3-540-70565-9.

Nistri:2008:NOC

[NS08g] Paolo Nistri and Gianna Stefani, editors. Nonlinear and Optimal Control Theory: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy June 19–29, 2004, volume 1932 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CO-DEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1932. URL http://link.springer.com/book/
10.1007/978-3-540-77653-6;   http://www.springerlink.
com/content/978-3-540-77653-6.

Nagasawa:2000:TDS

[NT00] Masao Nagasawa and Hiroshi Tanaka. Time dependent subordination and Markov processes with jumps. Lecture Notes in Mathematics, 1729:257–288, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103807/.

Nagasawa:2001:PVR

[NT01] Masao Nagasawa and Hiroshi Tanaka. The principle of variation for relativistic quantum particles. Lecture Notes in Mathematics, 1755:1–27, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44671-2_1/.

Nualart:2009:AMC

[Nua09] David Nualart. Application of Malliavin calculus to stochastic partial differential equations. Lecture Notes in Mathematics, 1962:73–109, 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-85994-9_3.pdf.

Nussbaum:2003:LPS

[Nus03] Roger D. Nussbaum. Limiting profiles for solutions of differentialdelay equations. Lecture Notes in Mathematics, 1822:299–342, 2003. CODEN LNMAA2. ISBN 3-540-40786-3 (print), 3-540-45204-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45204-1_5/.

Nastasescu:2004:CGR

[NV04a] Constantin Nastasescu and Freddy Van Oystaeyen. 1. The category of graded rings. Lecture Notes in Mathematics, 1836:1– 18, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_1/.

Nastasescu:2004:CGM

[NV04b] Constantin Nastasescu and Freddy Van Oystaeyen. 2. The category of graded modules. Lecture Notes in Mathematics, 1836: 19–79, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_2/.

Nastasescu:2004:MSG

[NV04c] Constantin Nastasescu and Freddy Van Oystaeyen. 3. Modules over strongly graded rings. Lecture Notes in Mathematics, 1836: 81–113, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_3/.

Nastasescu:2004:GCT

[NV04d] Constantin Nastasescu and Freddy Van Oystaeyen. 4. Graded Clifford theory. Lecture Notes in Mathematics, 1836:115–145, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_4/.

Nastasescu:2004:IH

[NV04e] Constantin Nastasescu and Freddy Van Oystaeyen. 5. Internal homogenization. Lecture Notes in Mathematics, 1836:147– 165, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_5/.

Nastasescu:2004:EH

[NV04f] Constantin Nastasescu and Freddy Van Oystaeyen. 6. External homogenization. Lecture Notes in Mathematics, 1836:167– 185, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_6/.

Nastasescu:2004:SP

[NV04g] Constantin Nastasescu and Freddy Van Oystaeyen. 7. Smash products. Lecture Notes in Mathematics, 1836:187–221, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-40998-4_7/.

Nastasescu:2004:LGR

[NV04h] Constantin Nastasescu and Freddy Van Oystaeyen. 8. Localization of graded rings. Lecture Notes in Mathematics, 1836:223– 240, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_8/.

Nastasescu:2004:AG

[NV04i] Constantin Nastasescu and Freddy Van Oystaeyen. 9. Application to gradability. Lecture Notes in Mathematics, 1836:241– 276, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40998-4_9/.

Nastasescu:2004:A

[NV04j] Constantin Nastasescu and Freddy Van Oystaeyen. Appendix. Lecture Notes in Mathematics, 1836:277–289, 2004. CODEN LN-MAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40998-4_
10/.

Nastasescu:2004:B

[NV04k] Constantin Nastasescu and Freddy Van Oystaeyen. Bibliography. Lecture Notes in Mathematics, 1836:291–302, 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40998-4_
11/.

Nastasescu:2004:MGR

[NV04l] Constantin Nastasescu and Freddy Van Oystaeyen. Methods of Graded Rings, volume 1836 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-20746-5 (print), 3-540-40998-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 304 pp. LCCN QA3 .L28 no. 1836. URL

http://link.springer.com/book/10.1007/b94904;   http:
//www.springerlink.com/content/978-3-540-40998-4.

Obloj:2007:MPR

[Obł07] Jan Obłój. The maximality principle revisited: On certain optimal stopping problems. Lecture Notes in Mathematics, 1899:309– 328, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_16.pdf.

OConnell:2003:RMN

[O’C03] Neil O’Connell. Random matrices, non-colliding processes and queues. Lecture Notes in Mathematics, 1801:165–182, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_3/.

Okounkov:2003:RTM

[Oko03] Andrei Okounkov. Random trees and moduli of curves. Lecture Notes in Mathematics, 1815:89–126, 2003. CODEN LN-MAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44890-X_5/.

Oleszkiewicz:2003:PRV

[Ole03] Krzysztof Oleszkiewicz. On p-pseudostable random variables, Rosenthal spaces and l_{p}^{n} ball slicing. Lecture Notes in Mathematics, 1807:188–210, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36428-3_16/.

Oliva:2002:DM

[Oli02a] Waldyr Muniz Oliva. 1. Differentiable manifolds. Lecture Notes in Mathematics, 1798:3–12, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45795-4_2/.

Oliva:2002:VFD

[Oli02b] Waldyr Muniz Oliva. 2. Vector fields, differential forms and tensor fields. Lecture Notes in Mathematics, 1798:13–21, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45795-4_3/.

Oliva:2002:PRM

[Oli02c] Waldyr Muniz Oliva. 3. Pseudo-Riemannian manifolds. Lecture Notes in Mathematics, 1798:23–53, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45795-4_4/.

Oliva:2002:NM

[Oli02d] Waldyr Muniz Oliva. 4. Newtonian mechanics. Lecture Notes in Mathematics, 1798:55–60, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45795-4_5/.

Oliva:2002:MSR

[Oli02e] Waldyr Muniz Oliva. 5. Mechanical systems on Riemannian manifolds. Lecture Notes in Mathematics, 1798:61–110, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45795-4_6/.

Oliva:2002:MSN

[Oli02f] Waldyr Muniz Oliva. 6. Mechanical systems with non-holonomic constraints. Lecture Notes in Mathematics, 1798:111–126, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45795-4_7/.

Oliva:2002:HAS

[Oli02g] Waldyr Muniz Oliva. 7. Hyperbolicity and Anosov systems. Vakonomic mechanics. Lecture Notes in Mathematics, 1798:127– 143, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45795-4_8/.

Oliva:2002:SR

[Oli02h] Waldyr Muniz Oliva. 8. Special relativity. Lecture Notes in Mathematics, 1798:145–163, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45795-4_9/.

Oliva:2002:GR

[Oli02i] Waldyr Muniz Oliva. 9. General relativity. Lecture Notes in Mathematics, 1798:165–181, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45795-4_10/.

Oliva:2002:HLF

[Oli02j] Waldyr Muniz Oliva. A. Hamiltonian and Lagrangian formalisms. Lecture Notes in Mathematics, 1798:183–193, 2002. CODEN LN-MAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45795-4_
11/.

Oliva:2002:BMT

[Oli02k] Waldyr Muniz Oliva. B. Möbius transformations and the Lorentz group. Lecture Notes in Mathematics, 1798:195–221, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45795-4_12/.

Oliva:2002:CQM

[Oli02l] Waldyr Muniz Oliva. C. Quasi-Maxwell form of Einstein’s equation. Lecture Notes in Mathematics, 1798:223–244, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45795-4_

Oliva:2002:DVS

[Oli02m] Waldyr Muniz Oliva. D. Viscosity solutions and Aubry– Mather theory. Lecture Notes in Mathematics, 1798:245–257, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45795-4_14/.

Oliva:2002:GM

[Oli02n] Waldyr Muniz Oliva. Geometric Mechanics, volume 1798 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 270 pp. LCCN QA3 .L28 no. 1798. URL http:/

/link.springer.com/book/10.1007/b84214;   http://www.
springerlink.com/content/978-3-540-45795-4.

Oliva:2002:Ib

[Oli02o] Waldyr Muniz Oliva. Index. Lecture Notes in Mathematics, 1798: 263–270, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45795-4_16/.

Oliva:2002:Ia

[Oli02p] Waldyr Muniz Oliva. Introduction. Lecture Notes in Mathematics, 1798:1–2, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45795-4_1.

Oliva:2002:R

[Oli02q] Waldyr Muniz Oliva. References. Lecture Notes in Mathematics, 1798:259–261, 2002. CODEN LNMAA2. ISBN 3-540-44242-1 (print), 3-540-45795-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45795-4_15/.

Olshanski:2003:IHA

[Ols03] Grigori Olshanski. An introduction to harmonic analysis on the infinite symmetric group. Lecture Notes in Mathematics, 1815:

127–160, 2003. CODEN LNMAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44890-X_6/.

Olsson:2008:BM

[Ols08a] Martin C. Olsson. Back matter. Lecture Notes in Mathematics, 1958:273–278, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-70519-2/1.pdf.

Olsson:2008:BPA

[Ols08b] Martin C. Olsson. A brief primer on algebraic stacks. Lecture Notes in Mathematics, 1958:7–29, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-70519-2_2.
pdf.

Olsson:2008:CMS

[Ols08c] Martin C. Olsson. Compactifying Moduli Spaces for Abelian Varieties, volume 1958 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 278 pp. LCCN QA3 .L28 no. 1958. URL http:

//link.springer.com/book/10.1007/978-3-540-70519-2;
http://www.springerlink.com/content/978-3-540-70519-2.

Olsson:2008:FM

[Ols08d] Martin C. Olsson. Front matter. Lecture Notes in Mathematics, 1958:I–VII, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-70519-2/1.pdf.

Olsson:2008:I

[Ols08e] Martin C. Olsson. Introduction. Lecture Notes in Mathematics, 1958:1–5, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-70519-2_1.pdf.

Olsson:2008:LS

[Ols08f] Martin C. Olsson. Level structure. Lecture Notes in Mathematics, 1958:231–271, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-70519-2_7.pdf.

Olsson:2008:MAV

[Ols08g] Martin C. Olsson. Moduli of Abelian varieties with higher degree polarizations. Lecture Notes in Mathematics, 1958:135–230, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-70519-2_6.pdf.

Olsson:2008:MBT

[Ols08h] Martin C. Olsson. Moduli of broken toric varieties. Lecture Notes in Mathematics, 1958:57–83, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-70519-2_4.pdf.

Olsson:2008:MPP

[Ols08i] Martin C. Olsson. Moduli of principally polarized Abelian varieties. Lecture Notes in Mathematics, 1958:85–134, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-70519-2_5.pdf.

Olsson:2008:P

[Ols08j] Martin C. Olsson. Preliminaries. Lecture Notes in Mathematics, 1958:31–55, 2008. CODEN LNMAA2. ISBN 3-540-70518-X (print), 3-540-70519-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-70519-2_3.pdf.

Osipenko:2007:A

[Osi07a] Prof. George Osipenko. Attractors. Lecture Notes in Mathematics, 1889:65–83, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35595-3_7.pdf.

Osipenko:2007:BM

[Osi07b] Prof. George Osipenko. Back matter. Lecture Notes in Mathematics, 1889:233–287, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-35595-3/1.pdf.

Osipenko:2007:CRS

[Osi07c] Prof. George Osipenko. Chain recurrent set. Lecture Notes in Mathematics, 1889:55–63, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_6.pdf.

Osipenko:2007:C

[Osi07d] Prof. George Osipenko. Controllability. Lecture Notes in Mathematics, 1889:175–179, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_14.pdf.

Osipenko:2007:DSM

[Osi07e] Prof. George Osipenko. A dynamical system of mathematical biology. Lecture Notes in Mathematics, 1889:219–232, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35595-3_17.pdf.

Osipenko:2007:DSG

[Osi07f] Prof. George Osipenko. Dynamical Systems, Graphs, and Algorithms, volume 1889 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 283 pp. LCCN QA3 .L28 no. 1889. URL http://

link.springer.com/book/10.1007/3-540-35593-6;   http:
//www.springerlink.com/content/978-3-540-35595-3.

Osipenko:2007:E

[Osi07g] Prof. George Osipenko. Entropy. Lecture Notes in Mathematics, 1889:107–121, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35595-3_10.pdf.

Osipenko:2007:F

[Osi07h] Prof. George Osipenko. Filtration. Lecture Notes in Mathematics, 1889:85–95, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35595-3_8.pdf.

Osipenko:2007:FM

[Osi07i] Prof. George Osipenko. Front matter. Lecture Notes in Mathematics, 1889:I–XII, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-35595-3/1.pdf.

Osipenko:2007:HSS

[Osi07j] Prof. George Osipenko. Hyperbolicity and structural stability. Lecture Notes in Mathematics, 1889:161–174, 2007. CO-DEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35595-3_13.pdf.

Osipenko:2007:IMD

[Osi07k] Prof. George Osipenko. Ikeda mapping dynamics. Lecture Notes in Mathematics, 1889:197–218, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_16.pdf.

Osipenko:2007:I

[Osi07l] Prof. George Osipenko. Introduction. Lecture Notes in Mathematics, 1889:1–14, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35595-3_1.pdf.

Osipenko:2007:IM

[Osi07m] Prof. George Osipenko. Invariant manifolds. Lecture Notes in Mathematics, 1889:181–195, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_15.pdf.

Osipenko:2007:IS

[Osi07n] Prof. George Osipenko. Invariant sets. Lecture Notes in Mathematics, 1889:43–54, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35595-3_5.pdf.

Osipenko:2007:MS

[Osi07o] Prof. George Osipenko. Morse spectrum. Lecture Notes in Mathematics, 1889:137–160, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_12.pdf.

Osipenko:2007:NM

[Osi07p] Prof. George Osipenko. Newton’s method. Lecture Notes in Mathematics, 1889:35–41, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_4.pdf.

Osipenko:2007:PT

[Osi07q] Prof. George Osipenko. Periodic trajectories. Lecture Notes in Mathematics, 1889:27–33, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_3.pdf.

Osipenko:2007:PSL

[Osi07r] Prof. George Osipenko. Projective space and Lyapunov exponents. Lecture Notes in Mathematics, 1889:123–136, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35595-3_11.pdf.

Osipenko:2007:SG

[Osi07s] Prof. George Osipenko. Structural graph. Lecture Notes in Mathematics, 1889:97–105, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35595-3_9.pdf.

Osipenko:2007:SI

[Osi07t] Prof. George Osipenko. Symbolic image. Lecture Notes in Mathematics, 1889:15–25, 2007. CODEN LNMAA2. ISBN 3-540-35593-6 (print), 3-540-35595-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35595-3_2.pdf.

Pajot:2002:SGM

[Paj02a] Hervé Pajot. 1. Some geometric measure theory. Lecture Notes in Mathematics, 1799:1–15, 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36074-2_1/.

Pajot:2002:PJT

[Paj02b] Hervé Pajot. 2. P. Jones’ traveling salesman theorem. Lecture Notes in Mathematics, 1799:17–27, 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36074-2_2/.

Pajot:2002:MC

[Paj02c] Hervé Pajot. 3. Menger curvature. Lecture Notes in Mathematics, 1799:29–54, 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36074-2_3/.

Pajot:2002:CSI

[Paj02d] Hervé Pajot. 4. The Cauchy singular integral operator on Ahlfors regular sets. Lecture Notes in Mathematics, 1799:55– 65, 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36074-2_4/.

Pajot:2002:ACP

[Paj02e] Hervé Pajot. 5. Analytic capacity and the Painlevé problem. Lecture Notes in Mathematics, 1799:67–79, 2002. CODEN LNMAA2.

ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36074-2_5/.

Pajot:2002:DVC

[Paj02f] Hervé Pajot. 6. The Denjoy and Vitushkin conjectures. Lecture Notes in Mathematics, 1799:81–103, 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36074-2_6/.

Pajot:2002:CPP

[Paj02g] Hervé Pajot. 7. The capacity γ_{+} and the Painlevé problem. Lecture Notes in Mathematics, 1799:105–114, 2002. CODEN LN-MAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36074-2_
7/.

Pajot:2002:ACR

[Paj02h] Hervé Pajot. Analytic Capacity, Rectifiability, Menger Curvature and the Cauchy Integral, volume 1799 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 119 pp. LCCN QA3 .L28 no. 1799.

URL http://link.springer.com/book/10.1007/b84244;
http://www.springerlink.com/content/978-3-540-36074-2.

Pajot:2002:B

[Paj02i] Hervé Pajot. Bibliography. Lecture Notes in Mathematics, 1799: 115–118, 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36074-2_8/.

Pajot:2002:I

[Paj02j] Hervé Pajot. Index. Lecture Notes in Mathematics, 1799: 119, 2002. CODEN LNMAA2. ISBN 3-540-00001-1 (print), 3-540-36074-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36074-2_9.

Paouris:2000:ICN

[Pao00] G. Paouris. On the isotropic constant of non-symmetric convex bodies. Lecture Notes in Mathematics, 1745:239–243, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0107217/.

Paouris:2003:ELF

[Pao03] G. Paouris. Ψ_{2}-estimates for linear functionals on zonoids. Lecture Notes in Mathematics, 1807:211–222, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36428-3_17/.

Pastur:2004:TLD

[Pas04] L. Pastur. On the thermodynamic limit for disordered spin systems. Lecture Notes in Mathematics, 1850:243–268, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44489-3_19/.

Pate:2005:MAG

[Pat05] E. Pate. Mathematical analysis of the generation of force and motion in contracting muscle. Lecture Notes in Mathematics, 1867:133–153, 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/11406501_5.pdf.

Pautrat:2006:KIR

[Pau06] Yan Pautrat. Kernel and integral representations of operators on infinite dimensional toy Fock spaces. Lecture Notes in Mathematics, 1874:47–60, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35513-7_5.pdf.

Peng:2004:NEN

[Pen04] Shige Peng. Nonlinear expectations, nonlinear evaluations and risk measures. Lecture Notes in Mathematics, 1856:165–253, 2004. CODEN LNMAA2. ISBN 3-540-22953-1 (print), 3-540-44644-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44644-6_4/.

Peskir:2007:CVF

[Pes07] Goran Peskir. A change-of-variable formula with local time on surfaces. Lecture Notes in Mathematics, 1899:70–96, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_2.pdf.

Petkovic:2008:BM

[Pet08a] Miodrag Petković. Back matter. Lecture Notes in Mathematics, 1933:197–210, 2008. CODEN LNMAA2. ISBN 3-540-77850-0 (print), 3-540-77851-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-77851-6/1.pdf.

Petkovic:2008:BC

[Pet08b] Miodrag Petković. Basic concepts. Lecture Notes in Mathematics, 1933:1–34, 2008. CODEN LNMAA2. ISBN 3-540-77850-0 (print), 3-540-77851-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77851-6_1.pdf.

Petkovic:2008:FSMa

[Pet08c] Miodrag Petković. Families of simultaneous methods of higher order: Part I. Lecture Notes in Mathematics, 1933:129–160, 2008. CODEN LNMAA2. ISBN 3-540-77850-0 (print), 3-540-77851-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77851-6_4.pdf.

Petkovic:2008:FSMb

[Pet08d] Miodrag Petković. Families of simultaneous methods of higher order: Part II. Lecture Notes in Mathematics, 1933:161–195, 2008. CODEN LNMAA2. ISBN 3-540-77850-0 (print), 3-540-77851-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77851-6_5.pdf.

Petkovic:2008:FM

[Pet08e] Miodrag Petković. Front matter. Lecture Notes in Mathematics, 1933:i–xii, 2008. CODEN LNMAA2. ISBN 3-540-77850-0 (print), 3-540-77851-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-77851-6/1.pdf.

Petkovic:2008:IPP

[Pet08f] Miodrag Petković. Iterative processes and point estimation theory. Lecture Notes in Mathematics, 1933:35–66, 2008. CO-DEN LNMAA2. ISBN 3-540-77850-0 (print), 3-540-77851-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77851-6_2.pdf.

Petkovic:2008:PER

[Pet08g] Miodrag Petković. Point Estimation of Root Finding Methods, volume 1933 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-77851-9. ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 210 pp. LCCN QA3 .L28 no.

1933; QA276.8 .P474 2008. URL http://link.springer.com/
book/10.1007/978-3-540-77851-6;   http://www.loc.gov/
catdir/enhancements/fy0906/2008921391-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0906/2008921391-t.
html.

Petkovic:2008:PES

[Pet08h] Miodrag Petković. Point estimation of simultaneous methods. Lecture Notes in Mathematics, 1933:67–127, 2008. CO-DEN LNMAA2. ISBN 3-540-77850-0 (print), 3-540-77851-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77851-6_3.pdf.

Pflaum:2001:AGS

[Pfl01a] Markus J. Pflaum. Analytic and Geometric Study of Stratified Spaces, volume 1768 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 230 pp. LCCN QA3 .L28 no. 1768. URL http://

link.springer.com/book/10.1007/3-540-45436-5;   http:
//www.springerlink.com/content/978-3-540-45436-6.

Pflaum:2001:CT

[Pfl01b] Markus J. Pflaum. Control theory. Lecture Notes in Mathematics, 1768:91–149, 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45436-5_5/.

Pflaum:2001:RC

[Pfl01c] Markus J. Pflaum. De Rham cohomology. Lecture Notes in Mathematics, 1768:169–181, 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-45436-5_7/.

Pflaum:2001:DGO

[Pfl01d] Markus J. Pflaum. Differential geometric objects on singular spaces. Lecture Notes in Mathematics, 1768:63–90, 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45436-5_4/.

Pflaum:2001:HAS

[Pfl01e] Markus J. Pflaum. Homology of algebras of smooth functions. Lecture Notes in Mathematics, 1768:183–199, 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45436-5_8/.

Pflaum:2001:I

[Pfl01f] Markus J. Pflaum. Introduction. Lecture Notes in Mathematics, 1768:1–9, 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45436-5_1/.

Pflaum:2001:N

[Pfl01g] Markus J. Pflaum. Notation. Lecture Notes in Mathematics, 1768: 11–14, 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45436-5_2/.

Pflaum:2001:OS

[Pfl01h] Markus J. Pflaum. Orbit spaces. Lecture Notes in Mathematics, 1768:151–168, 2001. CODEN LNMAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45436-5_6/.

Pflaum:2001:SSF

[Pfl01i] Markus J. Pflaum. Stratified spaces and functional structures. Lecture Notes in Mathematics, 1768:15–62, 2001. CODEN LN-MAA2. ISBN 3-540-42626-4 (print), 3-540-45436-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45436-5_3/.

Phan:2001:MAA

[Pha01] Anthony Phan. Martingales d’azéma asymétriques. description Élémentaire et unicité. (French) []. Lecture Notes in Mathematics, 1755:48–86, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44671-2_3/.

Pham:2007:BM

[Pha07a] Huyên Pham. Back matter. Lecture Notes in Mathematics, 1919: 245–249, 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-73327-0/1.pdf.

Pham:2007:FM

[Pha07b] Huyen Pham. Front matter. Lecture Notes in Mathematics, 1899: 201, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-71189-6/3/1.pdf.

Pham:2007:SFP

[Pha07c] Huyen Pham. On the smooth-fit property for one-dimensional optimal switching problem. Lecture Notes in Mathematics, 1899:

187–199, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-71189-6_8.pdf.

Pham:2007:SAM

[Pha07d] Huyên Pham. Some applications and methods of large deviations in finance and insurance. Lecture Notes in Mathematics, 1919:191–244, 2007. CODEN LNMAA2. ISBN 3-540-73326-4 (print), 3-540-73327-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-73327-0_5.pdf.

Phelps:2001:AT

[Phe01a] Robert R. Phelps. Additional topics. Lecture Notes in Mathematics, 1757:101–113, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-48719-0_16/.

Phelps:2001:AKM

[Phe01b] Robert R. Phelps. Application of the Krein–Milman theorem to completely monotonic functions. Lecture Notes in Mathematics, 1757:9–12, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-48719-0_2/.

Phelps:2001:AIE

[Phe01c] Robert R. Phelps. Application to invariant and ergodic measures. Lecture Notes in Mathematics, 1757:73–78, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_12/.

Phelps:2001:ACB

[Phe01d] Robert R. Phelps. Applications of the Choquet boundary to resolvents. Lecture Notes in Mathematics, 1757:35–38, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_7/.

Phelps:2001:ARH

[Phe01e] Robert R. Phelps. Applications to Rainwater’s and Haydon’s theorems. Lecture Notes in Mathematics, 1757:25–26, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_5.

Phelps:2001:CBL

[Phe01f] Robert R. Phelps. The Choquet–Bishop–de Leeuw existence theorem. Lecture Notes in Mathematics, 1757:17–23, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_4/.

Phelps:2001:CBA

[Phe01g] Robert R. Phelps. The Choquet boundary and approximation theory. Lecture Notes in Mathematics, 1757:47–49, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_9/.

Phelps:2001:CBU

[Phe01h] Robert R. Phelps. The Choquet boundary for uniform algebras. Lecture Notes in Mathematics, 1757:39–45, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_8/.

Phelps:2001:CTM

[Phe01i] Robert R. Phelps. Choquet’s theorem: The metrizable case. Lecture Notes in Mathematics, 1757:13–16, 2001. CODEN LN-MAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_3/.

Phelps:2001:DME

[Phe01j] Robert R. Phelps. A different method for extending the representation theorems. Lecture Notes in Mathematics, 1757:88– 91, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-48719-0_14/.

Phelps:2001:IKM

[Phe01k] Robert R. Phelps. Introduction. the Krein–Milman theorem as an integral representation theorem. Lecture Notes in Mathematics, 1757:1–8, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-48719-0_1/.

Phelps:2001:LCT

[Phe01l] Robert R. Phelps. Lectures on Choquet’s Theorem, volume 1757 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 124 pp. LCCN QA3 .L28 no. 1757. URL http:/

/link.springer.com/book/10.1007/b76887;   http://www.
springerlink.com/content/978-3-540-48719-7.

Phelps:2001:MER

[Phe01m] Robert R. Phelps. A method for extending the representation theorems: Caps. Lecture Notes in Mathematics, 1757:79– 87, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-48719-0_13/.

Phelps:2001:NSC

[Phe01n] Robert R. Phelps. A new setting: The Choquet boundary. Lecture Notes in Mathematics, 1757:27–33, 2001. CODEN LN-MAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_6/.

Phelps:2001:ODM

[Phe01o] Robert R. Phelps. Orderings and dilations of measures. Lecture Notes in Mathematics, 1757:93–99, 2001. CODEN LN-MAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_15/.

Phelps:2001:PRM

[Phe01p] Robert R. Phelps. Properties of the resultant map. Lecture Notes in Mathematics, 1757:65–71, 2001. CODEN LNMAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-48719-0_11/.

Phelps:2001:URM

[Phe01q] Robert R. Phelps. Uniqueness of representing measures. Lecture Notes in Mathematics, 1757:51–63, 2001. CODEN LN-MAA2. ISBN 3-540-41834-2 (print), 3-540-48719-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-48719-0_10/.

Picard:2005:LPT

[Pic05] Jean Picard, editor. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXXIII — 2003, volume 1869 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-26069-2 (print), 3-540-31537-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1869. URL http://

link.springer.com/book/10.1007/b136622;   http://www.
springerlink.com/content/978-3-540-31537-7.

Pilgrim:2003:I

[Pil03a] Kevin M. Pilgrim. 1. Introduction. Lecture Notes in Mathematics, 1827:1–35, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39936-0_1/.

Pilgrim:2003:CDT

[Pil03b] Kevin M. Pilgrim. 10. Canonical decomposition theorem. Lecture Notes in Mathematics, 1827:105–109, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39936-0_10/.

Pilgrim:2003:P

[Pil03c] Kevin M. Pilgrim. 2. Preliminaries. Lecture Notes in Mathematics, 1827:37–48, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39936-0_2/.

Pilgrim:2003:C

[Pil03d] Kevin M. Pilgrim. 3. Combinations. Lecture Notes in Mathematics, 1827:49–57, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39936-0_3/.

Pilgrim:2003:UC

[Pil03e] Kevin M. Pilgrim. 4. Uniqueness of combinations. Lecture Notes in Mathematics, 1827:59–68, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39936-0_4/.

Pilgrim:2003:D

[Pil03f] Kevin M. Pilgrim. 5. Decomposition. Lecture Notes in Mathematics, 1827:69–77, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39936-0_5/.

Pilgrim:2003:UD

[Pil03g] Kevin M. Pilgrim. 6. Uniqueness of decompositions. Lecture Notes in Mathematics, 1827:79–81, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-39936-0_6/.

Pilgrim:2003:CCA

[Pil03h] Kevin M. Pilgrim. 7. Counting classes of annulus maps. Lecture Notes in Mathematics, 1827:83–88, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39936-0_7/.

Pilgrim:2003:AMC

[Pil03i] Kevin M. Pilgrim. 8. Applications to mapping class groups. Lecture Notes in Mathematics, 1827:89–94, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39936-0_8/.

Pilgrim:2003:E

[Pil03j] Kevin M. Pilgrim. 9. Examples. Lecture Notes in Mathematics, 1827:95–103, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39936-0_9/.

Pilgrim:2003:CCD

[Pil03k] Kevin M. Pilgrim. Combinations of Complex Dynamical Systems, volume 1827 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 118 pp. LCCN QA3 .L28 no. 1827. URL http:/

/link.springer.com/book/10.1007/b14147;   http://www.
springerlink.com/content/978-3-540-39936-0.

Pilgrim:2003:R

[Pil03l] Kevin M. Pilgrim. References. Lecture Notes in Mathematics, 1827:111–116, 2003. CODEN LNMAA2. ISBN 3-540-20173-4 (print), 3-540-39936-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39936-0_11/.

Pillet:2006:QDS

[Pil06] Claude-Alain Pillet. Quantum dynamical systems. Lecture Notes in Mathematics, 1880:107–182, 2006. CODEN LNMAA2. ISBN 3-540-30991-8 (print), 3-540-33922-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-33922-1_4.pdf.

Pironneau:2000:OSD

[Pir00] Olivier Pironneau. Optimal shape design by local boundary variations. Lecture Notes in Mathematics, 1740:343–384, 2000. CO-DEN LNMAA2. ISBN 3-540-67971-5 (print), 3-540-44486-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0106744/.

Pisier:2001:IDC

[Pis01a] Gilles Pisier. 0. Introduction. Description of contents. Lecture Notes in Mathematics, 1618:1–13, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44563-0_1/.

Pisier:2001:NIA

[Pis01b] Gilles Pisier. 1. Von Neumann’s inequality and Ando’s generalization. Lecture Notes in Mathematics, 1618:14–30, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44563-0_
2/.

Pisier:2001:KSP

[Pis01c] Gilles Pisier. 10. The Kadison similarity problem. Lecture Notes in Mathematics, 1618:168–181, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44563-0_11/.

Pisier:2001:NUU

[Pis01d] Gilles Pisier. 2. Non-unitarizable uniformly bounded group representations. Lecture Notes in Mathematics, 1618:31–57, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44563-0_3/.

Pisier:2001:CBMa

[Pis01e] Gilles Pisier. 3. Completely bounded maps. Lecture Notes in Mathematics, 1618:58–74, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44563-0_4/.

Pisier:2001:CBH

[Pis01f] Gilles Pisier. 4. Completely bounded homomorphisms and derivations. Lecture Notes in Mathematics, 1618:75–98, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44563-0_
5/.

Pisier:2001:SMG

[Pis01g] Gilles Pisier. 5. Schur multipliers and Grothendieck’s inequality. Lecture Notes in Mathematics, 1618:99–113, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44563-0_
6/.

Pisier:2001:HSM

[Pis01h] Gilles Pisier. 6. Hankelian Schur multipliers. Herz–Schur multipliers. Lecture Notes in Mathematics, 1618:114–123, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44563-0_
7/.

Pisier:2001:SPCa

[Pis01i] Gilles Pisier. 7. The similarity problem for cyclic homomorphisms on a C^{∗}-algebra. Lecture Notes in Mathematics, 1618: 124–141, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44563-0_8/.

Pisier:2001:CBMb

[Pis01j] Gilles Pisier. 8. Completely bounded maps in the Banach space setting. Lecture Notes in Mathematics, 1618:142–151, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44563-0_9/.

Pisier:2001:SNH

[Pis01k] Gilles Pisier. 9. The Sz.-Nagy–Halmos similarity problem. Lecture Notes in Mathematics, 1618:152–167, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44563-0_10/.

Pisier:2001:R

[Pis01l] Gilles Pisier. References. Lecture Notes in Mathematics, 1618: 182–193, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44563-0_12/.

Pisier:2001:SPCb

[Pis01m] Gilles Pisier. Similarity Problems and Completely Bounded Maps, volume 1618 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second expanded edition, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 198 pp. LCCN QA3 .L28 no. 1618.

URL http://link.springer.com/book/10.1007/b55674;
http://www.springerlink.com/content/978-3-540-44563-0.

Pisier:2001:SNI

[Pis01n] Gilles Pisier. Subject and notation index. Lecture Notes in Mathematics, 1618:194–196, 2001. CODEN LNMAA2. ISBN 3-540-41524-6 (print), 3-540-44563-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44563-0_13/.

Pisier:2004:RPA

[Pis04] G. Pisier. On Read’s proof that B(ℓ_{1}) is not amenable. Lecture Notes in Mathematics, 1850:269–275, 2004. CODEN LNMAA2. ISBN 3-540-22360-6 (print), 3-540-44489-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44489-3_20/.

Pistorius:2005:PTR

[Pis05] Martijn R. Pistorius. A potential-theoretical review of some exit problems of spectrally negative Lévy processes. Lecture Notes in Mathematics, 1857:30–41, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31449-3_4.pdf.

Pistorius:2007:ETA

[Pis07] Martijn R. Pistorius. An excursion-theoretical approach to some boundary crossing problems and the Skorokhod embedding for reflected Lévy processes. Lecture Notes in Mathematics, 1899:287– 307, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_15.pdf.

Pitman:2006:BPC

[Pit06a] Jim Pitman. Bell polynomials, composite structures and Gibbs partitions. Lecture Notes in Mathematics, 1875:13–35, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-34266-4_2.pdf.

Pitman:2006:BBA

[Pit06b] Jim Pitman. Brownian bridge asymptotics for random mappings. Lecture Notes in Mathematics, 1875:193–206, 2006. CODEN LN-MAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/3-540-34266-4_
10.pdf.

Pitman:2006:BF

[Pit06c] Jim Pitman. The Brownian forest. Lecture Notes in Mathematics, 1875:143–175, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-34266-4_8.pdf.

Pitman:2006:BLT

[Pit06d] Jim Pitman. Brownian local times, branching and Bessel processes. Lecture Notes in Mathematics, 1875:177–191, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-34266-4_9.pdf.

Pitman:2006:CFP

[Pit06e] Jim Pitman. Coagulation and fragmentation processes. Lecture Notes in Mathematics, 1875:97–120, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-34266-4_6.pdf.

Pitman:2006:ERP

[Pit06f] Jim Pitman. Exchangeable random partitions. Lecture Notes in Mathematics, 1875:37–53, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-34266-4_3.pdf.

Pitman:2006:PCR

[Pit06g] Jim Pitman. Poisson constructions of random partitions. Lecture Notes in Mathematics, 1875:77–95, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-34266-4_5.pdf.

Pitman:2006:P

[Pit06h] Jim Pitman. Preliminaries. Lecture Notes in Mathematics, 1875: 1–11, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/3-540-34266-4_1.pdf.

Pitman:2006:RFA

[Pit06i] Jim Pitman. Random forests and the additive coalescent. Lecture Notes in Mathematics, 1875:207–221, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-34266-4_11.pdf.

Pitman:2006:RWR

[Pit06j] Jim Pitman. Random walks and random forests. Lecture Notes in Mathematics, 1875:121–141, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/3-540-34266-4_7.pdf.

Pitman:2006:SCR

[Pit06k] Jim Pitman. Sequential constructions of random partitions. Lecture Notes in Mathematics, 1875:55–75, 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/3-540-34266-4_4.pdf.

Pivovarov:2007:RCB

[Piv07] P. Pivovarov. Random convex bodies lacking symmetric projections, revisited through decoupling. Lecture Notes in Mathematics, 1910:255–263, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72053-9_14.pdf.

Pastur:2003:MAS

[PL03] Leonid Pastur and Antonie Lejay. Matrices aléatoires: Statistique asymptotique des valeurs propres. (French) []. Lecture Notes in Mathematics, 1801:135–164, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36107-7_2/.

Pitman:2006:CSP

[PP06] Jim Pitman and Jean Picard, editors. Combinatorial Stochastic Processes: École d’Été de Probabilités de Saint-Flour XXXII — 2002, volume 1875 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-30990-X (print), 3-540-34266-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1875; QA274.2 .P55 2006; QA3 .L28 no.1875. URL http:

//link.springer.com/book/10.1007/b11601500;   http://
www.loc.gov/catdir/enhancements/fy0663/2006921042-d.html;
 http://www.loc.gov/catdir/enhancements/fy0824/2006921042-
b.html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006921042-t.html;   http://www.springerlink.com/content/
978-3-540-34266-3.

Perez:2002:PEM

[PR02] Joaquín Pérez and Antonio Ros. Properly embedded minimal surfaces with finite total curvature. Lecture Notes in Mathematics, 1775:15–66, 2002. CODEN LNMAA2. ISBN 3-540-43120-9 (print), 3-540-45609-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45609-4_2/.

Prevot:2007:BM

[PR07a] Claudia Prévôt and Michael Röckner. Back matter. Lecture Notes in Mathematics, 1905:105–148, 2007. CODEN LNMAA2. ISBN 3-540-70780-8 (print), 3-540-70781-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-70781-3/1.pdf.

Prevot:2007:CSD

[PR07b] Claudia Prévôt and Michael Röckner. A class of stochastic differential equations. Lecture Notes in Mathematics, 1905:55–103, 2007. CODEN LNMAA2. ISBN 3-540-70780-8 (print), 3-540-70781-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-70781-3_4.pdf.

Prevot:2007:CCS

[PR07c] Claudia Prévôt and Michael Röckner. A Concise Course on Stochastic Partial Differential Equations, volume 1905 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-70780-8 (print), 3-540-70781-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 144 pp. LCCN

QA3 .L28 no. 1905. URL http://link.springer.com/book/
10.1007/978-3-540-70781-3;   http://www.springerlink.
com/content/978-3-540-70781-3.

Prevot:2007:FM

[PR07d] Claudia Prévôt and Michael Röckner. Front matter. Lecture Notes in Mathematics, 1905:V–VI, 2007. CODEN LNMAA2. ISBN 3-540-70780-8 (print), 3-540-70781-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-70781-3/1.pdf.

Prevot:2007:MAE

[PR07e] Claudia Prévôt and Michael Röckner. Motivation, aims and examples. Lecture Notes in Mathematics, 1905:1–4, 2007. CO-DEN LNMAA2. ISBN 3-540-70780-8 (print), 3-540-70781-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-70781-3_1.pdf.

Prevot:2007:SDE

[PR07f] Claudia Prévôt and Michael Röckner. Stochastic differential equations in finite dimensions. Lecture Notes in Mathematics, 1905:43– 54, 2007. CODEN LNMAA2. ISBN 3-540-70780-8 (print), 3-540-70781-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-70781-3_3.pdf.

Prevot:2007:SIH

[PR07g] Claudia Prévôt and Michael Röckner. Stochastic integral in Hilbert spaces. Lecture Notes in Mathematics, 1905:5–42, 2007. CODEN LNMAA2. ISBN 3-540-70780-8 (print), 3-540-70781-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-70781-3_2.pdf.

Privault:2001:QSC

[Pri01] Nicolas Privault. Quantum stochastic calculus for the uniform measure and Boolean convolution. Lecture Notes in Mathematics, 1755:28–47, 2001. CODEN LNMAA2. ISBN 3-540-41659-5 (print), 3-540-44671-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44671-2_2/.

Privault:2008:PTC

[Pri08a] Nicolas Privault. Potential theory in classical probability. Lecture Notes in Mathematics, 1954:3–59, 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-69365-9_2.
pdf.

Privault:2008:PIS

[Pri08b] Nicolas Privault. A probabilistic interpretation to the symmetries of a discrete heat equation. Lecture Notes in Mathematics, 1934:379–399, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77913-1_18.pdf.

Privault:2009:APS

[Pri09a] Nicolas Privault. Analysis on the Poisson space. Lecture Notes in Mathematics, 1982:195–246, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-02380-4_7.pdf.

Privault:2009:AWS

[Pri09b] Nicolas Privault. Analysis on the Wiener space. Lecture Notes in Mathematics, 1982:161–194, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-02380-4_6.pdf.

Privault:2009:ACO

[Pri09c] Nicolas Privault. Annihilation and creation operators. Lecture Notes in Mathematics, 1982:131–160, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-02380-4_5.
pdf.

Privault:2009:A

[Pri09d] Nicolas Privault. Appendix. Lecture Notes in Mathematics, 1982:295–300, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-02380-4_10.pdf.

Privault:2009:BM

[Pri09e] Nicolas Privault. Back matter. Lecture Notes in Mathematics, 1982:1–15, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-02380-4/1.pdf.

Privault:2009:CTN

[Pri09f] Nicolas Privault. Continuous time normal martingales. Lecture Notes in Mathematics, 1982:59–112, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-02380-4_3.
pdf.

Privault:2009:DTC

[Pri09g] Nicolas Privault. The discrete time case. Lecture Notes in Mathematics, 1982:7–58, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-02380-4_2.pdf.

Privault:2009:FM

[Pri09h] Nicolas Privault. Front matter. Lecture Notes in Mathematics, 1982:1–7, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-02380-4/1.pdf.

Privault:2009:GDO

[Pri09i] Nicolas Privault. Gradient and divergence operators. Lecture Notes in Mathematics, 1982:113–130, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-02380-4_4.
pdf.

Privault:2009:I

[Pri09j] Nicolas Privault. Introduction. Lecture Notes in Mathematics, 1982:1–6, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-02380-4_1.pdf.

Privault:2009:LGP

[Pri09k] Nicolas Privault. Local gradients on the Poisson space. Lecture Notes in Mathematics, 1982:247–280, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-02380-4_8.
pdf.

Privault:2009:OHC

[Pri09l] Nicolas Privault. Option hedging in continuous time. Lecture Notes in Mathematics, 1982:281–293, 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-02380-4_9.
pdf.

Privault:2009:SAD

[Pri09m] Nicolas Privault. Stochastic Analysis in Discrete and Continuous Settings: With Normal Martingales, volume 1982 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-02379-7 (print), 3-642-02380-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 310 pp. LCCN QA3 .L28 no.

1982. URL http://link.springer.com/book/10.1007/978-
3-642-02380-4;   http://www.springerlink.com/content/
978-3-642-02380-4.

Peccati:2005:GLV

[PT05] Giovanni Peccati and Ciprian A. Tudor. Gaussian limits for vector-valued multiple stochastic integrals. Lecture Notes in Mathematics, 1857:247–262, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-31449-3_17.pdf.

Perelli:2006:ANT

[PV06] Alberto Perelli and Carlo Viola, editors. Analytic Number Theory: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, July 11–18, 2002, volume 1891 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-36363-7 (print), 3-540-36364-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1891. URL http://

link.springer.com/book/10.1007/3-540-36363-7;   http:
//www.springerlink.com/content/978-3-540-36364-4.

Qian:2009:FM

[QXZ09a] Min Qian, Jian-Sheng Xie, and Shu Zhu. Front matter. Lecture Notes in Mathematics, 1978:1–11, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-01954-8/1.pdf.

Qian:2009:SET

[QXZ09b] Min Qian, Jian-Sheng Xie, and Shu Zhu. Smooth Ergodic Theory for Endomorphisms, volume 1978 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 277 pp. LCCN QA3 .L28 no.

1978. URL http://link.springer.com/book/10.1007/978-
3-642-01954-8;   http://www.springerlink.com/content/
978-3-642-01954-8.

Quian:2009:AE

[QXZ09c] Min Quian, Jian-Sheng Xie, and Shu Zhu. Axiom a endomorphisms. Lecture Notes in Mathematics, 1978:27–44, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01954-8_4.pdf.

Quian:2009:BM

[QXZ09d] Min Quian, Jian-Sheng Xie, and Shu Zhu. Back matter. Lecture Notes in Mathematics, 1978:1–38, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-01954-8/1.pdf.

Quian:2009:EPL

[QXZ09e] Min Quian, Jian-Sheng Xie, and Shu Zhu. Ergodic property of Lyapunov exponents. Lecture Notes in Mathematics, 1978:151– 171, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01954-8_8.pdf.

Quian:2009:EDH

[QXZ09f] Min Quian, Jian-Sheng Xie, and Shu Zhu. Exact dimensionality of hyperbolic measures. Lecture Notes in Mathematics, 1978:205– 244, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01954-8_10.pdf.

Quian:2009:EM

[QXZ09g] Min Quian, Jian-Sheng Xie, and Shu Zhu. Expanding maps. Lecture Notes in Mathematics, 1978:15–26, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-01954-8_3.
pdf.

Quian:2009:GEF

[QXZ09h] Min Quian, Jian-Sheng Xie, and Shu Zhu. Generalized entropy formula. Lecture Notes in Mathematics, 1978:173–204, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01954-8_9.pdf.

Quian:2009:MRI

[QXZ09i] Min Quian, Jian-Sheng Xie, and Shu Zhu. Margulis–Ruelle inequality. Lecture Notes in Mathematics, 1978:9–13, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01954-8_2.pdf.

Quian:2009:PEF

[QXZ09j] Min Quian, Jian-Sheng Xie, and Shu Zhu. Pesin’s entropy formula for endomorphisms. Lecture Notes in Mathematics, 1978:87–96, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01954-8_6.pdf.

Quian:2009:P

[QXZ09k] Min Quian, Jian-Sheng Xie, and Shu Zhu. Preliminaries. Lecture Notes in Mathematics, 1978:1–8, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01954-8_1.pdf.

Quian:2009:SMP

[QXZ09l] Min Quian, Jian-Sheng Xie, and Shu Zhu. SRB measures and Pesin’s entropy formula for endomorphisms. Lecture Notes in Mathematics, 1978:97–150, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01954-8_7.pdf.

Quian:2009:USM

[QXZ09m] Min Quian, Jian-Sheng Xie, and Shu Zhu. Unstable and stable manifolds for endomorphisms. Lecture Notes in Mathematics, 1978:45–86, 2009. CODEN LNMAA2. ISBN 3-642-01953-6 (print), 3-642-01954-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01954-8_5.pdf.

Rabetti:2001:DCH

[Rab01] Alessandra Rabetti. Dialgebra (co)homology with coefficients. Lecture Notes in Mathematics, 1763:67–103, 2001. CODEN LN-MAA2. ISBN 3-540-42194-7 (print), 3-540-45328-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45328-8_3/.

Rajeev:2006:PTB

[Raj06] Bhaskaran Rajeev. A path transformation of Brownian motion. Lecture Notes in Mathematics, 1874:259–267, 2006. CO-DEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_18.pdf.

Rasonyi:2003:RST

[Rás03] Miklós Rásonyi. A remark on the superhedging theorem under transaction costs. Lecture Notes in Mathematics, 1832:394– 398, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_17/.

Rasmussen:2007:ABN

[Ras07a] Martin Rasmussen. Attractivity and Bifurcation for Nonautonomous Dynamical Systems, volume 1907 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 212 pp. LCCN QA3 .L28 no.

1907. URL http://link.springer.com/book/10.1007/978-
3-540-71225-1;   http://www.springerlink.com/content/
978-3-540-71225-1.

Rasmussen:2007:BM

[Ras07b] Martin Rasmussen. Back matter. Lecture Notes in Mathematics, 1907:193–215, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-71225-1/1.pdf.

Rasmussen:2007:BDO

[Ras07c] Martin Rasmussen. Bifurcations in dimension one. Lecture Notes in Mathematics, 1907:137–152, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71225-1_6.pdf.

Rasmussen:2007:BAA

[Ras07d] Martin Rasmussen. Bifurcations of asymptotically autonomous systems. Lecture Notes in Mathematics, 1907:153–191, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71225-1_7.pdf.

Rasmussen:2007:FM

[Ras07e] Martin Rasmussen. Front matter. Lecture Notes in Mathematics, 1907:IX–XI, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-71225-1/1.pdf.

Rasmussen:2007:I

[Ras07f] Martin Rasmussen. Introduction. Lecture Notes in Mathematics, 1907:1–6, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71225-1_1.pdf.

Rasmussen:2007:LS

[Ras07g] Martin Rasmussen. Linear systems. Lecture Notes in Mathematics, 1907:81–113, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71225-1_4.pdf.

Rasmussen:2007:NMD

[Ras07h] Martin Rasmussen. Nonautonomous Morse decompositions. Lecture Notes in Mathematics, 1907:51–80, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71225-1_3.
pdf.

Rasmussen:2007:NS

[Ras07i] Martin Rasmussen. Nonlinear systems. Lecture Notes in Mathematics, 1907:115–135, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71225-1_5.pdf.

Rasmussen:2007:NAB

[Ras07j] Martin Rasmussen. Notions of attractivity and bifurcation. Lecture Notes in Mathematics, 1907:7–50, 2007. CODEN LNMAA2. ISBN 3-540-71224-0 (print), 3-540-71225-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71225-1_2.
pdf.

Rasonyi:2009:BM

[Rás09a] Mikós Rásonyi. Back matter. Lecture Notes in Mathematics, 1979:1–6, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-01763-6/1.pdf.

Rasonyi:2009:ENM

[Rás09b] Mikós Rásonyi. Erratum to: New methods in the arbitrage theory of financial markets with transaction costs, in Séminaire XLI. Lecture Notes in Mathematics, 1979:449, 2009. CO-DEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_17.pdf.

Rey-Bellet:2006:OCS

[RB06] Luc Rey-Bellet. Open classical systems. Lecture Notes in Mathematics, 1881:41–78, 2006. CODEN LNMAA2. ISBN 3-540-30992-6 (print), 3-540-33966-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-33966-3_2.pdf.

Rebolledo:2006:CPM

[Reb06] Rolando Rebolledo. Complete positivity and the Markov structure of open quantum systems. Lecture Notes in Mathematics, 1881:149–182, 2006. CODEN LNMAA2. ISBN 3-540-30992-6 (print), 3-540-33966-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/3-540-33966-3_4.pdf.

Reichel:2004:I

[Rei04a] Wolfgang Reichel. 1. Introduction. Lecture Notes in Mathematics, 1841:1–7, 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40915-1_1/.

Reichel:2004:UCPa

[Rei04b] Wolfgang Reichel. 2. Uniqueness of critical points (i). Lecture Notes in Mathematics, 1841:9–26, 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40915-1_2/.

Reichel:2004:UCPb

[Rei04c] Wolfgang Reichel. 3. Uniqueness of critical points (II). Lecture Notes in Mathematics, 1841:27–57, 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40915-1_3/.

Reichel:2004:VPR

[Rei04d] Wolfgang Reichel. 4. Variational problems on Riemannian manifolds. Lecture Notes in Mathematics, 1841:59–87, 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40915-1_
4/.

Reichel:2004:SPE

[Rei04e] Wolfgang Reichel. 5. Scalar problems in Euclidean space. Lecture Notes in Mathematics, 1841:89–125, 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40915-1_5/.

Reichel:2004:VPE

[Rei04f] Wolfgang Reichel. 6. Vector problems in Euclidean space. Lecture Notes in Mathematics, 1841:127–138, 2004. CODEN LNMAA2.

ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40915-1_6/.

Reichel:2004:A

[Rei04g] Wolfgang Reichel. Appendix. Lecture Notes in Mathematics, 1841:139–143, 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40915-1_7/.

Reichel:2004:R

[Rei04h] Wolfgang Reichel. References. Lecture Notes in Mathematics, 1841:145–149, 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40915-1_8/.

Reichel:2004:UTV

[Rei04i] Wolfgang Reichel. Uniqueness Theorems for Variational Problems by the Method of Transformation Groups, volume 1841 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-21839-4 (print), 3-540-40915-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 152 pp. LCCN QA3 .L28 no. 1841. URL http:/

/link.springer.com/book/10.1007/b96984;   http://www.
springerlink.com/content/978-3-540-40915-1.

Reisert:2005:STV

[Rei05] J. Reisert. Signal transduction in vertebrate olfactory receptor cells. Lecture Notes in Mathematics, 1867:155–171, 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11406501_6.pdf.

Rezakhanlou:2008:BM

[Rez08a] F. Rezakhanlou. Back matter. Lecture Notes in Mathematics, 1916:107–111, 2008. CODEN LNMAA2. ISBN 3-540-73704-9 (print), 3-540-73705-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-73705-6/1.pdf.

Rezakhanlou:2008:KLI

[Rez08b] F. Rezakhanlou. Kinetic limits for interacting particle systems. Lecture Notes in Mathematics, 1916:71–105, 2008. CO-DEN LNMAA2. ISBN 3-540-73704-9 (print), 3-540-73705-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73705-6_2.pdf.

Ricca:2009:SCD

[Ric09] Renzo L. Ricca. Structural complexity and dynamical systems. Lecture Notes in Mathematics, 1973:167–186, 2009. CO-DEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00837-5_6.pdf.

Ringhofer:2003:KGD

[Rin03] Christian Ringhofer. Kinetic and gas- dynamic models for semiconductor transport. Lecture Notes in Mathematics, 1823:97– 131, 2003. CODEN LNMAA2. ISBN 3-540-40802-9 (print), 3-540-45222-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45222-5_3/.

Ritter:2000:ACA

[Rit00a] Klaus Ritter. Average-Case Analysis of Numerical Problems, volume 1733 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 254 pp. LCCN QA3 .L28 no. 1733; QA3.L28 no.

1733 QA297. URL http://link.springer.com/book/10.1007/
BFb0103934;   http://www.loc.gov/catdir/enhancements/
fy0816/00030799-d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00030799-t.html;   http://www.springerlink.com/
content/978-3-540-45592-9.

Ritter:2000:IAM

[Rit00b] Klaus Ritter. Integration and approximation of multivariate functions. Lecture Notes in Mathematics, 1733:133–182, 2000. CO-DEN LNMAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103940/.

Ritter:2000:IAU

[Rit00c] Klaus Ritter. Integration and approximation of univariate functions. Lecture Notes in Mathematics, 1733:67–121, 2000. CODEN LNMAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103938/.

Ritter:2000:I

[Rit00d] Klaus Ritter. Introduction. Lecture Notes in Mathematics, 1733: 1–9, 2000. CODEN LNMAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0103935/.

Ritter:2000:LPD

[Rit00e] Klaus Ritter. Linear problems: Definitions and a classical example. Lecture Notes in Mathematics, 1733:11–31, 2000. CODEN LNMAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103936/.

Ritter:2000:LPU

[Rit00f] Klaus Ritter. Linear problems for univariate functions with noisy data. Lecture Notes in Mathematics, 1733:123–131, 2000. CO-DEN LNMAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103939/.

Ritter:2000:NML

[Rit00g] Klaus Ritter. Nonlinear methods for linear problems. Lecture Notes in Mathematics, 1733:183–211, 2000. CODEN LN-MAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103941/.

Ritter:2000:NP

[Rit00h] Klaus Ritter. Nonlinear problems. Lecture Notes in Mathematics, 1733:213–225, 2000. CODEN LNMAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103942/.

Ritter:2000:SOR

[Rit00i] Klaus Ritter. Second-order results for linear problems. Lecture Notes in Mathematics, 1733:33–65, 2000. CODEN LN-MAA2. ISBN 3-540-67449-7 (print), 3-540-45592-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0103937/.

Rivat:2002:UMD

[Riv02] J. Rivat. A. undecomposable matrices in dimension 3. Lecture Notes in Mathematics, 1794:375–376, 2002. CODEN LN-MAA2. ISBN 3-540-44141-7 (print), 3-540-45714-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45714-3_13.

Rosen:2000:VFL

[RK00] Jay Rosen and Haya Kaspi. p-Variation for families of local times on lines. Lecture Notes in Mathematics, 1729:171–184, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103802/.

Roberts:2004:BM

[Rob04a] John E. Roberts. Back matter. Lecture Notes in Mathematics, 1831:343–354, 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-39702-1/1.pdf.

Roberts:2004:MLA

[Rob04b] John E. Roberts. More lectures on algebraic quantum field theory. Lecture Notes in Mathematics, 1831:263–342, 2004. CODEN LNMAA2. ISBN 3-540-20357-5 (print), 3-540-39702-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39702-1_
5/.

Rogers:2003:DCO

[Rog03] L. C. G. Rogers. Duality in constrained optimal investment and consumption problems: a synthesis. Lecture Notes in Mathematics, 1814:95–131, 2003. CODEN LNMAA2. ISBN 3-540-40193-8 (print), 3-540-44859-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44859-4_3/.

Rohde:2009:CCC

[Roh09a] Christian Rohde. Cyclic Coverings, Calabi–Yau Manifolds and Complex Multiplication, volume 1975 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 228 pp. LCCN QA3 .L28 no.

1975. URL http://link.springer.com/book/10.1007/978-
3-642-00639-5;   http://www.springerlink.com/content/
978-3-642-00639-5.

Rohde:2009:FM

[Roh09b] Christian Rohde. Front matter. Lecture Notes in Mathematics, 1975:1–7, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-00639-5/1.pdf.

Rohde:2009:BM

[Roh09c] Jan Christian Rohde. Back matter. Lecture Notes in Mathematics, 1975:1–24, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-00639-5/1.pdf.

Rohde:2009:CHG

[Roh09d] Jan Christian Rohde. The computation of the Hodge group. Lecture Notes in Mathematics, 1975:91–119, 2009. CO-DEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00639-5_6.pdf.

Rohde:2009:CCY

[Roh09e] Jan Christian Rohde. The construction of Calabi–Yau manifolds with complex multiplication. Lecture Notes in Mathematics, 1975:143–156, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00639-5_8.pdf.

Rohde:2009:CCP

[Roh09f] Jan Christian Rohde. Cyclic covers of the projective line. Lecture Notes in Mathematics, 1975:59–69, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-00639-5_3.
pdf.

Rohde:2009:DC

[Roh09g] Jan Christian Rohde. The degree 3 case. Lecture Notes in Mathematics, 1975:157–167, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-00639-5_9.pdf.

Rohde:2009:ECF

[Roh09h] Jan Christian Rohde. Examples of CMCY families of 3-manifolds and their invariants. Lecture Notes in Mathematics, 1975:187– 198, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00639-5_11.pdf.

Rohde:2009:EFD

[Roh09i] Jan Christian Rohde. Examples of families with dense sets of complex multiplication fibers. Lecture Notes in Mathematics, 1975:121–142, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00639-5_7.pdf.

Rohde:2009:GGD

[Roh09j] Jan Christian Rohde. The Galois group decomposition of the Hodge structure. Lecture Notes in Mathematics, 1975:79–89, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00639-5_5.pdf.

Rohde:2009:I

[Roh09k] Jan Christian Rohde. Introduction. Lecture Notes in Mathematics, 1975:1–9, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-00639-5_1.pdf.

Rohde:2009:IHS

[Roh09l] Jan Christian Rohde. An introduction to Hodge structures and Shimura varieties. Lecture Notes in Mathematics, 1975:11–57, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00639-5_2.pdf.

Rohde:2009:MFC

[Roh09m] Jan Christian Rohde. Maximal families of CMCY type. Lecture Notes in Mathematics, 1975:199–208, 2009. CO-DEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00639-5_12.pdf.

Rohde:2009:OEV

[Roh09n] Jan Christian Rohde. Other examples and variations. Lecture Notes in Mathematics, 1975:169–186, 2009. CODEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-00639-5_10.
pdf.

Rohde:2009:SPF

[Roh09o] Jan Christian Rohde. Some preliminaries for families of cyclic covers. Lecture Notes in Mathematics, 1975:71–78, 2009. CO-DEN LNMAA2. ISBN 3-642-00638-8 (print), 3-642-00639-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00639-5_4.pdf.

Rosenberg:2002:BS

[Ros02] Harold Rosenberg. Bryant surfaces. Lecture Notes in Mathematics, 1775:67–111, 2002. CODEN LNMAA2. ISBN 3-540-43120-9 (print), 3-540-45609-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45609-4_3/.

Rosler:2003:DOT

[Rös03] Margit Rösler. Dunkl operators: Theory and applications. Lecture Notes in Mathematics, 1817:93–135, 2003. CODEN LN-MAA2. ISBN 3-540-40375-2 (print), 3-540-44945-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44945-0_3/.

Rosay:2004:IGT

[Ros04] Jean-Pierre Rosay. Introduction to a general theory of boundary values. Lecture Notes in Mathematics, 1848:165–189, 2004. CODEN LNMAA2. ISBN 3-540-22358-4 (print), 3-540-44487-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44487-9_4/.

Rosen:2005:DSI

[Ros05] Jay Rosen. Derivatives of self-intersection local times. Lecture Notes in Mathematics, 1857:263–281, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-31449-3_18.
pdf.

Roth:2009:BM

[Rot09a] Walter Roth. Back matter. Lecture Notes in Mathematics, 1964: 341–362, 2009. CODEN LNMAA2. ISBN 3-540-87564-6 (print), 3-540-87565-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-87565-9/1.pdf.

Roth:2009:FM

[Rot09b] Walter Roth. Front matter. Lecture Notes in Mathematics, 1964: i–x, 2009. CODEN LNMAA2. ISBN 3-540-87564-6 (print), 3-540-87565-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-87565-9/1.pdf.

Roth:2009:I

[Rot09c] Walter Roth. Introduction. Lecture Notes in Mathematics, 1964: 1–7, 2009. CODEN LNMAA2. ISBN 3-540-87564-6 (print), 3-540-87565-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-87565-9_1.pdf.

Roth:2009:LCC

[Rot09d] Walter Roth. Locally convex cones. Lecture Notes in Mathematics, 1964:9–117, 2009. CODEN LNMAA2. ISBN 3-540-87564-6 (print), 3-540-87565-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-87565-9_2.pdf.

Roth:2009:MIG

[Rot09e] Walter Roth. Measures and integrals. the general theory. Lecture Notes in Mathematics, 1964:119–248, 2009. CO-DEN LNMAA2. ISBN 3-540-87564-6 (print), 3-540-87565-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-87565-9_3.pdf.

Roth:2009:MLC

[Rot09f] Walter Roth. Measures on locally compact spaces. Lecture Notes in Mathematics, 1964:249–340, 2009. CODEN LNMAA2. ISBN 3-540-87564-6 (print), 3-540-87565-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-87565-9_4.pdf.

Roth:2009:OVM

[Rot09g] Walter Roth. Operator-Valued Measures and Integrals for Cone-Valued Functions, volume 1964 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-87564-6 (print), 3-540-87565-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 356 pp. LCCN QA3 .L28 no. 1964.

URL http://link.springer.com/book/10.1007/978-3-540-
87565-9;   http://www.springerlink.com/content/978-3-
540-87565-9.

Rosu:2003:DBS

[RS03] Ioanid Rosu and Dan Stroock. On the derivation of the Black– Scholes formula. Lecture Notes in Mathematics, 1832:399–414, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_18/.

Roberts:2007:BM

[RS07a] Brooks Roberts and Ralf Schmidt. Back matter. Lecture Notes in Mathematics, 1918:269–307, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-73324-9/1.pdf.

Roberts:2007:FM

[RS07b] Brooks Roberts and Ralf Schmidt. Front matter. Lecture Notes in Mathematics, 1918:I–VIII, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-73324-9/1.pdf.

Roberts:2007:HO

[RS07c] Brooks Roberts and Ralf Schmidt. Hecke operators. Lecture Notes in Mathematics, 1918:187–237, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-73324-9_6.pdf.

Roberts:2007:LN

[RS07d] Brooks Roberts and Ralf Schmidt. Local Newforms for GSp(4), volume 1918 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 307 pp. LCCN QA3 .L28 no. 1918. URL http://link.

springer.com/book/10.1007/978-3-540-73324-9;   http:/
/www.springerlink.com/content/978-3-540-73324-9.

Roberts:2007:NSR

[RS07e] Brooks Roberts and Ralf Schmidt. Non-supercuspidal representations. Lecture Notes in Mathematics, 1918:151–186, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73324-9_5.pdf.

Roberts:2007:PV

[RS07f] Brooks Roberts and Ralf Schmidt. Paramodular vectors. Lecture Notes in Mathematics, 1918:85–122, 2007. CODEN LNMAA2.

ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-73324-9_3.
pdf.

Roberts:2007:PMT

[RS07g] Brooks Roberts and Ralf Schmidt. Proofs of the main theorems. Lecture Notes in Mathematics, 1918:239–267, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73324-9_7.pdf.

Roberts:2007:RT

[RS07h] Brooks Roberts and Ralf Schmidt. Representation theory. Lecture Notes in Mathematics, 1918:27–83, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-73324-9_2.
pdf.

Roberts:2007:S

[RS07i] Brooks Roberts and Ralf Schmidt. A summary. Lecture Notes in Mathematics, 1918:1–25, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-73324-9_1.pdf.

Roberts:2007:ZI

[RS07j] Brooks Roberts and Ralf Schmidt. Zeta integrals. Lecture Notes in Mathematics, 1918:123–149, 2007. CODEN LNMAA2. ISBN 3-540-73323-X (print), 3-540-73324-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-73324-9_4.pdf.

Ruggeri:2008:ETG

[Rug08] Tommaso Ruggeri. From extended thermodynamics to granular materials. Lecture Notes in Mathematics, 1937:91–107, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78277-3_5.pdf.

Runde:2002:PD

[Run02a] Volker Runde. 0. Paradoxical decompositions. Lecture Notes in Mathematics, 1774:1–15, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45560-8_1/.

Runde:2002:ALC

[Run02b] Volker Runde. 1. Amenable, locally compact groups. Lecture Notes in Mathematics, 1774:17–36, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45560-8_2/.

Runde:2002:ABA

[Run02c] Volker Runde. 2. Amenable Banach algebras. Lecture Notes in Mathematics, 1774:37–61, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45560-8_3/.

Runde:2002:EAB

[Run02d] Volker Runde. 3. Examples of amenable Banach algebras. Lecture Notes in Mathematics, 1774:63–81, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45560-8_4/.

Runde:2002:ALP

[Run02e] Volker Runde. 4. Amenability-like properties. Lecture Notes in Mathematics, 1774:83–117, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45560-8_5/.

Runde:2002:BH

[Run02f] Volker Runde. 5. Banach homology. Lecture Notes in Mathematics, 1774:119–139, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45560-8_6/.

Runde:2002:A

[Run02g] Volker Runde. 6. C^{∗}- and W^{∗}-algebras. Lecture Notes in Mathematics, 1774:141–190, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45560-8_7/.

Runde:2002:OA

[Run02h] Volker Runde. 7. Operator amenability. Lecture Notes in Mathematics, 1774:191–207, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45560-8_8/.

Runde:2002:GSH

[Run02i] Volker Runde. 8. Geometry of spaces of homomorphisms. Lecture Notes in Mathematics, 1774:209–219, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45560-8_9/.

Runde:2002:AHA

[Run02j] Volker Runde. A abstract harmonic analysis. Lecture Notes in Mathematics, 1774:231–241, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45560-8_11/.

Runde:2002:BTP

[Run02k] Volker Runde. B tensor products. Lecture Notes in Mathematics, 1774:243–254, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45560-8_12/.

Runde:2002:CBS

[Run02l] Volker Runde. C Banach space properties. Lecture Notes in Mathematics, 1774:255–263, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45560-8_13/.

Runde:2002:DOS

[Run02m] Volker Runde. D operator spaces. Lecture Notes in Mathematics, 1774:265–274, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45560-8_14/.

Runde:2002:I

[Run02n] Volker Runde. Index. Lecture Notes in Mathematics, 1774:289– 296, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45560-8_17/.

Runde:2002:LA

[Run02o] Volker Runde. Lectures on Amenability, volume 1774 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 296 pp. LCCN QA3 .L28 no. 1774. URL http:/

/link.springer.com/book/10.1007/b82937;   http://www.
springerlink.com/content/978-3-540-45560-8.

Runde:2002:LS

[Run02p] Volker Runde. List of symbols. Lecture Notes in Mathematics, 1774:275–280, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45560-8_15/.

Runde:2002:OP

[Run02q] Volker Runde. Open problems. Lecture Notes in Mathematics, 1774:221–229, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45560-8_10/.

Runde:2002:R

[Run02r] Volker Runde. References. Lecture Notes in Mathematics, 1774: 281–288, 2002. CODEN LNMAA2. ISBN 3-540-42852-6 (print), 3-540-45560-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45560-8_16/.

Rruzicka:2000:EFM

[Růž00a] Michael Růžička. Electrorheological Fluids: Modeling and Mathematical Theory, volume 1748 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-41385-5 (print), 3-540-44427-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 151 (est.) pp. LCCN QA3 .L28 no. 1748.

URL http://link.springer.com/book/10.1007/BFb0104029;
 http://www.springerlink.com/content/978-3-540-44427-
5.

Rocircuzicka:2000:EFSa

[Růž00b] Michael Růžička. Electrorheological fluids with shear dependent viscosities: Steady flows. Lecture Notes in Mathematics, 1748: 61–103, 2000. CODEN LNMAA2. ISBN 3-540-41385-5 (print), 3-540-44427-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104032/.

Rocircuzicka:2000:EFSb

[Růž00c] Michael Růžička. Electrorheological fluids with shear dependent viscosities: Unsteady flows. Lecture Notes in Mathematics, 1748: 105–151, 2000. CODEN LNMAA2. ISBN 3-540-41385-5 (print), 3-540-44427-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104033/.

Rocircuzicka:2000:MF

[Růž00d] Michael Růžička. Mathematical framework. Lecture Notes in Mathematics, 1748:39–59, 2000. CODEN LNMAA2. ISBN 3-540-41385-5 (print), 3-540-44427-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104031/.

Rocircuzicka:2000:MEF

[Růž00e] Michael Růžička. Modeling of electrorheological fluids. Lecture Notes in Mathematics, 1748:1–37, 2000. CODEN LN-MAA2. ISBN 3-540-41385-5 (print), 3-540-44427-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104030/.

Russo:2007:ESC

[RV07] Francesco Russo and Pierre Vallois. Elements of stochastic calculus via regularization. Lecture Notes in Mathematics, 1899:147– 185, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_7.pdf.

Rezakhanlou:2008:FM

[RV08] Fraydoun Rezakhanlou and Cédric Villani. Front matter. Lecture Notes in Mathematics, 1916:i–xii, 2008. CODEN LNMAA2. ISBN 3-540-73704-9 (print), 3-540-73705-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-73705-6/1.pdf.

Roynette:2006:PQE

[RVY06] Bernard Roynette, Pierre Vallois, and Marc Yor. Pénalisations et quelques extensions du théorème de Pitman, relatives au mouvement brownien et à son. (French) []. Lecture Notes in Mathematics, 1874:305–336, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35513-7_20.pdf.

Rodino:2008:PDO

[RW08] Luigi Rodino and M. W. Wong. Pseudo-Differential Operators: Quantization and Signals, volume 1949 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxiv + 204 pp. LCCN QA3 .L28 no.

1949. URL http://link.springer.com/book/10.1007/978-
3-540-68268-4;   http://www.springerlink.com/content/
978-3-540-68268-4.

Roynette:2009:BM

[RY09a] Bernard Roynette and Marc Yor. Back matter. Lecture Notes in Mathematics, 1969:1–21, 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-89699-9/1.pdf.

Roynette:2009:FKP

[RY09b] Bernard Roynette and Marc Yor. Feynman–Kac penalisations for Brownian motion. Lecture Notes in Mathematics, 1969:1–64, 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89699-9_3.pdf.

Roynette:2009:FM

[RY09c] Bernard Roynette and Marc Yor. Front matter. Lecture Notes in Mathematics, 1969:1–11, 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-89699-9/1.pdf.

Roynette:2009:GPS

[RY09d] Bernard Roynette and Marc Yor. A general principle and some questions about penalisations. Lecture Notes in Mathematics, 1969:1–36, 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-89699-9_5.pdf.

Roynette:2009:I

[RY09e] Bernard Roynette and Marc Yor. Introduction. Lecture Notes in Mathematics, 1969:1–34, 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89699-9_1.pdf.

Roynette:2009:PBPa

[RY09f] Bernard Roynette and Marc Yor. Penalisations of a Bessel process with dimension d(0d2) by a function of the ranked lengths of its excursions. Lecture Notes in Mathematics, 1969:1–93, 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89699-9_4.pdf.

Roynette:2009:PBPb

[RY09g] Bernard Roynette and Marc Yor. Penalising Brownian Paths, volume 1969 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 275 pp. LCCN QA3 .L28 no. 1969. URL http://link.

springer.com/book/10.1007/978-3-540-89699-9;   http:/
/www.springerlink.com/content/978-3-540-89699-9.

Roynette:2009:SPW

[RY09h] Bernard Roynette and Marc Yor. Some penalisations of the Wiener measure. Lecture Notes in Mathematics, 1969:1–31, 2009. CODEN LNMAA2. ISBN 3-540-89698-8 (print), 3-540-89699-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89699-9_2.pdf.

Ryznar:2009:NCH

[Ryz09] M. Ryznar. Nontangential convergence for α-harmonic functions. Lecture Notes in Mathematics, 1980:57–72, 2009. CO-DEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-02141-1_3.pdf.

Saeki:2004:I

[Sae04a] Osamu Saeki. Introduction. Lecture Notes in Mathematics, 1854: 1–5, 2004. CODEN LNMAA2. ISBN 3-540-23021-1 (print), 3-540-44648-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44648-4_1/.

Saeki:2004:LSI

[Sae04b] Osamu Saeki. List of symbols and index. Lecture Notes in Mathematics, 1854:135–145, 2004. CODEN LNMAA2. ISBN 3-540-23021-1 (print), 3-540-44648-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-44648-4_6/.

Saeki:2004:PCS

[Sae04c] Osamu Saeki. Part I: Classification of singular fibers. Lecture Notes in Mathematics, 1854:7–57, 2004. CODEN LNMAA2. ISBN 3-540-23021-1 (print), 3-540-44648-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-44648-4_2/.

Saeki:2004:PIU

[Sae04d] Osamu Saeki. Part II: Universal complex of singular fibers. Lecture Notes in Mathematics, 1854:59–120, 2004. CODEN LN-MAA2. ISBN 3-540-23021-1 (print), 3-540-44648-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44648-4_
3/.

Saeki:2004:PIE

[Sae04e] Osamu Saeki. Part III: Epilogue. Lecture Notes in Mathematics, 1854:121–129, 2004. CODEN LNMAA2. ISBN 3-540-23021-1 (print), 3-540-44648-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44648-4_4/.

Saeki:2004:R

[Sae04f] Osamu Saeki. References. Lecture Notes in Mathematics, 1854: 131–134, 2004. CODEN LNMAA2. ISBN 3-540-23021-1 (print), 3-540-44648-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44648-4_5/.

Saeki:2004:TSF

[Sae04g] Osamu Saeki. Topology of Singular Fibers of Differentiable Maps, volume 1854 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-23021-1 (print), 3-540-44648-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 146 pp. LCCN QA3 .L28 no. 1854. URL http://

link.springer.com/book/10.1007/b100393;   http://www.
springerlink.com/content/978-3-540-44648-4.

Sampson:2008:SFN

[Sam08] W. W. Sampson. Stochastic fibre networks. Lecture Notes in Mathematics, 1953:161–194, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69393-2_9.pdf.

Sanderson:2005:BCC

[San05] M. J. Sanderson. Basic concepts of Ca^{2+} signaling in cells and tissues. Lecture Notes in Mathematics, 1867:1–13, 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11406501_1.pdf.

Schechtman:2000:ECP

[Sch00] G. Schechtman. An editorial comment on the preceding paper. Lecture Notes in Mathematics, 1745:19–20, 2000. CODEN LN-MAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107204.

Schmidt:2002:Ia

[Sch02a] Bernhard Schmidt. 1. Introduction. Lecture Notes in Mathematics, 1797:1–25, 2002. CODEN LNMAA2. ISBN 3-540-44243-X (print), 3-540-45797-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45797-8_1/.

Schmidt:2002:FD

[Sch02b] Bernhard Schmidt. 2. The field descent. Lecture Notes in Mathematics, 1797:27–51, 2002. CODEN LNMAA2. ISBN 3-540-44243-X (print), 3-540-45797-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45797-8_2/.

Schmidt:2002:EB

[Sch02c] Bernhard Schmidt. 3. Exponent bounds. Lecture Notes in Mathematics, 1797:53–78, 2002. CODEN LNMAA2. ISBN 3-540-44243-X (print), 3-540-45797-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45797-8_3/.

Schmidt:2002:TWI

[Sch02d] Bernhard Schmidt. 4. Two-weight irreducible cyclic codes. Lecture Notes in Mathematics, 1797:79–90, 2002. CODEN LNMAA2. ISBN 3-540-44243-X (print), 3-540-45797-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45797-8_4/.

Schmidt:2002:B

[Sch02e] Bernhard Schmidt. Bibliography. Lecture Notes in Mathematics, 1797:91–98, 2002. CODEN LNMAA2. ISBN 3-540-44243-X (print), 3-540-45797-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45797-8_5/.

Schmidt:2002:CCF

[Sch02f] Bernhard Schmidt. Characters and Cyclotomic Fields in Finite Geometry, volume 1797 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-44243-X (print), 3-540-45797-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 100 pp. LCCN QA3 .L28 no. 1797. URL

http://link.springer.com/book/10.1007/b84213;   http:
//www.springerlink.com/content/978-3-540-45797-8.

Schmidt:2002:Ib

[Sch02g] Bernhard Schmidt. Index. Lecture Notes in Mathematics, 1797: 99–100, 2002. CODEN LNMAA2. ISBN 3-540-44243-X (print), 3-540-45797-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45797-8_6.

Schindler:2003:I

[Sch03a] Werner Schindler. 1. Introduction. Lecture Notes in Mathematics, 1808:1–9, 2003. CODEN LNMAA2. ISBN 3-540-00235-9 (print), 3-540-36210-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36210-4_1/.

Schindler:2003:MT

[Sch03b] Werner Schindler. 2. Main theorems. Lecture Notes in Mathematics, 1808:11–54, 2003. CODEN LNMAA2. ISBN 3-540-00235-9 (print), 3-540-36210-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36210-4_2/.

Schindler:2003:SAA

[Sch03c] Werner Schindler. 3. Significance, applicability and advantages. Lecture Notes in Mathematics, 1808:55–62, 2003. CODEN LN-MAA2. ISBN 3-540-00235-9 (print), 3-540-36210-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36210-4_

Schindler:2003:A

[Sch03d] Werner Schindler. 4. Applications. Lecture Notes in Mathematics, 1808:63–153, 2003. CODEN LNMAA2. ISBN 3-540-00235-9 (print), 3-540-36210-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36210-4_4/.

Schindler:2003:GI

[Sch03e] Werner Schindler. Glossary and index. Lecture Notes in Mathematics, 1808:159–167, 2003. CODEN LNMAA2. ISBN 3-540-00235-9 (print), 3-540-36210-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36210-4_6/.

Schindler:2003:MSP

[Sch03f] Werner Schindler. Measures with Symmetry Properties, volume 1808 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00235-9 (print), 3-540-36210-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 167 pp. LCCN QA3 .L28 no. 1808. URL http:/

/link.springer.com/book/10.1007/b80163;   http://www.
springerlink.com/content/978-3-540-36210-4.

Schindler:2003:R

[Sch03g] Werner Schindler. References. Lecture Notes in Mathematics, 1808:155–158, 2003. CODEN LNMAA2. ISBN 3-540-00235-9 (print), 3-540-36210-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36210-4_5/.

Schlickewei:2003:AAN

[Sch03h] Hans Peter Schlickewei. Approximation of algebraic numbers. Lecture Notes in Mathematics, 1819:107–170, 2003. CODEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44979-5_3/.

Schmidt:2003:LRS

[Sch03i] Wolfgang M. Schmidt. Linear recurrence sequences. Lecture Notes in Mathematics, 1819:171–247, 2003. CODEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44979-5_4/.

Schachermayer:2004:UMI

[Sch04a] Walter Schachermayer. Utility maximisation in incomplete markets. Lecture Notes in Mathematics, 1856:255–293, 2004. CODEN LNMAA2. ISBN 3-540-22953-1 (print), 3-540-44644-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-44644-6_
5/.

Schnaubelt:2004:ABP

[Sch04b] Roland Schnaubelt. Asymptotic behaviour of parabolic nonautonomous evolution equations. Lecture Notes in Mathematics, 1855:401–472, 2004. CODEN LNMAA2. ISBN 3-540-23030-0 (print), 3-540-44653-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44653-8_5/.

Schechtman:2007:RVD

[Sch07a] Gideon Schechtman. The random version of Dvoretzky’s Theorem in l_{∞}^{n}. Lecture Notes in Mathematics, 1910:265–270, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_15.pdf.

Schuster:2007:ADC

[Sch07b] Thomas Schuster. Approaches for defect correction. Lecture Notes in Mathematics, 1906:89–103, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_9.pdf.

Schuster:2007:AIX

[Sch07c] Thomas Schuster. Approximate inverse and X-ray diffractometry. Lecture Notes in Mathematics, 1906:151–164, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71227-5_16.pdf.

Schuster:2007:AIS

[Sch07d] Thomas Schuster. Approximate inverse in L^{2}-spaces. Lecture Notes in Mathematics, 1906:11–24, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71227-5_2.
pdf.

Schuster:2007:AID

[Sch07e] Thomas Schuster. Approximate inverse in distribution spaces. Lecture Notes in Mathematics, 1906:39–47, 2007. CODEN LN-MAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-540-71227-
5_4.pdf.

Schuster:2007:AIH

[Sch07f] Thomas Schuster. Approximate inverse in Hilbert spaces. Lecture Notes in Mathematics, 1906:25–38, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71227-5_3.
pdf.

Schuster:2007:BM

[Sch07g] Thomas Schuster. Back matter. Lecture Notes in Mathematics, 1906:189–202, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-71227-5/1.pdf.

Schuster:2007:CRKa

[Sch07h] Thomas Schuster. Computation of reconstruction kernels. Lecture Notes in Mathematics, 1906:133–137, 2007. CO-DEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71227-5_13.pdf.

Schuster:2007:CRKb

[Sch07i] Thomas Schuster. Computation of reconstruction kernels in 3D computerized tomography. Lecture Notes in Mathematics, 1906:181–185, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-71227-5_18.pdf.

Schuster:2007:CPa

[Sch07j] Thomas Schuster. Conclusion and perspectives. Lecture Notes in Mathematics, 1906:49, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_5.pdf.

Schuster:2007:CPb

[Sch07k] Thomas Schuster. Conclusion and perspectives. Lecture Notes in Mathematics, 1906:105–106, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_10.pdf.

Schuster:2007:CPc

[Sch07l] Thomas Schuster. Conclusion and perspectives. Lecture Notes in Mathematics, 1906:145, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_15.pdf.

Schuster:2007:CPd

[Sch07m] Thomas Schuster. Conclusion and perspectives. Lecture Notes in Mathematics, 1906:187, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_19.pdf.

Schuster:2007:CS

[Sch07n] Thomas Schuster. Convergence and stability. Lecture Notes in Mathematics, 1906:81–87, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_8.pdf.

Schuster:2007:DM

[Sch07o] Thomas Schuster. Design of a mollifier. Lecture Notes in Mathematics, 1906:123–131, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_12.pdf.

Schuster:2007:FBA

[Sch07p] Thomas Schuster. A filtered backprojection algorithm. Lecture Notes in Mathematics, 1906:165–179, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71227-5_17.
pdf.

Schuster:2007:FMa

[Sch07q] Thomas Schuster. Front matter. Lecture Notes in Mathematics, 1906:I–XIII, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-71227-5/1.pdf.

Schuster:2007:FMb

[Sch07r] Thomas Schuster. Front matter. Lecture Notes in Mathematics, 1906:1–4, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-71227-5/1/1.pdf.

Schuster:2007:FMc

[Sch07s] Thomas Schuster. Front matter. Lecture Notes in Mathematics, 1906:51–54, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-71227-5/2/1.pdf.

Schuster:2007:FMd

[Sch07t] Thomas Schuster. Front matter. Lecture Notes in Mathematics, 1906:107–110, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-71227-5/3/1.pdf.

Schuster:2007:FMe

[Sch07u] Thomas Schuster. Front matter. Lecture Notes in Mathematics, 1906:147–149, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-71227-5/4/1.pdf.

Schuster:2007:IPP

[Sch07v] Thomas Schuster. Ill-posed problems and regularization methods. Lecture Notes in Mathematics, 1906:5–9, 2007. CO-DEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71227-5_1.pdf.

Schuster:2007:MAI

[Sch07w] Thomas Schuster. The Method of Approximate Inverse: Theory and Applications, volume 1906 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 198 pp. LCCN QA3 .L28 no.

1906. URL http://link.springer.com/book/10.1007/978-
3-540-71227-5;   http://www.springerlink.com/content/
978-3-540-71227-5.

Schuster:2007:NE

[Sch07x] Thomas Schuster. Numerical experiments. Lecture Notes in Mathematics, 1906:139–144, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_14.pdf.

Schuster:2007:SDS

[Sch07y] Thomas Schuster. A semi-discrete setup for Doppler tomography. Lecture Notes in Mathematics, 1906:55–61, 2007. CO-DEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71227-5_6.pdf.

Schuster:2007:SSD

[Sch07z] Thomas Schuster. Solving the semi-discrete problem. Lecture Notes in Mathematics, 1906:63–79, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-71227-5_7.
pdf.

Schuster:2007:SMO

[Sch07-27] Thomas Schuster. The spherical mean operator. Lecture Notes in Mathematics, 1906:111–121, 2007. CODEN LNMAA2. ISBN 3-540-71226-7 (print), 3-540-71227-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-71227-5_11.pdf.

Seidel:2008:LFD

[Sei08] Paul Seidel. Lectures on four-dimensional Dehn twists. Lecture Notes in Mathematics, 1938:231–267, 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-78279-7_4.
pdf.

Serlet:2000:LIL

[Ser00] Laurent Serlet. Laws of the iterated logarithm for the Brownian snake. Lecture Notes in Mathematics, 1729:302–312, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103809/.

Serlet:2007:RMB

[Ser07a] Laurent Serlet. Representation of the martingales for the Brownian snake. Lecture Notes in Mathematics, 1899:343–354, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_18.pdf.

Serre:2007:DSP

[Ser07b] Denis Serre. Discrete shock profiles: Existence and stability. Lecture Notes in Mathematics, 1911:79–158, 2007. CO-DEN LNMAA2. ISBN 3-540-72186-X (print), 3-540-72187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72187-1_2.pdf.

Serlet:2008:CDD

[Ser08] Laurent Serlet. Creation or deletion of a drift on a Brownian trajectory. Lecture Notes in Mathematics, 1934:215–232, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_11.pdf.

Schurmann:2005:QII

[SF05] Michael Schürmann and Uwe Franz. Quantum Independent Increment Processes I: From Classical Probability to Quantum Stochastic Calculus, volume 1865 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-24406-9 (print), 3-540-31450-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 299 pp. LCCN QA3 .L28 no. 1865.

URL http://link.springer.com/book/10.1007/b105131;
http://www.springerlink.com/content/978-3-540-31450-9.

Schuermann:2006:QII

[SF06] Michael Schüermann and Uwe Franz. Quantum Independent Increment Processes II: Structure of Quantum Levy Processes, Classical Probability, and Physics, volume 1866 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-24407-7 (print), 3-540-32385-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 336 pp. LCCN

QA3 .L28 no. 1866. URL http://link.springer.com/book/10.
1007/11376637;   http://www.springerlink.com/content/
978-3-540-32385-3.

Scharcanski:2008:BM

[SF08a] J. Scharcanski and S. Felipussi. Back matter. Lecture Notes in Mathematics, 1953:235–253, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-69393-2/1.pdf.

Scharcanski:2008:QC

[SF08b] J. Scharcanski and S. Felipussi. Quantum chaology. Lecture Notes in Mathematics, 1953:223–233, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69393-2_11.pdf.

Scharcanski:2008:SPM

[SF08c] J. Scharcanski and S. Felipussi. Stochastic porous media and hydrology. Lecture Notes in Mathematics, 1953:195–222, 2008. CODEN LNMAA2. ISBN 3-540-69391-2 (print), 3-540-69393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69393-2_10.pdf.

Serre:2002:ALM

[SG02] Jean-Pierre Serre and Pierre Gabriel, editors. Algèbre locale, multiplicités; cours au Collège de France, 1957–1958. (French) [Local algebra, multiplicities; courses at the Collège of France, 1957–1958], Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-07028-1 (Berlin), 0-387-07028-1 (New York). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN ???? URL http:/

/link.springer.com/book/10.1007/BFb0087610;   http://
www.springerlink.com/openurl.asp?genre=book&isbn=978-
3-540-07028-3.

Shannon:2005:ICM

[Sha05] T. R. Shannon. Integrated calcium management in cardiac myocytes. Lecture Notes in Mathematics, 1867:63–95, 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/11406501_3.pdf.

Sherrington:2007:SGP

[She07] David Sherrington. Spin glasses: a perspective. Lecture Notes in Mathematics, 1900:45–62, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-40908-3_2.pdf.

Shigekawa:2005:MTH

[Shi05] Ichiro Shigekawa. L^{p} multiplier theorem for the Hodge–Kodaira operator. Lecture Notes in Mathematics, 1857:226–246, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_16.pdf.

Siburg:2004:AMT

[Sib04a] Karl Friedrich Siburg. 1. Aubry–Mather theory. Lecture Notes in Mathematics, 1844:1–13, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40985-4_1/.

Siburg:2004:MMT

[Sib04b] Karl Friedrich Siburg. 2. Mather–Mañé theory. Lecture Notes in Mathematics, 1844:15–35, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40985-4_2/.

Siburg:2004:MAC

[Sib04c] Karl Friedrich Siburg. 3. The minimal action and convex billiards. Lecture Notes in Mathematics, 1844:37–57, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40985-4_3/.

Siburg:2004:MAN

[Sib04d] Karl Friedrich Siburg. 4. The minimal action near fixed points and invariant tori. Lecture Notes in Mathematics, 1844:59– 80, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40985-4_4/.

Siburg:2004:MAH

[Sib04e] Karl Friedrich Siburg. 5. The minimal action and Hofer’s geometry. Lecture Notes in Mathematics, 1844:81–95, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40985-4_
5/.

Siburg:2004:MAS

[Sib04f] Karl Friedrich Siburg. 6. The minimal action and symplectic geometry. Lecture Notes in Mathematics, 1844:97–119, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40985-4_6/.

Siburg:2004:PLA

[Sib04g] Karl Friedrich Siburg. The Principle of Least Action in Geometry and Dynamics, volume 1844 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LN-MAA2. ISBN 3-540-21944-7 (paperback), 3-540-40985-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 128 pp. LCCN QA3 .L28 no. 1844. URL http:/

/link.springer.com/book/10.1007/b97327;   http://www.
loc.gov/catdir/enhancements/fy0813/2004104313-d.html;
 http://www.loc.gov/catdir/enhancements/fy0813/2004104313-
t.html;   http://www.springerlink.com/content/978-3-540-
40985-4;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0075-8434&volume=1844.

Siburg:2004:R

[Sib04h] Karl Friedrich Siburg. References. Lecture Notes in Mathematics, 1844:121–125, 2004. CODEN LNMAA2. ISBN 3-540-21944-7 (print), 3-540-40985-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-40985-4_7/.

Siegert:2009:BM

[Sie09a] Wolfgang Siegert. Back matter. Lecture Notes in Mathematics, 1963:231–260, 2009. CODEN LNMAA2. ISBN 3-540-85963-2 (print), 3-540-85964-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-85964-2/1.pdf.

Siegert:2009:EPD

[Sie09b] Wolfgang Siegert. Exit probabilities for degenerate systems. Lecture Notes in Mathematics, 1963:125–142, 2009. CO-DEN LNMAA2. ISBN 3-540-85963-2 (print), 3-540-85964-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85964-2_3.pdf.

Siegert:2009:FM

[Sie09c] Wolfgang Siegert. Front matter. Lecture Notes in Mathematics, 1963:i–xvi, 2009. CODEN LNMAA2. ISBN 3-540-85963-2 (print), 3-540-85964-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-85964-2/1.pdf.

Siegert:2009:LDS

[Sie09d] Wolfgang Siegert. Linear differential systems with parameter excitation. Lecture Notes in Mathematics, 1963:9–51, 2009. CODEN LNMAA2. ISBN 3-540-85963-2 (print), 3-540-85964-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85964-2_1.pdf.

Siegert:2009:LLEa

[Sie09e] Wolfgang Siegert. Local Lyapunov exponents. Lecture Notes in Mathematics, 1963:143–229, 2009. CODEN LNMAA2. ISBN 3-540-85963-2 (print), 3-540-85964-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-85964-2_4.pdf.

Siegert:2009:LLEb

[Sie09f] Wolfgang Siegert. Local Lyapunov Exponents: Sublimiting Growth Rates of Linear Random Differential Equations, volume 1963 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CO-DEN LNMAA2. ISBN 3-540-85963-2 (print), 3-540-85964-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 254 pp. LCCN QA3 .L28 no. 1963. URL http://link.

springer.com/book/10.1007/978-3-540-85964-2;   http:/
/www.springerlink.com/content/978-3-540-85964-2.

Siegert:2009:LTS

[Sie09g] Wolfgang Siegert. Locality and time scales of the underlying non-degenerate stochastic system: Freidlin–Wentzell theory. Lecture Notes in Mathematics, 1963:53–123, 2009. CO-DEN LNMAA2. ISBN 3-540-85963-2 (print), 3-540-85964-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85964-2_2.pdf.

Simon:2003:SEI

[Sim03] Thomas Simon. Support d’une équation d’Itô avec sauts en dimension 1. (French) []. Lecture Notes in Mathematics, 1801:314– 330, 2003. CODEN LNMAA2. ISBN 3-540-00072-0 (print), 3-540-36107-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36107-7_12/.

Simons:2008:BM

[Sim08a] Stephen Simons. Back matter. Lecture Notes in Mathematics, 1693:233–248, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-1-4020-6919-2/1.pdf.

Simons:2008:FD

[Sim08b] Stephen Simons. Fenchel duality. Lecture Notes in Mathematics, 1693:41–69, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-4020-6919-2_3.pdf.

Simons:2008:HBM

[Sim08c] Stephen Simons. From Hahn–Banach to Monotonicity, volume 1693 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 244 pp. LCCN QA3 .L28 no. 1693. URL http://link.

springer.com/book/10.1007/978-1-4020-6919-2;   http:/
/www.springerlink.com/content/978-1-4020-6919-2.

Simons:2008:FM

[Sim08d] Stephen Simons. Front matter. Lecture Notes in Mathematics, 1693:I–XIV, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-1-4020-6919-2/1.pdf.

Simons:2008:GCM

[Sim08e] Stephen Simons. Glossary of classes of multifunctions. Lecture Notes in Mathematics, 1693:205–206, 2008. CODEN LNMAA2.

ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-1-4020-6919-2_10.
pdf.

Simons:2008:HBL

[Sim08f] Stephen Simons. The Hahn–Banach–Lagrange theorem and some consequences. Lecture Notes in Mathematics, 1693:15–39, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-4020-6919-2_2.pdf.

Simons:2008:I

[Sim08g] Stephen Simons. Introduction. Lecture Notes in Mathematics, 1693:1–13, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-4020-6919-2_1.pdf.

Simons:2008:MMG

[Sim08h] Stephen Simons. Monotone multifunctions on general Banach spaces. Lecture Notes in Mathematics, 1693:107–115, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-4020-6919-2_5.pdf.

Simons:2008:MMR

[Sim08i] Stephen Simons. Monotone multifunctions on reflexive Banach spaces. Lecture Notes in Mathematics, 1693:117–138, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-4020-6919-2_6.pdf.

Simons:2008:MSS

[Sim08j] Stephen Simons. Multifunctions, SSD spaces, monotonicity and Fitzpatrick functions. Lecture Notes in Mathematics, 1693:71– 105, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-1-4020-6919-2_4.pdf.

Simons:2008:OP

[Sim08k] Stephen Simons. Open problems. Lecture Notes in Mathematics, 1693:203–204, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-1-4020-6919-2_9.pdf.

Simons:2008:SR

[Sim08l] Stephen Simons. A selection of results. Lecture Notes in Mathematics, 1693:207–231, 2008. CODEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-1-4020-6919-2_11.pdf.

Simons:2008:SMM

[Sim08m] Stephen Simons. Special maximally monotone multifunctions. Lecture Notes in Mathematics, 1693:139–195, 2008. CO-DEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-4020-6919-2_7.pdf.

Simons:2008:SPG

[Sim08n] Stephen Simons. The sum problem for general Banach spaces. Lecture Notes in Mathematics, 1693:197–201, 2008. CO-DEN LNMAA2. ISBN 1-4020-6918-9 (print), 1-4020-6919-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
1-4020-6919-2_8.pdf.

Sinai:2008:BM

[Sin08a] Yakov G. Sinai. Back matter. Lecture Notes in Mathematics, 1942:165–174, 2008. CODEN LNMAA2. ISBN 3-540-78492-6 (print), 3-540-78493-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-78493-7/1.pdf.

Sinai:2008:MRR

[Sin08b] Yakov G. Sinai. Mathematical results related to the Navier– Stokes system. Lecture Notes in Mathematics, 1942:151–164, 2008. CODEN LNMAA2. ISBN 3-540-78492-6 (print), 3-540-78493-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78493-7_3.pdf.

Slade:2006:BM

[Sla06a] Gordon Slade. Back matter. Lecture Notes in Mathematics, 1879: 211–232, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-35518-2/1.pdf.

Slade:2006:BRW

[Sla06b] Gordon Slade. Branching random walk. Lecture Notes in Mathematics, 1879:171–182, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35518-2_15.pdf.

Slade:2006:CP

[Sla06c] Gordon Slade. The contact process. Lecture Notes in Mathematics, 1879:161–170, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35518-2_14.pdf.

Slade:2006:CSA

[Sla06d] Gordon Slade. Convergence for the self-avoiding walk. Lecture Notes in Mathematics, 1879:41–55, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35518-2_5.
pdf.

Slade:2006:DES

[Sla06e] Gordon Slade. Diagrammatic estimates for the self-avoiding walk. Lecture Notes in Mathematics, 1879:31–40, 2006. CO-DEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35518-2_4.pdf.

Slade:2006:EP

[Sla06f] Gordon Slade. The expansion for percolation. Lecture Notes in Mathematics, 1879:109–123, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35518-2_10.pdf.

Slade:2006:EOP

[Sla06g] Gordon Slade. Expansions for oriented percolation. Lecture Notes in Mathematics, 1879:151–159, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35518-2_13.pdf.

Slade:2006:FM

[Sla06h] Gordon Slade. Front matter. Lecture Notes in Mathematics, 1879: i–xiii, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-35518-2/1.pdf.

Slade:2006:FRS

[Sla06i] Gordon Slade. Further results for the self-avoiding walk. Lecture Notes in Mathematics, 1879:57–65, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35518-2_6.
pdf.

Slade:2006:ISB

[Sla06j] Gordon Slade. Integrated super-Brownian excursion. Lecture Notes in Mathematics, 1879:183–200, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35518-2_16.
pdf.

Slade:2006:LEL

[Sla06k] Gordon Slade. The lace expansion for lattice trees. Lecture Notes in Mathematics, 1879:77–86, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35518-2_8.pdf.

Slade:2006:LES

[Sla06l] Gordon Slade. The lace expansion for the self-avoiding walk. Lecture Notes in Mathematics, 1879:19–29, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35518-2_3.
pdf.

Slade:2006:LT

[Sla06m] Gordon Slade. Lattice trees. Lecture Notes in Mathematics, 1879:67–75, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35518-2_7.pdf.

Slade:2006:OP

[Sla06n] Gordon Slade. Oriented percolation. Lecture Notes in Mathematics, 1879:141–149, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35518-2_12.pdf.

Slade:2006:P

[Sla06o] Gordon Slade. Percolation. Lecture Notes in Mathematics, 1879:87–108, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35518-2_9.pdf.

Slade:2006:RP

[Sla06p] Gordon Slade. Results for percolation. Lecture Notes in Mathematics, 1879:125–139, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35518-2_11.pdf.

Slade:2006:SAW

[Sla06q] Gordon Slade. The self-avoiding walk. Lecture Notes in Mathematics, 1879:7–17, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35518-2_2.pdf.

Slade:2006:SRW

[Sla06r] Gordon Slade. Simple random walk. Lecture Notes in Mathematics, 1879:1–6, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35518-2_1.pdf.

Slade:2006:SBM

[Sla06s] Gordon Slade. Super-Brownian motion. Lecture Notes in Mathematics, 1879:201–210, 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35518-2_17.pdf.

Sneyd:2005:MDC

[Sne05a] J. Sneyd. Modeling IP_{3}-dependent calcium dynamics in nonexcitable cells. Lecture Notes in Mathematics, 1867:15–61, 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/11406501_2.pdf.

Sneyd:2005:TMB

[Sne05b] James Sneyd. Tutorials in Mathematical Biosciences II: Mathematical Modeling of Calcium Dynamics and Signal Transduction, volume 1867 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 202 pp. LCCN QA3 .L28 no. 1867. URL http://

link.springer.com/book/10.1007/b107088;   http://www.
springerlink.com/content/978-3-540-31438-7.

Sodin:2007:TSG

[Sod07] S. Sodin. Tail-sensitive Gaussian asymptotics for marginals of concentrated measures in high dimension. Lecture Notes in Mathematics, 1910:271–295, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72053-9_16.pdf.

Solonnikov:2003:LEF

[Sol03] Vsevolod A. Solonnikov. Lectures on evolution free boundary problems: Classical solutions. Lecture Notes in Mathematics, 1812:123–175, 2003. CODEN LNMAA2. ISBN 3-540-14033-6 (print), 3-540-39189-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39189-0_4/.

Sommerhauser:2002:P

[Som02a] Yorck Sommerhäuser. 1. Preliminaries. Lecture Notes in Mathematics, 1789:7–24, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45423-6_2/.

Sommerhauser:2002:A

[Som02b] Yorck Sommerhäuser. 10. Applications. Lecture Notes in Mathematics, 1789:141–145, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45423-6_11/.

Sommerhauser:2002:CT

[Som02c] Yorck Sommerhäuser. 2. Clifford theory. Lecture Notes in Mathematics, 1789:25–33, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45423-6_3/.

Sommerhauser:2002:E

[Som02d] Yorck Sommerhäuser. 3. Examples. Lecture Notes in Mathematics, 1789:35–47, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45423-6_4/.

Sommerhauser:2002:Ib

[Som02e] Yorck Sommerhäuser. 4. Isomorphisms. Lecture Notes in Mathematics, 1789:49–65, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45423-6_5/.

Sommerhauser:2002:C

[Som02f] Yorck Sommerhäuser. 5. Constructions. Lecture Notes in Mathematics, 1789:67–89, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45423-6_6/.

Sommerhauser:2002:CYDa

[Som02g] Yorck Sommerhäuser. 6. Commutative Yetter–Drinfel’d Hopf algebras. Lecture Notes in Mathematics, 1789:91–102, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-45423-6_7/.

Sommerhauser:2002:CYDb

[Som02h] Yorck Sommerhäuser. 7. Cocommutative Yetter–Drinfel’d Hopf algebras. Lecture Notes in Mathematics, 1789:103–113, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45423-6_8/.

Sommerhauser:2002:SHAa

[Som02i] Yorck Sommerhäuser. 8. Semisimple Hopf algebras of dimension p^{3}. Lecture Notes in Mathematics, 1789:115–129, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45423-6_9/.

Sommerhauser:2002:SHAb

[Som02j] Yorck Sommerhäuser. 9. Semisimple Hopf algebras of dimension pq. Lecture Notes in Mathematics, 1789:131–140, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45423-6_10/.

Sommerhauser:2002:Ia

[Som02k] Yorck Sommerhäuser. Introduction. Lecture Notes in Mathematics, 1789:1–6, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45423-6_1/.

Sommerhauser:2002:R

[Som02l] Yorck Sommerhäuser. References. Lecture Notes in Mathematics, 1789:147–150, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45423-6_12/.

Sommerhauser:2002:SSI

[Som02m] Yorck Sommerhäuser. Subject and symbol index. Lecture Notes in Mathematics, 1789:151–157, 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45423-6_13/.

Sommerhauser:2002:YDH

[Som02n] Yorck Sommerhäuser. Yetter–Drinfel’d Hopf Algebras over Groups of Prime Order, volume 1789 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43799-1 (print), 3-540-45423-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 158 pp. LCCN QA3 .L28 no. 1789.

URL http://link.springer.com/book/10.1007/b83955;
http://www.springerlink.com/content/978-3-540-45423-6.

Soner:2003:VDP

[Son03] Halil Mete Soner. Variational and dynamic problems for the Ginzburg–Landau functional. Lecture Notes in Mathematics, 1812:177–233, 2003. CODEN LNMAA2. ISBN 3-540-14033-6 (print), 3-540-39189-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-39189-0_5/.

Sontag:2008:ISS

[Son08] Eduardo D. Sontag. Input to state stability: Basic concepts and results. Lecture Notes in Mathematics, 1932:163–220, 2008. CODEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77653-6_3.pdf.

Slade:2006:LEA

[SP06] G. (Gordon) Slade and Jean Picard, editors. The Lace Expansion and its Applications: École d’Été de Probabilités de Saint-Flour XXXIV — 2004, volume 1879 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-31189-0 (print), 3-540-35518-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1879; QA3 .L28 no. 1879; QC174.85.P45 S57 2006. URL http://

link.springer.com/book/10.1007/b128444;   http://www.
loc.gov/catdir/enhancements/fy0663/2006921789-d.html;
 http://www.loc.gov/catdir/enhancements/fy0824/2006921789-
b.html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006921789-t.html;   http://www.springerlink.com/content/
978-3-540-35518-2.

Speicher:2003:FPT

[Spe03] R. Speicher. Free probability theory and random matrices. Lecture Notes in Mathematics, 1815:53–73, 2003. CODEN LN-MAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44890-X_3/.

Saint-Raymond:2009:BM

[SR09a] Laure Saint-Raymond. Back matter. Lecture Notes in Mathematics, 1971:1–25, 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-92847-8/1.pdf.

Saint-Raymond:2009:BEF

[SR09b] Laure Saint-Raymond. The Boltzmann equation and its formal hydrodynamic limits. Lecture Notes in Mathematics, 1971:1–34, 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-92847-8_2.pdf.

Saint-Raymond:2009:CEL

[SR09c] Laure Saint-Raymond. The compressible Euler limit. Lecture Notes in Mathematics, 1971:1–6, 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-92847-8_6.pdf.

Saint-Raymond:2009:FM

[SR09d] Laure Saint-Raymond. Front matter. Lecture Notes in Mathematics, 1971:1–10, 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-92847-8/1.pdf.

Saint-Raymond:2009:HLB

[SR09e] Laure Saint-Raymond. Hydrodynamic Limits of the Boltzmann Equation, volume 1971 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 188 pp. LCCN QA3 .L28 no. 1971.

URL http://link.springer.com/book/10.1007/978-3-540-
92847-8;   http://www.springerlink.com/content/978-3-
540-92847-8.

Saint-Raymond:2009:IEL

[SR09f] Laure Saint-Raymond. The incompressible Euler limit. Lecture Notes in Mathematics, 1971:1–42, 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-92847-8_5.
pdf.

Saint-Raymond:2009:INS

[SR09g] Laure Saint-Raymond. The incompressible Navier–Stokes limit. Lecture Notes in Mathematics, 1971:1–42, 2009. CO-DEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-92847-8_4.pdf.

Saint-Raymond:2009:I

[SR09h] Laure Saint-Raymond. Introduction. Lecture Notes in Mathematics, 1971:1–11, 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-92847-8_1.pdf.

Saint-Raymond:2009:MTD

[SR09i] Laure Saint-Raymond. Mathematical tools for the derivation of hydrodynamic limits. Lecture Notes in Mathematics, 1971:1–32, 2009. CODEN LNMAA2. ISBN 3-540-92846-4 (print), 3-540-92847-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-92847-8_3.pdf.

Schachermayer:2009:NMC

[SST09] Walter Schachermayer, Uwe Schmock, and Josef Teichmann. Non-monotone convergence in the quadratic Wasserstein distance. Lecture Notes in Mathematics, 1979:131–136, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_3.pdf.

Soner:2003:PSR

[ST03] H. Mete Soner and Nizar Touzi. The problem of super-replication under constraints. Lecture Notes in Mathematics, 1814:133– 172, 2003. CODEN LNMAA2. ISBN 3-540-40193-8 (print), 3-540-44859-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44859-4_4/.

Siebert:2008:BM

[ST08a] Bernd Siebert and Gang Tian. Back matter. Lecture Notes in Mathematics, 1938:343–345, 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-78279-7/1.pdf.

Siebert:2008:LPH

[ST08b] Bernd Siebert and Gang Tian. Lectures on pseudo-holomorphic curves and the symplectic isotopy problem. Lecture Notes in Mathematics, 1938:269–341, 2008. CODEN LNMAA2. ISBN 3-540-78278-8 (print), 3-540-78279-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-78279-7_5.pdf.

Stanton:2003:ESF

[Sta03] Dennis Stanton. Enumeration and special functions. Lecture Notes in Mathematics, 1817:137–166, 2003. CODEN LN-MAA2. ISBN 3-540-40375-2 (print), 3-540-44945-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44945-0_4/.

Steinbach:2003:P

[Ste03a] Olaf Steinbach. 1. Preliminaries. Lecture Notes in Mathematics, 1809:7–24, 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36250-0_2/.

Steinbach:2003:SR

[Ste03b] Olaf Steinbach. 2. Stability results. Lecture Notes in Mathematics, 1809:25–51, 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36250-0_3/.

Steinbach:2003:DNM

[Ste03c] Olaf Steinbach. 3. The Dirichlet–Neumann map for elliptic boundary value problems. Lecture Notes in Mathematics, 1809:53– 70, 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36250-0_4/.

Steinbach:2003:MDS

[Ste03d] Olaf Steinbach. 4. Mixed discretization schemes. Lecture Notes in Mathematics, 1809:71–83, 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36250-0_5/.

Steinbach:2003:HCD

[Ste03e] Olaf Steinbach. 5. Hybrid coupled domain decomposition methods. Lecture Notes in Mathematics, 1809:85–115, 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36250-0_
6/.

Steinbach:2003:I

[Ste03f] Olaf Steinbach. Introduction. Lecture Notes in Mathematics, 1809:1–5, 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36250-0_1/.

Steinbach:2003:R

[Ste03g] Olaf Steinbach. References. Lecture Notes in Mathematics, 1809: 117–120, 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36250-0_7/.

Steinbach:2003:SEH

[Ste03h] Olaf Steinbach. Stability Estimates for Hybrid Coupled Domain Decomposition Methods, volume 1809 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00277-4 (print), 3-540-36250-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 120 pp. LCCN QA3 .L28 no. 1809.

URL http://link.springer.com/book/10.1007/b80164;
http://www.springerlink.com/content/978-3-540-36250-0.

Stevens:2003:DSa

[Ste03i] Jan Stevens. 1. Deformations of singularities. Lecture Notes in Mathematics, 1811:5–14, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36464-1_2/.

Stevens:2003:QS

[Ste03j] Jan Stevens. 10. Quotient singularities. Lecture Notes in Mathematics, 1811:71–77, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36464-1_11/.

Stevens:2003:PM

[Ste03k] Jan Stevens. 11. The projection method. Lecture Notes in Mathematics, 1811:79–92, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36464-1_12/.

Stevens:2003:F

[Ste03l] Jan Stevens. 12. Formats. Lecture Notes in Mathematics, 1811: 93–104, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36464-1_13/.

Stevens:2003:SCC

[Ste03m] Jan Stevens. 13. Smoothing components of curves. Lecture Notes in Mathematics, 1811:105–111, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36464-1_14/.

Stevens:2003:KC

[Ste03n] Jan Stevens. 14. Kollár’s conjectures. Lecture Notes in Mathematics, 1811:113–124, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36464-1_15/.

Stevens:2003:CC

[Ste03o] Jan Stevens. 15. Cones over curves. Lecture Notes in Mathematics, 1811:125–136, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36464-1_16/.

Stevens:2003:VDH

[Ste03p] Jan Stevens. 16. The versal deformation of hyperelliptic cones. Lecture Notes in Mathematics, 1811:137–146, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36464-1_
17/.

Stevens:2003:SB

[Ste03q] Jan Stevens. 2. Standard bases. Lecture Notes in Mathematics, 1811:15–22, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36464-1_3/.

Stevens:2003:ID

[Ste03r] Jan Stevens. 3. Infinitesimal deformations. Lecture Notes in Mathematics, 1811:23–31, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36464-1_4/.

Stevens:2003:EFP

[Ste03s] Jan Stevens. 4. Example: the fat point of multiplicity four. Lecture Notes in Mathematics, 1811:33–38, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36464-1_5/.

Stevens:2003:DA

[Ste03t] Jan Stevens. 5. Deformations of algebras. Lecture Notes in Mathematics, 1811:39–44, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36464-1_6/.

Stevens:2003:FDT

[Ste03u] Jan Stevens. 6. Formal deformation theory. Lecture Notes in Mathematics, 1811:45–53, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36464-1_7/.

Stevens:2003:DCM

[Ste03v] Jan Stevens. 7. Deformations of compact manifolds. Lecture Notes in Mathematics, 1811:55–61, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36464-1_8/.

Stevens:2003:HSD

[Ste03w] Jan Stevens. 8. How to solve the deformation equation. Lecture Notes in Mathematics, 1811:63–66, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36464-1_9/.

Stevens:2003:CIS

[Ste03x] Jan Stevens. 9. Convergence for isolated singularities. Lecture Notes in Mathematics, 1811:67–70, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36464-1_10/.

Stevens:2003:DSb

[Ste03y] Jan Stevens. Deformations of Singularities, volume 1811 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 157 pp. LCCN QA3 .L28 no. 1811. URL http:/

/link.springer.com/book/10.1007/b10723;   http://www.
springerlink.com/content/978-3-540-36464-1.

Stevens:2003:Ib

[Ste03z] Jan Stevens. Index. Lecture Notes in Mathematics, 1811:155– 157, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36464-1_19/.

Stevens:2003:Ia

[Ste03-27] Jan Stevens. Introduction. Lecture Notes in Mathematics, 1811: 1–4, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36464-1_1/.

Stevens:2003:R

[Ste03-28] Jan Stevens. References. Lecture Notes in Mathematics, 1811: 147–153, 2003. CODEN LNMAA2. ISBN 3-540-00560-9 (print), 3-540-36464-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36464-1_18/.

Steuding:2007:BM

[Ste07a] Jörn Steuding. Back matter. Lecture Notes in Mathematics, 1877: 285–322, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-44822-8/1.pdf.

Steuding:2007:CU

[Ste07b] Jörn Steuding. Consequences of universality. Lecture Notes in Mathematics, 1877:193–207, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-44822-8_10.pdf.

Steuding:2007:DSPa

[Ste07c] Jörn Steuding. Dirichlet series and polynomial Euler products. Lecture Notes in Mathematics, 1877:35–47, 2007. CO-DEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-44822-8_2.pdf.

Steuding:2007:DSPb

[Ste07d] Jörn Steuding. Dirichlet series with periodic coefficients. Lecture Notes in Mathematics, 1877:209–227, 2007. CO-DEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-44822-8_11.pdf.

Steuding:2007:ER

[Ste07e] Jörn Steuding. Effective results. Lecture Notes in Mathematics, 1877:167–191, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-44822-8_9.pdf.

Steuding:2007:FM

[Ste07f] Jörn Steuding. Front matter. Lecture Notes in Mathematics, 1877: I–XIII, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-44822-8/1.pdf.

Steuding:2007:IRP

[Ste07g] Jörn Steuding. Interlude: Results from probability theory. Lecture Notes in Mathematics, 1877:49–61, 2007. CO-DEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-44822-8_3.pdf.

Steuding:2007:I

[Ste07h] Jörn Steuding. Introduction. Lecture Notes in Mathematics, 1877: 1–33, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-44822-8_1.pdf.

Steuding:2007:JU

[Ste07i] Jörn Steuding. Joint universality. Lecture Notes in Mathematics, 1877:229–248, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-44822-8_12.pdf.

Steuding:2007:FNF

[Ste07j] Jörn Steuding. L-functions of number fields. Lecture Notes in Mathematics, 1877:249–283, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-44822-8_13.pdf.

Steuding:2007:LT

[Ste07k] Jörn Steuding. Limit theorems. Lecture Notes in Mathematics, 1877:63–85, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-44822-8_4.pdf.

Steuding:2007:RH

[Ste07l] Jörn Steuding. The Riemann hypothesis. Lecture Notes in Mathematics, 1877:155–165, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-44822-8_8.pdf.

Steuding:2007:SC

[Ste07m] Jörn Steuding. The Selberg class. Lecture Notes in Mathematics, 1877:111–135, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-44822-8_6.pdf.

Steuding:2007:U

[Ste07n] Jörn Steuding. Universality. Lecture Notes in Mathematics, 1877:87–110, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-44822-8_5.pdf.

Steuding:2007:VDC

[Ste07o] Jörn Steuding. Value-distribution in the complex plane. Lecture Notes in Mathematics, 1877:137–154, 2007. CODEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-44822-8_7.
pdf.

Steuding:2007:VDF

[Ste07p] Jörn Steuding. Value-Distribution of L-Functions, volume 1877 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CO-DEN LNMAA2. ISBN 3-540-26526-0 (print), 3-540-44822-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 317 pp. LCCN QA3 .L28 no. 1877. URL http://link.

springer.com/book/10.1007/978-3-540-44822-8;   http:/
/www.springerlink.com/content/978-3-540-44822-8.

Szarek:2007:DWD

[STJ07] S. J. Szarek and N. Tomczak-Jaegermann. Decoupling weakly dependent events. Lecture Notes in Mathematics, 1910:297–303, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_17.pdf.

Schechtman:2003:MSS

[STJV03] G. Schechtman, N. Tomczak-Jaegermann, and R. Vershynin. Maximal ℓ_{p}^{n}-structures in spaces with extremal parameters. Lecture Notes in Mathematics, 1807:223–240, 2003. CODEN LN-MAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-36428-3_
18/.

Strang:2000:SPE

[Str00] Gilbert Strang. Signal processing for everyone. Lecture Notes in Mathematics, 1739:366–412, 2000. CODEN LNMAA2. ISBN 3-540-67782-8 (print), 3-540-44976-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103924/.

Strasser:2003:NSC

[Str03] Eva Strasser. Necessary and sufficient conditions for the supermartingale property of a stochastic integral with respect to a local martingale. Lecture Notes in Mathematics, 1832:385– 393, 2003. CODEN LNMAA2. ISBN 3-540-20520-9 (print), 3-540-40004-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40004-2_16/.

Sumners:2009:BM

[Sum09a] De Witt Sumners. Back matter. Lecture Notes in Mathematics, 1973:219–231, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-00837-5/1.pdf.

Sumners:2009:RKT

[Sum09b] De Witt Sumners. Random knotting: Theorems, simulations and applications. Lecture Notes in Mathematics, 1973:187–217, 2009. CODEN LNMAA2. ISBN 3-642-00836-4 (print), 3-642-00837-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-00837-5_7.pdf.

Sussmann:2008:GDV

[Sus08] Héctor J. Sussmann. Generalized differentials, variational generators, and the maximum principle with state constraints. Lecture Notes in Mathematics, 1932:221–287, 2008. CO-DEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77653-6_4.pdf.

Szarek:2000:SEP

[SV00] S. J. Szarek and D. Voiculescu. Shannon’s entropy power inequality via restricted Minkowski sums. Lecture Notes in Mathematics, 1745:257–262, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0107219/.

Song:2009:BM

[SV09a] R. Song and Z. Vondraček. Back matter. Lecture Notes in Mathematics, 1980:1–16, 2009. CODEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-02141-1/1.pdf.

Song:2009:PTS

[SV09b] R. Song and Z. Vondraček. Potential theory of subordinate Brownian motion. Lecture Notes in Mathematics, 1980:87–176, 2009. CODEN LNMAA2. ISBN 3-642-02140-9 (print), 3-642-02141-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-02141-1_5.pdf.

Schutt:2003:PVC

[SW03] Carsten Schütt and Elisabeth Werner. Polytopes with vertices chosen randomly from the boundary of a convex body. Lecture Notes in Mathematics, 1807:241–422, 2003. CODEN LNMAA2. ISBN 3-540-00485-8 (print), 3-540-36428-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36428-3_19/.

Seo:2009:MFE

[SW09] Jin Keun Seo and Eung Je Woo. Multi-frequency electrical impedance tomography and magnetic resonance electrical impedance tomography. Lecture Notes in Mathematics, 1983: 1–71, 2009. CODEN LNMAA2. ISBN 3-642-03443-8 (print), 3-642-03444-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-03444-2_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-642-03444-2/1.pdf.

Scheinkman:2004:HBS

[SX04] José Scheinkman and Wei Xiong. Heterogeneous beliefs, speculation and trading in financial markets. Lecture Notes in Mathematics, 1847:217–250, 2004. CODEN LNMAA2. ISBN 3-540-22266-9 (print), 3-540-44468-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-44468-8_3/.

Salminen:2007:TFS

[SY07] Paavo Salminen and Marc Yor. Tanaka formula for symmetric Lévy processes. Lecture Notes in Mathematics, 1899:265–285, 2007. CODEN LNMAA2. ISBN 3-540-71188-0 (print), 3-540-71189-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-71189-6_14.pdf.

Schechtman:2000:CB

[SZ00] G. Schechtman and J. Zinn. Concentration on the ℓ_{p}^{n} ball. Lecture Notes in Mathematics, 1745:245–256, 2000. CODEN LN-MAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107218/.

Taira:2009:AEB

[Tai09a] Kazuaki Taira. L^{p} approach to elliptic boundary value problems. Lecture Notes in Mathematics, 1499:77–85, 2009. CO-DEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01677-6_4.pdf.

Taira:2009:TPD

[Tai09b] Kazuaki Taira. L^{p} theory of pseudo-differential operators. Lecture Notes in Mathematics, 1499:55–75, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-01677-6_3.
pdf.

Taira:2009:ASI

[Tai09c] Kazuaki Taira. Application to semilinear initial-boundary value problems. Lecture Notes in Mathematics, 1499:161–168, 2009.

CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01677-6_10.pdf.

Taira:2009:BM

[Tai09d] Kazuaki Taira. Back matter. Lecture Notes in Mathematics, 1499: 1–17, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-01677-6/1.pdf.

Taira:2009:BVP

[Tai09e] Kazuaki Taira. Boundary Value Problems and Markov Processes, volume 1499 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 186 pp. LCCN QA3 .L28 no. 1499. URL http://link.

springer.com/book/10.1007/978-3-642-01677-6;   http:/
/www.springerlink.com/content/978-3-642-01677-6.

Taira:2009:CR

[Tai09f] Kazuaki Taira. Concluding remarks. Lecture Notes in Mathematics, 1499:169–174, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01677-6_11.pdf.

Taira:2009:FM

[Tai09g] Kazuaki Taira. Front matter. Lecture Notes in Mathematics, 1499:1–9, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-01677-6/1.pdf.

Taira:2009:IMR

[Tai09h] Kazuaki Taira. Introduction and main results. Lecture Notes in Mathematics, 1499:1–12, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01677-6_1.pdf.

Taira:2009:PE

[Tai09i] Kazuaki Taira. A priori estimates. Lecture Notes in Mathematics, 1499:95–100, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01677-6_6.pdf.

Taira:2009:PTa

[Tai09j] Kazuaki Taira. Proof of theorem 1.1. Lecture Notes in Mathematics, 1499:87–93, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01677-6_5.pdf.

Taira:2009:PTb

[Tai09k] Kazuaki Taira. Proof of theorem 1.2. Lecture Notes in Mathematics, 1499:101–111, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01677-6_7.pdf.

Taira:2009:PTPa

[Tai09l] Kazuaki Taira. Proof of Theorem 1.3 — Part (i). Lecture Notes in Mathematics, 1499:113–124, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01677-6_8.pdf.

Taira:2009:PTPb

[Tai09m] Kazuaki Taira. Proof of Theorem 1.3, Part (II). Lecture Notes in Mathematics, 1499:125–159, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01677-6_9.pdf.

Taira:2009:ST

[Tai09n] Kazuaki Taira. Semigroup theory. Lecture Notes in Mathematics, 1499:13–54, 2009. CODEN LNMAA2. ISBN 3-642-01676-6 (print), 3-642-01677-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01677-6_2.pdf.

Takamura:2006:AFSa

[Tak06a] Shigeru Takamura. Arrangement functions and singularities, I. Lecture Notes in Mathematics, 1886:421–438, 2006. CO-DEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_22.pdf.

Takamura:2006:AFSb

[Tak06b] Shigeru Takamura. Arrangement functions and singularities, II. Lecture Notes in Mathematics, 1886:439–459, 2006. CO-DEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_23.pdf.

Takamura:2006:BM

[Tak06c] Shigeru Takamura. Back matter. Lecture Notes in Mathematics, 1886:581–594, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-33364-7/1.pdf.

Takamura:2006:CT

[Tak06d] Shigeru Takamura. Classification theorem. Lecture Notes in Mathematics, 1886:483–485, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33364-7_25.pdf.

Takamura:2006:CBS

[Tak06e] Shigeru Takamura. Compound barking (stellar case). Lecture Notes in Mathematics, 1886:303–307, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-33364-7_17.
pdf.

Takamura:2006:CBDb

[Tak06f] Shigeru Takamura. Construction of barking deformations (constellar case). Lecture Notes in Mathematics, 1886:327–347, 2006.

CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_19.pdf.

Takamura:2006:CBDa

[Tak06g] Shigeru Takamura. Construction of barking deformations (stellar case). Lecture Notes in Mathematics, 1886:265–278, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_15.pdf.

Takamura:2006:CDB

[Tak06h] Shigeru Takamura. Construction of deformations by bunches of subbranches. Lecture Notes in Mathematics, 1886:253–262, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_14.pdf.

Takamura:2006:CDTa

[Tak06i] Shigeru Takamura. Construction of deformations by tame subbranches. Lecture Notes in Mathematics, 1886:99–117, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_6.pdf.

Takamura:2006:CDW

[Tak06j] Shigeru Takamura. Construction of deformations by wild subbranches. Lecture Notes in Mathematics, 1886:143–152, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_8.pdf.

Takamura:2006:CDTb

[Tak06k] Shigeru Takamura. Construction of deformations of type A_{l}. Lecture Notes in Mathematics, 1886:119–141, 2006. CO-DEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_7.pdf.

Takamura:2006:CDTc

[Tak06l] Shigeru Takamura. Construction of deformations of type B_{l}. Lecture Notes in Mathematics, 1886:177–181, 2006. CO-DEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_10.pdf.

Takamura:2006:CDTd

[Tak06m] Shigeru Takamura. Construction of deformations of type C_{l}. Lecture Notes in Mathematics, 1886:183–207, 2006. CO-DEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_11.pdf.

Takamura:2006:DTNa

[Tak06n] Shigeru Takamura. Deformations of tubular neighborhoods of branches (preparation). Lecture Notes in Mathematics, 1886:85– 98, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_5.pdf.

Takamura:2006:DTNb

[Tak06o] Shigeru Takamura. Deformations of tubular neighborhoods of trunks. Lecture Notes in Mathematics, 1886:309–326, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_18.pdf.

Takamura:2006:FMa

[Tak06p] Shigeru Takamura. Front matter. Lecture Notes in Mathematics, 1886:i–xxix, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-33364-7/1.pdf.

Takamura:2006:FMb

[Tak06q] Shigeru Takamura. Front matter. Lecture Notes in Mathematics, 1886:22, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33364-7/1/1.pdf.

Takamura:2006:FMc

[Tak06r] Shigeru Takamura. Front matter. Lecture Notes in Mathematics, 1886:84, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33364-7/2/1.pdf.

Takamura:2006:FMd

[Tak06s] Shigeru Takamura. Front matter. Lecture Notes in Mathematics, 1886:264, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33364-7/3/1.pdf.

Takamura:2006:FMe

[Tak06t] Shigeru Takamura. Front matter. Lecture Notes in Mathematics, 1886:382, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33364-7/4/1.pdf.

Takamura:2006:FMf

[Tak06u] Shigeru Takamura. Front matter. Lecture Notes in Mathematics, 1886:482, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33364-7/5/1.pdf.

Takamura:2006:FE

[Tak06v] Shigeru Takamura. Further examples. Lecture Notes in Mathematics, 1886:349–379, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33364-7_20.pdf.

Takamura:2006:GBD

[Tak06w] Shigeru Takamura. Global barking deformations: Ideas and examples. Lecture Notes in Mathematics, 1886:57–81, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_4.pdf.

Takamura:2006:LWC

[Tak06x] Shigeru Takamura. List of weighted crustal sets for singular fibers of genus ≤ 5. Lecture Notes in Mathematics, 1886:487–580, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_26.pdf.

Takamura:2006:RCD

[Tak06y] Shigeru Takamura. Recursive construction of deformations of type C_{l}. Lecture Notes in Mathematics, 1886:209–234, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_12.pdf.

Takamura:2006:SLB

[Tak06z] Shigeru Takamura. Semi-local barking deformations: Ideas and examples. Lecture Notes in Mathematics, 1886:41–56, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_3.pdf.

Takamura:2006:SCS

[Tak06-27] Shigeru Takamura. Simple crusts (stellar case). Lecture Notes in Mathematics, 1886:279–302, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33364-7_16.pdf.

Takamura:2006:SFA

[Tak06-28] Shigeru Takamura. Singularities of fibers around cores. Lecture Notes in Mathematics, 1886:383–419, 2006. CO-DEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_21.pdf.

Takamura:2006:SDDa

[Tak06-29] Shigeru Takamura. Splitting deformations of degenerations. Lecture Notes in Mathematics, 1886:23–31, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-33364-7_1.
pdf.

Takamura:2006:SDDb

[Tak06-30] Shigeru Takamura. Splitting Deformations of Degenerations of Complex Curves: Towards the Classification of Atoms of Degenerations, III, volume 1886 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (paperback), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 590 pp. LCCN QA3 .L28 no. 1886;

QA565.T36 2006. URL http://link.springer.com/book/10.
1007/978-3-540-33364-7.

Takamura:2006:STB

[Tak06-31] Shigeru Takamura. Subbranches of types A_{l}, B_{l}, C_{l}. Lecture Notes in Mathematics, 1886:153–176, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33364-7_9.pdf.

Takamura:2006:S

[Tak06-32] Shigeru Takamura. Supplement. Lecture Notes in Mathematics, 1886:461–479, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33364-7_24.pdf.

Takamura:2006:TBC

[Tak06-33] Shigeru Takamura. Types A_{l}, B_{l}, and C_{l} exhaust all cases. Lecture Notes in Mathematics, 1886:235–251, 2006. CO-DEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33364-7_13.pdf.

Takamura:2006:WB

[Tak06-34] Shigeru Takamura. What is a barking? Lecture Notes in Mathematics, 1886:33–39, 2006. CODEN LNMAA2. ISBN 3-540-33363-0 (print), 3-540-33364-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33364-7_2.pdf.

Talagrand:2007:MFM

[Tal07] Michel Talagrand. Mean field models for spin glasses: Some obnoxious problems. Lecture Notes in Mathematics, 1900:63–80, 2007. CODEN LNMAA2. ISBN 3-540-40902-5 (print), 3-540-40908-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-40908-3_3.pdf.

Tartar:2000:IHM

[Tar00] Luc Tartar. An introduction to the homogenization method in optimal design. Lecture Notes in Mathematics, 1740:47–156, 2000. CODEN LNMAA2. ISBN 3-540-67971-5 (print), 3-540-44486-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0106742/.

Tarasov:2002:HFR

[Tar02] Vitaly Tarasov. q-Hypergeometric functions and representation theory. Lecture Notes in Mathematics, 1776:193–267, 2002. CODEN LNMAA2. ISBN 3-540-43121-7 (print), 3-540-45617-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45617-9_4/.

Tarantello:2009:SEP

[Tar09] Gabriella Tarantello. On some elliptic problems in the study of selfdual Chern–Simons vortices. Lecture Notes in Mathematics, 1977:117–175, 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-01674-5_4.pdf.

Tavare:2004:PAI

[Tav04] Simon Tavaré. Part I: Ancestral inference in population genetics. Lecture Notes in Mathematics, 1837:1–188, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39874-5_
1/.

Tarabusi:2008:RTC

[TDP08] Enrico Casadio Tarabusi, Andrea D’Agnolo, and Massimo Picardello, editors. Representation Theory and Complex Analysis: Lectures given at the C.I.M.E. Summer School held in Venice, Italy June 10–17, 2004, volume 1931 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1931. URL http:

//link.springer.com/book/10.1007/978-3-540-76892-0;
http://www.springerlink.com/content/978-3-540-76892-0.

Telcs:2006:ARW

[Tel06a] András Telcs. The Art of Random Walks, volume 1885 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 195 pp. LCCN QA3 .L28 no. 1885. URL http://

link.springer.com/book/10.1007/b134090;   http://www.
springerlink.com/content/978-3-540-33028-8.

Telcs:2006:BM

[Tel06b] András Telcs. Back matter. Lecture Notes in Mathematics, 1885: 187–199, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-33028-8/1.pdf.

Telcs:2006:BDP

[Tel06c] András Telcs. Basic definitions and preliminaries. Lecture Notes in Mathematics, 1885:7–21, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33028-8_2.pdf.

Telcs:2006:CR

[Tel06d] András Telcs. Closing remarks. Lecture Notes in Mathematics, 1885:165–168, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33028-8_11.pdf.

Telcs:2006:ER

[Tel06e] András Telcs. Einstein relation. Lecture Notes in Mathematics, 1885:83–93, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33028-8_7.pdf.

Telcs:2006:FMa

[Tel06f] András Telcs. Front matter. Lecture Notes in Mathematics, 1885: i–vii, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33028-8/1.pdf.

Telcs:2006:FMb

[Tel06g] András Telcs. Front matter. Lecture Notes in Mathematics, 1885: 24, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33028-8/1/1.pdf.

Telcs:2006:FMc

[Tel06h] András Telcs. Front matter. Lecture Notes in Mathematics, 1885: 70, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-33028-8/2/1.pdf.

Telcs:2006:I

[Tel06i] András Telcs. Introduction. Lecture Notes in Mathematics, 1885: 1–6, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-33028-8_1.pdf.

Telcs:2006:II

[Tel06j] András Telcs. Isoperimetric inequalities. Lecture Notes in Mathematics, 1885:49–60, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33028-8_4.pdf.

Telcs:2006:LE

[Tel06k] András Telcs. Lower estimates. Lecture Notes in Mathematics, 1885:131–151, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33028-8_9.pdf.

Telcs:2006:MLA

[Tel06l] András Telcs. Motivation of the local approach. Lecture Notes in Mathematics, 1885:71–81, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33028-8_6.pdf.

Telcs:2006:PHI

[Tel06m] András Telcs. Parabolic Harnack inequality. Lecture Notes in Mathematics, 1885:169–179, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33028-8_12.pdf.

Telcs:2006:PVG

[Tel06n] András Telcs. Polynomial volume growth. Lecture Notes in Mathematics, 1885:61–67, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33028-8_5.pdf.

Telcs:2006:SLT

[Tel06o] András Telcs. Semi-local theory. Lecture Notes in Mathematics, 1885:181–185, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33028-8_13.pdf.

Telcs:2006:SEP

[Tel06p] András Telcs. Some elements of potential theory. Lecture Notes in Mathematics, 1885:25–47, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-33028-8_3.pdf.

Telcs:2006:TSE

[Tel06q] András Telcs. Two-sided estimates. Lecture Notes in Mathematics, 1885:153–163, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33028-8_10.pdf.

Telcs:2006:UE

[Tel06r] András Telcs. Upper estimates. Lecture Notes in Mathematics, 1885:95–129, 2006. CODEN LNMAA2. ISBN 3-540-33027-5 (print), 3-540-33028-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-33028-8_8.pdf.

Terman:2005:IDS

[Ter05] David Terman. An introduction to dynamical systems and neuronal dynamics. Lecture Notes in Mathematics, 1860:21–68, 2005. CODEN LNMAA2. ISBN 3-540-23858-1 (print), 3-540-31544-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31544-5_2.pdf.

Terwilliger:2006:AAA

[Ter06] Paul Terwilliger. An algebraic approach to the Askey scheme of orthogonal polynomials. Lecture Notes in Mathematics, 1883:255– 330, 2006. CODEN LNMAA2. ISBN 3-540-31062-2 (print), 3-540-36716-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-36716-1_6.pdf.

Teufel:2003:I

[Teu03a] Stefan Teufel. 1. Introduction. Lecture Notes in Mathematics, 1821:1–31, 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45171-6_1/.

Teufel:2003:FOA

[Teu03b] Stefan Teufel. 2. First order adiabatic theory. Lecture Notes in Mathematics, 1821:33–69, 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45171-6_2/.

Teufel:2003:SAP

[Teu03c] Stefan Teufel. 3. Space-adiabatic perturbation theory. Lecture Notes in Mathematics, 1821:71–104, 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45171-6_3/.

Teufel:2003:AE

[Teu03d] Stefan Teufel. 4. Applications and extensions. Lecture Notes in Mathematics, 1821:105–140, 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45171-6_4/.

Teufel:2003:QDP

[Teu03e] Stefan Teufel. 5. Quantum dynamics in periodic media. Lecture Notes in Mathematics, 1821:141–171, 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45171-6_5/.

Teufel:2003:ADS

[Teu03f] Stefan Teufel. 6. Adiabatic decoupling without spectral gap. Lecture Notes in Mathematics, 1821:173–201, 2003. CODEN LN-MAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45171-6_
6/.

Teufel:2003:APT

[Teu03g] Stefan Teufel. Adiabatic Perturbation Theory in Quantum Dynamics, volume 1821 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 236 pp. LCCN QA3 .L28 no. 1821. URL http:/

/link.springer.com/book/10.1007/b13355;   http://www.
springerlink.com/content/978-3-540-45171-6.

Teufel:2003:A

[Teu03h] Stefan Teufel. Appendix. Lecture Notes in Mathematics, 1821: 203–224, 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45171-6_7/.

Teufel:2003:LSR

[Teu03i] Stefan Teufel. List of symbols and references. Lecture Notes in Mathematics, 1821:225–234, 2003. CODEN LNMAA2. ISBN 3-540-40723-5 (print), 3-540-45171-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45171-6_8/.

Tanter:2009:TRW

[TF09] Mickael Tanter and Mathias Fink. Time reversing waves for biomedical applications. Lecture Notes in Mathematics, 1983:73– 97, 2009. CODEN LNMAA2. ISBN 3-642-03443-8 (print), 3-642-03444-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-03444-2_2.pdf.

Thaddeus:2008:FCG

[Tha08] M. Thaddeus. Floer cohomology with Gerbes. Lecture Notes in Mathematics, 1947:105–141, 2008. CODEN LNMAA2. ISBN 3-540-79813-7 (print), 3-540-79814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-79814-9_3.pdf.

Tian:2002:CSI

[Tia02] Gang Tian. Constructing symplectic invariants. Lecture Notes in Mathematics, 1776:269–311, 2002. CODEN LNMAA2. ISBN 3-540-43121-7 (print), 3-540-45617-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45617-9_5/.

Tian:2008:BM

[Tia08a] Jianjun Paul Tian. Back matter. Lecture Notes in Mathematics, 1921:119–129, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-74284-5/1.pdf.

Tian:2008:EA

[Tia08b] Jianjun Paul Tian. Evolution algebras. Lecture Notes in Mathematics, 1921:17–52, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74284-5_3.pdf.

Tian:2008:EAM

[Tia08c] Jianjun Paul Tian. Evolution algebras and Markov chains. Lecture Notes in Mathematics, 1921:53–90, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74284-5_4.
pdf.

Tian:2008:EAN

[Tia08d] Jianjun Paul Tian. Evolution algebras and non-Mendelian genetics. Lecture Notes in Mathematics, 1921:91–107, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-74284-5_5.pdf.

Tian:2008:EAT

[Tia08e] Jianjun Paul Tian. Evolution Algebras and their Applications, volume 1921 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 125 pp. LCCN QA3 .L28 no. 1921. URL http://link.

springer.com/book/10.1007/978-3-540-74284-5;   http:/
/www.springerlink.com/content/978-3-540-74284-5.

Tian:2008:FM

[Tia08f] Jianjun Paul Tian. Front matter. Lecture Notes in Mathematics, 1921:I–XI, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-74284-5/1.pdf.

Tian:2008:FRR

[Tia08g] Jianjun Paul Tian. Further results and research topics. Lecture Notes in Mathematics, 1921:109–118, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74284-5_6.
pdf.

Tian:2008:I

[Tia08h] Jianjun Paul Tian. Introduction. Lecture Notes in Mathematics, 1921:1–7, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74284-5_1.pdf.

Tian:2008:M

[Tia08i] Jianjun Paul Tian. Motivations. Lecture Notes in Mathematics, 1921:9–16, 2008. CODEN LNMAA2. ISBN 3-540-74283-2 (print), 3-540-74284-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74284-5_2.pdf.

Tignol:2004:GMA

[Tig04] Jean-Pierre Tignol, editor. Geometric Methods in the Algebraic Theory of Quadratic Forms: Summer School, Lens, 2000, volume 1835 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1835. URL http:/

/link.springer.com/book/10.1007/b94827;   http://www.
springerlink.com/content/978-3-540-40990-8.

Toft:2008:BM

[Tof08a] J. Toft. Back matter. Lecture Notes in Mathematics, 1949:203– 204, 2008. CODEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-68268-4/1.pdf.

Toft:2008:SPP

[Tof08b] J. Toft. Schatten properties for pseudo-differential operators on modulation spaces. Lecture Notes in Mathematics, 1949:175–202, 2008. CODEN LNMAA2. ISBN 3-540-68266-X (print), 3-540-68268-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-68268-4_5.pdf.

Toninelli:2009:LTD

[Ton09] Fabio Lucio Toninelli. Localization transition in disordered pinning models. Lecture Notes in Mathematics, 1970:1–48, 2009. CODEN LNMAA2. ISBN 3-540-92795-6 (print), 3-540-92796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-92796-9_3.pdf.

Toscani:2008:HDB

[Tos08] Giuseppe Toscani. Hydrodynamics from the dissipative Boltzmann equation. Lecture Notes in Mathematics, 1937:59–75, 2008. CODEN LNMAA2. ISBN 3-540-78276-1 (print), 3-540-78277-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78277-3_3.pdf.

Tsirelson:2004:PSL

[Tsi04] Boris Tsirelson. Part I: Scaling limit, noise, stability. Lecture Notes in Mathematics, 1840:1–106, 2004. CODEN LNMAA2. ISBN 3-540-21316-3 (print), 3-540-39982-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-39982-7_1/.

Tumanov:2004:EDG

[Tum04] Alexander Tumanov. Extremal discs and the geometry of CR manifolds. Lecture Notes in Mathematics, 1848:191–212, 2004. CODEN LNMAA2. ISBN 3-540-22358-4 (print), 3-540-44487-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44487-9_5/.

Turesson:2000:APT

[Tur00a] Bengt Ove Turesson. Applications of potential theory to Sobolev spaces. Lecture Notes in Mathematics, 1736:141–162, 2000. CO-DEN LNMAA2. ISBN 3-540-67588-4 (print), 3-540-45168-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0103912/.

Turesson:2000:NPT

[Tur00b] Bengt Ove Turesson. Nonlinear Potential Theory and Weighted Sobolev Spaces, volume 1736 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67588-4 (print), 3-540-45168-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 173 pp. LCCN QA3 .L28 no. 1736.

URL http://link.springer.com/book/10.1007/BFb0103908;
 http://www.springerlink.com/content/978-3-540-45168-
6.

Turesson:2000:PT

[Tur00c] Bengt Ove Turesson. Potential theory. Lecture Notes in Mathematics, 1736:69–140, 2000. CODEN LNMAA2. ISBN 3-540-67588-4 (print), 3-540-45168-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0103911/.

Turesson:2000:P

[Tur00d] Bengt Ove Turesson. Preliminaries. Lecture Notes in Mathematics, 1736:1–14, 2000. CODEN LNMAA2. ISBN 3-540-67588-4 (print), 3-540-45168-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103909/.

Turesson:2000:SS

[Tur00e] Bengt Ove Turesson. Sobolev spaces. Lecture Notes in Mathematics, 1736:15–68, 2000. CODEN LNMAA2. ISBN 3-540-67588-4 (print), 3-540-45168-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103910/.

Tsirelson:2004:LPT

[TW04] Boris Tsirelson and Wendelin Werner, editors. Lectures on probability theory and statistics: École d’Été de Probabilités de Saint-Flour XXXII — 2002, volume 1840 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-21316-3 (paperback), 3-540-39982-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1840; QA273.A1 T78 2004; QA3 .L28 no.1840. URL http:/

/link.springer.com/book/10.1007/b96719;   http://www.
loc.gov/catdir/enhancements/fy0813/2006299567-d.html;

 http://www.loc.gov/catdir/enhancements/fy0813/2006299567-
t.html;   http://www.springerlink.com/content/978-3-540-
39982-7;   http://www.springerlink.com/openurl.asp?genre=
issue&eissn=0075-8434&volume=1840.

Tavare:2004:LPT

[TZP04] Simon Tavaré, Ofer Zeitouni, and Jean Picard, editors. Lectures on Probability Theory and Statistics: École d’Été de Probabilités de Saint-Flour XXXI — 2001, volume 1837 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNMAA2. ISBN 3-540-20832-1 (paperback), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1837; QH455 .T37 2001. URL http:/

/link.springer.com/book/10.1007/b95197;   http://www.
loc.gov/catdir/enhancements/fy0813/2007279726-d.html;
 http://www.loc.gov/catdir/toc/fy0802/2007279726.html;
 http://www.springerlink.com/content/978-3-540-39874-
5;   http://www.springerlink.com/openurl.asp?genre=issue&
eissn=0075-8434&volume=1837.

Ukai:2006:AAF

[Uka06a] Seiji Ukai. Asymptotic analysis of fluid equations. Lecture Notes in Mathematics, 1871:189–250, 2006. CODEN LNMAA2. ISBN 3-540-28586-5 (print), 3-540-32454-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/11545989_5.pdf.

Ukai:2006:BM

[Uka06b] Seiji Ukai. Back matter. Lecture Notes in Mathematics, 1871: ??, 2006. CODEN LNMAA2. ISBN 3-540-28586-5 (print), 3-540-32454-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-32454-6/1.pdf.

Unterberger:2000:ALC

[Unt00a] André Unterberger. Another look at the composition of Weyl symbols. Lecture Notes in Mathematics, 1742:33–44, 2000. CO-DEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0104041/.

Unterberger:2000:AD

[Unt00b] André Unterberger. Automorphic distributions on R^{2}. Lecture Notes in Mathematics, 1742:119–130, 2000. CODEN LN-MAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104049/.

Unterberger:2000:AFA

[Unt00c] André Unterberger. Automorphic functions associated with quadratic PSL(2, Z)-orbits in P_{1}(R). Lecture Notes in Mathematics, 1742:213–230, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0104055/.

Unterberger:2000:DKS

[Unt00d] André Unterberger. A digression on Kloosterman sums. Lecture Notes in Mathematics, 1742:91–96, 2000. CODEN LN-MAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104046/.

Unterberger:2000:DAN

[Unt00e] André Unterberger. Distributions associated with the non-unitary principal series. Lecture Notes in Mathematics, 1742:11–15, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0104038/.

Unterberger:2000:EPB

[Unt00f] André Unterberger. The expansion of the Poisson bracket of two Eisenstein series. Lecture Notes in Mathematics, 1742:111– 117, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104048/.

Unterberger:2000:GSS

[Unt00g] André Unterberger. A generating series of sorts for Maass cuspforms. Lecture Notes in Mathematics, 1742:149–161, 2000. CO-DEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0104051/.

Unterberger:2000:HDP

[Unt00h] André Unterberger. The Hecke decomposition of products or Poisson brackets of two Eisenstein series. Lecture Notes in Mathematics, 1742:131–147, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0104050/.

Unterberger:2000:I

[Unt00i] André Unterberger. Introduction. Lecture Notes in Mathematics, 1742:1–9, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104037/.

Unterberger:2000:MD

[Unt00j] André Unterberger. Modular distributions. Lecture Notes in Mathematics, 1742:17–23, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104039/.

Unterberger:2000:MFL

[Unt00k] André Unterberger. Moving to the forward light-cone: the Lax– Phillips theory revisited. Lecture Notes in Mathematics, 1742: 191–212, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104054/.

Unterberger:2000:PSR

[Unt00l] André Unterberger. The principal series of SL(2, R) and the Radon transform. Lecture Notes in Mathematics, 1742:25–31, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104040/.

Unterberger:2000:PTE

[Unt00m] André Unterberger. The “product” of two Eisenstein distributions. Lecture Notes in Mathematics, 1742:69–75, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104044/.

Unterberger:2000:QOD

[Unt00n] André Unterberger. Quadratic orbits: a dual problem. Lecture Notes in Mathematics, 1742:231–246, 2000. CODEN LN-MAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104056/.

Unterberger:2000:QNH

[Unt00o] André Unterberger. Quantization and Non-holomorphic Modular Forms, volume 1742 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 253 pp. LCCN QA3 .L28 no. 1742. URL http:/

/link.springer.com/book/10.1007/BFb0104036;   http://
www.springerlink.com/content/978-3-540-44660-6.

Unterberger:2000:QPP

[Unt00p] André Unterberger. Quantization, products and Poisson brackets. Lecture Notes in Mathematics, 1742:177–190, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0104053/.

Unterberger:2000:RRS

[Unt00q] André Unterberger. Recovering the Roelcke–Selberg coefficients of a function in L^{2}(Γ \ Π). Lecture Notes in Mathematics, 1742: 61–68, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104043/.

Unterberger:2000:RSD

[Unt00r] André Unterberger. The Roelcke–Selberg decomposition and the Radon transform. Lecture Notes in Mathematics, 1742:45– 59, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0104042/.

Unterberger:2000:RSEa

[Unt00s] André Unterberger. The Roelcke–Selberg expansion of the product of two Eisenstein series: the continuous part. Lecture Notes in Mathematics, 1742:77–90, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104045/.

Unterberger:2000:RSEb

[Unt00t] André Unterberger. The Roelcke–Selberg expansion of the product of two Eisenstein series: the discrete part. Lecture Notes in Mathematics, 1742:97–109, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104047/.

Unterberger:2000:SAD

[Unt00u] André Unterberger. Some arithmetic distributions. Lecture Notes in Mathematics, 1742:163–176, 2000. CODEN LNMAA2. ISBN 3-540-67861-1 (print), 3-540-44660-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104052/.

Unterberger:2008:APA

[Unt08a] André Unterberger. Alternative Pseudodifferential Analysis: With an Application to Modular Forms, volume 1935 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 118 pp. LCCN

QA3 .L28 no. 1935. URL http://link.springer.com/book/
10.1007/978-3-540-77911-7;   http://www.springerlink.
com/content/978-3-540-77911-7.

Unterberger:2008:BM

[Unt08b] André Unterberger. Back matter. Lecture Notes in Mathematics, 1935:115–122, 2008. CODEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-77911-7/1.pdf.

Unterberger:2008:AAU

[Unt08c] André Unterberger. From anaplectic analysis to usual analysis. Lecture Notes in Mathematics, 1935:75–91, 2008. CO-DEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77911-7_4.pdf.

Unterberger:2008:FM

[Unt08d] André Unterberger. Front matter. Lecture Notes in Mathematics, 1935:i–ix, 2008. CODEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-77911-7/1.pdf.

Unterberger:2008:I

[Unt08e] André Unterberger. Introduction. Lecture Notes in Mathematics, 1935:1–9, 2008. CODEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-77911-7_1.pdf.

Unterberger:2008:MAR

[Unt08f] André Unterberger. The metaplectic and anaplectic representations. Lecture Notes in Mathematics, 1935:11–26, 2008. CODEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77911-7_2.pdf.

Unterberger:2008:ODA

[Unt08g] André Unterberger. The one-dimensional alternative pseudodifferential analysis. Lecture Notes in Mathematics, 1935:27–74, 2008. CODEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77911-7_3.pdf.

Unterberger:2008:PAM

[Unt08h] André Unterberger. Pseudodifferential analysis and modular forms. Lecture Notes in Mathematics, 1935:93–114, 2008.

CODEN LNMAA2. ISBN 3-540-77910-8 (print), 3-540-77911-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77911-7_5.pdf.

Urbano:2008:BM

[Urb08a] José Miguel Urbano. Back matter. Lecture Notes in Mathematics, 1930:145–150, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-75932-4/1.pdf.

Urbano:2008:FPM

[Urb08b] José Miguel Urbano. Flows in porous media: The variable exponent case. Lecture Notes in Mathematics, 1930:87–105, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75932-4_6.pdf.

Urbano:2008:FM

[Urb08c] José Miguel Urbano. Front matter. Lecture Notes in Mathematics, 1930:i–x, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-75932-4/1.pdf.

Urbano:2008:GSA

[Urb08d] José Miguel Urbano. The geometric setting and an alternative. Lecture Notes in Mathematics, 1930:21–34, 2008. CO-DEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75932-4_3.pdf.

Urbano:2008:IFC

[Urb08e] José Miguel Urbano. Immiscible fluids and chemotaxis. Lecture Notes in Mathematics, 1930:51–86, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-75932-4_5.
pdf.

Urbano:2008:I

[Urb08f] José Miguel Urbano. Introduction. Lecture Notes in Mathematics, 1930:1–8, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-75932-4_1.pdf.

Urbano:2008:MIS

[Urb08g] José Miguel Urbano. The Method of Intrinsic Scaling: A Systematic Approach to Regularity for Degenerate and Singular PDEs, volume 1930 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 150 pp. LCCN QA3 .L28 no. 1930. URL http://link.

springer.com/book/10.1007/978-3-540-75932-4;   http:/
/www.springerlink.com/content/978-3-540-75932-4.

Urbano:2008:PTD

[Urb08h] José Miguel Urbano. Phase transitions: The doubly singular Stefan problem. Lecture Notes in Mathematics, 1930:107–143, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-75932-4_7.pdf.

Urbano:2008:THC

[Urb08i] José Miguel Urbano. Towards the hölder continuity. Lecture Notes in Mathematics, 1930:35–48, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-75932-4_4.pdf.

Urbano:2008:WSP

[Urb08j] José Miguel Urbano. Weak solutions and a priori estimates. Lecture Notes in Mathematics, 1930:11–19, 2008. CODEN LNMAA2. ISBN 3-540-75931-X (print), 3-540-75932-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-75932-4_2.
pdf.

Utkin:2008:BM

[Utk08a] Vadim I. Utkin. Back matter. Lecture Notes in Mathematics, 1932:349–351, 2008. CODEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-77653-6/1.pdf.

Utkin:2008:SMC

[Utk08b] Vadim I. Utkin. Sliding mode control: Mathematical tools, design and applications. Lecture Notes in Mathematics, 1932:289–347, 2008. CODEN LNMAA2. ISBN 3-540-77644-3 (print), 3-540-77653-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77653-6_5.pdf.

Vakil:2008:BM

[Vak08a] R. Vakil. Back matter. Lecture Notes in Mathematics, 1947: 199–210, 2008. CODEN LNMAA2. ISBN 3-540-79813-7 (print), 3-540-79814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-79814-9/1.pdf.

Vakil:2008:MSC

[Vak08b] R. Vakil. The moduli space of curves and Gromov–Witten theory. Lecture Notes in Mathematics, 1947:143–198, 2008. CODEN LNMAA2. ISBN 3-540-79813-7 (print), 3-540-79814-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-79814-9_4.pdf.

Valette:2008:AMS

[Val08] Alain Valette. Amenability and Margulis super-rigidity. Lecture Notes in Mathematics, 1931:235–258, 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-76892-0_4.
pdf.

Vanhaecke:2001:ACI

[Van01a] Pol Vanhaecke. Algebraic completely integrable Hamiltonian systems. Lecture Notes in Mathematics, 1638:127–142, 2001. CO-DEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/3-540-44576-
5_5/.

Vanhaecke:2001:IHSb

[Van01b] Pol Vanhaecke. Integrable Hamiltonian systems and symmetric products of curves. Lecture Notes in Mathematics, 1638:71– 96, 2001. CODEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44576-5_3/.

Vanhaecke:2001:IHSa

[Van01c] Pol Vanhaecke. Integrable Hamiltonian systems on affine Poisson varieties. Lecture Notes in Mathematics, 1638:17–70, 2001. CODEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44576-5_2/.

Vanhaecke:2001:ISR

[Van01d] Pol Vanhaecke. Integrable Systems in the realm of Algebraic Geometry, volume 1638 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 256 pp. LCCN QA3 .L28 no. 1638. URL http://

link.springer.com/book/10.1007/3-540-44576-5;   http:
//www.springerlink.com/content/978-3-540-44576-0.

Vanhaecke:2001:IGA

[Van01e] Pol Vanhaecke. Interludium: the geometry of Abelian varieties. Lecture Notes in Mathematics, 1638:97–125, 2001. CODEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44576-5_4/.

Vanhaecke:2001:I

[Van01f] Pol Vanhaecke. Introduction. Lecture Notes in Mathematics, 1638:1–16, 2001. CODEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44576-5_1/.

Vanhaecke:2001:MS

[Van01g] Pol Vanhaecke. The Mumford systems. Lecture Notes in Mathematics, 1638:143–173, 2001. CODEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-44576-5_6/.

Vanhaecke:2001:TDC

[Van01h] Pol Vanhaecke. Two-dimensional a.c.i. systems and applications. Lecture Notes in Mathematics, 1638:175–241, 2001. CODEN LNMAA2. ISBN 3-540-42337-0 (print), 3-540-44576-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-44576-5_7/.

VanderJeugt:2003:COP

[Van03] Joris Van der Jeugt. 3nj-Coefficients and orthogonal polynomials of hypergeometric type. Lecture Notes in Mathematics, 1817: 25–92, 2003. CODEN LNMAA2. ISBN 3-540-40375-2 (print), 3-540-44945-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44945-0_2/.

Vasilev:2002:I

[Vas02a] Alexander Vasil’ev. 1. Introduction. Lecture Notes in Mathematics, 1788:1–5, 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45437-3_1/.

Vasilev:2002:MFCa

[Vas02b] Alexander Vasil’ev. 2. Moduli of families of curves and extremal partitions. Lecture Notes in Mathematics, 1788:7–55, 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45437-3_2/.

Vasilev:2002:MEPa

[Vas02c] Alexander Vasil’ev. 3. Moduli in extremal problems for conformal mapping. Lecture Notes in Mathematics, 1788:57–139, 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45437-3_3/.

Vasilev:2002:MEPb

[Vas02d] Alexander Vasil’ev. 4. Moduli in extremal problems for quasiconformal mapping. Lecture Notes in Mathematics, 1788:141– 174, 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45437-3_4/.

Vasilev:2002:MTS

[Vas02e] Alexander Vasil’ev. 5. Moduli on Teichmüller spaces. Lecture Notes in Mathematics, 1788:175–196, 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45437-3_5/.

Vasilev:2002:LSI

[Vas02f] Alexander Vasil’ev. List of symbols and index. Lecture Notes in Mathematics, 1788:207–211, 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45437-3_7/.

Vasilev:2002:MFCb

[Vas02g] Alexander Vasil’ev. Moduli of Families of Curves for Conformal and Quasiconformal Mappings, volume 1788 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 211 pp. LCCN QA3 .L28 no. 1788. URL http:/

/link.springer.com/book/10.1007/b83857;   http://www.
springerlink.com/content/978-3-540-45437-3.

Vasilev:2002:R

[Vas02h] Alexander Vasil’ev. References. Lecture Notes in Mathematics, 1788:197–206, 2002. CODEN LNMAA2. ISBN 3-540-43846-7 (print), 3-540-45437-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45437-3_6/.

vandenDriessche:2008:DCM

[vdD08a] P. van den Driessche. Deterministic compartmental models: Extensions of basic models. Lecture Notes in Mathematics, 1945: 147–157, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-78911-6_5.pdf.

vandenDriessche:2008:SSP

[vdD08b] P. van den Driessche. Spatial structure: Patch models. Lecture Notes in Mathematics, 1945:179–189, 2008. CO-DEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_7.pdf.

vandenDriessche:2008:FNB

[vdDW08] P. van den Driessche and James Watmough. Further notes on the basic reproduction number. Lecture Notes in Mathematics, 1945:159–178, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-78911-6_6.pdf.

vanderHoeven:2006:ADE

[vdH06a] Joris van der Hoeven. Algebraic differential equations. Lecture Notes in Mathematics, 1888:165–200, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35591-5_8.
pdf.

vanderHoeven:2006:BM

[vdH06b] Joris van der Hoeven. Back matter. Lecture Notes in Mathematics, 1888:235–259, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-35591-5/1.pdf.

vanderHoeven:2006:FM

[vdH06c] Joris van der Hoeven. Front matter. Lecture Notes in Mathematics, 1888:I–XXII, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-35591-5/1.pdf.

vanderHoeven:2006:GBO

[vdH06d] Joris van der Hoeven. Grid-based operators. Lecture Notes in Mathematics, 1888:115–133, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35591-5_6.pdf.

vanderHoeven:2006:GBS

[vdH06e] Joris van der Hoeven. Grid-based series. Lecture Notes in Mathematics, 1888:33–55, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35591-5_2.pdf.

vanderHoeven:2006:IVT

[vdH06f] Joris van der Hoeven. The intermediate value theorem. Lecture Notes in Mathematics, 1888:201–233, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35591-5_9.
pdf.

vanderHoeven:2006:LDE

[vdH06g] Joris van der Hoeven. Linear differential equations. Lecture Notes in Mathematics, 1888:135–164, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35591-5_7.pdf.

vanderHoeven:2006:NPM

[vdH06h] Joris van der Hoeven. The Newton polygon method. Lecture Notes in Mathematics, 1888:57–77, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-35591-5_3.
pdf.

vanderHoeven:2006:OT

[vdH06i] Joris van der Hoeven. Operations on transseries. Lecture Notes in Mathematics, 1888:97–113, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-35591-5_5.pdf.

vanderHoeven:2006:O

[vdH06j] Joris van der Hoeven. Orderings. Lecture Notes in Mathematics, 1888:11–32, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35591-5_1.pdf.

vanderHoeven:2006:T

[vdH06k] Joris van der Hoeven. Transseries. Lecture Notes in Mathematics, 1888:79–96, 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35591-5_4.pdf.

vanderHoeven:2006:TRD

[vdH06l] Joris van der Hoeven. Transseries and Real Differential Algebra, volume 1888 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNMAA2. ISBN 3-540-35590-1 (print), 3-540-35591-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 255 pp. LCCN QA3 .L28 no. 1888. URL http://

link.springer.com/book/10.1007/3-540-35590-1;   http:
//www.springerlink.com/content/978-3-540-35591-5.

Vershik:2003:TLA

[Ver03] A. Vershik. Two lectures on the asymptotic representation theory and statistics of Young diagrams. Lecture Notes in Mathematics, 1815:161–182, 2003. CODEN LNMAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-44890-X_7/.

Veselic:2008:BM

[Ves08a] Ivan Veselić. Back matter. Lecture Notes in Mathematics, 1917: 99–146, 2008. CODEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-72691-3/1.pdf.

Veselic:2008:ERP

[Ves08b] Ivan Veselić. Existence and Regularity Properties of the Integrated Density of States of Random Schrödinger Operators, volume 1917 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 142 pp. LCCN QA3 .L28 no. 1917. URL http://link.

springer.com/book/10.1007/978-3-540-72691-3;   http:/
/www.springerlink.com/content/978-3-540-72691-3.

Veselic:2008:EID

[Ves08c] Ivan Veselić. Existence of the integrated density of states. Lecture Notes in Mathematics, 1917:13–43, 2008. CODEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-72691-3_2.
pdf.

Veselic:2008:FM

[Ves08d] Ivan Veselić. Front matter. Lecture Notes in Mathematics, 1917: I–X, 2008. CODEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-72691-3/1.pdf.

Veselic:2008:LCI

[Ves08e] Ivan Veselić. Lipschitz continuity of the IDS. Lecture Notes in Mathematics, 1917:79–97, 2008. CODEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-72691-3_5.pdf.

Veselic:2008:RO

[Ves08f] Ivan Veselić. Random operators. Lecture Notes in Mathematics, 1917:1–11, 2008. CODEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72691-3_1.pdf.

Veselic:2008:WE

[Ves08g] Ivan Veselić. Wegner estimate. Lecture Notes in Mathematics, 1917:45–56, 2008. CODEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72691-3_3.pdf.

Veselic:2008:WOI

[Ves08h] Ivan Veselić. Wegner’s original idea. Rigorous implementation. Lecture Notes in Mathematics, 1917:57–77, 2008. CO-DEN LNMAA2. ISBN 3-540-72689-6 (print), 3-540-72691-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72691-3_4.pdf.

Villani:2003:OTD

[Vil03] Cedric Villani. Optimal transportation, dissipative PDE’s and functional inequalities. Lecture Notes in Mathematics, 1813:53– 89, 2003. CODEN LNMAA2. ISBN 3-540-40192-X (print), 3-540-44857-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-44857-0_3/.

Villani:2008:EPC

[Vil08] C. Villani. Entropy production and convergence to equilibrium. Lecture Notes in Mathematics, 1916:1–70, 2008. CO-DEN LNMAA2. ISBN 3-540-73704-9 (print), 3-540-73705-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-73705-6_1.pdf.

Vishik:2004:MQA

[Vis04] Alexander Vishik. Motives of quadrics with applications to the theory of quadratic forms. Lecture Notes in Mathematics, 1835: 25–101, 2004. CODEN LNMAA2. ISBN 3-540-20728-7 (print), 3-540-40990-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40990-8_2/.

Vogan:2008:URC

[Vog08] David A. Vogan, Jr. Unitary representations and complex analysis. Lecture Notes in Mathematics, 1931:259–344, 2008.

CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76892-0_5.pdf.

Voiculescu:2000:AFI

[Voi00a] Dan Voiculescu. Addition of freely independent noncommutative random variables. Lecture Notes in Mathematics, 1738:294– 308, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106721/.

Voiculescu:2000:GCF

[Voi00b] Dan Voiculescu. Generalized canonical form, noncrossing partitions. Lecture Notes in Mathematics, 1738:313–316, 2000. CO-DEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0106723/.

Voiculescu:2000:I

[Voi00c] Dan Voiculescu. Introduction. Lecture Notes in Mathematics, 1738:283–284, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0106719.

Voiculescu:2000:MFI

[Voi00d] Dan Voiculescu. Multiplication of freely independent noncommutative random variables. Lecture Notes in Mathematics, 1738: 308–313, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106722/.

Voiculescu:2000:NPO

[Voi00e] Dan Voiculescu. Noncommutative probability and operator algebra background. Lecture Notes in Mathematics, 1738:284– 294, 2000. CODEN LNMAA2. ISBN 3-540-67736-4 (print), 3-540-45029-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0106720/.

Vostrikova:2000:SIP

[VY00] L. Vostrikova and M. Yor. Some invariance properties (of the laws) of Ocone’s martingales. Lecture Notes in Mathematics, 1729:417–431, 2000. CODEN LNMAA2. ISBN 3-540-67314-8 (print), 3-540-46413-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0103817/.

Vershik:2003:ACA

[VY03] Anatoly M. Vershik and Yuri Yakubovich, editors. Asymptotic Combinatorics with Applications to Mathematical Physics: A European Mathematical Summer School held at the Euler Institute, St. Petersburg, Russia July 9–20, 2001, volume 1815 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CO-DEN LNMAA2. ISBN 3-540-40312-4 (print), 3-540-44890-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 1815. URL http://link.springer.com/book/
10.1007/3-540-44890-X;   http://www.springerlink.com/
content/978-3-540-44890-7.

Wagner:2000:NIP

[Wag00] R. Wagner. Notes on an inequality by Pisier for functions on the discrete cube. Lecture Notes in Mathematics, 1745:263–268, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0107220/.

Wakabayashi:2000:AWF

[Wak00a] Seiichiro Wakabayashi. Analytic wave front sets and microfunctions. Lecture Notes in Mathematics, 1737:115–204, 2000. CO-DEN LNMAA2. ISBN 3-540-67603-1 (softcover), 3-540-45161-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Wakabayashi:2000:BCF

[Wak00b] Seiichiro Wakabayashi. Basic calculus of Fourier integral operators and pseudodifferential operators. Lecture Notes in Mathematics, 1737:41–114, 2000. CODEN LNMAA2. ISBN 3-540-67603-1 (softcover), 3-540-45161-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Wakabayashi:2000:CMA

[Wak00c] Seiichiro Wakabayashi. Classical microlocal analysis in the space of hyperfunctions, volume 1737 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67603-1 (softcover), 3-540-45161-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 367 pp. LCCN QA3 .L28 no. 1737; QA3.L28 no.

1737 QA324. URL http://link.springer.com/book/10.1007/
BFb0103767;   http://www.loc.gov/catdir/enhancements/
fy0816/00040014-d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00040014-t.html.

Wakabayashi:2000:H

[Wak00d] Seiichiro Wakabayashi. Hyperfunctions. Lecture Notes in Mathematics, 1737:5–39, 2000. CODEN LNMAA2. ISBN 3-540-67603-1 (softcover), 3-540-45161-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Wakabayashi:2000:LS

[Wak00e] Seiichiro Wakabayashi. Local solvability. Lecture Notes in Mathematics, 1737:259–293, 2000. CODEN LNMAA2. ISBN 3-540-67603-1 (softcover), 3-540-45161-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Wakabayashi:2000:MU

[Wak00f] Seiichiro Wakabayashi. Microlocal uniqueness. Lecture Notes in Mathematics, 1737:205–258, 2000. CODEN LNMAA2. ISBN 3-540-67603-1 (softcover), 3-540-45161-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Waldschmidt:2003:BM

[Wal03a] Michel Waldschmidt. Back matter. Lecture Notes in Mathematics, 1819:345–351, 2003. CODEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-44979-9/1.pdf.

Waldschmidt:2003:LIM

[Wal03b] Michel Waldschmidt. Linear independence measures for logarithms of algebraic numbers. Lecture Notes in Mathematics, 1819: 249–344, 2003. CODEN LNMAA2. ISBN 3-540-40392-2 (print), 3-540-44979-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-44979-5_5/.

Walsh:2006:LS

[Wal06] John B. Walsh. A lost scroll. Lecture Notes in Mathematics, 1874:117–118, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-35513-7_9.pdf.

Wallach:2008:BM

[Wal08a] Nolan R. Wallach. Back matter. Lecture Notes in Mathematics, 1931:377–388, 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-76892-0/1.pdf.

Wallach:2008:QCE

[Wal08b] Nolan R. Wallach. Quantum computing and entanglement for mathematicians. Lecture Notes in Mathematics, 1931:345–376, 2008. CODEN LNMAA2. ISBN 3-540-76891-2 (print), 3-540-76892-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-76892-0_6.pdf.

Wang:2009:HE

[Wan09] Xu-Jia Wang. The k-Hessian equation. Lecture Notes in Mathematics, 1977:177–252, 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01674-5_5.pdf.

Webb:2008:PMS

[Web08] G. F. Webb. Population models structured by age, size, and spatial position. Lecture Notes in Mathematics, 1936:1–49, 2008. CODEN LNMAA2. ISBN 3-540-78272-9 (print), 3-540-78273-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-78273-5_1.pdf;   http://link.springer.
com/content/pdf/bfm:978-3-540-78273-5/1.pdf.

Weissauer:2009:ADC

[Wei09a] Rainer Weissauer. Appendix on double cosets. Lecture Notes in Mathematics, 1968:1–16, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89306-6_12.pdf.

Weissauer:2009:AGC

[Wei09b] Rainer Weissauer. Appendix on Galois cohomology. Lecture Notes in Mathematics, 1968:1–4, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89306-6_11.pdf.

Weissauer:2009:AHL

[Wei09c] Rainer Weissauer. An application of the hard Lefschetz theorem. Lecture Notes in Mathematics, 1968:1–17, 2009. CO-DEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89306-6_1.pdf.

Weissauer:2009:BM

[Wei09d] Rainer Weissauer. Back matter. Lecture Notes in Mathematics, 1968:1–18, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-89306-6/1.pdf.

Weissauer:2009:CL

[Wei09e] Rainer Weissauer. CAP-localization. Lecture Notes in Mathematics, 1968:1–34, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-89306-6_2.pdf.

Weissauer:2009:CIG

[Wei09f] Rainer Weissauer. Character identities and Galois representations related to the group GSp(4). Lecture Notes in Mathematics, 1968:1–99, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-89306-6_4.pdf.

Weissauer:2009:ECS

[Wei09g] Rainer Weissauer. Endoscopy for GSp(4) and the Cohomology of Siegel Modular Threefolds, volume 1968 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 368 pp. LCCN QA3 .L28 no.

1968. URL http://link.springer.com/book/10.1007/978-
3-540-89306-6;   http://www.springerlink.com/content/
978-3-540-89306-6.

Weissauer:2009:FM

[Wei09h] Rainer Weissauer. Front matter. Lecture Notes in Mathematics, 1968:1–16, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-89306-6/1.pdf.

Weissauer:2009:FLT

[Wei09i] Rainer Weissauer. Fundamental lemma (twisted case). Lecture Notes in Mathematics, 1968:1–30, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89306-6_9.
pdf.

Weissauer:2009:LST

[Wei09j] Rainer Weissauer. The Langlands–Shelstad transfer factor. Lecture Notes in Mathematics, 1968:1–20, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89306-6_8.
pdf.

Weissauer:2009:LGE

[Wei09k] Rainer Weissauer. Local and global endoscopy for GSp(4). Lecture Notes in Mathematics, 1968:1–36, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-89306-6_5.
pdf.

Weissauer:2009:RCG

[Wei09l] Rainer Weissauer. The Ramanujan conjecture for genus two Siegel modular forms. Lecture Notes in Mathematics, 1968:1– 21, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89306-6_3.pdf.

Weissauer:2009:RUE

[Wei09m] Rainer Weissauer. Reduction to unit elements. Lecture Notes in Mathematics, 1968:1–13, 2009. CODEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-89306-6_10.pdf.

Weissauer:2009:SCFa

[Wei09n] Rainer Weissauer. A special case of the fundamental lemma I. Lecture Notes in Mathematics, 1968:1–28, 2009. CO-DEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89306-6_6.pdf.

Weissauer:2009:SCFb

[Wei09o] Rainer Weissauer. A special case of the fundamental lemma II. Lecture Notes in Mathematics, 1968:1–31, 2009. CO-DEN LNMAA2. ISBN 3-540-89305-9 (print), 3-540-89306-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-89306-6_7.pdf.

Wengenroth:2003:Ia

[Wen03a] Jochen Wengenroth. 1. Introduction. Lecture Notes in Mathematics, 1810:1–6, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36211-1_1/.

Wengenroth:2003:NHA

[Wen03b] Jochen Wengenroth. 2. Notions from homological algebra. Lecture Notes in Mathematics, 1810:7–15, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36211-1_2/.

Wengenroth:2003:PLF

[Wen03c] Jochen Wengenroth. 3. The projective limit functor for countable spectra. Lecture Notes in Mathematics, 1810:17–57, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36211-1_3/.

Wengenroth:2003:UPS

[Wen03d] Jochen Wengenroth. 4. Uncountable projective spectra. Lecture Notes in Mathematics, 1810:59–76, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36211-1_4/.

Wengenroth:2003:DFH

[Wen03e] Jochen Wengenroth. 5. The derived functors of hom. Lecture Notes in Mathematics, 1810:77–107, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-36211-1_5/.

Wengenroth:2003:ISL

[Wen03f] Jochen Wengenroth. 6. Inductive spectra of locally convex spaces. Lecture Notes in Mathematics, 1810:109–118, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36211-1_6/.

Wengenroth:2003:DF

[Wen03g] Jochen Wengenroth. 7. The duality functor. Lecture Notes in Mathematics, 1810:119–127, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36211-1_7/.

Wengenroth:2003:DFF

[Wen03h] Jochen Wengenroth. Derived Functors in Functional Analysis, volume 1810 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 134 pp. LCCN QA3 .L28 no. 1810. URL http:/

/link.springer.com/book/10.1007/b80165;   http://www.
springerlink.com/content/978-3-540-36211-1.

Wengenroth:2003:Ib

[Wen03i] Jochen Wengenroth. Index. Lecture Notes in Mathematics, 1810: 133–134, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-36211-1_9.

Wengenroth:2003:R

[Wen03j] Jochen Wengenroth. References. Lecture Notes in Mathematics, 1810:129–132, 2003. CODEN LNMAA2. ISBN 3-540-00236-7 (print), 3-540-36211-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36211-1_8/.

Werner:2004:PIR

[Wer04] Wendelin Werner. Part II: Random planar curves and Schramm– Loewner evolutions. Lecture Notes in Mathematics, 1840:107– 195, 2004. CODEN LNMAA2. ISBN 3-540-21316-3 (print), 3-540-39982-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-39982-7_2/.

Wolf-Gladrow:2000:I

[WG00a] Dieter A. Wolf-Gladrow. 1. Introduction. Lecture Notes in Mathematics, 1725:1–13, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46586-7_1/.

Wolf-Gladrow:2000:CA

[WG00b] Dieter A. Wolf-Gladrow. 2. Cellular automata. Lecture Notes in Mathematics, 1725:15–37, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46586-7_2/.

Wolf-Gladrow:2000:LGCa

[WG00c] Dieter A. Wolf-Gladrow. 3. Lattice-gas cellular automata. Lecture Notes in Mathematics, 1725:39–138, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-46586-7_3/.

Wolf-Gladrow:2000:SSM

[WG00d] Dieter A. Wolf-Gladrow. 4. Some statistical mechanics. Lecture Notes in Mathematics, 1725:139–158, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-46586-7_4/.

Wolf-Gladrow:2000:LBM

[WG00e] Dieter A. Wolf-Gladrow. 5. Lattice Boltzmann models. Lecture Notes in Mathematics, 1725:159–246, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-46586-7_5/.

Wolf-Gladrow:2000:A

[WG00f] Dieter A. Wolf-Gladrow. 6. Appendix. Lecture Notes in Mathematics, 1725:247–270, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46586-7_6/.

Wolf-Gladrow:2000:LGCb

[WG00g] Dieter A. Wolf-Gladrow. Lattice Gas Cellular Automata and Lattice Boltzmann Models: An Introduction, volume 1725 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 308 pp. LCCN QA3 .L28 no. 1725. URL http:/

/link.springer.com/book/10.1007/b72010;   http://www.
springerlink.com/content/978-3-540-46586-7.

Wolf-Gladrow:2000:R

[WG00h] Dieter A. Wolf-Gladrow. References. Lecture Notes in Mathematics, 1725:275–308, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46586-7_8/.

Wolf-Gladrow:2000:SI

[WG00i] Dieter A. Wolf-Gladrow. Subject index. Lecture Notes in Mathematics, 1725:271–274, 2000. CODEN LNMAA2. ISBN 3-540-66973-6 (print), 3-540-46586-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46586-7_7/.

Winslow:2005:MMC

[WHG05] R. L. Winslow, R. Hinch, and J. L. Greenstein. Mechanisms and models of cardiac excitation-contraction coupling. Lecture Notes in Mathematics, 1867:97–131, 2005. CODEN LNMAA2. ISBN 3-540-25439-0 (print), 3-540-31438-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/11406501_4.pdf.

Williams:2007:SMV

[Wil07] Mark Williams. Stability of multidimensional viscous shocks. Lecture Notes in Mathematics, 1911:159–227, 2007. CO-DEN LNMAA2. ISBN 3-540-72186-X (print), 3-540-72187-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72187-1_3.pdf.

Williams:2008:NLM

[Wil08a] David Williams. A new look at ‘Markovian’ Wiener–Hopf theory. Lecture Notes in Mathematics, 1934:349–369, 2008. CODEN LNMAA2. ISBN 3-540-77912-4 (print), 3-540-77913-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-77913-1_16.pdf.

Wilson:2008:BM

[Wil08b] Michael Wilson. Back matter. Lecture Notes in Mathematics, 1924:219–228, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-74587-7/1.pdf.

Wilson:2008:CRFa

[Wil08c] Michael Wilson. The Calderón reproducing formula I. Lecture Notes in Mathematics, 1924:85–100, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74587-7_5.
pdf.

Wilson:2008:CRFb

[Wil08d] Michael Wilson. The Calderón reproducing formula II. Lecture Notes in Mathematics, 1924:101–127, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74587-7_6.
pdf.

Wilson:2008:CRFc

[Wil08e] Michael Wilson. The Calderón reproducing formula III. Lecture Notes in Mathematics, 1924:129–143, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-540-74587-7_7.
pdf.

Wilson:2008:EI

[Wil08f] Michael Wilson. An elementary introduction. Lecture Notes in Mathematics, 1924:9–37, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74587-7_2.pdf.

Wilson:2008:ES

[Wil08g] Michael Wilson. Exponential square. Lecture Notes in Mathematics, 1924:39–68, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74587-7_3.pdf.

Wilson:2008:FMT

[Wil08h] Michael Wilson. A Fourier multiplier theorem. Lecture Notes in Mathematics, 1924:197–202, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74587-7_12.pdf.

Wilson:2008:FM

[Wil08i] Michael Wilson. Front matter. Lecture Notes in Mathematics, 1924:I–XII, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-74587-7/1.pdf.

Wilson:2008:GG

[Wil08j] Michael Wilson. Goodbye to good-λ. Lecture Notes in Mathematics, 1924:189–195, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74587-7_11.pdf.

Wilson:2008:MDS

[Wil08k] Michael Wilson. Many dimensions; smoothing. Lecture Notes in Mathematics, 1924:69–84, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74587-7_4.pdf.

Wilson:2008:OS

[Wil08l] Michael Wilson. Orlicz spaces. Lecture Notes in Mathematics, 1924:161–188, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74587-7_10.pdf.

Wilson:2008:RPE

[Wil08m] Michael Wilson. Random pointwise errors. Lecture Notes in Mathematics, 1924:213–218, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74587-7_14.pdf.

Wilson:2008:SO

[Wil08n] Michael Wilson. Schrödinger operators. Lecture Notes in Mathematics, 1924:145–150, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74587-7_8.pdf.

Wilson:2008:SA

[Wil08o] Michael Wilson. Some assumptions. Lecture Notes in Mathematics, 1924:1–7, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74587-7_1.pdf.

Wilson:2008:SSI

[Wil08p] Michael Wilson. Some singular integrals. Lecture Notes in Mathematics, 1924:151–160, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74587-7_9.pdf.

Wilson:2008:VVI

[Wil08q] Michael Wilson. Vector-valued inequalities. Lecture Notes in Mathematics, 1924:203–212, 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-74587-7_13.pdf.

Wilson:2008:WLP

[Wil08r]  Michael Wilson.  Weighted Littlewood–Paley Theory and

Exponential-Square Integrability, volume 1924 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNMAA2. ISBN 3-540-74582-3 (print), 3-540-74587-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 224 pp. LCCN QA3 .L28 no.

1924. URL http://link.springer.com/book/10.1007/978-
3-540-74587-7;   http://www.springerlink.com/content/
978-3-540-74587-7.

Wittenberg:2007:APSa

[Wit07a] Olivier Wittenberg. Arithmétique des pinceaux semi-stables de courbes de genre 1 (première partie). (French) []. Lecture Notes in Mathematics, 1901:19–72, 2007. CODEN LNMAA2. ISBN 3-540-69137-5 (print), 3-540-69141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69141-9_1.pdf.

Wittenberg:2007:APSb

[Wit07b] Olivier Wittenberg. Arithmétique des pinceaux semi-stables de courbes de genre 1 (seconde partie). (French) []. Lecture Notes in Mathematics, 1901:73–108, 2007. CODEN LNMAA2. ISBN 3-540-69137-5 (print), 3-540-69141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-540-69141-9_2.pdf.

Wittenberg:2007:BM

[Wit07c] Olivier Wittenberg. Back matter. Lecture Notes in Mathematics, 1901:201–222, 2007. CODEN LNMAA2. ISBN 3-540-69137-5 (print), 3-540-69141-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-69141-9/1.pdf.

Wittenberg:2007:FM

[Wit07d] Olivier Wittenberg. Front matter. Lecture Notes in Mathematics, 1901:I–XXIV, 2007. CODEN LNMAA2. ISBN 3-540-69137-5 (print), 3-540-69141-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-69141-9/1.pdf.

Wittenberg:2007:IDQ

[Wit07e] Olivier Wittenberg. Intersections de deux quadriques et pinceaux de courbes de genre 1. (French) [Intersections of two quadrics and pencils of curves of genus 1], volume 1901 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNMAA2. ISBN 3-540-69137-5 (print), 3-540-69141-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 218 pp. LCCN QA3 .L28 no.

1901. URL http://link.springer.com/book/10.1007/3-540-
69137-5;   http://www.springerlink.com/content/978-3-
540-69141-9.

Wittenberg:2007:PHP

[Wit07f] Olivier Wittenberg. Principe de Hasse pour les surfaces de del Pezzo de degré 4. (French) []. Lecture Notes in Mathematics, 1901:109–200, 2007. CODEN LNMAA2. ISBN 3-540-69137-5 (print), 3-540-69141-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-69141-9_3.pdf.

Wonham:2008:BM

[WL08a] M. J. Wonham and M. A. Lewis. Back matter. Lecture Notes in Mathematics, 1945:391–412, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-540-78911-6/1.pdf.

Wonham:2008:CAM

[WL08b] M. J. Wonham and M. A. Lewis. A comparative analysis of models for West Nile virus. Lecture Notes in Mathematics, 1945:365–390, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_14.pdf.

Wojtaszczyk:2007:BM

[Woj07a] J. O. Wojtaszczyk. Back matter. Lecture Notes in Mathematics, 1910:315–332, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-72053-9/1.pdf.

Wojtaszczyk:2007:SNC

[Woj07b] J. O. Wojtaszczyk. The square negative correlation property for generalized Orlicz balls. Lecture Notes in Mathematics, 1910:305– 313, 2007. CODEN LNMAA2. ISBN 3-540-72052-9 (print), 3-540-72053-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-72053-9_18.pdf.

Wu:2008:SSP

[Wu08] Jianhong Wu. Spatial structure: Partial differential equations models. Lecture Notes in Mathematics, 1945:191–203, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_8.pdf.

Xiao:2001:CE

[Xia01a] Jie Xiao. Composite embedding. Lecture Notes in Mathematics, 1767:13–22, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45434-9_2/.

Xiao:2001:DL

[Xia01b] Jie Xiao. Dyadic localization. Lecture Notes in Mathematics, 1767:87–104, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45434-9_8/.

Xiao:2001:FM

[Xia01c] Jie Xiao. Fundamental material. Lecture Notes in Mathematics, 1767:1–12, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45434-9_1/.

Xiao:2001:HC

[Xia01d] Jie Xiao. Holomorphic Q Classes, volume 1767 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LN-MAA2. ISBN 3-540-42625-6 (paperback). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 112 pp. LCCN QA3 .L28 no. 1767; QA3 .L28 no. 1767 QA331.7. URL http:/

/link.springer.com/book/10.1007/b87877;   http://www.
loc.gov/catdir/enhancements/fy0817/2001049621-d.html;
 http://www.loc.gov/catdir/enhancements/fy0817/2001049621-
t.html.

Xiao:2001:IOS

[Xia01e] Jie Xiao. Inner-outer structure. Lecture Notes in Mathematics, 1767:45–56, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45434-9_5/.

Xiao:2001:MCM

[Xia01f] Jie Xiao. Modified Carleson measures. Lecture Notes in Mathematics, 1767:35–44, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45434-9_4/.

Xiao:2001:PHE

[Xia01g] Jie Xiao. Pseudo-holomorphic extension. Lecture Notes in Mathematics, 1767:57–66, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45434-9_6/.

Xiao:2001:RTE

[Xia01h] Jie Xiao. Representation via ∂-equation. Lecture Notes in Mathematics, 1767:67–86, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45434-9_7/.

Xiao:2001:SE

[Xia01i] Jie Xiao. Series expansions. Lecture Notes in Mathematics, 1767: 23–34, 2001. CODEN LNMAA2. ISBN 3-540-42625-6 (print), 3-540-45434-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45434-9_3/.

Xiao:2009:BM

[Xia09a] Yimin Xiao. Back matter. Lecture Notes in Mathematics, 1962: 213–222, 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-85994-9/1.pdf.

Xiao:2009:SPP

[Xia09b] Yimin Xiao. Sample path properties of anisotropic Gaussian random fields. Lecture Notes in Mathematics, 1962:145–212, 2009. CODEN LNMAA2. ISBN 3-540-85993-4 (print), 3-540-85994-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-85994-9_5.pdf.

Xu:2008:IBQ

[Xu08] Quanhua Xu. Interactions between quantum probability and operator space theory. Lecture Notes in Mathematics, 1954:117–159, 2008. CODEN LNMAA2. ISBN 3-540-69364-5 (print), 3-540-69365-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-69365-9_4.pdf.

Xu:2009:E

[Xu09] Fangjun Xu. On the equation μ = S_{t}μ ∗ μ_{t}. Lecture Notes in Mathematics, 1979:137–145, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01763-6_4.pdf.

Yafaev:2000:ACS

[Yaf00a] Dmitri R. Yafaev. Asymptotic completeness. a sketch of proof. Lecture Notes in Mathematics, 1735:106–117, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0105544/.

Yafaev:2000:BC

[Yaf00b] Dmitri R. Yafaev. Basic concepts. Lecture Notes in Mathematics, 1735:1–13, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0105532/.

Yafaev:2000:GFT

[Yaf00c] Dmitri R. Yafaev. The generalized Fourier transform. Lecture Notes in Mathematics, 1735:40–46, 2000. CODEN LN-MAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0105536/.

Yafaev:2000:HM

[Yaf00d] Dmitri R. Yafaev. The Heisenberg model. Lecture Notes in Mathematics, 1735:137–144, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105547/.

Yafaev:2000:IOS

[Yaf00e] Dmitri R. Yafaev. Infinite obstacle scattering. Lecture Notes in Mathematics, 1735:145–153, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105548/.

Yafaev:2000:LRC

[Yaf00f] Dmitri R. Yafaev. The long-range case. Lecture Notes in Mathematics, 1735:67–79, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0105540/.

Yafaev:2000:LRI

[Yaf00g] Dmitri R. Yafaev. Long-range interactions. the scheme of smooth perturbations. Lecture Notes in Mathematics, 1735:30–39, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0105535/.

Yafaev:2000:LRM

[Yaf00h] Dmitri R. Yafaev. Long-range matrix potentials. Lecture Notes in Mathematics, 1735:47–52, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105537/.

Yafaev:2000:NCS

[Yaf00i] Dmitri R. Yafaev. New channels of scattering. Lecture Notes in Mathematics, 1735:128–136, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105546/.

Yafaev:2000:RSM

[Yaf00j] Dmitri R. Yafaev. The relative scattering matrix. Lecture Notes in Mathematics, 1735:80–85, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105541/.

Yafaev:2000:RET

[Yaf00k] Dmitri R. Yafaev. Resolvent equations for three-particle systems. Lecture Notes in Mathematics, 1735:96–105, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0105543/.

Yafaev:2000:SME

[Yaf00l] Dmitri R. Yafaev. The scattering matrix and eigenfunctions for multiparticle systems. Lecture Notes in Mathematics, 1735:118– 127, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0105545/.

Yafaev:2000:STS

[Yaf00m] Dmitri R. Yafaev. Scattering Theory: Some Old and New Problems, volume 1735 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 169 pp. LCCN QA3 .L28 no. 1735. URL http:/

/link.springer.com/book/10.1007/BFb0105531;   http://
www.springerlink.com/content/978-3-540-45170-9.

Yafaev:2000:SSP

[Yaf00n] Dmitri R. Yafaev. Setting the scattering problem. Lecture Notes in Mathematics, 1735:86–95, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105542/.

Yafaev:2000:SRC

[Yaf00o] Dmitri R. Yafaev. The short-range case. Lecture Notes in Mathematics, 1735:59–66, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0105539/.

Yafaev:2000:SRIa

[Yaf00p] Dmitri R. Yafaev. Short-range interactions. asymptotic completeness. Lecture Notes in Mathematics, 1735:14–23, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0105533/.

Yafaev:2000:SRIb

[Yaf00q] Dmitri R. Yafaev. Short-range interactions. miscellaneous. Lecture Notes in Mathematics, 1735:24–29, 2000. CODEN LN-MAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0105534/.

Yafaev:2000:SR

[Yaf00r] Dmitri R. Yafaev. A stationary representation. Lecture Notes in Mathematics, 1735:53–58, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105538/.

Yan:2006:SPT

[Yan06] Jia-An Yan. A simple proof of two generalized Borel–Cantelli lemmas. Lecture Notes in Mathematics, 1874:77–79, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_7.pdf.

Yan:2008:DTS

[Yan08] Ping Yan. Distribution theory, stochastic processes and infectious disease modelling. Lecture Notes in Mathematics, 1945:229–293, 2008. CODEN LNMAA2. ISBN 3-540-78910-3 (print), 3-540-78911-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-78911-6_10.pdf.

Yang:2009:BM

[Yan09a] Paul Yang. Back matter. Lecture Notes in Mathematics, 1977: 1–14, 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-01674-5/1.pdf.

Yang:2009:MSC

[Yan09b] Paul Yang. Minimal surfaces in CR geometry. Lecture Notes in Mathematics, 1977:253–273, 2009. CODEN LNMAA2. ISBN 3-642-01673-1 (print), 3-642-01674-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-01674-5_6.pdf.

Yomdin:2004:IC

[YC04a] Yosef Yomdin and Georges Comte. 1. Introduction and content. Lecture Notes in Mathematics, 1834:1–22, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40960-1_
1/.

Yomdin:2004:SAR

[YC04b] Yosef Yomdin and Georges Comte. 10. Some applications and related topics. Lecture Notes in Mathematics, 1834:131–169, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40960-1_10/.

Yomdin:2004:E

[YC04c] Yosef Yomdin and Georges Comte. 2. Entropy. Lecture Notes in Mathematics, 1834:23–32, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-40960-1_2/.

Yomdin:2004:MV

[YC04d] Yosef Yomdin and Georges Comte. 3. Multidimensional variations. Lecture Notes in Mathematics, 1834:33–45, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40960-1_
3/.

Yomdin:2004:STS

[YC04e] Yosef Yomdin and Georges Comte. 4. Semialgebraic and tame sets. Lecture Notes in Mathematics, 1834:47–58, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40960-1_
4/.

Yomdin:2004:VST

[YC04f] Yosef Yomdin and Georges Comte. 5. Variations of semialgebraic and tame sets. Lecture Notes in Mathematics, 1834:59– 73, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40960-1_5/.

Yomdin:2004:SEA

[YC04g] Yosef Yomdin and Georges Comte. 6. Some exterior algebra. Lecture Notes in Mathematics, 1834:75–82, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-40960-1_6/.

Yomdin:2004:BVU

[YC04h] Yosef Yomdin and Georges Comte. 7. Behaviour of variations under polynomial mappings. Lecture Notes in Mathematics, 1834: 83–98, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40960-1_7/.

Yomdin:2004:QTC

[YC04i] Yosef Yomdin and Georges Comte. 8. Quantitative transversality and cuspidal values. Lecture Notes in Mathematics, 1834: 99–107, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40960-1_8/.

Yomdin:2004:MFS

[YC04j] Yosef Yomdin and Georges Comte. 9. Mappings of finite smoothness. Lecture Notes in Mathematics, 1834:109–130, 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-40960-1_9/.

Yomdin:2004:GR

[YC04k] Yosef Yomdin and Georges Comte. Glossary and references. Lecture Notes in Mathematics, 1834:171–186, 2004. CODEN LN-MAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-40960-1_
11/.

Yomdin:2004:TGA

[YC04l] Yosef Yomdin and Georges Comte. Tame Geometry with Application in Smooth Analysis, volume 1834 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-20612-4 (print), 3-540-40960-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 186 pp. LCCN QA3 .L28 no. 1834.

URL http://link.springer.com/book/10.1007/b94624;
http://www.springerlink.com/content/978-3-540-40960-1.

Yoccoz:2002:ALC

[Yoc02] Jean-Christophe Yoccoz. Analytic linearization of circle diffeomorphisms. Lecture Notes in Mathematics, 1784:125–173, 2002. CODEN LNMAA2. ISBN 3-540-43726-6 (print), 3-540-47928-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-47928-4_3/.

Yor:2002:GFS

[Yor02a] Marc Yor. Grossissement d’une filtration et semi-martingales: Théorèmes généraux. (French) []. Lecture Notes in Mathematics, 1771:88–96, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45530-1_7/.

Yor:2002:SED

[Yor02b] Marc Yor. Sous-espaces denses dans L^{1} ou H^{1} et représentation des martingales. (French) []. Lecture Notes in Mathematics, 1771: 428–472, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_19/.

Yor:2002:ISO

[Yor02c] Marc Yor. Sur les intégrales stochastiques optionelles et une suite remarquable de formules exponentielles. (French) []. Lecture Notes in Mathematics, 1771:341–360, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45530-1_13/.

Yor:2002:QAI

[Yor02d] Marc Yor. Sur quelques approximations d’intégrales stochastiques. (French) []. Lecture Notes in Mathematics, 1771:330– 340, 2002. CODEN LNMAA2. ISBN 3-540-42813-5 (print), 3-540-45530-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45530-1_12/.

Yor:2006:LSW

[Yor06] Marc Yor. The life and scientific work of Paul André Meyer (August 21st, 1934–January 30th, 2003) “Un modèle pour nous tous”. Lecture Notes in Mathematics, 1874:13–26, 2006. CO-DEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_2.pdf.

Yano:2009:LFH

[YYY09] Kouji Yano, Yuko Yano, and Marc Yor. On the laws of first hitting times of points for one-dimensional symmetric stable Lévy processes. (French) []. Lecture Notes in Mathematics, 1979:187– 227, 2009. CODEN LNMAA2. ISBN 3-642-01762-2 (print), 3-642-01763-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-01763-6_8.pdf.

Zakai:2005:RTP

[Zak05] Moshe Zakai. Rotations and tangent processes on Wiener space. Lecture Notes in Mathematics, 1857:205–225, 2005. CODEN LNMAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-31449-3_15.pdf.

Zeitouni:2004:PIR

[Zei04] Ofer Zeitouni. Part II: Random walks in random environment. Lecture Notes in Mathematics, 1837:189–312, 2004. CODEN LNMAA2. ISBN 3-540-20832-1 (print), 3-540-39874-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-39874-5_
2/.

Zhan:2002:ILP

[Zha02a] Xingzhi Zhan. 1. Inequalities in the Löwner partial order. Lecture Notes in Mathematics, 1790:1–15, 2002. CODEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-45421-2_1/.

Zhan:2002:ME

[Zha02b] Xingzhi Zhan. 2. Majorization and eigenvalues. Lecture Notes in Mathematics, 1790:17–25, 2002. CODEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-45421-2_2/.

Zhan:2002:SV

[Zha02c] Xingzhi Zhan. 3. Singular values. Lecture Notes in Mathematics, 1790:27–54, 2002. CODEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45421-2_3/.

Zhan:2002:NI

[Zha02d] Xingzhi Zhan. 4. Norm inequalities. Lecture Notes in Mathematics, 1790:55–98, 2002. CODEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-45421-2_4/.

Zhan:2002:SVW

[Zha02e] Xingzhi Zhan. 5. Solution of the van der Waerden conjecture. Lecture Notes in Mathematics, 1790:99–109, 2002. CODEN LN-MAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-45421-2_
5/.

Zhan:2002:I

[Zha02f] Xingzhi Zhan. Index. Lecture Notes in Mathematics, 1790:115– 116, 2002. CODEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45421-2_7.

Zhan:2002:MI

[Zha02g] Xingzhi Zhan. Matrix Inequalities, volume 1790 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CO-DEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 116 pp. LCCN QA3 .L28 no. 1790. URL http:/

/link.springer.com/book/10.1007/b83956;   http://www.
loc.gov/catdir/enhancements/fy0817/2002070800-d.html;
 http://www.loc.gov/catdir/enhancements/fy0817/2002070800-
t.html;   http://www.springerlink.com/content/978-3-540-
45421-2;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0075-8434&volume=1790.

Zhan:2002:R

[Zha02h] Xingzhi Zhan. References. Lecture Notes in Mathematics, 1790: 110–114, 2002. CODEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-45421-2_6/.

Zhidkov:2001:EER

[Zhi01a] Peter E. Zhidkov. Evolutionary equations. results on existence. Lecture Notes in Mathematics, 1756:9–38, 2001. CODEN LN-MAA2. ISBN 3-540-41833-4 (print), 3-540-45276-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/3-540-45276-1_3/.

Zhidkov:2001:I

[Zhi01b] Peter E. Zhidkov. Introduction. Lecture Notes in Mathematics, 1756:1–4, 2001. CODEN LNMAA2. ISBN 3-540-41833-4 (print), 3-540-45276-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45276-1_1/.

Zhidkov:2001:IM

[Zhi01c] Peter E. Zhidkov. Invariant measures. Lecture Notes in Mathematics, 1756:105–136, 2001. CODEN LNMAA2. ISBN 3-540-41833-4 (print), 3-540-45276-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-45276-1_6/.

Zhidkov:2001:KVN

[Zhi01d] Peter E. Zhidkov. Korteweg–de Vries and Nonlinear Schrödinger Equations: Qualitative Theory, volume 1756 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNMAA2. ISBN 3-540-41833-4 (print), 3-540-45276-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 147 pp. LCCN QA3 .L28 no.

1756. URL http://link.springer.com/book/10.1007/3-540-
45276-1;   http://www.springerlink.com/content/978-3-
540-45276-8.

Zhidkov:2001:N

[Zhi01e] Peter E. Zhidkov. Notation. Lecture Notes in Mathematics, 1756: 5–7, 2001. CODEN LNMAA2. ISBN 3-540-41833-4 (print), 3-540-45276-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
3-540-45276-1_2/.

Zhidkov:2001:SS

[Zhi01f] Peter E. Zhidkov. Stability of solutions. Lecture Notes in Mathematics, 1756:79–104, 2001. CODEN LNMAA2. ISBN 3-540-41833-4 (print), 3-540-45276-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/3-540-45276-1_5/.

Zhidkov:2001:SP

[Zhi01g] Peter E. Zhidkov. Stationary problems. Lecture Notes in Mathematics, 1756:39–78, 2001. CODEN LNMAA2. ISBN 3-540-41833-4 (print), 3-540-45276-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/3-540-45276-1_4/.

Zolesio:2000:SAW

[Zol00] Jean-Paul Zolésio. Shape analysis and weak flow. Lecture Notes in Mathematics, 1740:157–341, 2000. CODEN LNMAA2. ISBN 3-540-67971-5 (print), 3-540-44486-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0106743/.

Zumbrun:2007:BM

[Zum07a] Kevin Zumbrun. Back matter. Lecture Notes in Mathematics, 1911:327–356, 2007. CODEN LNMAA2. ISBN 3-540-72186-X (print), 3-540-72187-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-72187-1/1.pdf.

Zumbrun:2007:PSC

[Zum07b] Kevin Zumbrun. Planar stability criteria for viscous shock waves of systems with real viscosity. Lecture Notes in Mathematics, 1911:229–326, 2007. CODEN LNMAA2. ISBN 3-540-72186-X (print), 3-540-72187-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-72187-1_4.pdf.

Zvavitch:2000:MES

[Zva00] A. Zvavitch. More on embedding subspaces of L_{p} into ℓ_{p}^{N}, 0 < p < 1. Lecture Notes in Mathematics, 1745:269–280, 2000. CODEN LNMAA2. ISBN 3-540-41070-8 (print), 3-540-45392-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0107221/.

Zaitsev:2004:RMC

[ZZ04] Dmitri Zaitsev and Giuseppe Zampieri, editors. Real Methods in Complex and CR Geometry: Lectures given at the C.I.M.E. Summer School held in Martina Franca, Italy, June 30–July 6, 2002, volume 1848 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNMAA2. ISBN 3-540-22358-4 (print), 3-540-44487-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1848. URL http:/

/link.springer.com/book/10.1007/b98482;   http://www.
springerlink.com/content/978-3-540-44487-9.